### v9

# SPECIAL RELATIVITY

SPECIAL RELATIVITY
Evgeny Saldin Draft from May 2024
Applications to astronomy and the accelerator physics
Evgeny Saldin Draft from May 2024
New material has been added to the Chapter 1, Chapter 3, Chapter 4, Chapter 5, Chapter 9, Chapter 10, Chapter 11, Chapter 13.
If you have any comments, please send an email to evgueni.saldin@desy.de.
*Preface
There are many books on the classical subject of special relativity. However, after having spent a number of years, both in relativistic engineering and research with relativity, I have come to the conclusion that there exist a place for a new book. I do believe that the present book is not quite the same as the others, mainly due to attempt to cast light on dark corners.
I should make it clear what this little book is not. It is not a textbook on relativity theory. What the book is about is the nature of special relativistic kinematics, its relation to space and time, and the operational interpretation of coordinate transformations. Every theory contains a number of quantities that can be measured by experiment and an expressions that cannot possibly be observed. Whenever we have a theory containing an arbitrary convention, we should examine what parts of the theory depend on the choice of that convention and what parts do not. The distinction is not always made and many authors claim some data to be observable, according to arbitrary conventions, which do not correspond to any physical experiment. This leads to inconsistencies and paradoxes that should be avoided at all cost.
The practical approach used in the book should be acceptable to astronomers, space engineers, accelerator engineers, and more generally, relativistic engineers. This approach, unusual in the relativistic literature, may be clarified by quoting one of the problems discussed in the text: the new light beam kinematics for rotating frame of references. Since we live on such a rotating (earth-based) frame of reference, difference in relativistic kinematics between rotating and non-rotating frames of reference is of great practical as well as theoretical significance. A correct solution of this problem requires the use of relativistic principles even at low velocities since the first-order terms in $(v/c)$ play a fundamental role in the non-inertial relativistic kinematics of light propagation.
All the results presented here are derived from the ”first principles”, and all steps involving physical principles are given. To preserve a self-consistent style, I place the derivation of auxiliary results in appendices. To help readers form their own opinion on the topics discussed, the end of each chapter has a suggested bibliography together with relevant remarks. The list of references includes only the papers I have consulted directly. A lot of papers remain unmentioned, and for this I apologize.
I am grateful to my longtime friends Gianluca Geloni and Vitaly Kocharyan for discussions over many years about much of the material in this book. I should also like to express my thanks to DESY (Deutsches Electronen-Synchrotron) for enabling me to work in this interesting field.
1 Introduction
The standard books on special relativity do not usually address the questions of the physical meaning of relativistic effects and the nature of space-time. The presentation of the subject in the present book differs somewhat from the usual one in that the four-dimensional geometric formulation of the theory plays a dominant role than in most of the current textbooks.
The book begins with a critical survey of the present approaches to special relativity. The established way of looking at special relativity is based on Einstein postulates: the principle of relativity and the constancy of the velocity of light. In the most general geometric approach to the theory of special relativity, the principle of relativity, in contrast to Einstein formulation, is only a consequence of the geometry of space-time. We point that the essence of the special theory of relativity consists in the following postulate: all physical processes proceed in four-dimensional space-time, the geometry of which is pseudo-Euclidean.
The space-time geometric approach to special relativity deals with all possible choices of coordinates of the chosen reference frames, and therefore the second Einstein postulate, referred to as the constancy of the coordinate speed of light, does not have a place in this more general formulation. Only in Lorentz coordinates, when Einstein’s synchronization of distant clocks and Cartesian space coordinates are used, the coordinate speed of light is isotropic and constant. Thus, the basic elements of the space-time geometric formulation of the special relativity and the usual Einstein’s formulation, are quite different.
It should be emphasized that in practical applications there are two choices of clock synchronization convention useful to consider:
(a) Einstein’s convention, leading to the Lorentz transformations between frames.
(b) Absolute time convention, leading to the Galilean transformations between frames.
Absolute time (or simultaneity) can be introduced in special relativity without affecting neither the logical structure, no the (convention-independent) predictions of the theory. In the theory of relativity, this choice may seem quite unusual, but it is usually most convenient when one wants to connect to laboratory reality.
There is a widespread view that only philosophers of physics discuss the issue of distant clock synchronization. Indeed, a typical physical laboratory contains no space-time grid. It should be clear that a rule-clock structure exist only in our mind and manipulations with non existing clocks in the special relativity are an indispensable prerequisite for the application of dynamics and electrodynamics theory in the coordinate representation. Such situation usually forces physicist to believe that the application of the theory of relativity to the study of physical processes is possible without detailed knowledge of the clocks synchronization procedure.
However, many problems in special relativity can be adequately treated only by an approach which uses the non-standard absolute time synchronization. One of the features that is unique to this book is its treatment of the absolute time coordinatization. No other books deals with Galilean transformations in the framework of special relativity.
Third chapter presents an ”operational interpretation” of the Lorentz and absolute time coordinatizations. This is probably the most important chapter of this book. Today the statement about correctness of Galilean transformations is a ”shocking heresy”, which offends the ”relativistic” intuition and the generally accepted way of looking at special relativity of most physicists. The difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view will be an important discovery for every special relativity expert. To our knowledge, neither operational interpretation of the absolute time coordinatization nor the difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view, are given elsewhere in the literature.
We have tried to keep the mathematical complexity of the discussion to a minimum. The derivation of relativistic kinematics in the first part of the book (Chapter 3-11) is fairly elementary from a mathematical point of view, but it is conceptually subtle. We use synthetic approach to present the material: some simple models are studied first, and more complicated ones are introduced gradually.
We start with aberration of light phenomena. Light, being a special case of electromagnetic waves, is described by the electrodynamics theory. It is well known that the electrodynamics theory meets all requirements of the theory of relativity and therefore must accurately describe the properties of such a typical relativistic object as light.
In the Chapter 4-5, we present a critical reexamination of the existing aberration of light theory. The phenomenon of aberration of light is by no means simple to describe, even in the first order in $v/c$: a large number of incorrect results can be found in the literature. The utilization of the electrodynamics in the absolute time coordinatization becomes indispensable when we consider optical phenomena associated with a relative motion of two (or more) light sources. There is no problem to introduce Lorentz coordinates when these sources are independent. The possibility to introduce Lorentz coordinates in this situation consists in introducing individual coordinate system (i.e. individual set of clocks) for each sources. Suppose now that the second (e.g. secondary) source interacts with the light beam radiated from the first source. The peculiarity of this problem with the viewpoint of the special relativity is that here we cannot prepare for sources a common Lorentz coordinatization in the single inertial frame.
It is generally believed that for a mirror moving tangentially to its surface the law of reflection which holds for the stationary mirror is preserved. This statement, presented in most textbooks, is incorrect. The first error in the textbook argument is conceptual in nature. The observer describes the light beam from the emitter at rest using Minkowski metric. According to textbooks, the metric of the moving mirror is the same (i.e. Minkowski metric). The authors of textbooks incorrectly belief that the common Lorentz coordinatization for emitter and mirror in the single inertial frame can be assigned. This is misconception. The second error in the textbooks follows from the fact that the concept of a plane wave and an infinite plane mirror is used in the case of the tangential motion. In this case the ”mirror motion” is not a real observable effect. Indeed, if the mirror is infinite the problem is not time dependent. It is obvious that only the motion of a finite mirror is meaningful. All experimental methods for measuring the aberration increment operate with light signals, and hence do not measure the phase velocity (i.e. frequency and wavenumber vector) but the group velocity. What authors of textbooks generally overlooked is the fact that the energy transport problem is well-defined only if the source and mirror apertures have already been specified.
Suppose that we have a point source and observer with his measuring devices are at rest in the inertial lab frame. We demonstrate that when a finite aperture mirror is moving tangentially with a constant velocity and interacts in the far-field zone with the radiated light, there is a deviation of the energy transport for the reflected light. The result is extremely important from the applied point of view. According to the Babinet’s principle, this remarkable prediction of our aberration of light theory is correct also for light transmitted through a hole in the moving opaque screen or, consequently, through a moving open end of the telescope barrel.
Questions related to transmission through the transversely moving pupil detection (e.g. a moving end of the telescope barrel) lead to serious misunderstanding, which is actually due to an inadequate understanding of several complicated aspects of the statistical optics. Today one is told that the phenomenon of aberration of light could be interpreted, using ray model of light. One case of rather great interest is that which corresponds to transmission through a moving telescope tube. The rays of light coming from the star are falling on the telescope tube and, according to literature, not interact with its sides. One could naively expect that the region of applicability of ray optics, following from the textbooks reasoning, should be identified with any spatially incoherent radiation. This is misconception. In particular, a spatially completely incoherent source (e.g. an incandescent lamp or a star) is actually a system of elementary (statistically independent) point sources with different offsets. The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$, where $\lambda$ is the (visible) radiation wavelength. Radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individually incoherent sources. An elementary source produces in front of pupil detection effectively a plane wave and the measuring instrument always influences the measured radiation (due to unavoidable diffraction of a plane wave by an aperture). It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single point source characteristics like independence on source motion.
The Chapter 6 deals with astronomical applications. The effect of stellar aberration seems to be one of the simplest phenomena in astronomical observations. In spite of its apparent simplicity, aberration seems to be one of the most intricate effects in special relativity. It is widely believed that stellar aberration depends on the relative velocity of the source (star) and observer. Observations show clearly that stellar aberration does not depend on the relative motion between star and telescope on the earth. The lack of symmetry, between the cases when either the source or telescope is moving is shown clearly on the basis of the separation of binary stars. The relative motion of these stars with respect to each other (and hence, with respect to the earth) is never followed by any aberration, although motion of these stars is, sometimes, much faster than that of the earth around the sun. It should be stressed that it is the telescope and not the star that must change its velocity (relative to the fixed stars) to cause aberration. Contradiction is so plane that some physicists use this fact to argue that stellar aberration contradicts the special theory of relativity. There is no available explanation for the fact that, while the observational data on stellar aberration are compatible with moving earth, the symmetric description, when the star possesses the relative transverse motion, does not apparently lead to observations compatible with predictions. We demonstrate that the fact that we do not see myriads of widely separated binaries in wild gyration does not require any fundamental change of outlook, but it does require that aberration of ”distant” stars is treated in the framework of space-time geometric approach.
It is generally believed that the phenomenon of aberration of light could be interpreted, using the corpuscular model of light, as being analogous of the observation of the oblique fall of raindrops by a moving observer. This is a classical kinematics method to the computation of the stellar aberration used in astronomy for about three hundred years. This book develops a relativistic theory of aberration of light in rotating frames. To see the consequences of relativistic modification of stellar aberration, we start with the classical aberration increment, or $\theta_{a}=v/V$. Here $v$ is the velocity of observer and $V$ is the velocity of particles. It is assumed that $v\ll V$. According to the conventional approach, if we neglect $v^{2}/c^{2}$ comparison with 1, stellar aberration increment is simplified to elementary formula $\theta_{a}=v/c$. This result is particularly remarkable because the study of stellar aberration is intimately connected with the old kinematics: the Galilean vectorial law of addition of velocities is actually used. We come to the conclusion that the standard analysis of the stellar aberration does not take into account the fundamental difference between velocity of light and velocity of raindrops.
The satisfactory treatment of the relativistic modification should be based on the two relativistic parameters. For an observer on the earth, it is, with respect to the solar referential frame, of about 30 km/s, corresponding to the earth motion around the sun. Clearly, in the theory of stellar aberration, we consider the small expansion parameter $v/c\simeq 10^{-4}$ neglecting terms of order of $v^{2}/c^{2}$. In the theory of stellar aberration, we, however, must also consider the relativistic parameter $V/c$. For light $V/c=1$ and the special theory of relativity says that stellar aberration cannot be discussed as a topic in the classical theory. Light is always a (”ultra”) relativistic object, no matter how small the ratio $v/c$ may be. According to the special theory of relativity, physical processes takes place in a metric space-time geometry, so the geodesic interval must be used to describe aberration of light phenomenon.
The problem of the earth-based measurements is solved with the discovery of the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars. We demonstrate that for explanation of the aberration of light effect in the rotating frame of reference one does not need neither modify the special theory of relativity, nor apply general theory of relativity. It is only necessary to strictly follow the special relativity.
All phenomena in non-inertial reference system should be considered in the framework of the space-time geometric approach using metric tensor. Using Langevin metric in a frame of reference attending rotation enables us to explain all earth-based experiments. A correct solution of this problem in the rotating frame requires the use of metric tensor even in experiments of first order in $v/c$ since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of a light beam propagation. The historical rate of the studies of the optical effects in rotating frames is very unusual. The Sagnac effect was discovered in 1913 and described by Langevin in 1921. It is interesting that the Langevin metric has in the past hardly ever been applied in the stellar aberration theory.
We demonstrate that the aberration of light is a complex phenomenon which must be branched out into a number of varieties according to their origin. These branching out takes place depending on what is the cause of aberration - whether it is a jump in the velocity (with respect to the fixed stars) of the observer or of the light source. Aberration could undergo further splitting - depending on the physical influence of the optical instrument on the measurement.
The main facts which a theory of stellar aberration in the earth-based frame of reference must explain are (1) the annual apparent motion of the fixed stars about their locations and (2) the null apparent aberration of rotating binary systems (they exhibit aberrations not different from other stars). We present here a theory which accounts for all these, and in addition gives new results. All earth-based experiments can be explained on the basis of the effect of the measuring instrument (i.e. physical unavoidable influence of the telescope on the measurement) and the acceleration of traveling earth-based observer relative to the fixed stars.
In the Chapter 8 we analyze the potential for exploiting earth-based sources in order to confirm the predictions of the relativistic aberration of light theory. We discussed a simple scaling model for the stellar aberration. The motion of the stars with respect to the earth is never followed by any aberration. The aberration shift exist even in the case when star moves with the same velocity as the earth. We obtained a condition for optical similarity between the aberration of light from a distant star (which is moving with the earth velocity) and from the earth-based incoherent source. The proposed method of measuring the angle of aberration involves the use of earth-based sources and have a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth rotation axis. The apparent position of the source image is thus always a little displaced in the direction of the earth motion around the sun at that moment, and hence describe a small elliptical figure during the annular revolution the earth around its axis. In principle, records could be taken over a period of one day.
It is generally believed that the special relativity is reciprocal theory. The aberration of light problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers. In fact, without looking at anything external to the earth-based accelerated frame, one could determine the speed of the earth with respect to the sun-based frame by means of aberration of earth-based incoherent source measurements. No one has ever done such experiment, but we know what would happen from the astronomical observations.
Now many people who learn special relativity in the usual way find this disturbing. The argument that aberration shift must be symmetrical runs some thing like this:
(1) It is a basic principle of the special theory of relativity that the metric supplies all information about the physics of the situation, as described in the given coordinates.
(2) Information about the direction of observer acceleration is recorded in the crossed term of the Langevin metric.
(3) It is always possible to chose such variables, in which metric of the accelerated frame will be diagonal. This is a consequence of the (pseudo-Euclidean) geometry of space-time.
(4) The validity of Maxwell’s laws in all inertial frames implies that light propagates with the same velocity $c$ in all inertial frames. It implies the elimination of the privilege frame, so meaning that the notion of absolute motion does not have physical content at all.
We discuss this subject in the Chapter 9. In the argument above, steps (2) and (3) are correct, but step (1), and its consequence (4), are wrong. Argument (3) says that all inertial frames are equivalent with regard to physical laws, not with regard to physical facts. At first glance, after the diagonalization of the Langevin metric, we have the symmetry between inertial frames. Where does the asymmetry comes from? The electrodynamics equations are not supplies all information about the physics of the situation. To solve the electrodynamics equations it is necessary to determine the initial conditions. The time after diagonalization is readily obtained by introducing the time offset factor. This time shift has the effect of rotation of the plane of simultaneity. As a consequence of this, the plane wavefront rotates in the accelerated frame after metric diagonalization. Now the information about the direction of observer acceleration is recorded in the initial conditions (radiation wavefront orientation).
One can conclude that not all is relative in relativity, because this theory also contains some features which are absolute. Many people would like to think that since in the Lorentz coordinatization the time and distance have direct physical meaning, there should be some physical (dynamical) reason for this wavefront rotation. They would think that dynamics, based on the physical fields, is actually hidden in the language of pseudo-Euclidean geometry. However, it is a dynamical line of arguments that explains this paradoxical situation with the asymmetry between inertial and accelerated reference frames. A resolution of the asymmetry paradox identifies inertial (pseudo-gravitational) force within the accelerated system as the agency of asymmetry. The principle of equivalence can be applied to solve non-inertial kinematic problems with dynamical methods. Wavefront rotation associated with the transformation from the inertial frame to the accelerated (with respect to the fixed stars) frame may be regarded as a result of the action of pseudo-gravitational potential gradient during the acceleration process. See Chapter 9 for more ideas along this line.
There are several point to be made about the above results. One interesting question is, why we are not discuss about the aberration of light radiated by a laser? In regard to light aberration one should differentiate between that from the incoherent source and that from the laser source. One could naively expect that the rotation of the earth around the sun produces aberration in an amount large enough to be taken into account in precise observation work using laser as a earth-based light source. Let us stress that the aberration of light cannot be measured using laser. In fact, the electromagnetic wave travels in the laser resonator forward and back reflecting from the mirrors. It could be said that the asymmetry cancels during the (round-trip-to-round-trip) evolution of the radiation in the optical resonator. We would like to emphasize a very important difference between laser and incoherent source. The radiation of incoherent source clearly depends on an initial conditions and there is influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer. In contrast, when we are dealing with optical resonator the (3-space) boundary conditions define the energy propagation direction.
One finds many books which say that we cannot distinguish by any experiment which observer remains ”at rest” during the acceleration, because according to the principle of relativity only relative motion has any physical meaning. It follows that any observed asymmetry would lead to a contradiction with the principle of relativity. This argument is wrong. We discuss the apparent conflict between aberration of light and the principle of relativity in the Chapter 12. It is demonstrated that there is no conflict between the fundamental structure of special relativity on the one hand, and the aberration of light phenomena. Not only in Einstein writings, but in every textbook on special theory of relativity can be found the formulation of principle of special relativity as follows: the laws of physics are the same in all inertial frames. Principle of special relativity is irrelevancy of velocity with regard to physical laws, not with regard to anything.
It is generally accepted that covariant equations must have covariant solutions. This statement presented in most textbooks and is obviously incorrect. For instance, from the mathematical viewpoint, the Lorentz covariance of electrodynamics equations is insufficient to guarantee the covariant solutions. We only wish to emphasize here the following point. The principle of special relativity, which says that all physical equations must be invariant under Lorentz transformation (i.e. the pseudo-Euclidean geometry of space-time as an axiom of the theory), does not dictate the reciprocal symmetry of nature.
Authors of textbooks are mistaken in their belief that all inertial frames are equivalent. Relativistic (i.e. reciprocal) symmetry is usually identifies with symmetry under Lorentz boosts. At the same time it is common knowledge that Lorentz boosts alone do not make up a group. The product of two Lorentz boosts in different directions is equal to the product of a pure boost and the spatial (Wigner) rotation. From the mathematical viewpoint, when we are dealing with transformations and set of these transformations does not form a group we have no symmetry between inertial frames. The inertial frames are characterized by motion at constant velocity relative to the fixed stars. There is a real difference between accelerated ( with respect to the fixed stars) inertial frame and the inertial frame without accelerational history and this difference ( in the case of the aberration of light phenomena) is closely associated with the Wigner rotation effect.
It should be note that all well known methods to test the special relativity are second order optical (interference) measurements . The cardinal example is given by the Michelson-Morley experiment. The light from a source is split by a suitable apparatus into two beams and the two overlapping beams produce an interference pattern. The two beam originate in the same source. It is not possible to detect the orbital velocity using second order (interference) experiments. Phase is 4D invariant (i.e. it is simply number). It is independent on the chosen inertial frame and on the chosen coordinatization. This is a consequence of the geometry of space-time. In the interference experiments neither the Doppler effect nor the aberration of light exist separately as well defined physical phenomena and only phase is meaningful quantity. This discussion clearly show that the special relativity calculations yields the observed null fringe shift in the earth-based frame. A close examination of all optical second order experiments inside an inertial frame shows that phenomena appear to be independent of the uniform motion relative to the fixed stars. The usual formulation of the theory of relativity deals with formulated above principle of irrelevancy of velocity, but it should be understood only in a limiting sense when we are dealing with optical second order measurements.
There are a number of other second order experiments. They are not connected with light interference phenomena. The question arises whether it is possible to determine experimentally the state of motion of the initial inertial frame in the accelerated frame by means of time dilation effect for moving reference clock. We demonstrate that slowly down of the accelerated (with respect to the fixed stars) physical clock does not depend on the choice of reference system, in which this effect is measured. In other words, the time dilation effect is absolute, but not relative. It is a fundamental property of the pseudo-Euclidean space-time.
Now let us return to metric diagonalization. In textbooks, time dilation is usually derived from the diagonal (Minkowski) metric. After diagonalization, according to textbooks, the coordinates in inertial frames should be related by the Lorentz transformation. At first site the problem is complete symmetrical and the time dilation effect is relative. Something is evidently wrong. In the Chapter 9 we investigate the connection between time dilation effect and the Langevin metric. Perhaps the only way to demonstrate that the time dilation effect is absolute is to take a mathematical view. From a mathematical viewpoint, the argument that time dilation effect is absolute looks something like the following. It is obvious that metric tensor must be a continuous quantity. The Langevin metric is found by matching the accelerated frame and inertial frame metric tensors. For smooth tailoring of the metric tensor, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the Galilean boost. In this case the metric of inertial frame would pass into the Langevin metric of the accelerated frame. After diagonalization, the metric tensor in the accelerated frame must abruptly change from the Langevin value to the Minkowski (diagonal ) value. We must conclude that when we are dealing with Lorentz transformation and this transformation does not preserve continuity of the metric tensor we have no symmetry between inertial frames. Any change of velocity, or any acceleration (with respect to the fixed stars) has an absolute meaning.
In the Chapter 9-12 we will take up in more detail the absoluteness of acceleration; here we only wished to show how naturally the special relativity leads to the concept of absolute velocity. The formalism of relativistic physics is associated with absolute space-time structure. This structure must be associated with initial inertial frame, which incidentally provides an absolute rest frame. There is no relativity of the uniform motions and we can determine the particle absolute velocity in the initial inertial frame. The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of unified space-time model. In special relativity the absolute space and the absolute time of Newton are fused into absolute space-time.
The proper time interval may be identified with particle lifetime, atomic transition periods, nuclear half-life. Until now, all test of time dilation with atomic clocks in the earth-based frame have involved circle path geometry. Indeed, the clocks flown in the commercial jets. When the clocks returned to the laboratory they were compared with the reference clock in the same point at rest with respect to the earth. According to the special relativity, in any time dilation experiment which probe the influence of the orbital velocity one must use the linear (one-way) path geometry. These are experiments of decays of rapidly moving unstable particles. However, the modern experimental techniques are not sensitive enough to register the effect of orbital velocity on the average distance traveled by a disintegrated particle.
In contrast, we describe (see Chapter 8) a simple first order optical experiment within the means of most undergraduate departments that illustrates many of the principles of current research in the foundations of special relativity. It directly demonstrates that one could determine the speed of the earth with respect to the sun-based frame by means of aberration of earth-based incoherent source measurements.
The second part of the book (Chapter 13 - 18) deals with accelerator physics. In the Chapter 13 we describe the particle dynamics method in the Lorentz lab frame using lab time $t$ as evolution parameter. The study of relativistic particle motion in a constant magnetic field according to usual accelerator engineering, is intimately connected with the old (Newtonian) kinematics: the Galilean vectorial law of addition of velocities is actually used. A non-covariant approach to relativistic particle dynamics is based on the absolute time coordinatization, but this is actually a hidden coordinatization. The absolute time synchronization convention is self-evident and this is the reason why this subject is not discussed in accelerator physics.
According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity. In other words, in this standard approach, introducing as only modification to the classical case the relativistic mass, time differ from space. In convention particle tracking a particle trajectory can be seen from the lab frame as result of successive Galileo boosts that track the motion of the accelerated particle. There is a reason to prefer the non-covariant way within the framework of dynamics only. In this approach we have no mixture of positions and time. This (3+1) dimensional non-covariant particle tracking method is simple, self-evident, and adequate to the laboratory reality. It can be demonstrated that there is no principle difficulty with the non-covariant approach in mechanics and electrodynamics. It is perfectly satisfactory. It does not matter which transformation is used to describe the same reality. What matter is that, once fixed, such convention should be applied and kept in a consistent way for both dynamics and electrodynamics. In fact, the light (group) velocity transforms as a particle velocity. We cannot take old Newton’s kinematics for mechanics and Einstein’s kinematics for electrodynamics.
The common mistake made in accelerator physics is connected with the incorrect algorithm for solving the electromagnetic field equations. If one wants to use the usual Maxwell’s equations, only the solution of the dynamics equations in covariant form (i.e. in Lorentz coordinates) gives the correct coupling between the Maxwell’s equations and particle trajectories in the lab frame.
It is interesting to note that, the use of the conventional coupling of Maxwell’s equations and corrected Newton’s equation for the calculation of the radiation from a moving source does not necessarily leads to mistake. For rectilinear motion of the source and the emitted light beam, the non-covariant and covariant approaches produce the same trajectories, and Maxwell’ equations are compatible with the result of conventional particle tracking. This method was incorrectly extended to the case of non-collinear geometry.
First, we examine the reasoning presented in textbooks. It is generally believed that the magnetic field, in the Lorentz coordinatization, is only capable altering the direction of motion, but not the speed of the electron. Nevertheless, there is argument against this commonly accepted derivation of the composition of velocities. Let us to consider composition of two velocities when they are not lying on a straight line. In conventional particle tracking (similar to the Newton mechanics), one composes them geometrically by the method of parallelogram with velocities regarded as the usual vectors.
Only solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle trajectories in the lab frame. There are two difficulties in solving problems concerning covariant particle tracking. First of all, conventional particle tracking in the single inertial frame is actually based on classical Newton mechanics without reference to Lorentz transformations. Another difficulty is the confusion surrounding the combination of non-parallel velocities in special relativity. The standard textbook presentation of the relativistic velocity transformation is based on the hidden assumption that $(x^{\prime},y^{\prime},z^{\prime})$ axes of the moving observer and $(x,y,z)$ axes of the lab observer are parallel. In other words, it is assumed that observers have common 3-space. This is misconception. The fact that two observers with different trajectories in the Lorentz coordinatization have different 3-spaces is not considered in textbooks. First of all, there is no absolute notion of simultaneity in the theory of special relativity. In the case of the relativity of simultaneity we have a mixture - of positions and time. In other words, in the space measurement of one observer there is mixing of space and time, as seen by the other. Therefore, there is no notion of an instantaneous 3-space.
In textbooks on special relativity, the relativistic composition of velocities is studied analytically from the composition of Lorentz boosts. Authors of textbooks believe that the relativity principle requires that velocity addition law gives the composition law of a group. The non-associativity of the composition law of the non-colinear velocities is not widely known, although it is an exercise in elementary algebra to prove non-associativity. It is therefore impossible to get covariant trajectory, if we have nothing but Lorentz boosts. Something else has to be added. Lorentz boosts do not form a group due to presence of the Wigner rotation. It is comes out quite natural that breakdown of commutativity and associativity in Einstein addition, turn out to be Wigner rotation.
We emphasize that non-collinear velocities addition in the Lorentz coordinatization is regulated by the Wigner rotation. One of the consequences of non-commutativity of non-collinear Lorentz boosts is not the existence of a common ordinary space. Our result is at odds with the prediction from textbooks. The commonly accepted derivation of the composition of perpendicular velocities based on the use the relation $dt^{\prime}=dt/\gamma$ and does not account for the Wigner rotation (i.e. mixture of transverse positions and time). Only the solution of the dynamics equations in covariant form (accounting for the Wigner rotation in the transformation of non-collinear velocities) gives the correct coupling between the Maxwell’s equations and particles trajectories in the lab frame. We illustrate error in standard coupling fields and particles by considering the relatively simple example, wherein the essential physical features are not obscured by unnecessary mathematical difficulties. We hope that the present example will draw wide attention to the central role that Wigner rotation plays in the transformation of non-collinear velocities.
In the Chapter 13 we examine what parts of the dynamics theory depend on the choice of synchronization convention and what parts do not. The present analysis shows the difference between the notions of path and trajectory. The path is rather a purely geometrical notion. The trajectory of a particle conveys more information about its motion because every position is described additionally by the corresponding time instant. The path has exact objective meaning. In fact, the curvature radius of the path in the magnetic field (and consequently the three-momentum) has obviously an objective meaning, i.e. is convention-invariant. In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle is convention dependent and has no exact objective meaning. Dynamics theory contains a particle trajectory that we do not need to check directly, but which is used in the analysis of electrodynamics problem.
In the Chapter 14 we present a critical reexamination of radiation theory. To evaluate radiation fields arising from an external sources, we need to know the particle velocity $\vec{v}$ and the position $\vec{x}$ as a function of the lab frame time $t$. As discussed above, one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particles moving along covariant trajectory. For an arbitrary parameter $v/c$ covariant calculations of the radiation process is very difficult. There are, however, circumstances in which calculations can be greatly simplified. As example of such circumstance is a non-relativistic radiation setup. The reason is that the non-relativistic assumption implies the dipole approximation which is of great practical significance. In accounting only for the dipole part of the radiation we neglect all information about the electron trajectory. That means that the dipole radiation does not show any sensitivity to the difference between covariant and non-covariant particle trajectories. But that is only the first and most practically important term. The calculation of higher order corrections to the dipole radiation approximation requires detailed information about the electron trajectory. Obviously, in order to calculate the correction to the dipole radiation, we will have to use the covariant trajectory and not be satisfied with the non-covariant approach.
Accelerator physicists, who try to understand the situation related to the use of the theory of relativity in the synchrotron radiation phenomena, are often troubled by the fact that the difference between covariant and non-covariant particle trajectories was never understood, and that nobody realized that there was a contribution to the synchrotron radiation from relativistic kinematics effects. Accelerator physics was always thought in terms of the old (Newtonian) kinematics that is not compatible with Maxwell’s equations. At this point, a reasonable question arises: since storage rings are designed without accounting for the relativistic kinematics effects, how can they actually operate? In fact, electron dynamics in storage ring is greatly influenced by the emission of radiation. We answer this question in great detail in the Chapter 15. The main emphasis of this chapter is on spontaneous synchrotron radiation from bending magnets and undulators.
Similar to the non-relativistic asymptote, the ultrarelativistic asymptote also provides the essential simplicity of the covariant calculation. The reason is that the ultrarelativistic assumption implies the paraxial approximation. Note that the geometry of the electron motion in a bending magnet has a cylindrical symmetry. There are a number of remarkable effects which are a consequence of the cylindrical symmetry and the paraxial approximation. We demonstrate that there is no needs to use covariant particle tracking for derivation of the bending magnet radiation. However, there is one situation where the conventional theory fails. The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion along the magnetic field.
One way to demonstrate incompatibility between the standard approach to relativistic electrodynamics, which deals with the usual Maxwell’s equations, and particle trajectories calculated by using non-covariant particle tracking, is to make a direct laboratory test of synchrotron radiation theory. In other words, we are stating that, despite the many measurements done during decades, synchrotron radiation theory is not an experimentally well-confirmed theory. We analyze the potential for exploiting synchrotron radiation sources in order to confirm the predictions of corrected synchrotron radiation theory. Proposed experiments have never been performed at synchrotron facilities.
The theory of relativity shows that the relativity of simultaneity, which is a most fundamental relativistic kinematics effect, is related with extended relativistic objects. But up to 21 st century there were no macroscopic objects possessing relativistic velocities, and there was a general belief that only microscopic particles in experiments can travel at velocities close to that of light. The 2010s saw a rapid development of new laser light sources in the X-ray wavelength range. An X-ray free electron laser (XFEL) is an example where improvements in accelerator technology makes it possible to develop ultrarelativistic macroscopic objects with an internal structure (modulated electron bunches), and the relativistic kinematics plays an essential role in their description.
In Chapter 16 we present a critical reexamination of existing XFEL theory. It is mainly addressed to readers with limiting knowledge of accelerator physics. We study the production of coherent undulator radiation by modulated ultrarelativistic electron beam. Fortunately, the principle of production of coherent undulator radiation does not require specific knowledge of undulator radiation theory presented in the Chapter 15 and can be explained in a very simple way. Relativistic kinematics enters XFEL physics in a most fundamental way through the rotation of the modulation wavefront, which, in ultrarelativistic approximation, is closely associated to the relativity of simultaneity. When the trajectories of particles calculated in the Lorentz reference frame (i.e. an inertial frame where Einstein synchronization procedure is used to assign values to the time coordinate) they must include relativistic kinematics effects such as relativity of simultaneity. In the ultrarelativistic asymptote, the orientation of the modulation wavefront, i.e the orientation of the plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the modulated electron beam is treated using Lorentz coordinates.
We should remark that Maxwell’s equations are valid only in Lorentz reference frames. Einstein’s time order should obviously be applied and kept in consistent way both in dynamics and electrodynamics. It is important at this point to emphasize that the theory of relativity dictates that a modulated electron beam in the ultrarelativistic asymptote has the same kinematics, in Lorentz coordinates, as a laser beam. According to Maxwell’s equations, the wavefront of a laser beam is always orthogonal to the propagation direction. Experiments show that this prediction is, in fact, true. The theory of relativity as a theory of space-time with pseudo-Euclidean geometry has had more than hundred years of development, and rather suddenly it has begun to be fully exploited in practical ways in XFEL physics.
As known, a composition of noncollinear Lorentz boosts does not result in a different boost but in a Lorentz transformation involving a boost and a spatial rotation, the Wigner rotation (it is often called the ”Thomas rotation”). The results for the Wigner rotation in the Lorentz lab frame obtained by many experts on special relativity are incorrect. They overestimate the angle of the Wigner rotation by a relativistic factor $\gamma$ compared to its real value, and the direction of the rotation is also determined incorrectly. It should be note that the correct result was obtained in the works of several authors, which were published more than half century ago but remained unnoticed against the background of numerous incorrect works. To deduce the expression for Wigner rotation is the main purpose of Chapter 17. We demonstrate that what are usually considered the advanced parts of the theory of relativity are, in fact, quite simple.
The Wigner rotation is also relates the angular rotation velocity (Thomas precession) of the spin of an elementary particles following a curvalinear orbit. The kinematics tools needed to study the motion of charged elementary spinning particles in the storage ring are the main topic of final chapter of the book. In 1959, Bargman, Michel, and Telegdi (BMT) proposed a consistent relativistic theory for the dynamics of the spin as observed in the lab frame, which was successfully tested in experiments. It is commonly believed that the BMT equation contains the standard (and incorrect) result for the Wigner rotation in the Lorentz lab frame. There are many physicists who have already received knowledge about the Wigner (Thomas) rotation from well-known textbooks and who would say, ”The extremely precise measurements of the magnetic-moment anomaly of the electron made on highly relativistic electrons are based on the BMT equation, of which the Wigner rotation is an integral part, and can be taken as experimental confirmations of the standard expression for the Wigner rotation” This misconception about experimental test of the incorrect expression for the Wigner rotation in the lab frame is common and pernicious. We should remark that the results in the Bargmann-Michel-Telegdi paper were obtained by the method of semi-classical approximation of the Dirac equation. The Wigner rotation was not considered at all, because the Dirac equation allow obtaining the solution for the total particle’s spin motion without an explicit splitting it into the Larmor and Wigner parts. We demonstrate that the notion that the standard (incorrect) result for the Wigner rotation as an integral part of the BMT equation in most texts is based, in turn, on an incorrect physical argument.
A large number of incorrect expressions for the Wigner (Thomas) rotation can be found in the literature. Let us discuss the reasons that underlie the derivation of the incorrect expressions. The situation relating to the use of the Wigner (Thomas) rotation theory in the relativistic spin dynamics is complex. The starting point of Bargman, Michel and Telegdi was the particle rest frame and the equation of motion for particle angular momentum, which they generalized to the Lorentz lab frame and then transformed back to the rest frame. This brings up an interesting question: Why it is convenient to transform this equation to the rest frame as of that instant? The approach in which we deal with the proper spin is much preferred in the experimental practice due to clear physical meaning of the (3D) spin vector $\vec{s}$. Unlike (3D) momentum vector, which has definite components in each reference frame, angular momentum is defined only in one particular reference frame. It does not transform. Any statement about it refers to the rest frame as of that instant. If we say that in the lab frame the spin of a particle makes the angle $\phi$ with its velocity, we mean that in the particle’s rest frame the spin makes this angle with the line motion of the lab frame. The analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard (Einstein) special relativity taking advantage of well known relativistic kinematic effects. In the case of spin dynamics one has no occasion to make measurements with non-inertial devices. Common textbooks presentations of the spin dynamics use the approach which deals only with observations of the inertial (lab) observer.
The conventional method used to explain the spin dynamics in the lab frame is very unusual. First, we already pointed that the laws of physics in the lab frame are able to account for all physical phenomena, including the observation made by non-inertial observer. In particular, it is possible to determine the result of spin orientation measurements with respect to the lab frame axes in the proper frame. In this case, when viewed from the lab frame, the interpretation of the spin rotation experiments in the lab frame is simple and may be derived from the prediction of the proper observer as concerns the measurement made by the lab observer. Therefore, the spin rotation measurement by a polarimeter in the lab frame is interpreted with viewpoint of the proper observer as viewing this of the lab observer. This complexity is one of the reason why authors of textbooks obtained an incorrect expression for the Wigner rotation.
It is also important to elucidate the reason why did the error in the Wigner rotation theory remain so long undetected. The main conclusion drawn from the review of the literature is as follows. Wigner (Thomas) rotation is generally presented only in the context of spin kinematics as a peculiar effect of special relativity. This is supposedly the reason why until recently no researches have noticed so serious a discrepancy between the results of different works.
We emphasize that the Wigner rotation is very basic relativistic phenomenon. It is as basic as the time dilation and length contraction. For example, we demonstrate by elementary means that relativistic (non-collinear) velocities addition is regulated by the Wigner rotation. Now let us consider the aberration theory. The theory of special relativity shows that the aberration effect in the rotating earth-based frame, as viewed from the inertial sun-based frame (in the Lorentz coordinatization), presents a kinematic effect of special theory of relativity. In this case, when viewed from the sun-based frame, the aberration image shift in the earth-based frame is regulated by the Wigner rotation. According to the Wigner rotation theory, two observers with different trajectories have different 3-spaces. In contrast, the accepted in previous literature incorrect assumption that an earth-based observer and an inertial (e.g. sun-based) observer have common 3-space has always been considered obvious.
2 A Critical Survey of the Present Approaches to Special Relativity
2.1 What is Special Relativity?
The laws of physics are invariant with respect to Lorentz transformations. This is a restrictive principle and does not determine the exact form of the dynamics in question. Understanding the postulates of the theory of relativity is similar to understanding energy conservation: at first we learn this as a principle and later on we study microscopic interpretations that must be consistent with this principle. For any system to which the energy conservation principle can be applied, a deeper theory should exist which yields insight into the detailed physical processes involved. Of course, this deeper theory must lead to energy conservation.
The principle of conservation of energy is very useful in making analyses without knowing all the formulas of the fundamental theory. A methodological analogy with the postulates of the special relativity emerges by itself. Suppose we do not know why a muon disintegrates, but we know the law of decay in the Lorentz rest frame. This law would then be a phenomenological law. The relativistic generalization of this law to any Lorentz frame allows us to make a prediction on the average distance traveled by a muon. In particular, when a Lorentz transformation of the decay law is tried, one obtains the prediction that after the travel distance $\gamma v\tau_{0}$, the population in the lab frame would be reduced to 1/2 of the origin population. We may interpret this result by saying that, in the lab frame, the characteristic lifetime of a particle has increased from $\tau_{0}$ to $\gamma\tau_{0}$.
However, the theory of relativity is necessary incomplete. Constructive (microscopic) theories like electrodynamics or quantum field theory provide more insight into the nature of things than restrictive theories like special relativity. Relativistic kinematics is only an interpretation of the behavior of the dynamical matter fields in the view of different observers. The point is that one can, in principle compute any relativistic quantity directly from the underlying theories of matter without involving relativity at all. For example, muons in motion behave relativistically because the field forces that are responsible for the muon disintegration satisfy quantum field equations that are Lorentz covariant. Of course, in the ”microscopic” approach to relativistic phenomena, Lorentz covariance of all the fundamental laws of physics remains, similarly to energy conservation, an unexplained fact, but all explanation must stop somewhere.
2.2 Different Approaches to Special Relativity
In literature, three approaches to special relativity are discussed: Einstein’s approach, the usual covariant approach, and the space-time geometric approach.
Einstein formulation is based on postulates: the principle of relativity and the constancy of the velocity of light. The usual covariant formulation of the theory of relativity deals with the pseudo-Eucledian space-time geometry and with the invariance of interval $ds$, but it is understood only in a limited sense when the metric is strictly diagonal. Assuming diagonality of the metric we also automatically assume Lorentz coordinates, and that different inertial frames are related by Lorentz transformations.
In space-time geometric approach, primary importance is attributed to the geometry of space-time; it is supposed that the geometry of space-time is a pseudo-Euclidean geometry in which only 4-tensors quantities do have real physical meaning. In this most general approach the principle of relativity in contrast to Einstein formulation of the special relativity is a simple consequence of the space-time geometry. Since the space-time geometric approach deals with all possible choices of coordinates of the chosen reference frames, the second Einstein postulate referred to the constancy of the coordinate velocity of light does not hold in this formulation of the theory of relativity. Only in Lorentz coordinates, when Einstein’s synchronization of distant clocks and Cartesian space coordinates are used, the coordinate speed of light is isotropic and constant.
2.2.1 The Usual Einstein’s Approach
Traditionally, the special theory of relativity is built on the principle of relativity and on a second additional postulate concerning the velocity of light:
1. Principle of relativity. The laws of nature are the same (or take the same form) in all inertial frames
2. Constancy of the speed of light. Light propagates with constant velocity $c$ independently of the direction of propagation, and of the velocity of its source.
The constancy of the light velocity in all inertial systems of reference is not a fundamental statement of the theory of relativity. The central principle of special relativity is the Lorentz covariance of all the fundamental laws of physics. It it important to stress at this point that the second ”postulate”, contrary to the view presented in textbooks, is not a separate physical assumption, but a convention that cannot be the subject of experimental tests.
Assuming postulate 2 on the constancy of the speed of light in all inertial frames we also automatically assume Lorentz coordinates, and that different inertial frames are related by Lorentz transformations. In other words, according to such limiting understanding of the theory of relativity it is assumed that only Lorentz transformations must be used to map the coordinates of events between inertial observers.
2.2.2 The Usual Covariant Approach
In the usual covariant approach the special of relativity is understood as the theory of space-time with pseudo-Euclidean geometry. Quantities of physical interest are represented by tensors in a four-dimensional space-time, i.e. by covariant quantities, and the laws of physics are written in manifestly covariant way as four-tensor equations.
In order to develop space-time geometry, it is necessary to introduce a metric or a measure $ds$ of space-time intervals. The type of measure determines the nature of the geometry. Any event in the usual covariant approach is mathematically represented by a point in space-time, called world-point. The evolution of a particle is, instead, represented by a curve in space-time, called world-line. If $ds$ is the infinitesimal displacement along a particle world-line, then
$ds^{2}=c^{2}dT^{2}-dX^{2}-dY^{2}-dZ^{2}~{},$ (1)
where we have selected a special type of coordinate system (a Lorentz coordinate system), defined by the requirement that Eq. (1) holds.
To simplify our writing we will use, instead of variables $T,X,Y,Z$, variables $X^{0}=cT,~{}X^{1}=X,~{}X^{2}=Y,~{}X^{3}=Z$. Then, by adopting the tensor notation, Eq. (1) becomes $ds^{2}=\eta_{ij}dX^{i}dX^{j}$, where Einstein summation is understood. Here $\eta_{ij}$ are the Cartesian components of the metric tensor and by definition, in any Lorentz system, they are given by $\eta_{ij}=\mathrm{diag}[1,-1,-1,-1]$, which is the metric canonical, diagonal form. As a consequence of the space-time geometry, Lorentz coordinates systems are connected by Lorentz transformations.
Physical quantities are represented by space-time geometric (tensor) quantities. When some basis is introduced, the representation of a tensor as geometric quantity comprise both components and basis. In usual covariant approach, one only deals with the basis components of tensors in the Lorentz coordinates i.e. with the case when the basis four-vectors are orthogonal. As a result one deals only with four-tensor equations of physics written out in the component form.
However, the concept of a tensor in the usual covariant approach is given in terms of the transformation properties of its components. For example in the usual covariant approach the electromagnetic ”tensor” $F^{\mu\nu}$ is actually not a tensor since $F^{\mu\nu}$ are only components implicitly taken in standard (orthogonal) basis. The components are coordinate quantities and they do not contain the whole information about the physical quantity, since a basis of the space-time is not included. This is no problem only in the limiting case when transformations from one orthogonal basis to another orthogonal basis are selected i.e. only assuming that Lorentz transformations must be used to map the coordinates of events. According to the usual covariant approach, another transformations from standard to non standard (not orthogonal) basis, like Galilean transformations, are ”incorrect”.
2.2.3 The Space-Time Geometric Approach
We emphasize the great freedom one has in the choice of a Minkovski space-time coordinatization. The space-time continuum, determined by the interval Eq. (1) can be described in arbitrary coordinates and not only in Lorentz coordinates. In the transition to arbitrary coordinates, the geometry of four-dimensional space-time obviously does not change, and in the special theory of relativity we are not limited in any way in the choice of a coordinates system. The space coordinates $x^{1},x^{2},x^{3}$ can be any quantities defining the position of particles in space, and the time coordinate $x^{0}$ can be defined by an arbitrary running clock. The components of the metric tensor in the coordinate system $x^{i}$ can be determined by performing the transformation from the Lorentz coordinates $X^{i}$ to the arbitrary variables $x^{j}$, which are fixed as $X^{i}=f^{i}(x^{j})$. One then obtains
$ds^{2}=\eta_{ij}dX^{i}dX^{j}=\eta_{ij}\frac{\partial X^{i}}{\partial x^{k}}\frac{\partial X^{j}}{\partial x^{m}}dx^{k}dx^{m}=g_{km}dx^{k}dx^{m}~{},$ (2)
This expression represents the general form of the pseudo-Euclidean metric. In textbooks and monographs, the special theory of relativity is generally presented in relation to an interval $ds$ in the Minkowski form Eq.(1), while Eq.(2) is ascribed to the theory of general relativity.
However, in the space-time geometric approach, special relativity is understood as a theory of four-dimensional space-time with pseudo-Euclidean geometry. In this formulation of the theory of relativity the space-time continuum can be described equally well from the point of view of any coordinate system, which cannot possibly change $ds$. At variance, the usual formulation of the theory of relativity also deals with the invariance of $ds$, but it is understood only in a limited sense when the metric is strictly diagonal.
Common textbook presentations of the special theory of relativity use the Einstein approach or, as generalization, the usual covariant approach which deals, as discussed above, only with components of the 4-tensors in specific (orthogonal) Lorentz basis. The fact that in the process of transition to arbitrary coordinates the geometry of the space-time does not change, is not considered in textbooks. As a consequence there is a widespread belief among experts that a transformation from an orthogonal Lorentz basis to a non orthogonal basis is incorrect, while a Lorentz transformation, which is a transformation from an orthogonal Lorentz basis to another orthogonal Lorentz basis, is correct. This is not true. We can describe physics in any arbitrary coordinates system. The different transformations of coordinates only correspond to a change in the way of components of 4-tensors are written, but not influence of 4-tensors themselves. Although the Einstein synchronization i.e. Lorentz coordinates choice, is preferred by physicists due to its simplicity and symmetry, it is nothing more ”physical” than any other. A particularly very unusual choice of coordinates, the absolute time coordinate choice, will be considered and exploited in this book.
2.3 Myth About the Incorrectness of Galilean Transformations
The use of Galilean transformations within the theory of relativity requires some special discussion. Many physicists still tend to think of Galilean transformations (which is actually a transformations from an orthogonal Lorentz basis to a non-orthogonal basis) as old, incorrect transformations between spatial coordinates and time. A widespread argument used to support the incorrectness of Galilean transformations is that they do not preserve the form-invariance of Maxwell’s equations under a change of inertial frame. This idea is a part of the material in well-known books and monographs (1).
Authors of textbooks are mistaken in their belief about the incorrectness of Galilean transformations. The special theory of relativity is the theory of four-dimensional space-time with pseudo-Euclidean geometry. From this viewpoint, the principle of relativity is a simple consequence of the space-time geometry, and the space-time continuum can be described in arbitrary coordinates (2). Therefore, contrary to the view presented in most textbooks, Galilean transformations are actually compatible with the principle of relativity although, of course, they alter the form of Maxwell’s equations.
2.4 Myth About the Non-Relativistic Limit of the Lorentz Transformations
It is generally believed that a Lorentz transformation reduces to a Galilean transformation in the non-relativistic limit. We state that this typical textbook statement is incorrect and misleading. Kinematics is a comparative study which requires two coordinate systems, and one needs to assign time coordinates to the two systems. Different types of clock synchronization provide different time coordinates. The convention on the clock synchronization amounts to nothing more than a definite choice of the coordinate system in an inertial frame of reference in Minkowski space-time. Pragmatic arguments for choosing one coordinate system over another may therefore lead to different choices in different situations. Usually, in relativistic engineering, we have a choice between absolute time coordinate and Lorentz time coordinate. The space-time continuum can be described equally well in both coordinate systems. This means that for arbitrary particle speed, the Galilean coordinate transformations well characterize a change in the reference frame from the lab inertial observer to a comoving inertial observer in the context of the theory of relativity. Let us consider the non relativistic limit. The Lorentz transformation, for $v/c$ so small that $v^{2}/c^{2}$ is neglected can be written as $x^{\prime}=x-vt$, $t^{\prime}=t-xv/c^{2}$. This infinitesimal Lorentz transformation differs from the infinitesimal Galilean transformation $x^{\prime}=x-vt$, $t^{\prime}=t$. The difference is in the term $xv/c^{2}$ in the Lorentz transformation for time, which is a first order term.
We only wish to emphasize here the following point. An infinitesimal Lorentz transformation differs from Galilean transformation only by the inclusion of the relativity of simultaneity, which is the only relativistic effect that appears in the first order in $v/c$. All other higher order effects, that are Lorentz-Fitzgerald contraction, time dilation, and relativistic correction in the law of composition of velocities, can be derived mathematically, by iterating this infinitesimal transformation (3).
The main difference between the Lorentz coordinatization and the absolute time coordinatization is that the transformation laws connecting coordinates and times between relatively moving systems are different. It is impossible to agree with the textbook statement that there is reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to Galilean relation $t^{\prime}=t$ in the non-relativistic limit. This would mean that in the non-relativistic limit infinitesimal Lorentz transformations are identical to infinitesimal Galilean transformations. This statement is absurd conclusion from a mathematical standpoint. The essence of Lorentz (or Galilean) transformations consists in their infinitesimal form: relativistic kinematics effects cannot be found by the mathematical procedure of iterating the infinitesimal Galilean transformations.
2.5 Myth about the Constancy of the Speed of Light
It is generally believed that experiments show that the speed of light in vacuum is independent of the source or observer (4). This statement presented in most textbooks and is obviously incorrect. The constancy of the speed of light is related to the choice of synchronization convention, and cannot be subject to experimental tests (5).
In fact, in order to measure the one-way speed of light one has first to synchronize the infinity of clocks assumed attached to every position in space, which allows us to perform time measurements. Obviously, an unavoidable deadlock appears if one synchronizes the clocks by assuming a-priori that the one-way speed of light is $c$. In fact, in that case, the one-way speed of light measured with these clocks (that is the Einstein speed of light) cannot be anything else but $c$: this is because the clocks have been set assuming that particular one-way speed in advance.
Therefore, it can be said that the value of the one-way speed of light is just a matter of convention without physical meaning. In contrast to this, the two-way speed of light, directly measurable along a round-trip, has physical meaning, because round-trip experiments rely upon the observation of simultaneity or non-simultaneity of events at a single point in space and not depends on clock synchronization convention. All well known methods to measure the speed of light are, indeed, round-trip measurements. The cardinal example is given by the Michelson-Morley experiment: this experiment uses, indeed, an interferometer where light beams are compared in a two-way fashion.
Because of our usage of Galilean transformations within electrodynamics we have some apparent paradoxes. One finds many books which say that a Galilean transformations of the velocity of light is not consistent with the electron-theoretical explanation of refraction and reflection (6). It is widely accepted that if we consider a moving source and a stationary glass, the incident light wave and wave scattered by the dipoles of the glass cannot interfere as required by the electron theory of dispersion since their velocity are different. This is misconception. It is clear that an incident wave with a certain frequency, no matter what its velocity, excites the electrons of a glass into oscillations of the same frequency. They then emit radiation with the same frequency. Thus, the incident and scattered wave at any given point have the same frequency and can interfere. The effect of the different velocities is to produce a relative phase which varies with position in space.
2.6 Convention-Dependent and Convention-Invariant Parts of the Theory
Consider the motion of charged particle in a given magnetic field. The theory of relativity says that the particle trajectory $\vec{x}(t)$ in the lab frame depends on the choice of a convention, namely the synchronization convention of clocks in the lab frame. Whenever we have a theory containing an arbitrary convention, we should examine what parts of the theory depend on the choice of that convention and what parts do not. We may call the former convention-dependent, and the latter convention-invariant parts. Clearly, physically meaningful measurement results must be convention-invariant.
Consider the motion of two charged particles in a given magnetic field, which is used to control the particle trajectories. Suppose there are two apertures at point $A$ and at point $A^{\prime}$. From the solution of the dynamics equation of motion we may conclude that the first particle gets through the aperture at $A$ and the second particle gets through the aperture at $A^{\prime}$ simultaneity. The two events, i.e. the passage of particles at point $A$ and point $A^{\prime}$ have exact objective meaning i.e. convention-invariant. However, the simultaneity of these two events is convention-dependent and has no exact objective meaning. It is important at this point to emphasize that, consistently with the conventionality of simultaneity, also the value of the speed of particle is a matter of convention and has no definite objective meaning.
In order to examine what parts of the dynamics theory depend on the choice of that convention and what parts do not, we want to show the difference between the notions of path and trajectory. Let us consider the motion of a particle in three-dimensional space using the vector-valued function $\vec{x}(t)$. We have a prescribed curve (path) along which the particle moves. The motion along the path is described by $l(t)$, where $l$ is a certain parameter (in our case of interest the length of the arc). The trajectory of a particle conveys more information about its motion because every position is described additionally by the corresponding time instant. The path is rather a purely geometrical notion. If we take the origin of the (Cartesian) coordinate system and we connect the point to the point laying on the path and describing the motion of the particle, then the creating vector will be a position vector $\vec{x}(l)$ (7).
The difference between trajectory $\vec{x}(t)$ and path $\vec{x}(l)$ is very interesting. The path has exact objective meaning i.e. it is convention-invariant. In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle is convention dependent and has no exact objective meaning.
In order to avoid being to abstract for to long we have given some examples: just think of the experiments related with accelerator physics. Suppose we want to perform a particle momentum measurement. A uniform magnetic field can be used in making a ”momentum analyzer” for high-energy charge particles, and it must be recognized that this method for determining the particle’s momentum is convention-independent. In fact, the curvature radius of the path in the magnetic field (and consequently the three-momentum) has obviously an objective meaning, i.e. is convention-invariant. Dynamics theory contains a particle trajectory that we do not need to check directly, but which is used in the analysis of electrodynamics problem.
2.7 Myth about the Reality of Relativistic Time Dilation and Length Contraction
Generally, experts on the theory of relativity erroneously identify the properties of Minkovski space-time with the familiar form that certain convention-dependent quantities assume under the standard Lorentz coordinatization. These quantities usually are called ”relativistic kinematics effects”. There is a widespread belief that the convention-dependent quantities like the time dilation, length contraction, and Einstein’s addition of velocities have direct physical meaning. We found that statement like ” moving clocks run slow” is not true under the adopted absolute time clock synchronization, and, hence, are by no means intrinsic features of Minkowski space-time. Relativistic kinematic effects are coordinate (i.e. convention-dependent) effects and have no exact objective meaning (8). In the case of Lorentz coordinatization, one will experience e.g. the time dilation phenomenon. In contrast to this, in the case of absolute time coordinatization there are no relativistic kinematics effects and no time dilation will be found. However, all coordinate-independent quantities like the particle path $\vec{x}(l)$ and momentum $|\vec{p}|$ remain independent of such a change in clock synchronization.
2.8 Relativistic Particle Dynamics
2.8.1 Covariant Approach
The accelerated motion is described by a covariant equation of motion for a relativistic charged particle under the action of the four-force in the Lorentz lab frame. The trajectory of a particle $\vec{x}_{cov}(t)$ is viewed from the Lorentz lab frame as a result of successive infinitesimal Lorentz transformations. The lab frame time $t$ in the equation of motion cannot be independent from the space variables. This is because Lorentz transformations lead to a mixture of positions and time, and the relativistic kinematics effects are considered to be a manifestation of the relativity of simultaneity.
2.8.2 Noncovariant (”Single Frame”) Approach
Let us consider the conventional particle tracking approach. It is generally accepted that in order to describe dynamics of relativistic particles in the lab reference frame, which we assume inertial, one only needs to take into account the relativistic dependence of the particle momentum on the velocity. The treatment of relativistic particle dynamics involves only corrected Newton’s second law. In a given lab frame, there is an electric field $\vec{E}$ and magnetic field $\vec{B}$. They push on a particle in accordance with
$\frac{d\vec{p}}{dt}=e\left(\vec{E}+\frac{\vec{v}}{c}\times\vec{B}\right)~{},$ (3)
$\vec{p}=m\vec{v}\left(1-\frac{v^{2}}{c^{2}}\right)^{-1/2}~{},$ (4)
where here the particle’s mass, charge, and velocity are denoted by $m$, $e$, and $\vec{v}$ respectively. The Lorentz force law, plus measurements on the components of acceleration of test particles, can be viewed as defining the components of the electric and magnetic fields. Once field components are known from the acceleration of test particles, they can be used to predict the accelerations of other particles.
This solution of the dynamics problem in the lab frame makes no reference to Lorentz transformations. Conventional particle tracking treats the space-time continuum in a non-relativistic format, as a (3+1) manifold. In other words, in this approach, introducing as only modification to the classical case the relativistic mass, time differ from space. In fact, we have no mixture of positions and time (9).
Most of the interesting phenomena in which charges move under the action of electromagnetic fields occur in very complicated situations. But here we just want to discuss the simple problem of the accelerated motion of particles in a constant magnetic field. According to the non-covariant treatment, the magnetic field is only capable of altering the direction of motion, but not the speed (i.e. mass) of a particle. This study of relativistic particle motion in a constant magnetic field, usual for accelerator engineering, looks precisely the same as in nonrelativistic Newtonian dynamics and kinematics. The trajectory of a particle $\vec{x}(t)$, which follows from the solution of the corrected Newton’s second law, does not include relativistic kinematics effects as relativity of simultaneity and the Galilean vectorial law of addition of velocities is actually used.
Let us discuss the important problem of the addition of velocities in relativity. Suppose that in the case of accelerated motion one introduces an infinite sequence of co-moving frames. At each instant, the rest frame is a Lorentz frame centered on the particle and moving with it. Suppose that in inertial frame where particle is at rest at a given time, the traveler was observing light itself. In other words measured speed of light $v=c$, and yet the frame is moving relative the lab frame. How will it look to the observer in the lab frame? According to relativistic law of addition of velocities the answer will be $c$. Maxwell’s equations remain in the same form when Lorentz transformations are applied to them, but Lorentz transformations give rise to non-Galilean transformation rules for velocities, and therefore the theory of relativity shows that, if Maxwell’s equations is to be valid in the lab frame, the trajectories of the particles must include relativistic kinematics effects. In other words, Maxwell,s equations can be applied in the lab frame only in the case when particle trajectories are viewed, from the lab frame as the result of successive infinitesimal Lorentz transformations.
The absence of relativistic kinematics effects is the prediction of conventional non-covariant theory and is obviously absurd from the viewpoint of Maxwell’s electrodynamics. Therefore, something is fundamentally, powerfully, and absolutely wrong in coupling fields and particles within a ”single inertial frame”.
2.9 Mistake in Commonly Used Method of Coupling Fields and Particles
It is generally believed that the electrodynamics problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations. In all standard derivations it is assumed that usual Maxwell’s equations and corrected Newton’s second law can explain all experiments that are performed in a single inertial frame, for instance the lab reference frame.
Going to electrodynamics problem, the differential form of Maxwell’s equations describing electromagnetic phenomena in the same inertial lab frame (in cgs units) is given by the following expressions:
$\vec{\nabla}\cdot\vec{E}=4\pi\rho~{},$ (5)
$\vec{\nabla}\cdot\vec{B}=0~{},$ (6)
×$\vec{\nabla}\times\vec{E}=-\frac{1}{c}\frac{\partial\vec{B}}{\partial t}~{},$ (7)
×$\vec{\nabla}\times\vec{B}=\frac{4\pi}{c}\vec{j}+\frac{1}{c}\frac{\partial\vec{E}}{\partial t}~{}.$ (8)
Here the charge density $\rho$ and current density $\vec{j}$ are written as
$\rho(\vec{x},t)=\sum_{n}e_{n}\delta(\vec{x}-\vec{x}_{n}(t))~{},$ (9)
$\vec{j}(\vec{x},t)=\sum_{n}e_{n}\vec{v}_{n}(t)\delta(\vec{x}-\vec{x}_{n}(t))~{},$ (10)
where $\delta(\vec{x}-\vec{x}_{n}(t))$ is three-dimensional delta function, $m_{n},e_{n},\vec{x}_{n}(t)$, and $\vec{v}_{n}=d\vec{x}_{n}(t)/dt$ denote the mass, charge, position, and the velocity of the $n$th particle, respectively. To evaluate radiation fields arising from an external sources Eq. (8) we need the velocity $\vec{v}_{n}$ and the position $\vec{x}_{n}$ as a function of lab frame time $t$. It is generally accepted by physics community that equation of motion, which describes how coordinates of the particle carrying the charge change with time $t$, is described by corrected Newton’s second law Eq. (4).
This coupling of Maxwell’s equations and corrected Newton’s equation is commonly accepted as useful method in accelerator physics and, in particular, in analytical and numerical calculations of radiation properties. Such approach to relativistic dynamics and electrodynamics usually forces the accelerator physicist to believe that the design of particle accelerators possible without detailed knowledge of the theory of relativity.
However, there is a common mistake made in accelerator physics connected with the difference between $\vec{x}(t)$ and $\vec{x}_{cov}(t)$ trajectories. Let us look at this difference from the point of view of electrodynamics of relativistically moving charges. To evaluate fields arising from external sources we need to know their velocity and positions as a function of the lab frame time $t$. Suppose one wants to calculate properties of radiation. Given our previous discussion the question arises, whether one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particle moving along non-covariant trajectories like $\vec{x}(t)$. We claim that the answer to this question is negative. This algorithm for solving usual Maxwell’s equations in the lab frame, which is considered in all standard treatments as relativistically correct, is at odds with the theory of special relativity. This essential point has never received attention in the physical community. Only the solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle trajectories in the lab frame.
2.10 Clarification of the True Content of the ”Single Frame” Theory
Let us now examine the logical content of the concept of a ”single inertial frame”. If a traveler in a co moving frame, similar to an observer in the lab frame, introduces a definite coordinate-time grid, there is always a definite transformation between these two four-dimensional coordinate systems. Thus, particle trajectories are always viewed from the lab frame as a result of successive transformations, and the form of these transformations depends on the choice of coordinate systems in the comoving and the lab frame.
One might well wonder why it is necessary to discuss how different inertial frames are related to one another. The point is that all natural phenomena follow the principle of relativity, which is a restrictive principle: it says that the laws of nature are the same (or take the same form) in all inertial frames. In agreement with this principle, usual Maxwell’s equations can always be exploited in any inertial frame where electromagnetic sources are at rest using Einstein synchronization procedure in the rest frame of the source. The fact that one can deduce electromagnetic field equations for arbitrary moving sources by studying the form taken by Maxwell’s equations under the transformation between rest frame of the source and the frame where the source is moving is a practical application of the principle of relativity. The question now arises how to assign a time coordinate to the lab frame.
Coordinates serve the purpose of labeling events in an unambiguous way, and this can be done in infinitely many different ways. The principle of relativity dictates that Maxwell’s equations can be applied in the lab frame only in the case when Lorentz coordinates are assigned and particle trajectories are viewing from the lab frame as a result of successive infinitesimal Lorentz transformations between the lab and comoving inertial frames.
A ”single frame” (non-covariant 3+1) approach to relativistic particle dynamics has been used in particle tracking calculations for about seventy years. However, the type of clock synchronization which provides the time coordinate $t$ in the corrected Newton’s equation has never been discussed in literature. It is clear that without an answer to the question about the method of synchronization used, not only the concept of velocity, but also the dynamics law has no physical meaning. A ”single frame” approach to relativistic particle dynamics is forcefully based on a definite synchronization assumption but this is actually a hidden assumption. According to conventional particle tracking, the dynamical evolution in the lab frame is based on the use of the lab frame time $t$ as an independent variable. Such approach to relativistic particle dynamics is actually based on the use of a not standard (not Einstein) clock synchronization assumption in the lab frame.
In fact, the usual for accelerator engineering study of relativistic particle motion in a constant magnetic field looks precisely the same as in nonrelativistic Newtonian mechanics and the trajectories of the electrons does not include relativistic kinematics effects. According to textbooks, this is no problem. If no more than one frame is involved, one does not need to use (and does not need to know) the theory of relativity. Only when one passes from one reference frame to another the relativistic context is important. Conventional particle tracking in a constant magnetic field is actually based on classical Newton mechanics. It is generally believed that the electrodynamics problem, similar to conventional particle tracking, can be treated within a description involving a single inertial frame and one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particles moving along the non-covariant (single frame) trajectories.
This is misconception. The situation when only one frame is involved and the relativistic context is unimportant cannot be realized. The lab observer may argue, ”I don’t care about other frames.” Perhaps the lab observer doesn’t, but nature knows that, according to the principle of relativity, Maxwell’s equations are always valid in the Lorentz comoving frame. Electrodynamics equations can be written down in the lab frame only when a space-time coordinate system has been specified. An observer in the lab frame has only one freedom. This is the choice of a coordinate system (i.e. the choice of clock synchronization convention) in the lab frame. After this, the theory of relativity states that the electrodynamics equations in the lab frame are the result of transformation of Maxwell’s equations from the Lorentz comoving frame to the lab frame.
2.11 Bibliography and Notes
1. Many physicists tend to think of Galilean transformations as pre-relativistic transformations between spatial coordinates and time that are not compatible with the special theory of relativity. To quote e.g. Bohm [1] ”… the Galilean law of addition of velocities implies that the speed of light should vary with the speed of the observing equipment. Since this predicted variation is contrary to the fact, the Galilean transformations evidently cannot be the correct one.”. Similar statements can also be found in recently published pedagogical papers. To quote e.g. Drake and Purvis [2] ”One of the great insights to come relativity theory was the realization that Galilean transformations are wrong. The correct way to translate the space-time measure of events between inertial frames is with the Lorentz transformations” However, this is not true. Galilean transformations are simply transformations relating a given coordinate set to another coordinate set. The space-time continuum can be described in arbitrary coordinates, and choice of this set of coordinates cannot change the geometry of space-time.
2. The mathematical argument that in the process of transition to arbitrary coordinates the geometry of the space-time does not change, is considered in textbooks as erroneous. To quote L. Landau and E. Lifshitz [3]: ”This formula is called the Galileo transformation. It is easily to verify that this transformation, as was to be expected, does not satisfy the requirements of the theory of relativity; it does not leave the interval between events invariant.”. This fact is ascribed to a lack of understanding of the difference between convention-dependent and convention-invariant parts of the theory. In pseudo-Euclidean geometry the interval between events is an invariant in arbitrary coordinates. A comparison with three-dimensional Euclidean space might help here. In the usual 3D Euclidean space, one can consider a Cartesian coordinate system ($x,y,z)$, a cylindrical coordinate system ($r,\phi,z$), a spherical coordinate system ($\rho,\theta,\phi$), or any other. Depending on the choice of the coordinate system one respectively has $ds^{2}=dx^{2}+dy^{2}+dz^{2}$, $ds^{2}=dr^{2}+r^{2}d\phi^{2}+dz^{2}$, $ds^{2}=d\rho^{2}+\rho^{2}d\theta^{2}+\rho^{2}\sin\theta^{2}d\phi^{2}$. The metric actually does not change, but the components of the metric do, depending only on the choice of coordinates. In general, in fact, we write $ds^{2}=g_{ik}dx^{i}dx^{k}$. Considering Cartesian coordinates we will always have $g_{ij}=\mathrm{diag}(1,1,1)$. Similarly, Lorentz transformations between inertial frames with Einstein coordinates leave the components of the metric tensor unvaried.
3. There is a widespread belief that a Lorentz transformation reduces to a Galilean transformation in the non-relativistic limit. To quote e.g. French [4]: ”The reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to Galilean relation $t^{\prime}=t$ requires $x\ll ct$ as well as $v/c\ll 1$”. Similar statements can also be found in recently published textbooks. To quote Rafelski [5]: ”A wealth of daily experience shows that the Galilean coordinate transformation (GT) is correct in the nonrelativistic limit in which the speed of light is so large that it plays no physical role. Any coordinate transformation replacing the GT must also agree with this experience, and thus must contain the GT in the nonrelativistic limit.” This statement is absurd conclusion even from a mathematical standpoint. This has been recognized by some expects, perhaps most explicitly by Baierlein who states [6] that ”If the Lorentz transformation for infinitesimal $v/c$ were to reduce to the Galilean transformation, then iterative process could never generate a finite Lorentz transformation that is radically different from the Galilean transformation. But the finite transformations are radically different, and so -however subtly-the infinitesimal Lorentz transformation must differ significantly from the Galilean transformation.”
4. Many authors of textbooks still attribute a measurable status to the conventional quantities. To quote e.g. Cristodoulides [7] ”The fact that Galilean transformation does not leave Maxwell’s equations has already been mentioned […] On the other hand, experiments show that the speed of light in vacuum is independent of the source or observer.”.
5. Because we have empirical access only to the round-trip average speed of light, statements about the magnitude and isotrophy of the one-way speed of light must reflect the assumptions made in the choice of time coordinatization, and such entities change as the theory is re-synchronized. To quote C. Anderson, I. Vetharaniam, and G. Stedman [8]: ”No experiment, then, is a ”one-way” experiment. An empirical test of any property of the one-way speed of light is not possible. Such quantities as the one-way speed of light are irreducibly conventional in nature, and recognizing this aspect is to recognize a profound feature of nature”.
6. The peculiarity of the kinematic consequence of using Galilean transformations is that the speed of light emitted by a moving source depends on the relative velocity between source and observer. A widespread theoretical argument used to support the incorrectness of Galilean transformations is the conclusion that a Galilean transformation of the velocity of light is not consistent with the explanation of reflection and refraction. This idea is a part of the material in well-known books. To quote Pauli [9] ”[…] it is essential that the spherical waves emitted by the dipoles in the body should interfere with the incident wave. If we now think of the body as at rest, and the light source moving relative to it, then […] the wave emitted by the dipoles will have velocity different from that of the incident wave. Interference is therefore not possible.” This conclusion is incorrect. It is clear that the incident and scattered wave at any given point have the same frequency and can interfere.
7. For a general discussion of the difference between path and trajectory we suggest reading the book [10].
8. The standard textbooks erroneously identify the properties of space-time with the familiar form that certain synchronization-dependent quantities assume under the Einstein’s clock synchronization. These quantities are, for example, the formulae for ”time dilation” and ”length contraction”. Fortunately, it has also been occasionally stressed in the literature that the forms of the above ”relativistic effects” are coordinate-dependent, their forms depend in turn on the kind of synchronization procedure adopted. To quote Leubner, K. Auflinger, and P. Krumm [11]: ”… there is a widespread belief among students that the familiar form of coordinate-dependent quantities like the measured velocity of light, the Lorentz transformation between two observers, ”addition of velocities”, ”time dilation”, ”length contraction” which they assume under the standard clock synchronization, is relativity proper. In order to demonstrate that this is by no means so, this paper studies the consequences of a non-standard synchronization, and it is shown that drastic changes in the appearance of all these quantities are thus induced.”
9. Clarification of the true content of the non-covariant theory can be found in various advanced textbooks. To quote e.g. Ferrarese and Bini [12]: ”… within a single inertial frame, the time is an absolute quantity in special relativity also. As a consequence, if no more than one frame is involved, one would not expect differences between classical and relativistic kinematics. But in the relativistic context there are differences in the transformation laws of the various relative quantities (of kinematics or dynamics), when passing from one reference frame to another.” We see that authors give a special role to concept of a ”single inertial frame”. The name ”single inertial frame” tends to suggest that a distinctive trait of non-covariant theory is the absence of relativistic kinematics in the description of particle motion. This point was expressed by Friedman [15]: ”Within any single inertial frame, things looks precisely the same as in Newtonian kinematics: there is an enduring Euclidean three-space, a global (i.e. absolute) time $t$, and law of motion. But different inertial frames are related to one another in a non-Newtonian fashion.” According to conventional particle tracking, within the ”single” frame there is no relativistic kinematics effects. This, as we already mentioned, contradicts the Maxwell’s electrodynamics.
3 Space-Time and Its Coordinatization
3.1 Introductory Remarks
Let us discuss an ”operational interpretation” of the Lorentz and absolute time coordinatizations. We should underline that we claim the non covariant approach to relativistic particle dynamics is actually based on the use of a not standard and unusual clock synchronization assumption within the theory of relativity. It is important to know how to operationally interpret the absolute time convention i.e. how one should perform the clock synchronization in the lab frame. The result is very interesting, since it tell us about difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view.
3.2 Choice of Coordinates System in an Inertial Frame
Each physical phenomenon occurs in space and time. A concrete method for representing space and time is a frame of reference. One-and-the same space and time can be represented by various coordinate-time grids, i.e., by various frames of references. Even the simplest space-time coordinate systems require carefully description.
Clocks reveal the motion of a particle through the coordinate-time grid. The general approach to the determination of the motion of a particle is the following: at any instant a particle has a well-defined velocity $\vec{v}$ as measured in a laboratory frame of reference. How is a velocity of a particle found? The velocity is determined once the coordinates in the lab frame are chosen, and is then measured at appropriate time intervals along the particle’s trajectory. But how to measure a time interval between events occurring at different points in space? In order to do so, and hence measure the velocity of a particle within a single inertial lab frame, one first has to synchronize distant clocks. The concept of synchronization is a key concept in the understanding of special relativity. It is possible to think of various methods to synchronize distant clocks (1). The choice of a convention on clock synchronization is nothing more than a definite choice of coordinates system in an inertial frame of reference of the Minkowski space-time.
The space-time continuum can be described in arbitrary coordinates. By changing these arbitrary coordinates, the geometry of the four-dimensional space-time obviously does not change, and in the special theory of relativity we are not limited in any way in the choice of a coordinates system. Relying on the geometric structure of Minkowski space-time, one defines the class of inertial frames and adopts a Lorentz frame with orthonormal basis vectors. Within the chosen Lorentz frame, Einstein’s synchronization procedure of distant clocks (which based on the constancy of the speed of light in all inertial framers) and Cartesian space coordinates are enforced.
3.3 Inertial Frame where a Source of Light is at Rest
Let us give an ”operational interpretation” of the Lorentz coordinatizations. The fundamental laws of electrodynamics are expressed by Maxwell’s equations, according to which, as well-known, light propagates with the same velocity $c$ in all directions. This is because Maxwell’s theory has no intrinsic anisotropy. It has been stated that in their original form Maxwell’s equations are only valid in inertial frames. However, Maxwell’s equations can be written down in coordinate representation only if the space-time coordinate system has already been specified.
The problem of assigning Lorentz coordinates to the lab frame in the case of accelerated motion is complicated. We would like to start with the simpler question of how to assign space-time coordinates to an inertial frame, where a source of light is at rest. We need to give a ”practical”, ”operational” answer to this question. The most natural method of synchronization consists in putting all the ideal clocks together at the same point in space, where they can be synchronized. Then, they can be transported slowly to their original places (slow clock transport) (2).
The usual Maxwell’s equations are valid in any inertial frame where sources are at rest and the procedure of slow clock transport is used to assign values to the time coordinate. The same considerations apply when charged particles are moving in non-relativistic manner. In particular, when oscillating, charged particles emit radiation, and in the non-relativistic case, when charges oscillate with velocities much smaller than $c$, dipole radiation is generated and described with the help of the Maxwell’s equations in their usual form.
Let’s examine in a more detail how the dipole radiation term comes about. The retardation time in the integrands of the expression for the radiation field amplitude, can be neglected in the cases where the trajectory of the charge changes little during this time. It is easy to find the conditions for satisfying this requirement. Let us denote by $a$ the order of magnitude of the dimensions of the system. Then the retardation time $\sim a/c$. In order to ensure that the distribution of the charges in the system does not undergo a significant change during this time, it is necessary that $a\ll\lambda$, where $\lambda$ is the radiation wavelength. Thus, the dimensions of the system must be small compared to radiation wavelength. This condition can be written in still another form $v\ll c$, where $v$ is of the order of magnitude of the velocities of the charges. In accounting only for the dipole part of the radiation we neglect all information about the electron trajectory. Therefore, one should not be surprised to find that dipole radiation theory gives fields very much like the instantaneous theory.
The theory of relativity offers an alternative procedure of clocks synchronization based on the constancy of the speed of light in all inertial frames. This is usually considered a postulate but, as we have seen, it is just a convention. The synchronization procedure that follows is the usual Einstein synchronization procedure. Suppose we have a dipole radiation source. When the dipole light source is at rest, the field equations are constituted by the usual Maxwell’s equations. Indeed, in dipole radiation theory we consider the small expansion parameter $v/c\ll 1$ neglecting terms of order $v/c$. In other words, in dipole radiation theory we use zero order non relativistic approximation. Einstein synchronization is defined in terms of light signals emitted by the dipole source at rest, assuming that light propagate with the same velocity $c$ in all directions. Using Einstein synchronization procedure in the rest frame of the dipole source, we actually select the Lorentz coordinate system.
Slow transport synchronization is equivalent to Einstein synchronization in inertial system where the dipole light source is at rest (3). In other words, suppose we have two sets of synchronized clocks spaced along the $x$ axis. Suppose that one set of clocks is synchronized by using the slow clock transport procedure and the other by light signals. If we would ride together with any clock in either set, we could see that it has the same time as the adjacent clocks, with which its reading is compared. This is because in our case of interest, when light source is at rest, field equations are the usual Maxwell’s equations and Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions. Using any of these synchronization procedures in the rest frame we actually select a Lorentz coordinate system. In this coordinate system the metric of the light source has Minkowski form $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$. We want now to turn to the subject of electromagnetic waves. In free space, the electric field $\vec{E}$ of an electromagnetic wave emitted by the dipole source at rest satisfies the equation $\Box^{2}\vec{E}=\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$.
3.4 Motion of a Light Source With Respect to the Inertial Frame
We now consider the case when the light source in the lab frame is accelerated from rest up to velocity $v$ along the $x$-axis. In other words, we consider so-called active (physical) boost. When we state that the source undergoes an acceleration, and the inertial observer (with measuring devices) does not, the acceleration means acceleration relative to the fixed stars. Any acceleration relative to the fixed stars (i.e. any active boost of velocity) has an absolute meaning.
In the case of an active boost we are dealing with motion of the same physical system, evolving in time and treated from the point of view of the same reference system. A fundamental question to ask is whether our lab clock synchronization method depends on the state of motion of the light source or not. The answer simply fixes a convention. The simplest method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the light source was at rest, i.e. we still enforce the clock transport synchronization ( or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest). This choice is usually the most convenient one from the viewpoint of connection to laboratory reality. This synchronization convention preserves simultaneity and is actually based on the absolute time (or absolute simultaneity) convention.
Absolute simultaneity can be introduced in special relativity without affecting neither the logical structure, no the (convention-independent) predictions of the theory. We begin with the metric as true measure of space-time intervals for an inertial observer with coordinates $(t,x,y,z)$. Using the Galilean boost: we substitute $x\to x-vt$, while leaving time unchanged into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain
$ds^{2}=c^{2}(1-v^{2}/c^{2})dt^{2}+2vdxdt-dx^{2}-dy^{2}-dz^{2}~{}.$ (11)
This metric describes the electrodynamics of the moving light source with viewpoint of the inertial observer measurements. From the formula itself it is easy to see that the result of measurement of the electromagnetic field configuration in the moving system in terms of coordinate $x$ and time $t$ of the measured device at rest in the inertial frame is the same as the measured result of this configuration at point $x-vt$ and time $t=0$.
To understand the physical interpretation of the Eq. 11, it is important to note that there are infinitely many ways to label points in space-time by coordinates. The different coordinate systems are connected by coordinate transformations. It is always possible to create a new frame of reference by relabeling coordinates, and then discussing physical phenomena in terms of the new coordinate labels - a passive (coordinate) transformation. For example, it is always possible to create a so-called comoving coordinate system in the lab frame, and then discussing radiation from the moving source in therm of the new (comoving) coordinate labels. In the comoving coordinate system, fields are expressed as a function of the independent variables $x^{\prime},y^{\prime},z^{\prime}$, and $t^{\prime}$. The variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$ can be expressed in terms of the independent variables $x,y,z,t$ by means of a passive Galilean transformation, so that fields can be written in terms of $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$. After the passive transformation, the Cartesian coordinates of the source transform as $x^{\prime}=x-vt,~{}y^{\prime}=y,~{}z^{\prime}=z$. This transformation completes with the invariance of simultaneity, $t^{\prime}=t$. The transformation of time and spatial coordinates of any event has the form of a Galilean transformation. One must take into account that $(t,x,y,z)$ are coordinates of the rule clock structure at rest in the inertial frame.
The equivalence of the active and passive pictures is due to the fact that moving system one way is equivalent to moving the coordinate system the other way by an equal amount. Obviously, there is no a priory guarantee that the world behave like this. For instance, Newton mechanics is not invariant under passive rotation as demonstrated Newton’s bucket experiment. The empirical principle implies that within a single inertial frame every passive Galilean boost is postulated to corresponds to an active Galilean boost. According to this equivalence of passive and active boosts, in our case of interest the Maxwell’s equations always valid in the comoving frame. In the comoving frame, fields are expressed as a function of the independent variables $x^{\prime},y^{\prime},z^{\prime}$, and $t^{\prime}$. The electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$. However, the variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$ can be expressed in terms of the independent variables $x,y,z,t$ by means of a Galilean transformation, so that fields can be written in terms of $x,y,z,t$. From the Galilean transformation $x^{\prime}=x-vt,~{}y^{\prime}=y,~{}z^{\prime}=z,~{}t^{\prime}=t$, after partial differentiation, one obtains $\partial/{\partial t}=\partial/{\partial t^{\prime}}-v\partial/{\partial x^{\prime}}$, $\partial/{\partial x}=\partial/{\partial x^{\prime}}$. Hence the wave equation transforms into
$\Box^{2}\vec{E}=\left(1-\frac{v^{2}}{c^{2}}\right)\frac{\partial^{2}\vec{E}}{\partial x^{2}}-2\left(\frac{v}{c}\right)\frac{\partial^{2}\vec{E}}{\partial ct\partial x}+\frac{\partial^{2}\vec{E}}{\partial y^{2}}+\frac{\partial^{2}\vec{E}}{\partial z^{2}}-\frac{1}{c^{2}}\frac{\partial^{2}\vec{E}}{\partial t^{2}}=0~{},$ (12)
This last result is closely related to the metric Eq. 11.
The solution of this equation $F[x-(c+v)t]+G[x+(-c+v)t]$ is the sum of two arbitrary functions, one of argument $x-(c+v)t$ and the other of argument $x+(-c+v)t$. Here we obtained the solution for waves which move in the $x$ direction by supposing that the field does not depend on $y$ and $z$. The first term represents a wave traveling forward in the positive $x$ direction, and the second term a wave traveling backwards in the negative direction.
We conclude that the speed of light emitted by a moving source measured in the lab frame $(t,x)$ depends on the relative velocity of source and observer, in our example $v$. In other words, the speed of light is compatible with the Galilean law of addition of velocities. In fact, the coordinate velocity of light parallel to the $x$-axis is given by $dx/dt=c+v$ in the positive direction, and $dx/dt=-c+v$ in the negative direction. The reason why it is different from the electrodynamics constant $c$ is due to the fact that the clocks are synchronized following the absolute time convention. One fixed, such convention should be applied and kept in a consistent way for both dynamics and electrodynamics. In fact, the light velocity transforms as a particle velocity in the absolute time coordinatization as it mast be.
3.5 Discussion
So far we have considered the Galilean transformation of the electrodynamics equations. We would like to emphasize a very important difference between fundamental and phenomenological theory. For example, we can characterize Newton’s equation in the co-moving frame as a phenomenological law. The microscopic interpretation of the inertial mass of a particle is not given and the rest mass is introduced in an ad hoc manner. It can be said that the particle has some inner variables - that we do not yet know about. Any phenomenological law, which is valid in the Lorentz rest frame, can be embedded in the four dimensional space-time only by using Lorentz coordinatization. Now we wish to consider the electrodynamics. It is important to stress at this point that there is no machinery (i.e. inner variables) behind the laws of electrodynamics. It is obvious that electromagnetic fields are fundamental and the electrodynamics theory meets all requirements of the theory of relativity. Consequently, one can use any synchronization to describe the electromagnetic field dynamics. The variety of methods of description physical systems makes it possible to select for each problem the representation most suitable for its solution. It is always possible to create a new frame of reference by relabeling coordinates, and then discussing physical phenomena in terms of the new coordinate labels - a passive transformation. An active transformation of a physical system is its motion, i.e., a variation in its characteristics under the effect of some internal or external interactions. We are dealing with motion of the same physical system, evolving in time and treated from the point of view of the same reference system.
Now let us discuss correspondence between passive and active transformations. A comparison between the special relativity and the Newton theory might help here. The formalism of non-relativistic Newton’s theory is associated with absolute space structure. However, there is relativity of the uniform motions and we cannot determine the particle absolute velocity. Passive Galilean boosts are dynamical symmetries of Newtonian mechanics. The equivalence of the active and passive Galilean boost is due to the fact that moving system one way is equivalent to moving the coordinate system the other way by an equal amount as described by the famous (Galileo) ship experiment . The reason of this (Galilean group) symmetry is simply that equations of motion do not explicitly depend on velocity. The internal dynamics of the system does not change under the Galilean boost. In the same time, Newtonian mechanics is not invariant under passive rotation. In fact, the form of equations of motion change due to pseudo-force terms (4). The Newton bucket argument by Newton was based on comparison of two different situations: a vessel of water at rest, vs the vessel after it has gradually been rotated from rest to a state of uniform angular velocity. The argument against the equivalence of the active and passive rotation is based on notion that in both cases there is no relative motion between the water and the bucket, while the shape of the water turns out to be different. The passive rotation transformation is no more than a change of representation. There is no relativity of accelerations and we can determine the particle absolute acceleration in the inertial frame. The absolute acceleration means acceleration relative to the fixed stars.
Let us now move to consider correspondence between active and passive transformations in special relativity. Within in single inertial frame the kinematical symmetry is corresponds to dynamical symmetry and the symmetry principle (i.e. equivalence of the active and passive Galilean boost) is satisfied. It should be note that passive Galilean boost within a single inertial frame is simply another parametrization of the observations of the inertial observer. According to passive transformation ($x\to x-vt$, $t\to t$), the motion of fixed stars with respect to the observer and his devices not changes. It is important to note that $(t,x,y,z)$ are coordinates of the rule clock structure at rest in the inertial frame.
In the Chapter 5 we will consider the problem of aberration of light in accelerated systems on the basis of the theory of special relativity. This problem can be adequately treated only by an approach which uses the absolute time coordinatization. In order to write electrodynamics laws in an accelerated frame, we define the metric of the accelerated frame using the inverse Galilean boost. We want to emphasize that in the case of an inverse Galilean boost of velocity we consider the effect of interaction which defined in terms of acceleration motion relative to the fixed stars. Any acceleration relative to the fixed stars has an absolute meaning.
This means that the passive Galilean transformation within a single inertial frame is quite distinct from the (passive) inverse Galilean transformation, $x\to x_{n}+vt$, $t\to t_{n}$, i.e acceleration with respect to fixed stars of an observer (with his measuring instruments) from inertial frame to the accelerate frame. One must take into account that $(t_{n},x_{n},y_{n},z_{n})$ are coordinates of the rule clock structure at rest in the accelerated frame. Such transition of the coordinates of the light source does not stand for a change of representation, but for a change in the absolute motion of the light source with respect to the fixed stars. The fact that in the real process of transmission to an accelerated (with respect to the fixed stars) frame the observer will experience the inertial force, is not accounted in the standard treatments. We argue that reciprocity of the inertial frames, which is considered in textbooks as relativistically correct, is at odds with special relativity. The arguments concerning the relativity of motion in our case of interest cannot be applied, since the inertial and accelerated reference systems are not equitable.
The formalism of relativistic physics is associated with absolute space-time structure. There is no relativity of the uniform motions and we can determine the system absolute velocity in the initial inertial frame. The reason of this asymmetry is simply that electrodynamics equations are fundamental and explicitly depend on velocity. According to the special relativity, there is a real difference between accelerated (with respect to the fixed stars) inertial frame and the inertial frame without accelerational history and this difference is closely associated with the inertial forces. It should be note that the duration of the acceleration period has a negligible effect on the anisotropy in the accelerated frame. Nevertheless, the integral effect of inertial forces completely determined the problem. A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer. We can treat a uniformly accelerating frame as if it were an in inertial frame with the addition of a uniform pseudo-gravitational field. This is basic content of equivalence principle. The principle of equivalence can be applied to solve non-inertial kinematics problems with dynamics methods. For instance, when the accelerated system starts moving with constance velocity the pseudo-gravitational field is zero. For the gravitational potential difference between the two frames we get (in the simplified second order approximation) $\phi=-v^{2}/2$. The fact that the inertial observer always turns out to have aged more that accelerated one, in the equivalence principle analysis is elementary consequence of the slowly down of the rate of the processes in the pseudo-gravitational potential. In fact, for clock at rest in the accelerated frame we have $dt_{prop}=dt_{n}(1+\phi/c^{2})=dt_{n}[1-v^{2}/(2c^{2})]$. In other words, the time dilation effect is absolute, but not relative.
There is another way to solve the problem. From a space-time geometrical viewpoint, the explanation consists in using the fact that metric tensor must be continuous quantity. When the accelerated system starts moving with constant velocity the procedure of smooth tailoring of metric tensor can be performed. The (non-diagonal) Langevin metric is found by matching the accelerated frame and initial inertial frame metric tensors. Many experts who learned the theory of relativity using textbooks will find this statement disturbing. Indeed, according to textbooks, the problem is complete symmetric and both frames are equivalent. The popular belief remains that coordinates of these frames should be related by the Lorentz transformations and the metric in both frames will be diagonal, although it is an exercise in elementary algebra to prove discontinuity of metric tensor. With mathematical viewpoint, we must conclude that when authors of textbooks are dealing with transformation and this transformation does not preserve continuity of the metric tensor there are no symmetry between inertial frames.
The initial inertial frame is not physically identical to the Newton’s absolute space. The special relativity taught us to think in terms of pseudo-Euclidean space-time geometry. In special relativity the absolute space and the absolute time of Newton are fused into absolute space-time. We will discuss this subject further in the Chapter 5, Chapter 9.
3.6 A Clock Resynchronization Procedure
After the Galilean boost in the case of the absolute time coordinatization we can see that the homogeneous wave equation for the field in the lab frame has nearly but not quite the usual, standard form that takes when there is no acceleration up to velocity $v$. The main difference consists in the crossed term $\partial^{2}/\partial t\partial x$, which complicates the solution of the equation. To get around this difficulty, we observe that simplification is always possible. The trick needed here is to further make a change of the time variable. Suppose that we describe the effect of radiation by working only up to the first order in $v/c$. In this approximation, it is always possible to introduce a time transformation $t\to t+xv/c^{2}$. After this time shift we obtain the usual wave equation.
Let us now consider the case of an arbitrary velocity along the $x$-axis. We have, then, a general method for finding solution of electrodynamics problem in the case of the absolute time coordinatization. The new independent variables $(x_{L},t_{L})$ can be expressed in terms of the old independent variables $(x,t)$:
$ct_{L}=t\sqrt{1-v^{2}/c^{2}}+vx/[c^{2}\sqrt{1-v^{2}/c^{2}}]~{},~{}x_{L}=x/\sqrt{1-v^{2}/c^{2}}~{},$ (13)
Since the change variables completed by the Galilean transformation is mathematically equivalent to the Lorentz transformation, it obviously follows that transforming to new variables $x_{L},t_{L}$ leads to the usual Maxwell’s equations. In particular, the wave equation Eq. (12) transforms into
$\Box_{L}^{2}E=\nabla_{L}^{2}E-\partial^{2}E/\partial(ct_{L})^{2}=0~{}.$ (14)
In the new variables the velocity of light emitted by a moving source is constant in all directions, and equal to the electrodynamics constant $c$.
The overall combination of the Galilean boost and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the lab frame, but in this context this transformation are only to be understood as useful mathematical device, which allow one to solve the electrodynamics problem in the choice of absolute time synchronization with minimal effort.
We can now rise an interesting question: do we need to transform the results of the electrodynamics problem solution into the original variables? We state that the variable changes performed above have no intrinsic meaning - their meaning only being assigned by a convention. In particular, one can see the connection between the time shift $t\to t+xv/c^{2}$ and the issue of clock synchrony. Note that the final change in the scale of time and spatial coordinates is unrecognizable also from a physical viewpoint. It is clear that the convention-independent results of calculations are precisely the same in the new variables. As a consequence, we should not care to transform the results of the electrodynamics problem solution into the original variables.
Consider now two light sources say ”1” and ”2”. Suppose that in the lab frame the velocities of ”1” and ”2” are $\vec{v}_{1}$, $\vec{v}_{2}$ and $\vec{v}_{1}\neq\vec{v}_{2}$. The question now arises how to assign a time coordinate to the lab reference frame. We have a choice between an absolute time coordinate and a Lorentz time coordinate. The most natural choice, from the point of view of connecting to the laboratory reality, is the absolute time synchronization. In this case simultaneity is absolute, and for this we should prepare, for two sources, only one set of synchronized clocks in the lab frame. On the other hand, Maxwell’s equations are not form-invariant under Galilean transformations, that is, their form is different on the lab frame. In fact, the use of the absolute time convention, implies the use of much more complicated field equations, and these equations are different for each source. Now we are in the position to assign Lorentz coordinates. The only possibility to introduce Lorentz coordinates in this situation consists in introducing individual coordinate systems (i.e. individual set of clocks) for each source. It is clear that if operational methods are at hand to fix the coordinates for the first source, the same methods can be used to assign values to the coordinates for the second source and these will be two different Lorentz coordinate systems.
Suppose now that the second secondary source (e.g. mirror or pupil detection) interacts with the light beam radiated from the first source. The peculiarity of this problem with the viewpoint of the special relativity is that here we cannot prepare for sources a common Lorentz coordinatization in the single inertial frame.
3.7 Radiation by a Moving Source. Peculiarity of the Collinear Geometry
The utilization of the theory of relativity becomes indispensable when we consider optical phenomena associated with a motion of light sources. So far we have considered the covariant and non covariant ways to solve the problem of radiation by a moving source in an inertial frame. Let us consider the acceleration of a dipole source in the lab inertial frame up to velocity $v$ along the $x$-axis. The question arises how to assign synchronization in the lab frame after the source acceleration. Before acceleration we picked a Lorentz coordinate system. Without changing synchronization in the lab frame after the source acceleration we have a complicated situation as concerns electrodynamics of moving charges. As a result of such boost, the transformation of time and spatial coordinates has no form of a Lorentz transformation. The Maxwell’s equations can be applied in the lab inertial frame only in the case when Lorentz coordinates are assigned. In order to keep the Lorentz coordinatiyation after acceleration, one needs to change the space-time coordinate system by introducing new variables $(x_{L},y_{L},z_{L},t_{L})$, Eq. (13). In the new variables the velocity of light emitted by a moving source is equal to the electrodynamics constant $c$ in all directions.
The aberration of light and Doppler effect are practical cases of study for illustration the difference between covariant and non-covariant approaches. The present treatment is designed to try and bring out the role of the Lorentz time transformation. The most convenient way for calculating radiation is to make use of the formulae for Lorentz transition from one inertial frame of reference to another. It is convenient to introduce the reference frame $S^{\prime}$ fixed to the source. Then the problem solved as follows. In the frame $S^{\prime}$ the 4-vector of the light beam is specified, i.e. the frequency and the propagation direction of light are known. The frequency of light in the frame $S$ is easily to find using the Lorentz transformation formulae. Specifically, let the radiation have a dipole character and frequency $\omega_{0}$ in the frame of reference in which the source is at rest. Then in the laboratory frame of reference $S$ in which the source as a whole moves with velocity $v$ and the radiation comes along the velocity direction we observe the so-called radial Doppler effect. It is described by the well-known formula $\omega=\omega_{0}\sqrt{1-v^{2}/c^{2}}/(1-v/c)$. The effect of the factor $\sqrt{1-v^{2}/c^{2}}$ can be summarized in the following statement: on the moving object time is flowing slower than expected (time dilation).
It is interesting to note that, the use of the conventional coupling of Maxwell’s equations and corrected Newton’s equation for the calculation of the radiation from a moving source does not necessarily leads to mistake. For rectilinear motion of the source and the emitted light beam, the non-covariant and covariant approaches produce the same trajectories, and Maxwell’ equations are compatible with the result of conventional particle tracking. This method was incorrectly extended to the case of non-collinear geometry.
At this point a reasonable question arises: why the same method gives incorrect result in the case of radiation of light from a source moving perpendicular to its radiated light beam propagation. It is not difficult to see that the peculiarity of the collinear geometry where there is a source moving along the same line as the radiated beam is that here the velocity is perpendicular to the plane of radiation wavefront (i.e plane of simultaneity). Thus, for collinear motion, the plane of simultaneity in the absolute time coordinatization will have the same orientation for the Lorentz coordinatization.
So far we have dealt with the longitudinal effect when the motion takes place along the straight line through the source and the receiver. In the transverse case the motion is perpendicular to this direction and the source is accelerated along the radiation wavefront (i.e. initial plane of phase simultaneity). The orientation of the radiation wavefront is therefore no longer something absolute and depends on the clock synchronization convention. Let us first consider a Lorentz transformation. The explanation consists in using a Lorentz boost to describe the uniform translation motion of the light source in the lab frame. It is clear, that wavefront phases which are simultaneous in $S^{\prime}$, but occur at different $x^{\prime}$ locations, are not simultaneous in $S$. If make a Lorentz boost, we automatically introduce a time transformation $t^{\prime}=t-xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation wavefront in the lab frame on the angle $v/c$ in the first order approximation. According to Maxwell’s electrodynamics, radiation is always emitted in the direction normal to the radiation wavefront. Then, the radiated light beam is propagated at the angle $v/c$, yielding the phenomenon of the aberration of light. This remarkable phenomenon is what we shall discuss in the next chapter. According to the conventional (3+1) approach, the simultaneity is absolute and there is no mixture of positions and time when sources change their velocities in the inertial frame. Thus it seems as if the conventional (3+1) approach is unable to account for the geometric phenomenon of aberration. In making use of the Maxwell’s equations and the absolute time transformation we obtain no deflection at all of the wavefront plane and the energy transport direction.
Let us now return to our consideration of the conventional coupling between Maxwell’s equations and particle trajectories in the ”single lab frame”. According to the theory of special relativity, there is no objection to the standard description of the optical phenomena associated with a motion of light source in the case when the radiation comes along the velocity direction. For example, for collinear geometry, the conventional coupling fields and particles may be a useful approach for analysis of the Doppler effect.
In order to understand the relativistic red shifting of the light source in a moving system, we have to watch the machinery of the source and see what happens when it is moving. Since that is rather difficult, we shall take a very simple a kind of source, but it will work in principle.
[…]
According to covariant approach, the various relativistic kinematics effects concerning to the dipole radiation setup, turn up in successive orders of approximation. In the first order $(v/c)$. - relativity of simultaneity. In the second order $(v/c)^{2}$. - time dilation. The first order kinematics term $(v/c)$ plays an essential role only in the description of the dipole radiation in the perpendicular geometry. In the case of collinear geometry, a motion of the dipole source, according to the theory of relativity, influences the kinematics terms of the second order $(v/c)^{2}$ only. The ignorance of this distinction between collinear and non collinear boosts is the source of much confusion in the literature.
In the non-covariant approach, a solution of the dynamics problem in the lab frame makes no reference to Lorentz transformations. This means that, for instance, within the lab frame the motion of particles looks precisely the same as predicted by classical mechanics, with its absolute time. The relativity of simultaneously (i.e. mixture of positions and time) do not have a place in this description. Such (the absolute time) method is suitable to account for the outcome of the experiments in collinear geometry. We argue that this algorithm for solving usual Maxwell’s equations in the lab frame is not applicable in the transverse case. However, the Maxwell’s equations in the lab frame are compatible only with covariant trajectories calculated by using Lorentz coordinates, therefore including such relativistic features as relativity of simultaneity.
3.8 Absolute Time Synchronization in a Laboratory Frame
The utilization of the electrodynamics in the absolute time coordinatization becomes indispensable when we use non-covariant (3+1) dimensional approach (i.e. old kinematics) to relativistic particle dynamics. According to conventional particle tracking, the dynamical evolution in the lab frame is based on the use of the lab time $t$ as independent variable. The relativistic kinematics effects do not have a place in this description. Such approach to relativistic particle dynamics is actually based on the use of the absolute time synchronization assumption in the lab frame. If one wants to use the corrected Newton equations (in other words, using time as independent variable), only solution of the electrodynamics equations in non-covariant form (i.e. in the absolute time coordinatization) gives the correct coupling between electrodynamics equations and particle trajectories in the ”single frame”. That is the reason why the absolute time synchronization convention is extremely important from the applied point of view.
3.9 Experimental Test of Coupling between the Electrodynamics and Dynamics
In the Chapter 16 we shall discuss one of the consequences of our aberration of light theory, the radiation from an X-ray Free Electron Laser. Using XFEL example we can demonstrate that the coupling fields and particles in the conventional approach is intrinsically incorrect. The usual XFEL theory based on the use of old Newtonian kinematics for particle dynamics and the Einstein’s kinematics for the electrodynamics. In fact, the usual theoretical treatment of relativistic particle dynamics involves only a corrected Newton’s second law and is based on the use Galilean edition of velocities. This coupling of Maxwell’s equations and corrected Newton’s equation is commonly accepted as useful method in accelerator physics and, in particular, in numerical calculations of radiation properties. For rectilinear motion of the (ultra-relativistic) modulated electron beam, non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking.
Let us now move on to consider the predictions of the existing XFEL theory in the case of non-collinear electron beam motion. As well-known result of conventional particle tracking states that after an electron beam is kicked by a weak dipole magnet there is a change in the trajectory of the electron beam, while the orientation of the modulated wavefront remains as before. In the absolute time coordinatization the simultaneity of ivents have absolute character. The absolute character of temporal coincidence of events is a consequence of the (hidden) absolute time synchronization convention. According to this old kinematics, the modulation wavefront remains unvaried. In other words, the kick results in a difference between the directions of the electron motion and the normal to the modulation wavefront (i.e. the kick results in a wavefront tilt).
In existing XFEL theory the wavefront tilt is considered as real. However, in the ultra-relativistic limit, the wavefront tilt has no exact objective meaning at all since, due to finiteness of the speed of light, we cannot specify any experimental method by which this tilt could be ascertained. According to conventional (3+1) space-time treatment, a transverse kick does not change the orientation of a modulation wavefront, and hence suppresses the radiation emitted in the direction of the electron beam motion. In fact, according to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the modulation wavefront. This is because Maxwell’s theory has no intrinsic anisotropy. The electrons lying on the plane of simultaneity (i.e. on the modulation wavefront) gives rise to spherical radiation wavelets, and these combine according to Huygens’ principle to form what what is effective a radiated wave. If the plane of simultaneyty is $xy$-plane (i.e. beam modulation wavefront is perpendicular to the $z$-axis), then the Huygens’s construction shows that plane radiation wavefront will be emitted along $z$-direction.
Only solution of the electrodynamics equations in non-covariant form gives correct coupling between the usual particle trajectories and the fields in the lab frame. The non-covariant approach within the framework of both mechanics and electrodynamics predicts an effect in complete contrast to the conventional treatment.
In the Chapter 16 we demonstrate that our Galilean transformed electrodynamics within the lab frame (see Eq.11) says that by making measurement on the coherent radiation, one can observe only radiation in the kicked direction. Experiments (involving specially kicked modulated electron beam) show that this prediction is, in fact, true. The results of XFEL experiments demonstrated that even the direction of emission of coherent undulator radiation is beyond the predictive power of the conventional theory. A starting point in our study of relativity was this complete disagreement between XFEL theory and experiment.
At this point, a reasonable question arises: since XFELs are designed without accounting for the special relativity, how can they actually operate? All existing XFEL codes are based on a model in which the modulated electron beam moves only along the undulator axis. In a typical configuration for an XFEL, the orbit of a modulated electron beam is controlled to avoid large excursions from the undulator axis. However, random errors in the focusing system can cause angular trajectory errors (or ”kicks”). Analysis of the trajectory errors on the XFEL amplification process showed that any XFEL undulator magnetic field must satisfy stringent requirements. However, studies of this critical aspect in the design of a XFEL sources are based on an incorrect coupling of fields and particles. The pleasant surprise is that the tolerances predicted are more stringent than they need be according to our corrected XFEL theory. This can be considered as one of the reason for the exceptional progress in XFEL developments over last decades.
It is generally assumed that the usual Maxwell’s equations and corrected Newton’s second law can explain all experiments that are performed in a single inertial (lab) frame. According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity. There are many physicists who have already received knowledge about special relativity from textbooks and who would say, ”For those who want to learn just enough about special relativity so they can solve problems - it just changes Newton’s laws by introducing a correction factor to the mass.” This is misconception. We cannot take old kinematics for mechanics and Einstein’s kinematics for electrodynamics.
3.10 Bibliography and Notes
1. According to the thesis of conventionality of simultaneity [13, 14, 15, 16, 17], simultaneity of distant events is a conventional matter, as it can be legitimately fixed in different manners in any given inertial reference frame. To quote e.g. Moeller [13]: ”All methods for the regulation of clocks meet with the same fundamental difficulty. The concept of simultaneity between two events in different places obviously has no exact objective meaning at all, since we cannot specify any experimental method by which this simultaneity could be ascertained. The same is therefore true also for concept of velocity.” Let us illustrate a particular special example of the simultaneity convention. Let the synchronization of clocks in different spatial points be provided by light signals having, respectively, velocity $c_{1}$ in the direction parallel to the positive axis $x$, and velocity $c_{2}$ in the opposite direction. Then, a light signal sent from point $A$ at time $t_{A}$ will arrive at point $B$ at time $t_{B}=t_{A}+x_{AB}/c_{1}$. The reflected signal will arrive back at point $A$ at time $t^{\prime}_{A}=t_{B}+x_{AB}/c_{2}$. From these two expressions we get $t^{\prime}_{A}-t_{A}=x_{AB}(c_{2}+c_{1})/(c_{1}c_{2})$. Summing up we have $t_{B}=t_{A}+c_{2}(t^{\prime}_{A}-t_{A})/(c_{1}+c_{2})$. So we come to the synchronization first proposed in [14]: $t_{B}=t_{A}+\epsilon(t^{\prime}_{A}-t_{A})$, where $\epsilon=c_{2}/(c_{2}-c_{1})$. After substituting in this expression Einstein choice $c_{1}=-c_{2}=c$ we get $\epsilon=1/2$. This time order, as fixed by the standard synchronization, is frame dependent. This is the well-known thesis of relativity of simultaneity. The conventional nature of distant simultaneity in special relativity is not to be confused with the relativity of simultaneity. Clearly, the conventionality of simultaneity within a single inertial frame is quite distinct from the relativity of simultaneity in Einstein’s synchronization. If $\epsilon\neq 1/2$, then speed of light from $A$ to $B$ differs from the speed of light from $B$ to $A$. Although the Einstein synchronization choice, is preferred by physicists, it is nothing more ”physical” than any other. A nice pedagogical (but artificial) example of a non-standard synchronization is clock synchronization [11] in which $\epsilon=0$. To any convention on the simultaneity there will correspond a definite choice of the coordinate system in an inertial coordinate system of reference of the Minkowski space-time. A particular very unusual choice of coordinates, the absolute time coordinate choice, is exploited in this book. In our case, when time is absolute, we have $c_{1}=c+v$, $c_{2}=-c+v$. Substituting into the expression for $\epsilon$ we obtain $\epsilon=(1/2)(1-v/c)$.
2. In the text first published in 1923, Eddington discussed, apparently for the first time, a procedure for synchronization using slow transport of clocks [18]. The details can be found in review [8] ( see also [19]).
3. We already pointed that we have empirical access only to the round-trip average speed of light. An empirical test of any property of the one-way speed of light is not possible. Many authors of textbooks still attribute a reality status of the one-way speed of light. To quote Hrasko [20]: ”It is sometimes claimed that Einstein synchronization of distant clocks $A$ ans $B$ is circular. The argument is very simple: Einstein synchronization is based on the equality of light velocity on the path from $A$ to $B$ and back from $B$ to $A$, but measurement of light velocity in one direction between to distant points is impossible unless the clocks at these points have already been synchronized. This argument is, however, fallacious. It is true that one-way measurement of light velocity can be performed only if clocks at the endpoints are synchronized correctly. But since they need not show the correct coordinate time, they can be synchronized without light signals by transporting them from common site in a symmetrical manner. The procedure consists of following steps: … . As we see, the thought experiment described is capable to prove constancy of light speed if it is true, or to disprove it if it is false. It provided, therefore, solid logical foundation for Einstein’s synchronization prescription.” This logical argument is incorrect. Slow clock transport synchronization is equivalent to Einstein’s synchronization in inertial system where the light source is at rest.
4. For a general discussion of correspondence between passive and active transformations we suggest reading the paper [21].
4 Aberration of Light Phenomenon. Inertial Frame of Reference
4.1 Introductory Remarks
The effect of light aberration in an inertial frame of reference is usually understood as a change in the direction of light propagation ascribed to boosted light sources. The phenomenon of aberration of light in an inertial frame of reference by no means simple to describe, even in the first order in $v/c$: a large number of incorrect results can be found in the textbooks. Thus, in order to keep the mathematical complexity of the discussion to a minimum, we will describe the effect of aberration by working only up to the first order. The appearance of relativistic effects in optical phenomena does not depend on a large speed of the radiation sources. Light is always a relativistic object, no matter how small the ratio $v/c$ may be.
The subject of this chapter is the physical influence of the optical instrument on the aberration of light measurement. First, we will demonstrate that when one has finite-aperture mirror moving transversely and the plane wave of light is falling normally on the mirror, there is the aberration (deviation of the energy transport) for light reflected from the mirror. The next problem to be discussed in this chapter is of more practical importance. According to the Babinet’s principle, the problem of reflection from a transversely moving (finite-aperture) mirror complementary to the problem of transmission through a hole in the moving opaque screen. One case of rather great interest is that which corresponds to transmission through a moving end of the telescope barrel. It is shown that the (electromagnetic) wave theory approach to the aberration of light in an inertial frame leads to a substantially different result for energy transport than the predicted by conventional theory.
4.2 The ”Plane Wave” Emitter
We shall try to understand the effect in a very simple case. The explanation of the effect of aberration is actually based on the use of the model of single plane-wave emitter. As a simple model of a plane-wave emitter, we use a two-dimensional array of identical coherent elementary sources (dipoles), uniformly distributed on a given ($x-y$) plane $P$. We take the elementary sources to start radiating waves simultaneously with respect to the lab reference frame where the plane $P$ is at rest. Therefore we have a plane full of sources, oscillating together, with their motion in the plane and all having the same amplitude and phase. Let us suppose that the elementary sources are oscillating at frequency $\omega$. By letting the distance between each two adjacent elementary sources approach zero (i.e. much smaller with respect to the radiation wavelength $\lambda=2\pi c/\omega$), we may consider this two-dimensional arrangement as an ideal plane-wave emitter.
The concept of an (infinite) plane wave is widely used in physics. It is an analytically well-behaved solution of Maxwell’s equations. However, it is not a physically realizable solution, because the total energy content of such a wave is infinite. In any physically realizable situation, one will have to consider a finite source aperture. It is always hiddenly assumed that the detector for the direction of the radiation is an energy propagation detector and the size of the detector aperture is sufficiently large compared with the radiation beam size. Indeed, what is usually considered as an aberration is, in fact, an apparent deviation of the energy transport direction.
4.3 A Moving Emitter. The ”Single Frame” Description
Let us consider the case when the $(x-y)$ plane emitter in the lab inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis. Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the energy transport measurement.
There could be two approaches to the analysis of the aberration o light radiated by a single moving emitter. The first one is the covariant approach. The explanation of the effect of aberration of light in the case of a single moving emitter presented in the literature is actually based on the use of a Lorentz boost to describe how the direction of a beam of light depends on the velocity of the light source relative to the inertial frame of reference. Another non-covariant approach consists in using a ”single frame” description without reference to Lorentz transformations.
The two approaches, treated according to Einstein’s or absolute time synchronization conventions give the same result in the case of a single moving emitter. The choice between these two different approaches in this case is a matter of pragmatics.
In this section we demonstrate both approaches. Let us start with non-covariant approach. We must emphasize that there is no principle difficulty with the a non-covariant (3+1) approach in relativistic electrodynamics. It is perfectly satisfactory. The aberration of light problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations. Within a single inertial frame, the time is an absolute quantity in special relativity also. What does ”absolute” time mean? It means that simultaneity is absolute and there is no mixture of positions and time when sources change their velocities in the inertial frame. A distinctive trait of our non-covariant theory is the absence of relativistic kinematics in the description of aberration of light phenomena. When one has a transversely moving emitter there is the deviation of the energy transport for radiated light. According to the ”single frame” approach, this effect is a consequence of the fact that the Doppler effect is responsible for the angular frequency dispersion of the radiated light waves (Fig. 1 left).
Figure 1: The case when the $(x-y)$ plane emitter in an inertial lab frame is accelerated from rest up to velocity $v$ along the $x$-axis. It is assumed that there is no physical influence of the detector for the direction of the radiation on the measurement. The aberration increment is connected with the physical parameters by the relation: $\theta_{a}=v/c$. The explanation of the effect of aberration is based on the use of a Galilean (left) and Lorentz (right) boost to describe how the direction of a light beam depends on the velocity of light source relative to the lab frame. If make a Lorentz boost, we automatically introduce a time transformation $t\to t+xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation phase front in the lab frame.
Let us consider the electrodynamics of the moving source. The explanation of the phenomenon of radiation in our case of interest consists in using a (active) Galileo boost to describe the uniform translational motion of the source in the inertial lab frame. Maxwell’s equations are not preserved in form by the Galilean transformation, i.e. Maxwell’s equations are not invariant under Galilean transformation. The new terms that have to be put into the field equations due to use of Galilean transformation lead to the prediction of the Doppler effect.
One should not to be surprised to find that electrodynamics problem of moving emitter has intrinsic anisotropy. If fact, anisontropy results directly from the time-dependence of the transverse position of the moving emitter with finite aperture. What must be recognized is that in the time-dependent emitter problem, the results will depend on the direction of the velocity vector.
The present approach to moving emitter problem uses the Fourier transform methods. When we are dealing with linear systems it is useful to decompose a complicated input into a number of more simple inputs, to calculate the response of the system to each of these elementary functions, and to superimpose the individual responses to find the total response. Fourier analysis provides the basic means for performing such a decomposition (1).
[…]
The emitter with finite aperture is a kind of active medium which breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. Let us assume that the dipole density of the elementary source varies according to the law $\rho_{dip}=g(K_{\perp})\cos{({K}_{\perp}x)}$. We conclude, then, that the active medium of the elementary source is sinusoidally space-modulated.
[…]
There is a different physical viewpoint on the Doppler effect of the light beam radiated from a moving emitter that is equivalent to the presented above. The phase of the wave at the world point $(\vec{r},t)$ is invariant at a change of the reference system. This invariance is independent of the coordinate transformation, by which we describe the change of the reference system. Therefore, the phase $\vec{k}\cdot\vec{r}-\omega t$ must be invariant of the Galilean transformation. Consequently, $\vec{k}^{\prime}\cdot\vec{r}^{\prime}-\omega^{\prime}t^{\prime}=\vec{k}\cdot\vec{r}-\omega t$. Substituting the Galilean transformation formulae $x^{\prime}=x-vt,t^{\prime}=t$ into the phase equality formula we obtain $\Delta\omega=K_{\perp}v$. This frequency change coincides with the result derived directly from the dispersion equation, as must be.
Let us first remind the reader of the fact that the usual velocity of waves is defined as given the phase difference between the oscillations observed at two different points in a free plane wave. It is primary used for computing interference fringes that makes phase differences visible. In a plane wave we observe the phase velocity $\omega/k$. Another (group, or energy propagation) velocity can be defined, if we consider the propagation of a peculiarity, that is change in amplitude impressed on a train of waves. A simple combination of groups obtains when two waves $\omega_{1}=\omega+\Delta\omega,~{}k_{1}=k+\Delta k$ and $\omega_{2}=\omega-\Delta\omega,~{}k_{2}=k-\Delta k$ are superimposed. This represents a carrier with frequency $\omega$ and a modulation with frequency $\Delta\omega$. The wave may be described as a succession of moving beats (or groups). The carrier’s velocity is $\omega/k$, while the group velocity is given by $v_{g}=\Delta\omega/\Delta k\to d\omega/dk$.
Many textbooks on electromagnetic theory discuss the aberration of light phenomena in the context of plane wave. However, in dealing with plane wave one will have an incorrect model of the aberration of light. When an infinite sinusoidal wave travels, there is a uniform average energy density throughout the space. Does this energy remain where it is, or does it propagate through the space? It is impossible to know this. All experimental methods for measuring the aberration of light operate with light signals, and hence do not measure the phase velocity but the signal velocity and this velocity coincides with group velocity.
In our example, the plane waves with different wavenumber vectors propagate out from the moving emitter with the different frequencies. Then equation $\Delta\omega/\Delta k_{x}=v$ holds for each scattered waves independently on the sign and the magnitude of the radiated angle. In fact, the $\Delta\omega$ in our case of interest is the Doppler shift $\Delta\omega=\vec{K}_{\perp}\cdot\vec{v}$ and the $\Delta k_{x}$ is simply the transverse component of the radiated wavenumber vector $\Delta k_{x}=K_{\perp}$. The last equations state that radiated light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=v$.
4.4 A Moving Emitter. Explanation on the Basis of the Lorentz Transformations
It should be note, however, that there is another satisfactory way of explaining the effect of aberration of light from the moving source. The explanation consists in using a Lorentz boost to describe the uniform translation motion of the light source in the lab frame. On the one hand, the Maxwell’s equations remain invariant with respect to Lorentz transformations. On the other hand, if make a Lorentz boost, we automatically introduce a time transformation $t^{\prime}=t-xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation phase front in the inertial frame. This is because the effect of this time transformation is just a dislocation in the timing of processes, which has the effect of rotating the plane of simultaneity on the angle $v/c$ in the first order approximation.
According to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the radiation phase front. This is because Maxwell’s equations have no intrinsic anisotropy. In other words, when a uniform translational motion of the source is treated according to Lorentz transformations, the aberration of light effect is described in the language of relativistic kinematics. According to the relativistic kinematics, the extra phase chirp $d\phi/dx=k_{x}=v\omega_{i}/c^{2}$ is introduced and the array of identical elementary sources of the moving emitter now have different phases. As a consequence of this, the plane wave wavefront rotates after the Lorentz transformation. Then, the radiated light beam is propagated at the angle $v/c$, yielding the phenomenon of the aberration of light (Fig. 1 right).
We now ask about the group velocity of the radiated beam. With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ dispersion equation in the case of Maxwell’s electrodynamics is reduced to $k_{z}^{2}+k_{x}^{2}-\omega^{2}/c^{2}=0$. From the initial conditions and the Lorentz transformation we find that $\omega=(\omega_{i}+vK_{\perp})$, $k_{z}=\sqrt{\omega^{2}_{i}/c^{2}-K_{\perp}^{2}}$, $k_{x}=(v\omega_{i}/c^{2}+K_{\perp})$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated dipole density, $\omega_{i}$ is the frequency of the emitter radiation before the acceleration. Substituting these expressions in dispersion equation we find that the latter is satisfied, as must be.
As one of the consequences of the Doppler effect in the Lorentz coordinatization, we find an angular frequency dispersion of the light waves radiated from the moving emitter with finite aperture. The Doppler shift, $\Delta\omega$, of radiated light wave (in the first order approximation) is given by $\Delta\omega=\vec{K}_{\perp}\cdot\vec{v}$, where $K_{\perp}$ is the transverse component of the radiated wavenumber vector. The last equation state that radiated light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=v$.
It is interesting to discuss what it means that there are two different (covariant and non-covariant) approaches that produce the same group velocity. The point is that both approaches describe correctly the same physical reality and since the group velocity has obviously an objective meaning (i.e. convention-invariant), both approaches yield the same physical results.
The difference between the absolute time coordinatization and the Lorentz coordinatization is very interesting. In the Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization. We can interpret manipulations with rule-clock structure in the lab frame simply as a change of the time variable according to the transformation $t\to t+xv/c^{2}$. The overall combination of Galileo transformation and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the inertial frame. This variable change has no intrinsic meaning. One can see the connection between the time shift and the issue of clock synchrony. The convention-independent results of calculations are precisely the same in the new variables. As a consequence, we should not care to transform the results of the electrodynamics problem solution into the original (3+1) variables.
An idea of studying the relativistic electrodynamics using technique involving a change of variables is useful from a pedagogical point of view. It is worth remarking that the absent of a dynamical explanation for wavefront rotation in the Lorentz coordinatization has disturbed some physicists. It should be clear from the discussion in the Chapter 3 that a good way to think of the wavefront rotation is to regard it as a result of transformation to a new time variable in the framework of the Galilean (”single frame”) electrodynamics.
4.5 Reflection from a Mirror Moving Transversely. Mistake in Existing Theory
It is generally believed that for a mirror moving tangentially to its surface the law of reflection which holds for the stationary mirror is preserved, as shown in Fig.2. In other words, the velocity of the energy transport is equal to the phase velocity. This statement, presented in most textbooks, is incorrect.
First, we examine the reasoning presented in textbooks (2). The reflection from the mirror is analyzed in two Lorentz reference frames. The fixed (lab) frame $S$ is at rest with respect to the plane-wave emitter. The moving frame $S^{\prime}$ has velocity $v$. In this frame, the mirror is at rest. In both frames, we use a Cartesian coordinate system in which $x-y$ plane is tangent to the reflection surface. The $x$ direction coincides with the direction of $v$. For simplicity, we consider the case in which light is incident from the $z$ direction in the lab frame. Incident light is described by its four-dimensional wave vector, whose time like component is the angular frequency $\omega$ and whose space like components define the direction of propagation. In the lab frame, $(t,x,y,z)$, this vector has components $k_{1}=(\omega,0,0,-\omega/c)$, where the negative sign indicates propagation towards the mirror (Fig. 3a). Our task is to determine the wave vector for the reflected beam.
Figure 2: Transversely moving mirror with small aperture at normal incidence. According to textbooks, there is no deviation of the energy transport for the reflected light beam. A monochromatic plane wave of light is falling normally on the small moving aperture mirror, and generates a reflected oblique beam.
Figure 3: The effect of aberration of light is described in the language of relativistic kinematics, in terms of the wavenumber four vector. Geometry of the reflection as seen from (a) lab frame, (b) inertial frame moving with the same velocity as the mirror. According to textbooks, there is no aberration for light reflected from a tangentially moving mirror.
The argument that there is no aberration for light reflected from a transversely moving mirror runs something like this. It is easiest to consider the reflection in the moving frame (Fig. 3b). In this frame, the surface is at rest, so the usual laws of optical reflection apply. We will describe the effect of aberration of light by working only up to the first order $v/c$. An observer moving with the mirror surface sees the wave vector $k_{1}^{\prime}=(\omega,-v\omega/c^{2},0,-\omega/c)$. The effect of reflection is to reverse the sign of the $z^{\prime}$ component of the wave vector, $k^{\prime}_{2}=(\omega,-v\omega/c^{2},0,\omega/c)$. We now obtain the reflected wave vector in the lab frame by applying the inverse Lorentz transformation: $k_{2}=(\omega,0,0,\omega/c)$.This vector represents a light beam traveling away from the mirror, having the same frequency as the incoming beam. This shows that the beam is reflected according to the usual geometrical optics laws, and the beam suffers no aberration.
There are two errors in solving problems concerning reflection from a mirror moving transversely. The first error in the last argument follows from the fact that the concept of a plane wave and an infinite plane mirror is used in the case of the tangential motion. The absurdity of this, i.e. the impossibility of such a motion ever been detected is put forward as an obvious argument against the reasoning presented in textbooks. Indeed, if the mirror is infinite the problem is not time dependent (we must conclude that when we are dealing with hidden assumption that problem is not time dependent we have no aberration and that is not too surprising). It is obvious that only the motion of a finite mirror is meaningful. In view of the electrodynamics, only a velocity of the finite mirror has a physical meaning. The reasoning in textbooks ignores completely the interaction of the light and the moving mirror edges.
We would also noticed that the concept of plane wave is used in textbooks. There is a common mistake made in the electrodynamics connected with the energy transport direction in the case of an (infinite) plane wave. When an infinite plane wave travels, there is a uniform average energy density throughout the space. It is impossible to know the energy transport direction when one has deal with a plane wave. All experimental methods for measuring the aberration increment operate with light signals, and hence do not measure the phase velocity (i.e. frequency and wavenumber vector) but the group velocity. It can be defined only if we consider the propagation of a peculiarity, that is change in amplitude impressed on a train of waves. What authors of textbooks generally overlooked is the fact that the energy transport problem is well-defined only if the source and mirror apertures have already been specified.
[…]
The second error in the presented textbook argument is conceptual in nature. Suppose that an observer, which is at rest with respect to the inertial frame performs the direction of the energy transport measurement. The observer describes the light beam from the emitter at rest using diagonal (Minkowski) metric. According to textbooks, the metric of the moving mirror is the same (i.e. Minkowski metric) as metric of the emitter at rest. The Maxwell’s equations can be applied in the lab inertial frame only in the case when Lorentz coordinates are assigned. The authors of textbooks incorrectly belief that the common Lorentz time coordinate axis for emitter and mirror can be assigned. This is misconception. The question arises how to assign synchronization in the lab frame after the mirror acceleration.
The peculiarity of this problem with the viewpoint of relativistic kinematics is that here the emitter (and observer with his measuring devices) is at rest in the lab inertial frame and the mirror is moving with the constant speed with respect to the lab frame and interacts with the radiated light beam. How can we solve a problem involving the emitter-mirror relative velocity? Suppose that we assign the Lorentz time coordinate for the description of the emitter radiation. But this will be the absolute time coordinatization for the boosted mirror and this boost will be described in such coordinatization by the Galilean transformation. Without changing synchronization in the lab frame we can prepare, for mirror and emitter, a common set of synchronized clocks only in the case of absolute time coordinatization i.e in the case when simultaneity is absolute.
Suppose that we re-synchronize clocks in the lab frame in order to assign the Lorentz time coordinate for the boosted mirror. In this coordinatization, we describe the reflection of light using the usual Maxwell’s equations. On the other hand, the effect of this time transformation is just a rotation of the radiation phase front of the incoming plane wave on the angle $v/c$. As a result, this new time coordinate in the lab frame is interpreted by saying that Maxwell’s equations are not applicable to the emitter radiation description. It is not difficult to see that the peculiarity of this coordinatization is that here the energy transport velocity is not equal to the phase velocity for incoming light beam.
This problem in special relativity can be adequately treated by an approach which uses the absolute time coordinatization. We will demonstrate that when one has a transversely moving mirror with finite aperture and a plane wave of light is falling normally onto the mirror, there is a deviation of the energy transport for the reflected light. In this case, the effect of aberration of light results directly from the time-dependence of the position of the moving mirror, since a finite aperture has to be considered. What must be recognized is that in the time-dependent emitter-mirror problem, the solution involves light beams with different frequencies. It comes out quite natural that the result will depend on the velocity vector: in other words, we account to the influence of the Doppler effect. For a transversely moving mirror with finite aperture we cannot neglect the angular frequency dispersion, which is an effect of the first order in $v/c$.
We already pointed that Chapter 3 is probably the most important chapter in this book. To our knowledge, neither operational interpretation of the absolute time coordinatization nor the difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view, are given elsewhere in the literature. Using the Galilean boost: we substitute $x\to x-vt$, while leaving time unchanged into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain metric Eq. 11. This metric describes the electrodynamics of the moving mirror with viewpoint of the inertial observer measurements.
There is an intuitively plausible way to understanding why authors of textbooks conclude that there is no aberration for light reflected from a transversely moving mirror. A comparison between the conventional aberration of light theory and the sound theory might help here. Bellow in the Section 4.15 we shall consider the sound emitter at rest in the atmosphere frame. We shall consider the case of a mirror, in the atmosphere frame, moving with velocity $v$ along its surface. According to the theory of sound, there is no deviation of the energy transport for reflected beam. The group velocity is equal to the phase velocity. A plane wave is falling normally on the mirror, and generates an oblique sound beam. The (diagonal) wave equation of the moving mirror in the atmosphere frame is the same as wave equation of the emitter at rest. According to textbooks, there are important parallels between aberration of sound and light in the first order approximation. This is misconception. There is a fundamental difference between velocity of light and velocity of sound. The satisfactory treatment of the aberration of light phenomena should be based on the pseudo-Euclidean space-time geometry.
4.6 Way to Solve the Emitter-Mirror Problem in the (3+1) Space and Time
The aberration of light problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations. When the illumination of the object originates from a monochromatic spatially coherent source there exists a method for calculating the reflected intensity that has the special appeal of conceptual simplicity. It uses Fourier transforms of spatial filtering theory that is the Abbe diffraction theory.
The essence of Abbe’s approach, in our case of interest, is that one regards the mirror as a kind of a diffraction grating which breaks up the incident beam of the plane wave into a number of diffracted beams constituted by plane waves. Each of these beams corresponds to one of the Fourier components into which the reflected power of the mirror can be resolved. The finite-aperture mirror is a non-periodic object. It gives an infinite number of diffracted beams forming a continuum.
A simple example of a diffraction grating is shown in Fig.4. Let us assume that the reflectance of the grating varies according to the law $R=g(K_{\perp})\cos{(\vec{K}_{\perp}\cdot\vec{r})}$. The reflectance is sinusoidally space-modulated. It should be noted that the permanent reflectance distribution grating discussed here is only our mathematical model and we do not need to discuss how it can be created.
Figure 4: The Bragg diffraction grating at normal incidence. The reflectance is sinusoidally space-modulated.
The $\vec{k}$ vectors shown in Fig.4 represent the propagation vector of the incident plane wave $\vec{k}_{i}$, which is assumed to be directed perpendicularly to the surface. The vectors $\vec{k_{s}}^{(+)}$ and $\vec{k_{s}}^{(-)}$ are added to indicate the scattered light. The Bragg condition $\vec{k}_{s}=\vec{k}_{i}\pm\vec{K}_{\perp}$ shows how the direction of the incident and scattered wave are related. The first-order maxima dominate due to the fact that light is being scattered from a sinusoidal grating, rather than a set of discrete planes (grooves).
We assume that the $\vec{K}_{\perp}$ vector is directed parallel to the side of the diffraction grating with the incident wave impinging on the grating perpendicularly, as shown in Fig.4. The length of the vectors $\vec{k}_{s}$ and $\vec{k}_{i}$ must, of course, be the same, but the vector diagram does not quite match up. The Bragg conditions are then not satisfied precisely. For small angles, $\vec{k}_{s}=\vec{k}_{i}\pm\vec{K}_{\perp}$ still holds approximately, so that we obtain the scattering angle $\theta=K_{\perp}/k_{i}$.
When the scatterer wave is a progressive wave rather than a fixed modulation, the frequency of the scattered wave is different from that of the incident wave. This fact is interpreted as a Doppler effect, since the reflection is from a moving, rather than a stationary, set of waves. In the case of a transversely moving grating, light in the diffraction maxima undergoes a Doppler shift resulting from the fact that it has been reflected from moving waves with wavenumber vectors $\vec{K}_{\perp}$.
Let us demonstrate that the new terms that have to put into the field equations due to use of absolute time coordinatization lead to the prediction of the Doppler effect. We recall that with a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ dispersion equation in the absolute time coordinatization becomes: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c+k_{z}^{2}-\omega^{2}/c^{2}=0$. The wavenumber vector of the radiated plane wave is fixed by initial conditions. In fact, $k_{z}=\sqrt{\omega^{2}_{i}/c^{2}-k_{x}^{2}}$, $k_{x}=K_{\perp}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated reflectance. From this dispersion equation, we find the requirement that the wavenumber $K_{\perp}$ and the frequency change $\Delta\omega$ are related by $\Delta\omega=K_{\perp}v$.
The following important detail of such ”single inertial frame” description can hardly be emphasize enough. If the source of light is at rest and the mirror is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves. In other words, the dispersion equation in the absolute time coordinatization should be applied and kept in a consistent way for both incoming and scattered waves (Fig. 4). In our previous discussion of absolute time coordinatization we learned that the emitter at rest must be in the same time described by Maxwell’s electrodynamics. A dispersion equation in the case of Maxwell’s electrodynamics is reduced to $k_{i}^{2}-\omega_{i}^{2}=0$. From the initial conditions we find that $\vec{k}_{i}=\vec{e}_{z}k_{z}$, $\omega_{i}=ck_{z}$. The contradiction, however disappears if we perform geometrical analysis of light reflection. The peculiarity of the discussed geometry is that even after the Galilean transformation along the $x$-axis the dispersion equation in the absolute time coordinatization will have the same (diagonal) form $k_{z}^{2}-\omega^{2}=0$ for the incident wave. We worked out a very simple case in order to understand all the physical principle very clearly. In Section 5.11 we shall analyze the aberration of light with incoming wave which has rate of change of phase with respect to $x$-axis.
[…]
Figure 5: Transversely moving mirror with a small aperture at normal incidence. When a plane wave of light is falling normally on the mirror, there is a deviation of the energy transport for the reflected light beam. This effect is a consequence of the fact that the Doppler effect is responsible for angular frequency dispersion of the light waves reflected from the mirror. As a result, the velocity of the energy transport is not equal to the phase velocity.
One of the most important conclusions of the foregoing discussion is a remarkable prediction on the theory of the aberration of light, concerning the deviation of the energy transport for light reflected from a mirror moving transversely. Namely, when a plane wave of light is falling normally on the mirror, there is a deviation of the energy transport for reflected light beam (see Fig. 5). This phenomenon can be regarded as a simple consequence of the Doppler effect.
The argument that in the process of reflection from a transversely moving mirror the direction of propagation is not given by the normal to the wavefront is considered erroneous in literature (4). This fact is ascribed to a lack of understanding of the difference between convention-dependent and convention-invariant parts of the theory. The direction of the energy transport has an exact objective meaning i.e. is convention-invariant. However, the phase front orientation (i.e. the plane of simultaneity in the judgement of an observer) has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this orientation could be ascertained.
Since the phase front orientation does not exist as physical reality within the angular range $v/c$, a question arises: why do we need to account for the exact phase front orientation in our electrodynamics calculations? The answer is that when the evolution of the radiation beam is treated according to the single inertial frame of reference, one will experience that phase front orientation remains unvaried: this has no objective meaning but is used in the analysis of the electrodynamics problem. A comparison with a gauge transformation in Maxwell’s electrodynamics might help here. Even if the phenomena are quite different, the common mathematical formulation permits us to draw this analogue.
4.7 Discussion
Physicists who try to understand the situation related to the use of the textbook approach in the aberration of light phenomena, are often troubled by the fact that in the situation where there is a mirror moving normally to its surface the reasoning presented in textbooks is correct. The reflection from the mirror is analyzed in two Lorentz reference frames. It is interesting to note that, for this case, relativistic kinematics correctly predicts the light frequency variation on reflection from a moving mirror. At this point a reasonable question arises: why the same method gives incorrect result in the case of reflection of light from a mirror moving tangentially to its surface?
It is not difficult to see that the peculiarity of the situation where there is an infinite mirror moving normally to its surface is that here the ”mirror motion” is a real observable effect. Indeed, even if the mirror is infinite the problem is time dependent. In this situation, a velocity of the infinite mirror has physical meaning. Thus, for normal motion, the infinite mirror may be a useful concept for analysis of the Doppler effect. This method was incorrectly extended to the case of tangential motion.
The second peculiarity of the collinear geometry where there is a mirror moving normally along the same line as the radiated beam is that here the velocity is perpendicular to the plane of radiation wavefront (i.e plane of simultaneity). Thus, for collinear motion, the plane of simultaneity in the absolute time coordinatization will have the same orientation for the Lorentz coordinatization. We shall discuss this subject further in the Chapter 10.
4.8 Large Aperture Mirror
In this chapter our discussion is limited to the region of problem parameters, in which we forget about the emitter edges. Although this aberration of light theory is just an approximation, it is a very great importance practically. We shall also discuss the situation where the transverse size of the emitter is very small relative to the transverse size of the moving mirror. It should be note that this situation is not realized in the stellar aberration measurements. We shall work out this case in order to understand all the physical principles very clearly. It is easy to show that the deviation of the energy transport is absent in this case, Fig. 6. The direct approach to moving mirror problem uses the Fourier transform methods. We now consider another method of calculating the aberration of light effect - we want to illustrate the great variety of possibilities. The way of thinking that made the law about the behavior of light reflected from a large aperture mirror evident is based on Babinet’s principle. It is well known that, when light comes through a hole of a given shape, made in an opaque screen, the distribution of intensity after the hole (i.e. the diffraction pattern) is the same as in the case when the hole is replaced by sources (dipoles) uniformly distributed over the hole. In other words, the diffracted plane wave from a hole, or from a source with the same shape of the hole are the same.
Figure 6: Transversely moving mirror with a large aperture at normal incident. When a beam of light is falling normally on the mirror, there is no deviation of the energy transport for the reflected light beam. The velocity of the energy transport is equal to the phase velocity.
This is a particular case of Babinet’s principle, which states that the sum of diffraction fields behind two complementary opaque screens is the incident wave. We know from this principle, that the solution we have found using Abbe’s approach also corresponds to that for large aperture hole in a moving opaque screen. We see clearly that there is no electromagnetic interaction of a light beam with screen. Indeed light beam is not scattered by the hole edges. Does this discussion about large aperture hole have any meaning? To see whether it does, we should remember about the Babinet’s principle. Here we only wish to show how easy the law of reflection from a large aperture mirror can be found with the help of the Babinet’s principle.
4.9 Analysis of Transmission through a Hole in a Opaque Screen
Above we demonstrated that when one has a small aperture mirror moving transversely and the plane wave of light is falling normally on the mirror, there is the aberration (deviation of the energy transport) for light reflected from the mirror. The problem to be considered in this section is of more practical importance. We now consider the case of a screen, in the lab frame, moving with velocity $v$ along its surface. It is generally believed that there is no deviation of the energy transport for light transmitted through a hole in the moving opaque screen, Fig. 7.
However, there is a common mistake made in relativistic optics, connected with aberration effects from a transversely moving screen containing a hole. We describe the system using, again, a Fourier transform method similar to that considered above. The screen containing a hole is a kind of diffraction grating which breaks up the incident beam of the plane wave into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which a transmitted light beam can be resolved.
The gratings discussed so far modulate the amplitude of the incident plane wave by a periodic reflection function. However, we can immediately extend the range of validity of our analysis to gratings that modulate the amplitude of the incident light by a periodic transmission function. Let us assume that the transmittance of the grating varies according to the law $T=g(K_{\perp})\cos{(\vec{K}_{\perp}\cdot\vec{r})}$, Fig.8. The transmittance is sinusoidally space-modulated. All the equations that we derived so far hold immediately for the forward scattered beams.
Figure 7: Aberration of light in an inertial frame of reference. Transversely moving screen which has a hole in it. According to textbooks, a monochromatic plane wave of light is falling normally on the screen and generates a transmitted oblique beam. There is no deviation of the energy transport for the transmitted oblique light beam. The velocity of the energy transport is equal to the phase velocity.
Figure 8: The Bragg diffraction grating at the normal incident. The transmittance is sinusoidally space-modulated.
Figure 9: Aberration of light in an inertial frame of reference. Transversely moving screen which has a hole in it. A monochromatic plane wave of light is falling normally on the screen and generates a transmitted light beam. The Doppler effect is responsible for the angular frequency dispersion of the light waves transmitted through the hole. As a result, the velocity of the energy transport is not equal to the phase velocity.
According to our approach, there is a remarkable prediction of the theory of aberration of light concerning the deviation of the energy transport for light transmitted through a hole in a moving screen. Namely, when one has a transversely moving screen with a hole in it and a plane wave of light is falling normally on the screen, there is a deviation of the energy transport for light transmitted through the hole (see Fig. 9).
4.10 Spatiotemporal Transformation of the Transmitted Light Beam
Let us suppose that transmitted light pulse propagates in the $x-z$ plane. Now we are interested in the space-time intensity distribution in this plane. Spatiotemporal coupling arises naturally in transmitted radiation behind the screen, because the transmission process involves the introduction of an angular-frequency dispersion of the transmitted radiation. The emitted light beam is represented with sufficient accuracy as the product of factors separately depending on space and time. However, when the manipulation of the emitted light requires the transmission through a hole in a moving opaque screen, such assumption fails.
[…]
4.11 Applicability of the Ray Optics
Let us move on to consider the predictions of the existing aberration of light theory in the case of a transversely moving screen which has a hole in it (Fig. 7). According to conventional theory, the spot of the transmitted light beam on the observer screen also moves with the same velocity as the moving screen. It is important at this point to emphasize that the electrodynamics dictates that this would also lead to a consequent introduction of an angular-frequency dispersion. However, such angular spectrum change would mean a correction to a deviation of the energy transport direction of the transmitted light beam so that there is a glaring conflict with the prediction of the energy transport direction according to conventional aberration of light theory. The absence of the group velocity along the moving direction is the prediction of conventional aberration of light theory and is obviously absurd from the viewpoint of electrodynamics.
Figure 10: Transverse moving screen which has a hole in it. Light corpuscles are falling normally on the screen and, according to literature, generates an oblique light beam.
This incorrect statement is a straightforward consequence of the generally accepted way of looking at the aberration of light phenomena of most authors of the texbooks. Today one is told that the phenomenon of aberration of light could be interpreted, using corpuscular model of light. Light corpuscles are falling normally on the moving screen and, according to literature, generate oblique light beam as Fig. 10 shows. This wrong argument persists to this day. If the optical system is spatially coherently illuminated, then a satisfactory treatment of the aberration of light should be based on the electromagnetic wave theory.
Some experts believe that the applicability of corpuscular model in the theory of light should be reinterpreted as the applicability of ray optics. Let us see what happens, according to the ray optics, in our case of interest. Light rays are falling normally on the moving screen and generate oblique ray beam as Fig. 10 shows (6). In this situation, we just treat light rays like little particles and the effect is entirely familiar. We would like now to discuss the region of applicability of ray optics. The situation relating to use the ray optics in the theory of aberration of light is complicated. One could naively expect that the region of applicability of ray optics, following from the textbooks reasoning, should be identified with any spatially incoherent radiation. However, incorrect results are obtained by doing so. In particular, a spatially completely incoherent source (e.g. an incandescent lamp or a star) is actually a system of elementary (statistically independent) point sources with different offsets. Radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individual elementary point sources. An elementary source produces in front of a hole aperture effectively a plane wave. In other words, the transmission process involves the introduction of an angular frequency dispersion of the transmitted radiation. It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single point source characteristics like a deviation of the energy transport direction. This argument gives reason why ray optics is not applicable in the theory of aberration of light from (spatially) completely incoherent sources.
We will illustrate the applicability of ray optics in the theory of aberration of light for a particular class of spatially incoherent light beams. In fact, the ray beam shown in Fig. 10 can be realized as follows. Such beam may be produced by many commonly used lasers with a random spread of phases. One of the very useful properties of laser sources is their ability to produce fields that are highly directional.The intensity of such fields is concentrated in a very narrow solid angle. A method can be proposed for generating ray beam from primary sources by the use of array of randomly phased lasers. Such planar source generates rays which are falling normally on the moving screen (within the laser Rayleigh range) and generate oblique transmitted ray beam, Fig. 10. Intuitively, a transversely moving screen containing a hole acts like a switcher for lasers. Surely, a luminous spot moving at velocity $v$ can be realized simpler, so to speak, ”manually”. We can arrange the laser-like sources along the $x$ axis and switch them on one after another (independently) from left to right with a given time lag. Naturally, we can get a luminous spot moving at any velocity (even at $v>c$). From this example it is seen that in this process no information can be transmitted (along the $x$ axis) since each source radiates independently.
4.12 Measure of Aberration
[…]
4.13 Moving Large Aperture Emitter
Let us now consider the case when a ”plane-wave” emitter in the lab inertial frame is accelerated from rest up to velocity $v$ along the $x$ axis. An emitter with finite aperture is a kind of active medium which breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. We already know from our discussion from very beginning of this chapter that there is a deviation of the energy transport for the coherent light radiated by the transversely moving emitter, which is nothing else but well-known result: there is the aberration of light from the transversely moving emitter in the inertial frame of reference (Fig. 1).
The specific of our case of interest with the viewpoint of kinematics is that here the screen is at rest in the lab frame and the emitter is moving tangentially to its surface with constant speed with respect to the lab frame. For simplicity we shall assume that the transverse size of the moving ”plane-wave” emitter is very large relative to the transverse size of the hole in the screen. Suppose that an observer, which is at rest with respect to the screen performs the direction of the energy transport measurement, Fig.11.
Figure 11: Aberration of light in an inertial frame of reference. The large aperture ”plane-wave” emitter moving tangentially to its surface. As one of the consequences of the Doppler effect, we find group velocity of the light waves radiated off a large aperture moving emitter. The screen is at rest and we have actually the problem of steady-state transmission. The transmitted light beam is going vertically because it has lost its horizontal (group) velocity component.
If the screen is at rest and the source of light is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves. In our previous discussion we learned that the screen at rest must be described by Maxwell’s electrodynamics. In other words, the dispersion equation in the Lorentz coordinatization should be applied and kept in a consistent way for both incoming and scattered waves. A dispersion equation in the case of Maxwell’s electrodynamics is $k_{x}^{2}+k^{2}_{y}+k^{2}_{z}-\omega^{2}=0$. The peculiarity of the discussed geometry is that even after the Galilean transformation along the $x$ axis the dispersion equation in the absolute time coordinatization will have the same diagonal form $k^{2}_{z}-\omega^{2}=0$ for the (normal) incident wave. At close look at the physics of this subject shows that in the inertial lab frame, where the screen is at rest, we have actually the problem of steady-state transmission. The Doppler effect is absent and the transmitted beam is going vertically because it has lost its horizontal (group velocity) component. That is the transmission appears as shown in Fig.11. We only wish to emphasize here the following point. When the light passes through the small aperture hole we have a light beam whose fields have been perturbed by diffraction, and now not include information about emitter motion.
4.14 A Point Source in an Inertial Frame of Reference
Above we considered a single moving ”plane wave” emitter in an inertial frame of reference. In the description of the aberration of light in an inertial frame there are two choices of sources useful to consider:
(a) A ”plane wave”-like emitter
(b) A point-like ( or, more generally, spatially completely incoherent) source
Source field diffraction can be divided into categories - the Fresnel (near-zone) diffraction and Fraunhofer (far-zone) diffraction. In Fraunhofer diffraction, the phase of the wave is assumed to vary linearly across the detector aperture. This would occur if, for example, a plane wave were incident on the aperture at an angle with respect to the optical axis. In the Fresnel diffraction, we replace the assumption of a linear phase variation with quadratic phase variation. In the far zone, both types of sources produce in front of pupil detection effectively a plane wave. In other words, there is always the physical influence of the instrument on the measurement of the aberration of light in the Fraunhofer zone.
[…]
Figure 12: Aberration of light in an inertial frame of reference. A point source produces in front of a hole aperture effectively a plane wave. If the motion of the point source is parallel to the screen, transmitted beam is going vertically. The aberration of light phenomenon is absent in this situation.
Figure 13: A drawing of the modified source of the ”plane wave” emitter type. A point source of light is placed in the front focal plane of the lens.
One of the important conclusions of the discussion presented above (see Fig. 11 ) is that the aberration of point source is absent in an inertial frame, Fig. 12. There are a number of remarkable effects which are a consequence of the fact that the information about a point source motion is not included into the light beam transmitted through the hole. In fact, this is the key to the binary star paradox discussed in the Chapter 6.
For the rest of our discussion of the aberration of light phenomena in an inertial frame, it will be more convenient if we consider a somewhat modified source of the ”plane wave” emitter type. We have chosen the model described in the Section 4.2 because it is relatively simple. It is sufficiently complicated that it can stand as a prototype which can be generalized for the description of the aberration of light phenomena. No one has ever done all of ”thought experiments” we described in just this way, but we know what would happen from the laws of special relativity which are, of course, based on other experiments.
Fig. 13 shows a drawing of the modified source of the ”plane wave” emitter type. The setup looks more complicated at first, but it is practically possible to make. A point source of light is placed in the front focal plane of the lens. It is assumed that the order of magnitude of the dimensions of the ”point” source is about $\lambda$, where $\lambda$ is the optical radiation wavelength.
At this point, a reasonable question arises: why we are not discuss about the aberration of light radiated by a single laser. At first glance, the output light beam shown in Fig. 13 may be produced by a laser. Indeed, one of the specific properties of a laser is ability to produce coherent light beams that are highly directional. We will not be able actually to get into the subject until later. We will analyze the aberration of light transmitted from the laser resonator which is accelerated in an inertial frame in the Section 8.6 and 10.4.
4.15 The Difference between the Aberration of Light and Sound
We are now finished with the subject of the aberration of light in the initial inertial frame, and will soon go on to discuss the optical phenomena in the accelerated frame. But before doing so, we would like to discuss the optical-acoustic analogies. Sound is a periodic motion of the air caused by vibrations of parts of the source. The explanation of the effect of sound aberration is actually based on the use of the model of single plane-wave emitter. As a model, we use a two-dimensional array of identical coherent elementary sources uniformly distributed on a given $(x-y)$ plane $P$. We take the elementary sources to start radiation waves simultaneous.
We shall discuss the situation where there is a finite aperture mirror moving tangentially to its surface. For simplicity, we shall assume that the transverse size of the moving mirror is very small relative to the transverse size of the ”plane-wave” emitter. The observer describes the sound beam from the emitter at rest (in the atmosphere frame) using usual wave equation $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$, where $v_{s}$ is the sound speed. This emitter will radiate plane wavefront in a vertical $z$ direction. The wave equation of the moving mirror in the atmosphere frame is the same as wave equation of the emitter at rest. We may consider the amplitude of the beam radiated by plane of oscillating sources as a whole to be resultant of radiated spherical waves. This is because diagonal wave equation has no intrinsic anisotropy. The elementary sources lying on the plane of simultaneity gives rise to spherical radiated wavelets, and these combine according to Huygens principle to form what is effectively a radiated wave. The wavefront emitted by the plane-wave emitter at given instant time is constructed as an envelope of all the spherical wavelets from the point sources on the emitter that have radiated until this instant. According to the theory of sound, there is no deviation of the energy transport for the reflected sound beam. A plane wave is falling normally on the finite aperture mirror, and generates a reflected oblique beam. Note that the phase and group velocities are always equal in the atmosphere frame.
According to textbooks, there is no difference between the aberration of light (in the first order in $v/c$) and sound. The geometry of moving mirror reflection is identical, Fig. 2. We now consider the case of a screen, in the atmosphere frame, moving with velocity $v$ along its surface. According to the theory of sound, there is no deviation of the energy transport for sound transmitted through a hole in the moving screen. Consequently, it is generally believed that the geometry of the energy transport for light transmitted through a hole in the moving opaque screen is the same, Fig. 7. In the next chapter (see section 5.14) we shall continue our discussion of the aberration of sound theory.
The textbook approach to the aberration of light usually forces the physicist to believe that there are important parallels between the aberration of sound and light in the first order approximation. In fact, the wave equation of the moving mirror (screen) is the same (i.e. diagonal wave equation) as wave equation of the emitter at rest. On the other hand, in the sound theory we talk about Newton space and time rather than about space-time with pseudo-Euclidean geometry. We come to the conclusion that the standard analysis of the aberration of light in the single inertial frame does not take into account the fundamental difference between velocity of light and velocity of sound.
The satisfactory treatment of the relativistic modification should be based on the space-time geometric approach. The usual Minkowski metric is valid in inertial frame where observer is at rest and the procedure of slow clock transport (or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest) is used to assign values to the time coordinate. In this (Lorentz) coordinate system Minkowski metric describes the electrodynamics of the light source at rest with viewpoint of the inertial observer measurements. Now we consider the case when mirror in the inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis. The method of synchronization consists in keeping the same set of synchronized clocks used. This synchronization convention preserves the Minkowski metric of the inertial observer and the light source at rest. In this case we are dealing with the Galilean boost of the mirror. We substitute $x\to x-vt$, while leaving time unchanged into the Minkowski metric of the inertial observer to obtain Eq. 11. This metric describes the electrodynamics of the moving mirror with viewpoint of the inertial observer measurements. In contrast, the accepted in the previous literature incorrect assumption that an inertial observer (i.e. observer at rest in the inertial frame) and the accelerated ( in the same inertial frame) mirror have common Minkowski metric has always been considered obvious. So that the commonly accepted approach to special relativity in the first order approximation and sound theory are really just variants of the same classical theory.
4.16 Bibliography and Notes
1. For a general discussion of the Fourier transform methods of spatial filtering theory or Abbe diffraction theory we suggest reading the book [22].
2. It is generally believed that there is no aberration for light reflected from mirrors moving transversely. To quote e.g. Sommerfeld [23]: ”Thus, for a mirror moving tangentially to its surface the law of reflection which holds for the stationary mirror is preserved.” Similar statements can also be found in other textbooks. To quote e.g. Ugarov [24]: ” Hence, when the mirror moves parallel to itself the frequency of incident light is equal to that of reflected light, and the angle of incidence is equal to the angle of reflection.”
3. There is another interpretation that we can give to the scattering process. We note the very close correspondence between angular and frequency response of moving diffraction gratings and the Raman scattering [25]. Since the theory of Bragg diffraction applies to light scattering by sound waves in liquids and solids, it is not surprising that we can obtain our results from the quantum theory of light scattering by phonon. This scattering process is also known as Brillouin scattering. The quantum theory leads to equation $\vec{k}_{s}=\vec{k}_{i}\pm\vec{K}_{\perp}$ by the requirement that the momentum is conserved between interacting photon and phonon. The process of Brillouin scattering is a special case of Raman scattering. The momentum balance for scattering involving the emission or absorption of phonon leads to the Bragg condition. The requirement of conservation of energy leads to the equation $\omega_{s}=\omega_{i}\pm\omega_{phon}$ with $\omega_{s}$, $\omega_{i}$, and $\omega_{phon}$ being the radian frequencies, of the scattered photon, the incident photon, and the phonon. The frequency shift that occurs for light scattering from sound waves comes about as a result of the Doppler effect. Quantum mechanically, it is a consequence of the conservation of energy between the participating particles. We come to the conclusion that, according to the hidden choice of absolute time coordinatization in non-relativistic quantum mechanics, the lab observer actually sees the radiation after reflection as a result of a Galilean boost rather than a Lorentz boost.
4. There is a common misconception that the radiation wavefront orientation has objective meaning. To quote Norton [26]: ”One might try to escape the problem by supposing that the direction of propagation is not always given by the normal to the wavefront. We might identify the direction of propagation with the direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation. Whatever may be the merits of such proposals, they are unavailable to some trying to implement a principle of relativity. If the direction of propagation of a plane wave is normal to the wavefronts in one inertial frame then that must be true in all inertial frame.” This incorrect statement is a straightforward consequence of the generally accepted way of looking at special relativity of most physicists. Accepting the postulate on the constancy of the speed of light one also automatically assumes Lorentz coordinates. According to such limiting understanding of the theory of relativity, it is assumed that only Lorentz coordinatization must be used to map the coordinates of events.
5. Note that spatiotemporal coupling is discussed in literature usually in relation with ultrashort laser pulse propagation through a grating monochromator. Ultrashort laser pulses are usually represented as a products of electric field factors separately dependent on space and time. However, when the manipulation of ultrashort laser pulses requires the propagation through a grating monochromator, such assumption fails. In this situation one has to consider in addition (to phase fronts), planes of constant intensity, that is pulse fronts. A pulse-front tilt can be present in the beam due to propagation through an optical setup incorporating dispersive optical element. In the grating monochromator case, the different spectral components of the out-coming pulse travel in different directions. The electric field of a pulse including angular dispersion can be expressed in the Fourier domain $[k_{x},\omega]$ as $\bar{E}(k_{x}-p\omega,\omega)$, while the inverse Fourier transform from the $[k_{x},\omega]$ to the space-time domain $[x,t]$ can be expressed as $E(x,t+px)$, which is the electric field of pulse with a pulse-front tilt. The tilt angle $\theta_{tilt}$ is given by $\tan\theta_{tilt}=cp$. More specifically $p=dk_{x}/d\omega=kd\theta_{D}/d\omega=\lambda/(c\theta_{D}d)$, where $\lambda=2\pi c/\omega$, $\theta_{D}$ is the diffracted angle, and $d$ is the groove spacing. The diffracted angle $\theta_{D}$ is a function of frequency, according to the well-known plane grating equation. Assuming diffraction into the first order, one has $\lambda=(\cos\theta_{i}-\cos\theta_{D})d$, where $\theta_{i}$ is the incident angle. By differentiating this equation one obtains $d\theta_{D}/d\lambda=1/(\theta_{D}d)$, where we assume for simplisity grazing incidence geometry, $\theta_{i}\ll 1$ and $\theta_{D}\ll 1$. The physical meaning of the former equation is that different spectral components of the out-coming pulse travel in different directions. Therefore one concludes that the pulse-front tilt is invariably accompanied by angular dispersion. It follows that any device like a grating monochromator, that produces an angular dispersion, also introduces significant pulse-front tilt. In our case of interest we have deal with light diffracted by a (transversely) moving set of gratings and the Doppler effect is responsible for frequency dispersion $d\omega/dk_{x}=v$. Thus the spatitemporal coupling due to the light beam transmission through a hole in a moving screen and the usual pulse-front tilt distortion are quite different.
6. To quote Brillouin [27]: ” Fig. 5 explain the situation assuming a simplified device consisting of parallel plate moving with uniform velocity $v$ in the horizontal direction. Monochromatic light is falling normally on the plate and generates an oblique ray.” This oblique effect is demonstrated in our Fig. 10.
5 Aberration of Light Phenomenon. Noninertial Frame of Reference
5.1 Introductory Remarks
In this section we reexamine the issue of light transmission in non-inertial frames with particular reference to the aberration of light phenomena. We derive the aberration for a pulse of light traveling in the accelerated systems using the Langevin metric in general relativity (1). It comes out naturally if one writes the equation of the time transfer, from the inertial frame to the accelerated frame, in a generally covariant context. We only wish to emphasize here the following point. From a mathematical standpoint, there is no difference between calculations in the framework of the general theory of relativity and of the special theory of relativity in the absence of space-time curvature.
The aberration of light problem is solved with discovery of the essential asymmetry between the non-inertial and the inertial observers. Actually, in resent years it seems to be almost normally accepted in scientific community that the ”theory of relativity” is just a name, not to be taken literally. One can conclude that not all is relative in relativity, because this theory also contains some features which are absolute.
Has all our talk about asymmetry violated the relativity principle? At a first glance it might seem so, since the relativity principle is often interpreted as implying perfect symmetry among moving frames. The principle of relativity denied the possibility for an observer partaking in a uniform motion relative to an inertial frame of discovering by any measurement such a motion, of course, that one does not look outside. The arguments concerning the relativity of motion in our case of interest cannot be applied, since the inertial and non-inertial reference systems are not equitable.
A typical resolution of the asymmetry paradox identifies acceleration as the agency of asymmetry. It should be note that the duration of the acceleration period has a negligible effect on the anisotropy in the accelerated frame, so the acceleration need not be considered explicitly in working the problem. Nevertheless, the acceleration completely determines the problem. It is worth remarking that the absent of a dynamical explanation for the asymmetry in special relativity has disturbed some physicists. A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer. An idea of treating the accelerated frame using the equivalence principle is useful from a pedagogical point of view. See section 9.1 for more ideas along this line.
However, there remains an intriguing puzzle to solve: how can the (e.g. earth-based and sun-based) observers tell which observer took the acceleration? The surprising fact is, the determination of which observer took the acceleration can be made only by observation of the ”fixed stars”. The acceleration is in principle defined in terms of motion relative to the fixed stars, and they must be consulted in order to determine whether an acceleration occurred. Thus when we state that the earth-based observer undergoes an acceleration, and the sun-based observer does not, there is a hidden assumption concerning the distribution of mass in the universe. The implicit ”absolute” acceleration means acceleration relative to the fixed stars.
5.2 Absolute Time Coordinatization in the Accelerated Systems
We will consider the problem of aberration of light in the accelerated systems on the basis of the theory of special relativity. Let us demonstrate that for explanation of the optical effects in the rotating frame of reference one does not need neither modify the special theory of relativity, nor apply the general theory of relativity. It is only necessary to strictly follow the special theory of relativity. In order to write electrodynamics laws in the accelerated frame we have to go just one step further, and define the metric of the non-inertial frame. One must take into account that metric tensor must be a continuous quantity. The first mathematical idea is the smooth tailoring of the metric tensor. This problem in special relativity can be adequately treated only by an approach which uses the absolute time coordinatization. We describe the mathematical expansion of special theory of relativity onto accelerations in the Chapter 9. Also this can be done directly, it is some times possible to save time by getting the answer with some physical arguments. For example, in our case of interest one can take advantage of the following considerations, which are in agreement with rigorous mathematical derivation.
Suppose that the initial inertial system $S$ is at rest with respect to the fixed stars and a system $S_{n}$ (and an observer with his measuring instruments) in the frame $S$ is accelerated from the rest with respect to the fixed stars up to velocity $v$ along the $x$-axis. In accelerated systems, only the theory maintaining an absolute simultaneity is logically consistent with the natural behavior of clocks. The method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the system $S_{n}$ was at rest. It is well known that during the motion with acceleration (with respect to the fixed stars) the procedure of Einstein’s clock synchronization cannot be performed and the interval in the accelerated reference system $S_{n}$ will, by the moment when the system $S_{n}$ starts moving with constant velocity, have the non-diagonal form.
Absolute simultaneity can be introduced in special relativity without affecting neither the logical structure, no the (convention-independent) predictions of the theory. We begin with the metric as the true measure of space-time intervals for an non-inertial observer $S_{n}$ with coordinates $(t_{n},x_{n},y_{n},z_{n})$. We transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$, using the inverse Galilean transformation: we substitute $x=x_{n}+vt$, while leaving time unchanged $t_{n}=t$ into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain (1)
$ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}~{}.$ (15)
This metric describes the electrodynamics of the accelerated light source with viewpoint of of the accelerated observer measurements as viewing this of the inertial observer. Inspecting Eq. (15) we can find the components of the metric tensor $g_{\mu\nu}$ in the coordinate system of $S_{n}$. We obtain $g_{00}=1-v^{2}/c^{2}$, $g_{01}=-v/c$, $g_{11}=-1$. Note that the metric in Eq. (15) is not diagonal, since, $g_{01}\neq 0$, and this implies that time is not orthogonal to space. This Langevin metric is found by matching the accelerated frame and (initial) inertial frame metric tensors. Taking into account the Langevin metric we obtain that the length of the physical rod $dx$ and the physical (proper) time $dt$ in the inertial frame will coincide with the coordinate distance and the coordinate time in the accelerated frame, i.e. $dx_{n}=dx$, $dt_{n}=dt$. (see Chapter 9 for more details).
The velocity of light emitted by a source at rest in the coordinate system $(t,x)$ for $S$ is $c$. In that case the Minkowski metric Eq.(1) associated with inertial frame $S$ predict a symmetry in the one-way speed of light. In the coordinate system $(t_{n},x_{n})$, however, the speed of light emitted by the accelerated source (i.e. source which is at rest with respect to the accelerated frame $S_{n}$) cannot be equal $c$ anymore because $(t_{n},x_{n})$ is related to $(t,x)$ via a Galilean transformation. This is readily verified if one recalls that the velocity of light in the reference system $S$ is equal to $c$. If $ds$ is the infinitesimal displacement along the world line of a ray of light, then $ds^{2}=0$ and we obtain $c^{2}=(dx/dt)^{2}$. In the accelerated reference system, since $x_{n}=x-vt$ and $t=t_{n}$, this expression takes the form $c^{2}=(dx_{n}/dt_{n}+v)^{2}$, which can be seen by setting $ds^{2}=0$ in Eq. (15). This means that in the accelerated reference system of coordinates $(ct_{n},x_{n})$ the velocity of light parallel to the x-axis, is $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction as stated above. The reason why it is different from the electrodynamics constant $c$ is due to the fact that the clocks are synchronized following the absolute time convention, which is fixed because $(t_{n},x_{n})$ is related to $(t,x)$ via a Galilean transformation.
5.3 The Asymmetry between the Inertial and Accelerated Frames
We discovery of the essential asymmetry between the inertial and accelerated frames, namely, the Maxwell’s equations are not applicable from the viewpoint of an observer at rest with respect to accelerated frame $S_{n}$. In fact, the metric Eq.(15) associated with accelerated reference frame $S_{n}$ predicts an asymmetry in the one-way speed of light in the relative velocity direction. Accelerations (with respect to the fixed stars) have an effect on the propagation of light. On accelerated system $S_{n}$, the velocity of light emitted by a source at rest must be added to (or subtracted from) the speed due to acceleration and the velocity of light is different in opposite directions. In contrast, the Maxwell’s equations continue to hold from the viewpoint of an observer at rest with respect to inertial frame $S$. On inertial system $S$, the velocity of light emitted by a source at rest is $c$.
The following important detail of such acceleration frame description can hardly be emphasize enough. Metric applies to physical laws, not to physical facts. Indeed, it is always possible to chose such variables, in which metric of the accelerated source will be diagonal. We interpret the Langevin metric to mean that the law of electrodynamics is expressed by anisotropic field equations in the accelerated frame. By a physical facts in this context we mean the aberration of light radiated by a single ”plane wave” emitter in the accelerated frame. The electrodynamics equation needs to be integrated with initial condition for the radiation wavefront. After the boost we can see that acceleration has no effect on the wavefront orientation. In fact, the variables $(t,x,y,z)$ can be expressed in terms of the variables $(t_{n},x_{n},y_{n},z_{n})$ by means of Galilean transformation $x_{n}=x-vt$, $t_{n}=t$, so that the wavefront of the emitted light beam is perpendicular to the vertical direction $z_{n}$ after the acceleration, Fig. 14. We shall discuss this subject further in the Section 5.5.
It would be well to begin with bird’s view of some main results. Suppose that an observer in the accelerated frame $S_{n}$ performs an aberration measurement. How shall we describe the aberration of light from the ”plane wave” emitter which is at rest in the inertial frame $S$? In order to predict the result of the aberration measurement the accelerated observer should use the non-diagonal metric Eq.(15).
Figure 14: Aberration of light in an accelerated frame of reference $S_{n}$. Radiation wavefront orientation and the anisotropy of speed of light presented in the absolute time coordinatization ($t_{n}=t$) for the screen. The crossed term in metric Eq.(15) generates anisotropy in the accelerated frame that is responsible for the change of radiation direction (aberration).
Figure 15: Aberration of light in an accelerated frame of reference $S_{n}$. Screen is at rest in the accelerated frame. Radiation wavefront orientation and the anisotropy of speed of light presented in the absolute time coordinatization ($t_{n}=t$) for the screen. The aberration of the transmitted light is considered independent of the source speed and to have just a local origin exclusively based on the observer acceleration relative to the fixed stars.
According to the asymmetry between the inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of light. Namely, if the opaque screen with hole was at rest relative to the fixed stars and the screen started from rest to motion, then the apparent angular position of the ”plane-wave” emitter seen in the accelerated frame through the aperture would jump by angle $-v/c$. That is the transmission through the hole in the opaque screen in the frame $S_{n}$ appears as shown in Fig 14. It is important to note that the source of asymmetry is not the relative acceleration of two frames; rather, it is the difference in the accelerational history of each frame separately with respect to the fixed stars.
Above (see Chapter 4) we considered the emitter-screen problem in the frame of reference $S$ which is at rest (i.e. without accelerational history) with respect to the fixed stars. It is found that there is no aberration proceeding from the emitter, independently of their motion in the case of the small aperture hole. It is important to emphasize that the aberration of the light beam transmitted through the small aperture hole also does not dependent on the emitter motion in the accelerated system $S_{n}$, Fig. 15. The point is that the crossed term in metric Eq.(15), which generates aberration in this particular case, depends only on the velocity change of the accelerated frame relative to the initial inertial frame.
5.4 Explanation of the Aberration on the Basis of Electrodynamics
It is well known that the electrodynamics theory meets all requirements of the theory of relativity and therefore must accurately describe the properties of such a relativistic object as light. For example, we are able to demonstrate that velocity of light in the accelerated frame can be derived independently of a technique involving metric Eq. (15). In the inertial frame, fields are expressed as a function of the independent variables $x,y,z$, and $t$. According to the special relativity, the Maxwell’s equations always valid in the Lorentz reference frame. The electric field $\vec{E}$ of an electromagnetic wave radiated by the emitter at rest in the inertial frame satisfies the equation $\Box^{2}\vec{E}=\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. However, the variables $x,y,z,t$ can be expressed in terms of the variables $x_{n},y_{n},z_{n},t_{n}$ by means of a Galilean transformation, so that fields can be written in terms of $x_{n},y_{n},z_{n},t_{n}$. Hence the wave equation transforms into
$\Box^{2}\vec{E}_{n}=\left(1-\frac{v^{2}}{c^{2}}\right)\frac{\partial^{2}\vec{E}_{n}}{\partial x_{n}^{2}}+2\left(\frac{v}{c}\right)\frac{\partial^{2}\vec{E}_{n}}{\partial t_{n}\partial x_{n}}+\frac{\partial^{2}\vec{E}_{n}}{\partial y_{n}^{2}}+\frac{\partial^{2}\vec{E}_{n}}{\partial z_{n}^{2}}-\frac{1}{c^{2}}\frac{\partial^{2}\vec{E}_{n}}{\partial t_{n}^{2}}=0~{},$ (16)
where coordinates and time are transformed according to a Galilean transformation. The solution of this equation $F[x_{n}-(c-v)t_{n}]+G[x_{n}+(-c-v)t_{n}]$ is the sum of two arbitrary functions, one of argument $x-(c-v)t$ and the other of argument $x+(-c-v)t$. The first term represents a wave traveling forward in the positive $x_{n}$ direction, and the second term a wave traveling backwards in the negative $x_{n}$ direction. Acceleration has an effect on the propagation of light. In fact, the coordinate velocity of light parallel to the $x$-axis is given by $dx/dt=c-v$ in the positive direction, and $dx/dt=-c-v$ in the negative direction.
We now consider the transmission through the hole in the opaque screen in the accelerated frame $S_{n}$, Fig. 15. We describe the aberration of light based on the observations made by an observer in the same accelerated frame as the screen. It is assumed that the detector for the direction of the radiation is an energy propagation detector and the size of the detector aperture is sufficiently large compared with the radiation beam size. In other words, it is assumed that there is no physical influence of the detector (e.g. aperture) on the measurement.
We have already discussed the explanation of the aberration on the basis of electrodynamics in the previous chapter. Our earlier discussion is really about as far as anyone would normally need to go with the subject, but we are going to do it again. One reason is that one should know how to deal with what happens to waves in the accelerated frame from the viewpoint electrodynamics equations. First, we remind about the convenient way of describing diffraction by opaque screen we used in the previous chapter. The screen containing a hole is a kind of diffraction grating which breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an transmittance can be resolved. Let us assume that the transmittance of the grating varies according to the law $T=g(K_{\perp})\cos{({K}_{\perp}x)}$.
[…]
In our example, the plane waves with different wavenumber vectors propagate out from the screen with different frequencies. Then equation $\Delta\omega/\Delta k_{x}=-v$ holds for each transmitted waves independently on the sign and the magnitude of the transmitted angle. The last equations state that transmitted light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=-v$, Fig. 15.
We would now like to discuss an apparent paradox. We obtained the electrodynamics equations in accelerated frame using the Galilean transformation of the Maxwell’s equations. Intuition would seem to tell us that everything is at rest in the accelerated frame. But special relativity says that there is a derivative $\partial/\partial t_{n}=v\partial/\partial x_{n}$ because there is $\partial/\partial x_{n}$ that is not zero. Here we have a new kind of situation which is quite different from inertial frame electrodynamics. We will discuss this subject further in the Chapter 10.
5.5 A Resynchronization of the Accelerated Clocks
It should be noted, however, that there is another satisfactory way of explaining the effect of aberration of light in the accelerated frame $S_{n}$. The explanation consists in using a clock re-synchronization procedure. Well known that in their original form Maxwell’s equations are valid in the inertial frames. But Maxwell’s equations can be written down only if the Lorentz coordinates has already been specified.
When the system $S_{n}$ starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. The Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all direction. Using such synchronization procedure we actually select a Lorentz coordinate system for the screen. In this synchronization, we describe the transmission through the aperture using usual Maxwell’s equations. The interval in the accelerated reference system $S_{n}$ will have the diagonal form Eq.(1) for the transmitted light beam.
Figure 16: Aberration of light in the accelerated frame of reference $S_{n}$. The plane wavefront rotates in the accelerated frame after the accelerated clock re-synchronization. The Maxwell’s equations can now be used to describe the transmitted (through the aperture) light beam. According to the Maxwell’s electrodynamics, the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light.
The time $t^{\prime}_{n}$ under the Einstein’s synchronization in the $S_{n}$ frame is readily obtained by introducing the offset factor $x_{n}v/c^{2}$ and substituting $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first order approximation. This time shift has the effect of rotation the plane of simultaneity (that is emitter radiation wavefront) on the angle $-v/c$ in the first order approximation. As a consequence of this, the plane wavefront rotates in the accelerated frame after re-synchronization, Fig. 16. The new time coordinate in the accelerated frame is interpreted by saying that Maxwell’s equations are applicable to the light transmission (through the aperture) description. Then, the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light: the two approaches give the same result. The choice between these two different clock synchronizations is a matter of pragmatics. By changing the (four-dimensional) coordinate system, one cannot obtain a physics in which new physical phenomena appear. But we can obtain a more consistent description of these phenomena.
We had to restrict ourselves to finding the diagonalization procedure only for nonrelativistic velocities. The method can be easily generalized for an arbitrary parameter $v/c$. A full diagonalization can be obtained using transformation $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$. Physical time $t^{\prime}_{n}$ determines the flow of time in a physical process in the accelerated frame $S_{n}$. The $dx^{\prime}_{n}$ is the physical distance between to points of 3-space.
5.6 Composition of Motions
Until now we have considered only the case in which motion is nonrelativistic in speed. Now we shall be more general and study the case where the motion is at an arbitrary velocity. We already know from our discussion in this chapter that metric tensor must be a continuous quantity. The Langevin metric is found by matching the accelerated frame and inertial frame metric tensors. For smooth tailoring of the metric tensor, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the inverse Galilean boost. In this case the metric of inertial frame would pass into the Langevin metric of the accelerated frame. Let us consider the transformation of the direction of light propagation. The light (group) velocity transforms as a particle velocity. In this case we use the Galileo addition theorem of velocities for transforming the speed of propagation of light to the accelerated frame. Let us assume that the accelerated frame moves with velocity $v$, then the angle of aberration $\theta_{a}$ is such that   $\tan\theta_{a}=-v/c$. If $v/c\ll 1$ the angle of aberration is small, being approximated by $\theta_{a}=-v/c$.
It is always possible to chose such variables, in which metric of the accelerated source will be diagonal. In the Lorentz coordinatization, light propagates with constant velocity $c$ independently of the direction of propagation, and of the velocity of its source. If we look at transformation $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$, we see that velocity along the $x_{n}$-axis is the same in the Lorentz coordinatization too: $dx_{n}^{\prime}/dt_{n}^{\prime}=(dx_{n}/dt_{n})/[1-v^{2}/c^{2}-v(dx_{n}/dt_{n})]$. In fact, substituting $dx_{n}/dt_{n}=-v$ into this equation gives $dx_{n}^{\prime}/dt_{n}^{\prime}=-v$. Group velocity of light beam along the fixed stars motion is still given by $-v$, but when we use the new velocity of light this becomes $\sin\theta_{a}=-v/c$. This is modification of the aberration of light law in the Lorentz coordinatization. Suppose that accelerated observer performs a measurement of the light aberration. Since in the Lorentz coordinatization the time and distance have direct physical meaning, the accelerated observer finds that the angle of aberration is such that $\sin\theta_{a}=-v/c$.
5.7 Discussion
Now let us return to metric diagonalization and try to get better understanding of the relationship between the $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$ coordinate systems. We have already noted that the old coordinates $(x_{n},y_{n},z_{n},t_{n})$ are found by matching the accelerated frame and inertial frame metrics; what does mismatching of the coordinates $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$ mean, in terms of measurements made by accelerated observer? Here the metric in both reference frames will be diagonal, and, according to textbooks, the coordinates in this frames should be related by the Lorentz transformation. At first site the problem is complete symmetrical and both frames are equivalent. Above we demonstrated that there is a real difference between the accelerated inertial frame and the inertial frame without accelerational history. From a mathematical viewpoint , we suggest the following explanation. After diagonalization, the metric tensor in the accelerated frame must abruptly change from the value $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the value $g_{00}=1,g_{01}=0$. We must conclude that when we are dealing with (Lorentz ) transformation and this transformation does not preserve continuity of the metric tensor we have no symmetry between inertial frames.
Where does the asymmetry come from? The electrodynamics equation needs to be integrated with initial conditions. According to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the radiation wavefront. In the inertial frame, the wavefront of the emitted light beam is perpendicular to the vertical direction $z$. On the other hand, we know that when we are dealing with Lorentz coordinatization we have effect on the wavefront orientation in the accelerated frame $S_{n}$. According to relativistic kinematics, the extra phase chirp $d\phi/dx^{\prime}_{n}=k_{x}=-v\omega_{i}/c^{2}$ is introduced as a consequence of this, the plane wavefront rotates after the acceleration. Then, the radiated light beam is propagated at the angle $-v/c$ with respect to the $z_{n}$-axis, yielding the phenomenon of the aberration of light in the accelerated frame $S_{n}$.
It is generally accepted that this is purely kinematic effect, involving no forces. Many people will, we think, find this disturbing. They would like to think that since in the Lorentz coordinatization the time and distance have direct physical meaning, there should be some physical (dynamical) reason for this wavefront rotation. We discuss the key of this paradox in great detail in the Section 9.1. It is a dynamical line of arguments that explains this paradoxical situation with the asymmetry between inertial and accelerated reference frames. Without proof, we may state the results. A resolution of the asymmetry paradox identifies inertial (pseudo-gravitational) force within the system $S_{n}$ as the agency of asymmetry. The principle of equivalence can be applied to solve non-inertial kinematic problems with dynamical methods. Wavefront rotation associated with the transformation from the inertial frame to the accelerated (with respect to the fixed stars) frame may be regarded as a result of the action of pseudo-gravitational potential gradient during the acceleration process. The rate of a clock depends on the pseudo-gravitational potential at the place where the clock is situated. The time-offset relation can be interpreted as the accomulated time difference between spatially separated clocks. Suppose (for simplicity) that there is a constant acceleration $g$ along $x_{n}$-axis during the time $T=v/g$. The pseudo-gravitational acceleration is simply equal to the gradient of scalar potential: $g=-\partial\phi/\partial x_{n}$. The clock at higher gravitational potential (placed along the direction of acceleration) runs faster. For the pseudo-gravitational potential difference between two points along the $x_{n}$-axis we get $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$. When the system $S$ starts moving with constant velocity the gradient of potential in the system $S_{n}$ is zero. For accumulated time difference between two spatially separated clocks we have $t_{n}(1)-t_{n}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$. This time shift has the effect of rotation the plane of simultaneity on the angle $-v/c$.
5.8 A Single Moving Emitter in a Noninertial Frame of Reference
Above in the sections 4.3 - 4.4 we considered a single moving ”plane wave” emitter in the initial inertial frame of reference. Let us now analyze the aberration of light radiated by a single ”plane wave” emitter moving in the accelerated system. Before we go on to analyze an observations of a non-inertial observer, we should make one more remark about observations of an inertial observer. In the Chapter 4 we already emphasized that in the description of the aberration of light in the initial inertial frame of reference there are two choices of (four-dimensional) coordinatizations useful to consider:
(a) Non-standard (absolute time) coordinatization for a moving source
(b) Standard Lorentz coordinatization for a moving source
When the light source in an inertial frame is accelerated from rest up to velocity $v$ along the $x$ axis, the simplest ( absolute time) method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the light source was at rest, i.e. we still enforce the clock transport synchronization ( or Einstein synchronization which is defined in terms of light signals emitted by the light source at rest). This choice is usually the most convenient one from the viewpoint of connection to laboratory reality.
In fact, conventional particle tracking treats the space-time continuum in a non-relativistic format, as a (3 + 1) manifold. In other words, in this standard approach, introducing as only modification to the classical case the relativistic mass, time is absolute and we have no mixture of positions and time.
Let us, then, discuss (according to the Chapter 3) how to assign Lorentz coordinates in the case when a light source in an inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis. In order to assign a Lorentz coordinate system in the inertial frame after the Galilean boost, one needs to perform the distant clock resynchronization $t^{\prime}=t-xv/c^{2}$. This new space-time coordinates in the lab frame are interpreted, mathematically, by saying that the metric is now diagonal and the speed of light from the moving source is isotropic and equal to $c$.
Now let us return to observations of a non-inertial observer. In the description of the aberration of light radiated by a single moving emitter in an accelerated frame of reference there are also two choices of coordinatizations:
(a) Non-standard (absolute time) coordinatization for an accelerated source
(b) Standard Lorentz coordinatization for an accelerated source
(a) When the system $S_{n}$ in the initial inertial lab frame $S$ is accelerated from the rest up to velocity $v$ along the $x$-axis, the simplest ( absolute time) method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the system $S_{n}$ was at rest. Acceleration with respect to the fixed stars have an effect on the propagation of light. The velocity of light from the source which is at rest in the accelerated frame $S_{n}$ will, by the moment when the system $S_{n}$ starts moving with constant velocity, have anisotropy along the $x_{n}$ axis (see Eq. (15)).
Figure 17: Aberration of light in an accelerated frame of reference $S_{n}$. Emitter is at rest with respect to the fixed stars. Wave fronts orientation and the group velocity presented in the Einstein synchronization. The Einstein synchronization is defined in terms of light signals emitted by a source at rest (in the accelerated frame) assuming that light propagates with the same velocity $c$ in all direction. The aberration increment $\theta_{a}$ is connected with the physical parameters by the relation: $\theta_{a}=-2v/c$, where $v$ is the velocity of the fixed stars in the accelerated frame $S_{n}$.
(b) When the system $S_{n}$ starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. The Einstein synchronization is defined in terms of light signals emitted by a source at rest (in the accelerated frame) assuming that light propagates with the same velocity $c$ in all direction. Using such (Einstein time) synchronization procedure we actually describe the light from the source at rest (in the accelerated frame) using usual Maxwell’s equations.
Above we considered the single emitter problem in the initial inertial frame of reference $S$. It is found that there is a deviation of the energy transport for the light radiated by the transversely moving emitter, which is nothing else but well-known (from textbooks) result: there is the aberration of light from the transversely moving emitter in the inertial frame of reference (Fig. 1).
We now consider the case when the emitter is at rest in the initial inertial frame $S$ (i.e. at rest with respect to the fixed stars) and the observer, which is at rest with respect to the accelerated frame of reference $S_{n}$ performs the direction of the energy transport measurement. It is important to emphasize that the aberration of light radiated by the single emitter is also dependent of the emitter motion in the accelerated system $S_{n}$. According to the asymmetry between the inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of light. Namely, if the emitter is at rest relative to the fixed stars and the observer started from rest to motion relative to the fixed stars, then the apparent angular position of the ”plane-wave” emitter seen in the accelerated frame would jump by angle $-2v/c$. That is the aberration of light in the frame $S_{n}$ appears as shown in Fig 17.
We are going to discuss the asymmetry between inertial and accelerated frames. Now let us assume the absolute time coordinatization. The Langevin metric, together with initial condition, describes the aberration of light effect for the emitter at rest in the accelerated frame. The essential asymmetry is rooted in the electrodynamics equations. The laws of special relativity say that we can describe radiation of the moving with velocity $-v$ in the accelerated frame emitter in terms of the accelerated observer if we describe the electrodynamics by using metric
$ds^{2}=c^{2}(1-4v^{2}/c^{2})dt_{n}^{2}-4vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}~{}.$ (17)
This metric describes the electrodynamics of the moving light source with viewpoint of the accelerated observer measurements. From the formula itself it is easy to see that the result of measurement of the electromagnetic field configuration in the moving system in terms of coordinate $x_{n}$ and time $t_{n}$ of the measured device at rest in the accelerated frame is the same as the measured result of this configuration at point $x_{n}+vt_{n}$ and time $t_{n}=0$. In the inertial frame we have a similar situation. In the Chapter 3 we described the electrodynamics of the moving emitter by using metric Eq. (11). This is the same way we used in the accelerated frame.
So we begin to understand the basic machinery of the asymmetry between (initial) inertial and accelerated frames in the absolute time coordinatization: inertial observer describes the electrodynamics of the light source at rest using Minkowski metric. We substitute $x=x_{n}+vt,t_{n}=t$ into Minkowski metric to obtain Langevin metric Eq. 15 in the comoving frame. We may ask ourselves, what is the metric of the inertial observer with viewpoint of the accelerated observer? We transform coordinates $(t_{n},x_{n},y_{n},z_{n})$ that would be coordinates of an accelerated observer $S_{n}$ using the inverse Galilean transformation: we substitute $x_{n}=x-vt$, while leaving time unchanged $t_{n}=t$ into the Langevin metric Eq. 15 to obtain the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ in the initial inertial frame, as it must be. This metric describes the electrodynamics of the light source at rest in the inertial frame with viewpoint of the inertial observer measurements as viewing this of the accelerated observer. In particular this demonstrates the time dilation effect of a physical clock at rest in the accelerated frame. The slowly down of the accelerated clock does not depends on the choice of reference system, in which this effect is measured (see Chapter 9 for more details).
There is still another way of solving the problem. The explanation consists in using the Langevin metric diagonalization procedure. The new coordinatization in the accelerated frame is interpreted by saying that the accelerated observer should obtain diagonal metric $ds^{2}=c^{2}(dt^{\prime}_{n})^{2}-(dx_{n}^{\prime})^{2}-(dy_{n}^{\prime})^{2}-(dz_{n}^{\prime})^{2}$ for the emitter at rest in the accelerated frame and the metric $ds^{2}=c^{2}(1-v^{2}/c^{2})(dt^{\prime}_{n})^{2}-2vdx_{n}^{\prime}dt_{n}^{\prime}-(dx^{\prime}_{n})^{2}-(dy^{\prime}_{n})^{2}-(dz^{\prime}_{n})^{2}$ for the moving emitter. Now we see that there is a symmetry (reciprocity) in metrics (and consequently in electrodynamics equations) between the accelerated and inertial frames. This is consequence of the pseudo-Euclidean geometry of space-time. In fact, the difference between this metric and metric Eq. (11) is only the sign of velocity. Clearly, after the diagonalization, the asymmetry is rooted in the initial conditions. The plane radiation wavefront of moving emitter (independently of its velocity) rotates in the accelerated frame on the angle $-v/c$ after metric diagonalization. Of course in order to predict the result of aberration measurement the analysis can be done using directly the metric of the moving emitter in the absolute time coordinatization, Eq. (17). According to electrodynamics (see Section 5.4), the extra group velocity $-2v$ is introduced as a consequence of the crossed term $-4vdxdt$, the radiated beam of the moving emitter is propagated at the angle $-2v/c$.
Let us now consider the most general case when the system $S_{n}$ in the frame $S$ accelerated from the rest up to velocity $v$ along the $x$ axis and emitter in the accelerated frame $S_{n}$ is accelerated from rest up to velocity $u$ along the $x_{n}$ axis. Suppose that an observer in the accelerated frame $S_{n}$ performs an aberration measurement. We can describe radiation of the moving with velocity $u$ in the accelerated frame emitter in therms of the accelerated observer if we describe the electrodynamics by using metric $ds^{2}=c^{2}[1-(v-u)^{2}/c^{2}]dt_{n}^{2}-2(v-u)dx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. At close look at the physics of this subject shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-v/c+u/c$. The point is that the crossed term in metric Eq.(15), which generates the anisotropy in the accelerated frame, depends only on the velocity $v$ of the accelerated frame relative to the initial inertial frame.
Figure 18: Aberration of light in an accelerated frame of reference $S_{n}$. Emitter is at rest with respect to the fixed stars. The aberration of light might be calculated from classical theory, in which the light beam may be treated, say, as a corpuscular (raindrops) beam. On the basis of classical theory one gets a light aberration angle that is only one-half as big as that predicted by special relativity.
Figure 19: The deflection of starlight by the field of the sun. The light deflection might be calculated from classical theory, in which the light beam may be treated, say, as a comet approaches with the velocity of light. On the basis of classical theory one gets a deflection of light that is only one-half as big as that predicted by general relativity.
5.9 A Useful Analogy
We have presented theoretical evidence of an aberration increment $2v/c$ concerning the single emitter problem in a non-inertial frame of reference. First, we notice that there is an extra factor 2. On the basis of classical theory one gets aberration increment that is only one-half as big as that predicted by special relativity, Fig. 18. According to classical approach, there is no principle difference between the aberration of light and the aberration of raindrops.
Now let us consider a completely different situation which, in the classical case, gives deflection of light that is only one-half as big as that predicted by special relativity. There is a methodological analogy between the aberration of light effect and the deflection of light in general relativity. A ray of light which, coming from fixed star, passes close by the sun will thus be attracted to it and will describe a somewhat concave orbit with respect to the sun. This deflection might be calculated from Newton’s theory, in which the ray of light may be treated, say, as an asteroid which approaches with the velocity of light. We then get a formula similar to that of Einstein, but giving only half the value of the deflection, Fig. 19. Errors inherent in the classical method applied to the computation of the deflection of light are due to the use of the velocity of light simply as a classical velocity. According to classical approach, there is no principal difference between the deflection of light and the deflection of asteroid. A methodological analogy with the calculation of the aberration increment from classical (raindrops) theory emerges. Does this analogy have any physical meaning? To see whether it does, we should compare the motion of relativistic particles in the gravitation field with the aberration of particle effect. We shall do this in the Section 7.5.
Figure 20: Aberration of light in an accelerated frame of reference. Point source is at rest in the accelerated frame. Wavefront orientation presented in the absolute time coordinatization for the screen.
Figure 21: Aberration of light in an accelerated frame of reference. Point source is at rest with respect to the fixed stars.
5.10 A Point Source in a Non-inertial Frame of Reference
Now let us return to observations of a non-inertial observer. The peculiarity of point-like sources is that radiation emitted at one instant form a sphere around the source and the measuring instrument always influences the measured radiation, Fig. 20-Fig. 21 . The aberration of point source is considered independent of the source speed and to have just a local origin exclusively based on the observer (with his measuring instruments) speed relative to the fixed stars. The crossed term in metric Eq.(15) generates anisotropy in the accelerated frame that is responsible for the change of transmitted radiation direction (2). The point source follows the same pattern as fixed stars under the same aberration angle i.e. its apparent position changes with an angular displacement common to fixed stars.
In the framework of the conventional theory of the aberration of light, there is an outstanding puzzle concerning the stellar aberration. There are double star systems the components of which change their velocity on a time scale ranging from days to years. The components of such binary systems at some times can have velocities relative to the earth very different from one another; nevertheless it is well known that these components exhibits always the same aberration angle. Rotating binary systems follow the same pattern as all fixed stars and are observed within a period of a year under the same universal aberration angle, i.e. their apparent position changes with an annual period common to all distant stars. This argument suggests that results of the astronomical observations confirm our prediction for aberration of light from a point source in an non-inertial frame. In the next chapter we discuss this experimental test in more detail.
In regard to light aberration one should differentiate between that from the point source and that from the laser source. Let us analyze the aberration of light transmitted from the laser resonator which is at rest in the accelerated frame. A close examination of of all experiments inside the accelerated frame shows that all optical phenomena in an optical resonator appeared to be independent of the acceleration relative to the fixed stars. It should be note that there is no influence of the difference between Langevin and Minkowski metrics on the parameter of a laser source. It is not hard to understand this result. It is clear that the crossed term in the Langevin metric, which generates aberration, cancels during the (round-trip) evolution of the radiation in the laser resonator. We will analyze the aberration of light transmitted from the laser resonator in great detail in the Section 8.6
5.11 Tilted Incoming Plane Wave
We go back now to the situation shown in Fig.20 - 21. A point source in the far zone produces in front of a hole aperture effectively a plane wave. In this chapter we considered only the case in which the point source on the optical axis (which is parallel to the $z_{n}$ axis) generates normally incident plane wave. It is obvious that the electrodynamics equations must be identical for all electromagnetic waves. In other words, the dispersion equation in the accelerated frame should be applied and kept in a consistent way for both incoming and scattered waves.
A dispersion equation in the case of Maxwell’s electrodynamics is reduced to $k_{x}^{2}+k_{y}^{2}+k_{z}^{2}-\omega^{2}=0$. The simplicity of considered case is that even after the Galilean transformation along the $x$-axis the dispersion equation in the accelerated frame will have the same (diagonal) form $k_{z}^{2}-\omega_{i}^{2}=0$ for the normal incident wave. Now let us see what happens with incoming wave which has rate of change of phase with respect to $x_{n}$. In other words, there is the angle $\alpha$ between the wave number vector $\vec{k}$ and the $z_{n}$-axis. For the case when the angle is small we can write $\alpha\simeq k_{x}/k_{z}$. Our problem now is to work out the parameters of the incoming wave in terms of an offset $x_{o}$ of the point source with respect to the optical axis. We can write the angle of wave tilt as $\alpha\simeq x_{o}/z_{s}$, where $z_{s}$ is the distance from the source to the screen. Let us now analyze the aberration of light radiated by a point source which is at rest with respect to the accelerated frame. The wavenumber vector of the incoming plane wave is fixed by initial conditions. Let us assume that the wavefront of the emitted light beam is tilted by an angle $\alpha$ before the acceleration. Then we have $k_{x}=\alpha k$, $k_{z}=\sqrt{\omega^{2}_{i}/c^{2}-k_{x}^{2}}$. Acceleration has an effect on the field equations. With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ equation Eq.(16) becomes: $k_{x}^{2}-2vk_{x}\omega_{i}/c^{2}+k_{z}^{2}-(\omega_{i}-vk_{x})^{2}/c^{2}=0$ in the first order approximation with respect to a small parameter $v/c$. The screen containing a hole is a kind of diffraction grating which breaks up the incident plane wave into a number of diffracted plane waves. So the wave-number vector of the transmitted plane wave along the $x_{n}$ direction is $k_{x}+K_{\perp}$ where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated transmittance. From the dispersion equation, we find the requirement that the wavenumber $K_{\perp}$ and the frequency change $\Delta\omega$ are related by $\omega_{i}\Delta\omega=-\omega_{i}K_{\perp}v+c^{2}k_{x}K_{\perp}$. The last equations state that transmitted light beam moves along the $x_{n}$ direction with group velocity $d\omega/dk_{x}=-v+c^{2}k_{x}/\omega_{i}$. Then the apparent angular position of the point source is connected with the problem parameters by the relation $\theta=-v/c+\alpha$.
Let us now look again at our dispersion equation. At first site the Doppler effect problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers. The question arises whether it is possible to determine experimentally the state of motion of the accelerated frame $S_{n}$ in the initial inertial frame $S$ by means of the Doppler shift $\Delta\omega=vk_{x}$ for accelerated source. In the Chapter 10 we demonstrate that there exist a specific feature of the Doppler shift measurements which prevents us to determine the velocity of the accelerated frame with respect to the initial inertial frame.
5.12 Absence of Reciprocity in the Aberration of Light Theory
In the present section we shall continue our discussion of the aberration of light radiated by a single moving emitter. Imaging that there are two identical emitters. Let us consider the case when the first emitter is at rest in an inertial frame and the second emitter is accelerated from rest up to velocity $v$ along the $x$ axis. Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the energy transport measurement. Now we must be careful about initial phasing of these emitters. As example, we consider the case in which initially the velocity component of the light beam along the $x$-axis is equal to zero. Then how does the light beam from the moving emitter looks? The inertial observer would find that angular displacement is equal to $\theta_{a}=v/c$. That is the radiation appears as shown in Fig. 22. This is example what called the phenomenon of aberration of light and it is well known.
Now let us return to observation of an accelerated observer. When the accelerated system starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. In this synchronization, the accelerated observer describes the light beam from an emitter at rest using usual Maxwell’s equations. One may ask: ”Where is the information about the observer acceleration recorded in the case of the Lorentz clock synchronization?” When the accelerated observer performed the standard procedure of clock synchronization, the time shift has the effect of rotation the plane of simultaneity, that is source radiation wavefront, on the angle $-v/c$. It is easy to show that information about the observer acceleration is recorded in the phase front orientation (with respect to the coordinate axes of the accelerated frame) of the light beam radiated from the source at rest in the accelerated frame.
Figure 22: The aberration of light from stationary and moving sources. According to the asymmetry between the initial inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of light. Namely, if the emitter is at rest relative to the fixed stars and the observer with measuring devices started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the emitter seen in the accelerated frame would jump by angle $\theta_{a}=-2v/c$. This situation is not symmetrical with respect to the change of the reference frames. If the observer is at rest relative to the fixed stars and the emitter started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the emitter seen in the inertial frame would jump by angle $\theta_{a}=v/c$.
Figure 23: Reciprocity in a theory of the aberration of light. One must use the same source phasing to demonstrate the reciprocity. After the redirection procedure the inertial observer would find that angular displacement is $\theta_{a}=2v/c$.
Let us describe what happens when accelerated observer performed the redirection of the accelerated emitter. We consider the case in which finally the (group) velocity component of the light beam along the $x_{n}$-axis is equal to zero. After this redirection procedure the inertial observer would find that angular displacement is $\theta_{a}=2v/c$. Suppose that the accelerated observer also performs an aberration measurement. Fig. 23 shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-2v/c$.
The situation can be described quite naturally in the following way. In this case we define reference directions using light beams. We demonstrated that symmetry is a correct concept in the measurements of the angular displacement between the inertial $z$ and accelerated $z_{n}$ reference directions as it must be.
Figure 24: Procedure of the accelerated source redirection. The information about the source acceleration is recorded in the point source offset with respect to the optical axis (which is parallel to the $z_{n}$ axis). The wavefront orientation has no exact objective meaning. Because of this, the explanation of the effect of aberration (for sketch simplification) is based on the use of the absolute time (left) and Lorentz (right) coordinatization to describe how the direction of a light beam depends on the source offset.
Let us return to the redirection procedure of the accelerated source. Now the question is, where is the information about the source acceleration recorded in the case of redirecting light of the accelerated emitter. In order to understand the redirection of the light source in an accelerated system, we have to watch the machinery of the source and see what happens when the redirection takes place. Fig. 13 shows a drawing of the ”plane wave” emitter type. Electrodynamics offers a procedure of redirection based on the offset of the point source with respect to the optical axis (which is parallel to the $z_{n}$ axis), Fig. 24. The information about the source acceleration is recorded in the point source offset with respect to the optical axis.
It is generally believed that the special relativity is reciprocal theory. The aberration of light problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers. In fact, without looking at anything external to the accelerated frame, one could determine the speed of the accelerated frame with respect to the initial inertial frame by means of aberration of point source measurements. No one has ever done such experiment, but we know what would happen from the astronomical observations. The motion of the stars with respect to the earth is never followed by any aberration. The aberration shift exist even in the case when star moves with the same velocity as the earth. There is an optical similarity between aberration of light from distant star (which is moving with the earth velocity) and from the earth-based point source. We will discuss this subject further in the Chapter 8.
One finds many books which say that we cannot distinguish by any experiment which observer remains ”at rest” during the acceleration, because according to the principle of relativity only relative motion has any physical meaning. It follows that any observed asymmetry would lead to a contradiction with the principle of relativity. This argument is wrong. We discuss the apparent conflict between aberration of light and the principle of relativity in the Chapter 12. It is demonstrated that there is no conflict between the fundamental structure of special relativity on the one hand, and the aberration of light phenomena. Principle of special relativity is irrelevancy of velocity with regard to physical laws, not with regard to anything.
5.13 Two Metrics
There is an important theorem in the special relativity which is this: as a result of the Galilean boost, we obtain metric of the accelerated emitter in the initial inertial frame Eq. 11. In this coordinatization, we describe the non-accelerated light source using diagonal (Minkowski) metric. We transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer using passive Galilean transformation: we substitute $x\to x-vt,t\to t$ into Minkowski metric of the inertial observer to obtain Eq.11. This metric describes the electrodynamics of the moving emitter with viewpoint of the inertial observer measurements.
The Langevin metric describes the electrodynamics of the accelerated emitter with viewpoint of the accelerated observer measurements as viewing this of the inertial observer. We transform coordinates of the inertial observer using the inverse Galilean transformation: we substitute $x\to x_{n}+vt,t\to t_{n}$ into Minkowski metric of the inertial observer to obtain Langevin metric Eq.15. Now we see that the Langevin metric Eq.15 of the accelerated emitter in the accelerated frame is the same as metric Eq.11 of the accelerated emitter in the inertial frame. In fact the difference is only that the sign of velocity $v$. These two metrics are actually two sides of the same coin. The point is that we describe the electrodynamics of the same accelerated (with respect to the fixed stars) emitter from the point of view of both an inertial observer and an accelerated observer. One must take into account that metric tensor must be continuous quantity. This problem in special relativity can be adequately treated only by an approach uses absolute time coordinatization. For this purpose, it is sufficient to relate the coordinates and time by the Galilean boosts.
The commonly accepted approach to special relativity does not account for this theorem. The metric Eq.11 is represented here for the first time. It is difficult to see why this should be, since Langevin metric (which is found by matching the accelerated frame and inertial frame metric tensors by the Galilean boost) is widely used in the applications to the optical phenomena in a rotation frame.
In the problem discussed we have still a puzzle. We create a so-called comoving coordinate system in the initial inertial frame, and then discussing radiation from the moving source in terms of the new (comoving) coordinate labels. After the passive transformation, the coordinates of the source transform as $x^{\prime}=x-vt,t^{\prime}=t$. According to the equivalence of the active and passive pictures within a single inertial frame, in our case of interest the Maxwell’s equations are always valid in the comoving coordinatization. The electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$. In other words, Minkowski metric $ds^{2}=c^{2}(dt^{\prime})^{2}-(dx^{\prime})^{2}-(dy^{\prime})^{2}-(dz^{\prime})^{2}$ describes the electrodynamics of the accelerated light source in comoving coordinatization. We will then ask how it is possible to take Minkowski metric for comoving coordinatization in the inertial frame and the Langevin metric in the accelerated frame. This approach needs an explanation.
The difference between measurements made in both reference frames is very interesting. First, we must now discuss a certain feature of the correspondence between passive and active transformations. It should be note that passive Galilean boost within a single inertial frame, $x^{\prime}=x-vt,t^{\prime}=t$, is simply another parametrization of the observations of the inertial observer. The Eq.11 describes the inertial observer measurements. According to active transformations, the motion of fixed stars with respect to the inertial observe and his devices not changes and there is equivalence of the active and passive Galilean boost. In contrast, the Langevin metric describes the accelerated observer measurements. According to the inverse Galilean boost, the motion of fixed stars with respect to the space-time grid $(t_{n},x_{n},y_{n},z_{n})$ of the accelerated observer changes.
On the other hand, the standard textbook presentation of the relativistic transformation is based on the hidden assumption that $x_{n}y_{n}z_{n}$-axes of the accelerated observer and $xyz$-axes of the inertial observer are parallel. In other words, it is assumed that observers have common 3-space. It is generally believed that simultaneous acceleration of the (rigid) reference space grid has direct physical meaning. This is misconception. The commonly accepted derivation of the aberration of light effect does not account for the conventionality of distant simultaneity. There is an uncertainty of the relative position along the $z_{n}$-axis in the $x$-direction of amount $z_{n}v/c$ due to the uncertainty in the moments of acceleration. In fact in this case we have deal with space-like separated events. The $z_{n}$ axis orientation has no exact objective meaning, because the relativity of simultaneity takes place. The statement that the $z_{n}$-axis orientation with respect to the $x$-axis of the inertial observer has objective meaning to within a certain accuracy can be visualized by the picture of $z_{n}$ axis in the vertical orientation with approximate angle extension (blurring) given by $\Delta\theta=v/c$. This inclination angle between the inertial $z$ and accelerated $z_{n}$ reference directions has no exact objective meaning at all since, due to finiteness of the speed of light, we cannot specify any experimental method by which this displacement could be ascertained. We shall discuss this subject further in the Chapter 7.
5.14 Ether Theory and the Aberration of Light Effect
5.14.1 Introductory Remarks
Now we wish to continue in our analysis a little further. We will look for a different way of calculating the aberration effect. Another solution to the accelerated emitter problem will be discussed going back to ether theory. We accept the ether theory in the original form: there is ether, which rules the speed of light. Because the ether is immovable, it is causes anisotropy in every frames moving relative to the ether. Because the non-inertial frame is accelerated with respect to the fixed stars; consequently we must feel an ”ether wind” when measuring light propagation. Firstly we discuss the effect of the ether wind on the light speed. According to the hypothesis of an ether at rest in the initial inertia frame, the velocity of light judged from an accelerated reference system would be $c+u$ for the beam propagating in the same direction as the accelerated system and $-c+u$ for the beam propagating in the opposite direction, where $u=-v$ is the ether velocity in the accelerated reference system. The aberration effect is an effect of the ether-wind on the speed of light (Fig. 25).
Figure 25: Aberration of light in an accelerated frame. According to the ether theory, the ether wind generates anisotropy in the accelerated ( relative to the ether) frame that is responsible for change of radiation direction.
5.14.2 Lorentz’s Theory of Corresponding States
In 1895, Lorentz derived a first-order version of the so-called ”theorem of the corresponding states” and used it to give a very general explanation of the null results of the first order optical interference experiments. Lorentz’s theorem connects pairs of electromagnetic field configurations, one in a system at rest in the ether and one in a system in uniform motion through the ether, and say that, if one member of the pair is allowed by the laws of electrodynamics, so is the other. The theorem works as follows. The mathematical description of the configuration in the moving system in therms of ”local time” (instead of ”real time” ) is the same as the description of its corresponding state in the system at rest in terms of the real time. Lorentz showed that two field configurations related to one another in many of their observable properties, in particular, they give rise to identical interference patterns. Thus, it is not possible to detect the orbital velocity using interference (e.g. Michelson-Morley) experiments (2).
We now discuss the application of Lorentz’s pre-relativistic theory to the aberration of light effect in a non-inertial frame of reference. Above we considered the problem of aberration of light in the accelerated systems on the basis of theory of special relativity. We described the effect of aberration by working only up to the first order. In fact, the appearance of relativistic effects in optical phenomena does not depend on a large speed of the radiation sources. Light is always a relativistic object.
We have already discussed the asymmetry between initial inertial frame and accelerated frames in this chapter. Here’s what we found out there: We pointed that metric tensor must be a continuous quantity. This problem in special relativity can be adequately treated only by an approach which uses the absolute time coordinatization. The usual Maxwell’s equations are valid in the initial inertial frame where optical system (e.g. configuration of light source and lens, Fig. 13) is at rest and procedure of slow clock transport is used to assign values of time coordinate. We require that metric of a non inertial frame associated with accelerated (with respect to the initial inertial frame) optical system would pass continuosly into Langevin metric. For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer $(x_{n},y_{n},z_{n},t_{n})$ to the coordinates $(x,y,z,t)$ of the inertial observer by the inverse Galilean boost $x_{n}=x-vt,y_{n}=y,z_{n}=z$, $t_{n}=t$. Hence the wave equation transforms into Eq. 16 where coordinates and time are transformed according to the Galilean transformation. The Langevin metric, together with initial condition, describes optical effects in the accelerated frame. Above in this chapter we considered the aberration of light problem in the accelerated frame. The presented approach uses the Fourier transform method. The essential asymmetry is rooted in the electrodynamics equation Eq. 16. As one of the consequences of the crossed term in the wave equation we found the group velocity that is responsible for the aberration. However, there is another satisfactory way of explaining the effect of aberration of light in the accelerated frame. The explanation consists in using a procedure of clock resynchronization. Using Einstein’s clock synchronization we actually select a Lorentz coordinate system for the optical setup at rest in the accelerated frame. In this synchronization we describe the field configuration using usual Maxwell’s equations. The time $t^{\prime}_{n}$ under the Einstein’s synchronization is obtained by substituting $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first order approximation. Clearly, after the resynchronization, the asymmetry is rooted in the initial condition. The plane radiation wavefront of the emitter rotates in the accelerated frame on the angle $-v/c$ after resynchronization.
Lorentz’s treatment of optical setup in an inertial frame moving through the ether is similar to our treatment of such setup in the accelerated (with respect to the fixed stars) frame. Now we must analyze this similarities. One thing is clear: Our initial inertial frame (i.e. inertial frame without accelerational history with respect to the fixed stars) can be interpreted as what Lorentz thought of as a frame at rest in the ether. In the Lorentz’s theory, and more generally in Physics before the development of special relativity, the connection between an observer at rest who used $(x,y,z,t)$ coordinates and an observer bound to the moving body who used $(x_{n},y_{n},z_{n},t)$ coordinates was given by the Galileo relation: $x_{n}=x-vt,y_{n}=y,z_{n}=z,t_{n}=t$. Galilean boost in Lorentz’s theory is empirically equivalent to the smooth tailoring of the metric tensor. In this case the metric of initial inertial frame would pass into the Langevin metric of the accelerated frame. Lorentz of course would not have thought of his theory in this way, since he did not have the resources of Minkowski space-time to do so.
Because of this Galilean boost, the homogeneous wave equation for field in the moving frame, Eq. 16, has not standard form. The main difference consists in the crossed term, which complicates the solution of the equation. To get around of this difficulty, Lorentz observed that simplification is always possible. In solving wave equation, Lorentz found it convenient to introduce new variable $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$. It is called ”local time”. The substitution of this time shift in Lorentz theory is a mathematical device used to change the electrodynamics equations into Maxwell’s form.
At first cite, it seems that this Lorentz’s theorem of corresponding states for first order phenomena provides all the guidance we need to exact treatment of aberration of light in the accelerated frame. One could naively expect that Lorentz’s theory would predict in full generality that observation of aberration of light from the earth-based incoherent source would reveal the earth orbital motion, Fig. 24. At this point, a reasonable question arises: why Lorentz concluded from his theorem of corresponding states that no first order effects of the orbital motion of the earth could be detected?
5.14.3 Analogy with the Sound Theory
A comparison between the Lorentz ether theory and the sound theory might help here. Above in the Section 4.15 we considered the sound emitter problem in the atmosphere frame. We considered the case of a screen, in the atmosphere frame, moving with velocity $v$ along its surface. According to the theory of sound, there is no deviation of the energy transport for sound transmitted through a hole in the moving screen. A plane wave is falling normally on the screen, and generates a transmitted oblique sound beam, Fig. 7. The group velocity is equal to the phase velocity. The wave equation $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$ always valid in the atmosphere frame. This equation is not applicable from the viewpoint of an observer at rest with respect to the accelerated frame. Accelerations with respect to the atmosphere frame have an effect on the propagation of sound. In particular, the wave equation transforms into
$\left(1-\frac{v^{2}}{v_{s}^{2}}\right)\frac{\partial^{2}f}{\partial x_{n}^{2}}+2\left(\frac{v}{v_{s}}\right)\frac{\partial^{2}f}{\partial t\partial x_{n}}+\frac{\partial^{2}f}{\partial y_{n}^{2}}+\frac{\partial^{2}f}{\partial z_{n}^{2}}-\frac{1}{c^{2}}\frac{\partial^{2}f}{\partial t^{2}}=0~{},$ (18)
where coordinates and time are transformed according to the Galilean transformation $x_{n}=x-vt$, $t_{n}=t$. The wave equation needs to be integrated with initial condition for the sound wavefront. Acceleration has no effect on the wavefront orientation. So that wavefront of the emitted sound beam is perpendicular to the $z_{n}$ axis. Because the air is immovable, it is causes anisotropy in every frames moving relative to the atmosphere frame. Consequently we must feel an ”air wind” when measuring sound propagation. According to the wave equation Eq. 18, the velocity of sound judged from an accelerated reference system would be $v_{s}-v$ for the beam propagating in the same direction as the accelerated system and $-v_{s}-v$ for the beam propagating in the opposite direction, where $-v$ is the wind velocity in the accelerated reference system. The aberration effect is an effect of the air wind on the speed of sound (Fig. 25).
Let us now analyze the aberration of sound radiated by a single ”plane-wave” emitter in the atmosphere frame. The wavefront radiated by the plane-wave emitter at given instant time is constructed as an envelope of all the spherical wavelets from the point sources on the emitter that have radiated until this instant. It is not difficult to see that when the emitter is moving at a constant velocity $v$ along the $x$ axis, the wavefront will propagate at the speed $v_{s}$ at the angle $\theta=v/v_{s}$ from the vertical. In fact, due to the motion of the emitter, constructive interference occurs when the projection of phase velocity on the $x$ axis, is equal to the velocity of the emitter. According to classical kinematics, the extra phase chirp $d\phi/dx=k_{x}=vk/v_{s}=v\omega_{s}/v_{s}^{2}$ is introduced. As a consequence of this, the plane wavefront rotates after the emitter acceleration. Then, the radiated sound beam is propagated at the angle $v/v_{s}$ with respect to the $z$axis, yielding the phenomenon of the aberration of light in the atmosphere frame. Now let us return to observations of an accelerated observer. The wave equation needs to be integrated with initial condition for the sound wavefront. Galilean boost has no effect on the wavefront orientation. So that wavefront of the emitted sound beam is tiled at the angle $v/v_{s}$ with respect to the $z_{n}$ axis. The wave number vector of the radiated plane wave is fixed by initial condition. Then we have $k_{x}=vk/v_{s}=v\omega_{s}/v_{s}^{2}$. The non-diagonal wave equation Eq. 18 states that radiated sound beam moves along $x_{n}$ axis with group velocity $d\omega/dk_{x}=-v+v_{s}^{2}k_{x}/\omega_{s}=0$ (see Section 5.11 for more detail).
Thus we obtain the same addition theorem for sound beam velocities as for the particle velocities in the classical physics. So that the two (ether and sound) theories were really just variants of the same classical theory. According to Lorentz, the wave equation of the moving light (sound) emitter in the ether (atmosphere) frame is the same as wave equation of the emitter at rest. As a consequence of this, after the boost we can see that screen acceleration in the ether (atmosphere frame) has no effect on the wavefront orientation and the wave equation. Thus Lorentz concluded that one could not determine the orbital motion of the earth by means of aberration of earth-based source measurements. We come to the conclusion that the Lorentz analysis of the aberration of light does not take into account the fundamental difference between velocity of light and velocity of sound.
When Lorentz formulated his pre-relativistic theory he assumed a space-time background with a well defined geometrical structure: Newtonian space-time. At close look at the physics of this subject shows that there is a real difference between the inertial frame without accelerational history with respect to the fixed stars and the Lorentz classical ether at rest. The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of unified space-time model. In special relativity the absolute space and the absolute time of Newton are fused into absolute space-time.
5.15 Bibliography and Notes
1. The aberration of light problem in the accelerated systems is solved with discovery of the essential asymmetry between the non-inertial and the inertial observers. This asymmetry is of the same nature as that of the well-known Sagnac effect [28, 29, 30]. For instance Langevin’s 1921 explanation of the Sagnac effect rested upon the assertion that ”any change of velocity, or any acceleration has an absolute meaning.” [31].
2. For discussion of the relationship between the general accepted way of looking at special relativity and the Lorentz’s ether theory we suggest reading the paper [32].
6 Stellar Aberration
6.1 The Corpuscular Model of Light and the Stellar Aberration
It is generally believed that the phenomenon of aberration of light could be interpreted, using the corpuscular model of light, as being analogous of the observation of the oblique fall of raindrops by a moving observer. This is a classical kinematics method to the computation of the stellar aberration used in astronomy for about three hundred years (1). What had to be added in the 20th century, was that the dynamical laws of Newton for light were found to be all wrong, and electromagnetic wave theory had to be introduced to correct them.
As well-known, according to textbooks the physical basis of the stellar aberration is the fact that the velocity of light is finite and changes its direction when seen from another reference frame. It is a consequence of the formula for addition of velocities applied to a light beam when the observer is changing its reference frame.
For an observer on the earth, it is, with respect to the solar referential frame, of about $v=$ 30 km/s, corresponding to the earth motion around the sun. It is sufficient to describe the effect of stellar aberration by working only up to the first order $v/c=10^{-4}$. According to the conventional approach, the study of stellar aberration is intimately connected with the old (Newtonian) kinematics: the Galilean vectorial law of addition of velocities is actually used.
6.2 Wave Theory of the Stellar Aberration
Most authors treat light propagating through the telescope barrel as rain of photons, and not as a plane wave (2) . Questioning the validity of standard reasoning we argue that a satisfactory treatment of stellar aberration should be based on the coherent wave optics. It can be easily demonstrated that the light produced by a distant star is approximately coherent over a circular area whose diameter, in all practical cases, is much larger than the telescope diameter. Hence, we sample such a tiny portion of the coherent area of the starlight with our telescopes that the waveforms are effectively (flat) plane waves (3).
Let us verify that this assertion is correct. The spatial coherence of a light beam generally has to do with the coherence between two points in the field illuminated by the light source. The meaning of spatial coherence can best to understood with the help of Young’s two-pinhole experiment (see Fig.26). In its elementary sense, the degree of coherence between the two points simply describes the contrast of the interference fringes that are obtained when the two points are taken as secondary sources. Let a source $S$ illuminates the two pinholes $S_{1}$ and $S_{2}$, as shown in Fig.26. The source is perfectly non-coherent. That is to say, no interference fringes can be obtained by placing two pinholes in the plane of the source. It was shown, however, that if the two pinholes are placed far enough away from the non-coherent source, interference fringes of good contrast can be obtained. It is sometimes said that the spatial coherence in light beams increases with distance ”by mere propagation”. It would be nice to find an explanation which is elementary in the sense that we can see what is happened physically.
Suppose that a quasi-monochromatic wave is an incident on an aperture in an opaque screen, as illustrated in Fig. 26. In general, this wave may be partially coherent. The detailed structure of an optical wave undergoes changes as the wave propagates through space. In a similar fashion, the detailed structure of the spatial coherence undergoes changes, and in this sense, the transverse coherence function is said to propagate. Knowing the spatial coherence on the aperture, we wish to find the spatial coherence on the observing screen at distance $z$ beyond the aperture. The stellar radiation is a stochastic object and for any starlight beam there exist some characteristic linear dimension, $\Delta r$, which determines the scale of spatially random fluctuations. Fig.27 illustrates the type of spiky pattern on an aperture in an opaque screen. When $\Delta r\ll d$, the radiation beyond the aperture is partially coherent. This case is shown in Fig.27. Here $\Delta r$ may be estimated as the typical linear dimension of spikes.
[…]
Figure 26: Young’s interferometer.
Figure 27: Geometry for propagation of spatial coherence.
Figure 28: Reciprocal width relation of Fourier transform pair. Free space basically act as a Fourier transformation. This means that the field in the far zone is, phase factor and proportionality factor aside, the spatial Fourier transform of the field at the source plane.
A star is an incoherent source with the scale of spatially random fluctuations $\lambdabar\Delta r\sim\lambdabar$. Such a source emits radiation in all directions. Also, an angular spectrum of the starlight contains spikes with the typical width $(kd)^{-1}$, where $d$ is the star diameter. In the case of astronomical observations by telescope, each spike on the earth’s surface has a finite thickness which is the same order as the spatial coherence length of the starlight. It can be said that the coherent area of the observed starlight would be of the order of $z/(kd)$, where $z$ is the distance between the star and earth. The star Bradley chose was Draconis, which is one of the nearest stars. In this case the diameter of the coherent area on the earth surface can be estimated to be about 100 m.
6.3 Examination of Stellar aberration in the Context of Special Relativity
It is generally believed that the theory of relativity appears to conform to the phenomenon of stellar aberration discovered by Bradley by claiming it is a consequence of the motion of observers relative to light sources(4).
The lack of symmetry between the cases when either the star or telescope on the earth is moving is shown clearly on the basis of the separation of binary stars. Spectroscopic binaries have velocities exceeding the earth’s velocity round the sun. They revolve around their common center of gravity within days, a period during which the motion of the earth is practically constant. The components of the binary system should be easily separable, when their changing velocities are comparable to the earth’s velocity round the sun. This is, however, not observed (Fig.29-Fig.30) (5).
Figure 29: Aberration shift as inferred from astronomical observations. The plane wave fronts of starlight entering the telescope are imaged by the lens to a diffraction spot which lies in the focal plane. Two cases are selected with different velocities of a earth-based telescope and a star. In the first case (left) the star is at rest with respect to the fixed stars. In the second case (right) the star moves with the same velocity as the earth.
Figure 30: Aberration shift as predicted by the conventional theory of the stellar aberration. Two cases are selected with different velocities of a earth-based telescope and a star. In the first case (left) the star is at rest with respect to the fixed stars. In the second case (right) the star moves with the same velocity as the earth. The conventional theory predicts no aberration of the image in this case.
In Chaper 4, we presented a critical reexamination of the textbook statement that wavefronts and raindrops are to have the same aberration. We used the theory of relativity to show that when one has a transversely moving mirror and a plane wave of light is falling normally on the mirror, there is a deviation of the energy transport for light reflected from the mirror. This effect is a consequence of the fact that the Doppler effect is responsible for angular frequency dispersion of light waves reflected from the moving mirror with finite aperture. As a result, the velocity of the energy transport is not equal to the phase velocity. According to the Babinet’s principle, this remarkable prediction of our theory should be correct also for light transmitted through a hole punched in a moving opaque screen or, consequently, through the moving open end of a telescope barrel.
In principle the binary star paradox is resolved by noting that when light passes through the end of a telescope barrel we have a light beam whose fields have been perturbed by diffraction, and, therefore, do not include information about star motion relative to the fixed stars. If the telescope were at rest relative to the fixed stars and the star started to move from rest, then the apparent position of the star seen in the telescope would never jump by any angle.
However, other difficulties arise in the explanation of the earth-based observations, i.e the change in an apparent positions of the fixed stars, which happens when the earth-based telescope changes of its motion relative to the fixed stars. It should be stressed that it is the telescope and not the star that must change its velocity (relative to the fixed stars) to cause aberration.
An objective of this chapter is to consider non-relativistic interpretations of the stellar aberration, and clearly demonstrate that they are incorrect. Classical kinematics effects leads to serious mistakes if applied to the computation of stellar aberration as seen on the earth rotating around the sun (i.e. in a non-inertial geocentric frame of reference). The problem of the earth-based measurements is solved with the discovery of the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars (6). We derive the aberration for a pulse of light traveling on the surface of the earth using the Langevin metric. Though stellar aberration behaves asymmetrically, it does not contradict special relativity, because the heliocentric (i.e. sun-based) reference system is inertial and the geocentric reference system is non-inertial.
6.4 Heliocentric Inertial Frame of Reference
Above, we demonstrated that when one has some hole in the opaque screen at rest with respect to the fixed stars and the transversely moving point source, there is no aberration (deviation of the energy transport) for light transmitted through the hole. The absence of the effects of moving (relative to the fixed stars) source in this setup automatically implies the same problem for stellar aberration theory in the heliocentric frame of reference. How shall we solve it? It is like a hole-point source problem with the end of the telescope barrel as a hole.
Suppose that an observer, which is at rest relative to the telescope, performs the direction of the energy transport measurement. At close look at the physics of this subject shows that in the heliocentric frame of reference, where the telescope is at rest, we have actually the problem of steady-state transmission. Then how does the transmitted light beam looks? It looks as though the transmitted beam is going along the telescope axis because it has lost its horizontal group velocity component. That is the transmission appears as shown in Fig 12. It takes the case of a telescope positioned perpendicular to the plane phase front. In other words, the telescope pointed directly at the star. If the motion of the star is parallel to the phase front (i.e. perpendicular to the telescope axis), starlight entering the end of the telescope would be able to pass its full length.
One of the most important conclusions of the discussion presented above is that the aberration of starlight phenomenon is absent in this situation. In particular, the binary components remains unresolved which means that their velocity has no influence on aberration.
6.5 Earth-Based Non-Inertial Frame of Reference
The analogy between the obliquity of raindrops and the stellar aberration is incorrect. Only on the basis of the theory of relativity and the wave optics, we are able to describe all earth-based experimental observations of stellar aberration. Our theory predicts an effect of stellar aberration in complete agreement to the Bradley’s results, Fig. 31. According to the asymmetry between the inertial and rotating frames, there is a remarkable prediction on the theory of the aberration of light. Namely, if the telescope is at rest relative to the earth and the earth rotating relative to the fixed stars, then the direction of a star as seen from the earth is not the same as the direction when viewed by a hypothetical sun-based observer. Apparent angle is less than the actual angle. The difference between the actual angle and apparent angle $\theta_{a}$ is connected with the physical parameters by the relation: $\theta_{a}=v/c$, where $v$ is the velocity of the earth in its orbit around the sun. It could be said that the crossed term in metric Eq.(15) generates anisotropy in the rotating frame that is responsible for the change of radiation direction (aberration). That is the transmission through the telescope aperture in the rotating frame appears as shown in Fig. 20.
Figure 31: The direction of a star as seen from the earth is not the same as the direction when viewed by a hypothetical observer at the sun center. Apparent angle $\phi$ is less than the actual angle $\theta$. The difference between the actual angle and apparent angle is connected with the physical parameters by the relation: $\theta-\phi=v/c$, where $v$ is the velocity of the earth in its orbit around the sun.
The stellar aberration in the geocentric frame of reference is considered independent of the star speed and to have just a local origin exclusively based on the observer speed with respect to the sun.
The main facts which a theory of stellar aberration in the earth-based frame of reference must explain are (1) the annual apparent motion of the fixed stars about their locations and (2) the null apparent aberration of rotating binary systems. We presented here a theory which accounts for all these, and in addition gives new results. We demonstrated that the aberration of light is a complex phenomenon which must be branched out into a number of varieties according to their origin. These branching out takes place depending on what is the cause of aberration - whether it is a jump in the velocity (relative to the fixed stars) of the observer or of the light source. Aberration could undergo further splitting - depending on the physical influence of the optical instrument on the measurement. According to our interpretation, there are many kinds of aberration and the stellar aberration in the earth-based frame is only one of these. We demonstrated that all earth-based experiments can be explained on the basis of the effect of the measuring instrument (i.e. physical influence of the telescope on the measurement) and the acceleration of traveling earth-based observer relative to the fixed stars.
6.6 Physical Coordinate System in Space
The aberration of light is the geometric phenomenon. In order to detect the aberration effect inside the earth-based frame, it is obvious that some coordinate system with reference direction is needed. A conventional approach to the aberration of light effect is forcefully based on a definite assumption of reference direction, but this is actually a hidden assumption. Traditionally the physical interpretation of the aberration of light effect in terms of measurements performed with rods that are at rest in the observer’s frame. This (local frame of reference) convention is self-evident and this is the reason why it is never discussed in the aberration of light theory
Fortunately, it is possible to find the space description on a more fundamental level than that of measuring rods. The reference axis in the earth-based frame is formed by gravitation field vector. The plumb-line direction is known as the nadir, leading to the earth’s center. This is the most fundamental local earth-based coordinate system. For example, Bradley used the vertically-mounted telescope. The star he chose was Draconis because it transited almost exactly in zenith. The traditional plumb line provided a sufficiently accurate zenith-point for observations of the stellar aberration.
6.7 Region of Applicability
The region of applicability of our stellar aberration theory is more wider than one might think. Above we assumed for simplicity that a star image is actually a point spread function in the image plane of a telescope. In other words, it is assumed that the input signal is effectively a plane wave.
Suppose that a telescope is able to distinguish detail in the star image. A complete understanding of the relation between object and image can be obtained if the effect of diffraction are included. The effect of diffraction is to convolve that ideal image with the Fraunhofer diffraction pattern of the telescope pupil.
The star is a spatially completely incoherent source (7). This means that such source is actually a system of elementary (statistically independent) point sources with different offsets. An elementary source with a given offset produces in front of a telescope pupil effectively a plane wave. An elementary source offset tilts the far zone field. Radiation field generated by an completely incoherent source can be seen as a linear superposition of fields of individual elementary point sources. The image of an elementary point source is a point spread function. In other words, there is always the physical influence of the telescope on the measurement of a completely incoherent source. It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single point source characteristics like independence on source motion. This argument gives reason why our theory of stellar aberration is correct also for imaging of arbitrary completely incoherent sources.
6.8 Bibliography and Notes
1. The phenomenon of the annual apparent motion of celestial objects about their locations, named stellar aberration, was discovered by Bradley in 1727, who also explained it employing the corpuscular model of light [33].
2. There is a widespread belief that the rays of light coming from the star are falling on the telescope tube and not interact with its sides. To quote French [4]: ”Regarding light as being composed of rain of photons, we can easily calculate the change in apparent direction of a distant object such a star.” Similar statements can also be found in recently published textbooks. To quote Rafelski [5]: ”We consider a light ray originating in a distant star. … The following discussion is addressing the observation of well focusing light-rays and not (spherical) plane wave light. … This picture of the experimental situation is accurate since the light emitted by a star consists of an incoherent flux of photons produced in independent atomic processes.” This conclusion is incorrect. We emphasize that the design of a telescope optical system is based on the classical theory of diffraction (originated directly from the classical electromagnetic theory). The resolution of a telescope is limited by the diffraction of light waves and the telescope always influences the measured radiation (due to unavoidable diffraction of a star light by the telescope aperture). In a well corrected optical system and in a circular pupil, the size Airy disc of diffraction pattern (i.e. size of the image of a point source) is inversely proportional to the diameter of the pupil.
3. A star is a completely incoherent source. The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34]. Any star can be considered as very far away from the sun. In all cases of practical interest, telescopes are situated in the far-zone of the (distant star) source. Consider a star like Sirius, which is one of the the nearest stars. The coherent area of light observed from Sirius has a diameter of about 6 m. This correlation was observed by Brown and Twiss in 1956 [35].
4. It is widely believed that stellar aberration depends on the relative velocity of the source (star) and observer. In the paper on the theory of relativity Einstein deduced the aberration formula from the idea that the velocity of $v$ is the relative velocity of the star-earth system. The idea was represented by many authors of textbooks. To quote Moeller [13]: ”This phenomenon, which is called aberration, was observed … by Bradley who noticed that the stars seem to perform a collective annual motion in the sky. This apparent motion is simply due to the fact that the observed direction of a light ray coming from a star depends on the velocity of the earth relative to the star.”
5. In 1950 Ives [36] stressed for the first time that the presence of binaries in the sky gave rise to an important difficulty for the theory of relativity. It is stated that the idea that aberration may be described in terms of relative motions of the bodies concerned is immediately refuted by the existence of spectroscopic binaries with velocities comparable with that of the Earth in its orbit. Still this exhibit aberrations not different from other stars. For example, a spectroscopic binary, Mizar A, has well-known orbital parameters, from which can be calculated an observable angular separation of 1’10” if aberration were due to relative velocity. The empirical value is less than 0.01”, clearly incompatible with authors of textbooks point of view [37, 38]. There is no available explanation for the fact that, while the observational data on stellar aberration are compatible with moving earth, the symmetric description, when the star possesses the relative transverse motion, does not apparently lead to observations compatible with predictions.
6. Stellar aberration exist as observable phenomenon only in the presence of changing states of motion (i.e. acceleration). The problem of the earth-based measurements is related with the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars. This has been recognized by some expects, perhaps most explicitly by Selleri who states [38] that ”Thus a complete explanation of the aberration effect is given in terms of variations of the earth absolute velocity due to orbital motion, while the star/earth relative velocity is irrelevant. Thus acceleration (of planet, this time) enters once more into a game.”
7. From everything that authors of textbooks have written about the star light (”rain of photons”) it clearly follows that they did not understand the several aspects of the statistical optics. In the statistical optics for star emission, the fields are described classically at the level of Maxwell’s equations and the emitting medium is treated (as a ensemble of atoms) by quantum mechanics. The atomic polarization induced by the radiation field appears as a driving term in Maxwell’s equations and sustains the oscillations. The star is actually a system of elementary (statistically independent) point sources with different offsets. The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$, where $\lambda$ is the (visible) radiation wavelength. Within an elementary source volume ($\lambda^{3}$) there is an enormous number of atoms. Semi-classical theory is treating these atoms as coherent, interacting radiating dipoles. The induced macroscopic dipole moment in an elementary source leads to the classical electromagnetic radiation. The distinguishing characteristic of the statistical optics is the fact that electromagnetic fields are treated in a completely classical manner until they interact with the atoms of the photosensitive material on which they are incident. Thus there is no necessity to deal with quantization of the electromagnetic field; only the interaction of the classical field and matter is quantized. Thus in the semiclassical theory, where the atom is treated quantum mechanically but the electromagnetic field is treated classically, the photodetector has the effect of converting the continuous cycle-averaged classical intensity of the light beam into a succession of discrete photocounts. Light interact with matter in a fundamentally random way. As a consequence, any measurement of (star) light will be accompanied by certain unavoidable fluctuations in the detector. Stochastic fluctuations of the classical intensity can influence the statistical properties of the photoevents that are observed. Note in particular that the variance $\sigma^{2}_{K}$ of photoevents $K$ consists of two distinct terms, each of which has a physical interpretation: $\sigma^{2}_{K}=[K^{2}-<K>^{2}]/<K>^{2}=1/<K>+\sigma^{2}_{W}$. The first term is simply the variation of the counts that would be observed if classical intensities were constant and the photocounts were purely Poisson. We refer to this contribution to the count fluctuations as ”quantum noise”. The second therm, $\sigma^{2}_{W}$, is clearly zero if there are no fluctuations of the classical intensity. For instance, in the case of laser light, this component would be identically zero, and the count variance would be simply that arising from the photon ”shot noise”. This is intrinsic fluctuations associated with quantum fluctuations of vacuum and can give information only on the mean rate ot photons production. When star (thermal) light is incident on the photodetector, the classical fluctuations are nonzero, and the variance of the photocounts is larger than that expected for a Poisson distribution. The classical thermal fluctuations contains rich information about radiation source. In particular, thermal noise provide us with information for analyzing (using Van Cittert-Zernike theorem) the spatial source profile. The ratio of the classical variance to the ”photon shot noise” variance is equal to $\delta_{c}=<K>\sigma^{2}_{W}$. Parameter $\delta_{c}$ named as the photocount degeneracy parameter and can be interpreted as the average number of counts that occur in a single coherence interval of the incident radiation. When $\delta_{c}\ll 1$, it is highly probable that will be no more than one count per coherence interval of the wave, with the result that shot noise predominates over classically induced noise. On the other hand, when $\delta_{c}\gg 1$, there are many photoevents present in each coherence interval of the wave. The result is ”bunching” of the photoevents by the classical intensity fluctuations and an increase of the variance of the counts to the point where the classically induced fluctuations are far stronger than the shot noise variations.
7 Notion of Ordinary Space in Special Relativity
We introduce new approach to the aberration of light theory in non-inertial frames of reference, finding another way in which this complicated problem can be solved. For example, when viewed from the inertial frame, the aberration of light effect in the accelerated frame (Fig. 16) is easy to calculate in the framework of the special theory of relativity taking advantage of the relativity of simultaneity. In contrast, in order to compute the aberration of light effect in non-inertial systems we used a metric tensor.
7.1 Inertial Frame View of Observations of the Noninertial Observer
First we want to rise the following interesting and important point. The laws of physics in any inertial reference frame should be able to account for all physical phenomena, including the observations made by non-inertial observers. It is widely believed that all phenomena in non-inertial (e.g. rotating) reference system should be considered only in the framework of the space-time geometric approach using Langevin metric. However, the analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard (Einstein) special relativity taking advantage of well known relativistic kinematic effects.
For example, when viewed from the inertial lab system, the interpretation of the Sagnac effect is simple and the phase difference (attributed to the Sagnac effect) between counter-propagating waves may be derived from the relativistic law of velocity composition. In other words, Sagnac effect in the rotating frame, as viewed from the inertial lab frame, presents a kinematic effect (Einstein’s velocity addition) of special theory of relativity. It is implicitly assumed that the speed of light is $c$ in the lab frame which, in accordance with Einstein special relativity, is independent of motion of the source. In contrast, the Sagnac effect is not easy to calculate in a frame of reference attending rotation. In this case, authors of textbook used a metric tensor (Langevin metric) in a plane four-dimensional Minkowski space-time to calculate the propagation time difference between counter-running waves. Can we not look at the aberration of light effect in the same way? Here also we used the Langevin metric in a non-inertial frame of reference.
Let us first consider the application of the classical kinematic method to the computation of the aberration of light effect in a non-inertial frame, as viewed from an inertial frame. Authors of textbooks obtain the same addition theorem for light beam velocities as for particle velocities in the classical physics. This result is particularly remarkable because the study of light aberration is intimately connected with old kinematics. If we neglect $v^{2}/c^{2}$ comparison with unity, the Galilean vectorial law of addition of velocities is actually used. However, the classical kinematics method leads to a serious mistake if applied to the computation of the aberration of light effect. Let us consider the aberration of light in the inertial frame. According to our approach, there is a remarkable prediction concerning the deviation of the energy transport for light transmitted through a hole on a moving screen, Fig. 9. In order to predict the result of the aberration measurement the accelerated observer should use the Langevin metric. The crossed term in this metric generates anisotropy in the accelerated frame that is responsible for the change of radiation direction, Fig. 14. In contrast, based only on the (Galilean or Einstein’s) velocity addition, we arrive at the conclusion that there is no aberration of the transmitted light beam in the accelerated frame.
The error in the last argument follows from the fact that incorrect assumption that an inertial observer and accelerated observer have common 3-space is used. According to the special relativity, inertial observer and accelerated (with respect to the fixed stars) observer have different 3-spaces. Errors inherent in the classical kinematics method applied to the computation of the aberration of light effect are due to the use of the velocity of light simply as a classical velocity. Relativistic effects do not have a place in this description. According to classical approach, there is no principal difference between the aberration of light and the aberration of raindrops. We first notice that one of the postulate of special relativity states that if two distinct events cannot be connected by a causal signal that travels no faster than a light signal, then they cannot be connected by a causal signal at all. We will show that the aberration of light effect is a corollary to the relativistic kinematic effects. The appearance of relativistic effects in radiation phenomena does not depends on a large speed of the radiation sources. Light is always a relativistic object. In particular the relativity of simultaneity is responsible for aberrations to the first order in $v/c$.
Now, it is very interesting to show that the geometric effects in our ordinary space world are closely associated with the relativity of simultaneity. In Fig. 16 the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light. The question cannot be avoided relative to what a light beam propagated in the accelerated frame with angular displacement $-v/c$? Suppose that an observer in the accelerated frame performs the direction of the light beam measurement and the plane wavefront of transmitted light beam is imaged by a lens to a diffraction spot which lies in the focal plane on the optical axis. Measurement of the direction of the optical axis with respect to the frame axes is equivalent to the determination the angular displacement. In order to detect the aberration of light effect inside the accelerated frame, it is obvious that some coordinate system with reference direction is needed. We must inquire in detail by what method we assign coordinates. This method involves some sort of physical procedure; eventually it must be such that it will be give us coordinates in both (inertial and accelerated) frame of reference.
Figure 32: Aberration of light in the accelerated frame of reference. In the Einstein time synchronization for the screen, the accelerated observer describes the transmitted light beam and the reference beam from an emitter at rest using usual Maxwell’s equations. The $x_{n}$-axis is parallel to the wavefront of the reference beam. The aberration increment may be defined as an angle between the reference and transmitted light beams.
Figure 33: Aberration of light in the inertial frame of reference. In the absolute time synchronization for the screen, the inertial observer describes the reference beam from an emitter at rest using usual Maxwell’s equations. The $x$-axis is parallel to the wavefront of the reference beam. The aberration increment may be defined as an angle between the reference and transmitted light beams.
In ordinary space we find that the accelerated frame moves with respect to the inertial frame along the line motion and the inertial frame moves with respect to the accelerated frame along the line motion. The angle between the axis of the observer’s coordinate system and the line motion is a simple ordinary space geometric parameter. Using the line motion as a reference $x_{n}$-axis, the accelerated observer can then define the second reference axis. We need to give a ”practical”, ”operational” answer to the question of how to assign an axis perpendicular to the $x_{n}$-axis. Clearly, it is possible to define a reference direction using a light beam. We will define the second reference direction in the following way. Let us suppose that the aberration direction inside the accelerated frame is determined with reference to the fixed direction of light beam from a ”plane-wave” emitter which is at rest in the accelerated frame. In other words, a coordinate system is formed here by electromagnetic axis and line motion. The motion of the aberrated light beam are assumed for simplicity, to lie in the same plane and the angular position of the aberrated beam is described by one angle.
When the accelerated system starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. The Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all direction. In this synchronization, the accelerated observer describes the light beam from an emitter at rest using usual Maxwell’s equations. According to Maxwell’s electrodynamics, light is always emitted in the direction normal to the radiation wavefront. We consider the case in which the velocity component of the (reference) light beam along the $x_{n}$-axis is equal to zero. In other words, the $x_{n}$-axis is parallel to the wavefront of the reference ”plane wave”, Fig. 32. The number that specifies the aberration increment of the light beam transmitted through the hole may be defined as an angle between the reference and transmitted light beams.
Figure 34: Inertial frame view of observations of the accelerated observer. In the Lorentz coordinatization, the reference light beam in the accelerated frame is emitted in the direction normal to the wavefront. If make a Lorentz boost, we introduce a time transformation $t\to t+xv/c^{2}$ and the effect of this transformation is just an inclination of the electromagnetic axis in the accelerated frame.
Let us try to get an understanding of the relationship between the reference directions inside the inertial frame and the reference directions inside the accelerated frame. Such approach in our case of (aberration of light) interest is based on the use of local (reference) light sources. We determine the reference directions perpendicular to the line motion in the inertial frame and the accelerated frame using two local light sources. In effect, one source should be stay at rest in the accelerated frame (Fig. 32) while the second should be at rest in the inertial frame (Fig. 33). In other words, the reference electromagnetic axis in each frame is formed by an individual light beam. Of course, it is possible to find the space description on a more practical level than that of light beams. For instance, the reference axis in the earth-based frame may be formed by the gravitation field vector (e.g. a standard direction of plumb-line). The equivalence of all local physical frames of reference underlies the theory of relativity, so when the aberration angle emerges, it emerges in the same manner in all local physical reference systems. A comparison with a light clock might help here. When inertial observer looks at the light clock inside the accelerated frame, he sees that clock run slowly. Not only does this particular kind of clock run more slowly, but if the theory of relativity is correct, any other clock, operating on any principle whatsoever, would also appear to run slower.
Now let us return to observations of an accelerated observer, as viewed from an inertial frame. Let us start with non-covariant approach. The aberration of light problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations. According to our approach, when one has a transversely moving screen with a hole in it and a plane wave of light is falling normally on the screen, there is a deviation of the energy transport for light transmitted through a hole, Fig. 9. However, there is another satisfactory way of explaining the effect of aberration of light from the moving source. The explanation consists in using Lorentz coordinatization to describe the uniform translation motion of the light source in the inertial frame. In the Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization. We can interpret manipulations with rule clock structure in the inertial frame simply as a change of the time variable according to the transformation $t\to t+xv/c^{2}$. The overall combination of Galileo transformation and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the inertial frame. Now speed of light is independent of motion of the source in the inertial frame. This means that the velocity of light relative the moving system is $c-v$ in the positive $x$-direction, and $-c-v$ in the negative direction. It is immediately understood that, as a result of motion of accelerated system along the $x$-axis (i.e. along the plane of simultaneity of the reference beam in the accelerated frame) with velocity $v$, there is relativity of simultaneity. This reasoning is in analogy with Einstein’s train-embankment experiment. On the one hand, the wave equation remains invariant with respect to Lorentz transformations.On the other hand, if use the Lorentz coordinatization , we automatically introduce a time transformation $t\to t+xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation phase front in the accelerated frame on the angle $v/c$.
From what has preceded it is easy to see that in the accelerated system, from the point of view of the inertial observer, the electromagnetic axis, assigned by the accelerated observer, will not parallel to the $z$-axis of the inertial observer, Fig. 34. We arrive at the conclusion that in the accelerated frame transmitted light beam propagates along the $z$-axis of the inertial frame of reference with angular displacement $-v/c$ with respect to the reference beam (i.e. with respect to the $z_{n}$-axis) in the accelerated frame.
To continue our discussion of inertial frame view of observations of the non-inertial observer, let us consider the opposite situation. We determine the reference directions in the inertial frame and the accelerated frame using two local light sources. This has interesting consequences. For example, suppose that an observer in the inertial frame performs the direction of transmitted light measurement relative to the reference light beam from the emitter which is at rest with respect to the inertial frame (Fig. 33). The inertial observer would find that the angular displacement is positive and equal to $\theta_{a}=v/c$.
Now we ask about the angular displacement of transmitted light with respect to the reference beam inside the accelerated frame. On the one hand, we know that angular displacement of the transmitted light with respect to the reference light beam from the emitter which is at rest with respect to the accelerated frame is $-v/c$. On the other hand, the angular displacement of the inertial reference light beam is $-2v/c$, Fig. 17. We come to the conclusion that the accelerated observer can directly measure the angular displacement of the transmitted light with respect to the reference light beam of the inertial frame. In fact, the relative angular displacement is $\theta_{a}(\mathrm{transmitted~{}beam})-\theta_{a}(\mathrm{inertial~{}reference~{}beam})=-v/c+2v/c=v/c$. This result is consistent with the inertial observer measurements (Fig. 33), as must be.
7.2 The Non-Existence of an Instantaneous 3-Space
We derived the results for observations of the non-inertial observer with the help of Lorentz transformations. Let us assume, according to textbooks, that the line motion of observers is the same in the accelerated frame as in the inertial frame. At first site, if reference axes were orthogonal to the line motion they must be parallel to each other. Since there exist the angular displacement $v/c$, the situation seems paradoxical.
Figure 35: The lab observer sees the rotation of the moving axes with respect to the lab frame axes, which is closely associated with the Wigner rotation.
First of all, as noted above, the key idea in resolution of this paradox is the realization of the fact that there is no absolute notion of simultaneity in the theory of special relativity. In the case of the relativity of simultaneity we have a mixture - of positions and time. In other words, in the space measurement of one observer there is mixed a little bit of the time as seen by the other. The point, then, is not the existence of a common ordinary space. The fact that two observers with different trajectories have different 3-spaces is not considered in textbooks.
If we look at the situation carefully we see that the accepted in textbooks assumption of common (x) line motion of observers is based on the belief that observers have common 3-space. This is misconception. Right way would be to use the light beam as the common reference direction. We have already seen that the inertial observer can directly measure the angular displacement between the inertial ($z$) and accelerated ($z_{n}$) electromagnetic reference directions and this result is consistent with the accelerated observer measurement. In other words, the displacement between electromagnetic axes is a physical reality. Let us accept that the electromagnetic $z_{n}$ axis in the inertial frame is the reference light beam in the accelerated frame. Now the lab observer sees the rotation of the moving frame axes with respect to the lab frame axes, Fig. 35. This rotation is a relativistic kinematic effect. One can verify directly that the axes of accelerated frame are actually rotated with respect to the lab frame axes by the angle equal to $\theta_{w}=v/c$ in the first order in $v/c$.
Let us clear up the meaning of the parameter $\theta_{w}$. It can be associated with the angle of the so-called Wigner rotation (see the Section 7.5 for more details). Our results show that we cannot remain with the framework of parallel axes $z_{n},z$ when considering the inertial frame view of light beam observations of the non-inertial observer in perpendicular geometry. This inclination angle $v/c$ can be interpreted as a rotation angle $\theta_{w}=v/c$ of the moving axes with respect to the lab frame axes.
7.3 Acceleration of a Rigid Body and Special Relativity
Now let us look at some further consequences of relativity of simultaneity. The concept of rigid motion, as known from Newtonian mechanics, cannot be taken over unchanged to the special theory of relativity. This follows from the fact that in the space measurement of one observer there is mixing of space and time, as seen by other.
First, let us study the relativistic kinematics of rotation of the moving frame axes with respect to the inertial frame. Consider an emitter-detector setup in the lab inertial frame, Fig. 36. Suppose that the setup is accelerated from the rest with respect to the lab frame up to velocity $v$ along the $x$-axis. The main difference between the relativistic kinematics and the kinematics of Newton is that in the lab frame, the accelerated emitter-detector setup exhibits a shearing motion. This counterintuitive fact is not a paradox. In fact, this shearing contribution adds up to the Wigner rotation. The shearing (3-space geometric) effect is closely associated with the relativity of simultaneity.
Above we demonstrated that aberration of light effect in the accelerated frame, as viewed from the inertial lab frame, may be derived from the Lorentz transformations. In other words, aberration of light effect in the accelerated frame, as viewed from the inertial lab frame, presents a kinematics effect of special theory of relativity, Fig. 37. It is shown that the relativity of simultaneity is responsible for aberrations in the first order of $v/c$. We discovered of the essential asymmetry between the inertial and accelerated frame, namely, an acceleration (with respect to the inertial frame) have an effect on the propagation of light in the accelerated frame. On accelerated system, the accelerated detector does not receive light radiated by the accelerated emitter, Fig. 37. In contrast, the inertial detector at rest continues to receive light radiated by the inertial emitter at rest, Fig. 36 left.
Figure 36: Aberration of light in an inertial frame of reference. The aberration of light effect can be described within the standard special relativity taking advantage of the Wigner rotation theory. When one has a transversely moving emitter there is the deviation of the energy transport for radiated light. This effect is a consequence of the fact that the Doppler effect is responsible for the angular frequency dispersion of the radiated light waves (see Fig. 1). As viewed from the Lorentz lab frame, the coordinate axes of the accelerated frame is rotated through the angle $v/c$ with respect to the coordinate axes of the lab frame. In the first order approximation in $v/c$, the accelerated emitter-detector setup exhibits a shearing motion (during acceleration). This shearing contribution adds up to the Wigner rotation. The wavefront orientation has no exact objective meaning. Because of this, only the wavefront orientation in the absolute time coordinatization for the moving source is presented here for sketch simplification.
Figure 37: Aberration of light in an accelerated frame of reference. In order to predict result of the aberration measurement, the accelerated observer should use the Langevin metric. On accelerated system, the accelerated detector does not receive light radiated by the accelerated emitter.
Figure 38: Aberration of light in an inertial frame of reference as predicted by the conventional theory (see Fig. 1).
Figure 39: Aberration of light in an accelerated frame of reference as predicted by the conventional theory. An acceleration does not spoil the motional symmetry between the accelerated reference frame and inertial reference frame.
Common textbook presentations of the special relativity use the (3+1) approach which deals with hidden assumption that distant clocks are synchronized according to the absolute simultaneity convention. In fact, according to textbooks, when a reference frame at rest is put into motion, then all points in the 3D reference grid have a start to move at same time. Consequently, rigidity in the sense introduced by textbooks cannot be considered as a relativistic kinematics property. The standard presentation of the aberration of light effect is based on the hidden assumption that $(x_{x},y_{n},z_{n})$ axes of the accelerated frame and $(x,y,z)$ axes of the inertial frame are parallel. In other words, it is incorrectly assumed that accelerated and inertial observers have common 3-space, Fig. 38,39. According to the conventional theory, an acceleration does not spoil the motional symmetry between the accelerated reference frame and inertial reference frame. Our result (Fig. 36,37 ) is at odds with the prediction from textbooks. The commonly accepted derivation of the aberration of light effect does not account for the Wigner rotation, which in our case closely associated with the relativity of simultaneity.
7.4 Discussion
Let us return to observations of an inertial observer. Special theory of relativity says that the orthogonal emitter-detector setup exhibits a shearing motion during acceleration, Fig. 36. This is a purely kinematic effect, involving no forces. How can that be? That’s not an easy question, but here is one way of thinking about this.
It is clear that our ordinary intuition are quite wrong when we have deal with two space-like separated events. We see that the accepted in textbooks assumption of rigidity of accelerated frame is based on the belief that simultaneous acceleration of the reference space grid has direct physical meaning. This is misconception. In fact, when we have two distant events, we have deal with the conventionality of distant simultaneity within the time interval $L/c$, where $L$ is the space separation of these two events. The position along the $x$-axis of the accelerated emitter with respect to the accelerated detector has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this position could be ascertained . There is an uncertainty (blurring) of the relative position in the $x$-direction of amount $Lv/c$ (due to the uncertainty in the moments of the acceleration).
There are several points to be made about the above results. Above we discussed that the aberration of light effect in accelerated frames can be described in an inertial frame taking advantage of the Wigner rotation theory. To be complete we must add an analysis of the vertical displacement of the detector. We have considered only the case in which motion is non-relativistic in speed. In other words, we use a first order relativistic approximation. It is easy to see that for the vertical displacement we have $\Delta x\simeq\theta_{w}Lv/c\simeq Lv^{2}/c^{2}$. So special relativity says that there is no difference between the (kinematic) shearing motion and the (Wigner) rotation motion (during acceleration) as $v^{2}/c^{2}\to 0$.
7.5 Aberration of Relativistic Particles
We want now to discuss the phenomenon of the aberration of relativistic particles. We begin by recalling a very interesting (methodological) analogy between the aberration of light effect and the deflection of light in general relativity. Consider a ”point” mass $m$ moving with a velocity $V$ in the gravitation field of a point mass $M$. Both the motion of particles and the path of light in the gravitation field are described by the same equation of motion. The deflection angle for a particle is $\theta_{d}=(1+V^{2}/c^{2})\theta_{c}$, where $\theta_{c}$ is the particle deflection angle according to classical approach. The factor $(1+V^{2}/c^{2})$ is new from relativistic (Schwarzschild) metric, but reduced to familiar Newtonian limit as $V^{2}/c^{2}\to 0$. For light $V/c=1$ and there is an extra factor 2. On the basis of classical theory one gets a deflection of light that is only one-half as big as predicted by general relativity.
Suppose we have a particle source. All particles which come out of the source will have the same velocity $V$. Particles trace paths along the $z$-axis. We consider the case when the source is at rest in the initial inertial frame and the observer (with his measuring devices) is accelerated from rest up to velocity $v$ along the $x$ axis. It is assumed that $v\ll V$. Suppose that an observer in the accelerated frame performs an aberration measurement. How shall we describe the aberration of particle beam from the source which is at rest in the inertial frame? In order to predict the result of the aberration measurement the accelerated observer should use the Langevin metric Eq.(15). The aberration increment is connected with the problem parameters by the relation $\theta_{a}=-[1+(1-1/\gamma_{p})]\theta_{c}$, where $\theta_{c}=v/V$ is the particle deflection angle according to classical approach. Here $\gamma_{p}=1/\sqrt{1-V^{2}/c^{2}}$ is the relativistic factor. The term $[1+(1-1/\gamma_{p})]$ is new from relativistic (Langevin) metric, but reduced to familiar classical limit as $V^{2}/c^{2}\to 0$. For light $V/c=1$ and there is an extra factor 2. A correct solution of this problem in the rotating frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of a relativistic particle propagation. In the Chapter 8 (see Section 8.2) we will find that it is very easy to treat the problem by the use of dynamical line of arguments. The electromagnetic forces which govern the properties of an emitted electron beam must be affected by source acceleration in such a way that they lead to a deviation of the electron transport direction. We can obtain the electromagnetic fields in accelerated frame using Galilean transformation.
We want now to find the aberration of ultrarelativistic particles from stationary and moving sources. Imaging that there are two identical sources. For the sake of simplicity we assume that the relativistic factor of the beam $\gamma_{p}\gg 1$. Let us consider the case when the first source is at rest in an inertial frame and the second source is accelerated from rest up to velocity $v$ along the $x$-axis. Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the particle transport measurement. Now we must be careful about the initial direction of the velocity of a particle beam. As example, we consider the case in which particles are initially directed along the $z$-axis. Then how does the particle beam from the moving source looks? The inertial observer would find that angular displacement is equal to $\theta_{a}=v/c$. This is example what called aberration of relativistic particles and it is well known.
Figure 40: The aberration of ultrarelativistic ($\gamma_{p}\gg 1)$ particles from stationary and moving sources. If the particle source is at rest relative to the fixed stars and the observer with measuring devices started from rest to uniform motion with velocity $v\ll c$, then the apparent angular position of the source seen in the accelerated frame would jump by angle $\theta_{a}=-2v/c$. This situation is not symmetrical with respect to the change of the reference frames. If observer is at rest relative to the fixed stars and the source started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the source seen in the inertial frame would jump by angle $\theta_{a}=v/c$
Figure 41: Reciprocity in the theory of the aberration of particles. One must use the same beam directing to demonstrate the reciprocity. After the redirection procedure of the accelerated source the inertial observer would find that the angular displacement is $\theta_{a}=2v/c$ (for the sake of simplicity it is assumed that $\gamma_{p}\gg 1$).
The difference between measurements made in both reference frames is very interesting. According to the asymmetry between the inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of particles. Namely, if the source is at rest relative to the fixed stars and the observer with measuring devices started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the emitter seen in the accelerated frame would jump by angle $\theta_{a}=-2v/c$. This situation is not symmetrical with respect to the change of the reference frames, Fig. 40. It could be said that the crossed term in metric Eq.(15) generates anisotropy in the accelerated frame that is responsible for the difference between aberration increments.
When the accelerated system starts moving with constant velocity the procedure of beam redirection can be performed. Let us describe what happens when accelerated observer performed the redirection of the accelerated source. We consider the case in which finally the particles trace paths along the $z_{n}$-axis. After this redirection procedure the inertial observer would find that angular displacement is $\theta_{a}=2v/c$. Suppose that the accelerated observer also performs an aberration measurement. Fig. 41 shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-2v/c$.
Above in this chapter we demonstrated that aberration of light effect in the accelerated frame, as viewed from the inertial lab frame, presents a well known kinematic effect of special theory of relativity. What about particles? Can we not look at the aberration of particles in the accelerated frame in the same way? As we shall see, the aberration of particles effect in non-inertial frames of reference can be described in an inertial frame within standard special relativity taking advantage of the Wigner rotation theory.
The Wigner rotation is a relativistic kinematic effect, which consists in that a coordinate axes of a non-inertial reference frame is rotating about the axes of a (Lorentz) initial inertial frame. What is the formula for Wigner rotation? It turns out that this is complicated, and it takes a great deal of study and sophistication to appreciate it. (We shall do this in the Chapter 17). We only write it down to give the impression. The expression for the rotation angle in the inertial frame can be presented in vector form $\vec{\theta_{w}}=(1-1/\gamma_{p})\vec{V}\times\vec{v}/V^{2}$, where $\vec{V}$ is the vector of particle velocity in the initial Lorentz frame before acceleration, $\vec{v}$ is the vector of small velocity change ($v\ll V$) due to acceleration, $\theta_{w}$ is the Wigner rotation angle of the spatial coordinate axes of the system comoving with a particle relative to Lorentz lab frame. We shall discuss this expression further in the Chapter 17.
To find the rotation magnitude in the stated problem, we introduce a composition of boosts. Let $S$ be the initial inertial frame of reference, $S_{n}$ an accelerated frame with velocity $\vec{v}$ relative to the inertial frame, and $S^{\prime}$ is a particle comoving frame which moves (after the redirection of the accelerated source) relative to the $S_{n}$ with velocity $\vec{V}$ along the $z_{n}$-axis. Two sequential boosts from the inertial frame $S$ to $S_{n}$ and then to $S^{\prime}$ are equivalent to the boost from $S$ to $S^{\prime}$ and the subsequent rotation. The comoving frame $S^{\prime}$ is rotated through the angle $\theta_{w}$ with respect to the inertial frame $S$. Now we must be careful about direction of rotation. There is a good mnemonic rule. The rule says that the direction of the Wigner rotation in the Lorentz lab frame is the same as the direction of the velocity rotation in the Lorentz lab frame. We can easily understand that the coordinate axes of the particle comoving frame $(x^{\prime},y^{\prime},z^{\prime})$ will be parallel to the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$. As viewed from the inertial frame, the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$ is also rotated through the angle $\theta_{w}$ with respect to the coordinate axes of the inertial frame $(x,y,z)$, Fig. 41. Our results show that we cannot remain with the framework of parallel axes $z_{n},z$ when considering the inertial frame view of particle beam observations of the accelerated (up to the velocity $v$) observer in perpendicular geometry. For light $V/c\to 1$ and the axes of accelerated frame are then rotated with respect to the inertial frame axes by angle $\theta_{w}=v/c$.
8 Earth-Based Setups to Detect the Aberration Phenomena
8.1 Potential of Earth-Based Electron Microscopes
Above we discussed the new relativistic particle kinematics for rotating frame of references. Since we live on rotating earth frame, difference in relativistic kinematics between rotating and non-rotating frames of references is of practical as well as theoretical significance. We propose an aberration of particle experiment using an earth-based particle source. Experiments can be explained on the basis of special relativity. A correct solution of this problem in the earth-based frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of a relativistic particle beam propagation.
We propose an experiment using commercial 200-kV scanning transmission electron microscope (STEM) as a particle source. One could use a vertically oriented optical setup. We suppose that the electron beam is imaged by a lens to a spot which lies in the image plane. Measurement of the spot shift with respect to the optical axis (which is parallel to the zenith-nadir axis) is equivalent to the determination of the angular displacement. Due to the high stability, the aberration increment could be observed (1).
Consider the inertial sun-based reference system. In this system there is the earth which rotates around the sun with orbital velocity $v$. Developing into powers of $v/c$ we can classify effects for velocities $v\ll c$ as of the first order, of the second order or of higher orders. The essentially relativistic effects are of the second (or higher) orders. Clearly, in the case of the rotating (around the sun) earth-based frame, we consider the small expansion parameter $v/c\simeq 10^{-4}$ neglecting terms of order of $v^{2}/c^{2}$.
Suppose that an earth-based observer performs an aberration of particle measurement. The aberration angle varies with one-year temporal period. An approximate formula to express the aberration angle can be found by using Langevin metric in the earth-based frame or by using Wigner rotation theory in the sun-based frame. In the aberration of particles, the aberration angle is the apparent angular deviation of the position of the particle source relative to the real location of the source. The reference axis in the earth-based frame can be formed by plumb line, which is the most fundamental local earth-based coordinate system.
When measured with an earth-based particle source, the aberration angle is $\vec{\theta}_{a}=-\vec{\theta}_{w}=-(1-1/\gamma_{p})\vec{V}\times\vec{v}/V^{2}$, where $v$ is the orbital velocity of the earth relative to the sun, $V$ is the particle velocity, and $\gamma_{p}=1/\sqrt{1-V^{2}/c^{2}}$ is the particle relativistic factor. This means that the angle of Wigner rotation depends on $V^{2}/c^{2}$. It should be note that presented formula can be considered only as a first order approximation in $v/c$. In the theory of (infinitesimal) Wigner rotation we must also consider the relativistic parameter $V/c$. If a particle motion velocity is non-relativistic, the binomial expansion yields $\vec{\theta}_{a}=-[V^{2}/(2c^{2})]\vec{V}\times\vec{v}/V^{2}$.
Let us calculate the parameters for the STEM setup. The Wigner rotation angle parameter (constant of aberration) defines the scale of the aberration increment, and is equal to $\theta_{w}=(1-1/1/\gamma_{p})v/V\simeq 40\mu\mathrm{rad}$. Here $\gamma_{p}\simeq 1.39$, $V/c=0.7$, $v/c\simeq 100\mu\mathrm{rad}$. Earth moves around the sun, with a consequent change of the direction of $\vec{v}$. Therefore the angle of aberration changes. The total change of the order of $40\mu\mathrm{rad}$. It is oscillatory with a period of one year.
Figure 42: Definitions related observer’s position on the earth.
A specific aspect of our case study needs further investigations. The proposed method of measuring the angle of aberration involves the use of earth-based sources and have a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. Obviously, it is important that observation should be recorded in the shortest possible time. In principle, records could be taken over a period of one day. The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth rotation axis (2). In practice, the measurement procedure is complicated by a number of factors, Fig. 42. A corrections has to be made for the observer position on the earth’s surface and the obliquity of the ecliptic $\delta=23^{o}$. The image appears to move in an ellipse. The shape of the aberration ellipse obviously depends on the observer’s latitude $\phi$. For an observer on the equator the ellipse degenerates into a line segment and for an observer at the pole of the earth, the ellipse is a circle. At the value $l=2$ mm of the focal length (3), the major axis of the aberration ellipse of the order of 80 ($\cos\delta)$ nm. The value of minor axis depends on the observer’s position and is given by 80 ($\cos\delta\sin\phi))$ nm. When the latitude is $\phi=35^{o}$, the value of the scale of image shift per hour depends on the local time and various from a value about 10 nm/60 min to about 20 nm/60 min through the day (4).
8.2 Explanation of the Aberration of Electrons on the Basis of Electrodynamics
A large number of incorrect expression for the Wigner rotation (it is often called the ”Thomas rotation”) can be found in the literature. For those who have already received knowledge about the Wigner rotation from the well-known textbooks, it may be expedient to begin its aberration of particle (theoretical and experimental) development with the microscopic approach. The point is that one can compute any relativistic quantity directly from the underlying theories of matter without involving relativity at all. Wigner rotation associated with the transformation from the inertial (sun-based) coordinate system to the rotation (earth-base) coordinate system may be regarded as a result of the action of certain force. This problem has no relation to the problem of the ether and is already formulated within the framework of special relativity. The electromagnetic forces which govern the properties of an emitted electron beam must be affected by (earth-based) source acceleration around the sun in a such a way that they lead to a deviation of the electron transport direction. In fact, there is a machinery behind the electron momentum $\vec{p}$ changes when the source and observer (with his measuring instruments) are at rest in the rotating frame. Its origin is explained in the framework of the electrodynamics theory.
We can obtain the electrodynamics equations in rotating reference frame using the Galilean transformation. When we consider the electrodynamics equations under Galilean transformation, there is a question: what is the transformation law for the electromagnetic fields $\vec{E}$ and $\vec{B}$? The fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame where the particle source is at rest are not the same as those $\vec{E}$ and $\vec{B}$ in the inertial frame before the active Galilean boost.
We need to find the relation between $\vec{E}_{n}$ and $\vec{B}_{n}$ and $\vec{E}$ and $\vec{B}$. To do this, we transform coordinates $(t,x)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$, using a Galilean transformation: we substitute $x_{n}=x-vt$, while leaving time unchanged $t_{n}=t$ into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain Langevin metric $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. We can obtain the electrodynamics equations in rotating frame using the Galilean transformation (with velocity $-v$) of the Maxwell’s equations. With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. This expression describe first order (in $v/c$) effects only. Here $\vec{v}$ is the velocity vector of the earth in the sun-based frame.
[…]
This formula correspond to the result we demonstrated for observation of the non-inertial observer taking advantage of the Wigner rotation theory. The theory of special relativity shows that the aberration of particles effect in the rotating earth-based frame, as viewed from the inertial sun-based frame (in the Lorentz coordinatization), presents a kinematic effect of special theory of relativity. In this case, when viewed from the sun-based frame, the aberration image shift in the earth-based frame is regulated by the Wigner rotation. According to the Wigner rotation theory, two observers with different trajectories have different 3-spaces. In contrast, the accepted in previous literature incorrect assumption that an earth-based observer and an inertial (e.g. sun-based) observer have common 3-space has always been considered obvious. It is important to stress at this point that the dynamical line of arguments discussed here explains what the Wigner rotation physically means in the rotating frame. The relativistic kinematics effects in the rotating frame, as viewed from the inertial observer, is only an interpretation of the behavior of the electromagnetic fields.
8.3 Thought Experiment using Point-Like Source at Optical Frequencies
It is generally believed that the theory of relativity appears to conform to the phenomenon of stellar aberration claiming it is a consequence of the motion of observers relative to light sources. In particular, the conventional method used to explain the aberration phenomena is based on the belief that there is no aberration for the starlight radiated from the star moving with the same velocity as the earth. However, aberration shift, as inferred from astronomical observations, behaves asymmetrically. It should be stressed that it is the telescope and not the star that must change its velocity (relative to the fixed stars) to cause aberration.
We want now to discuss a simple scaling model for the stellar aberration, Fig. 43. Let us consider the scale transformation of the source linear dimension and the distance between the source and observer. We will obtain a condition for optical similarity between the aberration of light from a distant star and from the earth-based source. This scale transformation may help in the initial design of the earth-based setup using incandescent lamp as a point-like source instead of a star.
Figure 43: Scaling model for the stellar aberration. Earth-based setup using incandescent lamp. We know (without using our knowledge of special relativity) what would happen from the astronomical observations. Without looking at anything external to the earth, one could determine the speed of the earth around the sun by means of aberration of light measurements.
The peculiarity of point-like sources is that radiation emitted at one instant form a sphere around the source and the measuring instrument always influences the measured radiation. Source field diffraction can be divided into categories - the Fresnel (near-zone) diffraction and Fraunhofer ( or telescopic) diffraction. If the Fraunhofer (or equivalently, far zone) approximation $z>D^{2}/\lambda$ is satisfied, then quadratic phase factor is approximately unity over the detection aperture. Here $D$ is the aperture width, $\lambda$ is the wavelength of the radiation, and $z$ is the distance between source and observer. In the far zone a point source produces in front of pupil detection effectively a plane wave. The plane wave fronts of starlight entering the telescope are imaged by the lens to a diffraction spot which lies in the focal plane. It worth noting that we consider the divergence of the transmitted radiation, $\lambda/D$, that is relatively small compared to the aberration angle, $\theta_{a}=v/c$.
An incandescent lamp is a completely incoherent source. The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem (5). In our case of interest, imaging system is situated in the far-zone of the source (6). Under these circumstances, the Van Cittert-Zernike theorem takes its simplest form. It is the source linear dimension $d$ (source diameter) that determines the coherent area of the observed radiation $(z\lambda/d)^{2}$. The condition for neglecting the transverse size of the source, $d$, formulated as the requirement that the change of the correlation function along the aperture be less than unity, that is $d\ll D$.
Let us estimate the parameters for the earth-based setup using incandescent lamp. We want to describe some thought - but possible - experiment. At optical frequencies, the conditions required for validity of the Fraunhofer approximation can be sever ones. For example, at a wavelength of $\lambda=0.3\mu$m and an aperture width of $D=3$ cm, the observation distance must satisfy $z>3000$m. Obviously, the required conditions are met in a number of important problems. No one has ever done such experiment, but we know (without using our knowledge of special relativity) what would happen from the astronomical observations. The purpose of this thought experiment is to suggest that without looking at anything external to the earth, one could determine the speed of the earth around the sun by means of aberration of light measurements.
The aberration of light problem in the accelerated frame demonstrates the essential asymmetry between the non-inertial and inertial observers. There is a certain degree of analogy between the aberration of light and the Sagnac effect. Clearly, it is possible to detect a rotating velocity without looking outside using Sagnac phenomenon. A close look at the physics of these two subjects shows things which are common to these phenomena: In both situations these are experiments in the first order in $v/c$. Indeed, the aberration of light is the geometric effect in our ordinary space which is (similar to the Sagnac phenomenon) closely associated with the first order (crossed) term in the Langevin metric.
8.4 Simple Explanation of the Aberration of Light Effect in Rotating Frame
Above (see Chapter 5) we considered the aberration of light problem in the accelerated frame. The aberration of light effect is not easy to calculate in frame of reference attending rotation. In this case, we used a metric tensor to obtain the electrodynamics equations in rotating reference frame. The presented approach uses the Fourier transform method. As one of the consequences of the crossed term in the wave equation, we found the group velocity that is responsible for the aberration. Indeed, what is usually considered as an aberration is, in fact, a deviation of the energy transport direction.
In the Section 8.2 we presented a very simple explanation of the aberration of relativistic particles in the rotating frame on the basis of electrodynamics. To find the aberration increment, we simply used Galilean transformation law for the electric and magnetic fields. Can we not look at the aberration of light effect in the same way?
We have an expression for the energy flow vector of the electromagnetic field. This vector, $\vec{S}=c\vec{E}\times\vec{B}/(4\pi)$, is called ”Poynting’s vector”. It tell us the rate at which the field energy moves around in space. In order to detect the aberration of light effect inside the accelerated frame, it is obvious that some coordinate system with reference direction is needed. We determine the reference directions perpendicular to the line motion in the inertial frame and in the accelerated frame. The motion of the aberrated light beam are assumed for simplicity, to lie in the same plane and the angular position of the aberrated beam is described by one angle (as illustrated in Fig. 32, Fig. 33 ). The analysis is much simplified if we treat separately the case of a radiated beam with its $\vec{E}$-vector parallel to the ”plane of incidence” (that is $xz$-plane) and the case of a radiated beam with the $\vec{E}$-vector perpendicular to the $xz$-plane. We will carry through the analysis for an incoming beam polarized perpendicular to the plane of incidence, but the principle is the same for both. So we take that $\vec{E}$, has only $y$- component. For a magnetic field we get $\vec{B}=\vec{e}_{z}\times\vec{E}$, where $\vec{e}_{z}$ is the direction of Poynting vector $\vec{S}$ in the inertial frame.
In the previous chapters we demonstrated that we must match the (4D) coordinates in the accelerational and inertial parts of the trajectory. For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the Galilean boost. We can obtain the electrodynamics equations in rotating frame using the Galilean transformation (with velocity $-v$) of the Maxwell’s equations. The fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame where the emitter is at rest are not the same as those $\vec{E}$ and $\vec{B}$ in the inertial frame before the active Galilean boost. With the Galilean transformation, the field transformation is $\vec{E}_{n}=\vec{E}-\vec{v}\times\vec{B}/c$, $\vec{B}_{n}=\vec{B}-\vec{v}\times\vec{E}/c$. This expression describe first order (in $v/c$) effects only. Here $\vec{v}=v\vec{e}_{x}$ is the velocity vector of the earth in the sun-based frame. If the $\vec{E}$-vector perpendicular to the $xz$-plane, the magnetic field in the inertial frame has only a $x$-component and we have immediately $\vec{E}_{n}=\vec{E}=\vec{e}_{y}E$, $\vec{B}_{n}=(\vec{e}_{z}\times\vec{e}_{y})E-(\vec{e}_{x}\times\vec{e}_{y})vE/c=-\vec{e}_{x}E-\vec{e}_{z}vE/c$. Therefore there is a Poynting vector, $c\vec{E}_{n}\times\vec{B}_{n}/(4\pi)=[-(\vec{e}_{y}\times\vec{e}_{x})-(\vec{e}_{y}\times\vec{e}_{z})v/c]cE^{2}/(4\pi)=[\vec{e}_{z}-\vec{e}_{x}(v/c)]cE^{2}/(4\pi)$. Then, the radiated beam is propagated at the angle $-v/c$ with respect to the $z_{n}$-axis yielding the phenomenon of the aberration of light in the accelerated frame.
We have found that we get the same result whether we analyze the energy transport direction using electrodynamics equations in the rotating frame, or using Galilean transformation law for electromagnetic fields. In the first instance, the equations state that transmitted light beam with finite transverse size moves along $x$-direction with group velocity $d\omega/dk_{x}=-v$, in the second, the effect of light aberration is understood as a change in the direction of Poynting vector. The two approaches give the same result. In this section we demonstrate that the direction of group velocity and the direction of Poynting vector are actually two sides of the same coin.
When the observer are at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be modified. One can easily show this by transforming coordinates $(t,x,y,z)$ that would be coordinates of an inertial (sun-based) observer moving with velocity $-v$ with respect to the earth-based observer. From the Galilean transformation $x_{n}=x-vt,~{}y_{n}=y,~{}z_{n}=z,~{}t_{n}=t$, after partial differentiation, one obtains $\partial/{\partial t_{n}}=\partial/{\partial t}+v\partial/{\partial x}$, $\partial/{\partial x_{n}}=\partial/{\partial x}$. Acceleration has an effect on the field equations. In fact, the wave equation transforms into Eq. (16). In any measurement of the angular displacement of transmitted (through aperture) light beam, one will have to consider a finite radiation beam size. We have, then, a general method for finding the group velocity of transmitted light beam. We can solve the Eq. (16) by using Fourier transform method. The new terms that have to be put into the wave equation due to use Galilean transformation leads to prediction of the group velocity, $-v$, along the $x$-axis.
Our problem now is to work out the magnetic field of the transmitted light beam in terms of those of the incident plane wave. How can we do that? The magnetic field $\vec{B}_{n}$ satisfies Maxwell’s equation ×$\vec{\nabla}\times\vec{B}_{n}=(\partial/{\partial t}+v\partial/{\partial x})\vec{E}_{n}/c$, in the rotating frame. We are very specific and wright out explicitly the new component: $-\partial B_{z}/{\partial x}=(v/c)\partial E_{y}/{\partial x}$. If we integrate this equation with respect to $x$ across transmitted light beam, we conclude that $\vec{B}_{n}=-\vec{e}_{x}E-\vec{e}_{z}vE/c$. This last result is just what we got by a field transformation argument.
This is a good point to make a general remark about the physical reality of the energy flow (Poynting) vector. In any physically realizable aberration measurement, one will have to consider a finite radiation beam size. It is always hiddenly assumed that the detector for the direction of the radiation is an energy propagation detector and the size of the detector aperture is sufficiently large compared with the radiation beam size. The direction of the energy transport (and, hence, the direction of the Poynting vector) has an exact objective meaning i.e. convention-invariant. However, there is a common mistake made in electrodynamics connected with the energy transport direction in the case of a plane wave. It is impossible to know the energy transport direction when one has deal with plane wave.
8.5 Working Way of the Setup Using Incoherent Source at Optical Frequencies
It should be note that Fraunhofer diffraction patterns can be observed at distances much closer than implied by equation $z>D^{2}/\lambda$ provided the aperture is illuminated by a spherical wave converging towards the observer, or if a positive lens is properly situated between the observer and the aperture.
A working way of the earth-based optical setup to detect the aberration of light is schematically shown in Fig. 44. This is the well-known two-lens image formation scheme. This scheme allows for magnification by changing the focal distance of the second lens but for simplicity, in the following we will assume that the two focal distances are the same (i.e. we consider 1:1 imaging). This two-lens setup is usually employed for image-processing purposes, as it can be better used for image modification compared to the single-lens system. Given the two-lens setup discussed above, we consider the relatively simple problem of characterization of point source radiation in the image plane. We assume that the two lenses in Fig. 44 are identical.
To perform an aberration of light measurement, the source linear dimension $d$ should be radically decreased. We assume that the order of magnitude of the dimensions of the ”point” source is about $\lambda$, where $\lambda$ is the optical radiation wavelength. Within an elementary source volume ($\lambda^{3}$) there is an enormous number of atoms. Semi-classical theory is treating these atoms as coherent, interacting radiating dipoles. The induced macroscopic dipole moment in an elementary source leads to the classical electromagnetic radiation. In the optics for such ”point” source emission, the fields are described classically at the level of Maxwell’s equations and the emitting medium is treated (as a ensemble of atoms) by quantum mechanics. The atomic polarization induced by the radiation field appears as a driving term in Maxwell’s equations and sustains the oscillations. With the help of modern lithography it is not difficult to produce an unresolved point source at optical wavelengths and let sufficiently bright radiation from it (7).
Consider a diffracting aperture that is circular, and let the radius of the aperture be $a$. The physical properties of the lens can be combined in a single number $f$ called the focal length. We assume that the image produced by a diffraction-limited optical system (i.e. a system that is free from aberrations). Once the wavelength is fixed, the resolution only depends on the numerical aperture $NA=a/f\ll 1$ of the system. The response of the system at point $(x_{i},y_{i})$ of the image plane to a $\delta$ function input at coordinates $(x_{o},y_{o})$of the object plane is called the point spread function of the system. In our case of interest, one obtains the following amplitude point spread function: $A(x_{i},y_{i})=J_{1}(\alpha)/\alpha$, which is the diffraction pattern of a circular aperture, where $\alpha=2\pi ar_{i}/(f\lambda)$, $x_{0}=0,y_{0}=0$, $r_{i}=(x_{i}^{2}+y_{i}^{2})^{1/2}$. The first zero of the Airy’s pattern is at $r_{i}=0.6\lambda/NA$.
Let us estimate the parameters for the optical setup shown in Fig.44. Suppose $a=1.5$ cm and $\lambda=0.3\mu$m. We will use $NA=0.1$. The resolution analysis is reduced to the theoretical framework of standard imaging theory. One can take advantage of well-known resolution criteria like Rayleigh criteria i.e. $0.6\lambda/NA=2\mu$m. One could use a vertically oriented optical setup. Constant of aberration defines the scale of the aberration increment, and is equal to $v/c=100\mu$rad. Measurement of the spot shift with respect to the optical axis (which is parallel to the nadir direction) is equivalent to the determination of the angular displacement. As already pointed out, the proposed method of measuring the angle of aberration involves the use of earth-based sources and have a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. In principle, records could be taken over a period of one day. The image appears to move in an ellipse. The major axis of the ellipse is about 14 $\mu$m. The value of minor axis depends on the observer’s position and is given by $14(\sin\phi)$ $\mu$m (8).
Figure 44: A working way of the earth-based optical setup to detect the aberration of light. Two-lens image formation setup using the earth-based incoherent source at optical frequency.
We derived our results under the assumption that the order of magnitude of the dimensions of the source is about $\lambda$. It should be note that the resolution of proposed optical techniques is not limited by size of incoherent source. Radiation field generated by an incoherent source can be seen as a linear superposition of fields of individual elementary point sources. The image of an elementary point source is a point spread function. Optical system using incoherent light could be completely described in therm of convolution integral relating the object distribution to the image distribution. In such a description, the optical system is characterized by its impulse response, i.e. the point image. The rounding or softening of the source’s edges is characteristic of the convolution (”smoothing”) integral.
It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single point source characteristics like oscillatory image shift with a period of one day. This argument gives reason why our technique can be extended to allow for the reconstruction of the angle of aberration from a measurement of image shift of incoherent source. The edge object has become useful in aberration shift detection. An alternative application of a ”point” source is an incoherent source with (physically) sharp edges. The edge width of the source is strictly related to the practical resolution achievable. For our a particular example, the edge width should be about $1\mu$m .
8.6 Absent of the Aberration of Light from the Earth-based Laser Source
Let us make some comments on the aberration of light experiments using an earth-based source. In regard to light aberration one should differentiate between that from the incoherent source and that from the laser source. One could naively expect that the rotation of the earth around the sun produces aberration in an amount large enough to be taken into account in precise observation work using laser as a earth-based light source. Let us stress that the aberration of light cannot be measured using laser.
In order to understand this phenomenon, we consider the simple case when a laser resonator is equipped with plane mirrors. For oscillation to occur, the total loss in power due to diffraction and reflection loss must be less than the power gained by travel through the active medium. Thus diffraction loss is expected to be an important factor, both in determining the start-oscillation condition, and in determining the distribution of energy in the interferometer during oscillation. The purpose of our study is to investigate the effect of diffraction on the electromagnetic field in a Fabry-Perot interferometer in free space. The conclusion can be applied equally well to gaseous lasers provided the interferometer immersed in the active medium.
The electric field $\vec{E}$ of an electromagnetic wave in a passive resonator satisfies the equation $\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. For simplicity, we shall assume that the field in the resonator to be circularly polarized. Such an assumption does not violate the generality of the analysis because the polarization degeneracy takes place in the system under study. The electric field vector in the plane resonator is presented as an oscillation superposition with different longitudinal wave numbers, $E_{x}+iE_{y}=\sum_{m}\widetilde{E}_{m}(x,y,t)\exp(-i\omega_{m}t)\sin(K_{m}z+\delta)$, where $K_{m}=m\pi/l-i/(nl)$, $\omega_{m}=m\pi c/l$, $\delta=i/n$, $n$ is the (complex) refractive index of the mirror matter ($|n|\gg 1)$, $l$ is the distance between the mirrors, and $m$ is an integer number ($m\gg 1$). It is evident that this expression satisfies Leontovich’s boundary conduction (9) on the mirror surface when $z=0$ and $z=l$. We assume the field change per one resonator pass to be small i.e. $|\partial\widetilde{E}_{m}/\partial t|l/c\ll|\widetilde{E}_{m}|$, which means that the laser frequency is close to the natural frequencies $\omega_{m}$ of the longitudinal resonator modes. Let us find a solution for the field amplitude $\widetilde{E}_{m}$ in the form $\widetilde{E}_{m}=\Phi(x,y)\exp(\Lambda t)$. Substituting the ansatz in the wave equation we get $\nabla_{\perp}\Phi+2i\omega_{m}[\Lambda/c^{2}+2/(nl)]\Phi=0$.
In a plane resonator the diffraction effects at the edges of the mirrors can be taken into account if one considers the space between the mirrors as a waveguide, and uses diffraction theory at the open end of waveguide. If almost an integer number of half waves is laid along the length between the mirrors then at open resonator end we may put boundary conditions on the amplitude function $\Phi(x,y)$, equivalent by its effect to the open waveguide end action (so-called impedance boundary conditions of resonance type (10)): $\Phi+(1+i)\beta_{0}\sqrt{cl/(4\omega_{m})}\partial\Phi/\partial\zeta=0$, where $\zeta$ is the direction of the normal to the imaginary side surface of the resonator, and the parameter $\beta_{0}=0.824$ (11). The boundary condition do not depend on the polarization which means that polarization degeneracy takes place in the system under study. Therefore, the problem concerning the open plane resonator excitation by means of equivalent boundary condition can be reduced to the more usual classical formulation of the problem of closed resonator excitation.
[…]
The rigorous results of the three-dimensional theory of a plane Fabry-Perot resonator have shown that a radiation in the resonator may be presented by a set of modes. Each mode is characterized by the decrement and an eigenfunction of the field amplitude distribution along the mirror surface. For the eigenmode with transverse wave number $k_{\nu j}$, we obtain the following expression for eigenvalue: $\Lambda=-2c/(nl)-i\mu_{\nu j}^{2}(1-2\Delta)/(2R^{2})$. Due to the exponential dependence on $\mu_{\nu j}^{2}$, only the lower order modes tend to survive in the passive open resonator. We designate a normal mode of a circular plane mirrors as a $TEM_{\nu j}$ mode, with $\nu$ denoting the order of angular variations ( the angular variation are sinusoidal in form) and $j$ denoting the order of radial variation. The lowest order, of $TEM_{00}$ mode, we designate as the dominate mode for circular plane mirrors.
Suppose that an earth-based observer performs an aberration of light measurement using laser as a light source. One could use a vertically oriented optical setup. Above, we demonstrate that acceleration has an effect on the field equations. The electric field $\vec{E}$ of an electromagnetic wave into the resonator satisfies the wave equation transformed into Eq. (16), where coordinates and time are transformed according to a Galilean transformation. At first glance, measurement of the diffraction losses of the dominate mode is equivalent to the determination of the angular displacement.
Figure 45: Absent ot the aberration of light from the earth-based laser source. The electromagnetic wave travels in the resonator forward and back reflecting from the mirrors. The wavefront orientation in the Lorentz coordinatization for the laser source is presented here for sketch simplification.
It should be note, however, that there is no influence of the difference between the Langevin and Minkowski metrics on the parameters of a laser with optical resonator. In our previous discussion of the Langevin metric, we learned that after the Galilean transformation of the wave equation we obtain the complicated anisotropic equation Eq. (16). How shall we solve this electrodynamics equation together with boundary conditions? It is enough to assume the Lorentz covariance of the electrodynamics theory involved in the optical resonator calculations. We can make a mathematical trick for the solution of the electrodynamics field equations with anisotropic terms: in order to eliminate these terms in the transformed field equations, we make a change variables. The overall combination of the Langevin metric and variable changes actually yields the Minkowski metric in the accelerated frame, but in this context this transformation are only to be understood as useful mathematical device, which allow one to solve the electrodynamics problem in the case of the Langevin metric with minimal effort.
We state that the variable changes performed above have no intrinsic meaning - their meaning only being assigned by a convention. It is clear that the convention-independent results of calculations are precisely the same in the new variables. By changing the (four-dimensional) coordinate system, one cannot obtain a physics in which new physical phenomena appear. It is important to remark that a close examination of all experiments inside the uniformly moving accelerated $S_{n}$ frame, however, shows that all optical phenomena in an optical resonator appeared to be independent of the acceleration relative to the fixed stars. The point is that all methods to measure the radiation in an optical resonator, indeed, the standing wave (i.e. round-trip or two beams) measurements.
Let us analyze the aberration of light transmitted from the laser resonator which is at rest in the accelerated frame. When the measuring data is analyzed, the accelerated observer finds that the deviation of the energy transport direction is absent. It could be said that the crossed term in the Langevin metric, which generates aberration, cancels during the (round-trip-to-round-trip) evolution of the radiation in the optical resonator.
We would like to emphasize a very important difference between laser and incoherent source. The radiation of incoherent source clearly depends on an initial conditions and there is influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer. In contrast, when we are dealing with optical resonator the (3-space) boundary conditions define the energy propagation direction.
Now let us consider another problem in optical resonator operation that we should briefly discuss. This has to do with the optical experiments of second order. It is important to remark that a close examination of all experiments inside the uniformly moving accelerated $S_{n}$ frame, however, shows that all optical phenomena in the optical resonator appeared to be independent of the uniform motion relative to the fixed stars.
Imaging that there are two identical laser sources. We set up the lasers such that light is travels between mirrors parallel to the $z$-axis. Let us consider the case when the first source is at rest in an inertial frame and the second source is accelerated from rest up to velocity $v$ along the $x$-axis (i.e. transverse to the direction of optical beam). Suppose that an observer in the inertial frame performs a measurement of the frequency $\omega_{m}=m\pi c/l$ of the $TEM_{00}$ mode. In order to measure the frequency, the observer has to specify frequency standard. Suppose he took an ordinary atom, which had natural frequency $\omega_{o}$ at rest. Let us consider the opposite situation. Suppose that an observer in the accelerated frame performs a measurement of the $TEM_{00}$ mode. It is also assumed that the same atom at rest in the accelerated frame is used as frequency standard. In the case of Langevin metric Eq.(15), the speed of light propagated between the mirrors in the transverse $z_{n}$-axis direction cannot be equal $c$ anymore. Using Eq.(15) we obtain $(dz_{n}/dt_{n})^{2}=c^{2}(1-v^{2}/c^{2})$. Therefore, the frequency of the $TEM_{00}$ mode is $\omega_{m}=m\pi c\sqrt{1-v^{2}/c^{2}}/l$. Because $t_{n}=t$ (note that coordinate clock in the accelerated frame , although fixed in this frame , read the same time as the clocks in the inertial frame) and $\omega_{m}=m\pi c/l$ has the physical meaning of the frequency of the dominate mode at rest in the inertial frame, this implies that the frequency of the $TEM_{00}$ mode at rest in the accelerated frame is red shifted compared to $\omega_{m}$ in the original inertial frame. Taking into account the Langevin metric we obtain the time dilation of a physical clock at rest in the accelerated frame as compared to physical clock in the inertial frame. In the next chapter we shall discuss the asymmetry in the experiments involved the measurement of quantities of the second order in $v/c$.
An observer on the accelerated frame cannot observe the time dilation effect using optical resonator. In order to detect the red shift of the $TEM_{00}$ mode at rest in the accelerated frame, it is obvious that some reference frequency is needed. The point is that the natural frequency of the atom observed in the accelerated frame is $\omega_{0}\sqrt{1-v^{2}/c^{2}}$. We have found that the accelerated observer will measure the same mode frequency. It is not hard to understand this result. It is clear that an observer performs the measurement of standing wave mode of the optical resonator. If we analyze the geometry of the situation, we find that from the standing wave measurements in the optical resonator we can only extract information about two-way speed of light. In other words, the measurement of the two-way speed of light is universal. Although the optical cavity we have been describing seems to be quite different from a simple electromagnetic model for the light-clock of special relativity, the two systems are, of course closely related. The discussion of the behavior of an accelerated light clock is presented in the next chapter.
8.7 Bibliography and Notes
1. The dedicated STEM with high stability was made [39]. It was developed based on a 200-kV commercial instrument. The electron source was installed in the anti-seismic room. Temperature fluctuations in the room were within 0.2 ${}^{o}C$. The magnetic field in the room was less than $0.2mG$ near the source column. To prevent thermal drift, a shield box was placed over the high-voltage generator. The column was covered with rubber sheets to reduce the temperature fluctuations. The image drift measured during a period of 61 minutes was $\simeq 6~{}\mathrm{nm}$. The drift speed and direction during this period were approximately constant. The average drift speed was 0.12 nm/min. It is substantially smaller than that conventional commercial microscope, in which the drift speed is about 2 nm/min.
2. The rotation of the earth on its axis produces aberration in an amount large enough to be taken into account in precise observation work using electron microscope as a particle source. The Wigner rotation angle parameter (constant of diurnal aberration) defines the scale of the aberration increment, and is equal to $\theta_{w}=(1-1/1/\gamma_{p})v_{e}/V\simeq 0.6\mu\mathrm{rad}$. Here $\gamma_{p}\simeq 1.39$, $V/c=0.7$, $v_{e}/c\simeq 1,55\mu\mathrm{rad}$ where $v_{e}=0.45$km/s is the linear velocity of the earth rotation at equator.
3. Electron-optical system of the STEM contains an electron gun and several magnetic lenses, stacked vertically to form a lens column. The illumination system of the instrument comprises the electron gun, together with two condenser lenses that focus the electrons onto the spacemen. Electrons entering the lens column appear to come from a virtual source. The first condenser lens (C1) is a strong magnetic lens, with a focal length $f_{1}$ of typically 2 mm. Using a virtual electron source as its object, C1 produces a real image. Because the lens is located $D_{1}=15$ cm bellow the object (virtual source), its object distance is 15 cm $\gg f_{1}$ and the image distance $\simeq f_{1}$. The second condenser lens (C2) is a weak magnetic lens ($f_{2}\simeq 10$ cm) that provides little magnification but allows the diameter of illumination at the specimen to varied continuously over a wide range. Distance between lens centers typically 10 cm. Spacemen is located $D_{3}=25$ cm bellow the C2.
4. Experimental results (see Fig. 44 p.167 in [39]) shows that the image shift is quite close to the theoretical prediction $\simeq 10~{}\mathrm{nm}$. It is important to note that we used the typical focal length (2 mm) in our estimations. It is also assumed that the second condenser lens provides no magnification ($M\simeq 1$). The details of the optical setup may influence the image shift.
5. The original optical source (such as a star or an incandescent lamp) consists of an extended collection of independent radiators. Accordingly, it is of some special interest to know precisely how mutual intensity propagates away from an incoherent source. The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34].
6. In all cases of practical interest, telescopes are situated in the far zone of the distant stars. If the far zone approximation $z>2d^{2}/\lambda$ is satisfied, then the Van Cittert-Zernike theorem takes its simplest form. For example, the coherence area of the light emitted by the circular incoherent source of diameter $d$ at distance $z$ is $A_{c}=4\lambda^{2}z^{2}/(\pi d^{2})$. The maximal coherent area of light observed on the earth’s surface from a nearest star has diameter of about 10 m. In other words, any star can be considered as a point source and a star image is actually a point spread function in the image plane of a telescope. The star consists of an extended collection of independent ”point” sources. The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$. It is obvious that the star diameter $d$ is much larger than a telescope aperture $D$ and that far zone condition $z>D^{2}/\lambda$ for elementary ”point” source is automatically satisfied. Our scaling model is based on the far zone approximation. We have only to change the distance and the source diameter. With the limitation of the small far zone parameters ($d^{2}/(z\lambda)<1$, $D^{2}/(z\lambda)<1$) assumption, the proposed scaling method gives $d\ll D$.
7. It may be possible to use the masked spatially incoherent UV light source. One could use the light source onto (micron diameter) hole of specially designed mask. The mask should be placed at the distance $l<<d$, where $d$ is the transverse size of the spatially incoherent UV light source.
8. The important question is that of light detection. The small image size form rather sever restriction on the use of the CCD as a detector. In fact, for the case considered numerically image size is near 2 $\mu$m. One possibility to overcome this practical difficulty is the photon detection with an MCP (micro channel plate photo multiplier) detector. It may be possible to use a (commercially available) precision 2D translation stage with a resolution of 0.5 $\mu$m to controlling the motion of a hole in the opaque screen. The aperture diameter in this example may be set to 1 micron or smaller. The screen is fixed horizontally on the translation stage which is placed in the image plane. The MCP-PMT may realize the measurement of (transmitted through a hole) photons. The proposed design has the advantage of device compactness. The height of our proposed optical configuration is about one meter.
9. We have presented an example of introducing an approximate boundary conditions. This approximate boundary conditions were derived by Leontovich. They are valid on the surface of a material having a large value of the modulus of the refractive index. An extended theoretical treatment of Leontovich’s boundary conditions appears in the monograph [40].
10. Analytical method is based on the introduction of particular, complex boundary condition, which are called impedance boundary conditions [41]. Once these boundary conditions for the field are formulated on the virtual side surface of the open plane resonator, the original problem, which is an open resonator excitation, is simplified to the that of a closed resonator.
11. $\beta_{0}$ turns out to be related with one of the most famous mathematical functions, the Riemann zeta function $\zeta$. In fact, it is given by $-\zeta(1/2)/\sqrt{\pi}$, see [41].
9 Special Relativity and the Reciprocal Symmetry
In this chapter we will collect a number of useful facts concerning the space-time measurements made by different observers moving with respect to each other. The situation with measurements in special relativity is not symmetrical with respect to the change of the reference frames. The fact that in the real process of transmission to an accelerated reference frame (i.e in the process of an observer accelerating with respect to the fixed stars) the observer will experience the pseudo-gravitational force, is not considered in textbooks.
9.1 Pseudo-gravitation Field and the Langevin Metric
This section concentrates on the nature of special relativistic effects. Dynamics, based on the field equations, is usually hidden in the language of the relativistic kinematics. The metric structure of space-time is only an interpretation of the behavior of the dynamical matter fields in the view of different observers. It is worth remarking that the absent of a dynamical explanation for the Langevin metric in special relativity has disturbed some physicists. A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer.
Langevin metric associated with the transformation from the inertial coordinate system to the accelerated (with respect to the fixed stars) coordinate system may be regarded as a result of the action of certain force. This problem formulated within the framework of general relativity. Note that it is not correct to say that this is explanation of the effects discussed in the last four chapters. One of the aims in this section is to show that one can describe the non-inertial frames in the pseudo-Euclidean space-time using the language of general relativity.
Suppose that the reference frame $S$ is at rest with respect to the fixed stars. Let $(ct,x,y,z)$ again be Lorentz coordinates in a system of inertia $S$ and a system $S_{n}$ in the frame $S$ is uniformly accelerated from the rest with respect to the system $S$ up to velocity $v$ along the $x$-axis. The inverse Galilean boost $x\to x_{n}+vt$, $t\to t_{n}$ then defines a new system of reference which is the accelerated system $S_{n}$ moving in the direction of the $x$-axis with the velocity $v$ relative to $S$. After the acceleration, the system of inertia $S$ is (reciprocally) moving in the direction of $x_{n}$-axis with the velocity $-v$ relative to $S_{n}$.
When the system $S_{n}$ starts moving with constant velocity it will be inertial frame of reference. At first glance, we have the symmetry (reciprocity) between the inertial frame $S$ and inertial frame $S_{n}$. According to principle of covariance (i.e. principle of reciprocity), accelerated (with respect to the fixed stars) observer should obtain Minkowski metric $ds^{2}=c^{2}dt_{n}^{2}-dx_{n}^{2}$. This metric is reciprocal to the metric of the inertial observer $ds^{2}=c^{2}dt^{2}-dx^{2}$. Where does the asymmetry comes from? Equivalence principle analysis does not need any tricks to account for the asymmetry between an inertial frame and an accelerated (with respect to the fixed stars) frame. A resolution of the asymmetry paradox identifies a pseudo-gravitational potential within the system $S_{n}$ as the agency of asymmetry.
We can treat a uniformly accelerating frame as if it were an inertial frame with the addition of a uniform pseudo-gravitational field. By a ”pseudo-gravitational field”, we mean an apparent field (not a real gravitational field) that acts on all objects proportionally to their mass; by ”uniform” we mean that the force felt by each object is independent of its position. This is basic content of equivalence principle. The principle of equivalence can be applied to solve non-inertial kinematic problems with dynamics methods.
In order to keep the mathematical complexity of the discussion to a minimum, we will describe the effects by working only up to the second order in $v/c$. Suppose that each reference point in the system $S$ has a constant acceleration $g=-|g|$ in the negative direction along the $x_{n}$- axis during the time $T=v/|g|$. In other words, the system $S$ is accelerating (freely falling) in a pseudo-gravitational field. It is now immediately clear that the pseudo-gravitational acceleration is simply equal to the gradient of the scalar potential: $g=-\partial\phi/\partial x_{n}$. A simple calculation gives $x_{n}=gT^{2}/2+x$ at $t=T$. When the system $S$ starts moving with constant velocity the pseudo-gravitational field is zero. For the pseudo-gravitational potential difference between the two frames we get after some calculation $\phi=(\partial\phi/\partial x_{n})gT^{2}/2=-v^{2}/2$. It is assumed that $\phi=0$ at $x_{n}=x$.
The equivalence principle analysis of the accelerated frame does not use any real gravitational field, and so does not use the general relativity. Nevertheless, what general relativity does say about real gravitational fields does hold in the restrictive sense for pseudo-gravitational fields. One thing we need here is that time runs slower as we descend into the potential well of a pseudo force field. We can use that fact to our advantage when analyzing the metric in the non-inertial frame.
The equivalence principle implies that gravity can shift the frequency of an electromagnetic wave, and cause clocks to run slow. The light frequency increases with increase absolute value of the potential of the gravitational field. If a pulse of light emitted at a point where the gravitational potential is $\phi_{1}$, has at that point the frequency $\omega_{1}$, then upon arriving at a point where the potential is $\phi_{2}$, it will have a frequency measured in units of the proper time at that point equal to $\omega_{2}=\omega_{1}[1+(\phi_{1}-\phi_{2})/c^{2}]$.
Frequency is proportional to the inverse of the local proper time rate; the gravitational frequency shift formula can be converted to a time dilation formula. The rate of a coordinate clock thus depends on the gravitational potential at the place where the clock is situated. The relation between coordinate time and physical (proper) time can be written in the form $dt^{(p)}=dt_{n}(1+\phi/c^{2})$. Thus physical time elapses the more slowly the smaller the gravitation potential at a given point in a space, i.e. the lager the absolute value. If of two identical clocks is placed in a gravitation field for some time, the clock which has been in the field will thereafter appear to be slow. For clock at rest in the accelerated frame we have $dt^{(p)}=dt_{n}(1+\phi/c^{2})=dt_{n}[1-v^{2}/(2c^{2})]$. The fact that the inertial observer always turns out to have aged more than the non-inertial one, in the equivalence principle analysis is elementary consequence of the slowly down of the rate of the processes in the pseudo-gravitational potential, which can be shown by means of simple algebraic calculations.
Furthermore it is important to point out that the phenomenon of the relativity o simultaneity can be understood in terms of dynamical consideration. The time-offset relation can be interpreted as the accumulated time difference between two spatially separated clocks because of pseudo-gravity experienced by the accelerated observer. This suggests that the clock at higher gravitational potential (placed along the direction of acceleration) runs faster. For the pseudo-gravitational potential difference between two points along the $x_{n}$-axis we get $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$. When the system $S$ starts moving with constant velocity the gradient of potential in the system $S_{n}$ is zero. For accumulated time difference between two spatially separated clocks we have $t^{(p)}(1)-t^{(p)}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$.
The next thing we must investigate is the origin of the Lorentz deformation. Here we may point out that the Lorentz deformation (changes of the relative position in the $x_{n}$ direction) is closely related to the shift in the moments of the acceleration. With the help of time-offset relation, integrating with respect to the infinitesimal value $dv$ we find $dl_{n}=dx_{n}[1+v^{2}/(2c^{2})]$, which determines the spatial geometry in the accelerated frame. This equation relates the length $dl=dx_{n}$ of the rod in the inertial frame $S$ to the length $dl_{n}$ in the accelerated frame $S_{n}$. Our analysis based on using standard measuring rods in an accelerated frame to measure it geometrical properties shows that coordinate distance $dx_{n}$ has length $dx_{n}[1+v^{2}/(2c^{2})]$. This is a purely kinematic effect, involving no forces. We come to the conclusion that the various relativistic effects are not independent of each other, but that these effects can all be understood through a general equivalence principle.
We have now all quantities we wanted. Let us put them all together into the expression for the new space-time variables. The new independent variables $(x_{n}^{\mathrm{(p)}},t_{n}^{\mathrm{(p)}})$ can be expressed in terms of the old independent variables $(x_{n},t_{n})$: $ct_{n}^{\mathrm{(p)}}=t_{n}[1-v^{2}/(2c^{2})]-(vx_{n}/c^{2})[1+v^{2}/(2c^{2})]$, $x_{n}^{\mathrm{(p)}}=x_{n}[1+v^{2}/(2c^{2})]$. Above we made a simplification in calculating time and space transformations by considering only low velocities. The physical principles that produced the space-time transformation were, however, made clear. The next question is: What is the transformation in the case of an arbitrary velocity? That is easy: the answer is $t_{n}^{\mathrm{(p)}}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x_{n}^{\mathrm{(p)}}=x_{n}/\sqrt{1-v^{2}/c^{2}}$. Quantity $dt_{n}^{\mathrm{(p)}}$ characterizes physical time, which is independent on the choice of coordinate time. Let us note that $t_{n}$ in the expression for $t_{n}^{\mathrm{(p)}}$ is the coordinate time for the accelerated frame. Physical time determines the flow of time in a physical process. The $|dx_{n}^{\mathrm{(p)}}|$ is nothing, but the physical distance between two points of 3-space. In new physical coordinates we obtain interval in the following (Langevin) form $ds^{2}=(cdt_{n}^{\mathrm{(p)}})^{2}$ $(dx_{n}^{\mathrm{(p)}})^{2}$ $=(1-v^{2}/c^{2})c^{2}dt^{2}_{n}-2vdx_{n}dt_{n}-dx_{n}^{2}$.
The previous derivation of the Langevin metric in the accelerated frame from pseudo-gravitation field includes one delicate point. Metric applies to physical laws, not to physical facts. We interpret the Langevin metric to mean that the law of electrodynamics is expressed by equation that have the form Eq.(16) in the accelerated frame. By a physical facts in this context we mean, for example, the aberration of light radiated by a single ”plane wave” emitter in the accelerated frame. The electrodynamics equation needs to be integrated with initial condition for the radiation wavefront. After the inverse Galilean boost, the coordinates transform as $x_{n}=x-vt,t_{n}=t$, so that the wavefront of the emitted light beam is perpendicular to the vertical direction $z_{n}$ after the acceleration, Fig. 15. The Langevin metric, together with this initial condition, describes the aberration of light effect in the accelerated frame.
Although both observers consider themselves as at rest, they both predict that the Maxwell’s equations continue to hold in the inertial frame and the Minkowski metric is not applicable in the accelerated frame. But observers have totally different explanation for this difference of metrics. Inertial observer says that physical processes in the accelerated frame describes by the Langevin metric due to the active Galilean boost. In other words, the analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard pseudo-Euclidean geometry of space-time. Accelerated observer, on the other hand, says that Langevin metric in his frame can be understood in terms of dynamical consideration.
9.2 Equivalence Principle and the Gravitational Physics
We should make one further remark about the equivalence principle (1). The widely held opinion that such a principle is the basis of the general relativity is not entirely correct. The gravitational field is not simply a force that is universal and has the same action on all bodies. According to Einstein’s idea, the gravitational field is associated with the structure of space-time. The first thing we must understand is the Riemann tensor, since it is this tensor that is the gravitational field: if it vanishes (and only in this case) a field does not exist. Actually this argument is already fully sufficient to show that the equivalence principle in the usual formulation does not hold. The equivalence principle is valid for mechanical processes and does not hold for electrodynamics processes. This means that by means of measurements within the a system it is possible to establish whether the system is inertial or falls freely in a homogeneous gravitational field. In a homogeneous gravity field, a free charge is uniformly accelerated with respect to an inertial frame, and therefore the charge will radiate, and radiation will occur in both strong and weak fields. In the uniformly accelerated frame of reference, the free charge is not accelerated with respect to the inertial frame of reference. But in then follows in accordance with electrodynamics that it can never radiate, and therefore the equivalence principle is indeed violated .
With the special and general relativity we can discuss the case of earth-based experiments where pseudo-gravity and the real gravity are inextricably mixed. With this understanding, we next need to describe the gravitational fields near the sun due to the sun mass itself. Let us consider the simplest instance of a transformation from an inertial frame, in which the space-time is Minkowskian, to the earth-based frame. Newton gravity is known to be valid when gravitation fields are to weak to produce velocities near the speed of light. In other words, the gravitational potential energy of a particle is much less than its rest-mass energy. In such situation, general relativity must make the same predictions as Newton gravity. For weak gravitational field the ordinary Newtonian potential $\phi$ completely determines the metric, which has the form $ds^{2}=(1+2\phi/c^{2})c^{2}dt^{2}-(1-2\phi/c^{2})(dx^{2}+dy^{2}+dz^{2})$. The sign of $\phi$ is chosen negative, so that far from the source of Mass $M$, we have $\phi=-GM/r$. Now the condition above that the field be weak means that $|m\phi|<<mc^{2}$, so that $|\phi/c^{2}|<<1$. This metric is really only correct to first order in $\phi/c^{2}$. We say that space-time becomes flat asymptotically at large distances from the source. The Langevin metric for rotating frame could be used for rotating (around the sun) earth: $ds^{2}=(1-v^{2}/c^{2})d(ct)^{2}-2(\vec{v}\cdot d\vec{r})dt-(d\vec{r})^{2}$. Where $\vec{v}$ is the velocity of the earth, and $(d\vec{r})^{2}$ is the square of the coordinate distance. The absence of equivalence of the equations in an accelerated reference system and in an inertial frame with a static gravitational field is obvious, since in the first case the non-diagonal metric coefficients will be functions of the time, whereas in the second case the metric exhibits diagonal properties and does not depend on the time. Therefore, in the first case the charge at rest will radiate, whereas in the second case the charge at rest will not radiate. It is for this reason that one cannot deduce from the observations the validity of the equivalence principle as a physical principle. The relation between gravity and inertia is a subject that needs to be understood more deeply. Now let us return to observations of an earth-based observer. One additional expression for the invariant interval is needed: $ds^{2}=(1+2\phi/c^{2}-v^{2}/c^{2})d(ct)^{2}-2(\vec{v}\cdot d\vec{r})dt-(1-2\phi/c^{2})(d\vec{r})^{2}$. To the order of calculations, this result is a simple superposition of the Langevin metric with the metric in the sun centered inertial frame for the weak external gravitational field of the sun. Both orbit speed and gravitational field dilate time and contract space. The differences are small, however, and the Langevin metric represents an extremely good approximation for the metric in the rotating earth-based reference system.
9.3 Time Dilation
It is generally believed that the special relativity is a reciprocal theory. An excellent illustration of the absence of reciprocity is the time dilation effect. Consider the course of time in two inertial reference frames, one of which, $S$, will be considered to be at rest, while another one, $S_{n}$, will accelerate with respect to the first one up to velocity $v$ along the $x$- axis. As we already discussed, the transformation from the inertial frame to the accelerated frame apparently first used by Langevin. The mathematical idea that are used in the literature of time dilation effect is the smooth tailoring of the metric tensor. The transformation to the accelerated frame, with coordinates denoted by ”n”, has the Galilean form, $t_{n}=t$, $x_{n}=x-vt$. After the Galilean transformation, the metric of non-inertial frame goes to into the Langevin metric of the accelerated inertial frame $S_{n}$ in a continuous manner (see Section 9.6 for more detail).
In terms of these new coordinates the invariant interval given by Eq. (1) is Eq.(15). Note that coordinate clocks on the accelerated frame, although fixed in that frame, read the same time as the clocks in the inertial frame. With the cross-term, $dt_{n}dx_{n}$, clocks in the accelerated frame are not synchronized in the standard way (Einstein synchronization) of sending a light signal back and forth between the clocks.
Suppose we have two identical physical clocks at one and the same point of the inertial reference system $S$. Consider their readings to coincide at the initial moment $t=0$. Let first of these clocks always be at rest in the frame $S$. At moment $t=0$ the second clock accelerates and starts to move with a constant velocity $v$ along the $x$ axis. Now consider the reading of the clocks in the accelerated reference system, where the second clock is always at rest. Frames are not equivalent, since frame $S_{n}$ accelerated with respect to the fixed stars. In the accelerated frame the first clock moves with velocity $dx_{n}/dt_{n}=-v$. Taking into account the Langevin metric Eq.(15), $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we obtain $d\tau=ds/c=dt_{n}$, i.e. the time, shown by the first physical clock will coincide with the time $t=t_{n}$ of coordinate clock. Since the second physical clock is at rest, its reading of its proper time is $d\tau_{n}=\sqrt{1-v^{2}/c^{2}}dt_{n}$. This demonstrates the time dilation of a physical clock at rest in the accelerated frame as compared to physical clocks in the inertial frame. We have found that the slowly down of the accelerated physical clock does not depend on the choice of reference system, in which this effect is measured as it must be.
Now we go on to consider another situation, a very practical one. Suppose that we split the accelerated frame. Now frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ accelerate with respect to the inertial frame $S$, up to the same speed $v$ but move in opposite directions. The problem is complete symmetrical in the inertial frame $S$ and no difference in the time dilation of a physical clock at rest in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ (as compared to physical clock in the inertial frame $S$) can exist. The following simple analysis confirms this conclusion. Suppose we have two physical clocks. Let first of these clocks be at rest in the frame $S_{n}^{(1)}$. Now consider the reading of the second clock in the accelerated system $S_{n}^{(2)}$ where the second clock is at rest. In the accelerated frame $S_{n}^{(1)}$ second clock moves with velocity $dx_{n}/dt_{n}=-2v$. Taking into account the Langevin metric, $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we obtain $d\tau_{2}=ds/c=\sqrt{1-v^{2}/c^{2}}dt_{n}$, i.e. the time, shown by the first physical clock $d\tau_{1}$ coincide with the physical time $d\tau_{2}$ of the second physical clock. The situation seems paradoxical. Now the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ move relative to each other with a non zero speed. We see an interesting thing: the presence of some non zero relative velocity cannot be a cause for the effect of time dilation.
Above we discussed that one can describe the non-inertial frames using the language of general relativity. Time dilation effect associated with the transformation from inertial frame to the accelerated frame may be regarded as a result of the action of certain force. A resolution of the symmetry paradox identifies a pseudo-gravitational field within the systems $S_{n}^{(1)}$ and $S_{n}^{(2)}$ as the agency of time dilation effect. Proper time clocks having the same accelerational histories really do run at same rates and yield identical measurement results when at rest in different inertial systems.
Now we go on to consider another situation. Suppose that frame $S_{n}$ accelerates with respect to the inertial frame $S$ up to the speed $v$. Now frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ accelerate with respect to the frame $S_{n}$ up to the same speed $v$ but move in opposite direction. The problem is not symmetrical in the frame $S_{n}$ and there is a difference in the time dilation in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$. In the frame $S_{n}$ the frame $S_{n}^{(1)}$ moves with velocity $-v$ and we obtain proper time $d\tau_{n}^{(1)}=dt=dt_{n}$. In the frame $S_{n}$ second frame moves with velocity $v$. Taking into account the Langevin metric, $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we obtain $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt_{n}$. Thus we observe time dilation of the second frame with respect to the inertial frame $S$. We have made this analysis from the point of view of an observer at rest in the frame $S_{n}$: we would like to know how it would look to the observer who is at rest in the inertial frame. If we analyze time dilation in the frame $S$ we find that frame $S_{n}^{(2)}$ moves with velocity $dx/dt=2v$. Taking into account the Minkowski metric we obtain $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt$ as it must be. In our calculations we find that mathematics for relating proper times is extremely simple. That is not too surprising. We used Galilean boosts and the Galilean addition of velocities. We shall discuss the time dilation effect for moving reference (atomic) clock further in the Chapter 11.
9.4 Light-Clock. Observations of the Noninertial Observer
9.4.1 Introductory remarks
In order to understand the slowing of the clock in an accelerated system, we have to watch the machinery of the clock and see what happens when it is moving. We shall try to understand the effect in a very simple case. A clock which we shall call ”light-clock” is a rod (meter stick) with a mirror at each end. The tick of time is the reflection of light in the mirror. Let us now consider the parallel and orthogonal to motion orientations of the light clock. The equivalence of all clocks underlines the theory of relativity and we shall show that irrespective of the orientation of the light path with respect to the motion of the light-clock, the clock scores the same time. Two light-clocks are shown in Fig. 46
Figure 46: Parallel and orthogonal light-clock setups.
9.4.2 Parallel Light-Clock. Langevin Metric
Now we have to watch the machinery of the physical clock and see what happens to the accelerated light-clock. We describe the clock based on the observations made by an observer in the same accelerated frame as the clock. In a first step, we set up the light clock such that light travels between mirrors parallel to the direction of fixed stars motion. In the case of Langevin metric, the speed of light radiated by the accelerated emitter cannot be equal $c$ anymore. If $ds$ is the infinitesimal displacement along the world line of a beam of light, then $ds^{2}=0$ and using Eq.(15) we obtain $c^{2}=(dx_{n}/dt_{n}+v)^{2}$. This means that in the accelerated reference system of coordinates $(ct_{n},x_{n})$ the velocity of light radiated by the accelerated emitter parallel to the $x_{n}$-axis, is $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction. Designating the rod (meter stick) length, unknown so far, by $l_{n}$, we obtain the time in which light gets from the left mirror to the right mirror, $t_{1}=l_{n}/(c-v)$, and the time in which light gets from the right mirror to the left mirror, $t_{2}=l_{n}/(c+v)$. Therefore, the time interval between the sending and reception of the light signal is $\Delta t=t_{1}+t_{2}=2l_{n}/[c(1-v^{2}/c^{2})]$.
One must first determine the length of physical rods. For the measurement of distances between fixed points on the accelerated frame we shall use physical measuring rods of the same kind as those used in the inertial frame, but now at rest relative to the accelerated frame. In connection with the process of measuring distances in accelerated frames a problem arises which did not occur in inertial frames. For according to the special relativity, no rigid bodies can exist, since they would provide a means of transmitting signals with velocity greater than velocity of light.
With regard to spatial distances, the interpretive principle is that $\sqrt{-ds^{2}}$ gives the length of an infinitesimal rod whose endpoints are simultaneous according to standard simultaneity in the rod’s rest frame. Taking into account the Langevin metric Eq.(15), we obtain at $dt_{n}=0$ that $dx_{n}=dx$, i.e. the length of physical rod in the inertial frame will coincide with the coordinate distance in the accelerated frame. When we apply this rule to physical rods that are at rest in the accelerated frame, we encounter the complication that $dt_{n}=0$ does not automatically correspond to standard simultaneity in the accelerated frame. The length will surely be different for a physical rod at rest in the accelerated frame compared to a physical rod at rest in the inertial frame.
The spatial geometry in the accelerated frame is characterized by the differential spatial line element on it. It should be note that spatial geometry has a conventional character. However, the proper spatial line element $dl_{n}$ is of special significance, since it defines a coordinate-independent spatial geometry in the accelerated frame. The relation $dl_{n}^{2}=(-g_{\alpha\beta}+g_{0\alpha}g_{0\beta}/g_{00})dx_{\alpha}dx_{\beta}$ gives the connection between the 3-dimensional spatial line element and the metric of the four-dimensional space-time. It can be shown that the element $dl_{n}$ is invariant under a transformation connecting two different coordinate systems inside the same frame of reference. We note that the spatial line element $dl$ obtained by putting $dt_{n}=0$ in Eq.(15), $dl^{2}=dx_{n}^{2}$, is differ from $dl_{n}$.
With the help of Eq.(15), we find $dl_{n}^{2}=dx_{n}^{2}/(1-v^{2}/c^{2})$, which determines the spatial geometry in the accelerated frame. This equation is asymmetric with respect to the lengths $dx_{n}=dl$ and $dl_{n}$, since it relates the physical length $dl$ of the rod in the inertial frame $S$ to the physical length $dl_{n}$ in the accelerated frame $S_{n}$. Our analysis based on using standard measuring rods in an accelerated frame to measure it geometrical properties shows that coordinate distance $dx_{n}$ has physical length $dx_{n}/\sqrt{1-v^{2}/c^{2}}$. With reference to the accelerated frame it is interpreted as an effect of acceleration with respect to the fixed stars, i.e. the geometry of space is non-Euclidean in an accelerated frame.
This shows that the physical meter stick on which the mirrors are mounted undergoes compress on in the direction of fixed stars motion. Thus we have ascertained the asymmetry of meter stick’s length directly from the metric. Therefore, the time interval between the sending and reception of the light signal in the parallel light-clock in the accelerated frame is $\Delta t_{n}=(2l/c)/\sqrt{1-v^{2}/c^{2}}$. Because $t_{n}=t$ and $\Delta t=2l/c$ has the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame.
9.4.3 Orthogonal Light-Clock. Langevin Metric
Let us consider a light-clock accelerated up to velocity $v$ transverse to the direction of optical pulse. In other words, we set up light-clock such that the effect of a meter stick contraction is not present. We describe the orthogonal light-clock based on the observations made by an observer in the same accelerated frame as the light-clock. If $ds$ is the infinitesimal displacement along the world line of a beam of light, then $ds^{2}=0$ and using Eq.(15) we obtain $(dz_{n}/dt_{n})^{2}=c^{2}(1-v^{2}/c^{2})$. Therefore, the time interval between the sending and reception of the light signal is $\Delta t_{n}=2l/[c\sqrt{1-v^{2}/c^{2}}]$. Because $t_{n}=t$ and $\Delta t=2l/c$ has the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame. The theory of relativity shows us that irrespective of the orientation of the light path with respect to the motion of the light-clock, the clock scores the same time.
9.5 Metrics in the Accelerated Frame
Let us follow out the consequences of assuming that the inertial frame has no accelerational history and the (observer) space-time metric at rest has diagonal form. We start with the two emitters in the inertial frame. The first emitter is at rest and the second is moving with velocity $v$. As a result of Galilean boost, we obtained the metric of the accelerated emitter in the inertial frame Eq. (11). In this coordinatization, we describe the non-accelerated light source using diagonal metric, $ds^{2}=c^{2}dt^{2}-dx^{2}$.
Let us now return to our consideration of the accelerated frame. In the accelerated frame we have the reciprocal velocity. Second emitter is at rest and the first is moving with velocity $-v$. Now what about reciprocity in physics? According to the conventional theory, an acceleration does not spoil the motional symmetry between the accelerated frame and the inertial frame and the accelerated observer should obtain diagonal metric for the second emitter at rest in the accelerated frame. The principle of reciprocity implies that the metric Eq. (15) describes the aberration of light effect in the case of the first (moving) emitter in the accelerated frame. The metric of the moving emitter in the accelerated frame is the same as the metric Eq. (11) of the moving emitter in the inertial frame. According to principle of reciprocity, the difference is only that the sign of velocity $v$. The principle of reciprocity states that all inertial frames are equivalent. However, this relativity principle does not hold and the Lorentz covariance of electrodynamics equations does not dictate the Lorentz covariance of the solutions.
The commonly accepted approach to special relativity does not account for the inertial (pseudo-gravitational) force within the accelerated frame as the agency of asymmetry. Space-time geometrical analysis of the accelerated frame (see Section 9.1) leads to the asymmetry (non-reciprocity) between the inertial frame and the accelerated frame. We obtain in the accelerated frame the Langevin metric Eq. (15) of the second emitter. This metric describe the electrodynamics equations in the accelerated frame. We found of the essential asymmetry between inertial and accelerated frames, namely, the Maxwell’s equations are not applicable from the viewpoint of an observer at rest with respect to the accelerated frame. Because we are using absolute time coordinatization, the initial conditions (i.e. radiation wavefront orientations) are identical in both frames.
We are going to discuss the asymmetry between inertial and accelerated frames. In the discussion up to this point, we assumed absolute time coordinatization. The Langevin metric, together with initial condition, describes the aberration of light effect for the second emitter at rest in the accelerated frame. The essential asymmetry is rooted in the electrodynamics equations. However, there is another satisfactory way of explaining the effect of aberration of light in the accelerated frame. The explanation consists in using a metric diagonalization procedure. The new coordinatization in the accelerated frame is interpreted by saying that the accelerated observer should obtain diagonal metric for the second emitter at rest and the (Langevin) metric Eq. (15) for the first (moving) emitter. Now we see that there is a symmetry (reciprocity) in electrodynamics equations between the accelerated and the inertial frame. This is a consequence of the pseudo-Euclidean geometry of space-time. Our impossibility to detect orbital velocity using second order optical (interference) experiments is a result of this metric reciprocity, see Chapter 11 for more details. Now the question is, where asymmetry comes from? Clearly, after the diagonalization, the asymmetry is rooted in the initial conditions. The plane radiation wavefront of both emitters rotates in the accelerated frame on the angle $-v/c$ after metric diagonalization.
The previous derivation of the metrics in the accelerated frame includes one delicate point. According to a reciprocity in metrics, time dilation is symmetric. This contradiction is resolved by noting that metric diagonalization procedure does not preserve continuity of the metric tensor. In the previous chapters we learned that the accelerated frame is connected with the initial inertial frame by the Galilean transformation, but not the Lorentz one. In fact, the Langevin metric is found by matching the accelerated frame and inertial frame metric tensors. We shall discuss the smooth tailoring of the metric tensor further in the next section.
9.6 Mathematical Expansion of Special Relativity onto Acceleration
Before we begin the next topic, we would like to describe a number of mathematical ideas that are used in the literature of time dilation effect. Suppose we have two clocks in the inertial frame $S$ without accelerational history. Let first of these clocks always be at rest in the frame $S$ and the second clock accelerates and start to move with constant velocity $v$ along the $x$ axis. The proper time of the first clock coincides with the coordinate time $d\tau=dt$. Since the motion of both clocks is uniform, frames $S$ and $S_{n}$ will be inertial. According to textbooks, the coordinates and time in this frames should be related by the Lorentz transformation and the metrics of both frames will be diagonal. At first site the problem is complete symmetrical and there is no difference in the metrics in the frames $S$ and $S_{n}$. The situation seems paradoxical.
We know that independent of the chosen coordinatization the proper time of the second accelerated clock is always less than that of the first clock. The time dilation of the accelerated clock is an absolute effect but not relative. It is independent of the time coordinate definition at the inertial part of the trajectory. This is an intrinsic property of space-time geometry. A more careful analysis shows that one must take into account that the metric tensor must be a continuous quantity.
In the previous chapters we demonstrated that many problems in special relativity can be adequately treated only by an approach which uses the non-standard absolute time synchronization (i.e non diagonal metrics). We will keep the mathematical complexity to a minimum and assume that the second clock begins at $t=0$ to move with constant acceleration $a>0$ and does so till $t=T$. Now we would like to discuss an expression for the velocity from the viewpoint of inertial (non accelerated) observer. We know that uniformly accelerated motion must obey the equation $d(\gamma\vec{v})/dt=\vec{f}/m=\vec{a}$. This is readily integrated $\gamma\vec{v}=\vec{a}t$. We will then have $dx/dt=at/\sqrt{1+a^{2}t^{2}/c^{2}}$. We emphasize that $dx(t)/dt$ is the velocity of the second clock as seen by the inertial observer. Under the inverse Galilean transformation $dx=dx_{n}+atdt/\sqrt{1+a^{2}t^{2}/c^{2}}$, $dt_{n}=dt$ the metric of the uniformly accelerating observer becomes (3)
$ds^{2}=c^{2}dt_{n}^{2}(1+a^{2}t_{n}^{2}/c^{2})^{-1}-2at_{n}dt_{n}dx_{n}(1+a^{2}t_{n}^{2}/c^{2})^{-1/2}-dx_{n}^{2}~{}.$ (19)
At $t=0$ this metric naturally coincides with the $S$ frame metric and we have $g_{00}=1$. But here at the inertial part of the trajectory the metric of both frames will be diagonal in the Lorentz coordinatization and so at $t=T$ the metric tensor of the accelerated frame must abruptly change from the value $g_{00}=(1+a^{2}t_{n}^{2}/c^{2})^{-1}$ to the value $g_{00}=1$. This shows that we must match the coordinates in the accelerational and inertial parts of the (second clock) trajectory.
During acceleration we picked the absolute time coordinatization. Therefore, for the metric tensor to be continuous when the reference frame passes from non-inertial motion to inertial one we will make use of the metric Eq. 19. We will therefore require that at $t=T$ the metric of a non inertial frame associated with the second clock would pass continuously into the Langevin metric. For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer $(x_{n},t_{n})$ to the coordinates $(x,t)$ of the inertial observer by the Galilean boost $x_{n}=x-vt$, $t_{n}=t$, where $v=aT/\sqrt{1+a^{2}T^{2}/c^{2}}$. In this case, in the interval of coordinate time $T<t_{n}$ the reference frame of the second clock has the metric $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$. We emphasize that $v$ is the velocity of the second clock as seen by the inertial observer at the second (inertial) stage of motion. It becomes obvious that at $t=t_{n}=T$ the metric of non inertial frame goes into the metric of the (accelerated) inertial frame in a continuous manner.
9.7 Bibliography and Notes
1. For a general discussion of the equivalence principle we suggest reading the paper [42].
2. In the general case, the problem to assigning Lorentz coordinates in an inertial frame is complicated. Let us consider a source, arbitrary accelerating in the inertial frame, and let us analyze its evolution with a Lorentz coordinate system. The permanent rest frame of the source is obviously non inertial. To get around that difficulty, one introduces an infinite sequence of comoving coordinate systems. At each instant, the comoving coordinate system is a Lorentz coordinate system centered on the source and moving with it. As the source velocity changes to its new value at an infinitesimally latter instant, a new Lorentz coordinate system centered on the source and moving with it at the new velocity is used to observe the source. We should make one further remark about this covariant algorithm. An opinion is sometimes expressed that described above algorithm includes a hidden postulate. It seems necessary a dynamical assumption to justify attributing to an accelerated clock the same rate as a clock in inertial motion in relation to which it is momentary at rest. This is, in view of some authors, an extra condition that a clock must satisfy. It is assumed that the effect of the motion on the clock depends only on it instantaneous speed, not its acceleration. This condition often refereed to as a ”clock hypothesis”. To quote Brown [17]: ”If the accelerating forces are small in relation to the internal restorative forces of the clock, then the clock’s proper time will be proportional to the Minkowski distance along its world line. … This condition is often referred to as the clock hypothesis, and its justification, as we have seen, rests on accelerative forces being small in the appropriate sense.” We state that the clock hypothesis does not have status of independent hypothesis is not needed as an independent postulate in the theory of relativity. As discussed above, the transformations within a single inertial frame are simply another parametrization of the observations of the inertial observer. According to such transformations, the observer (and his clocks) motion relative to the fixed stars does not changes. In other words, when we discuss only the observations of an inertial observer the problem of accelerated clock does not exist at all.
3. For a general discussion of the mathematical expansion of special relativity onto acceleration we suggest reading the book [16].
10 Relativistic Measurement
In our approach to physics in general and of the theory of relativity in particular we think it very important always to remember that we are dealing with objective physical quantities measured by the measurements. We illustrate the problem of measurement in the special relativity by typical examples.
10.1 The Notion of Time in Special Relativity Theory
There are many changes in what concepts are important when we go from classical to relativistic theory. We have already discussed some of them in previous chapters. We will show the connection between the classical notion of time and the relativistic time. The theory of relativity show us that the notion of time is not what we would have expected on the basis of our intuitive ideas.
We attempt to give a definition of time on the basis of a principle ”Each physical magnitude is defined by the method it is measured.” Let us see how this definition works. In the case of length we know that there exist a length standard and we know that any body has its particular length. It should be note, however, that we are dealing with objective physical length only in the case of a stationary object. In contrast to this, the length of a moving object (e.g. moving rod) is convention-dependent (i.e. synchronization-dependent) and has no exact objective meaning. Now we are ready to consider the time problem in special relativity theory.
Here is one example which shows in a circumstance that is easy to understand. Suppose we know the law of muon decay in the Lorentz rest frame. When a Lorentz transformation of the decay law is tried, one obtains the prediction that, in the lab frame, the characteristic lifetime of a particle has increased from $\tau_{0}$ to $\gamma\tau_{0}$. We can interpreted this result by saying that after the travel distance $\gamma v\tau_{0}$, the population in the lab frame would reduced to 1/2 of the origin population. The time measured in this case is ”length” which is metrized as described above. That leads us to interesting question. Is the same statement perhaps also true for all cases of time measurements in the framework of special relativity? We claim that the answer to this question is positive.
Our next example has to do with frequency. What about the frequency measurements? Let us suppose that we have a (Fabry-Perot) interferometer. The frequency measured in this case is (standing wave) ”length”. Frequency measurements automatically obey the relationship between the frequency and interference. Another example of the relation between interference and frequency measurements is a grating spectrometer. The frequency measurement in this case is a position of the light spot along the dispersion direction. There is no specific, important physical difference between diffraction grating and interferometer. Grating spectrometer is also based on the interference of incident and reflected light waves.
10.2 Invariance of the Two-Way Speed of Light and the Interference Phenomena
We must now discuss a certain feature of the phenomenon of interference. Because of our usage of Galilean transformations within electrodynamics we have some apparent paradoxes. One finds many books which say that a Galilean transformation of the velocity of light is not consistent with the electron-theoretical explanation of refraction and reflection.
Let us now consider a reflection and refraction in the context of special relativity. We are now in a position to understand what is happening when light radiated by a moving emitter goes through a stationary material like glass. We see that what we have to do is to calculate the interference between the incident radiation and the radiation due to the material. At first glance, if one wants to calculate the refraction and the reflection in a single inertial frame one should take into account that the metrics, and consequently the electrodynamics equations, are different for a moving emitter and a stationary glass. So the possibility of calculating interference effects is not clear.
Figure 47: Light radiated by a moving emitter reflects from a stationary mirror in the collinear geometry. There is interference between the incident radiation and the radiation due to the material. Galilean transformation of the velocity of light is consistent with the electron-theoretical explanation of refraction and reflection. The Minkowski metric and the metric Eq. (11) give the same prediction for the wavelength of the standing wave.
If the mirror is at rest and the emitter is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves. In other words, the same metric should be applied and kept in a consistent way for both incident and scattered waves. The requirement that the electrodynamics equations must be identical for both incident and reflected waves appear to be a paradox here. In fact, we have two different metrics.
We have already made a few remarks about two different metrics in the aberration of light problem. In the Chapter 4, we considered only the first order approximation. We pointed out that the paradox is avoided because the peculiarity of the aberration of light (perpendicular) geometry is that even after the Galilean transformation the electrodynamics equations in the absolute time coordinatization will have Maxwell’s form for the incident plane wave. Let us now see what happens in the collinear geometry, Fig. 47.
Here we just summarize the general idea. We will show that the paradox is avoided because Minkowski metric and metric Eq. (11) give the same prediction. The point is that all methods to measure the interference, indeed, the standing wave (i.e. round-trip) measurements. In contrast, the deviation of the energy transport direction is a geometrical effect. Because we have empirical access only to the round-trip average speed of light the value of the one-way speed of light is just a matter of convention without physical meaning. In contrast to this, the two-way speed of light, directly measurable along a round-trip, has physical meaning.
We must now discuss a certain feature of the phenomenon of interference. One finds many books which say that a Galilean transformation of the velocity of light is not consistent with the electron-theoretical explanation of refraction and reflection. It is widely accepted that if we consider a moving source and a stationary glass, the incident light wave and wave scattered by the dipoles of the glass cannot interfere as required by the electron theory of dispersion since their velocity are different. This is simply not true in physics. It is clear that an incident wave with a certain frequency, no matter what its velocity, excites the electrons of a glass into oscillations of the same frequency. They then emit radiation with the same frequency. Thus, the incident and scattered wave at any given point have the same frequency and can interfere. The effect of the different velocities is to produce a relative phase which varies with position in space. This affects the velocity and amplitude envelope of the single wave which results from the superposition of the two separate waves.
The following simple analysis confirms these ideas. We will describe effect by assuming that the metric is non-diagonal, Eq. 11. Let $\exp i(\omega t-kx)$ represent an incoming wave whose velocity is $\omega/k=c+v$. Similarly, let $\exp i(\omega t+k^{\prime}x+\phi)$ represent another out-coming (scattered radiation) wave of the same amplitude and the same frequency, a different velocity $\omega/k^{\prime}=c-v$ and different phase. The superposition of these two waves is represented by $\exp i(\omega t-kx)+\exp i(\omega t+k^{\prime}x+\phi)=2[\cos[(k+k^{\prime})x/2+\phi/2]]\exp i[\omega t-(k-k^{\prime})x/2+\phi/2]$. There is the cosine factor representing an amplitude envelope which is stationary in space and whose periodicity is inversely proportional to the difference in the propagating constants $k$ and $k^{\prime}$ of the two component waves. This can be written in a simpler form $2[\cos[\omega x/[c(1-v^{2}/c^{2})]+\phi/2]]\exp i[\omega t-xv\omega/[c^{2}(1-v^{2}/c^{2})]+\phi/2]$.
Suppose that the source at rest is emitting waves at frequency $\omega_{0}$. In the lab frame after the Galilean transformation the velocity of incoming wave is $c+v$. Thus if $\omega_{0}$ is the natural frequency, the observed in the lab frame frequency would be $\omega=\omega_{0}/[1-v/(c+v)]$. Therefore, the observed in the lab frame frequency is $\omega=\omega_{0}(1+v/c)$. The shift in frequency observed in the above situation is the well known Doppler effect. Our equation for superposition of two waves now looks like $2[\cos[\omega_{0}x/[c(1-v/c)]+\phi/2]]\exp i[\omega_{0}(1+v/c)t-xv\omega_{0}/[c^{2}(1-v/c)]+\phi/2]$.
Suppose that an observer in the laboratory performs the standing wave measurement. We should examine what parts of the measured data depends on the choice of synchronization convention and what parts do not. We state that time oscillation has no intrinsic meaning - its meaning only being assigned by a convention. In particular, one can see the connection between the time shift $xv\omega_{0}(1+v/c)/c^{2}$ in $\exp i[\omega_{0}(1+v/c)t-xv\omega_{0}/[c^{2}(1-v/c)]+\phi/2]$ and the issue of distant clock synchrony. Note that the scale of time (frequency) is also unrecognizable from physical viewpoint.
Suppose that the laboratory observer performs a measurement of the wavelength of the standing wave. The relation $dl_{i}^{2}=(-g_{11}+g_{01}g_{01}/g_{00})dx^{2}$ gives the connection between the spatial line element and the metric. With the help of Eq.(11), we find $dl_{i}^{2}=dx^{2}/(1-v^{2}/c^{2})$, which determines the spatial geometry in the inertial frame. This is explained as due to Lorentz contraction of the measuring rods in the inertial frame. Then, when the measured data is analyzed, the laboratory observer finds that the wave number is equal to $\sqrt{1-v^{2}/c^{2}}\omega_{0}/[c(1-v/c)]=\omega_{0}\sqrt{1+v/c}/\sqrt{1-v/c}$. The observer finds that the wavelength of radiation from moving source (the source moves towards the observer) is decreased by the factor $\sqrt{1-v/c}/\sqrt{1+v/c}$. We see that it is the same factor that we can obtain by assuming that the metric is diagonal. In other words, the laboratory observer will measure the same two-way speed of light, irrespective of the choice of metric.
The theory of relativity shows us that there is a certain quantity, which we call two-way speed of light, that does not change under any reference frame transformations. Another way in which our complicated problem can be solved is the following interesting one. The reason why the laboratory observer will measure the same two-way speed of light is that the method is based on the phenomenon of interference. There is a fact governing all interference phenomena. The law is called the relativistic invariance of phase. It states that there is a certain quantity, which we call phase, that does not change under any reference frame transformations. The relativistic invariance of phase is enormously useful in making analysis (as we will illustrate in a few examples in the Chapter 11) without knowing all the formulas.
In the problem discussed we have still a puzzle. Let us consider the acceleration of a dipole source in the lab inertial frame up to velocity $v$ along the $x$-axis. Without changing synchronization in the lab frame after the source acceleration we have a complicated situation as concerns dynamics and electrodynamics of moving charges. Conventional particle tracking in the single inertial frame is actually based on classical Newton mechanics. It is not difficult to see that the peculiarity of this situation is that here a solution of the dynamics problem in the inertial frame makes no reference to Lorentz transformations. This means that, for instance, within the inertial frame the motion of particles looks precisely the same as predicted by classical mechanics, with its absolute time. The relativity of simultaneously (i.e. mixture of positions and time) do not have a place in this description. The dynamical evolution of the particle according to the non-covariant particle tracking may be considered as the result of successive infinitesimal Galilean transformations. The corrected Newton law is valid for each step.
On the other hand, the Maxwell’s equations can be applied in the inertial frame only in the case when Lorentz coordinates are assigned. We will then ask how it is possible to take old kinematics for mechanics and Einstein’s kinematics for electrodynamics. This approach needs an explanation. The resolution of this puzzle lies in the fact that such (coupling fields and particles) method is only suitable to account for the outcome of the experiments in collinear geometry. In particular, the usual Maxwell’s equations and corrected Newton’s second law can explain all interference experiments that are performed in a single inertial frame. The dynamical evolution of the particle according to the non-covariant particle tracking may be considered as the result of successive infinitesimal Lorentz transformations.
How can that be? We already know that the Galilean boosts commute. The collinear Lorentz boosts also commute. Now let us see what happens to the moving source. The peculiarity of the collinear geometry where there is a source moving along the same line as the radiated beam is that here the velocity is perpendicular to the plane of radiation wavefront (i.e plane of simultaneity). Thus, for collinear motion, the plane of simultaneity in the absolute time coordinatization will have the same orientation for the Lorentz coordinatization. In the case of collinear geometry, a motion of the source, according to the special relativity, influences the kinematic terms of the higher (than $v/c$) order only. At first site, the Newton’s particle dynamics (in the case of collinear geometry) does not involve the relativistic effects at all. But the higher order effects are there in the assumption that the mass of the moving particle is equal to its relativistic mass. We shall discuss this subject further in the Chapter 13.
10.3 Field Measurements in the Accelerated Frame
In special relativity, we are confronted with a few things that might seem counter intuitive. But this is still a mathematically based theory. It tell us a peculiar thing: that when we are accelerated electron source, the fields in the accelerated frame is not static; the problem is time dependent. That what special relativity says! When the electron source and observer are at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be nonzero. Intuition would seem to tell us that everything is at rest, so the electric field is not changing with time. But special relativity says that there is a derivative $\partial\vec{E}/\partial t_{n}=v\partial\vec{E}/\partial x_{n}$ because there is $\partial\vec{E}/\partial x_{n}$ that is not zero. This is just a feature of the 4D formalism with no special consequence. A comparison with a gauge transformation in classical electromagnetism might help here. In electrodynamics, it is often convenient to introduce the Coulomb gauge. In this case field propagates instantaneously. But this instantaneous propagation is just mathematical description in an intermediate step: when a complete calculation is made, proper cancellations of the instantaneous propagation take place.
10.3.1 The Aberration of Particles from a Moving Source
Let us consider what happens if the electron gun, an electron source for a relativistic electron beam, in the initial inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis, Fig. 48. In other words, we consider so-called active (physical) boost. In the case of an active boost we are dealing with motion of the same physical system, evolving in time and treated from the point of view of the same reference system. The simplest method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the particle source was at rest, i.e. we still enforce the clock transport synchronization ( or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest). This choice is usually the most convenient one from the viewpoint of the smooth tailoring of the metric tensor. This synchronization convention preserves simultaneity and is actually based on the absolute time (or absolute simultaneity) convention.
Figure 48: Aberration of electrons from a moving source in an inertial frame. The motion of the electron gun is perpendicular to the optical axis. If an observer is at rest relative to the fixed stars and the electron gun started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the source seen in the inertial frame would jump by angle $\theta_{a}=v/V$, where $V$ is the velocity of the accelerated electrons.
We can obtain the electric field $\vec{E}$ and the magnetic $\vec{B}$ observed in the inertial frame where the particle source is moving with velocity $v$ using the Galilean transformation. When we consider the electrodynamics equations under Galilean transformation, there is a question: what is the transformation law for the electromagnetic fields? The equivalence of the active and passive pictures is due to the fact that moving system one way is equivalent to moving the coordinate system the other way by an equal amount. It is for this reason that fields in the inertial frame before the active Galilean boost are the same as those $\vec{E^{\prime}}$ and $\vec{B^{\prime}}$ in the comoving coordinate system. With the Galilean transformation, the field transformation, for the case $\vec{B^{\prime}}=0$, is $\vec{E}=\vec{E^{\prime}}$, $\vec{B}=\vec{v}\times\vec{E^{\prime}}/c$. This expression describe first order (in $v/c$) effects only. Here $\vec{v}$ is the velocity vector of the source in the lab frame.
[…]
Figure 49: Aberration of electrons when the electron source and observer are at rest in the accelerated frame. The aberration increment is connected with the problem parameters by the relation $\theta_{a}=(1-1/\gamma)v/V$. Here $\gamma=1/\sqrt{1-V^{2}/c^{2}}$ is the relativistic factor.
We described the effect of aberration of particles by working only up to the first order $v/c$. An elementary explanation of this effect is well-known. The aberration of particles in the inertial lab frame can be easily explained in terms of Galilean transformation of velocities between different reference frames. It is important to stress at this point that the dynamical line of arguments discussed here explains what the aberration of particles physically means in the inertial frame. The kinematics effects is only an interpretation of the behavior of the electromagnetic fields.
In Chapter 8 we already discussed the aberration of electrons. We had to restrict ourselves to discussing the aberration only for the case when the electron source and observer are at rest in the accelerated frame. The physical principles that produces the aberration increment were, however, made clear. We have seen in Chapter 8 that the Wigner rotation associated with the transformation from inertial coordinate system to the accelerated coordinate system may be regarded as a result of the action of certain electromagnetic (Lorentz) force. In fact, the electromagnetic forces which govern the properties of an emitted electron beam must be affected by acceleration with respect to the fixed stars in such a way that they lead to a deviation of the electron transport direction, Fig. 49.
The microscopic approach for the motion of electrons in the accelerated frame gives a similar (for the motion of electrons in the inertial frame) answer for the total transverse momentum, $p_{\perp}$, of the accelerated electron beam except with the following modification. Because the electron source is at rest in the accelerated frame, the initial condition is modified. We can rewrite our formula for the transverse momentum in the inertial frame as $p_{\perp}=p_{0}+\Delta p_{\perp}$, where $\Delta p_{\perp}=vT/c^{2}=mv/\sqrt{1-V^{2}/c^{2}}-mv$. In the accelerated frame, the initial momentum of the emitted electron $p_{0}$ is equal to zero. Therefore, when the electron source is at rest in the accelerated frame, there is an aberration increment given by $\theta_{a}=p_{\perp}/p=\Delta p_{\perp}/p=(1-1/\gamma)v/V$. So we see that we could analyze this complicated situation either by the idea that there is an electromagnetic forces and that they lead to a deviation of the energy transport, or else by the Lorentz coordinatization and that the aberration of particles effect in the accelerated frame (as viewed from the inertial frame) presents a kinematics effect (Wigner rotation) of special theory of relativity.
10.3.2 Magnetic Field Measurements in the Accelerated Electron Source
Let us return to the case when the electron source and observer are at rest in the accelerated frame, Fig. 49. We would now like to describe an apparent paradox. We obtained the electrodynamics equations in accelerated frame using the inverse Galilean transformation (with velocity $-v$) of the Maxwell’s equations. With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. The electric $\vec{E}_{n}$ and magnetic $\vec{B}_{n}$ fields are not changing with time - all of the fields are static, so that the problem is not time dependent. There are many physicists who have already received knowledge about the electrodynamics from textbooks and who would say, ”There must be currents in order to get a static magnetic field at all - and currents can only from moving charges”. Since there are no moving charges , the situation seems indeed paradoxical. Here we have a new kind of situation which is quite different from (initial) inertial frame electrodynamics.
How can that be? When the electron source and observer are at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be nonzero. One can easily show this by transforming coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$. The variables $x,y,z,t$ can be expressed in terms of the independent variables $x_{n},y_{n},z_{n},t_{n}$ by means of a Galilean transformation, so that fields can be written in terms of $x_{n},y_{n},z_{n},t_{n}$. From the Galilean transformation $x_{n}=x-vt,~{}y_{n}=y,~{}z_{n}=z,~{}t_{n}=t$, after partial differentiation, one obtains $\partial/{\partial t_{n}}=\partial/{\partial t}+v\partial/{\partial x}$, $\partial/{\partial x_{n}}=\partial/{\partial x}$. So $\partial/\partial t_{n}=v\partial/\partial x_{n}$ at $\partial/\partial t=0$.
A natural assumption is to consider that time derivative $\partial/\partial t_{n}$ arises from the Galilean boost. Now we shall be more general and study the case of the Lorentz coordinatization. When the system $S_{n}$ starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. The time $t^{(L)}_{n}$ under the Einstein’s synchronization in the $S_{n}$ frame is readily obtained by introducing the offset factor $x_{n}v/c^{2}$ and substituting $t^{(L)}_{n}=t_{n}-x_{n}v/c^{2}$ in the first order approximation. The new time coordinate in the accelerated frame is interpreted by saying that Maxwell’s equations are applicable to the aberration of electrons description. After partial differentiation, one obtains $\partial/{\partial t^{(L)}_{n}}=v\partial/{\partial x_{n}}$, $\partial/{\partial x^{(L)}_{n}}=\partial/{\partial x_{n}}+(v/c^{2})\partial/{\partial t_{n}}$. So $\partial/\partial t^{(L)}_{n}=v\partial/\partial x^{(L)}_{n}$ neglecting terms of order of $v^{2}/c^{2}$. By changing the (four-dimensional) coordinate system, one cannot obtain a physics in which new physical phenomena appear.
When the situation is described as we have done it here, there doesn’t seem to be any paradox at all; it comes out quite natural that separation between space and time is no longer possible in general. The argument that the result is paradoxical runs some thing like this: If we could measuring magnetic field $\vec{B}_{n}$ by usual magnetometer there would indeed be a paradox. But we could not do that. We already discussed above that magnetic field $\vec{B}_{n}$ leads to change in the transverse momentum of test particles given by $\theta_{a}=p_{\perp}/p=(1-1/\gamma)v/V$. If a particle motion velocity is non-relativistic, the binomial expansion yields $\vec{\theta}_{a}=-[V^{2}/(2c^{2})]\vec{V}\times\vec{v}/V^{2}$ and the momentum perturbation presents a relativistic effect. So special relativity says that there are no changes in the (test) particle momentum as $V^{2}/c^{2}\to 0$. If we look at the usual (i.e. classical) magnetometer with test particles, we see no evidence of the magnetic field $B_{n}$ effect. Not only does this particular kind of classical magnetometer can not measure the magnetic field $B_{n}$, but if the theory of relativity is correct, any other classical magnetometer, operating on any principle whatsoever, would also appear to demonstrate no evidence of the magnetic field $B_{n}$ effect (1).
Now let us return to observations of an inertial observer, Fig. 48. In this case we are dealing with motion of the electron source, evolving in time and treated from point of view of the lab system. It is not difficult to see that the peculiarity of the situation is that here the ”source motion” is a real observable effect. Indeed, the problem is time dependent. In this situation, a velocity of the source has physical meaning. In the case of moving electron source the special relativity says that magnetic field $\vec{B}=\vec{v}\times\vec{E}$ leads to change in the transverse momentum of test particles given by $\theta_{a}=p_{\perp}/p=v/V$ and this result is valid also in the non-relativistic limit $V^{2}/c^{2}\to 0$. Now we can raise an interesting question. Suppose that lab observer is measuring the magnetic field with an usual magnetometer. Now we know that there are moving charges, so that means that when there is a current in the moving source, the source generates a magnetic field which can be measured with an usual magnetometer. We come to the conclusion that the magnetic field generated by a moving source in an inertial frame is in fact a real field in the sense we have defined it.
10.4 Inertial Frame View of Accelerated Light Clock
In the Chapter 9, we described the orthogonal light-clock based on observations made by an observer in the same accelerated frame as the light-clock. Now we want to extend the description in an important way - in a way that says something in detail about how light is propagated in the inertial frame. We begin by recalling the aberration of light from a moving laser source in an inertial frame of reference. Although the laser seems to be quite different from a simple electromagnetic model for the light-clock of special relativity, the two systems are, of course closely related.
10.4.1 The Aberration of Light from a Moving Laser Source
The aberration of light effect can be described within the standard special relativity taking advantage of the Wigner rotation theory. When one has a transversely (i.e. parallel to the mirrors) moving laser there is the deviation of the energy transport for light transmitted from the optical resonator. In the first order approximation in $v/c$, the inertial observer would find that angular displacement is $\theta_{a}=2v/c$, Fig. 50.
How shall we describe the aberration of light beam from the laser source which is accelerated from rest up to velocity $v$ in the inertial frame? First, we notice that there is an extra factor 2. On the basis of conventional theory one gets aberration increment only one-half as big as that predicted by special relativity. One-half of this increment is a consequence of the fact that the Doppler effect is responsible for the angular frequency dispersion of the radiated light waves. Another source of the aberration of light is the Wigner rotation. As viewed from the (Lorentz) initial inertial frame, the coordinate axes of the accelerated frame is rotated through the angle $v/c$ with respect to the coordinate axes of the inertial frame.
In regard to light aberration one should differentiate between that from the ”plane wave” emitter and that from the laser source. Suppose we have a ”plane wave” emitter. We consider the case when the ”plane wave” emitter is accelerated from rest up to velocity $v$ in the direction perpendicular to its optical axis. Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the light transport measurement. Then how does the light beam from the moving ”plane wave” emitter looks? The inertial observer would find that angular displacement is equal to $\theta_{a}=v/c$, Fig. 51.
The first thing we would say about inertial frame measurements is that there is an intuitively plausible way to understanding the aberration of light from a moving ”plane wave” emitter. An elementary explanation of this effect is well-known. This phenomenon is fully understandable in terms of transformation of velocities between different reference frames. The aberration of light in the initial inertial frame can also be easily explained on the basis of corpuscular theory of light. This is plausible if one keeps in mind that a light signal represents a certain amount of electromagnetic energy. Energy, like mass, is a quantity that is conserved, so that a light signal resembles, in many aspects, a material particles. Therefore, we should expect that group velocities of light signals obey the same addition theorem for particle velocities. A closer treatment based on wave theory of light confirm this expectation.
In the case of the moving laser source, intuition would seem to tell us that aberration increment would be the same. But the special relativity says that there is an extra factor 2. The commonly accepted derivation of the aberration of light effect does not account for the Wigner rotation. It is incorrectly assumed that accelerated and inertial observers have common 3-space.
Figure 50: Aberration of light from a moving laser source in an inertial frame of reference. The motion of the laser resonator is parallel to the mirrors. If an observer is at rest relative to the fixed stars and the laser source started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the laser source seen in the inertial frame would jump by angle $\theta_{a}=2v/c$. It is assumed that the accelerated laser operates in the steady-state regime.
Figure 51: Aberration of light from a moving ”plane wave” emitter in an inertial frame of reference. The motion of the source is perpendicular to the optical axis. If an observer is at rest relative to the fixed stars and the ”plane wave” emitter started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the source seen in the inertial frame would jump by angle $\theta_{a}=v/c$.
Figure 52: Aberration of light in an accelerated frame of reference. It is assumed that the accelerated laser operates in the steady-state regime. The aberration of light from the accelerated laser source is absent. The electromagnetic wave travels in the resonator forward and back reflecting from mirrors. The crossed term in the Langevin metric, which generates aberration, cancels during evolution of the radiation in the optical resonator.
Figure 53: Aberration of light from a ”plane wave” emitter in an accelerated frame of reference. If the ”plane wave” emitter is at rest in the accelerated frame, then the apparent angular position of the source seen in the accelerated frame would jump by angle $\theta_{a}=-v/c$.
Now we ask about the angular displacement of light transmitted from the laser resonator inside the accelerated frame. The accelerated observer would find that the deviation of the energy transport direction is absent, Fig. 52. According to asymmetry between the ”plane wave” emitter and the laser source, there is a remarkable prediction on the theory of the aberration of light. Namely, if the ”plane wave” emitter was at rest relative to the fixed stars and this source started from rest to motion relative to the fixed stars, then the apparent angular position of the ”plane wave” emitter seen in the accelerated frame would jump by angle $-v/c$, Fig. 53. Obviously, in order to understand the aberration of light, we will have to use the special theory of relativity and not be satisfied with classical theory.
How can such difference between a ”plane wave” emitter and a laser source come about? Let us first ask: ”What happens to accelerated optical resonator?” When we were dealing with optical resonators we used the electrodynamics boundary conditions in 3-space, and they defined the energy propagation direction along the optical axis (i.e. perpendicular to the mirror surface). According to special relativity, an acceleration does not change the direction of the optical axis (i.e mirrors orientation) with respect to the accelerated frame (3-space) axes. An observer in the rest frame of a laser source finds that the direction of the energy transport is independent of the motion with respect to the fixed stars, Fig. 52. We turn now to the radiation emitted by the accelerated laser source in the inertial frame. As viewed from the inertial frame, the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$ is rotated through the Wigner angle $\theta_{w}$ with respect to the axes of the inertial frame $(x,y,z)$. For light $V/c\to 1$ and the $z_{n}$-axis of accelerated frame is then rotated with respect to the $z$-axis by angle $\theta_{w}=v/c$. As a result, an extra factor of 2 appears in the aberration increment, in Fig. 50.
Let us now discuss more of the difference between laser and incoherent source. Suppose we have a ”plane wave” emitter which is at rest in an accelerated frame. How shall we find the general wave solution in the case of a point source? The answer is that electrodynamics equation needs to be integrated with initial condition for the radiation wavefront. According to the special relativity, physical process of light creation takes place in a metric space-time geometry, so the metric tensor must be used to describe the light creation and propagation in empty space. We only wish to emphasize here the following point. According to special relativity, accelerated and inertial observers have common absolute space-time but have different metrics and, consequently, different 3-spaces. When we are dealing with optical resonator the (3-space) boundary conditions define the energy propagation direction. In contrast, when we are dealing with a point source the combination of initial condition (i.e orientation of the radiation wavefront ) and metric tensor is important. We would like to emphasize a very important difference between laser and incoherent source. The radiation of incoherent source clearly depends on an initial conditions and there is influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer. The solution presented in terms of the absolute time coordinate, hence the wavefront of the emitted light beam is perpendicular to the direction $z_{n}$ after the acceleration. Using this initial condition and the Langevin metric in a frame of reference attending acceleration enables us to explain the aberration of light from a ”plane wave” emitter, Fig. 53.
10.4.2 Orthogonal Light Clock Operation
Slowing of the clock in a moving system is a very peculiar phenomenon. We have already discussed accelerated clocks in the previous chapter. We described the parallel and orthogonal light clocks based on the observations made by an observer in the same accelerated frame as the clock. Now let us return to observations of an inertial observer. We must say immediately that there is no objection to the standard description of the parallel light clock in the inertial lab frame. What must be recognized is that the concept of Wigner rotation is only introduced in the description of an orthogonal light clock.
Let us consider a light clock accelerated up to velocity $v$ transverse to the direction of optical pulse. First, we examine the reasoning presented in textbooks. Suppose that an inertial system $S$ (and an observer with his measuring instruments) is at rest with respect to the fixed stars and a system $S_{n}$ in the lab frame $S$ is accelerated from the rest with respect to the fixed stars up to velocity $v$ along the $x$-axis. A clock which we call ”light-clock” is a rod with mirror at each end. According to textbooks, when inertial observer looks at the orthogonal light clock going by, he sees that the light, in going from mirror to mirror is taking a zigzag path. If in a giving time the clock moves forward a distance proportional to $v$ in Fig. 54, the distance the light travels in the same time is proportional to $c$, and the vertical distance is therefore proportional to $\sqrt{c^{2}-v^{2}}$. That is it takes a longer time for light to go from end to end in the moving clock than in stationary clock. Therefore the apparent time between clicks is longer for the moving clock in the proportion $1/\sqrt{1-v^{2}/c^{2}}$.
Figure 54: The ”light clock” thought experiment. The calculation of the time interval between the sending and reception of the light signal in the reference frame $S$ relative to which a light clock moves. According to textbooks, it is assumed that $(x_{n},y_{n},z_{n})$ axes of the moving frame and $(x,y,z)$ axes of the lab observer are parallel.
Nevertheless, there is argument against this commonly accepted derivation of the slowing down of an orthogonal light clock in a moving system. The standard explanation is based on the hidden assumption that $(x_{n},y_{n},z_{n})$ axes of the moving observer and $(x,y,z)$ axes of the lab observer are parallel. In other words, it is assumed that observers have common 3-space. This is misconception. In fact, two observers with different trajectories in the Lorentz coordinatization have different 3-spaces.
The accepted in textbooks assumption of rigidity of accelerated frame is based on the belief that simultaneous acceleration of the reference space grid $(x_{n},y_{n},z_{n})$ has direct physical meaning. However, the position along the $x$-axis of the accelerated first mirror with respect to the accelerated second mirror has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this position could be ascertained. There is an uncertainty (blurring) of the relative position in the $x$-direction of amount $lv/c$, where $l$ is the rod length. This uncertainty in the relative position exist due to the uncertainty in the moments of the acceleration. The theory of relativity shows us that the relationship of positions and times are not what we would expect intuitively. When we have two distant events, we have deal with the conventionality of distant simultaneity within the time interval $l/c$, where $l$ is the space separation of these two events.
Now let us see how we can compare the clock rates in the frame $S$ and $S_{n}$. It is generally believed that observing clock rates in the two frames $S$ and $S_{n}$ moving relative to each other, one can only compare readings of one clock from one frame with readings of several clocks from another frame, because two clocks from different frames get together at the same point in space only once. In one of the frames there must be at least two clocks which are supposed to be synchronized.
We have already pointed out that in all cases of time measurements in the framework of special relativity the time measured in this case is actually ”length”. In the case of length measurement of resting objects we are dealing with objective (i.e. convention independent) physical quantity. Because we have empirical access only to the length measurements, it is impossible to agree with textbook statement that one can only compare readings of one clock from one frame with readings of several synchronized clocks from another frame. Let us try out our operational interpretation of time measurements on the light clock example, to see how it works.
Let us consider a specific thought experiment. We shall discuss the situation where time marks can be imprinted on the moving object (screen) by a clock, as sketched in Fig. 55. First, we describe the orthogonal light-clock based on the observations made by an observer in the same accelerated frame as the light-clock. In the case of Langevin metric Eq.(15), the speed of light radiated by the accelerated emitter in the transverse direction is $c\sqrt{1-v^{2}/c^{2}}$. Therefore, the time interval between the sending and reception of the light signal is $\Delta t_{n}=2l/[c\sqrt{1-v^{2}/c^{2}}]$. The spacing of the time marks, i.e the distance $\Delta x_{n}$ through which the object (screen in the lab frame) falls in the time is $\Delta x_{n}=\Delta x=2lv/[c\sqrt{1-v^{2}/c^{2}}]$.
Taking into account the Langevin metric Eq.(15), we obtain at $dt_{n}=dt=0$ that $dx_{n}=dx$, i.e. the length of physical rod in the inertial frame coincidence with the coordinate distance in the accelerated frame. Because $t_{n}=t$ and $\Delta t=2l/c$ has the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that, as physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame. The accelerated clock rate measured by the lab observer is actually spacing of the time marks on the lab screen.
Figure 55: The ”light clock” thought experiment. Time marks can be imprinted on the moving screen by clock. Observations made by an observer in the same accelerated frame as the light clock. The anisotropy of speed of light presented in the absolute time coordinatization. The distance $\Delta x_{n}$ through which the screen in the lab frame falls in the time (i.e. spacing of the time marks) is $\Delta x_{n}=\Delta x=2vl\gamma/c$.
Second, we describe the orthogonal light-clock based on the observations made by an observer in the same inertial frame as the light-clock, Fig. 56. Therefore, the time interval between the sending and reception of the light signal is $\Delta t=2l/c$. The spacing of the time marks, i.e the distance $\Delta x$ through which the screen in the accelerated frame falls in the time is $\Delta x=\Delta x_{n}=2lv/c$.
Analysis based on using standard measuring rods in an accelerated frame to measure it geometrical properties shows that coordinate distance $dx_{n}$ has physical length $dx_{n}/\sqrt{1-v^{2}/c^{2}}$. We have ascertained the asymmetry of meter stick’s length directly from the Langevin metric. This shows that physical spacing of the time marks measured in the accelerated frame undergoes compress on in the direction of fixed stars motion. Because $\Delta x_{n}=\Delta x$ and $\Delta x=2vl/c$ has the physical meaning of the time rate indicated by a physical clock at rest in the inertial frame, this implies that the accelerated observer can observe the shortness of mark spacing $\Delta l_{n}=2lv\sqrt{1-v^{2}/c^{2}}]/c$. Here $\Delta l_{n}$ is the physical mark spacing measured in the accelerated frame. In fact, we deal with the same time dilation effect. This thought experiment illustrates that there is no reciprocity in time dilation.
Figure 56: The ”light clock” thought experiment. Time marks can be imprinted on the moving screen by clock. Observations made by an observer in the same inertial frame as the light clock. The distance $\Delta x_{n}$ through which the screen in the accelerated frame falls in the time (i.e. spacing of the time marks) is $\Delta x_{n}=\Delta x=2vl/c$.
Figure 57: The ”light clock” thought experiment. Phase of the orthogonal light clock in the accelerated frame. Observation made by an observer in the inertial lab frame. The phase of the accelerated light clock cannot be measured.
Figure 58: The ”light clock” thought experiment. Observation made by an observer in the inertial lab frame. The position along $x$- axis of the accelerated first mirror with respect to the second mirror has no exact objective meaning. As viewed from the Lorentz lab frame, the orientation of coordinate axes of the accelerated frame with respect to coordinate axes of the lab frame is regulated by the Wigner rotation. The accepted in textbooks assumption of rigidity of accelerated orthogonal light clock is based on the incorrect belief that axes of the accelerated frame and axes of the lab frame are parallel. In order to keep the complexity of the discussion to a minimum, we describe the Wigner rotation effect by working only up to the first order in $v/c$.
Figure 59: The ”light clock” thought experiment. Phase of the orthogonal light clock in the accelerated frame. Observation made by an observer in the inertial lab frame. According to textbooks, it is incorrectly assumed that accelerated and inertial observer have common 3-space. The possibility of measurement of the phase is prediction of conventional light clock theory and is obviously absurd from the viewpoint of special relativity.
We demonstrated that apparent time between clicks is longer for accelerated clock. Now, continuing our discussion of the light clock operation in the frame $S$ and $S_{n}$, let us consider the phase of light clocks. The word ”phase” in this case has meaning the zigzag shape of the path followed by the light beam. We have already (Fig. 55 - 56) discussed how a clock rate can be measured when the clock is moving. Above we found that the rate of moving clock (i.e. the time interval between the sending and reception of the light signal) measured by the lab observer is actually spacing of the time marks on the lab screen. Is the same statement also true for phase measurements? Let us try to demonstrate that it is.
To see the idea, suppose that the lab observer has two screens and we shall discuss the up-and-down imprinting method (i.e. the situation where time marks can be imprinting on the both screens by a moving clock). We learned in the Chapter 7 that the orientation of coordinate axes of the accelerated frame with respect to coordinate axes of the Lorentz lab frame is regulated by the Wigner rotation. We can show that, as consequence of Wigner rotation, the shape of the zigzag pattern cannot be measured in the lab frame, Fig. 57.
The standard presentation of the orthogonal light clock operation is based on the hidden assumption that accelerated and inertial observers have common 3-space. According to the conventional theory, when a reference frame at rest is put into motion, then all points in the 3D reference grid have a start to move at same time. Consequently, rigidity of the orthogonal light clock in the sense introduced by textbooks cannot be considered as relativistic kinematics property, Fig. 58. Our result (Fig. 57) is at odds with the prediction from textbooks, Fig. 59. The commonly accepted description of the orthogonal light clock operation does not account for the Wigner rotation, which in our case closely associated with the relativity of simultaneity. The possibility of measurement of the phase is prediction of conventional light clock theory and is obviously absurd from the viewpoint of special relativity.
We want now to describe briefly the mechanism of initialization of the orthogonal light clock in the accelerated frame. In order to understand how light is going from mirror to mirror, we have to watch the machinery of the light clock and see what happens when it is moving in the inertial frame. We shall take a very simple a kind of light clock initialization, but it will work in principle. Although the Fabry-Perot interferometer we have been describing in the Chapter 8 seems to be quite different from the light clock of special relativity , the systems are closely related. We consider the simple case when an optical cavity is equipped with plane mirrors. Suppose that both mirrors are semitransparent for marking the both screens to measure the time between clock clicks. The light clock can be initialized using a short laser pulse (from an external laser in the accelerated frame) coupled in through the top (down) semitransparent mirror.
Suppose that the initialization laser pulse is impinging onto the top mirror of the light clock at point $A$, Fig.57. Special relativity tells us a peculiar thing: that light signal travels from the top mirror to the bottom mirror at the location denoted $B$. The position of the top mirror when it meets the returned signal is at $C$. This motion ($A\to B\to C$) constitutes the first tick of the clock, and subsequent ticks ($C\to D\to F$) are due to additional reflections of the pulse from the top and the bottom mirrors.
That’s what special relativity theory says. How can that be? It is clear that our ordinary intuition are quite wrong when we have deal with different 3D-spaces for different observers in the 4D space-time. In the space measurement of one observer there is mixing of space and time, as seen by other. Most paradoxes of the moving light clock operation are closely linked to this mixing. However, these paradoxes are only apparent, and can be removed by employing the pseudo-Euclidean space-time geometry. The position along the $x$-axis of the accelerated bottom mirror with respect to the accelerated top mirror has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this position could be ascertained . There is an uncertainty (blurring) of the relative position in the $x$-direction of amount $lv/c$ (due to the uncertainty in the moments of the acceleration), Fig.58.
10.5 Bibliography and Notes
1. In order to understand magnetic field measurements in the accelerated electron source, we have to watch the machinery of the classical magnetometer and see what happens when it is at rest in the accelerated frame. Since it is rather difficult, we shall take a simple a kind of usual magnetometer. It is relatively easy to observe the nuclear magnetism by the phenomena of ”nuclear magnetic resonance”. A proton resonance apparatus can be used as proton resonance magnetometer. We give a general analysis of this type of classical magnetometer independently of the experimental accuracy. In general a current loop, which has a magnetic moment $\vec{m}$ in the inertial lab frame in which it is at rest, has an electric dipole moment equal to $\vec{p}=\vec{v}\times\vec{m}/c$ when it is moving with velocity $\vec{v}$ relative to the lab frame. The magnetic properties of materials are attributed to atomic current loops. Then a magnetized body which is accelerated relative to the lab frame should have an electric polarization relative to the lab frame. Thus, in the lab frame, the additional potential energy of a moving particle with the proper magnetic moment $\vec{m}$ in the magnetic field $\vec{B}$ and the electrical field $\vec{E}$ becomes $U_{p}=-\vec{m}\cdot\vec{B}-\vec{p}\cdot\vec{E}$. We obtained the electrodynamics equations in accelerated frame using the Galilean transformation (with velocity $-\vec{v}$) of the Maxwell’s equations. With the Galilean transformation, the magnetic dipole transformation for the case $\vec{p}=0$, is $\vec{m}_{n}=\vec{m}$, $\vec{p}_{n}=-\vec{v}\times\vec{m}/c$. As we know from vector algebra, $(\vec{a}\times\vec{b})\cdot\vec{c}=\vec{a}\cdot(\vec{b}\times\vec{c})$; so our terms are also the same as $(-\vec{v}\times\vec{E})\cdot\vec{m}=-\vec{v}\cdot(\vec{E}\times\vec{m})$, $(-\vec{v}\times\vec{m})\cdot\vec{E}=-\vec{v}\cdot(\vec{m}\times\vec{E})$, as evaluated in the accelerated frame. Hence we obtain the additional potential energy equal to zero just as we expected. We could not measuring magnetic field $\vec{B}_{n}$ by usual magnetometer. Now let us return to observations of an inertial observer. In the case of moving electron source the special theory of relativity says that magnetic field $\vec{B}=\vec{v}\times\vec{E}$ can be measured with the usual magnetometer. Now that our magnetometer is at rest in the lab frame, so that means that the additional potential energy of a proton with the proper magnetic moment $\vec{m}$ in the magnetic field $\vec{B}$ is equal to $-\vec{m}\cdot\vec{B}\neq 0$. The moving electron source in the lab frame generates a magnetic field which can be measured with an usual magnetometer.
11 A Preferred Inertial Frame and the Second Order Experiments
It should be note that all well known methods to test the special relativity are second order optical measurements. The cardinal example is given by the Michelson-Morley interference experiment. A close examination of all second order optical experiments inside an inertial frame shows that phenomena appear to be independent of the uniform motion relative to the fixed stars. It is not possible to detect the orbital velocity of the earth using second order interference experiments. This is consequence of the geometry of space-time.
A number of other experiments were curried out, which all attempted to measure the velocity of the earth relative to the sun. They are not connected with light interference phenomena. We shall discuss experiments that compare the rates of different atomic clocks that move together relative to the earth-based frame. The question arises whether it is possible to determine experimentally the state of motion of the initial inertial frame $S$ in the accelerated frame $S_{n}$ by means of the time dilation effect for moving reference (atomic) clock. Until now, all test of time dilation with atomic clocks in the earth-based frame have involved circle path geometry. According to the special relativity, in any clock experiment which probe the influence of the orbital velocity one must use the linear (one-way) path geometry. These are experiments of decays of rapidly moving unstable particles. However, the modern experimental techniques are not sensitive enough to register the effect of orbital velocity on the average distance traveled by a disintegrated particle.
11.1 Second Order Optical Experiments and Special Relativity
The example of the time dilation effect is the resonance absorption of gamma rays in a Mossbauer rotating disk experiments.The measurements of the resonance absorption of gamma rays for the same relative velocity between the source and the observer give a blue shift when the source is at the center, and a red shift when it is at the tip. This experiment demonstrates that there is asymmetry in the transverse Doppler effect (1). These measurements also show that there is no a frequency shift when the source and observer are at the tip. The pulse from the source of one accelerated observer on the circle, which has red shift in the center of a circle due to time dilation, is perceived by diametrically opposite observer without frequency alteration due to his time dilation. This will take place despite the fact that each of the observer, having exchanged the pulse, is moving in the lab reference frame $S$ with velocity $v$, in opposing motion.
At first site the transverse Doppler effect problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers. The question arises whether it is possible to determine experimentally the state of motion of the inertial frame $S$ in the accelerated frame $S_{n}$ by means of the Doppler red-blue shift of second order for moving sources. In particular, the question arises whether it is possible to ascertain the state of motion of the earth relative to the sun-based frame.
Let us consider the measurements of the resonance absorption of gamma rays in a Mossbauer rotating disk experiments. Suppose that an absorber is placed in the center of a disk (which is at rest to the earth-based frame) and an emitter is placed on the edge of the disk. The emitter is moving round absorber along a circle with constant velocity $w$. In order to keep the mathematical complexity of the discussion to a minimum, we will describe the effects by working only up to the second order in $w/c$ and $v/c$, where $v$ is the earth orbital velocity. The frequency of the emitter changes with the velocity $\vec{w}$ in a manner as it suggest itself from the transverse Doppler effect. Indeed, the inner frequency of the emitter depends on its velocity $\vec{w}$, so that (using Langevin metric in the earth-based frame) $\omega(t_{n})=\omega_{0}\sqrt{1-|\vec{v}+\vec{w}|^{2}}$, then neglecting small terms $\omega(t_{n})=\omega_{0}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(t_{n})/c^{2}]$. Since the physical clock is at rest in the earth-based frame, its reading of its proper time is $d\tau=(1-v^{2}/2)dt_{n}$. We expect therefore the radiation emitted with the constant frequency $\omega_{0}$ from the emitter to arrive in the absorber with a varying frequency $\omega(\tau)=\omega_{0}[1-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(\tau)/c^{2}]$. If the emitter is moving along a circle the frequency of radiation received disk center should fluctuate periodically. Then the measure of the absorption would be expected to vary with the position of the emitter relative to the disk center.
In the actual experiment no effect on the absorption was found, when the disk was made to rotate. The negative result of the above experiment can be understood supposing that only a part of the whole frequency fluctuation is considered. A more careful analysis shows that one must take into account the change in frequency due to the so-called radial Doppler effect. We discuss a setup based on a perpendicular geometry and at first site the radial Doppler effect should be absent. We already know from our discussion in previous chapters that there is the aberration of light in the earth-based frame and the radial component of the emitter velocity is equal to $\vec{w}\cdot\vec{v}/c$. Thus we find that the emitted radiation falling on the disk center with periodically changing frequency $\delta\omega=\omega_{0}\vec{v}\cdot\vec{w}(\tau)/c^{2}$ will be exactly in resonance with the changing inner frequency. The two types of effects compensate each other and no observable effect remains.
The Doppler red shift of second order for moving atoms is a major prediction of special relativity. The first experimental confirmation of the transverse Doppler effect appeared in 1938. In this (Ives and Stilwell) experiment, ionized hydrogen molecules were accelerated in the ”cathode tube” up to 28 keV. The frequencies of the light emitted parallel and antiparallel to the beam direction were measured by a spectrograph (at rest in the laboratory). Now let us to analyze this experiment. We discuss a setup based on a parallel geometry. The radiation from the moving atoms, giving the Doppler shifted lines, was observed together with the radiation from the resting atoms existing in the same working volume, and giving an anshifted line. In that way Ives and Stilwell replaced the difficult problem of the determination of the asymmetry of shift of the ”red” and blue shifted lines with respect to the unshifted line.
At first site, the orbital motion of the earth relative to the sun can be observed through this type of experiment (2). Indeed, if the inner frequency of the atom depends on its velocity $\vec{w}$ so that (using Langevin metric in the earth-based frame) $\omega(t_{n})=\omega_{0}\sqrt{1-|\vec{v}-\vec{w}|^{2}}$ and the frequency of radiation received spectrograph should depend on the orbital velocity. In rotating disk experiments the effect which occurred on the ground of the transverse Doppler effect is compensated by the other effect. In the case of perpendicular geometry we explained of the results of the experiments taking into account the aberration of light together with the radial (i.e. classical) Doppler effect. We know that in the collinear geometry the aberration of light is absent and one must take into account only the radial Doppler effect. If we analyze the geometry of the situation, we find that the observed frequency (i.e frequency of the radiation falling on the detector) is $\omega=\omega_{0}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(t_{n})/c^{2}]/[1-w/(c-\vec{w}\cdot\vec{v}/w)]$, where (according to Langevin metric) $c^{*}=c-\vec{w}\cdot\vec{v}/w$ is the coordinate velocity of light in the earth-based frame. Thus we find just as in the case of the rotating disk experiment that also in measurements of the Doppler red shift of second order for moving atoms the two types of effects caused by the orbital motion of the earth compensate each other and no observable effect remains.
Usually, such a cancellation is found to stem from a deep underlying principle. From the series of negative results like that obtained by Michelson and Morley it seems reasonable to suppose that it is not accidental that various optical experiments trying to determine orbital velocity $v$ remained unsuccessful. And we should accept the view that there exist a general law of nature which prevents us to determine $v$ by laboratory experiments. Nevertheless, in this case there does not appear to be any such profound implication. This is a specific feature of the optical experiments of second order. Indeed, we already know from our discussion in previous chapters that it is possible to detect the orbital velocity without looking at anything external to the earth using aberration of light phenomenon.
Now many people who learn special relativity in the usual way find this disturbing. The argument that aberration shift must be symmetrical runs some thing like this (3): It is always possible to chose such variables, in which metric of the accelerated frame will be diagonal. The validity of Maxwell’s laws in all inertial frames implies that light propagates with the same velocity $c$ in all inertial frames. It implies the elimination of the privilege frame, so meaning that the notion of absolute motion does not have physical content at all.
What is wrong with a principle that the metric supplies all information about the physics of the situation, as described in the given coordinates? We already discussed this subject in the Chapter 5. At first glance, after the diagonalization of the Langevin metric, we have the symmetry between inertial frames. Where does the asymmetry comes from? The electrodynamics equations are not supplies all information about the physics of the situation. To solve the electrodynamics equations it is necessary to determine the initial conditions. We already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization.The time under the Einstein synchronization is readily obtained by introducing the time offset factor. This time shift has the effect of rotation of the plane of simultaneity. As a consequence of this, the plane wavefront rotates in the accelerated frame after metric diagonalization. Now the information about the direction of observer acceleration is recorded in the radiation wavefront orientation.
The electrodynamics equations are Lorentz covariant. This is a consequence of the (pseudo-Euclidean) geometry of space-time. However, from the mathematical viewpoint, the Lorentz covariance of electrodynamics equations is insufficient to guarantee the covariant solutions. In contrast, the accepted in previous literature incorrect assumption that covariant equations must have covariant solutions has always been considered obvious. We will discuss this subject further in the Chapter 12.
11.2 A Specific Feature of the Second Order Optical Experiments
We have seen that the principle of relativity is experimentally confirmed only in restricted sense. Presently we find that in the second order experiments the two types of effects caused by the orbital motion of the earth compensate each other and no observable effect remains. And we may accept the view that there exist a specific feature of the second order experiments which prevents us to determine the velocity of the earth relative to the sun by laboratory experiments. Let us verify that this assertion is correct. We can summarize all the effects that we shall now discuss by remarking that they have to do with the interference effects of two light beams. The light from a source is split by a suitable apparatus into two beams and the two overlapping beams produce an interference pattern. The two beam originate in the same source. Thus, it is not possible to detect the orbital velocity using second order (interference) experiments. Phase is 4D invariant (i.e. it is simply number). It is independent on the chosen inertial frame and on the chosen coordinatization. This is a consequence of the geometry of space-time. In the interference experiments neither the Doppler effect nor the aberration of light exist separately as well defined physical phenomena and only phase is meaningful quantity. This discussion clearly show that the special relativity calculations yields the observed null fringe shift in the earth-based frame.
Until now we have considered only the case in which we have deal with interference pattern as in Michelson interferometer. Now we shall more general and study the case of the Doppler shift of second order for moving atoms. The frequencies of the emitted light were measured by a grating spectrograph. There is no specific, important physical difference between diffraction grating and interferometer. Grating spectrometer is also based on the interference of incident and reflected light waves.
A more interesting example is the measurements of the resonance absorption of gamma rays in a Mossbauer rotating disk experiments. We shall try to understand the effect in a very simple case. A source is placed a large distance away from a thin plate of absorber. We inquire about the field at a large distance on the opposite side of the absorber. According to the electrodynamics, an electric field is the vector sum of the fields produced by the external source and the fields produced by each charges in the plane of absorber. We know that absorber consists of atoms which contains nuclear. When the electric field of the source acts on the nuclear it drives the nuclear charge. And moving charges generate a field - they constitute new radiators. This means that the two plane waves which interfere are propagate in the same direction. We have in this case interference of the incident and (coherently) forward scattered waves. Our model of the nuclear oscillator has some damping force. Because we put into take account of damping, the index of refraction is now a complex number and we have destructive interference. In this case we deal also with interference phenomenon.
11.3 Time Dilation Effect for Moving Reference Atomic Clock
The next thing that we have to discuss is the time dilation effect. We have already discussed this subject in Section 9.3. Here’s what we found there:
1. We have found that the slowly down of the accelerated (with respect to the fixed stars) physical clock does not depend on the choice of reference system, in which this effect is measured. In other words, the time dilation effect is absolute, but not relative.
2. According to the asymmetry between (initial) inertial and accelerated frames, there is a remarkable prediction on the theory of special relativity. Namely, suppose that frame $S$ is at rest with respect to the fixed stars and frame $S_{n}$ is accelerated with respect to the inertial frame $S$ up to the speed $v$. Now two clocks symmetrically accelerated in the frame $S_{n}$ up to the the same speed $v$ but moving in opposite direction. However, there is a difference in time dilation of accelerated clocks in the frame $S_{n}$. It is clear that clocks have different accelerational histories with respect to the fixed stars.
We would now like to describe an apparent paradox. The argument that all inertial frames are equivalent runs something like this. We know that in order to keep a Lorentz coordinates in the accelerated frame we need to perform a metric diagonalization. Here the metric in both reference frames will be diagonal, and, according to textbooks, the coordinates in this frames should be related by the Lorentz transformation. At first site the problem is complete symmetrical and both frames are equivalent. However, there is a real difference between the accelerated inertial frame and the inertial frame without accelerational history. Indeed, we should expect the metric tensor to be changed. The difference between the inertial and accelerated frames is understandable. After diagonalization, the metric tensor in the accelerated frame must abruptly change from the value $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the value $g_{00}=1,g_{01}=0$. We must conclude that when we are dealing with (Lorentz ) transformation and this transformation does not preserve continuity of the metric tensor we have no symmetry between inertial frames.
The question arises whether it is possible to determine experimentally the state of motion of the inertial frame $S$ in the accelerated frame $S_{n}$ by means of the time dilation effect for moving reference (atomic) clock. At first site it may appear as if orbital motion of the earth could be observed by time dilation experiments. Although in principle this can be done it is necessary to analyze the question in greater detail.
The proper time interval may be identified with particle lifetime, atomic transition periods, nuclear half-lives, but the crucial point here is that this interval is completely determined by his movement in the initial inertial frame. In 1971 Hafele and Keating synchronized four atomic clocks with reference clock at the laboratory, and then flew them around the earth in commercial jets (4) . The clocks were flown first from east to west and then from west to east. When the clocks returned to the laboratory they were compared with the reference clock. The clocks flown in the plane ran slower than the clocks that remainder on the earth. The agreement between the observed and predicted discrepancies confirms the predictions of special relativity for the clock acceleration with respect to the earth. In the actual experiment no effect of earth orbital velocity on the clock dilation was found.
[…]
11.4 Wigner Rotation and the Trouton-Noble Experiment
Now we shall discuss other second order experiments. They are not connected with the propagation of light and can be understood in terms of general dynamical consideration. The most important of these experiments was the experiment of Trouton and Noble.
The Trouton-Noble experiment looks for the turning motion of a charged parallel plate capacitor suspended at rest in the frame of the earth in order to measure the earth’s motion relative to the sun (i.e. relative to the fixed stars). The main idea of the experiment can be described as follows.
Consider two opposite point charges $+e$ and $-e$ in the earth-based frame; the radius-vector pointing from $-e$ to $+e$ be denoted by $\vec{r}_{n}$. If the charges are at rest the force acting upon $+e$ can be written $\vec{F}_{n}=e\vec{E}=-e^{2}\vec{r}_{n}/r^{3}_{n}$. As the force acts in the direction $\vec{r}_{n}$ the moment of force produced by the pair of charges vanishes, $\vec{M}_{n}=\vec{r}_{n}\times\vec{F}_{n}=0$.
If the pair of charges is made to move with a velocity $\vec{v}$ in the sun-based frame then the positive charge will be under the action of the Coulomb attraction of $-e$ and also under the influence of the magnetic field $\vec{B}=e(\vec{v}\times\vec{r})/(cr^{3})$. Thus the total force acting upon $e$ is given by $\vec{F}+e^{2}(\vec{v}\times(\vec{v}\times\vec{r}))/(c^{2}r^{3})$. Since $\vec{v}\times(\vec{v}\times\vec{r})=\vec{v}(\vec{v}\cdot\vec{r})-v^{2}\vec{r}$ we find that the moment of force produced by the pair of charges is equal to $\vec{M}=e^{2}(\vec{v}\cdot\vec{r})(\vec{r}\times\vec{v})/(c^{2}r^{3})$. Denoting the angle between $\vec{v}$ and $\vec{r}$ by $\theta$, we find for the absolute value of moment of force $M=e^{2}v^{2}\sin{\theta}\cos{\theta}/(c^{2}r)$.
In the actual experiment a charged condense was suspended on an elastic string. The condenser was placed so that $\theta=\pi/4$, i.e. so that the line perpendicular to the surface of the condenser plates subtended an angle $\pi/4$ with the supposed direction of the orbital velocity of the earth, the direction of $\vec{v}$. This experiment led to negative result. i.e. the expected effect proportional to $v^{2}/c^{2}$ was not found to occur.
The fact that capacitor in the earth-based frame produces no torque is in complete agreement with the special relativity and could be explained in a rather trivial way. We can obtain the electrodynamics equations in the earth-based frame using the Galilean transformation (with velocity $-v$) of the Maxwell’s equations. With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. Here $\vec{v}$ is the velocity vector of the earth in the sun-based frame. Thus we must conclude that this magnetic field produces no torque in the earth-based frame where capacitor is at rest.
In the experiment discussed we have still a puzzle. The existence of the magnetic field is responsible for the existence of the magnetic force and this force provides a torque on the charge in the sun-based frame. It is important to point out that experiment can be fully accounted for in terms of two inertial frames. We have the apparent paradox of different mechanical equations for force and torque governing the motion of a charged particle in different inertial frames; there is a torque and so a time rate of change of (3D) angular momentum in one inertial frame, but not in another relatively moving inertial frame. We will show that the real cause of the paradox is - the incorrect assumption that an earth-based observer and an sun-based observer have common 3-space. According to the special relativity, two observers with different trajectories have different 3-spaces. The standard presentation of the Trouton-Noble experiment is based on the hidden assumption that $(x_{n},y_{n},z_{n})$ axes of the moving earth-based observer and $(x,y,z)$ axes of the sun-based observer are parallel. We demonstrate that only solution of the dynamics equations in covariant form (accounting for the special relativity in the transformation of coordinate axes) gives the explanation of Trouton-Noble experiment in a complete agreement with the special relativity.
The key idea in resolution of the paradox arises in all usual explanation of the Trouton-Noble experiment is realization of the fact that in the space measurement of the earth-based observer there is mixing of space and time, as seen by the sun-based observer. While naively one should expect the angle between $\vec{v}$ and $\vec{r}_{n}$ in the earth-based frame to be the same angle between $\vec{v}$ and $\vec{r}$ in the sun-based frame it is not the case. The interpretation of the radius-vector rotation about the velocity vector is closely associated with the length contraction. Consider a rod directed along $x$-axis in the earth-based frame. After the Lorentz boost along the $x$-axis at velocity $v$ this rod will be contracted in the sun-based frame down to $l_{n}/\gamma$, where $l_{n}$ is the rod length in the earth-based frame. The relation $\gamma\tan\theta_{n}=\tan\theta$ gives the connection between the angle $\theta_{n}$ in the earth-based frame and the angle $\theta$ in the sun-based frame. Assuming $v^{2}/c^{2}\ll 1$, we find $\Delta\theta/\cos^{2}\theta=-[v^{2}/(2c^{2})]\tan\theta$, where $\Delta\theta=\theta_{n}-\theta$ is the rotation angle which is connected with the problem parameters by the relation $\Delta\theta=-[v^{2}/(2c^{2})]\sin\theta\cos\theta$. Thus we find that the moment of force produced by a pair of charges in the earth-based frame is reduced to $M_{n}=e^{2}v^{2}\sin\theta\cos\theta/(2c^{2}r)$.
[…]
11.5 Bibliography and Notes
1. The measurements of the resonance absorption of gamma rays in a Mossbouer rotating disk experiment [43] for the same relative velocity between the source and the observer give a blue shift when the source is at the center, and a red shift when it is at the tip.
2. In Ives and Stilwell experiment [44] the effect which occurred on the ground of the orbital motion of the earth relative to the sun was within the experimental limitations. This situation pertains because the high value of the emitters speed $w\gg v$ involved. Presently we give a general analysis of this type of experiments independently of the experimental accuracy.
3. Many physicists tend to think that the validity of Maxwell’s laws in all inertial frames implies the elimination of the privilege frame. To quote e.g. Dieks [45]: ” As we already mentioned, it is a basic principle of the special theory of relativity that the line element $ds$ supplies all information about the physics of the situation, as described in the given coordinates.”
4. The first experimental confirmation of the time dilation predicted in the circle path geometry with atomic clocks and aircraft was reported by Hafele and Keating in 1971 [46]. A very interesting result have been obtained by GPS team. The first example of the different inertial frames having the same accelerational histories is the global positioning system (GPS). The number of GPS satellites is around 80. The velocity of the GPS satellites with regard to the earth is 4 km/s. The experimental data of the GPS shows that the clocks in the satellites tick off time more slowly (7.1 $\mu$s every day) by the velocity. Asymmetry in clock run appears between the earth and the satellites. There are no asymmetries among the GPS satellites. That is, times are equal in every GPS satellites [47]. We see that the presence of some non zero relative velocity cannot be a cause for the effect of time dilation. In this experiment no effect of earth orbital velocity on the clock dilation was found. The negative result of the GPS experiment can be understood remembering that in this situation the clocks also are moving around the earth (i.e. around the reference clock) along the circular path and no observable effect of orbital velocity remains.
12 The Principle of Relativity and the Modern Cosmology
12.1 The Relativity Principle of Special Relativity
In the last few chapters we have treated the essential ideas necessary for an understanding of the aberration of light phenomena. We have not had to worry yet about the fact that the source-observer asymmetry associated with the stellar aberration phenomenon at first glance contradicts the principle of relativity.
The principle of relativity which we owe to Poincare, who first coined the term, denied the possibility for an observer partaking in a uniform motion relative to the fixed stars of discovering, by any measurement, such a motion, under the assumption, of course, that one does not look outside.
Now let us to analyze the stellar aberration experiment. Perhaps we should say: ” In the case of stellar aberration, the peculiarity of the aberration of light measurements is that the accelerated (earth-based) observer looks outside to the stars.” But no! The motion of the stars with respect to the earth is never followed by any aberration. The aberration shift (as inferred from astronomical observations) exists even in the case when star moves with the same velocity as the earth. We cannot remain with the framework of ”looking outside to the star” when considering the earth-based observations of the stellar aberration.
In the Section 8.3 we discussed a simple scaling model for the stellar aberration. We obtained a condition for optical similarity between the aberration of light from a distant star and from the earth-based incoherent source. We find that without looking outside, it is possible to determine the speed of the earth around the sun by means of aberration of light measurements. No one has ever done such experiment with earth-based incoherent source, but we know would happen from the astronomical observations.
It should be note that all well known methods to test the special relativity are second order measurements . The cardinal example is given by the Michelson-Morley experiment. A close examination of all second order experiments inside an inertial frame shows that phenomena appear to be independent of the uniform motion relative to the fixed stars. From the series of negative results like that obtained by Michelson and Morley it seems reasonable to suppose that it is not accidental that various optical experiments trying to determine orbital velocity remained unsuccessful. And we should accept the view that there is exist a general law of of nature which prevents us to determine orbital velocity by laboratory experiments. Nevertheless, in this case there does not appear to be any such profound implication. This is specific feature of the optical experiments of second order. All these methods are based on phenomenon of interference. There is a fact governing all interference phenomena. The law is called the relativistic invariance of phase (i.e. it is simply number). It is not change under any reference frame transformations. It is not possible to detect the orbital velocity using second order interference experiments. This is consequence of the geometry of space-time. We conclude the following: The usual formulation of the theory of relativity deals with formulated above principle of irrelevancy of velocity, but it should be understood only in a limiting sense when we are dealing with second order optical measurements.
There is no conflict between the fundamental structure of special theory of relativity on the one hand, and the aberration of light phenomena. First, from what is said in the previous chapters it should be clear that special relativity does not require the formulated above principle of irrelevancy of velocity. Principle of special relativity applies to physical laws, not to physical facts. Not only in Einstein writings, but in every textbook on special theory of relativity can be found the formulation of principle of special relativity as follows: the laws of physics are the same in all inertial frames (1).
The special principle of relativity says that the same laws should hold in all inertial frames. We interpret this to mean that the laws can be expressed by equations that have the same form in all inertial frames. According to special principle of relativity all inertial frames are equivalent with regard to physical laws, not with regard to physical facts. By a physical fact in this context we mean properties of existing objects - in our case of interest the properties of incoherent source. By mean a physical law we mean for example electrodynamics equations. Principle of special relativity is irrelevancy of velocity with regard to physical laws, not with regard to anything.
We point that the essence of the special theory of relativity consists in the following postulate: all physical process proceed in four-dimentional space-time, the geometry of which is pseudo-Euclidean. But give pseudo-Euclidean geometry of space-time does not dictate the irrelevancy of velocity with regard to anything. From a mathematical viewpoint, the argument looks something like the following. The inertial frames are characterized by motion at constant velocity relative to the fixed stars and different inertial frames are related by Lorentz boosts. Relativistic symmetry is usually identified with Lorentz symmetry, i.e. with symmetry under Lorentz boosts. At the same time it is common knowledge that Lorentz boosts alone do not make up a group. Lorentz boosts do not form a group due presence of the Wigner rotation. In general, two successive Lorentz boosts are not equivalent to boost. The product of two Lorentz boosts in different directions is equal to the product of a pure boost and a spatial rotation, the Wigner rotation. We must conclude that when we are dealing with transformations and set of these transformations does not form a group we have no symmetry between inertial frames. There is a real difference between the accelerated inertial frame and the inertial frame without accelerational history. It is comes out quite natural that this difference is closely associated with the Wigner rotation phenomenon.
The existence in nature of something corresponding to the concept of absolute acceleration entails the existence in nature of something corresponding to the concept of absolute velocity. If we look at the situation carefully we see that the time dilation effect problem in the accelerated frame demonstrate the essential asymmetry between accelerated and initial ( i.e. without acceleration history) inertial frame. The proper time of any particle moving arbitrary in the initial inertial frame is always slower than the physical time in this initial inertial frame. It is a fundamental property of the pseudo-Euclidean space-time. There is no relativity of the uniform motions and we can determine the particle absolute velocity in the initial inertial frame. The formalism of classical physics is associated with absolute space structure. This structure must be associated with some ”substrance”, the ether, which incidentally provides an absolute rest frame. So is space-time the new ether? It does have similar properties. It clearly does provide an ”absolute rest reference frame”. The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of unified space-time model, with no a priori splitting of space from time. The initial inertial frame is not physically identical to the absolute space. It is a concept of an absolute space-time.
The relativity principle has a long history in physics. Newtonian mechanics is invariant under Galilean transformations. However, from the mathematical viewpoint, the relativistic Lorentz transformation is qualitatively different from the non-relativistic Galilean one. The set of Galileo boosts form a group. So we see that principle of Galilean relativity states that all inertial frames are equivalent. Equivalence between two frames means that none of them preferred to the other. Galileo’s version of relativity principle was incorrectly extended to the whole of physics. The relativity principle does not hold for the whole range of validity of the Lorentz covariant physical laws and the Lorentz covariance is not a fundamental symmetry of physics (because the accelerated inertial frame is connected with the initial inertial frame by the Galilean transformation, but not the Lorentz one).
12.2 Aberration and the Modern Cosmology
Now we want to discuss a serious trouble - the failure of the aberration of light theory presented above. There are difficulties associated with the ideas of modern cosmology. The difficulty we speak of is associated with the concept of source-observer asymmetry, when applied to the motion relative to the cosmic microwave background (CMB). One may say that perhaps there is no use worrying about these difficulties since there are so many things about universe that we still don’t understand. It is obvious that we live in a universe ruled by the unknown. Indeed, dark energy and dark matter together contribute to almost 96 percents of the total energy-mass of the universe and we have no knowledge of what dark matter (or dark energy) is. It is important to realize that when we try to talk about working assumptions in modern cosmology, they are approximate and will change. Therefore what we discuss in this section will not be accurate in a certain sense.
We cannot go into details of the modern cosmology principles at this time. We shall assume that they are there, and go on to describe what some of the consequences are. According to the cosmological principle, the universe should appear isotropic, without any preferred direction, to a comoving observer, having no peculiar motion relative to the cosmic fluid of the expanding universe. However, a peculiar motion of such an observer might introduce a dipole anisotropy in the observed properties of a class of objects and which in turn, might be exploited to infer the peculiar velocity of the observer. For instance the CMB radiation shows an anisotropy distribution. The CMB dipole is almost ubiquitously assumed to be of kinematical origin, i.e. due to relative motion. It is widely believed that the dipole anisotropy is produced by the Doppler effect due to the relative motion between the earth, i.e. an observer, and the frame where cosmic microwave background looks nearly isotropic. By subtracting the dipole, the CMB is defined as the rest frame for the universe.
The observed dipole indicates that the solar system is moving at 370 km/s relative to the observed universe in the direction of galactic longitude $l=264^{o}$ and latitude $b=48^{O}$. This is quite far from the galactic rotation direction 250 km/s towards $l=90^{o}$ and $b=0$. The motion relative to the cosmic microwave background results from the sum of many components of velocity due to gravitational attraction of various mass concentrations. The existence of clusters and super clusters of galaxies and our motion is a natural consequence of the large scale organization of matter. The peculiar velocity consists of the five vector contributions: the motion of the earth around the sun (30 km/s), the hypothetical circular motion around the our galaxy (250 km/s), the motion of our galaxy in the local group, and the motion of the local group with respect to the cosmic microwave background. The origin of the velocity of the local group is still uncertain and has been under discussion over past two decades. This peculiar velocity is believed to be generated by the spatial inhomogeneties of mass (mainly dark matter) distribution in nearly large scale structures. The velocity of the local group with respect to the cosmic microwave background is estimated to be 500 km/s.
Now let us see how aberration of light various with the speed relative to the rest frame of the universe. Consider first the textbook explanation. It is very easy to understand how the aberration of light effect comes about from the point of view of the conventional aberration of light theory. The existence of an aberration of the line of sight by motion can be recognized by considering an observer in a car driving through rainstorm: raindrops falling vertically appear to be obliquely. An important application of the Galilean velocity transformation law is provided by stellar aberration, the change in the apparent direction of a star caused by the earth’s motion around the sun. The apparent positions of all fixed stars are thus always a little displaced in the direction of the earth’s motion at that moment, and hence describe a small elliptical figure during the annual revolution of the earth around the sun. What about the motion relative to the CMB rest frame? According to textbooks, if the earth motion were uniform, the aberration effect would be undetectable since the ”true” direction of the star is unknown. Indeed, who can say where a given star should be? On the other hand, the true direction of an earth-based source is known. But, according to conventional theory, the aberration of light phenomenon does not exist in an aberration of light experiments using an earth-based light source.
We now in position to understand the nature of our difficulty with the motion relative to the CMB. It is believed that the dipole anisotropy is produced by the Doppler effect due to acceleration (during the billions of years) of the solar system with respect to the rest frame of the universe. This acceleration is believed to be generated by the spatial inhomogenetics of (Dark matter) mass distribution in nearly large scale structures. A correct solution of the aberration problem in the earth-based frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric plays a fundamental role in the non-inertial kinematics of a light ( or relativistic particle ) beam produced by the earth-based source. According to the relativistic theory of aberration presented in this book, the proper rotation of the earth on its axis should produce a corresponding shift of the image. The aberration shift will depend also upon the value of $v_{\perp}^{CMB}$, the component of the solar system velocity (relative to the CMB) perpendicular to the earth rotation axis. The rotation of the earth produces aberration in an amount larger enough to be taken into account in precise observation work using electron microscope as an earth-based particle source. Experimental results show (see the Chapter 8) that the image shift is quite close to the theoretical prediction for the (30 km/s) orbital velocity and clearly indicate that the signal associated with motion (370 km/s) relative to the CMB does not exist. The simplest explanation is that the CMB dipole might be of non-kinematical origin.
More recent observations of astronomers also cast doubt on the CMB dipole, being the ultimate representatives of the solar peculiar velocity. In recent years observations have emerged hinting at an anisotropic universe. The discovery of the preferred direction in the universe was serendipitous. Until very recently the velocity of the solar system in the rest frame of universe is inferred from CMB temperature dipole anisotropy. Obviously, an independent measurements of this velocity is needed to fully establish the kinematical origin of the dipole. Another such quantity that could be employed to look for departures from isotropy is the angular distribution of distant radio sources in the sky. This could provide an independent check on the interpretation of CMB dipole. The radio data clearly indicate that significantly larger dipole exist in the rest frame of the radio galaxies. While the velocity of the solar system inferred from the CMB temperature dipole anisotropy is 370 km/s, the radio dipole measurements finds the speed of motion to be around 1000 km/s (i.e. to be around three times larger that that of CMB). From the Hubble diagram of quasars, motion of the solar system is derived, which out to be the largest value ever found, 8000 km/s (2).
On the other hand, a common direction for all these dipoles, determined from completely independent surveys by different groups employing different techniques, indicate that these dipoles are not resulting from some systematics in the observations or the data analysis, but could instead suggest an inherent anisotropy. This is totally unexpected in a standard model of universe. We have a preferred direction, aligned with the CMB dipole, in the universe. That is, going to the CMB rest frame, we see an anisotropic background. There is a difficulty with such some sort of an ”axis” of the universe which, in turn, would be against the cosmological principle. Three independent dipole vectors pointing along the same particular direction could imply an anisotropic universe, violating the cosmological principle, a cornerstone of the modern cosmology(2).
Let us be conservative and say that there are two kinds of the earth velocity that the total velocity with respect to the initial (i.e. privilege ) inertial frame could be the sum of the orbital velocity and the velocity of the sun. The velocity of the sun consists the two contributions: the circular motion around our galaxy and the motion of our galaxy. In earth-based experiments where we measure the earth velocity with respect to the privilege frame by seeing aberration of light (or particles), we are measuring the recorded (in the accelerational history) velocity. So the velocity with respect to the privilege frame consists of the two contributions: a recorded motion plus an unrecorded motion. We know that there is definitely a recorded (orbital) motion, and we have a formula for it. It is therefore impossible to get all the earth velocity to be record in the accelerational history in the way we hoped. It is not a legal theory if we have nothing but modern cosmology. Something else has to be added.
We would like to think a little more about why the earth is moving with respect to the initial inertial frame. There must have been a force pushing on the earth in order to get it going. So it may help our understanding if we look a little more closely at where the forces come from. We must say immediately that orbital motion is result of non-gravitational forces. It is well known that a full 98 percent of all the angular momentum in the solar system is concentrated in the planets, yet a staggering 99.8 percent of all the mass in our solar system is in our sun. Perhaps the first scientifically respectable theory of the origin of the solar system was given by Hoyle (1960). He invoked the action of magnetic field to transfer angular momentum from the central body, the sun, to the ejected matter which eventually formed the planets. In contrast, the unrecorded motion around our galactic and the motion of our galactic is result of the action of gravity.
We started by talking about the gravitational interaction of the earth with the sun. We believe the theory of gravity is so much that we allow it to tell us about the force of one galactic to another. Perhaps we are making to great an extrapolation of our limiting knowledge of gravity to galactic scales. Perhaps the entire difficulty is that a modification of gravity may be responsible for unrecorded velocities. We may someday find out one things we don’t understand today, for example dark matter, can, in fact be explained as modification of the gravitation theory. Today it doesn’t seem likely, but no one can say for sure. There are so many things about universe that we don’t understand.
12.3 Bibliography and Notes
1. By ”the same” is means form-invariance: the mathematical relations between the physical quantities remain identical after a coordinate transformation has been performed on these quantities. In special theory of relativity the (form-invariant) transformations between inertial frames are Lorentz-transformations [48].
2. Recent observations of the cosmic dipole anisotropies is revolutionized our understanding of the universe. Attempts to recover the CMB dipole from counts of later universe sources such as radio galaxies and quasars, which are assumed to be in the CMB frame, largely agree that CMB dipole direction is recovered, but not the magnitude. The various dipoles, included CMB dipole, all pointing along the same direction, suggest a preferred direction in the universe, raising thereby uncomfortable questions about the cosmological principle, the basis of the standard model in modern cosmology [49, 50, 51]. .
13 Relativistic Particle Dynamics
In previous chapters we considered the kinematics of the theory of relativity, which concerns the study of the four vectors of positions, velocity and acceleration. Kinematics studies trajectories as geometrical objects, independently of their causes. This means that it is not possible to predict the trajectory of a particle evolving under a given dynamical field using just a kinematic treatment. In dynamics we consider the effect of interaction on motion.
13.1 Manifestly Covariant Particle Dynamics
Dynamics equations can be expressed as tensor equations in Minkowski space-time. When coordinates are chosen, one may work with components, instead of geometric objects. Relying on the geometric structure of Minkowski space-time, one can define the class of inertial frames and can adopt a Lorentz frame with orthonormal basis vectors for any given inertial frame. In any Lorentz coordinate system the law of motion becomes
$m\frac{d^{2}x_{\mu}}{d\tau^{2}}=eF^{\mu\nu}\frac{dx_{\nu}}{d\tau}~{},$ (20)
where here the particle’s mass and charge are denoted by $m$ and $e$ respectively. The electromagnetic field is described by a second-rank, antisymmetric tensor with components $F^{\mu\nu}$. The coordinate-independent proper time $\tau$ is a parameter describing the evolution of physical system under the relativistic laws of motion, Eq. (20).
The covariant equation of motion for a relativistic charged particle under the action of the four-force $K_{\mu}=eF^{\mu\nu}dx_{\nu}/d\tau$ in the Lorentz lab frame, Eq.(20), is a relativistic ”generalization” of the Newton’s second law. The three-dimensional Newton second law $md\vec{v}/dt=\vec{f}$ can always be used in the instantaneous Lorentz comoving frame. Relativistic ”generalization” means that the previous three independent equations expressing Newton second law are be embedded into the four-dimensional Minkowski space-time (1).
The immediate generalization of $md\vec{v}/dt=\vec{f}$ to an arbitrary Lorentz frame is Eq.(20), as can be checked by reducing to the rest frame. In Lorentz coordinates there is a kinematics constraint $u^{\mu}u_{\mu}=c^{2}$ for the four-velocity $u_{\mu}=dx_{\mu}/d\tau$. Because of this constraint, the four-dimensional dynamics law, Eq.(20), actually includes only three independent equations of motion. Using explicit expression for Lorentz force we find that the four equations Eq.(20) automatically imply the constraint $u^{\mu}u_{\mu}=c^{2}$ as it must be. To prove this, we calculate the scalar product between both sides of the equation of motion and $u_{\mu}$. Using the fact that $F^{\mu\nu}$ is antisymmetric (i.e. $F^{\mu\nu}=-F^{\nu\mu}$), we find $u_{\mu}du^{\mu}/d\tau=eF^{\mu\nu}u_{\mu}u_{\nu}=0$. Thus, for the quantity $Y=(u^{2}-c^{2})$ we find $dY/d\tau=0$.
13.2 Conventional Particle Tracking. Hidden Absolute Time Coordinatization
Having written down the motion equation in a 4-vector form, Eq.(20), and determined the components of the 4-force, we satisfied the principle of relativity for one thing, and, for another, we obtained the four components of the equation of particle motion. This is covariant relativistic generalization of the three dimensional Newton’s equation of motion which is based on particle proper time as the evolution parameter.
We next wish to describe a particle motion in the Lorentz lab frame using the lab time $t$ as evolution parameter. First, we examine the reasoning presented in textbooks. Let us determine the first three spatial components of the 4-force. We consider for this the spatial part of the dynamics equation, Eq.(20): $\vec{Q}=(dt/d\tau)d(m\gamma\vec{v})/dt=\gamma d(m\gamma\vec{v})/dt$. The prefactor $\gamma$ arises from the change of the evolution variable from the proper time $\tau$, which is natural since $\vec{Q}$ is the space part of a four-vector, to the lab frame time $t$, which is needed to introduce the usual force three-vector $\vec{f}$: $\vec{Q}=\gamma\vec{f}$. Written explicitly, the relativistic form of the three-force is
$\frac{d}{dt}\left(\frac{m\vec{v}}{\sqrt{1-v^{2}/c^{2}}}\right)=e\left(\vec{E}+\frac{\vec{v}}{c}\times\vec{B}\right)~{}.$ (21)
The time component is
$\frac{d}{dt}\left(\frac{mc^{2}}{\sqrt{1-v^{2}/c^{2}}}\right)=e\vec{E}\cdot\vec{v}~{}.$ (22)
The evolution of the particle is subject to these four equations, but also to the constraint
$\mathcal{E}^{2}/c^{2}-|\vec{p}|^{2}=mc^{2}~{}.$ (23)
According to the non-covariant (3+1) approach we seek for the initial value solution to these equations. Using explicit expression for Lorentz force we find that the three equations Eq.(21) automatically imply the constraint Eq.(23), once this is satisfied initially at $t=0$. In the (3+1) approach, the four equations of motion ”split up” into (3+1) equations and we have no mixture of space and time parts of the dynamics equation Eq.(20). This approach to relativistic particle dynamics relies on the use of three independent equations of motion Eq.(21) for three independent coordinates and velocities, ”independent” meaning that equation Eq.(22) (and constraint Eq.(23)) are automatically satisfied.
One could expect that the particle’s trajectory in the lab frame, following from the previous reasoning $\vec{x}(t)$, should be identified with $\vec{x}_{cov}(t)$. However, paradoxical result are obtained by doing so. In particular, the trajectory $\vec{x}(t)$ does not include relativistic kinematics effects. In the non-covariant (3+1) approach, the solution of the dynamics problem in the lab frame makes no reference to Lorentz transformations. This means that, for instance, within the lab frame the motion of particles in constant magnetic field looks precisely the same as predicted by Newtonian kinematics: relativistic effects do not have a place in this description. In conventional particle tracking a particle trajectory $\vec{x}(t)$ can be seen from the lab frame as the result of successive Galileo boosts that track the motion of the accelerated (in a constant magnetic field) particle. The usual Galileo rule for addition of velocities is used to determine the Galileo boosts tracking a particular particle, instant after instant, along its motion along the curved trajectory.
The old kinematics is especially surprising, because we are based on the use of the covariant approach. Where does it comes from? The previous commonly accepted derivation of the equations for the particle motion in the three dimensional space from the covariant equation Eq.(20) includes one delicate point. In Eq.(21) and Eq.(22) the restriction $\vec{p}=m\vec{v}/\sqrt{1-v^{2}/c^{2}}$ has already been imposed. One might well wonder why, because in the accepted covariant approach, the solution of the dynamics problem for the momentum in the lab frame makes no reference to the three-dimensional velocity. In fact, equation Eq.(20) tells us that the force is the rate of change of the momentum $\vec{p}$, but does not tell us how momentum varies with speed. The four-velocity cannot be decomposed into $u=(c\gamma,\vec{v}\gamma)$ when we deal with a particle accelerating along a curved trajectory in the Lorentz lab frame.
Actually, the decomposition $u=(c\gamma,\vec{v}\gamma)$ comes from the relation $u_{\mu}=dx_{\mu}/d\tau=\gamma dx_{\mu}/dt=(c\gamma,\vec{v}\gamma)$. In other words, the presentation of the time component as the relation $d\tau=dt/\gamma$ between the $\tau$ and lab time $t$ is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
It is important to stress in this point that the situation when only one clock in the comoving frame is involved in dynamics cannot be realized. The Newton law can be written down in the proper frame only when a space-time coordinate system $(x^{\prime},y^{\prime},z^{\prime},\tau)$ has been specified. The type of clock synchronization which provides the time coordinate in the Newton equation has never been discussed in textbooks. It is clear that without an answer to the question about the method of distant clock synchronization used, not only the concept of acceleration, but also the dynamics law has no physical meaning. A proper frame approach to relativistic particle dynamics is forcefully based on a definite (Einstein) coordinatization assumption. After this, the dynamics theory states that the equation of motion in the proper frame is $md^{2}\vec{x^{\prime}}/d\tau^{2}=\vec{f}^{\prime}$. It should be stressed that in the case of velocity increment $\Delta\vec{v^{\prime}}=\Delta\vec{x^{\prime}}/\Delta\tau$ we also deal with the distant events. It can be said with some abuse of language that 3-velocity vector is always a ”spatially extended” object. Let us return to our consideration of the relation $d\tau=dt/\gamma$ between the $\tau$ and lab time $t$. The calculation carried out in the case of an spatially extended object shows that the temporal coincidence of two distant events has absolute character: $\Delta\tau=0$ implies $\Delta t=0$.
It should be note that usually the notation ”$\tau$” arose from the proper time. Now let us recall the standard concept of the object’s proper time. Let a point-like object move uniformly and rectilinearly relative to the lab frame $K$. The proper frame $K^{\prime}$ can be fixed to the moving object. The object is at rest in this frame, so that events happening with this object are registered by one clock. This clock counts the proper time at the point where the object is located. As we already know, the proper time on moving object flows slower than the time $t$. This phenomena was called time dilation. The proper time can also be introduced for a point particle moving with acceleration. This standard interpretation of the proper time $\tau$ has evidently nothing to do with dynamics and one may wonder where this contradiction existing between the name and the content of time in dynamic law come from.
In order to avoid being to abstract for to long we have given one example. Let us discuss the case in which the velocity increment is not in the direction of the uniform motion. For example, there may be a particle which is just moving ”upward” with velocity increment $\Delta y^{\prime}/\Delta\tau$ and the lab system is moving ”horizontally”. Now, according to textbooks, we simply using relation $d\tau=dt/\gamma$ with the standard result $\Delta v_{y}=\Delta v^{\prime}_{y}/\gamma$. However, paradoxical result are obtained by doing so. In particular, the trajectory of the particle does not include mixture of positions and time. One of the important conclusions of the discussion presented above in the Chapter 7 is that there is no absolute notion of simultaneity in the theory of special relativity. Therefore there is no notion of an instantaneous 3-space. A coordinate axes of reference frame $K^{\prime}$ rotating about the axes of a Lorentz lab frame. The standard presentation of the velocity increment transformation as the relation $\Delta v_{y}=\Delta v^{\prime}_{y}/\gamma$ is based on the hidden assumption that $y^{\prime}$-axis and $y$-axis are parallel. On the other hand Lorentz transformation makes axes ($y^{\prime}$ and $y$) oblique-angled due to Wigner rotation. A difference between non-covariant and covariant point particle trajectories follows from this immediately (We will discuss the transformation of velocities in greater detail in Section 13.5).
13.3 Incorrect Expansion of the Relation $d\tau=dt/\gamma$ to an Arbitrary Motion
Authors of textbooks are dramatically mistaken in their belief about the usual momentum-velocity relation. From the theory of relativity follows that the equation $\vec{p}_{cov}=m\vec{v}_{cov}/\sqrt{1-v^{2}_{cov}/c^{2}}$ does not hold for a curved trajectory in the Lorentz lab frame. Many experts who learned the theory of relativity using textbooks will find this statement disturbing at first sight.
How can such an unusual momentum-velocity relation come about? We know that the components of momentum four-vector $p_{\mu}=(E/c,\vec{p})$ behave under transformations from one Lorentz frame to another, exactly in the same manner as the component of the four-vector event $x=(x_{0},\vec{x})$. Surprises can surely be expected when we return from the four-vectors language to the three-dimensional velocity vector $\vec{v}$, which can be represented in terms of the components of four-vector as $\vec{v}=d\vec{x}/dx_{0}$. In contrast with the pseudo-Euclidean four-velocity space, the relativistic three-velocity space is a three-dimensional space with constant negative curvature, i.e. three-dimensional space with Lobachevsky geometry.
It is well known that for rectilinear accelerated motion the usual momentum-velocity relation holds. In fact, for the rectilinear motion the combination of the usual momentum-velocity relation and the covariant three-velocity transformation (according to Einstein’s law of velocity addition) is consistent with the covariant three-momentum transformation and both (non-covariant and covariant) approaches produce the same trajectory.
We can see why by examine the transformation of the three velocity in the theory of relativity. For a rectilinear motion, this transformation is performed as $v=(v^{\prime}+V)/(1+v^{\prime}V/c^{2})$. The relativistic factor $1/\sqrt{1-v^{2}/c^{2}}$ is given by: $1/\sqrt{1-v^{2}/c^{2}}=(1+v^{\prime}V/c^{2})/(\sqrt{1-v^{\prime 2}/c^{2}}\sqrt{1-V^{2}/c^{2}})$. The new momentum is then simply $mv$ times the above expression. But we want to express the new momentum in terms of the primed momentum and energy, and we note that $p=(p^{\prime}+\mathcal{E}^{\prime}V/c^{2})/\sqrt{1-V^{2}/c^{2}}$. Thus, for a rectilinear motion, the combination of Einstein addition law for parallel velocities and the usual momentum-velocity relation is consistent with the covariant momentum transformation. The dynamical evolution of the particle according to the non-covariant particle tracking may be considered as the result of successive infinitesimal Lorentz transformations. The corrected Newton law is valid for each step. In other words, the two observers with rectilinear trajectories have common 3-space. This result was incorrectly extended to an arbitrary trajectory.
We already know that the collinear Lorentz boosts (similar to the collinear Galilean boosts) form a commutative group. This means that the resultant of successive collinear Lorentz boosts is independent of the transformation order. On the contrary, Lorentz boosts in different directions do not commute and do not make up a group. In contrast, the set of Galilean boosts always form a group. As well-known, the composition of non-collinear Lorentz boosts is equivalent to a Lorentz boost followed by a spatial rotation. This Wigner rotation is relativistic effect that does not have a non-covariant analogue. One of the consequences of non-commutativity of non-collinear Lorentz boosts is the unusual momentum-velocity relation $\vec{p}_{cov}\neq m\vec{v}_{cov}/\sqrt{1-v^{2}_{cov}/c^{2}}$, which also does not have any non-covariant analogue.
The theory of relativity shows that the unusual momentum-velocity relation discussed above is related with the acceleration along curved trajectories. In this case there is no notion of a common ordinary space and there is a difference between covariant and non-covariant particle trajectories. Only the solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle trajectories in the lab frame. A closer analysis of the concept of velocity, i.e. a discussion of the methods by which a time coordinate can actually be assigned in the lab frame, opens up the possibility of a description of such physical phenomena as radiation from a relativistic electron accelerating along a curved trajectory in accordance with the theory of relativity.
13.4 The Usual Integration of the 4D Covariant Equation of Motion
Attempts to solve the dynamics equation Eq.(20) in manifestly covariant form can be found in the literature. The trajectory which is found does not include relativistic kinematics effects. Therefore, it cannot be identified with $\vec{x}_{cov}(t)$ even if, at first glance, it appears to be derived following covariant prescription.
First, we analyze textbooks presentation of the integration of the covariant equation of motion. Consider, for example, the motion of a particle in a given electromagnetic field. The simplest case, of great practical importance, is that of an uniform electromagnetic field meaning that $F^{\mu\nu}$ is constant on the whole space-time region of interest. In particular we consider the motion of a particle in a constant homogeneous magnetic field, specified by tensor components $F^{\mu\nu}=B(e^{\mu}_{2}e^{\nu}_{3}-e^{\nu}_{2}e^{\mu}_{3})$ where $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are orthonormal space like basis vectors $e^{2}_{2}=e^{2}_{3}=-1$, $e_{2}\cdot e_{3}=0$. In the lab frame of reference where $e^{\mu}_{0}$ is taken as the time axis, and $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are space vectors the field is indeed purely magnetic, of magnitude $B$ and parallel to the $e_{1}$ axis. Let us set the initial four-velocity $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$, where $v$ is the initial particle’s velocity relative to the lab observer along the axis $e_{2}$ at the instant $\tau=0$, and $\gamma=1/\sqrt{1-v^{2}/c^{2}}$. The components of the equation of motion are then $du^{(0)}/d\tau=0$, $du^{(1)}/d\tau=0$, $du^{(2)}/d\tau=-eBu^{(3)}/(mc)$, $du^{(3)}/d\tau=eBu^{(2)}/(mc)$. We seek for the initial value solution to these equations as done in the existing literature (2). A distinctive feature of the initial value problem in relativistic mechanics, is that the dynamics is always constrained. In fact, the evolution of the particle is subject to $mdu^{\mu}/d\tau=eF^{\mu\nu}u_{\nu}$, but also to the constraint $u^{2}=c^{2}$. However, such a condition can be weakened requiring its validity at certain values of $\tau$ only, let us say initially, at $\tau=0$. Therefore, if $Y(\tau)$ vanishes initially, i.e. $Y(0)=0$, then $Y(\tau)=0$ at any $\tau$. In other words, the differential Lorentz-force equation implies the constraint $u^{2}=c^{2}$ once this is satisfied initially. Integrating with respect to the proper time we have $u^{\mu}(\tau)=\gamma e^{\mu}_{0}+\gamma v[e^{\mu}_{2}\cos(\omega\tau)+e^{\mu}_{3}\sin(\omega\tau)]$ where $\omega=eB/(mc)$. We see that $\gamma$ is constant with time, meaning that the energy of a charged particle moving in a constant magnetic field is constant. After two successive integrations we have $X^{\mu}(\tau)=X^{\mu}(0)+\gamma c\tau e^{\mu}_{0}+R[e^{\mu}_{2}\sin(\omega\tau)-e^{\mu}_{3}\cos(\omega\tau)]$ where $R=\gamma v/\omega$. This enables us to find the time dependence $[0,X^{(2)}(t),X^{(3)}(t)]$ of the particle’s position since $t/\gamma=\tau$. From this solution of the equation of motion we conclude that the motion of a charged particle in a constant magnetic field is a uniform circular motion.
One could expect that the particle’s trajectory in the lab frame, following from the previous reasoning $[0,X^{(2)}(t),X^{(3)}(t)]$, should be identified with $\vec{x}_{cov}(t)$. However, paradoxical result are obtained by doing so. In particular, the trajectory $[0,X^{(2)}(t),X^{(3)}(t)]$ does not include relativistic kinematics effects. We found that the usual integration of the four-dimensional covariant equation of motion Eq.(20) gives particle trajectory which looks precisely the same as in Newton dynamics and kinematics. The trajectory of the electron does not include relativistic effects and the Galilean vectorial law of addition of velocities is actually used.
So we must have made a mistake. Notice that we are using textbooks to solve covariant equation of motion. We did not make a computational mistake in our integrations, but rather a conceptual one. We must say immediately that there is no objection to the first integration of Eq.(20) from initial conditions over proper time $\tau$. With this, we find the four-momentum. The momentum has exact objective meaning i.e. it is convention-invariant. What must be recognized is that the concept of velocity is only introduced in the second integration step. However, in accepted covariant approach, the solution of the dynamics problem for the momentum in the lab frame makes no reference to three-dimensional velocity. In fact, the initial condition which we used is $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$ and includes $\gamma c$ and $\gamma v$, which are actually notations for the time and space parts of the initial four-momentum. The three-dimensional trajectory and respectively velocity, which are convention-dependent, are only found after the second integration step. Then, where does the old kinematics comes from? The second integration was performed using the relation $d\tau=dt/\gamma$. It is only after we have made those replacement for $d\tau$ that we obtain the usual formula for conventional (non-covariant) trajectory for an electron in a constant magnetic field.
We should then expect to get results similar to those obtained in the case of the (3+1) non-covariant particle tracking. In fact, based on the structure of the four components of the equation of motion Eq.(20), we can arrive to another mathematically identical formulation of the dynamical problem. The fact that the evolution of the particle in the lab frame is subject to a constraint has already been mentioned. This means that the mathematical form of the dynamics law includes only three independent equations of motion. It is easy to see from the initial set of four equations, $du^{(0)}/d\tau=0$, $du^{(1)}/d\tau=0$, $du^{(2)}/d\tau=-eBu^{(3)}/(mc)$, $du^{(3)}/d\tau=eBu^{(2)}/(mc)$, that the presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is just a simple parametrization that yields the corrected Newton’s equation Eq.(21) as another equivalent form of these four equations in terms of absolute time $t$ instead of proper time of the particle. This approach to integrating dynamics equations from the initial conditions relies on the use of three independent spatial coordinates and velocities without constraint and is intimately connected with old kinematics. The presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
13.5 Covariant Particle Tracking
In conventional particle tracking a particle trajectory $\vec{x}(t)$ can be seen from the lab frame as the result of successive Galileo boosts that track the motion of the accelerated (in a constant magnetic field) particle. The usual Galileo rule for addition of velocities is used to determine the Galileo boosts tracking a particular particle, instant after instant, along its motion along the curved trajectory.
We cannot take old Newton’s kinematics for mechanics and Einstein’s kinematics for electrodynamics. Only solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle traectories in the lab frame. The trajectory $\vec{x}_{cov}(t)$ is viewed from the lab frame as the result of successive Lorentz transformations. In this case relativistic kinematics effects arise. Let us try out our algorithm for reconstructing $\vec{x}_{cov}(t)$ on some example, to see how it works. We will find formulas relating the velocity of a particle in one reference system to its velocity in a second reference system. First, we examine the reasoning presented in textbooks. Let us suppose that the $K^{\prime}$ system moves relative to the $K$ system with velocity $V$ along the $x$ axis. Let $\vec{v}=d\vec{x}/dt,$ be the vector of the particle velocity in the $K$ system and $\vec{v^{\prime}}=d\vec{x^{\prime}}/dt^{\prime}$ the velocity vector of the same particle in the $K^{\prime}$ system. From Lorentz transformation we have $dx=\gamma(dx^{\prime}+Vdt^{\prime})$, $dy=dy^{\prime}$, $dz=dz^{\prime}$, $dt=\gamma(dt^{\prime}+Vdx^{\prime}/c^{2})$, where $\gamma=1/\sqrt{1-V^{2}/c^{2}}$. Dividing the first three equations by the fourth we find $v_{x}=(v_{x}^{\prime}+V)/(1+Vv^{\prime}_{x}/c^{2})$, $v_{y}=v^{\prime}_{y}/[\gamma(1+Vv^{\prime}_{x}/c^{2})]$, $v_{z}=v^{\prime}_{z}/[\gamma(1+Vv^{\prime}_{x}/c^{2})]$. These formulae determine the transformation of velocities. According to textbooks, they describe the law of composition of velocities in the theory of relativity.
We begin with the simple case of relativistically combining perpendicular velocities. There is no length contraction for perpendicular distances, but time dilation still occur. Therefore the particle velocity in the frame $K$ is just $v_{x}=V$, $v_{y}=v_{y}^{\prime}/\gamma$. Note that we must be clear what we mean by ”perpendicular”. It only makes sense to say two velocities are perpendicular if they are measured in the same reference frame. Hence when authors of textbooks say that $\vec{e}_{x}^{\prime}v_{x}^{\prime}$ and $\vec{e}_{x}V$ are perpendicular they actually mean that $(x^{\prime},y^{\prime},z^{\prime})$ axes and $(x,y,z)$ axes are parallel. In other words, it is (hidden) assumed that $K$ and $K^{\prime}$ observers have common 3-space. This is misconception.
One of the consequences of non-commutativity of non-collinear Lorentz boosts is not the existence of a common ordinary space. Suppose that in the Lorentz frame $K^{\prime}$ the traveler was observing particle motion along the $y^{\prime}$ axis. In other words measured velocity of the particle $v^{\prime}_{y}$ , and yet the frame is moving relative to the frame $K$ along $x$ axis. Then to an observer at rest in the Lorentz frame $K$, the frame $K^{\prime}$ with the moving particle seems to rotate about the frame $K$. Let us consider the simple case when the velocity $V$ approaches to that of light and the velocity $v^{\prime}_{y}$ is small. This means that we take limit $\gamma\gg 1\gg v^{\prime}_{y}/c$. We consider the small parameter $v^{\prime}_{y}/c$ and use the second order approximation. For ultrarelativistic limit $\gamma\gg 1$ and the axes of $K^{\prime}$ frame are then rotated with respect to the $K$ frame axes by angle $\theta_{w}=v^{\prime}_{y}/(\gamma c)$. We want to look now at the consequences of this rotation. First we notice that there is a projection of the velocity $v^{\prime}_{y}$ on the $x$-axis, so the velocity component along the $x$-axis in the frame $K$ is going to be smaller than $V$. Our problem now is to work out the decrement of the horizontal velocity in the $K$ frame. How we can do that? First, the relativistic correction to the horizontal velocity appears only in the order $[v^{\prime}_{y}/(\gamma c)]^{2}$. The trajectory of the particle is viewed from the $K$ frame as the result of successive infinitesimal Lorentz transformations. Integrating with respect to infinitesimal value $dv^{\prime}_{y}$, we find that the horizontal velocity is $v_{x}=V-(v^{\prime}_{y})^{2}/(2\gamma^{2}c)$. In the ultrarelativistic approximation we find the simple result $v_{y}=v^{\prime}_{y}/\gamma$, $v_{x}=V-v_{y}^{2}/(2c)$, so that Lorentz transformations simply leads to a rotation of particle velocity $V$ of the angle $v_{y}/c=v_{y}/V$. In contrast, according to textbooks, it follows that the total particle speed in the Lorentz $K$ frame increase from $V$ to $V+v_{y}^{2}/(2c)$. Our result is at odds with the prediction from textbooks. The commonly accepted derivation of the composition of velocities based on the use the relation $dt^{\prime}=dt/\gamma$ and does not account for the mixture of (transverse) positions and time.
Now we go on to consider another situation, a very practical one. We want to study the case of relativistically combining perpendicular velocities where velocity $v_{y}^{\prime}$ approaches to that of light, $v_{y}^{\prime}\to c$, and the velocity $V$ is small, $V/c\ll 1$. According to textbooks, the velocity addition formula gives $v_{x}\to V$, $v_{y}=v_{y}^{\prime}/\gamma\to v_{y}^{\prime}-V^{2}/(2c)$. So that two Lorentz boosts simply leads to a rotation of particle velocity $v_{y}^{\prime}$ of the angle $V/c$. We can believed that it is right; it is example what called aberration of (ultra) relativistic particles and it is well known. How can that be? Perhaps it is not puzzling, through, when we remember that what we called inertial frame $K^{\prime}$ is really accelerated frame with respect to the inertial lab frame $K$. It is important to stress in this point that we consider the motion of a particle in the lab frame and we want to describe how to determine the covariant trajectory $\vec{x}(t)$. Above in the Chapter 7 we demonstrated that the aberration of particles in the accelerated frame $K^{\prime}$ can be described in the lab frame within standard special relativity taking advantage of the Wigner rotation theory. As viewed from the lab frame, the coordinate axes of the accelerated frame is rotated through the angle $V/c$ with respect to the coordinate axes of the lab frame. We can easily understand that the particle motion in the $K^{\prime}$ frame will be parallel to the $y$ coordinate axis of the lab frame. It is clear that we should then expect to get results similar to those obtained in the textbooks. When we studied asymptotic $V\to c$, $v_{y}^{\prime}/c\ll 1$, we saw that when particle is moving nonrelativistically in the frame $K^{\prime}$ (i.e. when the aberration effect is negligible), a particle motion will be parallel to the coordinate axis $y^{\prime}$ and we have no problem of demonstrating that there is the connection between the velocity transformation and the Wigner rotation.
It is hoped that presented example will draw wide attention to the central role that the Wigner rotation plays in the transformation of non-collinear velocities. To emphasize the physical concepts involved, rather than mathematical formalism, we consider not the most general case of the (corrected) velocity addition theorem. However, the special case considered already involves all necessary ingredients.
13.6 On the Relativistic Velocity Composition and the Wigner Rotation
There are two difficulties in solving problems concerning covariant particle tracking. First of all, conventional particle tracking in the single inertial frame is actually based on classical Newton mechanics without reference to Lorentz transformations. According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity. In other words, in this standard approach, introducing as only modification to the classical case the relativistic mass, time differ from space. It fact, we have no mixture positions and time. This essential point has never received attention in the physical community.
Another difficulty is the confusion surrounding the combination of non-parallel velocities in special relativity. Let us to consider composition of two velocities when they are not lying on a straight line. In conventional particle tracking (similar to the Newton mechanics), one composes them geometrically by the method of parallelogram with velocities regarded as the usual vectors. In relativistic kinematics, the composition of the velocities is studied analytically from the composition of Lorentz boosts. At the same time it is common knowledge that Lorentz boosts alone do not make a group. We must conclude that when we are dealing with transformations and set of these transformations does not form a group we have no legal theory. Authors of textbooks believe that the relativity principle requires that Einstein’s velocity addition law gives the composition law of a group. The non-associativity of the composition law of the non-colinear velocities is not widely known. Many experts who learned the theory of relativity using textbooks will find this statement disturbing. The popular belief remains that Einstein’s law of addition of velocities is associative, although it is an exercise in elementary algebra to prove non-associativity (3). It is therefore impossible to get covariant trajectory, $\vec{x}_{cov}(t)$, if we have nothing but Lorentz boosts. Something else has to be added. Lorentz boosts do not form a group due to presence of the Wigner rotation. It is comes out quite natural that breakdown of commutativity and associativity in Einstein addition, turn out to be Wigner rotation.
13.7 Convention-Invariant Particle Tracking
So far we have considered the motion of a particle in three-dimensional space using the vector-valued function $\vec{x}(t)$. We have a prescribed curve (path) along which the particle moves. The motion along the path is described by $l(t)$, where $l$ is a certain parameter (in our case of interest the length of the arc). Note the difference between the notions of path and trajectory. The trajectory of a particle conveys more information about its motion because every position is described additionally by the corresponding time instant. The path is rather a purely geometrical notion. Complete paths or their parts may consist of, e. g., line segments, arcs, circles, helical curves. If we take the origin of the (Cartesian) coordinate system and we connect the point to the point laying on the path and describing the motion of the particle, then the creating vector will be a position vector $\vec{x}(l)$. The derivative of a vector is the vector tangent to the curve described by the radius vector $\vec{x}(l)$. The sense of the $d\vec{x}(l)/dl$ is determined by the sense of the curve arc $l$.
We already know from our discussion in Introduction that the path $\vec{x}(l)$ has exact objective meaning i.e. it is convention-invariant. The components of the momentum four vector $mu=(\mathcal{E}/c,\vec{p})$ have also exact objective meaning. In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle in the lab frame is convention dependent and has no exact objective meaning.
We want now to describe how to determine the position vector $\vec{x}(l)_{cov}$ in covariant particle tracking. We consider the motion in a uniform magnetic field with zero electric field. Using the Eq.(20) we obtain
$\frac{d\vec{p}}{d\tau}=\frac{e}{mc}~{}\vec{p}\times\vec{B},~{}~{}\frac{d\mathcal{E}}{d\tau}=0~{}~{}.$ (24)
From $d\mathcal{E}/d\tau=0$ and from the constraint $\mathcal{E}^{2}/c^{2}-|\vec{p}|^{2}=mc^{2}$ we have $dp/d\tau=0$, where $p=|\vec{p}|=m|d\vec{x}_{cov}|/d\tau$. The unit vector $\vec{p}/p$ can be described by the equation $\vec{p}/p=d\vec{x}_{cov}/|d\vec{x}_{cov}|=d\vec{x}_{cov}/dl$, where $|d\vec{x}_{cov}|=dl$ is the differential of the path length. From the foregoing consideration follows that
$\frac{d^{2}\vec{x}_{cov}}{dl^{2}}=\frac{d\vec{x}_{cov}}{dl}\times\left(\frac{e\vec{B}}{pc}\right)~{}.$ (25)
These three equations corresponds exactly to the equations for the components of the position vector that can be found using the non-covariant particle tracking approach, and $\vec{x}(l)_{cov}$ is exactly equal to $\vec{x}(l)$ as it must be. The point is that both approaches describe correctly the same physical reality and since the curvature radius of the path in the magnetic field, and consequently the three-momentum, has obviously an objective meaning (i.e. is convention-invariant), both approaches yield the same physical results.
13.8 The Nature of Mass
This section concentrates on the nature of mass. The concept of mass is one of the most fundamental notions in physics, comparable in importance only to concept of space and time. With advent of the special theory of relativity, physicists and philosophers focused their attention on the concept of relativistic mass. The remarkable progress in experimental and theoretical physics made during the past few decades has considerably deepened our knowledge concerning the nature of mass. We conclude this section with a brief discussion of the physical meaning of the Minkowski geometry.
13.8.1 Phenomenology and Relativistic Extensions
In order fully to understand the meaning of the embedding of the Newton’s dynamics law in the Minkowski space-time, one must keep in mind that, above, we characterized Newton’s equation in the Lorentz comoving frame as a phenomenological law. The microscopic interpretation of the inertial mass of a particle is not given. In other words, it is generally accepted that Newton’s second law is a phenomenological law and the rest mass is introduced in an ad hoc manner. The system of coordinates in which the equations of Newton’s mechanics are valid can be defined as Lorentz rest frame. The relativistic generalization of the Newton’s second law to any Lorentz frame permits us to make correct predictions.
We are in the position to formulate the following general statement: any phenomenological law, which is valid in the Lorentz rest frame, can be embedded in the four dimensional space-time only by using Lorentz coordinatization (i.e. Einstein synchronization convention). Suppose we do not know why a muon disintegrates, but we know the law of decay in the Lorentz rest frame. This law would then be a phenomenological law. The relativistic generalization of this law to any Lorentz frame allows us to make a prediction on the average distance traveled by the muon. In contrast, in the non covariant (3+1) space and time approach there is no time dilation effect, since for Galilean transformations the time scales do not change. Therefore, in the (3+1) non covariant approach, there is no kinematics correction factor $\gamma$ to the travel distance of relativistically moving muons. The two approaches give, in fact, a different result for the travel-distance, which must be, however, convention-invariant. This glaring conflict between results of covariant and non covariant approaches can be explained as follows: it is a dynamical line of arguments that explains this paradoxical situation with the relativistic $\gamma$ factor. In fact, there is a machinery behind the muon disintegration. Its origin is explained in the framework of the Lorentz-covariant quantum field theory. In the microscopic approach to muon disintegration, Einstein and absolute time synchronization conventions give the same result for such convention-invariant observables like the average travel distance, and it does not matter which transformation (Galilean or Lorentz) is used.
13.8.2 The Relativistic Mass
In the non covariant (3+1) space and time approach, there is no time dilation nor length contraction, because for Galilean transformations time and spatial coordinates scales do not change. Moreover, it can easily be verified that Newton’s second law keeps its form under Galilean transformations. Therefore, in the (3+1) non covariant approach, there is no kinematics correction factor $\gamma$ to the mass in Newton’s second law. However, in contrast to kinematics effects like time dilation and length contraction, the correction factor $\gamma$ to the mass in the Newton’s second law has direct objective meaning. In fact, if we assign space-time coordinates to the lab frame using the absolute time convention, the equation of motion is still given by Newton’s second law corrected for the relativistic dependence of momentum on velocity even though, as just stated, it has no kinematical origin. Understanding this result of the theory of relativity is similar to understanding previously discussed results: at first we use Lorentz coordinates and later the (3+1) non covariant approach in terms of a microscopic interpretation that must be consistent with the principle of relativity.
It is well-known from classical electrodynamics that the electromagnetic field of an electron carries a momentum proportional to its velocity for $v\ll c$, while for an arbitrary velocity $v$, the momentum is altered by the relativistic $\gamma$ factor in the case when the absolute time convention is used. Many attempts have been made to explain the electron mass as fully originating from electromagnetic fields. However, these attempts have failed. In fact, it is impossible to have a stationary non-neutral charge distribution held together by purely electromagnetic forces. In other words, mass and momentum of an electron cannot be completely electromagnetic in origin and in order to grant stability there is a necessity for compensating electromagnetic forces with non electromagnetic fields. From this viewpoint, Newton’s second law is an empirical phenomenological law where the relativistic correction factor $\gamma$ to the mass is introduced in an ad hoc manner.
From a microscopic viewpoint, today accepted explanation of how structureless particles like leptons and quarks acquire mass is based on the coupling to the Higgs field, the Higgs boson having been recently experimentally observed at the LHC. This mechanism can be invoked to explain Newton’s second law from a microscopic viewpoint even for structureless particles like electrons. However, at larger scales, an interesting and intuitive concept of the origin of physical inertia is illustrated, without recurring to the Higgs field, by results of Quantum Chromodynamics (QCD) for protons and neutrons, which are not elementary and are composed of quarks and gluon fields. If an initial, unperturbed nuclear configuration is disturbed, the gluon field generates forces that tend to restore this unperturbed configuration. It is the distortion of the nuclear field that gives rise to the force in opposition to the one producing it, in analogy to the electromagnetic case. But in contrast to the electromagnetic model of an electron, the QCD model of a nucleon is stable, and other compensation fields are not needed. Now, the gluon field mass can be computed from the total energy (or momentum) stored in the field, and it turns out that the QCD version in which quark masses are taken as zero provides a remarkably good approximation to reality. Since this version of QCD is a theory whose basic building blocks have zero mass, the most of the mass of ordinary matter (more than 90 percent) arises from pure field energy. In other words, the mass of a nucleon can be explained almost entirely from a microscopic viewpoint, which automatically provides a microscopic explanation of Newton’s second law of motion. In order to predict, on dynamical grounds, the inertial mass of a relativistically moving nucleon one does not need to have access to the detailed dynamics of strong interactions. It is enough to assume Lorentz covariance (i.e. Lorentz form-invariance of field equations) of the complete QCD dynamics involved in nucleon mass calculations.
The previous discussion, results in a most general statement: it is enough to assume Lorentz covariance of the quantum field theory involved in micro-particle (elementary or not elementary) mass calculations in order to obtain the same result for the relativistic mass correction from the two synchronization conventions discussed here, and it does not matter which transformation (Galilean or Lorentz) is used.
13.8.3 What does Space-Time Geometry Explain?
It is important to stress at this point that the dynamical line of argument discussed here explains what the Minkowski geometry physically means. The pseudo-Euclidean geometric structure of space-time is only an interpretation of the behavior of the dynamical matter fields in the view of different observers, which is an observable, empirical fact. It should be clear that the relativistic properties of the dynamical matter fields are fundamental, while the geometric structure is not. Dynamics, based on the field equations, is actually hidden in the language of kinematics. The Lorentz covariance of the equations that govern the fundamental interactions of nature is an empirical fact, while the postulation of the pseudo-Euclidean geometry of space-time is a mathematical interpretation of it that yields the laws of relativistic kinematics: at a fundamental level this postulate is, however, based on the way fields behave dynamically.
13.9 Bibliography and Notes
1. Let us try to get a better understanding of the geometric restatement of Newton’s second law. To derive the covariant form of relativistic dynamics, we should embed the three-dimensional vector relation $md\vec{v}/dt=\vec{f}$ into the four-dimensional geometry of Minkowski space [52]. The idea of embedding is based on the principle of relativity i.e. on the fact that the usual Newton’s second law can always be used in any Lorentz frame where the particle, whose motion we want to describe, is at rest. In other words, if an instantaneously comoving Lorentz frame is given at some instant, one can precisely predict the evolution of the particle in this frame during an infinitesimal time interval. In geometric language, the Newton law is strict on a hyperplane perpendicular to the world line. However, the hyperplane tilts together with its normal $u_{\mu}$ as one moves along the world line. For the embedding we need an operator $\hat{P}_{\perp}$ that continually projects vectors of Minkowski space on hyperplanes perpendicular to world line. The desired operator is $(\hat{P}_{\perp})_{\mu\nu}=\eta_{\mu\nu}-u_{\mu}u_{\nu}/u^{2}$ [52]. In the instantaneously comoving frame one can unambiguously construct a four-force $K_{\mu}=[0,\vec{f}]$. Then, in an arbitrary Lorentz frame, the components $K_{\mu}$ can be found through the appropriate Lorentz transformation. In the rest frame obviously $u_{\mu}K^{\mu}=0$. It follows that, since $u_{\mu}K^{\mu}$ is an invariant, the four-force $K_{\mu}$ is perpendicular to the four-velocity $u_{\mu}$ in any Lorentz frame. The desired embedding of Newton’s second law in hyperplanes perpendicular to the world line is found by imposing $(\hat{P}_{\perp})_{\mu\nu}(mdu_{\nu}/d\tau-K_{\nu})=0$. This is a tensor equation in Minkowski space-time that relates geometric objects and does not need coordinates to be expressed. The evolution of a particle can be described in terms of world line $\sigma(\tau)$, and the 4-velocity by $u=d\sigma/d\tau$, having a meaning independently of any coordinate system. Similarly in geometric language, the electromagnetic field is described by the second-rank, antisymmetric tensor $F$, which also requires no coordinates for its definition. This tensor produces a 4-force on any charged particle given by $\hat{P}_{\perp}\cdot(mdu/d\tau-eF\cdot u)=0$ [52]. This is the basic dynamics law for relativistic charged particles expressed in terms of geometric objects and automatically included the principle of relativity. The presence of the projector operator $\hat{P}_{\perp}$ suggests that we have only three independent equations. In the case of Maxwell’s equation we are able to rewrite the equations in the relativistic form without any change in the meaning at all, just with a change notations. It is important to noticed that the situation with dynamics equations is more complicated. In order fully to understand the meaning of the embedding of the dynamics law in the hyperplanes perpendicular to the world line, one must keep in mind that, above, we characterized the Newton’s equation in the Lorentz comoving frame as a phenomenological law. The system of coordinates in which the equations of Newton’s mechanics are valid can be defined as Lorentz rest frame. The relativistic generalization of the Newton’s second law to any Lorentz frame permits us to make correct predictions. The projector operator guarantees that this coordinate system restriction will be satisfied.
2. In general, the covariant equation of motion can be solved only by numerical methods; however, it is always attractive to find instances where exact solutions can be obtained. The simplest case of great practical importance is that the motion of a particle in a constant homogeneous magnetic field. From the solution of covariant equation of motion authors of textbooks [5, 52, 53] conclude that the covariant motion of a charge particle in a constant magnetic field is a uniform circular motion. The trajectory of the particle does not include relativistic kinematics effects and the Galilean vectorial law of addition of velocities is actually used. Actually, the old kinematics comes from the relation $d\tau=dt/\gamma$. This relation between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate in the lab frame, is based on the use of the absolute time convention. It is only after the authors of textbooks have make those replacement for $d\tau$ that authors obtain the usual formula for non-covariant trajectory.
3. The non-commutativity and non-associativity of the relativistic composition of non-collinear velocities was first recognized by Mocanu [54] and latter resolved by Ungar [55], by showing role played by the Wigner rotation in the transformation.
14 Relativity and Electrodynamics
The differential form of Maxwell’s equations describing electromagnetic phenomena in the Lorentz lab frame is given by Eq.(8). Now let us use these equations to discuss the phenomena called radiation. To evaluate radiation fields arising from an external sources in Eq. (8), we need to know the velocity $\vec{v}$ and the position $\vec{x}$ as a function of the lab frame time $t$. As discussed above, it is generally accepted that one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particles moving along non-covariant trajectory like $\vec{x}(t)$. The trajectory $\vec{x}(t)$, which follow from the solution of the corrected Newton’s second law Eq. (4) under the absolute time convention, does not include, however, relativistic kinematics effects.
We argue that this algorithm for solving usual Maxwell’s equations in the lab frame, which is considered in all standard treatments as relativistically correct, is at odds with the principle of relativity. However, the usual Maxwell’s equations in the lab frame, Eq. (8), are compatible only with covariant trajectories calculated by using Lorentz coordinates, therefore including such relativistic features as the relativity of simultaneity and the Wigner rotation.
We now consider the particle motion in a giving magnetic field. In their usual form, Maxwell’s equations are valid only in Lorentz reference frames. According to correct coupling of fields and particles, from the Eq. 10 follows that
$\rho(\vec{x},t)=e\delta(\vec{x}-\vec{x}_{cov}(t))~{},$ (26)
$\vec{j}(\vec{x},t)=e\vec{v}_{cov}(t)\delta(\vec{x}-\vec{x}_{cov}(t))~{},$ (27)
where $\vec{v}_{cov}=d\vec{x}_{cov}/dt$. The covariant trajectory of a particle is viewed from the lab frame as the result of successive Lorentz transformations.
We find in this book that the Wigner rotation turns out to play an important role as the regulator of the velocity addition law. As we have already pointed, the four-velocity cannot be decomposed into $u=(c\gamma,\vec{v}\gamma)$ when we deal with a particle accelerating along a curved trajectory in the Lorentz lab frame (1). The presentation of the time component simply as the relation $d\tau=dt^{\prime}/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
One of the consequences of non-commutativity of non-collinear Lorentz boosts is a difference between covariant and non covariant particle trajectories in a giving magnetic field. One can see that this essential point has never received attention by the physical community. As a result, a correction of the conventional radiation theory is required.
14.1 Why did the Error in Radiation Theory Remain so Long Undetected?
The difference between covariant and non-covariant particle trajectories was never understood. So, physicists did not appreciate that there was a contribution to the radiation from relativistic kinematics effects. At this point, a reasonable question arises: why the error in radiation theory should have so long remained undetected?
In order to answer this question, we shall discuss the subject more mathematically. For an arbitrary parameter $v/c$ covariant calculations of the radiation process is very difficult. There are, however, circumstances in which calculations can be greatly simplified. As example of such circumstance is a non-relativistic radiation setup. The non-relativistic asymptote provides the essential simplicity of the covariant calculation. The reason is that the non-relativistic assumption implies the dipole approximation which is of great practical significance. In accounting only for the dipole part of the radiation we neglect all information about the electron trajectory That means that the dipole radiation does not show any sensitivity to the difference between covariant and non-covariant particle trajectories.
We want now to solve electrodynamics equations mathematically in a general way and consider the radiation associated with the succeeding terms in (multi-pole) expansion of the field in powers of the ratio $v/c$. Radiation theory is naturally developed in the space-frequency domain, as one is usually interested in radiation properties at a given position in space and at a certain frequency. In this book we define the relation between temporal and frequency domain via the following definition of Fourier transform pair:
[…]
Suppose we are interested in the radiation generated by an electron and observed far away from it. In this case it is possible to find a relatively simple expression for the electric field (see Appendix I). We indicate the electron velocity in units of $c$ with $\vec{\beta}$, the electron trajectory in three dimensions with $\vec{r}(t)$ and the observation position with $\vec{r}_{0}$. Finally, we introduce the unit vector
$\vec{n}=\frac{\vec{r}_{0}-\vec{r}(t)}{|\vec{r}_{0}-\vec{r}(t)|}$ (29)
pointing from the retarded position of the electron to the observer. In the far zone, by definition, the unit vector $\vec{n}$ is nearly constant in time. If the position of the observer is far away enough from the charge, one can make the expansion
$\left|\vec{r}_{0}-\vec{r}(t)\right|=r_{0}-\vec{n}\cdot\vec{r}(t)~{}.$ (30)
We then obtain the following approximate expression for the the radiation field in the space-frequency domain:
[…]
where $\omega$ is the frequency, $(-e)$ is the negative electron charge and we make use of Gaussian units. A different constant of proportionality in Eq.(LABEL:revwied) and in the well known textbooks is to be ascribed to the use of different units and definition of the Fourier transform.
First we will limit our consideration to the case of sources moving in a non-relativistic fashion. According to the principle of relativity, usual Maxwell’s equations can always be used in any Lorentz frame where sources are at rest. The same considerations apply where sources are moving in non-relativistic manner. In particular, when oscillating, charge particles emit radiation, and in the non-relativistic case, when the velocities of oscillating charges $v\ll c$, dipole radiation will be generated and described with the help of the Maxwell’s equations in their usual form, Eq. (8).
Let’s examine in a more detail how the dipole radiation term comes about. The time $\vec{r}(t)\cdot(\vec{n}/c)$ in the integrands of the expression for the radiation field amplitude, Eq. (LABEL:revwied), can be neglected in the cases where the trajectory of the charge changes little during this time. It is easy to find the conditions for satisfying this requirement. Let us denote by $a$ the order of magnitude of the dimensions of the system. Then the time $\vec{r}(t)\cdot(\vec{n}/c)\sim a/c$. In order to ensure that the distribution of the charges in the system does not undergo a significant change during this time, it is necessary that $a\ll\lambda$. This condition can be written in still another form $v\ll c$, where $v$ is of the order of magnitude of the velocities of the charges.
We consider the radiation associated with the first order term in the expansion of the Eq. (LABEL:revwied) in power of $\vec{r}(t)\cdot(\vec{n}/c)$. In doing so, we neglected all information about the electron trajectory $\vec{r}(t)$. In this dipole approximation the electron orbit scale is always much smaller than the radiation wavelength and Eq. (LABEL:revwied) gives fields very much like the instantaneous theory. So we are satisfied using the non-covariant approach when considering the dipole radiation theory.
But that is only the first and most practically important term. The other terms tell us that there are higher order corrections to the dipole radiation approximation. The calculation of this correction requires detailed information about the electron trajectory. Obviously, in order to calculate the correction to the dipole radiation, we will have to use the covariant trajectory and not be satisfied with the non-covariant approach. However, correction to the multipole radiation is expected to be small. For example, covariant (trajectory) correction for quadrupole radiation is treated as a correction of higher (than quadrupole) order.
14.2 An Illustrative Example
In the next chapter we present a critical reexamination of existing synchrotron radiation theory. But before the discussion of this topic it would be well to illustrate error in standard coupling fields and particles in accelerator and plasma physics by considering the relatively simple example, wherein the essential physical features are not obscured by unnecessary mathematical difficulties. This illustrative example is mainly addressed to readers with limiting knowledge of accelerator and synchrotron radiation physics. Fortunately, the error in standard coupling fields and particles can be explained in a very simple way.
14.2.1 The Velocity of an Electron Accelerated by the Kicker Field
An electron kicker setup is a practical case of study for illustrating the difference between covariant and non-covariant trajectories. Let us consider the simple case when an ultrarelativistic electron moving with the velocity $v$ along $z$-axis in the lab frame is kicked by a weak dipole magnetic field directed along $x$-axis. We assume for simplicity that the kick angle is small compare with $1/\gamma$, where $\gamma=1/\sqrt{1-v^{2}/c^{2}}$ is the relativistic factor. This means that we take the limit $\gamma\gg 1\gg\gamma v_{y}/v$. Let us start with non-covariant particle tracking calculations. The trajectory of the electron, which follows from the solution of the corrected Newton’s second law under the absolute time convention, does not include relativistic effects. Therefore, as usual for Newtonian kinematics, Galilean vectorial law of addition of velocities is actually used. Non-covariant particle dynamics shows that the electron direction changes after the kick, while the speed remains unvaried (Fig. 60). According to non-covariant particle tracking, the magnetic field $B\vec{e}_{x}$ is only capable of altering the direction of motion, but not the speed of the electron. After the kick, the beam velocity components are $(0,v_{y},v_{z})$, where $v_{z}=\sqrt{v^{2}-v_{y}^{2}}$. Taking the ultrarelativistic limit $v\simeq c$ and using the second order approximation we get $v_{z}=v[1-v_{y}^{2}/(2v^{2})]=v[1-v_{y}^{2}/(2c^{2})]$.
Figure 60: A setup for illustration the difference between covariant and non-covariant trajectories. The motion of a relativistic electron accelerated by a kicker field. It is assumed for simplicity that $\gamma\gg 1\gg\gamma v_{y}/v$. According to the non-covariant particle tracking, the magnetic field $B\vec{e}_{x}$ is only capable altering the direction of motion, but not the speed of the electron.
In contrast, covariant particle tracking, which is based on the use of Lorentz coordinates, yields different results for the velocity of the electron. Let us consider a composition of passive Lorentz transformations that track the motion of the relativistic electron accelerated by the kicker field. Let the $S$ be the lab frame of reference and $S^{\prime}$ a comoving frame with velocity $\vec{v}$ relative to $S$. Upstream of the kicker, the particle is at rest in the frame $S^{\prime}$. In order to have this, we impose that $S^{\prime}$ is connected to $S$ by the Lorentz boost $L(\vec{v})$, with $\vec{v}$ parallel to the $z$ axis, which transforms a given four vector event $X$ in a space-time into $X^{\prime}=L(\vec{v})X$. Let us analyze the particle evolution within $S^{\prime}$ frame. Our particle is at rest and the kicker is running towards it with velocity $-\vec{v}$. The moving magnetic field of the kicker produces an electric field orthogonal to it. When the kicker interacts with the particle in $S^{\prime}$ we thus deal with an electron moving in the combination of perpendicular electric and magnetic fields.
We consider the small expansion parameter $\gamma v_{y}/c\ll 1$, neglecting terms of order $(\gamma v_{y}/c)^{3}$, but not of order $(\gamma v_{x}/c)^{2}$. In other words, we use the second-order kick angle approximation. It is easy to see that the acceleration in the crossed fields yields a particle velocity $v^{\prime}_{y}=\gamma v_{y}$ parallel to the $y$-axis and $v^{\prime}_{z}=-v(\gamma v_{y}/c)^{2}/2$ parallel to the $z$-axis. If we neglect terms in $(\gamma v_{y}/c)^{3}$, the relativistic correction in the composition of velocities does not appear in this approximation.
First, we analyze textbook presentation of a composition of motions. Let $S"$ be a frame fixed with respect to the particle downstream the kicker. As is known, the non collinear Lorentz boosts does not commute. In our second order approximation we can neglect the difference between the $\gamma v_{y}/c$ and $\gamma_{z}v_{y}/c$, where $\gamma_{z}=1/\sqrt{1-v_{z}^{2}/c^{2}}$. Here $v_{z}=v(1-\theta_{k}^{2}/2)$ and $\theta_{k}=v_{y}/v=v_{y}/c$ in our (ultrarelativistic) case of interest. Therefore we can use a sequence of two commuting non-collinear Lorentz boosts linking $X^{\prime}$ in $S^{\prime}$ to $X^{\prime\prime}$ in $S^{\prime\prime}$ as $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e^{\prime}}_{z}v^{\prime}_{z})X^{\prime}=L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e^{\prime}}_{y}v^{\prime}_{y})X^{\prime}$ in order to discuss the beam motion in the frame $S^{\prime}$ after the kick. Here $\vec{e^{\prime}}_{y}$ and $\vec{e^{\prime}}_{z}$ are unit vectors directed, respectively, along the $x^{\prime}$ and $z^{\prime}$ axis. Note that as observed by an observer on $S^{\prime}$, the axes of the frame $S^{\prime\prime}$ are parallel to those of $S^{\prime}$. The relation $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e}_{z}v)X$ presents a step-by-step change from $S$ to $S^{\prime}$ and then to $S"$. For the simple case of parallel velocities, the addition law is $L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e}_{z}v)=L(\vec{e}_{z}v_{z})$. The resulting (passive) Lorentz boost composition can be represented as $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e}_{z}v_{z})X$ and textbooks state that the magnetic field, in the Lorentz coordinatization, is only capable altering the direction of motion, but not the speed of the electron.
According to textbooks, in relativistic kinematics, the composition of the perpendicular velocities is studied analytically from the composition of Lorentz boosts. We have already discussed the law of composition of velocities in Section 13.5. It was shown that velocities addition in the Lorentz coordinatization is regulated by the Wigner rotation. In the ultrarelativistic approximation we find the simple result $v_{tot}=v_{z}=v[1-v_{y}^{2}/(2c^{2})]$, so that a Lorentz transformation leads to a rotation of the particle velocity $v_{z}$ of the angle $v_{y}/c$ (Fig. 61). It becomes clear that the contradiction is related to the Wigner rotation. In fact, as we already discussed in the previous chapter, the relation $dt^{\prime}=dt/\gamma$ cannot be used when we deal with relativistic composition of perpendicular velocities.
It should be note, however, that there is another satisfactory way of the covariant electron tracking. Let us now return to our consideration of the motion of a relativistic electron accelerated by the kicker field and let us discuss the observation of an inertial lab observer without passing from one reference frame to another. We now consider the case when the electron in the lab frame is accelerated by the kicker field. In other words, we consider an active boost. In the case of an active boost we are dealing with motion of the same physical system, evolving in time and treated from the point of view of the same reference system. The simplest method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the electron was upstream the kicker. This synchronization convention preserves simultaneity and is actually based on the absolute time (or absolute simultaneity) convention. The usual Galileo rule of addition of velocities is used to determine the Galileo boosts tracking an electron, instant after instant, along its motion in a constant magnetic field.
In the Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization. The overall combination of Galileo transformation and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the lab frame. Let us analyze the resynchronization process of the lab distant clocks during the acceleration of the electron. This will allow us to demonstrate a direct relation between the decrease of the electron speed after the kick in Lorentz coordinates and the time dilation phenomenon. As we already remarked, the Lorentz coordinate system is only a mental construction: manipulations with non existing clocks are only needed for the application of the usual Maxwell’s equations for synchrotron radiation calculations.
Suppose that upstream the kicker we pick a Lorentz coordinates in the lab frame. Then, an instant after entering the magnetic field, the electron velocity changes of the infinitesimal value $d\vec{v}$ along the $y$-axis. At this first step, Eq.(21) allows us to express the differential $d\vec{v}$ through the differential $dt$ in the Lorentz coordinate system assigned upstream the kicker. If clock synchronization is fixed, this is equivalent to the application of the absolute time convention. In order to keep Lorentz coordinates in the lab frame, as discussed before, we need to perform a clock resynchronization by introducing an infinitesimal time shift. The simplest case is when the kick angle $\theta_{k}$ is very small, and we evaluate transformations, working only up to the order $(\theta_{k}\gamma)^{2}$. The restriction to this order provides an essential simplicity of calculations in our case of interest for two reasons. First, relativistic correction to compositions of non-collinear velocity increments does not appear in this expansion order, but only in the order $(\gamma\theta_{k})^{3}$. Second, the time dilation appears in the highest order we use. Thus, Eq.(21) allows us to express the small velocity change $\Delta\vec{v}$ after the kick in the initial Lorentz coordinates system, and to perform clock resynchronization only downstream the kicker. Therefore, after the kick we can consider the composition of two Lorentz boosts along the perpendicular $x$ and $z$ directions. The first boost imparts the velocity $v\theta_{k}\vec{e}_{y}$ to the electron along the $y$-axis and the second boost imparts the additional velocity $-(v\theta_{k}^{2}/2)\vec{e}_{z}$ along the $z$ axis, while the restriction to second order assures that the boosts commute.
Figure 61: The motion of a relativistic electron accelerated by the kicker field. In order to describe the kicker influence on the electron trajectory as Lorentz transformation, one needs to perform a clock resynchronization by introducing a time shift and change the scale of time, that is the rhythm of all clocks, from $t$ to $\gamma_{y}t$, with $\gamma_{y}=1+\theta_{k}^{2}/2$. It follows that the total electron speed in the Lorentz lab frame downstream the kicker decreases from $v$ to $v(1-\theta^{2}_{k}/2)$.
In order to keep a Lorentz coordinates system in the lab frame after the kick, that is equivalent to describe the kicker influence on the electron trajectory as Lorentz transformation, we need to perform a clock resynchronization by introducing a time shift and change the scale of time, that is the rhythm of all clocks, from $t$ to $\gamma_{y}t$, with $\gamma_{y}\simeq 1+\theta_{k}^{2}/2$. It is immediately understood that the speed of electron downstream the kicker is no longer independent of the electron motion in the magnetic field (Fig. 61). No relativistic correction to the velocity component along the $y$-axis appears in the second order, but a correction of the longitudinal velocity component, changing $v_{z}$ to $v_{z}/\gamma_{y}$ with $v_{z}=v(1-\theta_{k}^{2}/2)$ and $v_{z}/\gamma_{y}=v(1-\theta_{k}^{2})$. It follows that the total electron speed in the lab frame, after clock resynchronization downstream the kicker, decreases from $v$ to $v(1-\theta_{k}^{2}/2)$. The time dilation does not come into the calculation of the velocity increment, but appears in the correction of the initial (relativistic) velocity $\vec{v}=v\vec{e}_{z}$. We have found that we get the same result whether we using passive Lorentz transformations or using active Lorentz transformations as it must be.
Note that we discuss particle tracking in the limit of a small kick angle $\gamma v_{y}/c\ll 1$. However, even in this simple case and for a single electron we are able to demonstrate the difference between non-covariant and covariant particle trajectories. The electron speed decreases from $v$ to $v(1-\theta_{k}^{2}/2)$. This result is at odds with the prediction from non-covariant particle tracking.
14.2.2 Discussion
We now want to point out that when we accelerate the electron in the lab frame by a kicker, the information about this acceleration is included into the covariant trajectory. The result of covariant approach clearly depends on the absolute value of the kick angle. The argument concerning the relativity motion in our case of interest cannot be applied. Not all is relative in special relativity, because any change of velocity has an absolute meaning. The ”absolute” acceleration means acceleration with respect to the fixed stars.
There are several points to be made about the above results. One interesting question is, where physically is the information about the electron acceleration? We answer this question in great detail in the Section 15.6. It is a dynamical line of arguments that explains this paradoxical situation with the trajectory of the accelerated electron. Dynamics, based on the field equations, is actually hidden in the language of relativistic kinematics. Without proof, we may state the results. The distribution of the electromagnetic fields from a fast moving charge is described by Ginzburg-Frank formula (see Appendix IV). We remark first that for a rapidly moving electron we have nearly equal transverse and mutually perpendicular electric and magnetic fields. These are indistinguishable from the fields of a beam of radiation. From microscopic viewpoint, when the electron passes through a kicker, its fields are perturbed, and now include information about acceleration.
Let us quickly look at some other cases of interest. Above we consider the simple case when kick angle is small. When the electron passes through the weak kicker there is no synchrotron radiation (to be more precise, in this case radiation is indistinguishable from the self-electromagnetic fields of the electron). We expect that an electron that passes through the weak kicker is still ”field-dressed”. Suppose the ultrarelativistic electron in the lab frame is kicked by dipole magnet field, but the kick angle is large compare with $1/\gamma$. In other words, we now consider an electron moving along a standard synchrotron radiation setup. At the exit of strong kicker we have a ”naked” (or ”field-free”) electron i.e. an electron that is not accompanied by virtual radiation fields. There is a process of formation of the ”field dressed” electron (i.e. formation of the fields from a fast moving charge) within the formation length from the very beginning of the straight section downstream of the strong kicker. The information about acceleration is included now into the synchrotron radiation. But the acceleration history (together with the self fields of the ultrarelativistic electron) is washed out during the kicking process, and now electron fields not include information about acceleration. That is to say, the information about acceleration is not included into the covariant trajectory downstream the strong kicker (we will discuss this subject further in the Section 14.5).
14.2.3 The Motion of an Electron Accelerated by a Kicker in a Bending Magnet
In our relativistic but non-covariant study of electron motion in a given magnetic field, the electron has the same velocity and consequently the same relativistic factor $\gamma$ upstream and downstream of the kicker. Suppose we now put the electron through a bending magnet (i.e. a uniform magnetic field directed along the $y$-axis ), Fig. 62. The motion in the bending magnet we obtained is practically the same as in the case of non-relativistic dynamics, the only difference being the appearance of the relativistic factor $\gamma$ in the determination of cyclotron frequency $\omega_{c}=eB/(m\gamma)$. The curvature radius $R$ of the trajectory is derived from the relation $v_{\perp}/R=\omega_{c}$, where $v_{\perp}=v(1-\theta_{k}^{2}/2)$ is the component of the velocity normal to the field of the bending magnet $\vec{B}=B\vec{e}_{y}$. According to non-covariant particle tracking, after the kick, the correction to the radius $R$ is only of order $\theta_{k}^{2}$.
Figure 62: Geometry for radiation production from a bending magnet. The motion of a relativistic electron accelerated by a kicker field. According to the non-covariant particle tracking , the magnetic field $B\vec{e}_{x}$ of the kicker is only capable altering the direction of motion, but not the speed of the electron.
One could naively expect that according to covariant particle tracking, since the total speed of electron in the lab frame downstream of the kicker decreases from $v$ to $v(1-\theta_{k}^{2}/2)$, this would also lead to a consequent decrease of the three-momentum $|\vec{p}|$ from $m\gamma v$ to $m\gamma v(1-\gamma^{2}\theta_{k}^{2}/2)$ in our approximation. However, such a momentum change would mean a correction to the radius $R$ of order $\gamma^{2}\theta_{k}^{2}$ so that there is a glaring conflict with the calculation of the radius according to non covariant tracking. Since the curvature radius of the trajectory in the bending magnet has obviously an objective meaning, i.e. it is convention-invariant, this situation seems paradoxical. The paradox is solved taking into account the fact that in Lorentz coordinates the three-vector of momentum $\vec{p}$ is transformed, under Lorentz boosts, as the space part of the four vector $p_{\mu}$. Let us consider a composition of Lorentz boosts that track the motion of the relativistic electron accelerated by the kicker field. Under this composition of boosts the longitudinal momentum component remains unchanged (with accuracy $\theta_{k}^{2}$).
Let us verify that this assertion is correct. We have $p_{\mu}=[\mathcal{E}/c,\vec{p}]$. We consider the Lorentz frame $S^{\prime}$ fixed with respect to the electron upstream the kicker, and in the special case when electron is at rest $p_{\mu}^{\prime}=[mc,\vec{0}]$. We turn focus on what happens in $S^{\prime}$. Acceleration in the crossed kicker fields gives rise to an electron velocity $v_{y}^{\prime}=\gamma v_{y}$ parallel to the $y$-axis and $v_{z}^{\prime}=-v(\gamma v_{y})^{2}/2$ parallel to the $z$-axis. Downstream of the kicker the transformed four-momentum is $p_{\mu}^{\prime}=[mc+mv_{y}^{\prime 2}/(2c),0,mv_{y}^{\prime},mv_{z}^{\prime}]$, where we evaluate the transformation only up to the order $(\gamma v_{y}/c)^{2}$, as done above. We note that, due to the transverse boost, there is a contribution to the time-like part of the four-momentum vector i.e. to the energy of the electron. In fact, the energy increases from $mc^{2}$ to $mc^{2}+m(\gamma v_{y})^{2}/2$. We remind that $S^{\prime}$ is connected to the lab frame $S$ by a Lorentz boost. Now, with a boost to a frame moving at velocity $\vec{v}=-v\vec{e}_{z}$, the transformation of the longitudinal momentum component, normal to the magnetic field of the bend, is $p_{z}=\gamma(p_{z}^{\prime}+vp_{0}^{\prime}/c)=\gamma mv$. Therefore we can see that the momentum component along the $z$-axis remains unchanged in our approximation of the Lorentz transformation. We also have, from the transformation properties of four-vectors, that the time component $p_{0}=\gamma(p^{\prime}_{0}+vp^{\prime}_{z})=\gamma mc$
Let us now return to our consideration on the covariant electron trajectory calculation in the Lorentz lab frame when a constant magnetic field is applied. We analyzed a very simple (but very practical) kicker setup and we noticed that, in fact, the three-momentum is not changed; so we have already verified that this transformation is the same as the non covariant transformation for the three-momentum, i.e. $\vec{p}_{cov}=\vec{p}$. We also found that there is a difference between covariant and non covariant output velocities, $v_{cov}<v$. In these transformations we therefore demonstrated that $\vec{p}_{cov}\neq m\vec{v}_{cov}/\sqrt{1-v^{2}_{cov}/c^{2}}$ for curved trajectory in ultrarelativistic asymptotic. It is interesting to discuss what it means that there are two different (covariant and non covariant) approaches that produce the same particle three-momentum. The point is that both approaches describe correctly the same physical reality and the curvature radius of the trajectory in the magnetic field (and consequently the three-momentum) has obviously an objective meaning, i.e. is convention-invariant. In contrast to this, the velocity of the particle has objective meaning only up to a certain accuracy, because the finiteness of velocity of light takes place.
14.2.4 Red Shift of the Critical Frequency of the Synchrotron Radiation
Next we discuss the interesting problem of emission of synchrotron radiation in a bending magnet with and without kick. Let us consider the setup pictured in Fig. 62. Suppose that an ultrarelativistic electron moving along the $z$-axis in the lab frame is kicked by a weak dipole field directed along the $x$-axis before entering a uniform magnetic field directed along the $y$-axis, i.e. a bending magnet. An accelerated electron traveling on a curved trajectory emits radiation. When moving at relativistic speed, this radiation is emitted as a narrow cone tangent to the path of the electron. Moreover, the radiation amplitude becomes very large in this direction. This phenomenon is known as Doppler boosting. Synchrotron radiation is generated when a relativistic electron is accelerated in a bending magnet. Without going into the details of computation, it is possible to present intuitive arguments explaining why the characteristics of the spectrum of synchrotron radiation only depend, in the ultrarelativistic limit, on the difference between electron and light speed.
We turn now to the radiation emitted by an ultrarelativistic electron in a bending magnet. Let us discuss the case when the source is heading towards the observer (Fig. 63). An electromagnetic source propagates through the system as a function of time following a certain trajectory $\vec{x}(t^{\prime})$. However, an electromagnetic signal emitted at time $t^{\prime}$ at a given position $\vec{x}(t^{\prime})$ arrives at the observer position at a different time $t$, due to the finite speed of light. As a result, an observer sees the motion of the electromagnetic source as a function of $t$. The prime used here to indicate the retarded times should not be confused with the primes referring to a Lorentz transformed frame in the proceeding sections.
[…]
Figure 63: Geometry for synchrotron radiation from a bending magnet. Radiation from an electron passing through the setup is observed through a spectral filter by a fixed observer positioned on the tangent to the bend at point $P$, as shown in Fig. (a). Electromagnetic source propagates through the system, as a function of time, as shown in Fig. (b). However, electromagnetic signal emitted at time $t^{\prime}$ at a given position $x(t^{\prime})$ arrives at observer position at a different time $t$, due to finite speed of light. As a result, the observer in Fig. (a) sees the electromagnetic source motion as a function of $t$. The apparent motion is a hypocycloid, and not the real motion $x(t^{\prime})$. The observer sees a time-compressed motion of the sources, which go from point $A$ to point $B$ in an apparent time corresponding to an apparent distance $2R\theta dt/dt^{\prime}$.
Figure 64: Formation length for bend. Formation length $L_{f}$ can also be considered as a longitudinal size of a single-electron source. It does not make sense at all to talk about the position where electromagnetic signals are emitted within $L_{f}$.
It is clear from the above that, according to conventional synchrotron radiation theory, if we consider radiation, the introduction of the kick only amounts to a rigid rotation of the angular distribution along the new direction of the electron motion. This is plausible, if one keeps in mind that after the kick the electron has the same velocity and emits radiation in the kicked direction owing to the Doppler effect.
According to the correct coupling of fields and particles, there is a remarkable prediction of synchrotron radiation theory concerning the setup described above. Namely, there is a red shift of the critical frequency of the synchrotron radiation in the kicked direction. To show this, let us consider the covariant treatment, which makes explicit use of Lorentz transformations. When the kick is introduced, covariant particle tracking predicts a non-zero red shift of the critical frequency, which arises because in Lorentz coordinates the electron velocity decreases from $v$ to $v-v\theta_{k}^{2}/2$, while the velocity of light is unvaried and equal to the electrodynamics constant $c$. The red shift in the critical frequency can be expressed by the formula $\Delta\omega_{cr}/\omega_{cr}\simeq-\gamma^{2}v_{y}^{2}/c^{2}=-\gamma^{2}\theta_{k}^{2}$. We now see a second order correction $\theta_{k}^{2}$ that is, however, multiplied by a large factor $\gamma^{2}$.
It should be note, however, that there is another satisfactory way of explaining the red shift. We can reinterpret this result with the help of a non-covariant treatment, which deals with non- covariant particle trajectories, and with Galilean transformations of the electromagnetic field equations. According to non-covariant particle tracking the electron velocity is unvaried. However, Maxwell’s equations do not remain invariant with respect to Galilean transformation, and the velocity of light has increased from $c$, without kick, to $c(1+\theta_{k}^{2}/2)$ with kick. The reason for the velocity of light being different from the electrodynamics constant $c$ is due to the fact that, according to the absolute time convention, the clocks after the kick are not resynchronized. The ratio of the electron velocity to that of light is convention independent i.e. it does not depend on the distant clock synchronization or on the rhythm of the clocks. Our calculations show that covariant and non-covariant treatments (at the correct coupling fields and particles) give the same result for the red shift prediction, which is obviously convention-invariant and depends only on the (dimensionless) parameter $v_{coord}/c_{coord}=v(1-\theta^{2}_{k}/2)/c=v/[c(1+\theta^{2}_{k}/2)]$, where $v_{coord}$ is the coordinate electron velocity, and $c_{coord}$ is the coordinate velocity of light.
In order to confirm the predictions of our synchrotron radiation theory, we propose an experimental test at third generation synchrotron radiation sources. Synchrotron radiation from bending magnets is emitted within a wide range of frequencies. The possibility of using narrow bandwidth sources in an experimental study on the red shift in synchrotron radiation spectrum looks more attractive. This allows one to increase the sensitivity of the output intensity on the red shift, and to relax the requirement on beam kicker strength and photon beam line aperture. Undulators, as sources of quasi-monochromatic synchrotron radiation, produce light in a sufficiently narrow bandwidth for our purposes. They cause the electron beam to follow a periodic undulating trajectory with the consequence that interference effects occur. Undulators have typically many periods. The interference of radiation produced in different periods results in a bandwidth that scales as the inverse number of periods. Therefore, the use of insertion devices installed at third generation synchrotron radiation facilities would allow us to realize a straightforward increase in the sensitivity to the red shift at a relatively small kick angle, $\theta_{k}<1/\gamma$. In the next chapter we discuss this experimental test in more detail.
14.3 Bibliography and Notes
[…]
15 Synchrotron Radiation
15.1 Introductory Remarks
Accelerator physics was always thought in terms of the old (Newtonian) kinematics that is not compatible with Maxwell’s equations. We would like now to use our ideas about dynamics and electrodynamics to consider in some greater detail the question: ”Why did the error in the synchrotron radiation theory remain so long undetected?”
The phenomena of electromagnetic radiation which we want to study are relatively complicated. For an arbitrary setup covariant calculations of the radiation process is very difficult. There are, however, circumstances in which calculations can be greatly simplified. As example of such circumstance is a synchrotron radiation setup. Similar to the non-relativistic asymptote, the ultrarelativistic asymptote also provides the essential simplicity of the covariant calculation. The reason is that the ultrarelativistic assumption implies the paraxial approximation. Since the formation length of the radiation is much longer than the wavelength, the radiation is emitted at small angles of order $1/\gamma$ or even smaller, and we can therefore enforce the small angle approximation. We assume that the transverse velocity is small compared to the velocity of light. In other words, we use a second order relativistic approximation for the transverse motion. Instead of small (total) velocity parameter $(v/c)$ in the non-relativistic case, we use a small transverse velocity parameter $(v_{\perp}/c)$. The next step is to analyze the longitudinal motion, following the same method. We should remark that the analysis of the longitudinal motion in a synchrotron radiation setup is very simple. If we evaluate the transformations up to second order $(v_{\perp}/c)^{2}$, the relativistic correction in the longitudinal motion does not appear in this approximation.
According to covariant approach, the various relativistic kinematics effects concerning to the synchrotron radiation setup, turn up in successive orders of approximation.
In the first order $(v_{\perp}/c)$. - relativity of simultaneity. Wigner rotation, which in the ultrarelativistic approximation appears in the first order already, and results directly from the relativity of simultaneity.
In the second order $(v_{\perp}/c)^{2}$. - time dilation. Relativistic correction in law of composition of velocities, which already appears in the second order, and results directly from the time dilation.
The first order kinematics term $(v_{\perp}/c)$ plays an essential role only in the description of the coherent radiation from a modulated electron beam. In a storage ring the distribution of the longitudinal position of the electrons in a bunch is essentially uncorrelated. In this case, the radiated fields due to different electrons are also uncorrelated and the average power radiated is a simple sum of the radiated power from individual electrons; that is we sum intensities, not fields. In the case of incoherent synchrotron radiation, a motion of the single ultrarelativistic electron in a constant magnetic field, according to the theory of relativity, influences the kinematics terms of the second order $(v_{\perp}/c)^{2}$ only.
15.2 Paraxial Approximation for the Radiation Field
The general method to derive the frequency spectrum is to transform the electric field from the time domain to the frequency domain by the use of Fourier transform. First, let us rewrite Eq. (LABEL:revwied) as follows
[…]
Eq. (LABEL:revwied) and Eq. (LABEL:rrevwied) are equivalent but include different integrands. This is no mistake, as different integrands can lead to the same integral (see Appendix I).
We call $z_{0}$ the observation distance along the optical axis of the system, while $(x_{0},y_{0})$ fixes the transverse position of the observer. Using the complex notation, in this and in the following sections we assume, in agreement with Eq. (28), that the temporal dependence of fields with a certain frequency is of the form:
$\vec{E}\sim\vec{\bar{E}}(z_{0},x_{0},y_{0},\omega)\exp(-i\omega t)~{}.$ (35)
With this choice for the temporal dependence we can describe a plane wave traveling along the positive $z$-axis with
$\vec{E}=\vec{E}_{a}\exp\left(\frac{i\omega}{c}z_{0}-i\omega t\right)~{}.$ (36)
In the following we will always assume that the ultra-relativistic approximation is satisfied, which is the case for SR setups. As a consequence, the paraxial approximation applies too. The paraxial approximation implies a slowly varying envelope of the field with respect to the wavelength. It is therefore convenient to introduce the slowly varying envelope of the transverse field components as
$\vec{\widetilde{E}}(z_{0},x_{0},y_{0},\omega)=\vec{\bar{E}}(z_{0},x_{0},y_{0},\omega)\exp{\left(-i\omega z_{0}/c\right)}~{}.$ (37)
We will now replace all vectors by their components to obtain directional dependency of the synchrotron radiation. The emission angle $\theta=\sqrt{\theta_{x}^{2}+\theta_{y}^{2}}$ is taking with respect to the $z$-axis. Here $\theta_{x}=x_{0}/z_{0}$ is the observation angle projected onto the $x-z$ plane, $\theta_{y}=y_{0}/z_{0}$ is the observation angle projected onto the $y-z$ plane. The components of the unit vector $\vec{n}$ can be approximated by $n_{z}=1-\theta_{x}^{2}/2-\theta_{y}^{2}/2$, $n_{x}=\theta_{x}$, $n_{y}=\theta_{y}$, so $\vec{n}\cdot(\vec{r}_{0}-\vec{r})=(z_{0}-z)(1-\theta_{x}^{2}/2-\theta_{y}^{2}/2)-x\theta_{x}-y\theta_{y}$. We consider the motion in a static magnetic field. According to the conventional particle tracking the magnitude of the velocity is constant and is equal $v=ds/dt=\mathrm{const.}$, where $s(z)$ is the longitudinal coordinate along the path. The transverse components of the envelope of the field in Eq. (LABEL:rrevwied) in the far zone and in paraxial approximation finally becomes
[…]
where the total phase $\Phi_{T}$ is
$\Phi_{T}=\omega\left[{s(z^{\prime})\over{v}}-{z^{\prime}\over{c}}\right]$ (39)
$+\frac{\omega}{2c}\left[z_{0}(\theta_{x}^{2}+\theta_{y}^{2})-2\theta_{x}x(z^{\prime})-2\theta_{y}y(z^{\prime})+z^{\prime}(\theta_{x}^{2}+\theta_{y}^{2})\right]~{}.$ (40)
Here $v_{x}(z^{\prime})$ and $v_{y}(z^{\prime})$ are the horizontal and the vertical components of the transverse velocity of the electron, $x(z^{\prime})$ and $y(z^{\prime})$ specify the transverse position of the electron as a function of the longitudinal position, $\vec{e}_{x}$ and $\vec{e}_{y}$ are unit vectors along the transverse coordinate axis.
15.3 Undulator Radiation
Traditionally, all courses in synchrotron radiation theory have begun in the same way, retracing the path followed in the historical development of the subject. One first learns about synchrotron radiation from bending magnet. We will begin in this chapter by dealing with the ”advanced” subject of the synchrotron radiation theory. For undulator setup covariant calculations of the radiation process is quite simple.
To generate specific synchrotron radiation characteristics, radiation is often produced from special insertion devices called undulators. The resonance approximation, that can always be applied in the case of undulator radiation setups, yields simplifications of the theory. This approximation does not replace the paraxial one, but it is used together with it. It takes advantage of another parameter that is usually large, i.e. number of undulator periods $N_{w}\gg 1$. The frequency emitted by a particle going through an undulator can be obtained by considering the interference between the parts of the radiation created at successive periods (see Fig. 65). The frequency of this field is Doppler shifted. The shortest wavelength is observed on undulator axis. In resonance approximation, all undulator radiation at shortest wavelength is emitted within an angle much smaller than $1/\gamma$ (see Fig. 66). This automatically selects observation angles of interest. In fact, if we consider observation angles outside the diffraction angle, we obtain zero intensity with accuracy $1/N_{w}$.
We now have to ask: Why did the error in insertion device theory remain so long undetected? We answer this question in great detail bellow. In discussing this we will deal only with the radiation into the central cone (see Fig. 66). In doing so, we neglected all information about the electron trajectory. In this approximation the electron oscillation amplitude is always (independently of how large the undulator strength parameter) much smaller than the radiation diffraction size at the undulator exit, because $N_{w}\gg 1$. Thus, the undulator radiation theory gives fields very much like the (dipole-like) instantaneous theory. So, we are satisfied using the conventional (i.e. non-covariant) approach for describing the undulator radiation into the central cone, that is the practical situation of interest.
However, there is one situation where the conventional theory fails. The covariant approach predicts a non-zero red shift of the resonance frequency, which arises when there are perturbations (kicks with respect to the longitudinal axis) of the electron motion. Experimental results confirm our correction for spontaneous undulator emission.
15.3.1 Conventional Theory
Eq. (38) can be used to characterize the far field from an electron moving on any path. In this section we present a simple derivation of the frequency representation of the radiated field produced by an electron in the planar undulator. The magnetic field on the undulator axis has the form
$\vec{B}(z)=\vec{e}_{y}B_{w}\cos(k_{w}z)~{},$ (41)
Here $k_{w}=2\pi/\lambda_{w}$, and $\lambda_{w}$ is the undulator period. The Lorentz force is used to derive the equation of motion of the electron in the presence of a magnetic field. Integration of this equation gives
$v_{x}(z)=-{c\theta_{s}}\sin(k_{w}z)=-\frac{c\theta_{s}}{2i}\left[\exp(ik_{w}z)-\exp(-ik_{w}z)\right]~{}.$ (42)
Here $\theta_{s}=K/\gamma$, where $K$ is the deflection parameter defined as
$K=\frac{e\lambda_{w}B_{w}}{2\pi mc^{2}}~{},$ (43)
$m$ being the electron mass at rest and $B_{w}$ being the maximal magnetic field of the undulator on axis.
In this case the electron path is given by
$x(z)=r_{w}\cos(k_{w}z)~{},$ (44)
where $r_{w}=\theta_{s}/k_{w}$ is the oscillation amplitude.
We write the undulator length as $L=N_{w}\lambda_{w}$, where $N_{w}$ is the number of undulator periods. With the help of Eq. (38) we obtain an expression, valid in the far zone:
[…]
Here
$\Phi_{T}=\left({\omega\over{2c\bar{\gamma}_{z}^{2}}}+{\omega\theta^{2}\over{2c}}\right)z^{\prime}-{K\theta_{x}\over{\gamma}}{\omega\over{k_{w}c}}\cos(k_{w}z^{\prime})-{K^{2}\over{8\gamma^{2}}}{\omega\over{k_{w}c}}\sin(2k_{w}z^{\prime})~{},$ (47)
where the average longitudinal Lorentz factor $\bar{\gamma}_{z}$ is defined as
$\bar{\gamma}_{z}=\frac{\gamma}{\sqrt{1+K^{2}/2}}~{}.$ (49)
The choice of the integration limits in Eq. (46) implies that the reference system has its origin in the center of the undulator.
Figure 65: Constructive interference of radiation from the successive poles
Figure 66: Geometry for radiation production from an undulator
Usually, it does not make sense to calculate the intensity distribution from Eq. (46) alone, without extra-terms (both interfering and not) from the other parts of the electron path. This means that one should have complete information about the electron path and calculate extra-terms to be added to Eq. (46) in order to have the total field from a given setup. Yet, we can find particular situations for which the contribution from Eq. (46) is dominant with respect to others. In this case Eq. (46), alone, has independent physical meaning.
One of these situations is when the resonance approximation is valid. This approximation does not replace the paraxial one, based on $\gamma^{2}\gg 1$, but it is used together with it. It takes advantage of another parameter that is usually large, i.e. the number of undulator periods $N_{w}\gg 1$. In this case, the integral in $dz^{\prime}$ in Eq. (46) exhibits simplifications, independently of the frequency of interest due to the long integration range with respect to the scale of the undulator period.
A well known expression for the angular distribution of the first harmonic field in the far zone (see Appendix II for a detailed derivation) can be obtained from Eq. (46). Such expression is axis-symmetric, and can, therefore, be presented as a function of a single observation angle $\theta$, where $\theta^{2}=\theta_{x}^{2}+\theta_{y}^{2}$. One obtains the following distribution for the slowly varying envelope of the electric field:
[…]
Here $\omega=\omega_{r}+\Delta\omega$, $C=k_{w}\Delta\omega/\omega_{r}$ and
$\omega_{r}=2k_{w}c\bar{\gamma}_{z}^{2}~{},$ (52)
is the fundamental resonance frequency. Finally $A_{JJ}$ is defined as
$A_{JJ}=J_{0}\left(\frac{K^{2}}{4+2K^{2}}\right)-J_{1}\left(\frac{K^{2}}{4+2K^{2}}\right)~{},$ (53)
$J_{n}$ being the $n$-th order Bessel function of the first kind. The integration over longitudinal coordinate can be carried out leading to the well-known final result:
${\vec{\widetilde{E}}}(z_{0},\vec{\theta})=-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\mathrm{sinc}\left[\frac{L}{2}\left(C+\frac{\omega\theta^{2}}{2c}\right)\right]\vec{e}_{x}~{},$ (54)
where $\mathrm{sinc}(\cdot)\equiv\sin(\cdot)/(\cdot)$. Therefore, the field is horizontally polarized and azimuthal symmetric. Eq. (55) describes a field with spherical wavefront centered in the middle of the undulator.
15.3.2 Why did the Error in Insertion Device Theory Remain so Long Undetected?
We have seen that in all generality the expression for the undulator field in the far zone and in the ultrarelativistic (i.e. paraxial) approximation can be written as Eq. (152). Within the resonance approximation ($N_{w}\gg 1$) for the frequencies around the first harmonic it can be simplified to the well-known expression Eq. (55) where the field is horizontally polarized and azimuthal symmetric. The divergence of this radiation is much smaller compared to the angle $1/\bar{\gamma}_{z}$. The mathematical reason stems from the fact that the factor $\sin(\cdot)/(\cdot)$ represents the well-known resonance character of the undulator radiation. If we are interested in the angular width of the peak around the observation angle $\theta=0$, we can introduce an angular displacement $\Delta\theta$. Taking the first zero of the $\sin(\cdot)/(\cdot)$ function at $C=0$ we will be able to determine the natural angular width of the radiation for the first harmonic $\theta_{c}$. The cone with aperture $\theta_{c}$ is usually called central cone. It can be found that $\theta_{c}^{2}=1/(2N_{w}\bar{\gamma}_{z}^{2})\ll 1/\bar{\gamma}_{z}^{2}$.
Now we would like to understand what is the characteristic transverse size of the field distribution at the exist of the undulator. The radiation from magnetic poles always interferes coherently at zero angle with respect to undulator axis. This interference is constructive within an angle of about $\sqrt{c/(\omega L_{w})}$. We can estimate the interference size at the undulator exit as about $\sqrt{cL_{w}/\omega}$. On the other hand, the electron oscillating amplitude is given by $r_{w}=c\theta_{s}/k_{w}=cK/(\gamma k_{w})$. It follows that $r_{w}^{2}/(cL_{w}/\omega)=K^{2}\omega/(L_{w}k_{w}^{2}\gamma^{2})=K^{2}/[\pi N_{w}(1+K^{2}/2)]\ll 1$, where we use the fact that $\gamma^{2}=(1+K^{2}/2)\bar{\gamma}_{z}^{2}$. This inequality holds independently of the value of $K$, because $N_{w}\gg 1$. Thus, the electron oscillating amplitude is always much smaller than the radiation diffraction size at the undulator exit.
We consider the radiation associated with the first order term in the expansion of the Eq. (158) in power of small parameter $[K\theta_{x}\omega/(\gamma k_{w}c)]$ (see Eq. (162)). But in doing so we miss all information about transverse electron trajectory in the phase factor Eq. (48) since the term $K\theta_{x}\omega\cos(k_{w}z^{\prime})/(\gamma k_{w}c)$ is neglected. In this approximation the electron orbit scale is always much smaller than the radiation diffraction size and Eq. (55) gives fields very much in agreement with the dipole radiation theory. So we are satisfied using the non covariant approach when considering the transverse electron motion.
There are several points to be made about the above result. We have just explained that in accounting only for the radiation in the central cone, we miss all information about the transverse electron motion. To be complete we must add an analysis of the accelerated motion along the $z$-direction (i.e. along the undulator axis). We assume that the transverse velocity $v_{\perp}(z)$ is small compared to the velocity of light $c$. We consider the small expansion parameter $v_{\perp}/c$, neglecting terms of order $(v_{\perp}/c)^{3}$, but not of order $(v_{\perp}/c)^{2}$. In other words we use a second order relativistic approximation for transverse motion. We should remark that the analysis of the longitudinal motion in the ultrarelativistic approximation is much simpler than in the case of transverse motion. It is easy to see that the acceleration in the constant magnetic field yields an transverse electron velocity $v_{\perp}$ and $\Delta v_{z}=-v(v_{\perp}/c)^{2}/2$ parallel to the $z$-axis.
If we evaluate the transformations up to the second order $(v_{\perp}/c)^{2}$, the relativistic correction in the longitudinal motion does not appear. So one should not be surprised to find that, in this approximation, there is no influence of the difference between the non-covariant and covariant constrained electron trajectories on the undulator radiation in the central cone.
15.3.3 Influence of the Kick According to Conventional Theory
Eq. (55) can be generalized to the case of a particle with a given offset $\vec{l}$ and deflection angle $\vec{\eta}$ with respect to the longitudinal axis, assuming that the magnetic field in the undulator is independent of the transverse coordinate of the particle (see Appendix III). Although this can be done using Eq. (174) directly, it is sometimes possible to save time by getting the answer with some trick. For example, in the undulator case one takes advantage of the following geometrical considerations, which are in agreement with rigorous mathematical derivation. First, we consider the effect of an offset $\vec{l}$ on the transverse plane, with respect to the longitudinal axis $z$. Since the magnetic field experienced by the particle does not change, the far-zone field is simply shifted by a quantity $\vec{l}$. Eq. (55), can be immediately generalized by systematic substitution of the transverse coordinate of observation, $\vec{r}_{0}$ with $\vec{r}_{0}-\vec{l}$. This means that $\vec{\theta}=\vec{r}_{0}/z_{0}$ must be substituted by $\vec{\theta}-\vec{l}/z_{0}$, thus yielding
$\widetilde{{E}}\left(z_{0},\vec{l},\vec{\theta}\right)$ $=$ $-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega z_{0}}{2c}\left|\vec{\theta}-\frac{\vec{l}}{z_{0}}\right|^{2}\right]\mathrm{sinc}\left[\frac{\omega L\left|\vec{\theta}-\left({\vec{l}}/{z_{0}}\right)\right|^{2}}{4c}\right]~{}.$ (56)
Let us now discuss the effect of a deflection angle $\vec{\eta}$. Since the magnetic field experienced by the electron is assumed to be independent of its transverse coordinate, the path followed is still sinusoidal, but the effective undulator period is now given by $\lambda_{w}/\cos(\eta)\simeq(1+\eta^{2}/2)\lambda_{w}$. This induces a relative red shift in the resonant wavelength $\Delta\lambda/\lambda\sim\eta^{2}/2$. In practical cases of interest we may estimate $\eta\sim 1/\gamma$. Then, $\Delta\lambda/\lambda\sim 1/\gamma^{2}$ should be compared with the relative bandwidth of the resonance, that is $\Delta\lambda/\lambda\sim 1/N_{w}$, $N_{w}$ being the number of undulator periods. For example, if $\gamma>10^{3}$, the red shift due to the deflection angle can be neglected in all situations of practical relevance. As a result, the introduction of a deflection angle only amounts to a rigid rotation of the entire system. Performing such rotation we should account for the fact that the phase factor in Eq. (56) is indicative of a spherical wavefront propagating outwards from position $z=0$ and remains thus invariant under rotations. The argument in the $\mathrm{sinc}(\cdot)$ function in Eq. (56), instead, is modified because the rotation maps the point $(z_{0},0,0)$ into the point $(z_{0},-\eta_{x}z_{0},-\eta_{y}z_{0})$. As a result, after rotation, Eq. (56) transforms to
$\widetilde{{E}}\left(z_{0},\vec{\eta},\vec{l},\vec{\theta}\right)=-\frac{K\omega eLA_{JJ}}{2c^{2}z_{0}\gamma}\exp\left[i\frac{\omega z_{0}}{2c}\left|\vec{\theta}-\frac{\vec{l}}{z_{0}}\right|^{2}\right]\mathrm{sinc}\left[\frac{\omega L\left|\vec{\theta}-\left({\vec{l}}/{z_{0}}\right)-\vec{\eta}\right|^{2}}{4c}\right]$ (57)
Finally, in the far-zone case, we can always work in the limit for $l/z_{0}\ll 1$, that allows one to neglect the term ${\vec{l}}/{z_{0}}$ in the argument of the $\mathrm{sinc}(\cdot)$ function, as well as the quadratic term in $\omega l^{2}/(2cz_{0})$ in the phase. Thus Eq. (58) can be further simplified, giving the generalization of Eq. (55) in its final form:
$\widetilde{{E}}\left(z_{0},\vec{\eta},\vec{l},\vec{\theta}\right)$ $=$ $-\frac{K\omega eLA_{JJ}}{2c^{2}z_{0}\gamma}\exp\left[i\frac{\omega}{c}\left(\frac{z_{0}\theta^{2}}{2}-\vec{\theta}\cdot\vec{l}\right)\right]\mathrm{sinc}\left[\frac{\omega L\left|\vec{\theta}-\vec{\eta}\right|^{2}}{4c}\right]~{}.$ (59)
It is clear from the above that, according to conventional synchrotron radiation theory, if we consider radiation from one electron at detuning $C$ from resonance, the introduction of a kick only amounts to a rigid rotation of the angular distribution along the new direction of the electron motion. This is plausible, if one keeps in mind that after the kick the electron has the same velocity and emits radiation in the kicked direction owing to the Doppler effect. After such rotation, Eq. (55) transforms into Eq. (59)
15.3.4 Influence of the Kick According to Correct Coupling of Fields and Particles
According to the correct coupling of fields and particles, there is a remarkable prediction of undulator radiation theory concerning to the undulator radiation from the single electron with and without kick. Namely, when a kick is introduced, there is a red shift in the resonance wavelength of the undulator radiation in the velocity direction. To show this, let us consider the covariant treatment, which makes explicit use of Lorentz transformations.
When the kick is introduced, covariant particle tracking predicts a non-zero red shift of the resonance frequency, which arises because in Lorentz coordinates the electron velocity decreases from $v$ to $v-v\theta_{k}^{2}/2$ after the kick, while the velocity of light is unvaried and equal to the electrodynamics constant $c$.
Now the formula Eq. (181) is not quite right, because we should have used not the velocity of electron $v$ but $v-v\theta_{k}^{2}/2$. The shift in the total phase $\Phi_{T}$ under the integral Eq. (174) can be expressed by the formula $\Delta\Phi_{T}=\omega\theta_{k}^{2}z^{\prime}/(2c)$, where we account for that $v\simeq c$ in ultrarelativistic approximation.
Suppose that without kick the electron moves along the constrained trajectory parallel to the undulator axis. The field which produces this electron in the far zone is given by Eq. (55). Referring back to the Eq. (59), we see that the conventional undulator radiation theory gives the following expression for radiation field after the kick
${\vec{\widetilde{E}}}=-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\mathrm{sinc}\left[\frac{L}{2}\left(C+\frac{\omega\left|\vec{\theta}-\vec{\theta}_{k}\right|^{2}}{2c}\right)\right]\vec{e}_{x}~{}.$ (60)
The covariant equations say that, when the kick is introduced, the radiation field in question is given by the formula
${\vec{\widetilde{E}}}=-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\mathrm{sinc}\left[\frac{L}{2}\left(C+\frac{\omega\theta_{k}^{2}}{2c}+\frac{\omega\left|\vec{\theta}-\vec{\theta}_{k}\right|^{2}}{2c}\right)\right]\vec{e}_{x}~{},$ (62)
This formula has nearly, but not quite the same form as Eq. (61), the difference consisting in the term $\omega\theta_{k}^{2}/(2c)$ in the argument of $\mathrm{sinc}$ function. Attention must be called to the difference in resonance frequency between the undulator radiation setup with and without the kick. Remembering the definition of the detuning parameter $C=k_{w}\Delta\omega/\omega_{r}$, we can write the red shift in resonance frequency as $\Delta\omega/\omega_{r}=-\omega_{r}\theta_{k}^{2}/(2k_{w}c)$. With this we also pointed out that the red shift can be written as $\Delta\omega/\omega_{r}=-\gamma^{2}\theta_{k}^{2}/(1+K^{2}/2)$. We now see a second order correction $\theta_{k}^{2}$ that is, however, multiplied by the large factor $\bar{\gamma}_{z}^{2}$.
We are now ready to investigate, more generally, what form the field expression takes under the introduction of a kick. Suppose that, without kick, the electron moves along the trajectory with angle $\vec{\eta}$ with respect to the undulator axis. The field produced by this electron is given by Eq. (59). We let $\vec{\theta}_{k}$ be the kick angle of the electron with respect to its initial motion. The conventional approach gives the following expression for the field after the kick
${\vec{\widetilde{E}}}=-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\mathrm{sinc}\left[\frac{L}{2}\left(C+\frac{\omega\left|\vec{\theta}-\vec{\eta}-\vec{\theta}_{k}\right|^{2}}{2c}\right)\right]\vec{e}_{x}~{}.$ (64)
In contrast, the covariant approach gives
${\vec{\widetilde{E}}}=-\frac{K\omega eL}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\mathrm{sinc}\left[\frac{L}{2}\left(C+\frac{\omega\theta_{k}^{2}}{2c}+\frac{\omega\left|\vec{\theta}-\vec{\eta}-\vec{\theta}_{k}\right|^{2}}{2c}\right)\right]\vec{e}_{x}~{},$ (66)
Now this all leads to an interesting situation. According to the conventional theory, the resonance wavelength depends only on the observation angle with respect to the electron velocity direction. Equation (65) says that for any kick angle $\vec{\theta}_{k}$ and for any angle $\vec{\eta}$ between the undulator axis and the initial electron velocity direction, the radiation along the velocity direction has no red shift. We would like to emphasize a very important difference between conventional and covariant theory. The result of the covariant approach Eq. (67) clearly depends on the absolute value of the kick angle $\theta_{k}$ and the radiation along the velocity direction has the red shift only when the kick angle has nonzero value.
We must conclude that when we accelerate the electron in the lab frame upstream the undulator, the information about this acceleration is included into the covariant trajectory.
15.3.5 Experimental Test of SR Theory in 3rd Generation Light Source
One way to demonstrate incompatibility between the standard approach to relativistic electrodynamics, which deals with the usual Maxwell’s equations, and particle trajectories calculated by using non-covariant particle tracking, is to make a direct laboratory test of synchrotron radiation theory. In other words, we are stating here that, despite the many measurements done during decades, synchrotron radiation theory is not an experimentally well-confirmed theory.
Let us analyze the potential for exploiting synchrotron radiation sources in order to confirm the predictions of corrected synchrotron radiation theory. The emittance of the electron beam in new generation synchrotron radiation sources is small enough, so that one can neglect finite electron beam size and angular divergence in the soft X-ray wavelength range, and such synchrotron radiation source can be examined under the approximation of a filament electron beam. This allows us to take advantage of analytical presentations for single electron synchrotron radiation fields.
Figure 67: Basic setup for the proposed critical experimental test of synchrotron radiation theory with third generation light source. Top: the case for an electron beam without kick. Bottom: the case for the electron beam kicked by an angle $\eta$. In both cases, the X-ray pulse is filtered by a monochromator and the total energy recorded by a detector as a function of the undulator detuning.
The basic setup for a test experiment is sketched in Fig 67. The soft X-ray undulator beam line should be tuned to a minimum photon energy (typically this limit is related with the so called ”water window” wavelength range). The radiation pulse goes through a monochromator filter $F$ and its energy is subsequently measured by the detector. No precise monochromatization of the undulator radiation is required in this case: a monochromator line width $\Delta\omega/\omega\simeq 10^{-3}$ is sufficient. In order for proposed test experiment to be curried out, it is necessary to control the beam kicking e.g. by corrector magnet. In the case of no kick the maximum pulse energy registered by the detector will coincide with the monochromator line tuned to resonance. When the kick is introduced the conventional synchrotron radiation theory still predicts a zero red shift in the resonance wavelength. In contrast to this, one of the immediate consequences of the corrected theory is the occurrence of a non-zero red shift.
The proposed experimental procedure is relatively simple, because is based on relative measurements in the (electron beam) velocity direction with and without transverse kick. Such a measurement is critical, in the sense that the prediction of conventional theory is the absence of red shift and (to our knowledge) has never been performed at synchrotron radiation facilities. However, XFEL based experiment confirm our correction for spontaneous undulator emission (1).
15.4 Synchrotron Radiation from Bending Magnets
Consider a single relativistic electron moving on a circular orbit. The observer in the standard treatment is assumed to be located in a vertical plane tangent to the circular trajectory at the origin, at an angle $\theta$ above the level of the orbit. In other words, in this geometry the $z$ axis is not fixed, but depends on the observer’s position. Note that the geometry of the electron motion has a cylindrical symmetry, with the vertical axis going through the center of the circular orbit. Because of this symmetry, in order to calculate spectral and angular photon distributions, it is not necessary to consider an observer at a more general location.
There are a number of remarkable effects which are a consequence of the cylindrical symmetry and the paraxial approximation. In discussing this we will demonstrate that there is no needs to use covariant particle tracking for derivation of the bending magnet radiation. However, there is one situation where the conventional theory fails. The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion along the magnetic field (i.e with respect to nominal orbit).
It would be well to begin with bird’s view of the main results. We want to solve the electrodynamics problem based on Maxwell’s equations in their usual form. In this case we should analyze the particle evolution within the framework of special relativity, where the problem of assigning Lorentz coordinates to the lab frame in the case of accelerating motion is complicated. The only possibility to introduce Lorentz coordinates in this situation consists in introducing individual coordinate systems (i.e. individual rule-clock structure) for each point of the path.
In order to keep Lorentz coordinates in the lab frame, as discussed before, we need only to perform a clock resynchronization by introducing a time shift and change the scale of time, that is the rhythm of clocks. Because of cylindrical symmetry, the observer is assumed to be located in the vertical plane tangent to the circular trajectory at the origin. Because the formation length is very short, it is necessary to know the velocity and position over only a small arc of the trajectory. In the ultrarelativistic (paraxial) approximation, the observer sees a uniform acceleration of the electron $a=v^{2}/R\simeq c^{2}/R$ in the transverse direction. We can, then, write velocity and offset of the electron as follows $v_{\perp}=at$, $r_{\perp}=at^{2}/2$. Relativity of simultaneity gives time shift $\Delta t=v_{\perp}r_{\perp}/c^{2}=v_{\perp}^{2}t/(2c^{2})$. But a correction of the rhythm is $t\sqrt{1-v_{\perp}^{2}/c^{2}}\simeq t-v_{\perp}^{2}t/(2c^{2})$. We have a beautiful cancellation. This is a coincidence. It is because we have deal with uniform acceleration in the transverse direction using a second order (paraxial) approximation when an electron is moving along an arc of a circle.
15.4.1 Conventional Theory
Consider a single relativistic electron moving on a circular orbit and an observer. It is worth to underline the difference between the more general geometry which we use and the geometry used in most synchrotron radiation textbooks for the treatment of bending magnet radiation. In standard treatments of bending magnet radiation, the horizontal observation angle $\theta_{x}$ is always equal to zero. The reason for this is that most textbooks focus on the calculation of the intensity radiated by a single electron in the far zone, which involves the square modulus of the field amplitude, but do not analyze, for instance, situations like source imaging.
We can use Eq. (38) to calculate the far zone field of radiation from a relativistic electron moving along an arc of a circle. Assuming a geometry with a fixed $z$ we can write the transverse position of the electron as a function of the curvilinear abscissa $s$ as
$\vec{r}(s)=-R\left(1-\cos(s/R)\right)\vec{e_{x}}$ (68)
and
$z(s)=R\sin(s/R)$ (69)
where $R$ is the bending radius.
Since the integral in Eq. (38) is performed along $z$ we should invert $z(s)$ in Eq. (69) and find the explicit dependence $s(z)$:
$s(z)=R\arcsin(z/R)\simeq z+{z^{3}\over{6R^{2}}}$ (70)
so that
$\vec{r}(z)=-{z^{2}\over{2R}}\vec{e_{x}}~{},$ (71)
where the expansion in Eq. (70) and Eq. (71) is justified, once again, in the framework of the paraxial approximation.
With Eq. (38) we obtain the radiation field amplitude in the far zone:
[…]
where
$\Phi_{T}=\omega\left[\left({\theta_{x}^{2}+\theta_{y}^{2}\over{2c}}z_{0}\right)+\left({1\over{2\gamma^{2}c}}+{\theta_{x}^{2}+\theta_{y}^{2}\over{2c}}\right)z^{\prime}\right.$ (73)
$\left.+\left({\theta_{x}\over{2Rc}}\right)z^{\prime 2}+\left(1\over{6R^{2}c}\right)z^{\prime 3}\right]~{}.$ (74)
One can easily reorganize the terms in Eq. (74) to obtain
$\Phi_{T}=\omega\left[\left({\theta_{x}^{2}+\theta_{y}^{2}\over{2c}}z_{0}\right)-{R\theta_{x}\over{2c}}\left({1\over{\gamma^{2}}}+{\theta_{x}^{2}\over{3}}+\theta_{y}^{2}\right)\right.$ (75)
$\left.+\left({{1\over{\gamma^{2}}}+\theta_{y}^{2}}\right){\left(z^{\prime}+R\theta_{x}\right)\over{2c}}+{\left(z^{\prime}+R\theta_{x}\right)^{3}\over{6R^{2}c}}\right]~{}.$ (76)
With redefinition of $z^{\prime}$ as $z^{\prime}+R\theta_{x}$ under integral we obtain the final result:
[…]
×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\theta_{y}^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (78)
where
$\Phi_{s}={\omega z_{0}\over{2c}}\left(\theta_{x}^{2}+\theta_{y}^{2}\right)$ (79)
and
$\Phi_{0}=-{\omega R\theta_{x}\over{2c}}\left({1\over{\gamma^{2}}}+{\theta_{x}^{2}\over{3}}+\theta_{y}^{2}\right)~{}.$ (80)
In standard treatments of bending magnet radiation, the phase term $\exp(i\Phi_{0})$ is absent. In fact, the horizontal observation angle $\theta_{x}$ is always equal to zero.
15.4.2 Why did the Error in Synchrotron Radiation Remain so Long Undetected?
Our case of interest is an ultrarelativistic electron accelerating in a circle. As already remarked, in conventional (non-covariant) particle tracking the description of the dynamical evolution in the lab frame is based on the use of the absolute time convention. In this case simultaneity is absolute, and we only need one set of synchronized clocks in the lab frame, to be used for the description of the accelerated motion. However, the use of the absolute time convention automatically implies the use of much more complicated field equations, and these equations are different for each value of the particle velocity i.e. for each point along its path. This is the reason to prefer the covariant approach within the framework of both dynamics and electrodynamics.
We start by considering an electron moving along a circular trajectory that lies in the $(x,z)$-plane and tangent to the $z$ axis. Because of cylindrical symmetry, in order to calculate spectral and angular photon distributions, it is not necessary to consider an observer at general location. The observer is assumed to be located in the vertical plane tangent to the circular trajectory at the origin. In ultrarelativistic (paraxial) approximation we evaluate transformations working only up to the order of $v^{2}_{x}/c^{2}$. The restriction to this order provides an essential simplicity of calculations. We can interpret manipulation with rule-clock structure in the lab frame simply as a change of variables according to the transformation Eq. (13): $x_{L}=\gamma_{x}x$, $t_{L}=(t/\gamma_{x}+\gamma_{x}xv_{x}/c^{2})$. We are dealing with a second order approximation and $\gamma_{x}=1+v_{x}^{2}/(2c^{2})$. The overall combination of Galilean transformation and variable changes actually yields to the transverse Lorentz transformation. Since the Galilean transformation, completed by the introduction of the new variables, is mathematically equivalent to a Lorentz transformation, it obviously follows that transforming to new variables leads to the usual Maxwell’s equations.
In order to keep Lorentz coordinates in the lab frame, as discussed before, we need only to perform a clock resynchronization by introducing the time shift $\Delta t=t_{L}-t=-[v_{x}^{2}/(2c^{2})]t+xv_{x}/c^{2}$. The relativistic correction to the particle’s offset ”$x$” does not appear in this expansion order, but only in order of $v_{x}^{3}/c^{3}$ and $x_{L}=x$ in our case of interest. Although we have only shown that time shift in one rather special case, the result is right for any offset and (transverse) velocity direction: $\Delta t=t_{L}-t=-[|\vec{v}_{\perp}|^{2}/(2c^{2})]z^{\prime}/c+\vec{r}_{\perp}\cdot\vec{v}_{\perp}/c^{2}$. To finish our analysis we need only find a relativistic correction to the longitudinal motion. We remark again that if we evaluate the transformations up to the second order $(v_{\perp}/c)^{2}$, the relativistic correction in the longitudinal motion does not appear in this approximation. We have demonstrated the covariant method that can be used for any trajectory - a general way of funding what happens directly in space-frequency domain and in paraxial approximation.
Let us now see how to apply this covariant method to a special situation. Let’s use our knowledge of the relativistically correct method for calculating synchrotron radiation emission to find the photon angular-spectral density distributions from a bending magnet. In the ultrarelativistic approximation, we have a uniform acceleration of the electron $a=v^{2}/R\simeq c^{2}/R$ in the transverse direction. We can, then, write velocity and offset of the electron as follows $v_{x}=at=az^{\prime}/v=az^{\prime}/c$, $x=at^{2}/2=az^{\prime 2}/(2c^{2})$. We have now all quantities we wanted. Let us put them all together in relativistic time shift: $\Delta t=t_{L}-t=-a^{2}z^{\prime 3}/(2c^{5})+a^{2}z^{\prime 3}/(2c^{5})=0$. There is no difference! We do not need to use covariant particle tracking for derivation of the bending magnet radiation. Usually, such a beautiful cancellation is found to stem from a deep underlying principle. Nevertheless, in this case there does not appear to be any such profound implication. This is a coincidence.
This cancellation is not surprising, if one analyzes the general expression for the radiation field from bending magnet in the far zone Eq.(78). In our previous discussion of undulator radiation, we learned that the relativistic correction appears only when the transverse electron trajectory is included in the total phase $\Phi_{T}$ under the integral Eq.(38). Referring back to Eq.(40) for the phase factor $\Phi_{T}$ , we see that the term which depends on the transverse position of the electron can be written as $\exp i(\omega/c)[\theta_{x}x(z^{\prime})+\theta_{y}y(z^{\prime})]$. We conclude that the observation angle in the total phase factor under the integral must be related with the contribution of the transverse electron trajectory. Now look at Eq.(78). This equation includes only the observation angle $\theta_{y}$ in the phase factor under the integral. This means that the transverse constraint motion of the electron in the bending magnet does not affect synchrotron radiation. So we are justified using a non-covariant approach for considering the constrained electron motion along the nominal orbit in $(x,z)$-plane.
We point out that the cancellation in relativistic time shift and the independence of the Fraunhofer propagator (to be more precise, in space-frequency domain we are dealing with a paraxial approximation of Green’s function of nonhomogeneous Helmholtz equation) on the observation angle $\theta_{x}$ in the far zone can be regarded as the two sides of the same coin: they are manifestation of the cylindrical symmetry when an electron is moving along an arc of a circle. Because of cylindrical symmetry, in order to calculate spectral and angular photon distributions in the far zone, it is not necessary to consider an observer at a general location. The observer is assumed to be located in the vertical plane tangent to the circular trajectory at the origin. In this case observation angle $\theta_{x}=0$ and the observation angle $\theta_{y}$ is above the level of the orbit. In other words, in this very special geometry the $z$-axis is not fixed, but depends on the observer position. However, this way of proceeding can hardly help to obtain radiation fields in the near zone. Indeed, in the near zone we are dealing with the Fresnel propagator, which obviously depends on the constrained motion of the electron. We use far-zone arguments only to show that there is no influence of the difference between the non-covariant and covariant trajectories on the synchrotron radiation from bending magnets. The cancellation in the relativistic time shift leads to the same outcome in the near zone as it must be.
15.4.3 Influence of the Kick According to Conventional Theory
Up to this point we considered an electron moving along a circular trajectory that lies in the $(x,z)$-plane and tangent to the $z$ axis. The phase difference in the fields will be determined by the position of the observer and by the electron trajectory. Let us now discuss the bending magnet radiation from a single electron with arbitrary angular deflection and offset with respect to the nominal orbit.
Approximation for the electron path (see Eq. (168), Eq. (171) in the Appendix III) can be used to characterize the field from an electron moving on any trajectory. Using Eq. (70) and Eq. (71) an approximated expression for $s(z)$ can be found:
$s(z)=z+{z^{3}\over{6R^{2}}}+{z^{2}\eta_{x}\over{2R}}+{z\eta_{x}^{2}\over{2}}+{z\eta_{y}^{2}\over{2}}$ (81)
so that
$\vec{v}_{\perp}(z)=\left(-{vz\over{R}}+v\eta_{x}\right)\vec{e_{x}}+\left(v\eta_{y}\right)\vec{e_{y}}~{}$ (82)
and
$\vec{r}(z)=\left(-{z^{2}\over{2R}}+\eta_{x}z+l_{x}\right)\vec{e_{x}}+\left(\eta_{y}z+l_{y}\right)\vec{e_{y}}~{}.$ (83)
It is evident that the offsets $l_{x}$ and $l_{y}$ are always subtracted from $x_{0}$ and $y_{0}$ respectively: a shift in the particle trajectory on the vertical plane is equivalent to a shift of the observer in the opposite direction. With this in mind we introduce angles $\bar{\theta}_{x}=\theta_{x}-l_{x}/z_{0}$ and $\bar{\theta}_{y}=\theta_{y}-l_{y}/z_{0}$ to obtain
[…]
and
$\Phi_{T}=\omega\left({\bar{\theta}_{x}^{2}+\bar{\theta}_{y}^{2}\over{2c}}z_{0}\right)+{\omega\over{2c}}\left({1\over{\gamma^{2}}}+\left(\bar{\theta}_{x}-\eta_{x}\right)^{2}+\left(\bar{\theta}_{y}-\eta_{y}\right)^{2}\right)z^{\prime}$ (85)
$+\left({\omega(\bar{\theta}_{x}-\eta_{x})\over{2Rc}}\right)z^{\prime 2}+\left(\omega\over{6R^{2}c}\right)z^{\prime 3}~{}.$ (86)
One can easily reorganize the terms in Eq. (86) to obtain
$\Phi_{T}=\omega\left({\bar{\theta}_{x}^{2}+\bar{\theta}_{y}^{2}\over{2c}}z_{0}\right)-{\omega R(\bar{\theta}_{x}-\eta_{x})\over{2c}}$ (87)
×$\times\left({1\over{\gamma^{2}}}+(\bar{\theta}_{y}-\eta_{y})^{2}+{(\bar{\theta}_{x}-\eta_{x})^{2}\over{3}}\right)$ (88)
$+\left({{1\over{\gamma^{2}}}+(\bar{\theta}_{y}-\eta_{y})^{2}}\right){\omega\left(z^{\prime}+R(\bar{\theta}_{x}-\eta_{x})\right)\over{2c}}$ (89)
$+{\omega\left(z^{\prime}+R(\bar{\theta}_{x}-\eta_{x})\right)^{3}\over{6R^{2}c}}~{}.$ (90)
Redefinition of $z^{\prime}$ as $z^{\prime}+R(\bar{\theta}_{x}-\eta_{x})$ gives the result
[…]
×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}(\bar{\theta}_{y}-\eta_{y})^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (92)
where
$\Phi_{s}={\omega z_{0}\over{2c}}\left(\bar{\theta}_{x}^{2}+\bar{\theta}_{y}^{2}\right)$ (93)
and
$\Phi_{0}=-{\omega R(\bar{\theta}_{x}-\eta_{x})\over{2c}}\left({1\over{\gamma^{2}}}+(\bar{\theta}_{y}-\eta_{y})^{2}+{(\bar{\theta}_{x}-\eta_{x})^{2}\over{3}}\right)~{}.$ (94)
In the far zone we can neglect terms in $l_{x}/z_{0}$ and $l_{y}/z_{0}$, which leads to
[…]
×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\left(\theta_{y}-\eta_{y}\right)^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (96)
where
$\Phi_{s}={\omega z_{0}\over{2c}}\left(\theta_{x}^{2}+\theta_{y}^{2}\right)$ (97)
and
$\Phi_{o}\simeq-{\omega R({\theta_{x}}-\eta_{x})\over{2c}}\left({1\over{\gamma^{2}}}+(\theta_{y}-\eta_{y})^{2}+{(\theta_{x}-\eta_{x})^{2}\over{3}}\right)-{\omega\over{c}}(l_{x}\theta_{x}+l_{y}\theta_{y})~{}.$ (98)
It is clear from the above that the field distribution in the far zone depends only on the observation angle with respect to the electron velocity direction.
According to the conventional (incorrect) coupling of fields and particles, there is a prediction of radiation theory concerning to the bending magnet radiation from a single electron with and without kick. Namely, when a kick is introduced, there is a rigid rotation of the angular distribution in the far zone.
15.4.4 Influence of the Kick According to Correct Coupling of Fields and Particles
Let us discuss the covariant treatment, which makes explicit use of Lorentz transformations. Consider the bending magnet radiation from a single electron with a kick with respect to the nominal orbit in $(x,z)$-plane. In this case, we additionally have a translation along the $y$-axis with constant velocity $v_{y}=v\theta_{k}$. We can, then, write the offset of the electron as follows $y=\theta_{k}z^{\prime}$. Let’s put velocity and offset in the relativistic time shift: $\Delta t=t_{L}-t=-\theta_{k}^{2}z^{\prime}/(2c)+\theta_{k}^{2}z^{\prime}/c=\theta_{k}^{2}z^{\prime}/(2c)$. So, the shift in the total phase under the integral along the path can be expressed by the formula $\Delta\Phi_{T}=\omega\theta_{k}^{2}z^{\prime}/(2c)$. The result agrees with our red shift calculation in the undulator case when the kick is introduced, as it must be.
We would like to make a historical note. The difference between covariant and non-covariant particle trajectories was never understood. So, accelerator physicists did not appreciate that there was a contribution to the synchrotron radiation from relativistic kinematics effects. The question now arises how can storage rings actually operate. The point is that this example deals with a situation where electron beam kinetics is determined by the emission of synchrotron radiation from bending magnets. However, because of the cylindrical symmetry, covariant and non-covariant solutions for the electron motion along an arc of a circle yield similar properties of synchrotron radiation except the following modification. The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion in the vertical direction. But synchrotron radiation from bending magnets is emitted within a wide range of frequencies, and the output intensity is not sensitive on the red shift.
15.5 How to Solve Problems Involving Many Trajectory Kicks
We shall now discuss the situation where there are $n$ arbitrary spaced kickers, all different from one another in terms of the rotation angle introduced. Let us consider how we may apply covariant particle tracking in this circumstance, and try to understand what is happening when we have for example an undulator downstream of the kicker setup. Formally, if one wants to calculate the radiation from the undulator one should take into account all kicks in the electron trajectory, from the generation of the electron. However, this situation is not surprising, if one analyzes the general expression for the radiation field from a single electron Eq.(LABEL:revwied). In fact, we should note that, in general, one needs to know the entire history of the electron from $t^{\prime}=-\infty$ to $t^{\prime}=\infty$ since the integration in Eq.(LABEL:revwied) is performed (at least formally) between these limits. However, this statement should be interpreted physically, depending on the situation under study: integration should in fact be performed from and up to times when the electron does not contribute to the field anymore.
We should pointed out that it is the electrodynamics theory, which ultimately decides what part of the particle trajectory is important for calculating undulator radiation and what part can be neglected. The most important, general statement concerning the relevant part of the particle trajectory, is that it must be calculated according to the covariant method (if one wants to use the usual Maxwell’s equations).
Let us consider the ultrarelativistic assumption $1/\gamma^{2}\ll 1$, which is verified for synchrotron radiation setups. In general, the introduction of a small parameter in any theory brings simplifications. The ultrarelativistic approximation implies a paraxial approximation and Eq.(LABEL:revwied) can be simplified to Eq.(38). Suppose that we take a situation in which the rotation angle of the first bending magnet upstream of the undulator is much larger than $1/\gamma$. In other words, we now consider an electron moving along a standard synchrotron radiation setup. The electron enters the setup via a bending magnet, passes through a straight section, an undulator, and another straight section. Finally, it leaves the setup via another bend. Note that, although the integration in Eq.(38) is performed from $\infty$- ∞ to $\infty$, the only (edge) part of the trajectory into the bending magnets contributing to the integral is of order of the radiation formation length $L_{f}$. Mathematically, it is reflected in the fact that $\Phi_{T}(z^{\prime})$ in Eq.(38) exhibits more and more rapid oscillations as $z^{\prime}$ becomes larger than the formation length. At the critical wavelength the formation length is simply of order of $R/\gamma$, $R$ being the radius of the bend. That simply corresponds to an orbiting angular interval $\Delta\theta\simeq 1/\gamma$. Typically, the critical wavelength of the radiation from a bending magnet in synchrotron radiation source is about 0.1 nm and the formation length in this case is only few millimeters.
[…]
Figure 68: Standard undulator radiation setup. When the electron passes through a bending magnet there is the synchrotron radiation, washing out the (Frank- Ginzburg) fields of the fast moving charge. At the exit of the bending magnet we have ”naked” electron. There is a process of formation of the field-dressed electron within the formation length from the very beginning of the straight section downstream the bending magnet. The field-dressed ultrarelativistic electron has a visible transverse size of order a few microns for third generation synchrotron radiation sources.
[…]
The electron trajectory being divided into two essentially different parts: before and after the switcher. If we accelerate the electron in the lab frame upstream of the switcher, the information about this acceleration is included into the first part of the covariant trajectory. But this acceleration prehistory (together with the fields of the ultrarelativistic electron) is washed out during the switching process and at the entrance of the straight section we have a ”naked” electron.
We start with the description of the field formation process along the straight section downstream of the switcher, based on the covariant approach. First of all we have to synchronize distant clocks within the lab frame. The synchronization procedure that follows is the usual Einstein synchronization procedure. It is assumed that in the Lorentz lab frame the electron proceeds following a rectilinear trajectory with velocity $v$. This assumption is used as initial condition. Then we can analyze situation downstream the switcher by using the usual Maxwell’s equations.
When one analyzes the process of ”field-dressed” electron formation from the viewpoint of the non covariant approach, one assumes the same initial conditions (rectilinear trajectory with velocity $v$) for the electron motion. Then one solves the electrodynamics problem of fields formation by using the usual Maxwell’s equations. We already mentioned that the type of clock synchronization which results in time coordinate $t$ in an electron trajectory $\vec{x}(t)$ is never discussed in accelerator physics. However, we know that the usual Maxwell’s equations are only valid in the Lorentz frame. The non covariant approach is obviously based on a definite synchronization assumption, but this is actually a hidden assumption. In our case of interest, within the lab frame the Lorentz coordinates are then automatically enforced. So one should not be surprised to find that in this simple case of rectilinear motion there is no difference between covariant and non covariant calculations of the initial conditions at the undulator entrance.
Because of the characteristics of undulator radiation, in order to calculate the radiation field within the central cone, we simply use the instantaneous (i.e. dipole-like) theory of radiation. So we are satisfied using a non covariant approach for considering the constrained motion along the undulator. We conclude that it does not matter which approach is used to describe the standard synchrotron radiation setup. The two approaches, treated according to Einstein’s or absolute time synchronization conventions give the same result for the radiation within the central cone.
Let us now see what happens with a weak dipole magnet (a kicker), which is installed in the straight section upstream of the undulator and is characterized by a small kick angle $(\gamma\theta_{k})^{2}\ll 1$. What do we expect for the undulator radiation? At first glance the situation is similar to the switcher setup and the electron trajectory is again divided into two parts: before and after the kicker. The most important difference, however, is that electrodynamics now dictates that both trajectories are important for the calculation of the undulator radiation. When the electron passes through the kicker there is no synchrotron radiation (to be more precise, in this case radiation is indistinguishable from the self-electromagnetic fields of the electron), washing out the virtual radiation fields like in the switcher case. We expect that an electron that passes through a kicker is still ”field-dressed”, but we have an electron whose fields has been perturbed, and now include information about the acceleration with respect to an inertial frame.
According to the conventional theory, as usual for Newtonian kinematics, the Galilean vectorial law of addition of velocities is actually used. Non-covariant particle dynamics shows that the direction of the electron trajectory changes after the kick, while its speed remains unvaried. In contrast, covariant particle tracking, which is based on the use of Lorentz coordinates, yields different results for the trajectory of the electron. The electron speed decreases from $v$ to $v(1-\theta^{2}_{k}/2)$. This result is at odds with the prediction from non-covariant particle tracking, because the corrected relativistic addition law for non-parallel velocities is used to calculate the electron trajectory.
According to the conventional algorithm for solving electrodynamics field equations, which deals with the usual Maxwell’s equations, and particle trajectories calculated by using non-covariant particle tracking, the undulator radiation along the velocity direction has no red shift of resonance frequency for any kick angle $\theta_{k}$. According to the correct coupling of fields and particles, there is a remarkable prediction of synchrotron radiation theory concerning the setup described above. Namely, there is a red shift of the resonance frequency of the undulator radiation in the kicked direction. The red shift can be expressed by the formula $\Delta\omega_{r}/\omega_{r}=-\gamma^{2}\theta_{k}^{2}/(1+K^{2}/2)$.
15.6 Synchrotron Radiation in the Case of Particle Motion on a Helix
The presence of red shift in bending magnet radiation automatically implies the same problem for conventional cyclotron radiation theory. In the ultrarelativistic limit, there are well-known analytical formulas that describe the spectral and angular distribution of cyclotron radiation emitted by an electron moving in a constant magnetic field having a non-relativistic velocity component parallel to the field, and an ultrarelativistic velocity component perpendicular to it. According to the conventional approach, exactly as for the bending magnet case, the angular-spectral distribution of radiation is a function of the total velocity of the particle due, again, to the Doppler effect. In contrast, the covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion in the magnetic field direction. It should be note that cyclotron-synchrotron radiation emission is one of the most important processes in plasma physics and astrophysics and our corrections are important for a much wider part of physics than that of synchrotron sources.
15.6.1 Existing Theory
Let us discuss in some detail the relativistic cyclotron radiation. Here we shall only give some final results and discuss their relation with the conventional synchrotron radiation theory from bending magnet. In the case of an uniform translation motion with non-relativistic velocity along the magnetic field direction (Fig. 69), a widely accepted (in astrophysics) expression for the angular and spectral distributions of radiation from an ultra-relativistic electron on a helical orbit is given by (2)
${\vec{\widetilde{E}}}(\chi,\alpha)$$\sim\Bigg{\{}\vec{e}_{x}\left[(\xi^{2}+\psi^{2})K_{2/3}\left(\frac{\omega}{2\omega_{c}}\left(1+\frac{\psi^{2}}{\xi^{2}}\right)^{3/2}\right)\right]$ (100)
$-i\vec{e}_{y}\left[(\xi^{2}+\psi^{2})^{1/2}\psi K_{1/3}\left(\frac{\omega}{2\omega_{c}}\left(1+\frac{\psi^{2}}{\xi^{2}}\right)^{3/2}\right)\right]\Bigg{\}}~{},$
where $K_{1/3}$ and $K_{2/3}$ are the modified Bessel functions, $\xi=1/\gamma$, $\psi=\chi-\alpha$, ($\chi$ is the angle between $\vec{v}$ and $\vec{B}$ and $\alpha$ that between $\vec{n}$ and $\vec{B}$); the angle $\psi$ is clearly the angular distance between the direction of the electron velocity $\vec{v}$ and the direction of observation $\vec{n}$. Here the $\omega_{c}$ is defined by $3eB\gamma^{2}/(2mc)$.
Figure 69: Geometry for radiation production from helical motion.
Actually we have already discussed radiation from an ultrarelativistic electron on a helical orbit in the previous section. Equation Eq. (96) is the result we worked out above for the bending magnet radiation from a single electron with angular deflection with respect to nominal orbit. Eq. ( 100) does not look the same as Eq. (96). It will, however, if we now define the small deflection angle $\eta_{y}=\pi/2-\chi$ and the observation angle $\theta_{y}=\pi/2-\alpha$ (the observer is also assumed to be located in the vertical plane tangent to the trajectory i.e. $\theta_{x},\eta_{x}=0$). The integrals in Eq. (96) can be expressed in terms of the modified Bessel functions:
[…]
Then, making the necessary variable changes, the formula reduces to Eq.( 100).
15.6.2 Methodology of Solving Problems Involving Boosts
The calculation leading to Eq. ( 100) is rather elaborate. It is therefore desirable to have an independent derivation. The simplest way of analyzing the radiation for an ultrarelativistic helical motion makes use of the theory of relativity and involves practically no calculations. The way for computing the radiation in the case of uniform translation is simple. One describes a complicated situation by finding a coordinate system where the analysis is already done (radiation in the case of circular motion) and transforms back to the old coordinate system. The reference system $S^{\prime}$ in which the electron moves in circular motion can be transformed to a reference system $S$ in which the electron proceeds following a helical trajectory. Eq. ( 100) holds, indeed, in the frame $S$ for a particle whose velocity is $(v_{x},v_{y},v_{z})=(v_{0}\sin\chi\sin\phi,v_{0}\cos\chi,v_{0}\sin\chi\cos\phi)$. The Lorentz transformation, which leads to the value $v_{y}=v_{0}\cos\chi$ for the $y$-component of the velocity yields $(v_{x},v_{y},v_{z})=(v^{\prime}\sin\phi^{\prime}/\gamma_{y},v_{y},v^{\prime}\cos\phi^{\prime}/\gamma_{y})$, where $\gamma_{y}=1/\sqrt{1-v^{2}_{y}/c^{2}}$, $v^{\prime}$ is the velocity of the electron in the frame $S^{\prime}$ and the phase angle $\phi^{\prime}=\phi$ is invariant. This means that, in order to end up in $S$ with a transverse (to the magnetic field direction) velocity $v_{\perp}=v_{0}\sin\chi$, one must start in $S^{\prime}$ with $v^{\prime}=\gamma_{y}v_{0}\sin\chi$. In the ultrarelativistic approximation $\gamma_{\perp}^{2}=1/(1-v^{2}_{\perp}/c^{2})\gg 1$, and one finds the simple result $v_{0}=v^{\prime}$, so that a Lorentz boost with non-relativistic velocity $v_{y}$ leads to a rotation of the particle velocity $\vec{v}_{0}$ of the angle $\eta=\pi/2-\chi\simeq v_{y}/c\ll 1$ (if angle $\eta$ is small and $v_{0}\simeq c$, we would write $\gamma_{y}\sin\chi\simeq 1$). If one transforms the radiation field for a particle in a circular motion in the system $S^{\prime}$, one obtains the result that the effect of a boost amounts to a rigid rotation of the angular-spectral distribution of the radiation emitted by the electron moving with velocity $v_{0}$ on a circle that is, once more, Eq. ( 100) (3) . The notation ”” used here to indicate the velocity in the transverse to the magnetic field direction should not be confused with the ”” referring to an acceleration of the electron in the transverse direction in the proceeding sections.
It comes out quite naturally that the covariant way of analyzing the radiation for helical motion considered above is based on the Lorentz transformation. In other words, within the lab frame the Lorentz coordinates are automatically enforced. It assumed that in the Lorentz lab frame the electron proceeds following a helical trajectory with velocity $v_{0}$. This is employed as initial condition. In the ultrarelativistic approximation a Lorentz boost along the field direction with non relativistic velocity $v_{y}$ leads to the circular motion of the electron with the same velocity $v_{0}$. Thus the single boost along the field direction will leave the radiation properties unchanged.
Let us discuss the synchrotron radiation from a single electron with a kick with respect to the nominal orbit in $(x,z)$-plane. In this case, the acceleration yields a particle velocity increment $v_{y}=v\theta_{k}$ parallel to the $y$-axis and $\Delta v_{z}=-v\theta^{2}_{k}/2$ parallel to the $z$-axis. The restriction to the second order provides an essential simplicity of the calculations. We can use a sequence of two commuting non-collinear Lorentz boosts in order to discuss the particle motion downstream the kicker.
When the kick is introduced, the covariant particle tracking predicts a change of the initial condition at the entrance of the synchrotron radiation setup. If we will keep the Lorentz coordinate system in the lab frame downstream of the kicker, we will find that the covariant velocity on the helical orbit after the kick decreases from $v$ to $v-v\theta^{2}_{k}/2$ and the covariant way of analyzing the radiation for a helical motion with covariant velocity $v_{0}=v-v\theta^{2}_{k}/2$ considered above will leads to a red shift in the critical wavelength.
It should be note that the single passive Lorentz boost to the reference system $S^{\prime}$ discussed above is only a mathematical trick. A good way to think of this transformation is to regard it as a result of transformation to new variables. According to this kinematic transformation, the electron (and the observer) motion relative to the fixed stars does not changes. An absolute (i.e. physically real) acceleration is always a dynamical process. We must conclude that when we accelerate with respect to the fixed stars an electron in the lab inertial frame upstream the uniform magnetic field, this acceleration is absolute and the information about this acceleration is included into the covariant trajectory.
15.6.3 On the Advanced ”Paradox” Related to the Coupling Fields and Particles
We now want to point out that there are two different sets of initial conditions resulting in the same uniform translation along the magnetic field direction in the Lorentz lab frame. We start by considering an electron moving along a circular trajectory that lies in the $(x,z)$-plane. We then rotate the magnetic field vector $\vec{B}$ in the $(y,z)$-plane by an angle $\theta_{0}$, assuming that rotation angle is small $(\theta_{0}\gamma)^{2}\ll 1$. We consider a situation in which the electron is in uniform motion with velocity $v\theta_{0}$ along the magnetic field direction. It is clear that if we consider the radiation from an electron moving on a circular orbit, the introduction of the magnetic field vector rotation will leave the radiation properties unchanged. This is plausible, if one keeps in mind that after rotating the bending magnet, the electron has the same velocity and emits radiation in the velocity direction owing to the Doppler effect. After the rotation, correction to the curvature radius $R$ is only of order $\theta_{0}^{2}$ and can be neglected.
Now we consider another situation. Let us see what happens with a kicker, which is installed in the straight section upstream of the bending magnet and is characterized by a kick angle $\theta_{k}=\theta_{0}$. When the kick in the $y$ direction is introduced, there is a red shift of the critical wavelength which arise because, according to corrected addition velocities law, the electron velocity decreases from $v$ to $v-v\theta_{0}^{2}/2$ after the kick. The red shift of the critical frequency $\omega_{c}$ can be expressed by the formula $\Delta\omega_{c}/\omega_{c}=-(3/2)\gamma^{2}\theta_{0}^{2}$. We see a second order correction $\theta_{0}^{2}$ that is, however, multiplied by a large factor $\gamma^{2}$. The result of the covariant approach clearly depends on the absolute value of the kick angle $\theta_{0}$ and the radiation along the velocity direction has a red shift only when the kick angle has nonzero value. The implicit ”absolute” acceleration means acceleration relative to the fixed stars.
The difference between these two situations, ending with a final uniform translation along the direction of the magnetic field is very interesting. It comes about as the result of the difference between two Lorentz coordinate systems in the lab frame. By trying to accelerate the electron upstream the bending magnet we have changed Lorentz coordinates for that particular source. We know that in order to keep a Lorentz coordinates system in the lab frame after the kick we need to perform a clock resynchronization. So we should expect the electron velocity to be changed. The difference between the two setups is understandable. When we do not perturb the electron motion (relative to the fixed stars) upstream of the bending magnet, no clock resynchronization takes place, while when we do perturb the motion, clock resynchronization is introduced.
Figure 70: Two sets of initial conditions resulting in the same uniform motion along the magnetic field direction in the case of absolute time coordinatization in the lab frame. The magnitude of the electron velocity and the orientation of the velocity vector with respect to the magnetic field vector are identical in both setups.
We would now like to describe an apparent paradox. The argument that the difference between these two situations, ending with a final uniform translation along the magnetic fields direction, is paradoxical can be summarized in the following way: in the case of absolute time coordinatization in the lab frame, the initial conditions at the bending magnet entrance are apparently identical. In fact, the magnitude of the electron velocity and the orientation of the velocity vector with respect to the magnetic field vector are identical in both setups. We must conclude that when we accelerate the electron in the lab frame upstream the bending magnet, the information about this acceleration is not included into the non-covariant trajectory. Where is the information about the electron acceleration recorded in the case of absolute time coordinatization? Since an electron is a structureless particle, the situation seems indeed paradoxical.
The above statement includes one delicate point. It is not true that an ultrarelativistic electron is a ”structureless particle”. Electrodynamics deals with observable quantities. Let us consider the measurement of the red shift in the bending magnet radiation from our kicked electron. We can measure the accurate value of the red shift using a spectrometer in the lab frame, and this leads to a description of the setup in the space-frequency domain.
[…]
When the electron passes through a kicker, its fields are perturbed, and now include information about the acceleration. According to the old kinematics, the orientation of the virtual radiation phase front is unvaried. However, Maxwell’s equations do not remain invariant with respect to Galilean transformations and, as discuss throughout this book, the choice of the old kinematics implies using anisotropic field equations. As a result, the phase front remains plane but the direction of propagation is not perpendicular to the phase front. In other words, the radiation beam motion and the radiation phase front normal have different directions. Then, having chosen the absolute time synchronization, electrodynamics predicts that the virtual radiation beam propagates in the kicked direction with the phase front tilt $\theta_{k}$. This is the key to the ”paradox” discussed here. The information about the electron acceleration is recorded in the perturbation of the self-electromagnetic fields of the electron. Mathematically information is recorded in the phase front tilt of the virtual radiation beam.
15.7 Bibliography and Notes
1. It should be note that results of the beam splitting experiment at LCLS confirm our correction for spontaneous undulator emission [62]. It apparently demonstrated that after a modulated electron beam is kicked on a large angle compared to the divergence of the XFEL radiation, the modulation wavefront is readjusted along the new direction of the motion of the kicked beam. Therefore, coherent radiation from the undulator placed after the kicker is emitted along the kicked direction practically without suppression (see the Chapter 16 for more detail). In the framework of the conventional theory, there is a second outstanding puzzle concerning the beam splitting experiment at the LCLS. In accordance with conventional undulator radiation theory, if the modulated electron beam is at perfect resonance without kick, then after the kick the same modulated beam must be at perfect resonance in the velocity direction. However, experimental results clearly show that when the kick is introduced there is a red shift in the resonance wavelength. The maximum power of the coherent radiation is reached when the undulator is detuned to be resonant to the lower longitudinal velocity after the kick [62]. It should be remarked that any linear superposition of a given radiation field from single electrons conserves single-particle characteristics like parametric dependence on undulator parameters and polarization. Consider a modulated electron beam kicked by a weak dipole field before entering a downstream undulator. Radiation fields generated by this beam can be seen as a linear superposition of fields from individual electrons. Now experimental results clearly show that there is a red shift in the resonance wavelength for coherent undulator radiation when the kick is introduced. It follows that the undulator radiation from the single electron has red shift when the kick is introduced as well. This argument suggests that results of the beam splitting experiment in reference [62] confirm our correction for spontaneous undulator emission.
2. A widely accepted expression for the angular and spectral distributions of radiation from an ultra-relativistic electron on a helical orbit were calculated in [57, 58]. At present, relativistic cyclotron radiation results are textbook examples (see e.g. [59]) and do not require a detail description.
3. The covariant way of analyzing the radiation for helical motion was considered in [60]. It is generally believed that $\vec{x}(t)=\vec{x}(t)_{cov}$ and this is the reason why in the [60] there is no distinction between the two (non covariant and covariant) approaches to describe the electron motion on a helix downstream of the kicker setup.
16 Relativity and X-Ray Free Electron Lasers
16.1 Introductory Remarks
In the previous chapter we attempted to answer the question of why the error in radiation theory should have so long remained undetected. According to covariant approach, the various relativistic kinematics effects concerning to the synchrotron radiation setup, turn up in successive orders of approximation. Instead of small (total) velocity parameter $(v/c)$ in the non-relativistic case, we use a small transverse velocity parameter $(v_{\perp}/c)$. In our previous discussion of bending magnet radiation, we learned that a motion of the single ultrarelativistic electron in a constant magnetic field, according to the theory of relativity, influences the kinematics terms of the second order $(v_{\perp}/c)^{2}$ only. It is demonstrated that due to a combination of the ultrarelativistic (i.e. paraxial) approximation and a very special symmetry of the conventional synchrotron radiation setup there is a cancellation of the second order relativistic kinematics effects except the non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion in the (bending) magnetic field direction. But synchrotron radiation from a bending magnets is emitted within a wide range of frequencies, and the output intencity is not sensitive on the red shift. That means that the spontaneous synchrotron radiation does not show sensitivity to the difference between covariant and non-covariant particle trajectories.
But in the 21st century with the operation XFELs this situation changes. An XFEL is an example where the first order kinematics term $(v_{\perp}/c)$ plays an essential role in the description of the XFEL radiation and, in this case, covariant coupling of fields and particles predicts an effect in complete contrast to the conventional treatment. In this chapter we present a critical reexamination of existing XFEL theory. The main emphasize of this chapter is on coherent undulator radiation from the modulated electron beam. This chapter mainly addressed to readers with limiting knowledge of accelerator and XFEL physics.
Figure 71: A well-known result of conventional (non-covariant) particle tracking. A micro-bunching electron beam passing through a weak dipole magnet (kicker) and undergoes a kick of an angle $\theta_{k}$. The propagation axis of the electron beam is deflected, while the wavefront orientation is preserved.
The usual XFEL theory based on the use of old Newtonian kinematics for particle dynamics and the Einstein’s kinematics for the electrodynamics. In fact, the usual theoretical treatment of relativistic particle dynamics involves only a corrected Newton’s second law and is based on the use Galilean edition of velocities. For rectilinear motion of the modulated electron beam, non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking. However, one of the consequences of the relativity of simultaneity (i.e. mixture of positions and time) is a difference between covariant and non-covariant kinematics of a modulated electron beam in a given magnetic field. The theory of relativity shows that discussed above difference related with the acceleration along curved trajectories.
There are several cases where the first order relativistic effect can occur in XFELs, mainly through the introduction of an trajectory kick (1). The most elementary of the effect that represents a crucial test of the correct coupling fields and particles is a problem involves the production of coherent undulator radiation by modulated ultrarelativistic electron beam kicked by a weak dipole field before entering a downstream undulator.
It would be well to begin with bird’s view of some of the main results. Let us now move on to consider the predictions of the existing XFEL theory in the case of non-collinear electron beam motion. As well-known result of conventional particle tracking states that after an electron beam is kicked by a weak dipole magnet there is a change in the trajectory of the electron beam, while the orientation of the modulated wavefront remains as before (Fig. 71). In other words, the kick results in a difference between the directions of the electron motion and the normal to the modulation wavefront (i.e. in a wavefront tilt). In existing XFEL theory the wavefront tilt is considered as real. According to conventional treatment, a transverse kick does not change the orientation of a modulation wavefront, and hence suppresses the radiation emitted in the direction of the electron motion (2).
Figure 72: A result of covariant particle tracking. In the ultrarelativistic limit, the orientation of the modulation wavefront, i.e. the orientation of plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the modulation electron beam is treated using Lorentz coordinates. The theory of relativity dictates that a modulated electron beam in the ultrarelativistic limit has the same kinematics, in Lorentz coordinates, as laser beam. According to Maxwell’s equations, the wavefront of a laser beam is always orthogonal to the propagation direction.
The covariant approach within the framework of both mechanics and electrodynamics predicts an effect in complete contrast to the conventional treatment. Namely, in the ultrarelativistic limit, the wavefront of modulation, that is a plane of simultaneity, is always perpendicular to the electron beam velocity (Fig. 72). As a result, the Maxwell’s equations predict strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction. Experiments show that this prediction is, in fact, true (3). The results of XFEL experiments demonstrated that even the direction of emission of coherent undulator radiation is beyond the predictive power of the conventional theory (4).
It is worth remarking that the absent of a dynamical explanation for the modulation wavefront readjusting in the Lorentz coordinatization has disturbed some XFEL experts. We only wish to emphasize that a good way to think of the modulation wavefront readjusting is to regard it as a result of transformation to a new time variable in the framework of the Galilean (”single frame”) electrodynamics.
16.2 Modulation Wavefront Orientation
Let us suppose that a modulated electron beam moves along the $z$-axis of a Cartesian $(x,y,z)$ system in the lab frame. As an example, suppose that the modulation wavefront is perpendicular to the velocity $v_{z}$. How to measure this orientation? A moving electron bunch changes its position with time. The natural way to do this is to answer the question: when does each electron cross the $y$-axis of the reference system? If we have adopted a method for timing distant events (i.e. a synchronization convention), we can also specify a method for measuring the orientation of the modulation wavefront: if electrons located at the position with maximum density cross the $y$-axis simultaneously at certain position $z$, then the modulation wavefront is perpendicular to $z$-axis. In other words, the modulation wavefront is defined as a plane of simultaneous events (the events being the arrival of particles located at maximum density): in short, a ”plane of simultaneity”.
Let us formulate the initial conditions in the lab frame in terms of orientation of the modulation wavefront and beam velocity. Suppose that $v_{z}$ is the velocity of the comoving frame $R(\tau)$ with respect to the lab frame $K(\tau)$ along the common $z$-axis in positive direction. In the lab frame we select a special type of coordinate system, a Lorentz coordinate system to be precise. Within a Lorentz frame (i.e. inertial frame with Lorentz coordinates), Einstein’s synchronization of distant clocks and Cartesian space coordinates $(x,y,z)$ are enforced. In order to have this, we impose that $R$ is connected to $K$ by the Lorentz boost $L(\vec{v}_{z})$, with $\vec{v}_{z}$, which transforms a given four vector event $X$ in space-time into $X_{R}=L(\vec{v_{z}})X$.
We now consider the acceleration of the beam in the lab frame up to velocity $v_{y}$ along the $y$-axis. The question arises how to assign synchronization in the lab frame after the beam acceleration. Before acceleration we picked a Lorentz coordinate system. Then, after the acceleration, the beam velocity changes of an small value $v_{y}$ along the $y$-axis. Without changing synchronization in the lab frame after the particle acceleration we have a complicated situation as concerns electrodynamics of moving charges. As a result of such boost, the transformation of time and spatial coordinates has the form of a Galilean transformation. In order to keep a Lorentz coordinate system in the lab frame after acceleration, one needs to perform a clock resynchronization by introducing the time shift $t\to t+yv_{y}/c^{2}$. This form of time transformation is justified by the fact that we are dealing with a first order approximation. Therefore, $v_{y}/c$ is so small that $v_{y}^{2}/c^{2}$ can be neglected and one arrives at the coordinate transformation $y\to y+v_{y}t$, $t\to t+yv_{y}/c^{2}$. The Lorentz transformation just described differs from a Galilean transformation just by the inclusion of the relativity of simultaneity, which is only relativistic effect that appearing in the first order in $v_{y}/c$. The relation $X_{R}=L(\vec{v_{z}})L(\vec{v_{y}})X$ presents a step-by-step change from the lab reference frame $K(\tau+d\tau)$ to $K(\tau)$ and then to the proper reference frame $R$. The shift in the time when electrons located at the position with maximum density cross the $y$-axis of the lab frame $\Delta t=yv_{y}/c^{2}$ has the effect of a rotation the modulation wavefront on the angle $v_{z}\Delta t/y=v_{z}v_{y}/c^{2}$ in the first order approximation. In ultrarelativistic limits, $v_{z}\simeq c$, and the modulation wavefront rotates exactly as the velocity vector $\vec{v}$.
What does this wavefront readjustment mean in terms of measurements? In the absolute time coordinatization the simultaneity of a pair of events has absolute character. The absolute character of the temporal coincidence of two events is a consequence of the absolute time synchronization convention. According to this old kinematics, the modulation wavefront remains unvaried. However, according to the covariant approach we establish a criterion for the simultaneity of events, which is based on the invariance of the speed of light. It is immediately understood that, as a result of the motion of electrons along the $y$ axis (i.e. along the plane of simultaneity before the boost) with the velocity $v_{y}$, the simultaneity of different events is no longer absolute, i.e. independent of the kick angle $\theta=v_{y}/c$. This reasoning is in analogy with Einstein’s train-embankment thought experiment.
The wavefront orientation has no exact objective meaning, because the relativity of simultaneity takes place. The statement that the wavefront orientation has objective meaning to within a certain accuracy can be visualized by the picture of wavefront in the proper orientation with approximate angle extension (blurring) given by $\Delta\theta\simeq v_{z}(v_{y}/c^{2})$. This relation specifies the limits within which the non relativistic theory can be applied. In fact, it follows that for a very non relativistic electron beam for which $v_{z}^{2}/c^{2}$ is very small, the angle ”blurring” becomes very small too. In this case angle of wavefront tilt $\theta=v_{y}/v_{z}$ is practically sharp $\Delta\theta/\theta\simeq v_{z}^{2}/c^{2}\ll 1$. This is a limiting case of non-relativistic kinematics. The angle ”blurring” is a peculiarity of relativistic beam motion. In the ultrarelativistic limit when $v_{z}\simeq c$, the wavefront tilt has no exact objective meaning at all since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this tilt could be ascertained.
16.3 XFEL Radiation Setup
The most elementary of the effect that represents a crucial test of the correct coupling fields and particles is a problem involves the production of coherent undulator radiation by modulated ultrarelativistic electron beam kicked by a weak dipole field before entering a downstream undulator. We want to study the process of emission of coherent undulator radiation from such setup.
The key element of a XFEL source is the udulator, which forces the electrons to move along curved periodical trajectories. There are two popular undulator configurations: helical and planar. To understand the basic principles of undulator source operation, let us consider the helical undulator (it is interesting to note that the first XFEL experiment demonstrating the apparent wavefront readjusting used helical undulator (3)).
The magnetic field on the axis of the helical undulator is given by
$\vec{B}_{w}=\vec{e}_{x}B_{w}\cos(k_{w}z)-\vec{e}_{y}B_{w}\sin(k_{w}z)~{},$ (103)
where $k_{w}=2\pi/\lambda_{w}$ is the undulator wavenumber and $\vec{e}_{x,y}$ are unit vectors directed along the $x$ and $y$ axes. We neglected the transverse variation of the magnetic field. It is necessary to mention that in XFEL engineering we deal with a very high quality of the undulator systems, which have a sufficiently wide good-field-region, so that our studies, which refer to a simple model of undulator field nevertheless yields a correct quantitative description in large variety of practical problems. The Lorentz force $\vec{F}=-e\vec{v}\times\vec{B}_{w}/c$ is used to derive the equation of motion of electrons with charge $-e$ and mass $m$ in the presence of magnetic field
$m\gamma\frac{dv_{x}}{dt}=\frac{e}{c}v_{z}B_{y}=-\frac{e}{c}v_{z}B_{w}\sin(k_{w}z)~{},$ (104)
$m\gamma\frac{dv_{y}}{dt}=-\frac{e}{c}v_{z}B_{x}=-\frac{e}{c}v_{z}B_{w}\cos(k_{w}z)~{}.$ (105)
Introducing $\tilde{v}=v_{x}+iv_{y}$, $dz=v_{z}dt$ we obtain
$m\gamma\frac{d\tilde{v}}{dz}=-i\frac{e}{c}(B_{x}+iB_{y})=-i\frac{e}{c}B_{w}\exp(-ik_{w}z)~{}.$ (106)
Integration of the latter equation gives
$\frac{\tilde{v}}{c}=\theta_{w}\exp(-k_{w}z)~{},$ (107)
where $\theta_{w}=K/\gamma$ and $K=eB_{w}/(k_{w}mc^{2})$ is the undulator parameter. The explicit expression for the electron velocity in the field of the helical undulator has the form
$\vec{v}=c\theta_{w}[\vec{e}_{x}\cos(k_{w}z)-\vec{e}_{y}\sin(k_{w}z)]~{},$ (108)
This means that the reference electron in the undulator moves along the constrained helical trajectory parallel to the $z$ axis. As a rule, the electron rotation angle $\theta_{w}$ is small and the longitudinal electron velocity $v_{z}$ is close to the velocity of light, $v_{z}=\sqrt{v^{2}-v^{2}_{\perp}}\simeq v(1-\theta_{w}^{2}/2)\simeq c$.
Let us consider a modulated ultrarelativistic electron beam moving alone the $z$ axis in the field of the helical undulator. In the present study we introduce the following assumptions. First, without kick the electrons move along constrained helical trajectories in parallel with the $z$ axis. Second, electron beam density at the undulator entrance is simply
$n=n_{0}(\vec{r}_{\perp})[1+a\cos\omega(z/v_{z}-t)]~{},$ (109)
where $a=\mathrm{const.}$ In other words we consider the case in which there are no variation in amplitude and phase of the density modulation in the transverse plane. Under these assumptions the transverse current density may be written in the form
$\vec{j}_{\perp}=-e\vec{v}_{\perp}(z)n_{0}(\vec{r}_{\perp})[1+a\cos\omega(z/v_{z}-t)]~{}.$ (110)
Even through the measured quantities are real, it is generally more convenient to use complex representation, starting with real $\vec{j}_{\perp}$, one defines the complex transverse current density:
$j_{x}+ij_{y}=-ec\theta_{w}n_{0}(\vec{r}_{\perp})\exp(-ik_{w}z)[1+a\cos\omega(z/v_{z}-t)]~{}.$ (111)
The transverse current density has an angular frequency $\omega$ and two waves traveling in the same direction with variations $\exp i(\omega z/v_{z}-k_{w}z-\omega t)$ and $\exp-i(\omega z/v_{z}+k_{w}z-\omega t)$ will add to give a total current proportional to $\exp(-ik_{w}z)[1+a\cos\omega(z/v_{z}-t)]$. The factor $\exp i(\omega z/v_{z}-k_{w}z-\omega t)$ indicates a fast wave, while the factor $\exp-i(\omega z/v_{z}+k_{w}z-\omega t)$ indicates a slow wave. The use of the word ”fast” (”slow”) here implies a wave with a phase velocity faster (slower) than the beam velocity.
Having defined the sources, we now should consider the electrodynamics problem. Maxwell equations can be manipulated mathematically in many ways in order to yield derived equations more suitable for certain applications. For example, from Maxwell equations Eq.(8) we can obtain an equation which depends only on the electric field vector $\vec{E}$ (in Gaussian units):
$c^{2}\vec{\nabla}\times(\vec{\nabla}\times{\vec{E}})=-\partial^{2}\vec{E}/\partial t^{2}-4\pi\partial\vec{j}/\partial t~{}.$ (112)
With the help of the identity
×(×$\vec{\nabla}\times(\vec{\nabla}\times{\vec{E}})=\vec{\nabla}(\vec{\nabla}\cdot{\vec{E}})-\nabla^{2}\vec{E}$ (113)
and Poisson equation
$\vec{\nabla}\cdot\vec{E}=4\pi\rho$ (114)
we obtain the inhomogeneous wave equation for $\vec{E}$
$c^{2}\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial t^{2}=4\pi c^{2}\vec{\nabla}\rho+4\pi\partial\vec{j}/\partial t~{}.$ (115)
Once the charge and current densities $\rho$ and $\vec{j}$ are specified as a function of time and position, this equation allows one to calculate the electric field $\vec{E}$ at each point of space and time. Thus, this nonhomogeneous wave equation is the complete and correct formula for radiation. However we want to apply it to still simpler circumstance in which second term (or, the current term) in the right-hand side provides the main contribution to the value of the radiation field. It is relevant to remember that our case of interest is the coherent undulator radiation and the divergence of this radiation is much smaller compared to the angle $1/\gamma$. It can be shown that when this condition is fulfilled the gradient term, $4\pi c^{2}\vec{\nabla}\rho$, in the right-hand side of the nonhomoheneous wave equation can be neglected. Thus we consider the wave equation
$c^{2}\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial t^{2}=4\pi\partial\vec{j}_{\perp}/\partial t~{}.$ (116)
We wish to examine the case when the phase velocity of the current wave is close to the velocity of light. This requirement may be met under resonance condition $\omega/c=\omega/v_{z}-k_{w}$. This is the condition for synchronism between the transverse electromagnetic wave and the fast transverse current wave with the propagation constant $\omega/v_{z}-k_{w}$. With the current wave traveling with the same phase speed as electromagnetic wave, we have the possibility of obtaining a spatial resonance between electromagnetic wave and electrons. If this the case, a cumulative interaction between modulated electron beam and transverse electromagnetic wave in empty space takes place. We are therefore justified in considering the contributions of all the waves except the synchronous one to be negligible as long as the undulator is made of a large number of periods.
Here follows an explanation of the resonance condition which is elementary in the sense that we can see what is happening physically. The field of electromagnetic wave has only transverse components, so the energy exchange between the electron and electromagnetic wave is due to transverse component of the electron velocity. For effective energy exchange between the electron and the wave, the scalar product $-e\vec{v}_{\perp}\cdot\vec{E}$ should be kept nearly constant along the whole undulator length. We see that required synchronism $k_{w}+\omega/c-\omega/v_{z}=0$ takes place when the wave advances the electron beam by the wavelength at one undulator period $\lambda_{w}/v_{z}=\lambda/(c-v_{z})$, where $\lambda=2\pi/\omega$ is the radiation wavelength. This tells us that the angle between the transverse velocity of the particle $\vec{v}_{\perp}$ and the vector of the electric field $\vec{E}$ remains nearly constant. Since $v_{z}\simeq c$ this resonance condition may be written as $\lambda\simeq\lambda_{w}/(2\gamma_{z}^{2})=\lambda_{w}(1+K^{2})/(2\gamma^{2})$.
We will use an adiabatic approximation that can be taken advantage of, in all practical situations involving XFELs, where the XFEL modulation wavelength is much shorter than the electron bunch length $\sigma_{b}$, i.e. $\sigma_{b}\omega/c\gg 1$. Since we are interested in coherent emission around the modulation wavelength the theory of coherent undulator radiation is naturally developed in the space-frequency domain. In fact, in this case one is usually interested into radiation properties at fixed modulation frequency.
We first apply a temporal Fourier transformation to the inhomogeneous wave equation to obtain the inhomogeneous Helmholtz equation
$c^{2}\nabla^{2}\vec{\bar{E}}+\omega^{2}\vec{\bar{E}}=-4\pi i\omega\vec{\bar{j}}_{\perp}~{},$ (117)
where $\vec{\bar{j}}_{\perp}(\vec{r},\omega)$ is the Fourier transform of the current density $\vec{j}_{\perp}(\vec{r},t)$. The solution can be represented as a weighted superposition of solutions corresponding to a unit point source located at $\vec{r}^{\prime}$. The Green function for the inhomogeneous Helmholtz equation is given by (for unbounded space and outgoing waves)
$4\pi G(\vec{r},\vec{r^{\prime}},\omega)=\frac{1}{|\vec{r}-\vec{r^{\prime}}|}\exp\left[i\frac{\omega}{c}|\vec{r}-\vec{r^{\prime}}|\right]~{},$ (118)
with $|\vec{r}-\vec{r^{\prime}}|=\sqrt{(x^{\prime}-x)^{2}+(y^{\prime}-y)^{2}+(z^{\prime}-z)}^{2}$. With the help of this Green function we can write a formal solution for the field equation as:
[…]
This is just a mathematical description of the concept of Huygens’ secondary sources and waves, and is of course well-known, but we still recalled how it follows directly from the Maxwell’s equations. We may consider the amplitude of the beam radiated by plane of oscillating electrons as a whole to be the resultant of radiated spherical waves. This is because Maxwell’s theory has no intrinsic anisotropy. The electrons lying on the plane of simultaneity gives rise to spherical radiated wavelets, and these combine according to Huygens’ principle to form what is effectively a radiated wave. If the plane of simultaneity is the $xy$-plane (i.e. beam modulation wavefront is perpendicular to the $z$- axis), then the Huygens’ construction shows that plane wavefronts will be emitted along the $z$-axis.
In summary: according to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the modulation wavefront. We already stressed that Maxwell’s equations are valid only in a Lorentz reference frame, i.e. when an inertial frame where the Einstein synchronization procedure is used to assign values to the time coordinates. Einstein’s time order should be applied and kept in consistent way in both dynamics and electrodynamics. Our previous description implies quite naturally that Maxwell’s equations in the lab frame are compatible only with covariant trajectories $\vec{x}_{cov}(t)$, calculated by using Lorentz coordinates and, therefore, including relativistic kinematics effects.
Let us go back to the modulated electron beam, kicked transversely with respect to the direction of motion, that was discussed before. Conventional particle tracking shows that while the electron beam direction changes after the kick, the orientation of the modulation wavefront stays unvaried. In other words, the electron motion and the wavefront normal have different directions. Therefore, according to conventional coupling of fields and particles that we deem incorrect, the coherent undulator radiation in the kicked direction produced in a downstream undulator is expected to be dramatically suppressed as soon as the kick angle is larger than the divergence of the output coherent radiation.
In order to estimate the loss in radiation efficiency in the kicked direction according to the conventional coupling of fields and particles, we make the assumption that the spatial profile of the modulation is close to that of the electron beam and has a Gaussian shape with standard deviation $\sigma$. A modulated electron beam in an undulator can be considered as a sequence of periodically spaced oscillators. The radiation produced by these oscillators always interferes coherently at zero angle with respect to the undulator axis. When all the oscillators are in phase there is, therefore, strong emission in the direction $\theta=0$. If we have a triangle with a small altitude $r\simeq\theta z$ and long base $z$, than the diagonal $s$ is longer than the base. The difference is $\Delta=s-z\simeq z\theta^{2}/2$. When $\Delta$ is equal to one wavelength, we get a minimum in the emission. This is because in this case the contributions of various oscillators are uniformly distributed in phase from 00 to $2\pi$. In the limit for a small size of the electron beam, $\sigma\to 0$, the interference will be constructive within an angle of about $\Delta\theta\backsimeq\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$, where $L_{w}=\lambda_{w}N_{w}$ is the undulator length. In the limit for a large size of the electron beam, the angle of coherence is about $\Delta\theta\backsimeq c/(\omega\sigma)$ instead. The boundary between these two asymptotes is for sizes of about $\sigma_{\mathrm{dif}}\backsimeq\sqrt{cL_{w}/\omega}$. The parameter $\omega\sigma^{2}/(cL_{w})$ can be referred to as the electron beam Fresnel number. It is worth noting that, for XFELs, the transverse size of electron beam $\sigma$ is typically much larger than $\sigma_{\mathrm{dif}}$ (i.e electron beam Fresnel number is large). Thus, we can conclude that the angular distribution of the radiation power in the far zone has a Gaussian shape with standard deviation $\sigma_{\mathrm{\theta}}\backsimeq c/(\sqrt{2}\omega\sigma)$. However, still according to the conventional treatment, after the electron beam is kicked we have the already-mentioned discrepancy between direction of the electron motion and wavefront normal. Then, the radiation intensity along the new direction of the electron beam can be approximated as $I\backsimeq I_{0}\exp[-\theta_{k}^{2}/(2\sigma_{\mathrm{\theta}}^{2})]$, where $I_{0}$ is the on-axis intensity without kick and $\theta_{k}$ is the kick angle. The exponential suppression factor is due to the tilt of the modulation wavefront with respect to the direction of motion of the electrons.
We presented a study of very idealized situation for illustrating the difference between conventional and covariant coupling of fields and particles. We solved the dynamics problem of the motion of a relativistic electrons in the prescribed force field of weak kicker magnet by working only up to the order of $\gamma\theta_{k}$. This approximation is of particular theoretical interest because it is relatively simple and at the same time forms the basis for understanding relativistic kinematic effects such as relativity of simultaneity.
Let us discuss the region of validity of our small kick angle approximation $\theta_{k}\gamma\ll 1$. Since in XFELs the Fresnel number is rather large, we can always consider a kick angle which is relatively large compared to the divergence of the output coherent radiation, and, at the same time, it is relatively small compared to the angle $1/\gamma$. In fact, from $\omega\sigma^{2}/(cL_{w})\gg 1$, with some rearranging, we obtain $\sigma_{\theta}^{2}\simeq c^{2}/(\omega^{2}\sigma^{2})\ll c/(\omega L_{w})$. Then we recall that $\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$. Therefore, the first order approximation used to investigate the kicker setup in this chapter is of practical interest in XFEL engineering.
It is one of the aims of this chapter is to demonstrate the kind of experimental predictions we are expecting from our corrected radiation theory. We worked out a very simple case in order to illustrate all the essential physical principles very clearly. Surprisingly, the first order approximation used to investigate the kicker setup in this section has also important practical applications.
Above we have shown that our correct coupling of fields and particles predicts an effect in complete contrast to the conventional treatment. Namely, in the ultrarelativistic limit, the plane of simultaneity, that is wavefront orientation of the modulation, is always perpendicular to the electron beam velocity. As a result, we predict strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction, Fig. 73.
Figure 73: Basic setup for the experimental test of XFEL theory. The correct coupling fields and particles predicts an effect in complete contrast to the conventional treatment. According to the covariant approach, in the ultrarelativistic limit, the wavefront of modulation is always perpendicular to the electron beam velocity. As a result, the Maxwell’s equations predict strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction. Experiments show that this prediction is, in fact, true.
XFEL experts actually witnessed an apparent wavefront readjusting due to the relativistic kinematics effect, but they never drew this conclusion. In this book, we are actually first in considering the idea that results of the conventional theory of radiation by relativistically moving charges are not consistent with the special relativity. In previous literature, identification of the trajectories in the source part of the usual Maxwell’s equations with the trajectories calculated by conventional particle tracking in the (”single”) lab frame has always been considered obvious.
16.4 Modulation Wavefront Tilt in Maxwell’s Theory. Logical Inconsistency
In existing literature theoretical analysis is presented, of an XFEL driven by an electron beam with wavefront tilt, and this analysis is based on the exploitation of usual Maxwell’s equations and standard simulation codes. Using only a kicker setup (i.e. without undulator radiation setup) we can demonstrate that the coupling fields and particles in the conventional XFEL theory is intrinsically incorrect.
The existing XFEL theory based on the use of the absolute time convention (i.e. old kinematics) for particle dynamics. Here we will give a simple proof of the conflict between conventional particle tracking and Maxwell’s electrodynamics. The purpose is to show how one can demonstrate in a simple way that the conventional XFEL theory is absolutely incapable of correctly describing the distribution of the electromagnetic fields from a fast moving modulated electron beam downstream the kicker.
Under the Maxwell’s electrodynamics, the fields of a modulated electron beam moving with a constant velocity exhibit an interesting behavior when the velocity of charges approaches that of light: namely, in the space-time domain they resemble more and more close of a laser beam (see Appendix IV). In fact, for a rapidly moving modulated electron beam we have nearly equal transverse and mutually perpendicular electric and magnetic fields: in the limit $v\to c$ they become indistinguishable from the fields of a laser beam, and according to Maxwell’s equations, the wavefront of the laser beam is always perpendicular to the propagation direction. This is indeed the case for virtual laser-like radiation beam in the region upstream the kicker.
Let us now consider the effect of the kick on the electron modulation wavefront. If we rely on the conventional particle tracking, the kick results in a difference between the directions of electron motion and the normal to the modulation wavefront, i.e. in a tilt of the modulation wavefront.
This is already a conflict result, because we now conclude that, according to the conventional ”single frame” approach, the direction of propagation after the kick is not perpendicular to the radiation beam wavefront. In other words, the radiation beam motion and the radiation wavefront normal have different directions. The virtual radiation beam (which is indistinguishable from a real radiation beam in the ultrarelativistic asymptote) propagates in the kicked direction with a wavefront tilt. This is what we would get for the case when our analysis is based on the conventional particle tracking, and is obviously absurd from the viewpoint of Maxwell’s electrodynamics.
In existing literature theoretical analysis is presented, of an XFEL driven by an electron beam with wavefront tilt, and this analysis is based on the exploitation of usual Maxwell’s equations and standard simulation codes. Using only a kicker setup (i.e. without undulator radiation setup) we demonstrated that the coupling fields and particles in the conventional XFEL theory is intrinsically incorrect.
The difficulty above is a part of the continual problem of XFEL physics, which started with coherent undulator radiation from an ultrarelativistic modulated electron beam in the kicked direction, and now has been focused on the wavefront tilt of the self-electromagnetic fields of the modulated electron beam.
In existing XFEL theory the wavefront tilt is considered as real. However, there is a common mistake made in accelerator physics connected with the wavefront tilt. In the ultrarelativistic domain the wavefront tilt has no exact objective meaning. The angle of wavefront tilt depends on the choice of a procedure for clock synchronization in the lab frame, as a result of which it can be given any preassigned values within the interval $(0,\theta_{k})$. For instance, in the ultrarelativistic domain, the orientation of the modulation wavefront is always perpendicular to the electron beam velocity (i.e. $\theta_{tilt}=0$) when the evolution of the modulated electron beam is treated using Lorentz coordinates. No physical effects may depends on an arbitrary constant or an arbitrary function (5).
16.5 Final Remarks
Finally, we make some remarks on another way in which our complicated problem can be solved. We know the result for Lorentz coordinatization. Now let us understand why the emission of coherent undulator radiation from the modulated electron beam in the kicked direction exist in the framework of absolute time synchronization. The physics is actually very simple, and if we treat the production of coherent undulator radiation by kicked electron beam as the aberration of light we shall see that we can understand almost everything that happens in the undulator. Our Galilean transformed electrodynamics says that by making a measurement on the coherent radiation, one can observe only radiation in the kicked direction. This is exactly analogous to the aberration (deviation of the energy transport) for light radiated from the single moving emitter in an inertial frame, Fig. 1.
There is an intuitively plausible way to understanding the aberration of light from a kicked electron beam. The aberration of light in an inertial frame can be easily explained on the basis of corpuscular theory, Fig. 48. This phenomenon is fully understandable in terms of transformation of velocities between different reference frames. In the case of the transversely moving electron beam in an undulator, intuition would seem to tell us that aberration increment would be the same. This is plausible if one keeps in mind that a (undulator) radiation pulse represents a certain amount of electromagnetic energy. Energy, like mass, is a quantity that is conserved, so that a coherent X-ray pulse resembles, in many aspects, a material particles. Therefore, we should expect that group velocities of undulator radiation pulses obey the same addition theorem for particle velocities in the inertial lab frame. A closer treatment based on wave theory of light confirm this expectation.
Now let us demonstrate that the Galilean transformed electrodynamics will give a deviation of the energy transport direction for radiated X-ray pulse. The modulated electron beam with finite aperture in an undulator is a kind of active medium which breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. Let us assume that the density of the elementary source varies in the transverse direction according to the law $\rho_{e}=g(K_{\perp})\cos{({K}_{\perp}y)}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated electron beam density. From the Galilean transformation, after partial differentiation, one obtains wave equation Eq.(12). The new terms that have to be put into the field equations due to use of Galilean transformation lead to the prediction of the Doppler effect. As one of the consequences of the Doppler effect in the absolute time coordinatization, we find an angular frequency dispersion of the light waves radiated from the kicked (modulated) electron beam with finite aperture. The Doppler shift, $\Delta\omega$, of radiated light wave (in the first order approximation) is given by $\Delta\omega=K_{\perp}v_{y}$, where $K_{\perp}$ is the transverse component of the radiated wavenumber vector. The last equation states that radiated coherent X-ray beam with finite transverse size moves along the $y$ direction with group velocity $d\omega/dk_{y}=v_{y}$. Thus, according to correct coupling conventional particle tracking and electrodynamics, we identify the direction of X-ray beam propagation with direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation.
16.6 Bibliography and Notes
1. An angular kick is often an essential part of many XFEL related diagnostic or experimental procedures. The standard gain length measurement procedure in XFELs employs such kicks. Other applications include ”beam-splitting” schemes where different polarization components are separated by means of an angular kick to the modulated electron beam [61, 62].
2. In a typical configuration for an XFEL, the orbit of a modulated electron beam is controlled to avoid large excursions from the undulator axis. All existing XFEL codes are based on a model in which the modulated electron beam moves only along the undulator axis. However, random errors in the focusing system can cause angular trajectory errors (or ”kicks”). The discrepancy between directions of the electron motion and wavefront normal after the kick have been discussed in the literature. One particular consequence that received attention following the [63] is the effect of the trajectory error (single kick error) on the XFEL amplification process. It was pointed out that coherent radiation is emitted towards the wavefront normal of the beam modulation. Thus, according to conventional coupling of fields and particles (which we claimed incorrect), the discrepancy between the two directions decreases the radiation efficiency [63]. Analysis of the trajectory errors on the XFEL amplification process showed that any XFEL undulator magnetic field must satisfy stringent requirements. However, semi-analytical studies of this critical aspect in the design of a XFEL sources are based on an incorrect coupling of fields and particles. The pleasant surprise is that the tolerances predicted are more stringent than they need be according to the corrected XFEL theory. This can be considered as one of the reason for the exceptional progress in XFEL developments over last decades.
3. The fact that our theory predicts reality in a satisfactory way is well-illustrated by comparing the prediction we just made with the results of an experiment involving ”X-ray beam splitting” of a circularly-polarized XFEL pulse from the linearly-polarized XFEL background pulse, a technique used in order to maximize the degree of circular polarization. The XFEL experiments apparently demonstrated that after a modulated electron beam is kicked on a large angle compared to the divergence of the XFEL radiation, the modulation wavefront is readjusted along the new direction of motion of the kicked beam, Fig. 73. This is the only way to justify coherent radiation emission from the short undulator placed after the kicker and along the kicked direction, see Fig. 14 in [64]. These results came unexpectedly, but from a practical standpoint, the ”apparent wavefront readjusting” immediately led to the realization that the unwanted, linearly-polarized radiation background could be fully eliminated without extra-hardware.
4. In existing literature a theoretical analysis of XFELs driven by an electron beam with wavefront tilt was presented in [65, 66], based on the use the usual Maxwell’s equations. In fact, the Maxwell solver was used as a part of the standard simulation code. We state that this approach is fundamentally incorrect. In ultrarelativistic asymptote a modulation wavefront tilt is absurd with the viewpoint of Maxwell’s electrodynamics. In the case of an XFEL we deal with an ultrarelativistic electron beam and within the Lorentz lab frame (i.e. within the validity of the Maxwell’s equations) the tilted modulation wavefront is at odds with the principle of relativity.
5. In existing XFEL theory the wavefront tilt is considered as real. Let us consider one example. One finds some papers (see e.g. [63]) which say that a wavefront tilt leads to significant degradation of the electron beam modulation in XFELs. First, suppose that modulation wavefront is perpendicular to the beam velocity. The effect of betatron oscillations, which can influence the operation of the XFEL, has its origin in an additional longitudinal velocity spread. Particles with equal energies, but with different betatron angles, have different longitudinal velocities. In other words, on top of the longitudinal velocity spread due to the energy spread, there is an additional source of velocity spread. To estimate the importance of the last effect, we should calculate the dispersion of the longitudinal velocities due to both effects. The deviation of the longitudinal velocity from nominal value is $\Delta v_{z}=v\Delta\gamma/\gamma^{3}-v\Delta\theta^{2}/2$. The finite angular spread of the electron beam results in a difference in time when each electron arrives at the same longitudinal position. This is so called normal debunching effect. From the viewpoint of the existing XFEL theory, the time difference is enhanced by the kick angle $\theta_{k}$. In this case, according to conventional (non-covariant) particle tracking, the angle of wavefront tilt is $\theta_{tilt}=\theta_{k}$. It is a widespread belief that the wavefront tilt has physical meaning, and that the deviation of the longitudinal velocity component (i.e. velocity component which is perpendicular to the modulation wavefront within the framework of Galilean kinematics) is now given by the expression $\Delta v_{z}=-v|\Delta\vec{\theta}+\vec{\theta}_{k}|^{2}/2$. If such picture is correct, the crossed term $v\vec{\theta}_{k}\cdot\Delta\vec{\theta}$ leads to a significant degradation of the modulation amplitude. This mechanism is called smearing of modulation and should be distinguished from the normal debunching [63]. Many experts would like to think that any debunching process obviously has objective meaning. The theory of relativity says, however, that normal debunching has objective meaning, but smearing effects not exist at all. The explanation of the new debunching mechanism clearly demonstrates the essential dependence of the smearing effect on the choice of the coordinate system in the four-dimensional space, which from the physical point of view is meaningless. Now let us understand physically why the new debunching mechanism does not exist in framework of Galilean kinematics. In this old kinematics the crossed term $v\vec{\theta}_{k}\cdot\Delta\vec{\theta}$ leads to a degradation of modulation amplitude in the forward direction. Our Galilean transformed electrodynamics says, however, that by making a measurement on the coherent radiation, one can observe only radiation in the kicked direction. But the crossed term is absent in the expression for the deviation of the velocity component along the kicked direction. It comes out quite naturally that the smearing effect is not a real phenomenon.
17 Kinematics of the Wigner Rotation
17.1 Introductory Remarks
The subject of this chapter is the Wigner rotation (1). We have already discussed the Wigner rotation in Chapter 7. Our earlier discussion is really about aberration of light and particles. The Wigner rotation is also relates the angular rotation velocity of the spin of an elementary particle following a curvilinear orbit. The Wigner rotation is a relativistic kinematic effect, which consists in that a coordinate axes of a reference frame, moves along a curvilinear trajectory rotating about the axes of a Lorentz lab frame. The point is that two observers with different trajectories have different 3-spaces. In other words, in the space measurement of one observer there is mixing of space and time, as seen by the other. When we studied aberration of light and particles, we begun by writing down an expression for Wigner rotation angle. To deduce this expression is the main purpose of this chapter. We have tried to keep the mathematical complexity of the discussion to a minimum.
17.2 The Commutativity of Collinear Lorentz Boosts
Lorentz transformations are essential to the further mathematical development of the Wigner rotation theory, so the next two subsections detail the usual applications together with some physical discussion.
Let us now consider a relativistic particle, accelerating in the lab frame, and let us analyze its evolution within Lorentz coordinate systems. The permanent rest frame of the particle is obviously not inertial. To get around that difficulty one introduces an infinite sequence of comoving frames. At each instant, the rest frame is a Lorentz frame centered on the particle and moving with it. As the particle velocity changes to its new value at an infinitesimally later instant, a new Lorentz frame centered on the particle and moving with it at the new velocity is used to observing the particle.
Let us denote the three inertial frames by $K,R(\tau),R(\tau+d\tau)$. The lab frame is $K$, $R(\tau)$ is the rest frame with velocity $\vec{v}=\vec{v}(\tau)$ relative to $K$, and $R(\tau+d\tau)$ is the rest frame at the next instant of proper time $\tau+d\tau$, which moves relative to $R(\tau)$ with infinitesimal velocity $d\vec{v^{\prime}}$. All inertial reference frames are assumed to be Lorentz reference frames. In order to have this, we impose that $R(\tau)$ is connected to $K$ by the Lorentz boost $L(\vec{v})$, with $\vec{v}$, which transforms a given four vector event $X$ in a space-time into $X_{R}=L(\vec{v})X$. The relation $X_{R}=L(d\vec{v^{\prime}})L(\vec{v})X$ presents a step-by-step change from $K$ to $R(\tau)$ and then to $R(\tau+d\tau)$.
There is another composition of reference-frame transformations which describes the same particle evolution in the Minkowski space-time. Let $K(\tau)$ be an inertial frame with velocity $d\vec{v}$ relative to the lab frame $K(\tau+d\tau)$. We impose that $K(\tau)$ is connected to $K(\tau+d\tau)$ by the Lorentz boost $L(d\vec{v})$. The Lorentz rest frame $R$ is supposed to move relative to the Lorentz frame $K(\tau)$ with velocity $\vec{v}$. The relation $X_{R}=L(\vec{v})L(d\vec{v})X$ presents a step-by-step change from $K(\tau+d\tau)$ to $K(\tau)$ and then to the rest frame $R$.
Let us examine the transformation of the three-velocity in the theory of relativity. For a rectilinear motion along the $z$ axis it is performed in accordance with the following equation: $v_{z}(\tau+d\tau)=(dv_{z}+v_{z})/(1+v_{z}dv_{z}/c^{2})$. Like it happens with the composition of Galilean boosts, collinear Lorentz boosts commute: $L(dv_{z})L(v_{z})=L(v_{z})L(dv_{z})$. This means that the resultant of successive collinear Lorentz boosts is independent of which transformation applies first.
17.3 The Noncommutativity of Two Lorentz Boosts in Nonparallel Directions
In contrast with the case of Lorentz boosts in collinear directions, Lorentz boosts in different directions do not commute. While the successive application of two Galilean boosts is Galilean boost and the successive application of two rotations is a rotation, the successive application of two non-collinear Lorentz boosts is not a Lorentz boost. The composition of non-collinear boosts will results to be equivalent to a boost, followed by spatial rotation, the Wigner rotation.
Let us compare the succession $K\to R(\tau)\to R(\tau+d\tau)$ with the succession $K(\tau+d\tau)\to K(\tau)\to R$ in the case when the acceleration in the rest frame is perpendicular to the line of flight of the lab frame in the rest frame. The frame $R(\tau+d\tau)$ is supposed to move relative to $R(\tau)$ with velocity $d\vec{v^{\prime}}_{x}$. Now, since we can write the result in terms of succession $L(\vec{v}_{z})L(d\vec{v}_{x})$ as well as in terms of succession $L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})$, there is a need to clarify a number of questions associated with these compositions of Lorentz frames. We can easily understand that the operational interpretation of the succession $L(\vec{v}_{z})L(d\vec{v}_{x})$ is particular simple, involving physical operation used in the measurement of the particle’s velocity increment $d\vec{v}_{x}$ in the lab frame. We should be able to understand the operational interpretation of the succession $L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})$. We begin by making an important point: the laws of physics in any one inertial reference frame should be able to account for all physical phenomena, including the observations made by moving observers. The lab observer sees the time dilation in the Lorentz frame which moves with respect to the lab frame with velocity $\vec{v}_{z}$: $dt/\gamma=d\tau$. What velocity increment $d\vec{v}_{R}$ is measured by moving observer? As viewed from the lab frame the moving observer measures the increment $d\vec{v}_{R}=\gamma d\vec{v}_{x}$.
Because of time dilation in the moving frame, the velocity increment in the lab frame $dv_{x}$ corresponds to a velocity $\gamma dv_{x}$ and $-\gamma dv_{x}$ in the frames $R(\tau)$ and $R(\tau+d\tau)$ respectively. The resulting boost compositions can be represented as $X_{R}=L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})X=L(\vec{v}_{z})L(d\vec{v}_{x})X$. In other words, Lorentz boosts in different direction do not commute: $L(\vec{v}_{z})L(d\vec{v}_{x})\neq L(d\vec{v}_{x})L(\vec{v}_{z})$.
17.4 Wigner Rotation
17.4.1 Expression for the Wigner Rotation in the Case of an Arbitrary Velocity
A large number of incorrect expressions for the Wigner rotation can be found in the literature. Therefore, there is a need to consider the Wigner rotation in detail. Rather than working out all the transformation matrices for four-vector components, we would like to show a geometrical approach that is very useful for this problem.
[…]
We derived the exact relation $d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$ using only rudimentary knowledge of special relativity. In textbooks on the theory of relativity, the spatial rotation associated with the composition of two Lorentz boosts in non-parallel directions is often introduced using the algebraic approach. This is one of the reason why authors of textbooks obtained an incorrect expression for the Wigner rotation. They describe the rotation of a moving object without geometrical meaning of such rotation and encounter serious difficulties in the interpretation of the applied calculations and of the results.
Figure 74: The interpretation of the Wigner rotation about the proper frame axes. Lab frame view of the observation of the proper observer. The lab observer is able to account for the observation the velocity rotation angle in the proper frame ($\gamma dv_{x}/v_{z}$) and the observation of the rotation angle made in the lab frame ($dv_{x}/v_{z}$). In 3D space the proper frame moves with respect to the lab frame along the line motion. Using the line motion as reference line, the lab observer sees that the observer in the proper frame measures the rotating angle of the lab frame axes with respect to proper frame axes $\delta\theta_{w}=\delta\theta(\gamma-1)$.
17.5 How to Measure a Wigner rotation?
The Wigner rotation is a relativistic kinematic effect, which consists in that a coordinate axes of a reference frame, moves along a curvalinear trajectory rotating about the axes of a Lorentz lab frame. In the Lorentz lab frame, for infinitely small transformations (due to acceleration) we obtained the formula
$d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}=\left(1-\frac{1}{\gamma}\right)\vec{\delta\theta}~{},$ (120)
where $d\vec{v}$ is the vector of small velocity change due to acceleration, $\Phi$ is the Wigner rotation angle of the spatial coordinate axes of the proper system relative to Lorentz lab frame, and $\theta$ is the orbital angle. But how to measure this orientation? A moving coordinate system changes its position in time. The question arises whether it is possible to give an experimental interpretation of the rotation of a moving coordinate system. We illustrate the problem of how to represent orientation of the moving coordinate system with a simple example.
The execution of successive transformations from $K(\tau+d\tau)$ to $K(\tau)$ at velocity $d\vec{v_{x}}$ and from from $K(\tau)$ to $R$ at velocity $\vec{v_{z}}$ equivalent to the composition of boost $L(\vec{v}+d\vec{v})$ and rotation. The Wigner rotation Eq.(120) is performed additionally to the Lorentz boost at velocity $\vec{v_{z}}+d\vec{v_{x}}$. The interpretation of this rotation about the lab frame of reference is closely associated with the length contraction.
Figure 75: The interpretation of the Wigner rotation about the lab frame. A rod directed along $x$-axis in the comoving frame. After the first boost the motion takes place along the $z$-axis. The lab frame $(x_{r},z_{r})$ coordinate system rotated with respect to the initial lab frame $(x,z)$ coordinate system on the angle $\delta\theta$. The projection of the moving rod on the $z_{r}$-axis is simply $L\delta\theta$. After the second boost at velocity $dv_{x}$ along the $x$-axis of the lab frame this projection will be contracted down to $L\delta\theta/\gamma$. It is assumed that infinitesimal angle $\delta\theta=dv_{x}/v_{z}$. According to the contracted projection the comoving frame is rotated with respect to the initial lab frame axes by the Wigner angle equal to $\delta\theta(1-1/\gamma)$.
Suppose that the lab observer after the Lorentz boost at velocity $v_{z}$ rotates the coordinate system on the angle $\vec{\delta\theta}=\vec{v_{z}}\times d\vec{v_{x}}/v_{z}^{2}$ and now $x_{r}$ locates orthogonally to the vector $\vec{v_{z}}+d\vec{v_{x}}$. Similarly, the axis $z_{r}$ is parallel to the vector $\vec{v_{z}}+d\vec{v_{x}}$. Consider a rod directed along $x$-axis in the comoving frame. The motion takes place in the plane $(x,z)$ and the rod located perpendicularly to the velocity $\vec{v_{z}}$. After the rotation of the lab frame axes, the projection of the rod on the $z_{r}$ axes will be simply $l\delta\theta$, where $l$ is the rod length in the $R$ frame and also in the lab frame after the first boost along the $z$-axis. After the second Lorentz boost at velocity $d\vec{v_{x}}$ this projection will be contracted down to $L\delta\theta/\gamma$ (Fig. 75). Let the observer in the lab frame fix the position of the axes of the comoving frame. In ultrarelativistic limit $\gamma\to\infty$ these axes will be parallel to the rotated lab frame axes $(x_{r},z_{r})$. In fact, projection of the rod on the $z_{r}$ axis will be zero. In the case of an arbitrary velocity, axes of the comoving frame are not parallel to the rotated lab frame axes $(x_{r},z_{r})$. According to contracted projection, the angle will be $-\vec{\delta\theta}/\gamma$. And one can verify directly that the axes of the comoving frame are actually rotated with respect to the initial lab frame axes $(x,z)$ by the angle equal to $\delta\theta-\delta\theta/\gamma$, which is just the Wigner rotation angle in accordance with equation Eq.(120).
Here we only wished to show how naturally Lorentz transformations lead to the Wigner rotation phenomenon. We have come to the conclusion that what are usually considered advanced parts of the theory of relativity are, in fact, quite simple. Indeed, we demonstrated that the Wigner rotation results directly from the length contraction.
17.6 Wigner Rotation and the Reciprocity
There are several points to be made about the Wigner rotation theory. We wish to remark that the expression for the Wigner rotation angle in the proper frame (as viewed from the Lorentz lab frame), $d\vec{\Phi}_{R}=(\gamma-1)\vec{v}\times d\vec{v}/v^{2}$, can be presented in the form $d\vec{\Phi}_{R}=(1-1/\gamma)\vec{v}_{R}\times d\vec{v}_{R}/v_{R}^{2}$. Thus, the Wigner rotation angle in the proper frame is expressed in terms of the lab frame velocity $\vec{v}_{R}=\vec{v}$ and its increment $d\vec{v}_{R}=\gamma d\vec{v}$ in the proper frame. We have just demonstrated that the reciprocity is the correct concept in the case of infinitesimal Wigner rotation, as it must be.
17.7 Bibliography and Notes
1. As known, a composition of noncollinear Lorentz boosts does not results in a different boost but in a Lorentz transformation involving a boost and a spatial rotation, the latter being known as Wigner rotation [67, 68]. Wigner rotation is sometimes called Thomas rotation (see e.g. [13, 69]).
2. The correct expression for the Thomas (Wigner) rotation was first obtained by V. Ritus [70]. In deriving expressions for the Thomas (Wigner) rotation, the majority of authors (see e.g. [69]) were supposedly guided by the incorrect expression for Thomas (Wigner) rotation from Moeller’s monograph [13]. The expression obtained by Moeller is given by $\vec{\delta\Phi}=(1-\gamma)\vec{v}\times d\vec{v}/v^{2}=(1-\gamma)\vec{\delta\theta}$ (and subsequently expression for Thomas precession $\Omega_{\mathrm{T}}=(1-\gamma)\omega_{0}$). It should be note that, in his monograph, Moeller stated several times that this expression valid in the lab Lorentz frame. Clearly, this expression and correct result Eq. (120) differ both in sign and in magnitude. An analysis of the reason why Moeller obtained an incorrect expression for the Thomas (Wigner) rotation in the lab frame is the focus of Ritus paper [71]. As shown in [71], the Moeller’s mistake is not computational, but conceptual in nature. In review [72] it is shown that the correct result was obtained in the works of several authors, which were published more than half century ago but remained unnoticed against the background of numerous incorrect works. The authors of some papers believe that the incorrect result for Wigner rotation in the lab frame presented in textbooks $d\vec{\Phi}=-(\gamma-1)\vec{v}\times d\vec{v}/v^{2}$ is only incorrectly interpreted with the understanding that it should be reinterpreted as a Wigner rotation of the lab frame in the proper frame. We note that such reinterpreted expression for Wigner rotation in the proper frame $d\vec{\Phi}\to d\vec{\Phi}_{R}$ is also incorrect in sign.
3. We note that in 1986, M. Stranberg obtained an expression for the Thomas (Wigner) rotation correct both in the lab inertial frame and the comoving reference frame [73]. It is noteworthy that [73] is one of the few papers that explicitly states that the angle of the Thomas (Wigner) rotation in the comoving reference frame is $\gamma$ times higher than in the lab frame.
18 Relativistic Spin Dynamics
18.1 Introductory Remarks
In 1959, a paper by Bargmann, Michel, and Telegdi was published, which dealt with the motion of elementary charged spinning particles with an anomalous magnetic moment in electromagnetic field (1). The extremely precise measurements of the magnetic-moment anomaly of the electron made on highly relativistic electrons are based on the BMT equation. The anomalous magnetic moment can be calculated by use of quantum electrodynamics. The theoretical result agrees with experiments to within a very high accuracy. This can be regarded as a direct test of BMT equation.
The existing textbooks suggest that the experimental test of the BMT equation is a direct test of what we consider the incorrect expression for Wigner rotation in the Lorentz lab frame. We claim that the inclusion of this incorrect expression as an integral part of the BMT equation in most texts is based on an incorrect physical argument. In this chapter we will investigate in detail the reason why this is the case.
18.2 Magnetic Dipole at Rest in an Electromagnetic Field
Let us consider at first the spin precession for a non relativistic charge particle. The proportionality of magnetic moment $\vec{\mu}$ and angular momentum $\vec{s}$ has been confirmed in many ”gyromagnetic” experiments on many different systems. The constant of proportionality is one of the parameters charactering a particular system. It is normally specified by giving the gyromagnetic ratio or $g$ factor, defined by $\vec{\mu}=ge\vec{s}/(2mc)$. This formula says that the magnetic moment is parallel to the angular momentum and can have any magnitude. For an electron $g$ is very nearly 2.
Suppose that a particle is at rest in an external magnetic field $\vec{B}_{R}$. The equation of motion for the angular momentum in its rest frame is $d\vec{s}/d\tau=\vec{\mu}\times\vec{B}_{R}=eg\vec{s}\times\vec{B}_{R}/(2mc)=\vec{\omega}_{s}\times\vec{s}$. In other words, the spin precesses around the direction of magnetic field with the frequency $\omega_{s}=-eg\vec{B}_{R}/(2mc)$. In the same non relativistic limit the velocity processes around the direction of $\vec{B}_{R}$ with the frequency $\omega_{p}=-(e/mc)\vec{B}_{R}$: $d\vec{v}/d\tau=(e/mc)\vec{v}\times\vec{B}_{R}$. Thus, for $g=2$ spin and velocity precess with the same frequency, so that the angle between them is conserved.
18.3 Derivation of the Covariant (BMT) Equation of Motion of Spin
Spin dynamics equations can be expressed as tensor equations in Minkowski space-time. We shall limit ourselves to the case of a particle with a magnetic moment $\vec{\mu}$ in a microscopically homogeneous electromagnetic field. Evidently the torque affects only the spin and the force affects only the momentum. It follows that the motion of the system as a whole in any frame is determined entirely by its charge, independent of magnetic dipole moment. This part of the motion has been treated in the Chapter 12. We need now only consider the spin motion.
In seeking the equation for the spin motion, we shall be guided by the known dynamics in the rest frame and the known relativistic transformation laws. We emphasize that spin is defined in a particular frame (the rest frame). Therefore, to form expressions with known transformation behavior, we need to introduce a four-quantity related to the spin. A convenient choice is a four- (pseudo)-vector $S$ defined by the requirement that in the rest frame its space-like components are the spin components, while the time-like component is zero. We shall call $S$ four-spin; when normalized by dividing by its invariant length, it will be called polarization four-vector. It is space-like, and therefore in no frame does it space-like part vanish.
Let the spin of the particle be represented in the rest frame by $\vec{s}$. The four-vector $S^{\alpha}$ is by definition required to be purely spatial at time $\tau$ in an instantaneous Lorentz rest frame $R(\tau)$ of the particle and to coincide at this time with the spin $\vec{s}(\tau)$ of the particle; that is $S_{R}^{\alpha}(\tau)=(0,\vec{S}_{R}(\tau))=(0,\vec{s}(\tau))$. At a later instant $\tau+\Delta\tau$ in an instantaneous inertial rest frame $R(\tau+\Delta\tau)$, we have similarly $S_{R}^{\alpha}(\tau+\Delta\tau)=(0,\vec{S}_{R}(\tau+\Delta\tau))=(0,\vec{s}(\tau+\Delta\tau))$.
The BMT equation is manifestly covariant equation of motion for a four-vector spin $S^{\alpha}$ in an electromagnetic field $F^{\alpha\beta}$:
$\frac{dS^{\alpha}}{d\tau}=\frac{ge}{2mc}\left[F^{\alpha\beta}S_{\beta}+\frac{1}{c^{2}}u^{\alpha}\left(S_{\lambda}F^{\lambda\mu}u_{\mu}\right)\right]-\frac{1}{c^{2}}u^{\alpha}\left(S_{\lambda}\frac{du^{\lambda}}{d\tau}\right)~{},$ (121)
where $u_{\mu}=dx_{\mu}/d\tau$ is the four-dimensional particle velocity vector. With Eq.(20), one has (1)
$\frac{dS^{\alpha}}{d\tau}=\frac{e}{mc}\left[\frac{g}{2}F^{\alpha\beta}S_{\beta}+\frac{g-2}{2c^{2}}u^{\alpha}\left(S_{\lambda}F^{\lambda\mu}u_{\mu}\right)\right]~{},$ (122)
The BMT equation is valid for any given inertial frame, and consistently describes, together with the covariant-force law, the motion of a charged particle with spin and magnetic moment. If $F^{\mu\nu}\neq 0$, even with $g=0$, we see that $dS^{\mu}/d\tau\neq 0$. Thus, a spinning charged particle will precess in an electromagnetic field even if it has no magnetic moment. This precession is pure relativistic effect.
The covariant equation of spin motion for a relativistic particle under the action of the four-force $Q^{\mu}=eF^{\mu\nu}u_{\nu}$ in the Lorentz lab frame, Eq.(121), is a relativistic ”generalization” of the equation of motion for a particle angular momentum in its rest frame. Relativistic ”generalization” means that the three independent equations expressing the Larmor spin precession are be embedded into the four-dimensional Minkowski space-time. The idea of embedding is based on the principle of relativity i.e. on the fact that the classical equatuion of motion for particle angular momentum $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$ can always be used in any Lorentz frame where the particle, whose motion we want to describe, is at rest. In other words, if an instantaneously comoving Lorentz frame is given at some instant, one can precisely predict the evolution of the particle spin in this frame during an infinitesimal time interval.
In Lorentz coordinates there is a kinematics constraint $S^{\mu}u_{\mu}=0$, which is orthogonality condition of four-spin and four-velocity. Because of this constraint, the four-dimensional dynamics law, Eq.(121), actually includes only three independent equations of motion. Using the explicit expression for Lorentz force we find that the four equations Eq.(121) automatically imply the constraint $S^{\mu}u_{\mu}=0$ as it must be. To prove this we may point out that one has in every Lorentz frame $S^{0}=\vec{S}\cdot\vec{v}$. While $S^{0}$ vanishes in the rest frame, $dS^{0}/d\tau$ need not. In fact $d(S^{\mu}u_{\mu})/d\tau=0$ implies $dS^{0}/d\tau=\vec{S}\cdot d\vec{v}/d\tau$. The immediate generalization of $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$ and $dS^{0}/d\tau=\vec{S}\cdot d\vec{v}/d\tau$ to arbitrary Lorentz frames is Eq.(121) as can be checked by reducing to the rest frame. A methodological analogy with the relativistic generalization of the Newton’s second law emerges.
In order to fully understand the meaning of embedding of the spin dynamics law in the Minkowski space-time, one must keep in mind that, above, we characterized the spin dynamics equation in the Lorentz comoving frame as a phenomenological law. The microscopic interpretation of the magnetic moment of a particle is not given. In other words, it is generally accepted that the spin dynamics law is a phenomenological law and the magnetic moment is introduced in an ad hoc manner. The system of coordinates in which the classical equations of motion for particle angular momentum are valid can be defined as Lorentz rest frame. The relativistic generalization of the three-dimensional equation $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$ to any Lorentz frame permits us to make correct predictions.
18.4 Change Spin Variables
When Bargman, Michel and Telegdi first discovered the correct laws of spin dynamics, they wrote a manifestly covariant equation in Minkowski space-time, Eq.(122), which describes the motion of the four spin $S^{\mu}$. The derivation of this equation was very similar to the four-tensor equations that were already known to relativistic particle dynamics. The equation Eq.(122) is more complex than one might think. In fact, it is composed by a set of coupled differential equations. To find solution directly from the system seems quite difficult, even for a very symmetric, uniform magnetic field setup. How to solve this four-tensor equation is an interesting question. In relativistic spin dynamics it is done in one particular way, which is very convenient.
In order to apply Eq.(122) to specific problems it is convenient to introduce a three vector $\vec{s}~{}$ by the equation $\vec{s}=\vec{S}+S^{0}\vec{p}c/(\mathcal{E}+mc^{2})$. With the help of this relation one can work out the equation of motion for $\vec{s}~{}$. In the important case of a uniform magnetic field with no electric field in the lab frame one has, after a somewhat lengthly calculations:
$\frac{d\vec{s}}{d\tau}=-\frac{e}{2m}\left[\left(g-2+\frac{2}{\gamma}\right)\gamma\vec{B}-\frac{(g-2)\gamma}{\gamma+1}\frac{\vec{v}}{c}\left(\frac{\vec{v}}{c}\cdot\gamma\vec{B}\right)\right]\times\vec{s}~{},$ (123)
What must be recognized is that in the accepted covariant approach (indeed, Eq.(122) is obviously manifestly covariant), the solution of the dynamics problem for the spin in the lab frame makes no reference to the three-dimensional velocity. In fact, the Eq.(123) includes relativistic factor $\gamma$ and vector $\vec{v}/c$, which are actually notations: $\gamma=\mathcal{E}/(mc^{2})$, $\vec{v}/c=\vec{p}c/\mathcal{E}$. All quantities $\mathcal{E},\vec{p},\vec{B}$ are measured in the lab frame and have exact objective meaning i.e. they are convention-independent. The evolution parameter $\tau$ is also measured in the lab frame and has exact objective meaning . For instance, it is not hard to demonstrate that $d\tau=mdl/|\vec{p}|$, where $dl$ is the differential of the path length.
Spin vector $\vec{s}~{}$ is not part of a four-vector, and depends on both $\vec{S}$ and $S^{0}$. While not being a four-vector, it is effectively a three-dimensional object (having zero time component in the inertial frame in question) and the spatial part of this object undergoes pure rotation with constant rate for the example of motion along a circle in special relativity. If we perform an arbitrary velocity mapping, $\vec{s}$ will have to be recomputed from the transformed values $S^{\mu}$ and $p^{\mu}$. However, this new $\vec{s}$ will satisfy an equation of the form Eq.(123), with $\vec{B}$ computed from the transformed $F^{\mu\nu}$.
Let us restrict our treatment of spinning particle dynamics to purely transverse magnetic fields. This means that the magnetic field vector $\vec{B}$ is oriented normal to the particle line motion. If the field is transverse, then equation Eq.(123) is reduced to
$\frac{d\vec{s}}{d\tau}=\vec{\Omega}\times\vec{s}=-\frac{e}{2mc}\left[\left(g-2+\frac{2}{\gamma}\right)\gamma\vec{B}\right]\times\vec{s}~{},$ (124)
Now we have an equation in the most convenient form to be solved. Suppose we let the charged spinning particle in the lab frame through a bending magnet with the length $dl$. We know that $d\theta=-eBdl/(|\vec{p}|c)$ is the orbital angle of the particle in the lab frame. Note that $d\tau=mdl/|\vec{p}|$. Then, Eq.(124) tells us that we may write the spin rotation angle with respect to the lab frame axes $\Omega d\tau$ as $\Omega d\tau=[(g/2-1)\gamma d\theta+d\theta]$. This tell us that in the lab frame the spin of a particle $\vec{s}$ changes the angle $\phi$ with its line motion. The rate of change of the angle $\phi$ with the orbital angle is $(g/2-1)\gamma$, so we can write $d\phi=(g/2-1)\gamma d\theta$.
We would like to discuss the following question: Is the vector $\vec{s}$ merely a device which is useful in making calculations - or is it a real quantity ( i.e. a quantity which has direct physical meaning)? The starting point of Bargman, Michel and Telegdi was the particle rest frame and the equation of motion for particle angular momentum, which they generalized to the Lorentz lab frame and then transformed back to the rest frame, $\vec{s}=\vec{S}+S^{0}\vec{p}c/(\mathcal{E}+mc^{2})$. The spin vector $\vec{s}$ directly gives the spin as perceived in a comoving system. If we say that in the lab frame the spin of a particle makes the angle $\phi$ with its velocity, we mean that in the particle’s rest frame the spin makes this angle with the line motion of the lab frame.
This brings up an interesting question: Why it is convenient to transform equation Eq.(122) to the rest frame as of that instant? This is understandable. The approach in which we deal with the proper spin is much preferred in the experimental practice due to clear physical meaning of the spin vector $\vec{s}$. Unlike (3D) momentum vector $\vec{p}$, which has definite components in each reference frame, (3D) angular momentum vector $\vec{s}$ is defined only in one particular reference frame. It does not transform. Any statement about it refers to the rest frame as of that instant.
We must conclude that the conventional method used to explain the spin dynamics in the lab frame is very unusual. Indeed, the spin rotation measurement in the lab frame is interpreted with viewpoint of the proper observer as viewing this of the lab observer.
18.5 An Alternative Approach to the BMT Theory
18.5.1 BMT Equation Transformed to the Rest Frame
We want to emphasize that the equation Eq.(124) for the proper spin $\vec{s}$ and the BMT equation Eq.(121) for the four spin $S^{\mu}$ are completely equivalent, they both determine the behaviour of the spin from the point of view of the lab frame. With Eq.(124) have what we need to know - the evolution of the proper spin vector $\vec{s}$ with respect to the lab frame axes. Starting from the classical equation $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$, which describes the Larmor precession with respect to the proper frame axes, we have derived the equation Eq.(123), which describes the spin motion with respect to the lab frame axes in the proper frame and reduced to Eq.(124) in the case of purely transverse magnetic fields. That means that we know the orientation of the proper spin with respect to the lab coordinate system which is moving with velocity $-\vec{v}$ and acceleration $-\gamma d\vec{v}/d\tau$ in the proper frame.
Above we described the BMT equation, Eq.(124), in the standard manner. It uses a spin quantity defined in the proper frame but observed with respect to the lab frame axes. Let’s look at what the equation Eq.(124) says in a little more detail. It will be more convenient if we rewrite this equation as
$d\vec{s}=\vec{\Omega}d\tau\times\vec{s}=-eg\gamma\vec{B}d\tau/(2mc)\times\vec{s}+e(\gamma-1)\vec{B}d\tau/(mc)\times\vec{s}~{}.$ (125)
18.5.2 Relativistic Kinematic Addition to the Larmor Rotation
Now let’s see how we can write the right-hand side of Eq.(125) . The first term is that we would expect for the spin rotation due to a torque with respect to the proper frame axes $d\vec{\phi}_{L}=-eg\gamma\vec{B}d\tau/(2mc)=(g/2)\gamma d\vec{\theta}$. Here $d\vec{\theta}=-eBdl/(|\vec{p}|c)$ is the angle of the velocity rotation in the lab frame. It has also been made evident by our analysis in the previous Chapter 16 that angle of rotation $d\vec{\phi}_{W}=-e(\gamma-1)\vec{B}d\tau/(mc)=(\gamma-1)d\vec{\theta}$ corresponds to the Wigner rotation of the lab frame axes with respect to the proper frame axes. With this definitions, we have
$d\vec{s}=\vec{\Omega}d\tau\times\vec{s}=d\vec{\phi}_{L}\times\vec{s}-d\vec{\phi}_{W}\times\vec{s}~{},$ (126)
which begins to look interesting.
18.5.3 Wigner Rotation in the Proper Frame. First Practical Application
Now we introduce our new approach to the BMT theory, finding another way in which our complicated problem can be solved. We know that $d\vec{\phi}_{L}$ and $d\vec{\phi}_{W}$ are the rotations with respect to the proper frame axes. Actually we only need to find the spin motion with respect to the lab frame axes. Now we must be careful about signs of rotations.
There is a good mnemonic rule to learn the signs of different rotations. The rule says, first, that the direction of the velocity rotation in the proper frame is the same as the direction of the velocity rotation in the lab frame. Second, the direction of the lab frame rotation in the proper frame is the same as the direction of the velocity rotation in the proper frame. Third, the sign of the spin rotation due to a torque at $g>0$ (it is handy to remember that for an electron $g$ is positive and very nearly 2) means that the direction of the rotation in the proper frame is the same as the direction of the velocity rotation in the proper frame.
We now ask about the proper spin rotation with respect to the lab frame axes. This is easy to find. The relative rotation angle is $d\vec{\phi}_{L}-d\vec{\phi}_{W}$. So we begin to understand the basic machinery behind spin dynamics. We see why the Wigner rotation of the lab frame axes in the proper frame must be taken into account if we need to know the proper spin dynamics with respect to the lab frame axes.
Why the new derivation of the BMT equation is so simple? The reason is that the splitting of the particle spin motion with respect to the lab frame axes into the dynamic and kinematic parts can only be realized in the proper frame. In the proper frame, we do not need to know any more about a relativistic ”generalization” of the (phenomenological) classical equation of motion for the particle angular momentum. In this case, it is possible to separate the spin dynamics problem into the trivial (Larmor) dynamic problem and into the kinematic problem of Wigner rotation of the lab frame axes in the proper frame.
18.6 Spin Tracking
Having written down the spin motion equation in a 4-vector form, Eq.(122), and determined the components of the 4-force, we satisfied the principle of relativity for one thing, and, for another, we obtained the four components of the equation for the spin motion. This is a covariant relativistic generalization of the usual three dimensional equation of magnetic moment motion, which is based on the particle proper time as the evolution parameter. We next wish to describe the spin motion with respect to the Lorentz lab frame using the lab time $t$ as the evolution parameter.
18.6.1 Conventional spin tracking. Hidden absolute time coordinatization
When going from the proper time $\tau$ to the lab time $t$, the frequency of spin precession with respect to the lab frame can be obtained using the well-known formula $d\tau=dt/\gamma$. We then find
$\frac{d\vec{s}}{dt}=\vec{\varpi}\times\vec{s}=-\frac{e}{2mc}\left[\left(g-2+\frac{2}{\gamma}\right)\vec{B}\right]\times\vec{s}~{}.$ (127)
The frequency of spin precession can be written in the form
$\varpi=\omega_{0}[1+\gamma(g/2-1)]~{},$ (128)
where $\omega_{0}$ is the particle revolution frequency. Now the time-like part of the four-velocity is decomposed to $c\gamma=c/\sqrt{1-v^{2}/c^{2}}$ and the trajectory does not include relativistic kinematics effects. In particular, the Galilean vectorial law of addition of velocities is actually used. So we must have made a jump to the absolute time coordinatization.
The previous commonly accepted derivation of the equations for the spin precession in the lab frame from the covariant equation Eq.(121) has the same delicate point as the derivation of the equation of particle motion from the covariant equation Eq.(20). The four-velocity cannot be decomposed into $u=(c\gamma,\vec{v}\gamma)$ when we deal with a particle accelerating along a curved trajectory in the Lorentz lab frame. One of the consequences of non-commutativity of non-collinear Lorentz boosts is the unusual momentum-velocity relation. In this case there is a difference between covariant and non-covariant particle trajectories.
The old kinematics comes from the relation $d\tau=dt/\gamma$. The presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization that provides the time coordinate $t$ in the lab frame is based on the use of the absolute time convention.
18.6.2 Convention-Invariant Spin Tracking
In the Chapter 12 we saw that the particle path $\vec{x}(l)$ has an exact objective meaning i.e. it is convention-invariant. The spin orientation $\vec{s}$ at each point of the particle path $\vec{x}(l)$ has also exact objective meaning. In contrast to this, and consistently with the conventionality of the three-velocity, the function $\vec{s}(t)$ describing the spinning particle in the lab frame has no exact objective meaning.
We now want to describe how to determine the spin orientation along the path $\vec{s}(l)$ in covariant spin tracking. Using the covariant equation Eq.(121) we obtain Eq.(124). If we use the relation $d\tau=mdl/|\vec{p}|$ our convention-invariant equation of spin motion reads
$\frac{d\vec{s}}{dl}=-\frac{e\mathcal{E}}{m|\vec{p}|c^{3}}\left[\left(\frac{g}{2}-1+\frac{mc^{2}}{\mathcal{E}}\right)\vec{B}\right]\times\vec{s}=\left[\left(\frac{g}{2}-1\right)\frac{\mathcal{E}}{mc^{2}}+1\right]\frac{d\vec{\theta}}{dl}\times\vec{s}~{},$ (129)
which is based on the path length $l$ as the evolution parameter. These three equations corresponds exactly to the equations for components of the proper spin vector that can be found from the non-covariant spin tracking equation Eq.(127). We want to emphasize that there are two different (covariant and non covariant) approaches that produce the same spin orientation $\vec{s}(l)$ along the path. The point is that both approaches describe correctly the same physical reality and the orientation of the proper spin $\vec{s}$ at any point of particle path in the magnetic field has obviously an objective meaning, i.e. is convention-invariant.
Now we take an example, so it can be seen that we do not need to ask questions about the function $\vec{s}(t)$ of a spinning particle experimentally. Just think of experiments related with accelerator physics. Suppose we want to calibrate the beam energy in a storage ring based on measurement of spin precession frequency of polarized electrons. To measure the precession frequency $\varpi$, a method of beam resonance depolarization by an oscillating electromagnetic field can be used (2). There are many forms of depolarizers, but we will mention just one, which especially simple. It is a depolarizer whose operation depends on the radio-frequency longitudinal magnetic field which is produced by a current-curring loop around a ceramic section of the vacuum chamber.
Suppose the observer in the lab frame performs the beam energy measurement. We should examine what parts of the measured data depends on the choice of synchronization convention and what parts do not. Clearly, physically meaningful results must be convention-invariant. One might think that this is a typical time-depending measurement of function $\vec{s}(t)$. However, we state that the precession frequency $\varpi$ has no intrinsic meaning - its meaning is only being assigned by a convention. It is not possible to determine the precession frequency $\varpi$ uniquely, because there is always some arbitrariness in the $\vec{s}(t)$. For instance, it is always possible to make an arbitrary change in the rhythm of the clocks (i.e. scale of the time). But our problem is to determine the energy for an electron beam. So one needs to measure also the revolution frequency $\omega_{0}$ by using the same space-time grid. What this all means physically is very interesting. The ratio $\varpi/\omega_{0}$ is convention independent i.e it does not depend on the distant clocks synchronization or on the rhythm of the clocks. It means, for example, that if we observe the dimensionless frequency $\varpi/\omega_{0}$, we can find out the value of the convention-invariant beam energy $\mathcal{E}$. The $(g/2-1)$ factor can be calculated by use of quantum electrodynamics.
[…]
18.7 Spin Rotation in the Limit $g\to 0$
It is generally accepted that spin dynamics law is a phenomenological law and that the magnetic moment is introduced in an ad hoc manner. Let us consider the special case with $g\to 0$. The BMT equation for a particle with small $g$ factor is
$\frac{dS^{\alpha}}{d\tau}=-\frac{1}{c^{2}}u^{\alpha}\left(S_{\lambda}\frac{du^{\lambda}}{d\tau}\right)=-\frac{e}{mc^{3}}u^{\alpha}\left(S_{\lambda}F^{\lambda\mu}u_{\mu}\right)~{}.$ (130)
It is often more convenient to write this equation as the equation of motion for $\vec{s}$. If the field is transverse, then the equation Eq.(130) is reduced to
$\frac{d\vec{s}}{d\tau}=\left[\left(\frac{\mathcal{E}}{mc^{2}}-1\right)\frac{e}{mc}\vec{B}\right]\times\vec{s}~{},$ (131)
Note that the equation Eq.(131) for the proper spin $\vec{s}$ and the BMT equation Eq.(130) for four spin $S^{\mu}$ are completely equivalent. Eq.(131) is the result of transformation to new spin variables.
It will be more convenient if we rewrite this equation as
$d\vec{s}=-(\gamma-1)d\vec{\theta}\times\vec{s}~{}.$ (132)
Here $d\vec{\theta}$ is the angle of the velocity rotation in the lab frame, $(\gamma-1)d\vec{\theta}$ is the Wigner rotation of the lab frame axes with respect to the proper frame axes. BMT equation uses a spin quantity defined in the proper frame but observed with respect to the lab frame axes.
Conventional spin tracking treats the space-time continuum in a non relativistic format, as a (3+1) manifold. In the conventional spin tracking, we assign absolute time coordinate and we have no mixture of positions and time. This approach to relativistic spin dynamics relies on the use of three equations for the spin motion
$\frac{d\vec{s}}{dt}=\left[\left(1-\frac{1}{\gamma}\right)\frac{e}{mc}\vec{B}\right]\times\vec{s}=-\left[\left(\gamma-1\right)\vec{\omega_{0}}\right]\times\vec{s}~{},$ (133)
which are based on the use of the absolute time $t$ as the evolution parameter. Here, $\vec{\omega_{0}}=-e\vec{B}/(mc\gamma)$ is the particle angular frequency in the lab frame. Now the time-like part of the four-velocity is decomposed to $c\gamma=c/\sqrt{1-v^{2}/c^{2}}$. This decomposition is a manifestation of the absolute time convention.
It should be note that trick with $dt$ in Eq. (133) is only technical; this equation includes the differential $dt$ in the left and right parts. We can eliminate $dt$ from Eq.(133). If we make this elimination we find Eq.(132).
There are two different (covariant and non covariant) approaches that produce the same spin orientation $\vec{s}(l)$ along the path. Using the Eq.(131) or Eq.(133) we obtain
$\frac{d\vec{s}}{dl}=-\left[\frac{\mathcal{E}}{mc^{2}}-1\right]\frac{d\vec{\theta}}{dl}\times\vec{s}~{},$ (134)
Both approaches describe correctly the same physical reality, and the rotation of the proper spin $\vec{s}$ with respect to the lab frame axes at $g\to 0$ is convention-invariant.
The relativistic kinematic effects such as Wigner rotation, Lorentz-Fitzgerald contraction, time dilation and relativistic corrections to the law of composition of velocities are coordinate (i.e. convention-dependent) effects and have no exact objective meaning. In the case of the Lorentz coordinatization, one will experience e.g. the Wigner rotation phenomenon. In contrast to this, in the case of absolute time coordinatization there are no relativistic kinematics effects, and no Wigner rotation will be found.
However, the spin orientation at each point of the particle path has exact objective meaning. In fact, Eq.(134) is convention-invariant i.e includes only quantities which have exact objective meaning. Understanding this result of the theory of relativity is similar to understanding the previously discussed result for relativistic mass correction. We find that the evolution of a particle along its path is still given by the corrected Newton’s second law even though the relativistic mass correction has no kinematical origin. A methodological analogy with the spin dynamics equation Eq.(134) emerge by itself. The spin rotation in the lab frame at $g\to 0$ is relativistic effect (as the relativistic mass correction) but it has no kinematical origin.
18.7.1 Proper frame view of observations of the lab observer
Now we wish to continue in our analysis a little further. We will look for a different way of calculating the spin rotation. We found earlier that the easiest way to derive the BMT equation is to use the Lorentz proper frame. The Wigner rotation of the lab frame axes in the proper frame must be taken into account if we need to know the proper spin dynamics with respect to the lab frame axes. In contrast, in the case of the absolute time coordinatization in the proper frame, there is no kinematic Wigner rotation of the lab frame axes with respect to the proper frame axes. The two approaches give, in fact, a different result for spin rotation with respect to the lab frame axes, which must be, however, convention-invariant. This glaring conflict between results of covariant and non covariant approaches in the proper frame can actually explained. We will see that it is a dynamical line of arguments that explains this paradoxical situation with the relativistic spin rotation.
Suppose that an observer in the lab frame performs a spin rotation measurement. To measure the spin direction, a polarimeter at rest in the lab frame can be used. Suppose we put a charged spinning particle with small $g$ factor through a bending magnet. The lab observer can directly measure the angle of spin rotation at the magnet exit using the polarimeter. The result he observes is consistent with the spin dynamics equation Eq.(134). How does it happen that the construction of the polarimeter never came into discussion before? In the lab frame the polarimeter is at rest and the field theory involved in the polarimeter operation description is isotropic. We do not need to know any more about the polarimeter operation. In this sense we can discuss in the lab frame about the spin orientation with respect to the lab frame axes and any detail about the polarimeter is not needed. The proper observer sees that polarimeter is moving with a given acceleration and the lab observer, moving with the polarimeter, performs the spin direction measurement. Then, when the polarimeter measurement is analyzed, the proper observer finds that the measured spin rotation angle is nonzero and consistent with the spin dynamics equation Eq.(134), as must be.
How shell we describe the polarimeter operation after Galilean transformations? Suppose that the operation of the polarimeter depends on the electromagnetic field. After the Galilean transformations of the field equations we obtain the complicated anisotropic electrodynamics equations. The new terms that have to be put into the field equations due to the use of Galilean transformations lead to the same prediction as concerns experimental results: the spin of the particle is rotated with respect to the lab frame axes according to Eq.(134).
In order to predict the result of the moving polarimeter measurement one does not need to have access to the detailed dynamics of the particle into the polarimeter. It is enough to assume Lorentz covariance of the field theory involved in the description of the polarimeter operation. As before, we use a mathematical trick for solving the electromagnetic field equations with anisotropic terms: in order to eliminate these terms in the transformed field equations, we make a change of the variables. Using new variables we obtain the phenomenon of spin rotation with respect to the lab frame axes.
18.8 Incorrect Expression for Wigner Rotation. Myth About Experimental Test
18.8.1 Terminology. Thomas Precession: Correct and Incorrect Solutions
The expression for the Wigner rotation in the lab frame obtained by authors of textbooks is given by $\vec{\delta\Phi}=(1-\gamma)\vec{v}\times d\vec{v}/v^{2}=(1-\gamma)\vec{\delta\theta}$, which often presented as $\vec{\omega}_{Th}=d\vec{\Phi}/dt=(1-\gamma)[\vec{v}\times d\vec{v}/dt]/v^{2}$. In other words, the proper frame coordinate performs a precession relative to the lab frame with the velocity of precession $\vec{\omega}_{Th}$, where $d\vec{v}/dt$ is the acceleration of the spinning particle in the lab frame. This precession phenomenon is called Thomas precession. From the viewpoint of the generally accepted terminology, Thomas precession is actually a manifestation of the Wigner (Thomas) rotation. According to expression for Thomas precession in the lab frame presented in textbooks, the comoving frame precesses in the opposite direction with respect to the direction of the angular velocity of the precession $\vec{\omega}_{0}=\vec{v}\times d\vec{v}/dt/v^{2}$ and $\omega_{Th}\to-\infty$ in the limit $\gamma\longrightarrow\infty$. The theory of relativity shows us that the textbook expression for the Thomas precession in the lab frame and correct result $\vec{\omega}_{Th}=(1-1/\gamma)[\vec{v}\times d\vec{v}/dt]/v^{2}$ actually differ both in sign and magnitude.
18.8.2 Incorrect Interpretation of the Correct BMT Result
The existence of the usual incorrect expression for the Thomas precession in the lab frame has led to incorrect interpretation of the BMT result and, in particular, of the spin dynamics equation Eq.(133). Using the incorrect result for the Thomas precession, the BMT result for a small $g$ factor, Eq.(133), is usually presented as
$\frac{d\vec{s}}{dt}=-\left[\left(\gamma-1\right)\vec{\omega_{0}}\right]\times\vec{s}=\vec{\omega}_{Th}\times\vec{s}~{},$ (135)
Frequently, the first stumbling blocks in the process of understanding and accepting the correct Wigner (Thomas) rotation theory is a widespread belief that the experimental test of the BMT equation is a direct test of the incorrect expression for Thomas precession(3).
The interpretation of Eq.(133) as the Thomas precession Eq.(135) is presented in textbooks as alternative approach to the already developed BMT theory. Authors of textbooks got the correct BMT result by using the incorrect expression for the Thomas precession and an incorrect physical argument.
According to textbooks, equation Eq. (135) gives the rotation of an elementary particle spin in the lab frame $d\vec{s}=-(\gamma-1)d\vec{\theta}\times\vec{s}$. In other words, the spin rotation is recorded in the lab frame. In contrast, equation Eq. (132) $d\vec{s}=-(\gamma-1)d\vec{\theta}\times\vec{s}$ gives the rotation of the spin vector $\vec{s}$ with respect to the lab frame axes in the comoving frame. We emphasize once again: according to the BMT theory, the spin rotation is recorded in the comoving frame (as viewing this of the lab observer). The spin rotation measurement in the lab frame is interpreted with viewpoint of the proper observer due to clear physical meaning of the proper spin vector. Angular momentum vector $\vec{s}$ is defined only in one particular ( proper) reference frame, and any statement about it refers to the rest frame as of that instant.
18.9 Bibliography and Notes
1. The motion of the classical spin in an external electromagnetic field is presented by the Bargmann-Michel-Telegdi (BMT) equation [74]. The BMT equation is manifestly covariant and can be used in any inertial frame. It is the law of motion of the four-spin for a particle in a uniform electric and magnetic fields.
2. A method for measuring the particle energy in an electron-positron storage ring by means of resonance depolarization by a radio-frequency longitudinal magnetic field is described in [75].
3. The results in the Bargmann-Michel-Telegdi paper [74] were obtained by the method of semi-classical approximation of the Dirac equation. The Wigner rotation was not considered in [74] at all, because the Dirac equation allow obtaining the solution for the total particle’s spin motion without an explicit splitting it into the Larmor and Wigner parts.
Appendix I. Radiation by Moving Charges
We start with the solution of Maxwell’s equation in the space-time domain, the well-known Lienard-Wiechert expression, and we subsequently apply a Fourier transformation. The Lienard-Wiechert expression for the electric field of a point charge $(-e)$ reads (see, e.g. [69]):
$\vec{{E}}(\vec{r}_{o},t)$ $=$ $-e{\vec{n}-\vec{\beta}\over{\gamma^{2}(1-\vec{n}\cdot\vec{\beta})^{3}|\vec{r}_{o}-\vec{r^{\prime}}|^{2}}}-{e\over{c}}{\vec{n}\times[(\vec{n}-\vec{\beta})\times\dot{\vec{\beta}}]\over{(1-\vec{n}\cdot\vec{\beta})^{3}|\vec{r}_{o}-\vec{r^{\prime}}|}}.$ (136)
$R=|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|$ denotes the displacement vector from the retarded position of the charge to the point where the fields are calculated. Moreover, $\vec{\beta}=\vec{v}/c$, $\vec{\dot{\beta}}=\vec{\dot{v}}/c$, while $\vec{v}$ and $\vec{\dot{v}}$ denote the retarded velocity and acceleration of the electron. Finally, the observation time $t$ is linked with the retarded time $t^{\prime}$ by the retardation condition $R=c(t-t^{\prime})$. As is well-known, Eq. (136) serves as a basis for the decomposition of the electric field into a sum of two quantities. The first term on the right-hand side of Eq. (136) is independent of acceleration, while the second term linearly depends on it. For this reason, the first term is called ”velocity field”, and the second ”acceleration field” [69]. The velocity field differs from the acceleration field in several respects, one of which is the behavior in the limit for a very large distance from the electron. There one finds that the velocity field decreases like $R^{-2}$, while the acceleration field only decreases as $R^{-1}$. Let us apply a Fourier transformation:
[…]
As in Eq. (136) one may formally recognize a velocity and an acceleration term in Eq. (138) as well. Since Eq. (138) follows directly from Eq. (136), that is valid in the time domain, the magnitude of the velocity and acceleration parts in Eq. (138), that include terms in $1/R^{2}$ and $1/R$ respectively, do not depend on the wavelength $\lambda$. It is instructive to take advantage of integration by parts. With the help of
${1\over{c}}{d\over{dt^{\prime}}}|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|=-\vec{n}\cdot\vec{\beta}~{}~{}~{}~{}~{}\mathrm{and}~{}~{}~{}~{}{d\vec{n}\over{dt^{\prime}}}={c\over{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|}}\left[-\vec{\beta}+\vec{n}\left(\vec{n}\cdot\vec{\beta}\right)\right]~{},$ (139)
Eq. (138) can be written as
$\vec{\bar{E}}$ (((( $\vec{r}_{o},\omega)=-{e}\int_{-\infty}^{\infty}dt^{\prime}{\vec{n}\over{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|^{2}}}\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right]$ (141)
[…]
Eq. (LABEL:revtrasfbiss) may now be integrated by parts. When edge terms can be dropped one obtains [76]
[…]
×$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~{}.$
The only assumption made going from Eq. (138) to Eq. (144) is that edge terms in the integration by parts can be dropped. This assumption can be justified by means of physical arguments in the most general situation accounting for the fact that the integral in $dt^{\prime}$ has to be performed over the entire history of the particle and that at $t^{\prime}=-\infty$ and $t^{\prime}=+\infty$ the electron does not contribute to the field anymore. Let us give a concrete example for an ultra-relativistic electron. Imagine that bending magnets are placed at the beginning and at the end of a given setup, such that they deflect the electron trajectory of an angle much larger than the maximal observation angle of interest for radiation from a bending magnet. This means that the magnets would be longer than the formation length associated with the bends, i.e. $L_{\mathrm{fb}}\sim(c\rho^{2}/\omega)^{1/3}$, where $\rho$ is the bending radius. In this way, intuitively, the magnets act like switches: the first magnet switches the radiation on, the second switches it off. Then, what precedes the upstream bend and what follows the downstream bend does not contribute to the field detected at the screen position. With these caveat Eq. (144) is completely equivalent to Eq. (138).
The derivation of Eq. (144) is particularly instructive because shows that each term in Eq. (144) is due to a combination of velocity and acceleration terms in Eq. (138). In other words the terms in $1/R$ and in $1/R^{2}$ in Eq. (144) appear as a combination of the terms in $1/R$ (acceleration term) and $1/R^{2}$ (velocity term) in Eq. (138). As a result, one can say that there exist contributions to the radiation from the velocity part in Eq. (138). The presentation in Eq. (144) is more interesting with respect to that in Eq. (138) (although equivalent to it) because the magnitude of the $1/R^{2}$-term in Eq. (144) can directly be compared with the magnitude of the $1/R$-term inside the integral sign.
The bottom line is that physical sense can be ascribed only to the integral in Eq. (138) or Eq. (144). The integrand is, in fact, an artificial construction. In this regard, it is interesting to note that the integration by parts giving Eq. (144) is not unique. First, we find that [76]
${\vec{n}\times[(\vec{n}-\vec{\beta})\times\vec{\dot{\beta}}]\over{|\vec{r_{o}}-\vec{r^{\prime}}|(1-\vec{n}\cdot\vec{\beta})^{2}}}$ $=$ ${1\over{|\vec{r_{o}}-\vec{r^{\prime}}|}}{d\over{dt^{\prime}}}\left[{\vec{n}\times(\vec{n}\times\vec{\beta})\over{(1-\vec{n}\cdot\vec{\beta})}}\right]$ (146)
$-\left[{\vec{\dot{n}}(\vec{n}\cdot\vec{\beta})+\vec{n}(\vec{\dot{n}}\cdot\vec{\beta})-\vec{\dot{n}}(\vec{n}\cdot\vec{\beta})^{2}-\vec{\beta}(\vec{\dot{n}}\cdot\vec{\beta})\over{|\vec{r_{o}}-\vec{r^{\prime}}|(1-\vec{n}\cdot\vec{\beta})^{2}}}\right].$
Note that Eq. (146) accounts for the fact that $\vec{n}=(\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime}))/|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|$ is not a constant in time. Using Eq. (146) in the integration by parts, we obtain
[…]
×$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~{}.$
Similarly as before, the edge terms have been dropped. Eq. (138), Eq. (144) and Eq. (148) are equivalent but include different integrands. This is no mistake, as different integrands can lead to the same integral.
If the position of the observer is far away enough from the charge, one can make the expansion Eq. (30). Using Eq. (148), we obtain Eq. (LABEL:revwied).
Appendix II. Undulator Radiation in Resonance Approximation. Far Zone
Calculations pertaining undulator radiation are well established see e.g. [77]. In all generality, the field in Eq. (46) can be written as
${\vec{\widetilde{E}}}=\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\frac{i\omega e}{c^{2}z_{0}}$ (149)
×$\times\int_{-L/2}^{L/2}dz^{\prime}\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}$ (150)
×$\times\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}-{K\theta_{x}\over{\gamma}}{\omega\over{k_{w}c}}\cos(k_{w}z^{\prime})-{K^{2}\over{8\gamma^{2}}}{\omega\over{k_{w}c}}\sin(2k_{w}z^{\prime})\right]~{}.$ (151)
Here $\omega=\omega_{r}+\Delta\omega$, $C=k_{w}\Delta\omega/\omega_{r}$ and
$\omega_{r}=2k_{w}c\bar{\gamma}_{z}^{2}~{},$ (153)
is the fundamental resonance frequency.
Using the Anger-Jacobi expansion:
$\exp\left[ia\sin(\psi)\right]=\sum_{p=-\infty}^{\infty}J_{p}(a)\exp\left[ip\psi\right]~{},$ (154)
where $J_{p}(\cdot)$ indicates the Bessel function of the first kind of order $p$, to write the integral in Eq. (152) in a different way:
${\vec{\widetilde{E}}}=\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\frac{i\omega e}{c^{2}z_{0}}\sum_{m,n=-\infty}^{\infty}J_{m}(u)J_{n}(v)\exp\left[\frac{i\pi n}{2}\right]$ (155)
×$\times\int_{-L/2}^{L/2}dz^{\prime}\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}\right]$ (156)
×$\times\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}\exp\left[i(n+2m)k_{w}z^{\prime}\right]~{},$ (157)
where [^1] Here the parameter $v$ should not be confused with the velocity.
$u=-\frac{K^{2}\omega}{8\gamma^{2}k_{w}c}~{}~{}~{}~{}\mathrm{and}~{}~{}~{}v=-\frac{K\theta_{x}\omega}{\gamma k_{w}c}~{}.$ (159)
Up to now we just re-wrote Eq. (46) in a different way. Eq. (46) and Eq. (158) are equivalent. Of course, definition of $C$ is suited to investigate frequencies around the fundamental harmonic but no approximation is taken besides the paraxial approximation.
Whenever
$C+\frac{\omega\theta^{2}}{{2c}}\ll k_{w}~{},$ (160)
[…]
Within the resonance approximation we further select frequencies such that
$\frac{|\Delta\omega|}{\omega_{r}}\ll 1~{},~{}~{}~{}~{}\mathrm{i.e.}~{}~{}|C|\ll k_{w}~{}.$ (161)
Note that this condition on frequencies automatically selects observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$. In fact, if one considers observation angles outside the range $\theta^{2}\ll 1/\gamma_{z}^{2}$, condition (160) is not fulfilled, and the integrand in Eq. (158) exhibits fast oscillations on the integration scale $L$. As a result, one obtains zero transverse field, $\vec{\widetilde{E}}=0$, with accuracy $1/N_{w}$. Under the constraint imposed by (161), independently of the value of $K$ and for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, we have
$|v|={K|\theta_{x}|\over{\gamma}}{\omega\over{k_{w}c}}=\left(1+\frac{\Delta\omega}{\omega_{r}}\right)\frac{2\sqrt{2}K}{\sqrt{2+K^{2}}}\bar{\gamma}_{z}|\theta_{x}|\lesssim\bar{\gamma}_{z}|\theta_{x}|\ll 1~{}.$ (162)
This means that, independently of $K$, $|v|\ll 1$ and we may expand $J_{n}(v)$ in Eq. (158) according to $J_{n}(v)\simeq[2^{-n}/\Gamma(1+n)]~{}v^{n}$, $\Gamma(\cdot)$ being the Euler gamma function
[…]
Similar reasonings can be done for frequencies around higher harmonics with a different definition of the detuning parameter $C$. However, around odd harmonics, the before-mentioned expansion, together with the application of the resonance approximation for $N_{w}\gg 1$ (fast oscillating terms in powers of $\exp[ik_{w}z^{\prime}]$ effectively average to zero), yields extra-simplifications.
Here we are dealing specifically with the first harmonic. Therefore, these extra-simplifications apply. We neglect both the term in $\cos(k_{w}z^{\prime})$ in the phase of Eq. (152) and the term in $\vec{\theta}$ in Eq. (152). First, non-negligible terms in the expansion of $J_{n}(v)$ are those for small values of $n$, since $J_{n}(v)\sim v^{n}$, with $|v|\ll 1$. The value $n=0$ gives a non-negligible contribution $J_{0}(v)\sim 1$. Then, since the integration in $dz^{\prime}$ is performed over a large number of undulator periods $N_{w}\gg 1$, all terms of the expansion in Eq. (158) but those for $m=-1$ and $m=0$ average to zero due to resonance approximation. Note that surviving contributions are proportional to $K/\gamma$, and can be traced back to the term in $\vec{e}_{x}$ only, while the term in $\vec{\theta}$ in Eq. (158) averages to zero for $n=0$. Values $n=\pm 1$ already give negligible contributions. In fact, $J_{\pm 1}(v)\sim v$. Then, the term in $\vec{e}_{x}$ in Eq. (158) is $v$ times the term with $n=0$ and is immediately negligible, regardless of the values of $m$. The term in $\vec{\theta}$ would survive averaging when $n=1,~{}m=-1$ and when $n=-1,~{}m=0$. However, it scales as $\vec{\theta}v$. Now, using condition (161) we see that, for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, $|\vec{\theta}|~{}|v|\sim(\sqrt{2}~{}K~{}/\sqrt{2+K^{2}}~{})~{}\bar{\gamma}_{z}\theta^{2}\ll K/\gamma$. Therefore, the term in $\vec{\theta}$ is negligible with respect to the term in $\vec{e}_{x}$ for $n=0$, that scales as $K/\gamma$. All terms corresponding to larger values of $|n|$ are negligible.
Summing up, all terms of the expansion in Eq. (154) but those for $n=0$ and $m=-1$ or $m=0$ give negligible contribution. After definition of
$A_{JJ}=J_{0}\left(\frac{\omega K^{2}}{8k_{w}c\gamma^{2}}\right)-J_{1}\left(\frac{\omega K^{2}}{8k_{w}c\gamma^{2}}\right)~{},$ (164)
that can be calculated at $\omega=\omega_{r}$ since $|C|\ll k_{w}$, we have
[…]
Appendix III. Approximation for the Electron Path
Let us now discuss the case of the radiation from a single electron with an arbitrary angular deflection $\vec{\eta}$ and an arbitrary offset $\vec{l}$ with respect to a reference orbit defined as the path through the origin of the coordinate system, that is $x(0)=y(0)=0$.
If the magnetic field in the setup does not depend on the transverse coordinates, i.e. $B=B(z)$, an initial offset $x(0)=l_{x}$, $y(0)=l_{y}$ shifts the path of an electron of $\vec{l}$. Similarly, an angular deflection $\vec{\eta}=(\eta_{x},\eta_{y})$ at $z=0$ tilts the path without modifying it. Cases when the magnetic field of SR sources include focusing elements (or the natural focusing of insertion devices) are out of the scope of this paper. Assuming further that $\eta_{x}\ll 1$ and $\eta_{y}\ll 1$, which is typically justified for ultrarelativistic electron beams, one obtains the following approximation for the electron path:
$x(z)=x_{r}(z)+\eta_{x}z+l_{x}~{},y(z)=y_{r}(z)+\eta_{y}z+l_{y}~{},$ (167)
where the subscript ‘r’ refers to the reference path. This gives a parametric description of the path of a single electron with offset $\vec{l}$ and deflection $\vec{\eta}$. The curvilinear abscissa on the path can then be written as
[…]
$=s_{r}(z)+\frac{\eta^{2}z}{2}+x_{r}(z)\eta_{x}+y_{r}(z)\eta_{y}~{},$ (171)
[…]
We now substitute Eq. (168) and Eq. (171) into Eq. (38) to obtain:
[…]
×$\times\left[\left({v_{x}(z^{\prime})\over{c}}-(\theta_{x}-\eta_{x})\right){\vec{e_{x}}}+\left({v_{y}(z^{\prime})\over{c}}-(\theta_{y}-\eta_{y})\right){\vec{e_{y}}}\right]~{},$ (173)
where the total phase $\Phi_{T}$ is
$\Phi_{T}=\omega\left[\frac{s_{r}(z^{\prime})}{v}+\frac{\eta^{2}z^{\prime}}{2v}+\frac{1}{v}[x_{r}(z^{\prime})\eta_{x}+y_{r}(z^{\prime})\eta_{y}]-{z^{\prime}\over{c}}\right]$ (175)
$+\frac{\omega}{2c}\left[z_{0}(\theta_{x}^{2}+\theta_{y}^{2})-2\theta_{x}x_{r}(z^{\prime})-2\theta_{x}\eta_{x}z^{\prime}-2\theta_{x}l_{x}\right.$ (176)
$\left.-2\theta_{y}y(z^{\prime})-2\theta_{y}\eta_{y}z^{\prime}-2\theta_{y}l_{y}+z^{\prime}(\theta_{x}^{2}+\theta_{y}^{2})\right]~{},$ (177)
which can be rearranged as
$\Phi_{T}\simeq\omega\left[{s_{r}(z^{\prime})\over{v}}-{z^{\prime}\over{c}}\right]-\frac{\omega}{c}(\theta_{x}l_{x}+\theta_{y}l_{y})$ (178)
$+\frac{\omega}{2c}\left[z_{0}(\theta_{x}^{2}+\theta_{y}^{2})-2(\theta_{x}-\eta_{x})x_{r}(z^{\prime})\right.$ (179)
$\left.-2(\theta_{y}-\eta_{y})y_{r}(z^{\prime})+z^{\prime}\left((\theta_{x}-\eta_{x})^{2}+(\theta_{y}-\eta_{y})^{2}\right)\right]~{}.$ (180)
Appendix IV. Self-Electromagnetic Fields of the Modulated Electron Beam
In general, the electrodynamical theory is based on the exploitation, for the ultra-relativistic particles, of the small parameter $1/\gamma^{2}$. By this, Maxwell’s equations are reduced to much simpler equations with the help of paraxial approximation.
Whatever the method used to present results, one needs to solve Maxwell’s equations in unbounded space. We introduce a cartesian coordinate system, where a point in space is identified by a longitudinal coordinate $z$ and transverse position $\vec{r}_{\perp}$. Accounting for electromagnetic sources, i.e. in a region of space where current and charge densities are present, the following equation for the field in the space-frequency domain holds in all generality:
$c^{2}\nabla^{2}\vec{\bar{E}}+\omega^{2}\vec{\bar{E}}=4\pi c^{2}\vec{\nabla}\bar{\rho}-4\pi i\omega\vec{\bar{j}}~{},$ (182)
where $\bar{\rho}(z,\vec{r}_{\perp},\omega)$ and $\vec{\bar{j}}(z,\vec{r}_{\perp},\omega)$ are the Fourier transforms of the charge density $\rho(z,\vec{r}_{\perp},t)$ and of the current density $\vec{j}(z,\vec{r}_{\perp},t)$. Eq. (182) is the well-known Helmholtz equation. Here $\vec{\bar{E}}$ indicates the Fourier transform of the electric field in the space-time domain.
Eq. (182) can be solved with the help of an appropriate Green’s function $G(z-z^{\prime},\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot})$ yielding
[…]
the integration in $d\vec{r^{\prime}}_{\bot}$ being performed over the entire transverse plane. An explicit expression for the Green’s function to be used in Eq. (183) is given by
$G(z-z^{\prime},\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot})$ $=$ $-\frac{\exp\left\{i(\omega/c)\left[\left|\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\right|^{2}+\left(z-z^{\prime}\right)^{2}\right]^{1/2}\right\}}{4\pi\left[\left|\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\right|^{2}+\left(z-z^{\prime}\right)^{2}\right]^{1/2}}~{},$ (184)
that automatically includes the proper boundary conditions at infinity.
[…]
$\mathcal{D}\left[\vec{\widetilde{E}}_{\bot}(z,\vec{r}_{\bot},\omega)\right]=\vec{g}(z,\vec{r}_{\bot},\omega)~{},$ (185)
where the differential operator $\mathcal{D}$ is defined by
$\mathcal{D}\equiv\left({\nabla_{\bot}}^{2}+{2i\omega\over{c}}{\partial\over{\partial z}}\right)~{},$ (186)
${\nabla_{\bot}}^{2}$ being the Laplacian operator over transverse cartesian coordinates. Eq. (185) is Maxwell’s equation in paraxial approximation.
[…]
The source-term vector $\vec{g}(z,\vec{r}_{\bot})$ is specified by the trajectory of the source electrons, and can be written in terms of the Fourier transform of the transverse current density, $\vec{\bar{j}}_{\bot}(z,\vec{r}_{\bot},\omega)$, and of the charge density, $\bar{\rho}(z,\vec{r}_{\bot},\omega)$, as
$\vec{g}=$ $-{4\pi}\exp\left[-\frac{i\omega z}{c}\right]\left(\frac{i\omega}{c^{2}}\vec{\bar{j}}_{\bot}-\vec{\nabla}_{\bot}\bar{\rho}\right)~{}.$ (187)
$\vec{\bar{j}}_{\bot}$ and $\bar{\rho}$ are regarded as given data. We will treat $\vec{\bar{j}}_{\bot}$ and $\bar{\rho}$ as macroscopic quantities, without investigating individual electron contributions. In the time domain, we may write the charge density $\rho(\vec{r},t)$ and the current density $\vec{j}(\vec{r},t)$ as
$\rho(\vec{r},t)=\frac{1}{v}\rho_{\bot}(\vec{r}_{\bot})f\left(t-\frac{z}{v}\right)$ (188)
and
$\vec{j}(\vec{r},t)$ $=$ $\frac{1}{v}\vec{v}\rho_{\bot}(\vec{r}_{\bot})f\left(t-\frac{z}{v}\right)~{},$ (189)
where $v$ denote the velocity of the electron. The quantity $\rho_{\bot}$ has the meaning of transverse electron beam distribution, while $f$ is the longitudinal charge density distribution.
In the space-frequency domain, Eq. (188) and Eq. (LABEL:curr) transform to:
$\bar{\rho}(\vec{r}_{\bot},z,\omega)=\rho_{\bot}\left(\vec{r}_{\bot}\right)\bar{f}(\omega)\exp\left[{i\omega z)/v}\right]~{},$ (191)
and
$\vec{\bar{j}}(\vec{r}_{\bot},z,\omega)=\vec{v}\rho_{\bot}\left(\vec{r}_{\bot}\right)\bar{f}(\omega)\exp\left[{i\omega z/v}\right]~{}.$ (192)
It should be remarked that $\bar{\rho}$ and $\vec{\bar{j}}=\bar{\rho}\vec{v}$ satisfy the continuity equation. In other words, one can find $\vec{\nabla}\cdot\vec{\bar{j}}=i\omega\bar{\rho}$.
We find an exact solution of Eq. (185) without any other assumption about the parameters of the problem. A Green’s function for Eq. (185), namely the solution corresponding to the unit point source can be written as (see e.g. [76]):
$G(z-z^{\prime};\vec{r_{\bot}}-\vec{r^{\prime}_{\bot}})$ $=$ $-{1\over{4\pi(z-z^{\prime})}}\exp\left\{i\omega{\mid\vec{r_{\bot}}-\vec{r^{\prime}_{\bot}}\mid^{2}\over{2c(z-z^{\prime})}}\right\}~{},$ (193)
assuming $z-z^{\prime}>0$. When $z-z^{\prime}<0$ the paraxial approximation does not hold, and the paraxial wave equation Eq. (185) should be substituted, in the space-frequency domain, by a more general Helmholtz equation. Yet, the radiation formation length for $z-z^{\prime}<0$ is very short with respect to the case $z-z^{\prime}>0$, i.e. we can neglect contributions from sources located at $z-z^{\prime}<0$.
Since it is assumed that electrons are moving along the $z$-axis, we have $\vec{\bar{j}}_{\perp}=0$. Thus, after integration by parts, we obtain the solution
[…]
×$\times\frac{1}{z-z^{\prime}}{\rho_{\bot}}\left(\vec{r^{\prime}}_{\bot}\right)\left(\frac{\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}}{z-z^{\prime}}\right)~{}.$
Eq. (LABEL:ggeneralfin) describes the field at any position $z$.
[…]
$\vec{\widetilde{E}}_{\bot}(z,\vec{r}_{\bot})$ $=$ $-\frac{i\omega\bar{f}(\omega)}{c}\int d\vec{r^{\prime}}_{\bot}{\rho_{\bot}}\left(\vec{r^{\prime}}_{\bot}\right)\exp\left[\frac{i\omega z}{2c\gamma^{2}}\right]\Bigg{\{}\left[\frac{ic}{\omega}\frac{\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}}{\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid}\cdot\frac{d}{d\left[\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid\right]}\right]$ (198)
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\int_{0}^{\infty}\frac{d\xi}{\xi}\exp\left[+i\omega\frac{\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid^{2}}{2c\xi}-\frac{i\omega\xi}{2c\gamma^{2}}\right]\Bigg{\}}$
We now use the fact that, for any real number $\alpha>0$:
[…]
where $K_{0}$ is the zero order modified Bessel function of the second kind. Using Eq. (199) we can write Eq. (198) as
[…]
where $K_{1}(\cdot)$ is the modified Bessel function of the first order.
[…]
$\vec{\widetilde{E}}_{\bot}(z,\vec{r}_{\bot})=-\frac{2\omega e}{c^{2}\gamma}\exp\left[\frac{i\omega z}{2c\gamma^{2}}\right]\frac{\vec{r}_{\bot}}{r_{\bot}}K_{1}\left(\frac{\omega r_{\bot}}{c\gamma}\right)~{}.$ (202)
[…]

### v10

@@
 SPECIAL RELATIVITY
-Evgeny Saldin Draft from May 2024
+Evgeny Saldin Final draft
 Applications to astronomy and the accelerator physics
-Evgeny Saldin Draft from May 2024
-New material has been added to the Chapter 1, Chapter 3, Chapter 4, Chapter 5, Chapter 9, Chapter 10, Chapter 11, Chapter 13.
+Evgeny Saldin Final draft
 If you have any comments, please send an email to evgueni.saldin@desy.de.
@@
 There are many books on the classical subject of special relativity.
-However, after having spent a number of years, both in relativistic engineering and research with relativity, I have come to the conclusion that there exist a place for a new book.
-I do believe that the present book is not quite the same as the others, mainly due to attempt to cast light on dark corners.
+However, after having spent a number of years, both in relativistic engineering and research with relativity, I have come to the conclusion that there exists a place for a new book.
+I do believe that the present book is not quite the same as the others, mainly due to the attempt to cast light on dark corners.
 I should make it clear what this little book is not.
@@
 What the book is about is the nature of special relativistic kinematics, its relation to space and time, and the operational interpretation of coordinate transformations.
-Every theory contains a number of quantities that can be measured by experiment and an expressions that cannot possibly be observed.
+Every theory contains a number of quantities that can be measured by experiment and expressions that cannot possibly be observed.
 Whenever we have a theory containing an arbitrary convention, we should examine what parts of the theory depend on the choice of that convention and what parts do not.
 The distinction is not always made and many authors claim some data to be observable, according to arbitrary conventions, which do not correspond to any physical experiment.
-This leads to inconsistencies and paradoxes that should be avoided at all cost.
+This leads to inconsistencies and paradoxes that should be avoided at all costs.
 The practical approach used in the book should be acceptable to astronomers, space engineers, accelerator engineers, and more generally, relativistic engineers.
-This approach, unusual in the relativistic literature, may be clarified by quoting one of the problems discussed in the text: the new light beam kinematics for rotating frame of references.
-Since we live on such a rotating (earth-based) frame of reference, difference in relativistic kinematics between rotating and non-rotating frames of reference is of great practical as well as theoretical significance.
+This approach, unusual in the relativistic literature, may be clarified by quoting one of the problems discussed in the text: the new light beam kinematics for rotating frames of references.
+Since we live on such a rotating (earth-based) frame of reference, the difference in relativistic kinematics between rotating and non-rotating frames of reference is of great practical as well as theoretical significance.
 A correct solution of this problem requires the use of relativistic principles even at low velocities since the first-order terms in $(v/c)$ play a fundamental role in the non-inertial relativistic kinematics of light propagation.
@@
 To preserve a self-consistent style, I place the derivation of auxiliary results in appendices.
-To help readers form their own opinion on the topics discussed, the end of each chapter has a suggested bibliography together with relevant remarks.
+To help readers form their own opinions on the topics discussed, the end of each chapter has a suggested bibliography together with relevant remarks.
 The list of references includes only the papers I have consulted directly.
-A lot of papers remain unmentioned, and for this I apologize.
+A lot of papers remain unmentioned, and for this, I apologize.
 I am grateful to my longtime friends Gianluca Geloni and Vitaly Kocharyan for discussions over many years about much of the material in this book.
@@
 1 Introduction
-The standard books on special relativity do not usually address the questions of the physical meaning of relativistic effects and the nature of space-time.
-The presentation of the subject in the present book differs somewhat from the usual one in that the four-dimensional geometric formulation of the theory plays a dominant role than in most of the current textbooks.
 The book begins with a critical survey of the present approaches to special relativity.
-The established way of looking at special relativity is based on Einstein postulates: the principle of relativity and the constancy of the velocity of light.
-In the most general geometric approach to the theory of special relativity, the principle of relativity, in contrast to Einstein formulation, is only a consequence of the geometry of space-time.
-We point that the essence of the special theory of relativity consists in the following postulate: all physical processes proceed in four-dimensional space-time, the geometry of which is pseudo-Euclidean.
+The established way of looking at special relativity is based on Einstein’s postulates: the principle of relativity and the constancy of the velocity of light.
+In the most general geometric approach to the theory of special relativity, the principle of relativity, in contrast to Einstein’s formulation, is only a consequence of space-time geometry.
+We point out that the essence of the special theory of relativity consists in the following postulate: all physical processes proceed in four-dimensional space-time, the geometry of which is pseudo-Euclidean.
+The presentation of the subject in the present book differs somewhat from the usual one in that the four-dimensional geometric formulation of the theory plays a dominant role.
 The space-time geometric approach to special relativity deals with all possible choices of coordinates of the chosen reference frames, and therefore the second Einstein postulate, referred to as the constancy of the coordinate speed of light, does not have a place in this more general formulation.
@@
 Thus, the basic elements of the space-time geometric formulation of the special relativity and the usual Einstein’s formulation, are quite different.
-It should be emphasized that in practical applications there are two choices of clock synchronization convention useful to consider:
+It should be emphasized that in practical applications there are two choices of clock synchronization conventions useful to consider:
 (a) Einstein’s convention, leading to the Lorentz transformations between frames.
@@
 Indeed, a typical physical laboratory contains no space-time grid.
-It should be clear that a rule-clock structure exist only in our mind and manipulations with non existing clocks in the special relativity are an indispensable prerequisite for the application of dynamics and electrodynamics theory in the coordinate representation.
-Such situation usually forces physicist to believe that the application of the theory of relativity to the study of physical processes is possible without detailed knowledge of the clocks synchronization procedure.
-However, many problems in special relativity can be adequately treated only by an approach which uses the non-standard absolute time synchronization.
-One of the features that is unique to this book is its treatment of the absolute time coordinatization.
-No other books deals with Galilean transformations in the framework of special relativity.
-Third chapter presents an ”operational interpretation” of the Lorentz and absolute time coordinatizations.
-This is probably the most important chapter of this book.
-Today the statement about correctness of Galilean transformations is a ”shocking heresy”, which offends the ”relativistic” intuition and the generally accepted way of looking at special relativity of most physicists.
+It should be clear that a rule-clock structure exists only in our mind and manipulations with nonexisting clocks in special relativity are an indispensable prerequisite for the application of dynamics and electrodynamics theory in the coordinate representation.
+Such a situation usually forces physicists to believe that applying the theory of relativity to the study of physical processes is possible without detailed knowledge of the clock synchronization procedure.
+However, many problems in special relativity can be adequately treated only by an approach that uses the non-standard absolute time synchronization.
+One of the features that is unique to this book is its treatment of the absolute time coordinatization (i.e.
+Galilean transformations) in the framework of special relativity.
+The third chapter presents an ”operational interpretation” of the Lorentz and absolute time coordinatizations.
+This is probably the most important chapter of this article.
+Today the statement about the correctness of Galilean transformations is a ”shocking heresy”, which offends the ”relativistic” intuition and the generally accepted way of looking at special relativity of most physicists.
 The difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view will be an important discovery for every special relativity expert.
-To our knowledge, neither operational interpretation of the absolute time coordinatization nor the difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view, are given elsewhere in the literature.
-We have tried to keep the mathematical complexity of the discussion to a minimum.
-The derivation of relativistic kinematics in the first part of the book (Chapter 3-11) is fairly elementary from a mathematical point of view, but it is conceptually subtle.
-We use synthetic approach to present the material: some simple models are studied first, and more complicated ones are introduced gradually.
-We start with aberration of light phenomena.
+To our knowledge, neither the operational interpretation of absolute time synchronization nor the difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view, are given elsewhere in the literature.
+We start with the aberration of light phenomena.
 Light, being a special case of electromagnetic waves, is described by the electrodynamics theory.
 It is well known that the electrodynamics theory meets all requirements of the theory of relativity and therefore must accurately describe the properties of such a typical relativistic object as light.
-In the Chapter 4-5, we present a critical reexamination of the existing aberration of light theory.
+In Chapters 4-5, we present a critical reexamination of the existing aberration of light theory.
 The phenomenon of aberration of light is by no means simple to describe, even in the first order in $v/c$: a large number of incorrect results can be found in the literature.
 The utilization of the electrodynamics in the absolute time coordinatization becomes indispensable when we consider optical phenomena associated with a relative motion of two (or more) light sources.
-There is no problem to introduce Lorentz coordinates when these sources are independent.
-The possibility to introduce Lorentz coordinates in this situation consists in introducing individual coordinate system (i.e.
-individual set of clocks) for each sources.
+There is no problem with introducing Lorentz coordinates when these sources are independent.
+The possibility to introduce Lorentz coordinates in this situation consists of introducing an individual coordinate system (i.e.
+individual set of clocks) for each source.
 Suppose now that the second (e.g.
@@
 The first error in the textbook argument is conceptual in nature.
-The observer describes the light beam from the emitter at rest using Minkowski metric.
+The observer describes the light beam from the emitter at rest using the Minkowski metric.
 According to textbooks, the metric of the moving mirror is the same (i.e.
 Minkowski metric).
-The authors of textbooks incorrectly belief that the common Lorentz coordinatization for emitter and mirror in the single inertial frame can be assigned.
-This is misconception.
-The second error in the textbooks follows from the fact that the concept of a plane wave and an infinite plane mirror is used in the case of the tangential motion.
+The authors of textbooks incorrectly believe that the common Lorentz coordinatization for emitter and mirror in the single inertial frame can be assigned.
+This is a misconception.
+The second error in the textbooks follows from the fact that the concept of a plane wave and an infinite plane mirror is used in the case of tangential motion.
 In this case the ”mirror motion” is not a real observable effect.
-Indeed, if the mirror is infinite the problem is not time dependent.
-It is obvious that only the motion of a finite mirror is meaningful.
+Indeed, if the mirror is infinite the problem is not time-dependent.
+Only the motion of a finite mirror is meaningful.
 All experimental methods for measuring the aberration increment operate with light signals, and hence do not measure the phase velocity (i.e.
@@
 What authors of textbooks generally overlooked is the fact that the energy transport problem is well-defined only if the source and mirror apertures have already been specified.
-Suppose that we have a point source and observer with his measuring devices are at rest in the inertial lab frame.
+Suppose that we have a point source and observer with his measuring devices at rest in the inertial lab frame.
 We demonstrate that when a finite aperture mirror is moving tangentially with a constant velocity and interacts in the far-field zone with the radiated light, there is a deviation of the energy transport for the reflected light.
 The result is extremely important from the applied point of view.
-According to the Babinet’s principle, this remarkable prediction of our aberration of light theory is correct also for light transmitted through a hole in the moving opaque screen or, consequently, through a moving open end of the telescope barrel.
+According to Babinet’s principle, this remarkable prediction of our aberration of light theory is correct also for light transmitted through a hole in the moving opaque screen or, consequently, through a moving open end of the telescope barrel.
 Questions related to transmission through the transversely moving pupil detection (e.g.
-a moving end of the telescope barrel) lead to serious misunderstanding, which is actually due to an inadequate understanding of several complicated aspects of the statistical optics.
-Today one is told that the phenomenon of aberration of light could be interpreted, using ray model of light.
+a moving end of the telescope barrel) lead to serious misunderstanding, which is due to an inadequate understanding of several complicated aspects of the statistical optics.
+Today one is told that the phenomenon of aberration of light could be interpreted, using a ray model of light.
 One case of rather great interest is that which corresponds to transmission through a moving telescope tube.
-The rays of light coming from the star are falling on the telescope tube and, according to literature, not interact with its sides.
-One could naively expect that the region of applicability of ray optics, following from the textbooks reasoning, should be identified with any spatially incoherent radiation.
-This is misconception.
+The rays of light coming from the star are falling on the telescope tube and, according to the literature, do not interact with its sides.
+One could naively expect that the region of applicability of ray optics, following the textbook’s reasoning, should be identified with any spatially incoherent radiation.
+This is a misconception.
 In particular, a spatially completely incoherent source (e.g.
-an incandescent lamp or a star) is actually a system of elementary (statistically independent) point sources with different offsets.
+an incandescent lamp or a star) is a system of elementary (statistically independent) point sources with different offsets.
 The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$, where $\lambda$ is the (visible) radiation wavelength.
-Radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individually incoherent sources.
-An elementary source produces in front of pupil detection effectively a plane wave and the measuring instrument always influences the measured radiation (due to unavoidable diffraction of a plane wave by an aperture).
-It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single point source characteristics like independence on source motion.
-The Chapter 6 deals with astronomical applications.
+A radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individually incoherent sources.
+An elementary source produces in front of pupil detection effectively a plane wave and the measuring instrument always influences the measured radiation (due to the unavoidable diffraction of a plane wave by an aperture).
+It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single-point source characteristics like independence on source motion.
+Chapter 6 deals with astronomical applications.
 The effect of stellar aberration seems to be one of the simplest phenomena in astronomical observations.
-In spite of its apparent simplicity, aberration seems to be one of the most intricate effects in special relativity.
+Despite its apparent simplicity, aberration seems to be one of the most intricate effects of special relativity.
 It is widely believed that stellar aberration depends on the relative velocity of the source (star) and observer.
-Observations show clearly that stellar aberration does not depend on the relative motion between star and telescope on the earth.
+Observations show clearly that stellar aberration does not depend on the relative motion between stars and telescopes on the earth.
 The lack of symmetry, between the cases when either the source or telescope is moving is shown clearly on the basis of the separation of binary stars.
-The relative motion of these stars with respect to each other (and hence, with respect to the earth) is never followed by any aberration, although motion of these stars is, sometimes, much faster than that of the earth around the sun.
+The relative motion of these stars with respect to each other (and hence, with respect to the earth) is never followed by any aberration, although the motion of these stars is, sometimes, much faster than that of the earth around the sun.
 It should be stressed that it is the telescope and not the star that must change its velocity (relative to the fixed stars) to cause aberration.
-Contradiction is so plane that some physicists use this fact to argue that stellar aberration contradicts the special theory of relativity.
+Contradiction is so plain that some physicists use this fact to argue that stellar aberration contradicts the special theory of relativity.
 There is no available explanation for the fact that, while the observational data on stellar aberration are compatible with moving earth, the symmetric description, when the star possesses the relative transverse motion, does not apparently lead to observations compatible with predictions.
-We demonstrate that the fact that we do not see myriads of widely separated binaries in wild gyration does not require any fundamental change of outlook, but it does require that aberration of ”distant” stars is treated in the framework of space-time geometric approach.
-It is generally believed that the phenomenon of aberration of light could be interpreted, using the corpuscular model of light, as being analogous of the observation of the oblique fall of raindrops by a moving observer.
-This is a classical kinematics method to the computation of the stellar aberration used in astronomy for about three hundred years.
-This book develops a relativistic theory of aberration of light in rotating frames.
-To see the consequences of relativistic modification of stellar aberration, we start with the classical aberration increment, or $\theta_{a}=v/V$.
-Here $v$ is the velocity of observer and $V$ is the velocity of particles.
+We demonstrate that the fact that we do not see myriads of widely separated binaries in wild gyration does not require any fundamental change of outlook, but it does require that the aberration of ”distant” stars is treated in the framework of space-time geometric approach.
+It is generally believed that the phenomenon of aberration of light could be interpreted, using the corpuscular model of light, as being analogous to the observation of the oblique fall of raindrops by a moving observer.
+This is a classical kinematics method for the computation of the stellar aberration used in astronomy for about three hundred years.
+This article develops a relativistic theory of aberration of light in rotating frames.
+To see the consequences of relativistic modification of stellar aberration, we start with the classical aberration increment $\theta_{a}=v/V$.
+Here $v$ is the velocity of the observer and $V$ is the velocity of particles.
 It is assumed that $v\ll V$.
 According to the conventional approach, if we neglect $v^{2}/c^{2}$ comparison with 1, stellar aberration increment is simplified to elementary formula $\theta_{a}=v/c$.
-This result is particularly remarkable because the study of stellar aberration is intimately connected with the old kinematics: the Galilean vectorial law of addition of velocities is actually used.
-We come to the conclusion that the standard analysis of the stellar aberration does not take into account the fundamental difference between velocity of light and velocity of raindrops.
+This result is particularly remarkable because the study of stellar aberration is intimately connected with the old kinematics: the Galilean vectorial law of addition of velocities is used.
+We conclude that the standard analysis of the stellar aberration does not take into account the fundamental difference between the velocity of the light and the velocity of raindrops.
 The satisfactory treatment of the relativistic modification should be based on the two relativistic parameters.
-For an observer on the earth, it is, with respect to the solar referential frame, of about 30 km/s, corresponding to the earth motion around the sun.
+For an observer on the earth, it is, with respect to the solar referential frame, of about 30 km/s, corresponding to the earth’s motion around the sun.
 Clearly, in the theory of stellar aberration, we consider the small expansion parameter $v/c\simeq 10^{-4}$ neglecting terms of order of $v^{2}/c^{2}$.
@@
 Light is always a (”ultra”) relativistic object, no matter how small the ratio $v/c$ may be.
-According to the special theory of relativity, physical processes takes place in a metric space-time geometry, so the geodesic interval must be used to describe aberration of light phenomenon.
+According to the special theory of relativity, physical processes take place in a metric space-time geometry, so the geodesic interval must be used to describe the aberration of light phenomenon.
 The problem of the earth-based measurements is solved with the discovery of the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars.
-We demonstrate that for explanation of the aberration of light effect in the rotating frame of reference one does not need neither modify the special theory of relativity, nor apply general theory of relativity.
+We demonstrate that for an explanation of the aberration of light effect in the rotating frame of reference, one does not need to modify the special theory of relativity nor apply the general theory of relativity.
 It is only necessary to strictly follow the special relativity.
-All phenomena in non-inertial reference system should be considered in the framework of the space-time geometric approach using metric tensor.
+All phenomena in non-inertial reference systems should be considered in the framework of the space-time geometric approach using metric tensor.
 Using Langevin metric in a frame of reference attending rotation enables us to explain all earth-based experiments.
-A correct solution of this problem in the rotating frame requires the use of metric tensor even in experiments of first order in $v/c$ since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of a light beam propagation.
+A correct solution of this problem in the rotating frame requires the use of metric tensor even in experiments of the first order in $v/c$ since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of light beam propagation.
 The historical rate of the studies of the optical effects in rotating frames is very unusual.
@@
 It is interesting that the Langevin metric has in the past hardly ever been applied in the stellar aberration theory.
-We demonstrate that the aberration of light is a complex phenomenon which must be branched out into a number of varieties according to their origin.
+We demonstrate that the aberration of light is a complex phenomenon that must be branched out into a number of varieties according to their origin.
 These branching out takes place depending on what is the cause of aberration - whether it is a jump in the velocity (with respect to the fixed stars) of the observer or of the light source.
 Aberration could undergo further splitting - depending on the physical influence of the optical instrument on the measurement.
-The main facts which a theory of stellar aberration in the earth-based frame of reference must explain are (1) the annual apparent motion of the fixed stars about their locations and (2) the null apparent aberration of rotating binary systems (they exhibit aberrations not different from other stars).
-We present here a theory which accounts for all these, and in addition gives new results.
-All earth-based experiments can be explained on the basis of the effect of the measuring instrument (i.e.
+The main facts that a theory of stellar aberration in the earth-based frame of reference must explain are (1) the annual apparent motion of the fixed stars about their locations and (2) the null apparent aberration of rotating binary systems (they exhibit aberrations not different from other stars).
+We present here a theory that accounts for all these, and in addition, gives new results.
+All earth-based experiments can be explained based on the effect of the measuring instrument (i.e.
 physical unavoidable influence of the telescope on the measurement) and the acceleration of traveling earth-based observer relative to the fixed stars.
-In the Chapter 8 we analyze the potential for exploiting earth-based sources in order to confirm the predictions of the relativistic aberration of light theory.
+In Chapter 10 we analyze the potential for exploiting earth-based sources to confirm the predictions of the relativistic aberration of light theory.
 We discussed a simple scaling model for the stellar aberration.
 The motion of the stars with respect to the earth is never followed by any aberration.
-The aberration shift exist even in the case when star moves with the same velocity as the earth.
-We obtained a condition for optical similarity between the aberration of light from a distant star (which is moving with the earth velocity) and from the earth-based incoherent source.
-The proposed method of measuring the angle of aberration involves the use of earth-based sources and have a big advantage.
+The aberration shift exists even in the case when a star moves with the same velocity as the Earth.
+We obtained a condition for optical similarity between the aberration of light from a distant star (which is moving with the earth’s velocity) and from the earth-based incoherent source.
+The proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage.
 The rotation of the earth on its axis should produce a corresponding shift of the image.
-The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth rotation axis.
-The apparent position of the source image is thus always a little displaced in the direction of the earth motion around the sun at that moment, and hence describe a small elliptical figure during the annular revolution the earth around its axis.
-In principle, records could be taken over a period of one day.
-It is generally believed that the special relativity is reciprocal theory.
+The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth’s rotation axis.
+The apparent position of the source image is thus always a little displaced in the direction of the earth’s motion around the sun at that moment, and hence describes a small elliptical figure during the annular revolution of the earth around its axis.
+In principle, records could be taken over one day.
+It is generally believed that special relativity is a reciprocal theory.
 The aberration of light problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers.
-In fact, without looking at anything external to the earth-based accelerated frame, one could determine the speed of the earth with respect to the sun-based frame by means of aberration of earth-based incoherent source measurements.
-No one has ever done such experiment, but we know what would happen from the astronomical observations.
+In fact, without looking at anything external to the earth-based accelerated frame, one could determine the speed of the earth with respect to the sun-based frame using aberration of earth-based incoherent source measurements.
+No one has ever done such an experiment, but we know what would happen from the astronomical observations.
 Now many people who learn special relativity in the usual way find this disturbing. The argument that aberration shift must be symmetrical runs something like this:
@@
 (2) Information about the direction of observer acceleration is recorded in the crossed term of the Langevin metric.
-(3) It is always possible to chose such variables, in which metric of the accelerated frame will be diagonal. This is a consequence of the (pseudo-Euclidean) geometry of space-time.
+(3) It is always possible to choose such variables, in which the metric of the accelerated frame will be diagonal. This is a consequence of the (pseudo-Euclidean) geometry of space-time.
 (4) The validity of Maxwell’s laws in all inertial frames implies that light propagates with the same velocity $c$ in all inertial frames.
-It implies the elimination of the privilege frame, so meaning that the notion of absolute motion does not have physical content at all.
-We discuss this subject in the Chapter 9.
-In the argument above, steps (2) and (3) are correct, but step (1), and its consequence (4), are wrong.
+It implies the elimination of the privileged frame, so meaning that the notion of absolute motion does not have physical content at all.
+In the argument above, steps (2) and (3) are correct, but steps (1), and its consequence (4), are wrong.
 Argument (3) says that all inertial frames are equivalent with regard to physical laws, not with regard to physical facts.
 At first glance, after the diagonalization of the Langevin metric, we have the symmetry between inertial frames.
-Where does the asymmetry comes from?
-The electrodynamics equations are not supplies all information about the physics of the situation.
+Where does the asymmetry come from?
+The electrodynamics equations do not supply all information about the physics of the situation.
 To solve the electrodynamics equations it is necessary to determine the initial conditions.
@@
 Now the information about the direction of observer acceleration is recorded in the initial conditions (radiation wavefront orientation).
-One can conclude that not all is relative in relativity, because this theory also contains some features which are absolute.
-Many people would like to think that since in the Lorentz coordinatization the time and distance have direct physical meaning, there should be some physical (dynamical) reason for this wavefront rotation.
-They would think that dynamics, based on the physical fields, is actually hidden in the language of pseudo-Euclidean geometry.
+One can conclude that not all is relative in relativity because this theory also contains some features which are absolute.
+Many people would like to think that since in the Lorentz coordinatization, the time and distance have direct physical meaning, there should be some physical (dynamical) reason for this wavefront rotation.
+They would think that dynamics, based on the physical fields, is hidden in the language of pseudo-Euclidean geometry.
 However, it is a dynamical line of arguments that explains this paradoxical situation with the asymmetry between inertial and accelerated reference frames.
@@
 Wavefront rotation associated with the transformation from the inertial frame to the accelerated (with respect to the fixed stars) frame may be regarded as a result of the action of pseudo-gravitational potential gradient during the acceleration process.
-See Chapter 9 for more ideas along this line.
-There are several point to be made about the above results.
-One interesting question is, why we are not discuss about the aberration of light radiated by a laser?
-In regard to light aberration one should differentiate between that from the incoherent source and that from the laser source.
-One could naively expect that the rotation of the earth around the sun produces aberration in an amount large enough to be taken into account in precise observation work using laser as a earth-based light source.
-Let us stress that the aberration of light cannot be measured using laser.
-In fact, the electromagnetic wave travels in the laser resonator forward and back reflecting from the mirrors.
+There are several points to be made about the above results.
+One interesting question is, why we are not discussing the aberration of light radiated by a laser?
+Regarding light aberration, one should differentiate between that from the incoherent source and that from the laser source.
+One could naively expect that the rotation of the earth around the sun produces aberration in an amount large enough to be taken into account in precise observation work using a laser as an earth-based light source.
+Let us stress that the aberration of light cannot be measured using a laser.
+The electromagnetic wave travels in the laser resonator forward and back reflecting from the mirrors.
 It could be said that the asymmetry cancels during the (round-trip-to-round-trip) evolution of the radiation in the optical resonator.
 We would like to emphasize a very important difference between laser and incoherent source.
-The radiation of incoherent source clearly depends on an initial conditions and there is influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer.
-In contrast, when we are dealing with optical resonator the (3-space) boundary conditions define the energy propagation direction.
-One finds many books which say that we cannot distinguish by any experiment which observer remains ”at rest” during the acceleration, because according to the principle of relativity only relative motion has any physical meaning.
+The radiation of an incoherent source clearly depends on initial conditions and there is an influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer.
+In contrast, when we are dealing with an optical resonator the (3-space) boundary conditions define the energy propagation direction.
+One finds many books that say that we cannot distinguish by any experiment which observer remains ”at rest” during the acceleration because according to the principle of relativity, only relative motion has any physical meaning.
 It follows that any observed asymmetry would lead to a contradiction with the principle of relativity.
 This argument is wrong.
-We discuss the apparent conflict between aberration of light and the principle of relativity in the Chapter 12.
 It is demonstrated that there is no conflict between the fundamental structure of special relativity on the one hand, and the aberration of light phenomena.
-Not only in Einstein writings, but in every textbook on special theory of relativity can be found the formulation of principle of special relativity as follows: the laws of physics are the same in all inertial frames.
-Principle of special relativity is irrelevancy of velocity with regard to physical laws, not with regard to anything.
+Not only in Einstein’s writings, but in every textbook on the special theory of relativity can be found the formulation of the principle of special relativity as follows: the laws of physics are the same in all inertial frames.
+The principle of special relativity is irrelevancy of velocity with regard to physical laws, not with regard to anything.
 It is generally accepted that covariant equations must have covariant solutions.
-This statement presented in most textbooks and is obviously incorrect.
+This statement is presented in most textbooks and is incorrect.
 For instance, from the mathematical viewpoint, the Lorentz covariance of electrodynamics equations is insufficient to guarantee the covariant solutions.
@@
 Relativistic (i.e.
-reciprocal) symmetry is usually identifies with symmetry under Lorentz boosts.
+reciprocal) symmetry is usually identified with symmetry under Lorentz boosts.
 At the same time it is common knowledge that Lorentz boosts alone do not make up a group.
 The product of two Lorentz boosts in different directions is equal to the product of a pure boost and the spatial (Wigner) rotation.
-From the mathematical viewpoint, when we are dealing with transformations and set of these transformations does not form a group we have no symmetry between inertial frames.
+From the mathematical viewpoint, when we are dealing with transformations and a set of these transformations does not form a group we have no symmetry between inertial frames.
 The inertial frames are characterized by motion at constant velocity relative to the fixed stars.
 There is a real difference between accelerated ( with respect to the fixed stars) inertial frame and the inertial frame without accelerational history and this difference ( in the case of the aberration of light phenomena) is closely associated with the Wigner rotation effect.
-It should be note that all well known methods to test the special relativity are second order optical (interference) measurements .
+It should be noted that all well-known methods to test special relativity are second-order optical (interference) measurements.
 The cardinal example is given by the Michelson-Morley experiment.
 The light from a source is split by a suitable apparatus into two beams and the two overlapping beams produce an interference pattern.
-The two beam originate in the same source.
-It is not possible to detect the orbital velocity using second order (interference) experiments.
+The two beams originate from the same source.
+It is not possible to detect the orbital velocity using second-order (interference) experiments.
 Phase is 4D invariant (i.e.
-it is simply number).
-It is independent on the chosen inertial frame and on the chosen coordinatization.
+it is simply a number).
+It is independent of the chosen inertial frame and the chosen coordinatization.
 This is a consequence of the geometry of space-time.
-In the interference experiments neither the Doppler effect nor the aberration of light exist separately as well defined physical phenomena and only phase is meaningful quantity.
-This discussion clearly show that the special relativity calculations yields the observed null fringe shift in the earth-based frame.
-A close examination of all optical second order experiments inside an inertial frame shows that phenomena appear to be independent of the uniform motion relative to the fixed stars.
-The usual formulation of the theory of relativity deals with formulated above principle of irrelevancy of velocity, but it should be understood only in a limiting sense when we are dealing with optical second order measurements.
-There are a number of other second order experiments.
+In the interference experiments neither the Doppler effect nor the aberration of light exist separately as well-defined physical phenomena and only phase is a meaningful quantity.
+This discussion clearly shows that the special relativity calculations yield the observed null fringe shift in the earth-based frame.
+A close examination of all optical second-order experiments inside an inertial frame shows that phenomena appear to be independent of the uniform motion relative to the fixed stars.
+The usual formulation of the theory of relativity deals with formulated above principle of irrelevancy of velocity, but it should be understood only in a limiting sense when we are dealing with optical second-order measurements.
+There are a number of other second-order experiments.
 They are not connected with light interference phenomena.
-The question arises whether it is possible to determine experimentally the state of motion of the initial inertial frame in the accelerated frame by means of time dilation effect for moving reference clock.
-We demonstrate that slowly down of the accelerated (with respect to the fixed stars) physical clock does not depend on the choice of reference system, in which this effect is measured.
+The question arises whether it is possible to determine experimentally the state of motion of the initial inertial frame in the accelerated frame using the time dilation effect for a moving reference clock.
+We demonstrate that the slowdown of the accelerated (with respect to the fixed stars) physical clock does not depend on the choice of reference system, in which this effect is measured.
 In other words, the time dilation effect is absolute, but not relative.
@@
 In textbooks, time dilation is usually derived from the diagonal (Minkowski) metric.
-After diagonalization, according to textbooks, the coordinates in inertial frames should be related by the Lorentz transformation.
-At first site the problem is complete symmetrical and the time dilation effect is relative.
+After diagonalization, according to textbooks, the coordinates in inertial frames should be related to the Lorentz transformation.
+At first sight, the problem is completely symmetrical and the time dilation effect is relative.
 Something is evidently wrong.
-In the Chapter 9 we investigate the connection between time dilation effect and the Langevin metric.
+In Chapter 12 we investigate the connection between the time dilation effect and the Langevin metric.
 Perhaps the only way to demonstrate that the time dilation effect is absolute is to take a mathematical view.
-From a mathematical viewpoint, the argument that time dilation effect is absolute looks something like the following.
-It is obvious that metric tensor must be a continuous quantity.
+From a mathematical viewpoint, the argument that the time dilation effect is absolute looks something like the following.
+The metric tensor must be a continuous quantity.
 The Langevin metric is found by matching the accelerated frame and inertial frame metric tensors.
 For smooth tailoring of the metric tensor, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the Galilean boost.
-In this case the metric of inertial frame would pass into the Langevin metric of the accelerated frame.
+In this case, the metric of the inertial frame would pass into the Langevin metric of the accelerated frame.
 After diagonalization, the metric tensor in the accelerated frame must abruptly change from the Langevin value to the Minkowski (diagonal ) value.
-We must conclude that when we are dealing with Lorentz transformation and this transformation does not preserve continuity of the metric tensor we have no symmetry between inertial frames.
+We must conclude that when we are dealing with Lorentz transformation and this transformation does not preserve the continuity of the metric tensor we have no symmetry between inertial frames.
 Any change of velocity, or any acceleration (with respect to the fixed stars) has an absolute meaning.
-In the Chapter 9-12 we will take up in more detail the absoluteness of acceleration; here we only wished to show how naturally the special relativity leads to the concept of absolute velocity.
 The formalism of relativistic physics is associated with absolute space-time structure.
-This structure must be associated with initial inertial frame, which incidentally provides an absolute rest frame.
-There is no relativity of the uniform motions and we can determine the particle absolute velocity in the initial inertial frame.
-The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of unified space-time model.
-In special relativity the absolute space and the absolute time of Newton are fused into absolute space-time.
-The proper time interval may be identified with particle lifetime, atomic transition periods, nuclear half-life.
-Until now, all test of time dilation with atomic clocks in the earth-based frame have involved circle path geometry.
+This structure must be associated with an initial inertial frame, which incidentally provides an absolute rest frame.
+There is no relativity of the uniform motions and we can determine the particle’s absolute velocity in the initial inertial frame.
+The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of a unified space-time model.
+In special relativity, the absolute space and the absolute time of Newton are fused into absolute space-time.
+The proper time interval may be identified with particle lifetime, atomic transition periods, and nuclear half-life.
+Until now, all tests of time dilation with atomic clocks in the earth-based frame have involved circle path geometry.
 Indeed, the clocks flown in the commercial jets.
 When the clocks returned to the laboratory they were compared with the reference clock in the same point at rest with respect to the earth.
-According to the special relativity, in any time dilation experiment which probe the influence of the orbital velocity one must use the linear (one-way) path geometry.
+According to special relativity, in any time dilation experiment that probes the influence of the orbital velocity, one must use the linear (one-way) path geometry.
 These are experiments of decays of rapidly moving unstable particles.
-However, the modern experimental techniques are not sensitive enough to register the effect of orbital velocity on the average distance traveled by a disintegrated particle.
-In contrast, we describe (see Chapter 8) a simple first order optical experiment within the means of most undergraduate departments that illustrates many of the principles of current research in the foundations of special relativity.
-It directly demonstrates that one could determine the speed of the earth with respect to the sun-based frame by means of aberration of earth-based incoherent source measurements.
-The second part of the book (Chapter 13 - 18) deals with accelerator physics.
-In the Chapter 13 we describe the particle dynamics method in the Lorentz lab frame using lab time $t$ as evolution parameter.
-The study of relativistic particle motion in a constant magnetic field according to usual accelerator engineering, is intimately connected with the old (Newtonian) kinematics: the Galilean vectorial law of addition of velocities is actually used.
-A non-covariant approach to relativistic particle dynamics is based on the absolute time coordinatization, but this is actually a hidden coordinatization.
+However, modern experimental techniques are not sensitive enough to register the effect of orbital velocity on the average distance traveled by a disintegrated particle.
+In contrast, we describe (see Chapter 10) a simple first-order optical experiment within the means of most undergraduate departments that illustrates many of the principles of current research in the foundations of special relativity.
+It directly demonstrates that one could determine the speed of the earth with respect to the sun-based frame using aberration of earth-based incoherent source measurements.
+In Chapter 16 we describe the particle dynamics method in the Lorentz lab frame using lab time $t$ as the evolution parameter.
+The study of relativistic particle motion in a constant magnetic field according to usual accelerator engineering, is intimately connected with the old (Newtonian) kinematics: the Galilean vectorial law of addition of velocities is used.
+A non-covariant approach to relativistic particle dynamics is based on absolute time coordinatization, but this is a hidden coordinatization.
 The absolute time synchronization convention is self-evident and this is the reason why this subject is not discussed in accelerator physics.
 According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity.
-In other words, in this standard approach, introducing as only modification to the classical case the relativistic mass, time differ from space.
-In convention particle tracking a particle trajectory can be seen from the lab frame as result of successive Galileo boosts that track the motion of the accelerated particle.
+In other words, in this standard approach, introducing only modification to the classical case of the relativistic mass, the time differs from space.
+In conventional particle tracking a particle trajectory can be seen from the lab frame as a result of successive Galileo boosts that track the motion of the accelerated particle.
 There is a reason to prefer the non-covariant way within the framework of dynamics only.
-In this approach we have no mixture of positions and time.
+In this approach, we have no mixture of positions and time.
 This (3+1) dimensional non-covariant particle tracking method is simple, self-evident, and adequate to the laboratory reality.
-It can be demonstrated that there is no principle difficulty with the non-covariant approach in mechanics and electrodynamics.
+It can be demonstrated that there is no principal difficulty with the non-covariant approach in mechanics and electrodynamics.
 It is perfectly satisfactory.
 It does not matter which transformation is used to describe the same reality.
-What matter is that, once fixed, such convention should be applied and kept in a consistent way for both dynamics and electrodynamics.
-In fact, the light (group) velocity transforms as a particle velocity.
-We cannot take old Newton’s kinematics for mechanics and Einstein’s kinematics for electrodynamics.
+What matters is that once fixed, such a convention should be applied and kept in a consistent way for both dynamics and electrodynamics.
 The common mistake made in accelerator physics is connected with the incorrect algorithm for solving the electromagnetic field equations.
-If one wants to use the usual Maxwell’s equations, only the solution of the dynamics equations in covariant form (i.e.
+If one wants to use the Maxwell’s equations, only the solution of the dynamics equations in covariant form (i.e.
 in Lorentz coordinates) gives the correct coupling between the Maxwell’s equations and particle trajectories in the lab frame.
-It is interesting to note that, the use of the conventional coupling of Maxwell’s equations and corrected Newton’s equation for the calculation of the radiation from a moving source does not necessarily leads to mistake.
-For rectilinear motion of the source and the emitted light beam, the non-covariant and covariant approaches produce the same trajectories, and Maxwell’ equations are compatible with the result of conventional particle tracking.
+It is interesting to note that, the use of the conventional coupling of Maxwell’s equations and corrected Newton’s equation for the calculation of the radiation from a moving source does not necessarily lead to a mistake.
+For the rectilinear motion of the source and the emitted light beam, the non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking.
 This method was incorrectly extended to the case of non-collinear geometry.
 First, we examine the reasoning presented in textbooks.
-It is generally believed that the magnetic field, in the Lorentz coordinatization, is only capable altering the direction of motion, but not the speed of the electron.
-Nevertheless, there is argument against this commonly accepted derivation of the composition of velocities.
-Let us to consider composition of two velocities when they are not lying on a straight line.
-In conventional particle tracking (similar to the Newton mechanics), one composes them geometrically by the method of parallelogram with velocities regarded as the usual vectors.
-Only solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle trajectories in the lab frame.
+It is generally believed that the magnetic field, in the Lorentz coordinatization, is only capable of altering the direction of motion, but not the speed of the electron.
+Nevertheless, there is an argument against this commonly accepted derivation of the composition of velocities.
+Let us consider the composition of two velocities when they are not lying on a straight line.
+In conventional particle tracking (similar to Newton’s mechanics), one composes them geometrically by the method of a parallelogram with velocities regarded as the usual vectors.
+Only the solution of the dynamics equations in covariant form gives the correct coupling between the Maxwell’s equations and particle trajectories in the lab frame.
 There are two difficulties in solving problems concerning covariant particle tracking.
-First of all, conventional particle tracking in the single inertial frame is actually based on classical Newton mechanics without reference to Lorentz transformations.
+First of all, conventional particle tracking in the single inertial frame is based on classical Newton mechanics without reference to Lorentz transformations.
 Another difficulty is the confusion surrounding the combination of non-parallel velocities in special relativity.
 The standard textbook presentation of the relativistic velocity transformation is based on the hidden assumption that $(x^{\prime},y^{\prime},z^{\prime})$ axes of the moving observer and $(x,y,z)$ axes of the lab observer are parallel.
-In other words, it is assumed that observers have common 3-space.
-This is misconception.
+In other words, it is assumed that observers have a common 3-space.
+This is a misconception.
 The fact that two observers with different trajectories in the Lorentz coordinatization have different 3-spaces is not considered in textbooks.
 First of all, there is no absolute notion of simultaneity in the theory of special relativity.
-In the case of the relativity of simultaneity we have a mixture - of positions and time.
-In other words, in the space measurement of one observer there is mixing of space and time, as seen by the other.
+In the case of the relativity of simultaneity, we have a mixture - of positions and time.
+In other words, in the space measurement of one observer, there is the mixing of space and time, as seen by the other.
 Therefore, there is no notion of an instantaneous 3-space.
 In textbooks on special relativity, the relativistic composition of velocities is studied analytically from the composition of Lorentz boosts.
-Authors of textbooks believe that the relativity principle requires that velocity addition law gives the composition law of a group.
+Authors of textbooks believe that the relativity principle requires that the velocity addition law gives the composition law of a group.
 The non-associativity of the composition law of the non-colinear velocities is not widely known, although it is an exercise in elementary algebra to prove non-associativity.
-It is therefore impossible to get covariant trajectory, if we have nothing but Lorentz boosts.
+It is therefore impossible to get a covariant trajectory if we have nothing but Lorentz boosts.
 Something else has to be added.
-Lorentz boosts do not form a group due to presence of the Wigner rotation.
-It is comes out quite natural that breakdown of commutativity and associativity in Einstein addition, turn out to be Wigner rotation.
+Lorentz boosts do not form a group due to the presence of the Wigner rotation.
+It comes out quite natural that the breakdown of commutativity and associativity in Einstein’s addition, turns out to be Wigner rotation.
 We emphasize that non-collinear velocities addition in the Lorentz coordinatization is regulated by the Wigner rotation.
@@
 Our result is at odds with the prediction from textbooks.
-The commonly accepted derivation of the composition of perpendicular velocities based on the use the relation $dt^{\prime}=dt/\gamma$ and does not account for the Wigner rotation (i.e.
+The commonly accepted derivation of the composition of perpendicular velocities is based on the use of the relation $dt^{\prime}=dt/\gamma$ and does not account for the Wigner rotation (i.e.
 mixture of transverse positions and time).
-Only the solution of the dynamics equations in covariant form (accounting for the Wigner rotation in the transformation of non-collinear velocities) gives the correct coupling between the Maxwell’s equations and particles trajectories in the lab frame.
-We illustrate error in standard coupling fields and particles by considering the relatively simple example, wherein the essential physical features are not obscured by unnecessary mathematical difficulties.
-We hope that the present example will draw wide attention to the central role that Wigner rotation plays in the transformation of non-collinear velocities.
-In the Chapter 13 we examine what parts of the dynamics theory depend on the choice of synchronization convention and what parts do not.
+Only the solution of the dynamics equations in covariant form (accounting for the Wigner rotation in the transformation of non-collinear velocities) gives the correct coupling between Maxwell’s equations and particle trajectories in the lab frame.
+In Chapter 16 we examine what parts of the dynamics theory depend on the choice of synchronization convention and what parts do not.
 The present analysis shows the difference between the notions of path and trajectory.
@@
 The trajectory of a particle conveys more information about its motion because every position is described additionally by the corresponding time instant.
-The path has exact objective meaning.
-In fact, the curvature radius of the path in the magnetic field (and consequently the three-momentum) has obviously an objective meaning, i.e.
+The path has an exact objective meaning.
+The curvature radius of the path in the magnetic field (and consequently the three-momentum) has an objective meaning, i.e.
 is convention-invariant.
-In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle is convention dependent and has no exact objective meaning.
-Dynamics theory contains a particle trajectory that we do not need to check directly, but which is used in the analysis of electrodynamics problem.
-In the Chapter 14 we present a critical reexamination of radiation theory.
-To evaluate radiation fields arising from an external sources, we need to know the particle velocity $\vec{v}$ and the position $\vec{x}$ as a function of the lab frame time $t$.
-As discussed above, one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particles moving along covariant trajectory.
+In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle is convention-dependent and has no exact objective meaning.
+Dynamics theory contains a particle trajectory that we do not need to check directly, but which is used in the analysis of electrodynamics problems.
+In Chapter 18 we present a critical reexamination of radiation theory.
+To evaluate radiation fields arising from external sources, we need to know the particle velocity $\vec{v}$ and the position $\vec{x}$ as a function of the lab frame time $t$.
+As discussed above, one should solve the Maxwell’s equations in the lab frame with current and charge density created by particles moving along the covariant trajectory.
 For an arbitrary parameter $v/c$ covariant calculations of the radiation process is very difficult.
 There are, however, circumstances in which calculations can be greatly simplified.
-As example of such circumstance is a non-relativistic radiation setup.
+An example of such a circumstance is a non-relativistic radiation setup.
 The reason is that the non-relativistic assumption implies the dipole approximation which is of great practical significance.
@@
 But that is only the first and most practically important term.
-The calculation of higher order corrections to the dipole radiation approximation requires detailed information about the electron trajectory.
-Obviously, in order to calculate the correction to the dipole radiation, we will have to use the covariant trajectory and not be satisfied with the non-covariant approach.
-Accelerator physicists, who try to understand the situation related to the use of the theory of relativity in the synchrotron radiation phenomena, are often troubled by the fact that the difference between covariant and non-covariant particle trajectories was never understood, and that nobody realized that there was a contribution to the synchrotron radiation from relativistic kinematics effects.
-Accelerator physics was always thought in terms of the old (Newtonian) kinematics that is not compatible with Maxwell’s equations.
+The calculation of higher-order corrections to the dipole radiation approximation requires detailed information about the electron trajectory.
+In order to calculate the correction to the dipole radiation, we will have to use the covariant trajectory and not be satisfied with the non-covariant approach.
+It is generally assumed that the Maxwell’s equations and corrected Newton’s second law can explain all experiments that are performed in a single inertial (lab) frame.
+According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity.
+Many physicists have already received knowledge about special relativity from textbooks and who would say, ”For those who want to learn just enough about special relativity so they can solve problems - it just changes Newton’s laws by introducing a correction factor to the mass.” This is a misconception.
+The light velocity transforms as a particle velocity.
+We cannot take old kinematics for mechanics and Einstein’s kinematics for electrodynamics.
+We illustrate errors in standard coupling fields and particles by considering the relatively simple example, wherein the essential physical features are not obscured by unnecessary mathematical difficulties.
+We hope that the present example will draw wide attention to the central role that Wigner rotation plays in the transformation of non-collinear velocities.
+Accelerator physicists, who try to understand the situation related to the use of the theory of relativity in the synchrotron radiation phenomena, are often troubled by the fact that the difference between covariant and non-covariant particle trajectories was never understood and that nobody realized that there was a contribution to the synchrotron radiation from relativistic kinematics effects.
+Accelerator physics was always thought of in terms of the old (Newtonian) kinematics that is not compatible with Maxwell’s equations.
 At this point, a reasonable question arises: since storage rings are designed without accounting for the relativistic kinematics effects, how can they actually operate?
-In fact, electron dynamics in storage ring is greatly influenced by the emission of radiation.
-We answer this question in great detail in the Chapter 15.
+In fact, electron dynamics in the storage ring is greatly influenced by the emission of radiation.
+We answer this question in great detail in Chapter 19.
 The main emphasis of this chapter is on spontaneous synchrotron radiation from bending magnets and undulators.
@@
 There are a number of remarkable effects which are a consequence of the cylindrical symmetry and the paraxial approximation.
-We demonstrate that there is no needs to use covariant particle tracking for derivation of the bending magnet radiation.
+We demonstrate that there is no need to use covariant particle tracking for derivation of the bending magnet radiation.
 However, there is one situation where the conventional theory fails.
@@
 One way to demonstrate incompatibility between the standard approach to relativistic electrodynamics, which deals with the usual Maxwell’s equations, and particle trajectories calculated by using non-covariant particle tracking, is to make a direct laboratory test of synchrotron radiation theory.
-In other words, we are stating that, despite the many measurements done during decades, synchrotron radiation theory is not an experimentally well-confirmed theory.
-We analyze the potential for exploiting synchrotron radiation sources in order to confirm the predictions of corrected synchrotron radiation theory.
+In other words, we are stating that, despite the many measurements done over decades, synchrotron radiation theory is not an experimentally well-confirmed theory.
+We analyze the potential for exploiting synchrotron radiation sources to confirm the predictions of corrected synchrotron radiation theory.
 Proposed experiments have never been performed at synchrotron facilities.
-The theory of relativity shows that the relativity of simultaneity, which is a most fundamental relativistic kinematics effect, is related with extended relativistic objects.
-But up to 21 st century there were no macroscopic objects possessing relativistic velocities, and there was a general belief that only microscopic particles in experiments can travel at velocities close to that of light.
+The theory of relativity shows that the relativity of simultaneity, which is a most fundamental relativistic kinematics effect, is related to extended relativistic objects.
+But up to the 21 st century, there were no macroscopic objects possessing relativistic velocities, and there was a general belief that only microscopic particles in experiments could travel at velocities close to that of light.
 The 2010s saw a rapid development of new laser light sources in the X-ray wavelength range.
-An X-ray free electron laser (XFEL) is an example where improvements in accelerator technology makes it possible to develop ultrarelativistic macroscopic objects with an internal structure (modulated electron bunches), and the relativistic kinematics plays an essential role in their description.
-In Chapter 16 we present a critical reexamination of existing XFEL theory.
-It is mainly addressed to readers with limiting knowledge of accelerator physics.
-We study the production of coherent undulator radiation by modulated ultrarelativistic electron beam.
-Fortunately, the principle of production of coherent undulator radiation does not require specific knowledge of undulator radiation theory presented in the Chapter 15 and can be explained in a very simple way.
-Relativistic kinematics enters XFEL physics in a most fundamental way through the rotation of the modulation wavefront, which, in ultrarelativistic approximation, is closely associated to the relativity of simultaneity.
-When the trajectories of particles calculated in the Lorentz reference frame (i.e.
-an inertial frame where Einstein synchronization procedure is used to assign values to the time coordinate) they must include relativistic kinematics effects such as relativity of simultaneity.
-In the ultrarelativistic asymptote, the orientation of the modulation wavefront, i.e the orientation of the plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the modulated electron beam is treated using Lorentz coordinates.
+An X-ray free electron laser (XFEL) is an example where improvements in accelerator technology make it possible to develop ultrarelativistic macroscopic objects with an internal structure (modulated electron bunches), and the relativistic kinematics plays an essential role in their description.
+In Chapter 20 we present a critical reexamination of existing XFEL theory.
+It is mainly addressed to readers with limited knowledge of accelerator physics.
+We study the production of coherent undulator radiation by the modulated ultrarelativistic electron beam.
+Fortunately, the principle of production of coherent undulator radiation does not require specific knowledge of the undulator radiation theory presented in Chapter 19 and can be explained in a very simple way.
+Relativistic kinematics enters XFEL physics in a most fundamental way through the rotation of the modulation wavefront, which, in ultrarelativistic approximation, is closely associated with the relativity of simultaneity.
+When the trajectories of particles are calculated in the Lorentz reference frame (i.e.
+an inertial frame where the Einstein synchronization procedure is used to assign values to the time coordinate) they must include relativistic kinematics effects such as relativity of simultaneity.
+In the ultrarelativistic asymptote, the orientation of the modulation wavefront, i.e.
+the orientation of the plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the modulated electron beam is treated using Lorentz coordinates.
 We should remark that Maxwell’s equations are valid only in Lorentz reference frames.
@@
 The theory of relativity as a theory of space-time with pseudo-Euclidean geometry has had more than hundred years of development, and rather suddenly it has begun to be fully exploited in practical ways in XFEL physics.
-As known, a composition of noncollinear Lorentz boosts does not result in a different boost but in a Lorentz transformation involving a boost and a spatial rotation, the Wigner rotation (it is often called the ”Thomas rotation”).
-The results for the Wigner rotation in the Lorentz lab frame obtained by many experts on special relativity are incorrect.
-They overestimate the angle of the Wigner rotation by a relativistic factor $\gamma$ compared to its real value, and the direction of the rotation is also determined incorrectly.
-It should be note that the correct result was obtained in the works of several authors, which were published more than half century ago but remained unnoticed against the background of numerous incorrect works.
-To deduce the expression for Wigner rotation is the main purpose of Chapter 17.
-We demonstrate that what are usually considered the advanced parts of the theory of relativity are, in fact, quite simple.
-The Wigner rotation is also relates the angular rotation velocity (Thomas precession) of the spin of an elementary particles following a curvalinear orbit.
-The kinematics tools needed to study the motion of charged elementary spinning particles in the storage ring are the main topic of final chapter of the book.
-In 1959, Bargman, Michel, and Telegdi (BMT) proposed a consistent relativistic theory for the dynamics of the spin as observed in the lab frame, which was successfully tested in experiments.
-It is commonly believed that the BMT equation contains the standard (and incorrect) result for the Wigner rotation in the Lorentz lab frame.
-There are many physicists who have already received knowledge about the Wigner (Thomas) rotation from well-known textbooks and who would say, ”The extremely precise measurements of the magnetic-moment anomaly of the electron made on highly relativistic electrons are based on the BMT equation, of which the Wigner rotation is an integral part, and can be taken as experimental confirmations of the standard expression for the Wigner rotation” This misconception about experimental test of the incorrect expression for the Wigner rotation in the lab frame is common and pernicious.
-We should remark that the results in the Bargmann-Michel-Telegdi paper were obtained by the method of semi-classical approximation of the Dirac equation.
-The Wigner rotation was not considered at all, because the Dirac equation allow obtaining the solution for the total particle’s spin motion without an explicit splitting it into the Larmor and Wigner parts.
-We demonstrate that the notion that the standard (incorrect) result for the Wigner rotation as an integral part of the BMT equation in most texts is based, in turn, on an incorrect physical argument.
-A large number of incorrect expressions for the Wigner (Thomas) rotation can be found in the literature.
-Let us discuss the reasons that underlie the derivation of the incorrect expressions.
-The situation relating to the use of the Wigner (Thomas) rotation theory in the relativistic spin dynamics is complex.
-The starting point of Bargman, Michel and Telegdi was the particle rest frame and the equation of motion for particle angular momentum, which they generalized to the Lorentz lab frame and then transformed back to the rest frame.
-This brings up an interesting question: Why it is convenient to transform this equation to the rest frame as of that instant?
-The approach in which we deal with the proper spin is much preferred in the experimental practice due to clear physical meaning of the (3D) spin vector $\vec{s}$.
-Unlike (3D) momentum vector, which has definite components in each reference frame, angular momentum is defined only in one particular reference frame.
-It does not transform.
-Any statement about it refers to the rest frame as of that instant.
-If we say that in the lab frame the spin of a particle makes the angle $\phi$ with its velocity, we mean that in the particle’s rest frame the spin makes this angle with the line motion of the lab frame.
-The analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard (Einstein) special relativity taking advantage of well known relativistic kinematic effects.
-In the case of spin dynamics one has no occasion to make measurements with non-inertial devices.
-Common textbooks presentations of the spin dynamics use the approach which deals only with observations of the inertial (lab) observer.
-The conventional method used to explain the spin dynamics in the lab frame is very unusual.
-First, we already pointed that the laws of physics in the lab frame are able to account for all physical phenomena, including the observation made by non-inertial observer.
-In particular, it is possible to determine the result of spin orientation measurements with respect to the lab frame axes in the proper frame.
-In this case, when viewed from the lab frame, the interpretation of the spin rotation experiments in the lab frame is simple and may be derived from the prediction of the proper observer as concerns the measurement made by the lab observer.
-Therefore, the spin rotation measurement by a polarimeter in the lab frame is interpreted with viewpoint of the proper observer as viewing this of the lab observer.
-This complexity is one of the reason why authors of textbooks obtained an incorrect expression for the Wigner rotation.
-It is also important to elucidate the reason why did the error in the Wigner rotation theory remain so long undetected.
-The main conclusion drawn from the review of the literature is as follows.
-Wigner (Thomas) rotation is generally presented only in the context of spin kinematics as a peculiar effect of special relativity.
-This is supposedly the reason why until recently no researches have noticed so serious a discrepancy between the results of different works.
-We emphasize that the Wigner rotation is very basic relativistic phenomenon.
-It is as basic as the time dilation and length contraction.
-For example, we demonstrate by elementary means that relativistic (non-collinear) velocities addition is regulated by the Wigner rotation.
-Now let us consider the aberration theory.
-The theory of special relativity shows that the aberration effect in the rotating earth-based frame, as viewed from the inertial sun-based frame (in the Lorentz coordinatization), presents a kinematic effect of special theory of relativity.
-In this case, when viewed from the sun-based frame, the aberration image shift in the earth-based frame is regulated by the Wigner rotation.
-According to the Wigner rotation theory, two observers with different trajectories have different 3-spaces.
-In contrast, the accepted in previous literature incorrect assumption that an earth-based observer and an inertial (e.g.
-sun-based) observer have common 3-space has always been considered obvious.
 2 A Critical Survey of the Present Approaches to Special Relativity
@@
 This is a restrictive principle and does not determine the exact form of the dynamics in question.
-Understanding the postulates of the theory of relativity is similar to understanding energy conservation: at first we learn this as a principle and later on we study microscopic interpretations that must be consistent with this principle.
-For any system to which the energy conservation principle can be applied, a deeper theory should exist which yields insight into the detailed physical processes involved.
+Understanding the postulates of the theory of relativity is similar to understanding energy conservation: at first, we learn this as a principle, and later on, we study microscopic interpretations that must be consistent with this principle.
+For any system to which the energy conservation principle can be applied, a deeper theory should exist that yields insight into the detailed physical processes involved.
 Of course, this deeper theory must lead to energy conservation.
 The principle of conservation of energy is very useful in making analyses without knowing all the formulas of the fundamental theory.
-A methodological analogy with the postulates of the special relativity emerges by itself.
+A methodological analogy with the postulates of special relativity emerges by itself.
 Suppose we do not know why a muon disintegrates, but we know the law of decay in the Lorentz rest frame.
@@
 In particular, when a Lorentz transformation of the decay law is tried, one obtains the prediction that after the travel distance $\gamma v\tau_{0}$, the population in the lab frame would be reduced to 1/2 of the origin population.
+Here $\gamma$ is the relativistic factor and $\tau_{0}$ is the proper lifetime.
 We may interpret this result by saying that, in the lab frame, the characteristic lifetime of a particle has increased from $\tau_{0}$ to $\gamma\tau_{0}$.
-However, the theory of relativity is necessary incomplete.
+However, the theory of relativity is necessarily incomplete.
 Constructive (microscopic) theories like electrodynamics or quantum field theory provide more insight into the nature of things than restrictive theories like special relativity.
@@
 In literature, three approaches to special relativity are discussed: Einstein’s approach, the usual covariant approach, and the space-time geometric approach.
-Einstein formulation is based on postulates: the principle of relativity and the constancy of the velocity of light.
-The usual covariant formulation of the theory of relativity deals with the pseudo-Eucledian space-time geometry and with the invariance of interval $ds$, but it is understood only in a limited sense when the metric is strictly diagonal.
-Assuming diagonality of the metric we also automatically assume Lorentz coordinates, and that different inertial frames are related by Lorentz transformations.
-In space-time geometric approach, primary importance is attributed to the geometry of space-time; it is supposed that the geometry of space-time is a pseudo-Euclidean geometry in which only 4-tensors quantities do have real physical meaning.
-In this most general approach the principle of relativity in contrast to Einstein formulation of the special relativity is a simple consequence of the space-time geometry.
+Einstein’s formulation is based on postulates: the principle of relativity and the constancy of the velocity of light.
+The usual covariant formulation of the theory of relativity deals with the pseudo-Euclidean space-time geometry and with the invariance of interval $ds$, but it is understood only in a limited sense when the metric is strictly diagonal.
+Assuming the diagonality of the metric we also automatically assume Lorentz coordinates, and that different inertial frames are related by Lorentz transformations.
+In the space-time geometric approach, primary importance is attributed to the geometry of space-time; it is supposed that the geometry of space-time is a pseudo-Euclidean geometry in which only 4-tensors quantities do have real physical meaning.
+In this most general approach the principle of relativity in contrast to Einstein’s formulation of special relativity is a simple consequence of space-time geometry.
 Since the space-time geometric approach deals with all possible choices of coordinates of the chosen reference frames, the second Einstein postulate referred to the constancy of the coordinate velocity of light does not hold in this formulation of the theory of relativity.
@@
 The central principle of special relativity is the Lorentz covariance of all the fundamental laws of physics.
-It it important to stress at this point that the second ”postulate”, contrary to the view presented in textbooks, is not a separate physical assumption, but a convention that cannot be the subject of experimental tests.
+It is important to stress at this point that the second ”postulate”, contrary to the view presented in textbooks, is not a separate physical assumption, but a convention that cannot be the subject of experimental tests.
 Assuming postulate 2 on the constancy of the speed of light in all inertial frames we also automatically assume Lorentz coordinates, and that different inertial frames are related by Lorentz transformations.
-In other words, according to such limiting understanding of the theory of relativity it is assumed that only Lorentz transformations must be used to map the coordinates of events between inertial observers.
+In other words, according to such a limiting understanding of the theory of relativity, it is assumed that only Lorentz transformations must be used to map the coordinates of events between inertial observers.
 2.2.2 The Usual Covariant Approach
-In the usual covariant approach the special of relativity is understood as the theory of space-time with pseudo-Euclidean geometry.
-Quantities of physical interest are represented by tensors in a four-dimensional space-time, i.e.
-by covariant quantities, and the laws of physics are written in manifestly covariant way as four-tensor equations.
+In the usual covariant approach, the special of relativity is understood as the theory of space-time with pseudo-Euclidean geometry.
+Quantities of physical interest are represented by tensors in four-dimensional space-time, i.e.
+by covariant quantities, and the laws of physics are written in a manifestly covariant way as four-tensor equations.
 In order to develop space-time geometry, it is necessary to introduce a metric or a measure $ds$ of space-time intervals.
@@
 Here $\eta_{ij}$ are the Cartesian components of the metric tensor and by definition, in any Lorentz system, they are given by $\eta_{ij}=\mathrm{diag}[1,-1,-1,-1]$, which is the metric canonical, diagonal form.
-As a consequence of the space-time geometry, Lorentz coordinates systems are connected by Lorentz transformations.
+As a consequence of the space-time geometry, Lorentz coordinate systems are connected by Lorentz transformations.
 Physical quantities are represented by space-time geometric (tensor) quantities.
-When some basis is introduced, the representation of a tensor as geometric quantity comprise both components and basis.
-In usual covariant approach, one only deals with the basis components of tensors in the Lorentz coordinates i.e.
+When some basis is introduced, the representation of a tensor as a geometric quantity comprises both components and basis.
+In the usual covariant approach, one only deals with the basis components of tensors in the Lorentz coordinates i.e.
 with the case when the basis four-vectors are orthogonal.
-As a result one deals only with four-tensor equations of physics written out in the component form.
+As a result, one deals only with four-tensor equations of physics written out in the component form.
 However, the concept of a tensor in the usual covariant approach is given in terms of the transformation properties of its components.
@@
 only assuming that Lorentz transformations must be used to map the coordinates of events.
-According to the usual covariant approach, another transformations from standard to non standard (not orthogonal) basis, like Galilean transformations, are ”incorrect”.
+According to the usual covariant approach, other transformations from standard to nonstandard (not orthogonal) basis, like Galilean transformations, are ”incorrect”.
 2.2.3 The Space-Time Geometric Approach
@@
 (1) can be described in arbitrary coordinates and not only in Lorentz coordinates.
-In the transition to arbitrary coordinates, the geometry of four-dimensional space-time obviously does not change, and in the special theory of relativity we are not limited in any way in the choice of a coordinates system.
+In the transition to arbitrary coordinates, the geometry of four-dimensional space-time does not change, and in the special theory of relativity, we are not limited in any way in the choice of a coordinates system.
 The space coordinates $x^{1},x^{2},x^{3}$ can be any quantities defining the position of particles in space, and the time coordinate $x^{0}$ can be defined by an arbitrary running clock.
@@
 However, in the space-time geometric approach, special relativity is understood as a theory of four-dimensional space-time with pseudo-Euclidean geometry.
-In this formulation of the theory of relativity the space-time continuum can be described equally well from the point of view of any coordinate system, which cannot possibly change $ds$.
+In this formulation of the theory of relativity, the space-time continuum can be described equally well from the point of view of any coordinate system, which cannot possibly change $ds$.
 At variance, the usual formulation of the theory of relativity also deals with the invariance of $ds$, but it is understood only in a limited sense when the metric is strictly diagonal.
@@
 The fact that in the process of transition to arbitrary coordinates the geometry of the space-time does not change, is not considered in textbooks.
-As a consequence there is a widespread belief among experts that a transformation from an orthogonal Lorentz basis to a non orthogonal basis is incorrect, while a Lorentz transformation, which is a transformation from an orthogonal Lorentz basis to another orthogonal Lorentz basis, is correct.
+As a consequence, there is a widespread belief among experts that a transformation from an orthogonal Lorentz basis to a nonorthogonal basis is incorrect, while a Lorentz transformation, which is a transformation from an orthogonal Lorentz basis to another orthogonal Lorentz basis, is correct.
 This is not true.
-We can describe physics in any arbitrary coordinates system.
+We can describe physics in any arbitrary coordinate system.
 The different transformations of coordinates only correspond to a change in the way of components of 4-tensors are written, but not influence of 4-tensors themselves.
@@
 Lorentz coordinates choice, is preferred by physicists due to its simplicity and symmetry, it is nothing more ”physical” than any other.
-A particularly very unusual choice of coordinates, the absolute time coordinate choice, will be considered and exploited in this book.
+A particularly very unusual choice of coordinates, the absolute time coordinate choice, will be considered and exploited in this article.
 2.3 Myth About the Incorrectness of Galilean Transformations
 The use of Galilean transformations within the theory of relativity requires some special discussion.
-Many physicists still tend to think of Galilean transformations (which is actually a transformations from an orthogonal Lorentz basis to a non-orthogonal basis) as old, incorrect transformations between spatial coordinates and time.
+Many physicists still tend to think of Galilean transformations (which are transformations from an orthogonal Lorentz basis to a non-orthogonal basis) as old, incorrect transformations between spatial coordinates and time.
 A widespread argument used to support the incorrectness of Galilean transformations is that they do not preserve the form-invariance of Maxwell’s equations under a change of inertial frame.
-This idea is a part of the material in well-known books and monographs (1).
+This idea is a part of the material in well-known books and monographs 111 Many physicists tend to think of Galilean transformations as pre-relativistic transformations between spatial coordinates and time that are not compatible with the special theory of relativity.
+To quote e.g.
+Bohm [1] ”… the Galilean law of addition of velocities implies that the speed of light should vary with the speed of the observing equipment.
+Since this predicted variation is contrary to the fact, the Galilean transformations evidently cannot be the correct one.”.
+Similar statements can also be found in recently published pedagogical papers.
+To quote e.g.
+Drake and Purvis [2] ”One of the great insights to come relativity theory was the realization that Galilean transformations are wrong.
+The correct way to translate the space-time measure of events between inertial frames is with the Lorentz transformations”..
 Authors of textbooks are mistaken in their belief about the incorrectness of Galilean transformations.
 The special theory of relativity is the theory of four-dimensional space-time with pseudo-Euclidean geometry.
-From this viewpoint, the principle of relativity is a simple consequence of the space-time geometry, and the space-time continuum can be described in arbitrary coordinates (2).
-Therefore, contrary to the view presented in most textbooks, Galilean transformations are actually compatible with the principle of relativity although, of course, they alter the form of Maxwell’s equations.
+From this viewpoint, the principle of relativity is a simple consequence of the space-time geometry, and the space-time continuum can be described in arbitrary coordinates  [^2] The mathematical argument that in the process of transition to arbitrary coordinates the geometry of the space-time does not change, is considered in textbooks as erroneous.
+To quote L.
+Landau and E.
+Lifshitz [3]: ”This formula is called the Galileo transformation.
+It is easy to verify that this transformation, as was to be expected, does not satisfy the requirements of the theory of relativity; it does not leave the interval between events invariant.”.
+This fact is ascribed to a lack of understanding of the difference between convention-dependent and convention-invariant parts of the theory.
+In pseudo-Euclidean geometry the interval between events is an invariant in arbitrary coordinates..
+Therefore, contrary to the view presented in most textbooks, Galilean transformations are compatible with the principle of relativity although, of course, they alter the form of Maxwell’s equations.
 2.4 Myth About the Non-Relativistic Limit of the Lorentz Transformations
@@
 We state that this typical textbook statement is incorrect and misleading.
-Kinematics is a comparative study which requires two coordinate systems, and one needs to assign time coordinates to the two systems.
+Kinematics is a comparative study that requires two coordinate systems, and one needs to assign time coordinates to the two systems.
 Different types of clock synchronization provide different time coordinates.
@@
 This infinitesimal Lorentz transformation differs from the infinitesimal Galilean transformation $x^{\prime}=x-vt$, $t^{\prime}=t$.
-The difference is in the term $xv/c^{2}$ in the Lorentz transformation for time, which is a first order term.
+The difference is in the term $xv/c^{2}$ in the Lorentz transformation for time, which is a first-order term.
 We only wish to emphasize here the following point.
 An infinitesimal Lorentz transformation differs from Galilean transformation only by the inclusion of the relativity of simultaneity, which is the only relativistic effect that appears in the first order in $v/c$.
-All other higher order effects, that are Lorentz-Fitzgerald contraction, time dilation, and relativistic correction in the law of composition of velocities, can be derived mathematically, by iterating this infinitesimal transformation (3).
-The main difference between the Lorentz coordinatization and the absolute time coordinatization is that the transformation laws connecting coordinates and times between relatively moving systems are different.
-It is impossible to agree with the textbook statement that there is reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to Galilean relation $t^{\prime}=t$ in the non-relativistic limit.
-This would mean that in the non-relativistic limit infinitesimal Lorentz transformations are identical to infinitesimal Galilean transformations.
-This statement is absurd conclusion from a mathematical standpoint.
+All other higher order effects, that are Lorentz-Fitzgerald contraction, time dilation, and relativistic correction in the law of composition of velocities, can be derived mathematically, by iterating this infinitesimal transformation 333 There is a widespread belief that a Lorentz transformation reduces to a Galilean transformation in the non-relativistic limit.
+To quote e.g.
+French [4]: ”The reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to Galilean relation $t^{\prime}=t$ requires $x\ll ct$ as well as $v/c\ll 1$”.
+Similar statements can also be found in recently published textbooks.
+To quote Rafelski [5]: ”A wealth of daily experience shows that the Galilean coordinate transformation (GT) is correct in the nonrelativistic limit in which the speed of light is so large that it plays no physical role.
+Any coordinate transformation replacing the GT must also agree with this experience, and thus must contain the GT in the nonrelativistic limit.” This statement is an absurd conclusion even from a mathematical standpoint.
+This has been recognized by some experts, perhaps most explicitly by Baierlein who states [6] that ”If the Lorentz transformation for infinitesimal $v/c$ were to reduce to the Galilean transformation, then the iterative process could never generate a finite Lorentz transformation that is radically different from the Galilean transformation.
+But the finite transformations are radically different, and so -however subtly-the infinitesimal Lorentz transformation must differ significantly from the Galilean transformation.” .
+The main difference between Lorentz coordinatization and absolute time coordinatization is that the transformation laws connecting coordinates and times between relatively moving systems are different.
+It is impossible to agree with the textbook statement that there is the reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to Galilean relation $t^{\prime}=t$ in the non-relativistic limit.
+This would mean that in the non-relativistic limit, infinitesimal Lorentz transformations are identical to infinitesimal Galilean transformations.
 The essence of Lorentz (or Galilean) transformations consists in their infinitesimal form: relativistic kinematics effects cannot be found by the mathematical procedure of iterating the infinitesimal Galilean transformations.
 2.5 Myth about the Constancy of the Speed of Light
-It is generally believed that experiments show that the speed of light in vacuum is independent of the source or observer (4).
-This statement presented in most textbooks and is obviously incorrect.
-The constancy of the speed of light is related to the choice of synchronization convention, and cannot be subject to experimental tests (5).
-In fact, in order to measure the one-way speed of light one has first to synchronize the infinity of clocks assumed attached to every position in space, which allows us to perform time measurements.
+It is generally believed that experiments show that the speed of light in a vacuum is independent of the source or observer [^4] Many authors of textbooks still attribute a measurable status to the conventional quantities.
+To quote e.g.
+Cristodoulides [7] ”The fact that Galilean transformation does not leave Maxwell’s equations has already been mentioned […] On the other hand, experiments show that the speed of light in a vacuum is independent of the source or observer.”.
+This statement is presented in most textbooks and is incorrect.
+The constancy of the speed of light is related to the choice of synchronization convention, and cannot be subject to experimental tests  [^5] Because we have empirical access only to the round-trip average speed of light, statements about the magnitude and isotropy of the one-way speed of light must reflect the assumptions made in the choice of time coordinatization, and such entities change as the theory is re-synchronized.
+To quote C.
+Anderson, I.
+Vetharaniam, and G.
+Stedman [8]: ”No experiment, then, is a ”one-way” experiment.
+An empirical test of any property of the one-way speed of light is not possible.
+Such quantities as the one-way speed of light are irreducibly conventional in nature, and recognizing this aspect is to recognize a profound feature of nature”..
+To measure the one-way speed of light one has first to synchronize the infinity of clocks assumed attached to every position in space, which allows us to perform time measurements.
 Obviously, an unavoidable deadlock appears if one synchronizes the clocks by assuming a-priori that the one-way speed of light is $c$.
@@
 In contrast to this, the two-way speed of light, directly measurable along a round-trip, has physical meaning, because round-trip experiments rely upon the observation of simultaneity or non-simultaneity of events at a single point in space and not depends on clock synchronization convention.
-All well known methods to measure the speed of light are, indeed, round-trip measurements.
+All well-known methods to measure the speed of light are, indeed, round-trip measurements.
 The cardinal example is given by the Michelson-Morley experiment: this experiment uses, indeed, an interferometer where light beams are compared in a two-way fashion.
-Because of our usage of Galilean transformations within electrodynamics we have some apparent paradoxes.
-One finds many books which say that a Galilean transformations of the velocity of light is not consistent with the electron-theoretical explanation of refraction and reflection (6).
+Because of our usage of Galilean transformations within electrodynamics, we have some apparent paradoxes.
+One finds many books that say that Galilean transformations of the velocity of light are not consistent with the electron-theoretical explanation of refraction and reflection  [^6] The peculiarity of the kinematic consequence of using Galilean transformations is that the speed of light emitted by a moving source depends on the relative velocity between the source and the observer.
+A widespread theoretical argument used to support the incorrectness of Galilean transformations is the conclusion that a Galilean transformation of the velocity of light is not consistent with the explanation of reflection and refraction.
+This idea is a part of the material in well-known books.
+To quote Pauli [9] ”[…] it is essential that the spherical waves emitted by the dipoles in the body should interfere with the incident wave.
+If we now think of the body as at rest, and the light source moving relative to it, then […] the wave emitted by the dipoles will have a velocity different from that of the incident wave.
+Interference is therefore not possible.”.
 It is widely accepted that if we consider a moving source and a stationary glass, the incident light wave and wave scattered by the dipoles of the glass cannot interfere as required by the electron theory of dispersion since their velocity are different.
-This is misconception.
+This is a misconception.
 It is clear that an incident wave with a certain frequency, no matter what its velocity, excites the electrons of a glass into oscillations of the same frequency.
@@
 Thus, the incident and scattered wave at any given point have the same frequency and can interfere.
-The effect of the different velocities is to produce a relative phase which varies with position in space.
+The effect of the different velocities is to produce a relative phase that varies with position in space.
 2.6 Convention-Dependent and Convention-Invariant Parts of the Theory
-Consider the motion of charged particle in a given magnetic field.
+Consider the motion of a charged particle in a given magnetic field.
 The theory of relativity says that the particle trajectory $\vec{x}(t)$ in the lab frame depends on the choice of a convention, namely the synchronization convention of clocks in the lab frame.
 Whenever we have a theory containing an arbitrary convention, we should examine what parts of the theory depend on the choice of that convention and what parts do not.
-We may call the former convention-dependent, and the latter convention-invariant parts.
+We may call the former convention-dependent and the latter convention-invariant parts.
 Clearly, physically meaningful measurement results must be convention-invariant.
@@
 Suppose there are two apertures at point $A$ and at point $A^{\prime}$.
-From the solution of the dynamics equation of motion we may conclude that the first particle gets through the aperture at $A$ and the second particle gets through the aperture at $A^{\prime}$ simultaneity.
+From the solution of the dynamics equation of motion, we may conclude that the first particle gets through the aperture at $A$ and the second particle gets through the aperture at $A^{\prime}$ simultaneity.
 The two events, i.e.
@@
 However, the simultaneity of these two events is convention-dependent and has no exact objective meaning.
-It is important at this point to emphasize that, consistently with the conventionality of simultaneity, also the value of the speed of particle is a matter of convention and has no definite objective meaning.
-In order to examine what parts of the dynamics theory depend on the choice of that convention and what parts do not, we want to show the difference between the notions of path and trajectory.
+It is important at this point to emphasize that, consistently with the conventionality of simultaneity, also the value of the speed of the particle is a matter of convention and has no definite objective meaning.
+To examine what parts of the dynamics theory depend on the choice of that convention and what parts do not, we want to show the difference between the notions of path and trajectory.
 Let us consider the motion of a particle in three-dimensional space using the vector-valued function $\vec{x}(t)$.
@@
 The motion along the path is described by $l(t)$, where $l$ is a certain parameter (in our case of interest the length of the arc).
-The trajectory of a particle conveys more information about its motion because every position is described additionally by the corresponding time instant.
-The path is rather a purely geometrical notion.
-If we take the origin of the (Cartesian) coordinate system and we connect the point to the point laying on the path and describing the motion of the particle, then the creating vector will be a position vector $\vec{x}(l)$ (7).
+If we take the origin of the (Cartesian) coordinate system and connect the point to the point laying on the path, then the creating vector will be a position vector $\vec{x}(l)$  [^7] For a general discussion of the difference between path and trajectory we suggest reading the book [10]..
 The difference between trajectory $\vec{x}(t)$ and path $\vec{x}(l)$ is very interesting.
@@
 it is convention-invariant.
-In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle is convention dependent and has no exact objective meaning.
-In order to avoid being to abstract for to long we have given some examples: just think of the experiments related with accelerator physics.
+In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle is convention-dependent and has no exact objective meaning.
+Let us give some examples: just think of the experiments related to accelerator physics.
 Suppose we want to perform a particle momentum measurement.
-A uniform magnetic field can be used in making a ”momentum analyzer” for high-energy charge particles, and it must be recognized that this method for determining the particle’s momentum is convention-independent.
-In fact, the curvature radius of the path in the magnetic field (and consequently the three-momentum) has obviously an objective meaning, i.e.
+A uniform magnetic field can be used in making a ”momentum analyzer” for high-energy charged particles, and it must be recognized that this method for determining the particle’s momentum is convention-independent.
+The curvature radius of the path in the magnetic field (and consequently the three-momentum) has an objective meaning, i.e.
 is convention-invariant.
-Dynamics theory contains a particle trajectory that we do not need to check directly, but which is used in the analysis of electrodynamics problem.
+Dynamics theory contains a particle trajectory that we do not need to check directly, but which is used in the analysis of electrodynamics problems.
 2.7 Myth about the Reality of Relativistic Time Dilation and Length Contraction
-Generally, experts on the theory of relativity erroneously identify the properties of Minkovski space-time with the familiar form that certain convention-dependent quantities assume under the standard Lorentz coordinatization.
+Generally, experts on the theory of relativity erroneously identify the properties of Minkowski space-time with the familiar form that certain convention-dependent quantities assume under the standard Lorentz coordinatization.
 These quantities usually are called ”relativistic kinematics effects”.
-There is a widespread belief that the convention-dependent quantities like the time dilation, length contraction, and Einstein’s addition of velocities have direct physical meaning.
-We found that statement like ” moving clocks run slow” is not true under the adopted absolute time clock synchronization, and, hence, are by no means intrinsic features of Minkowski space-time.
+There is a widespread belief that convention-dependent quantities like time dilation, length contraction, and Einstein’s addition of velocities have direct physical meaning.
 Relativistic kinematic effects are coordinate (i.e.
-convention-dependent) effects and have no exact objective meaning (8).
+convention-dependent) effects and have no exact objective meaning  [^8] The standard textbooks erroneously identify the properties of space-time with the familiar form that certain synchronization-dependent quantities assume under Einstein’s clock synchronization.
+These quantities are, for example, the formulae for ”time dilation” and ”length contraction”.
+Fortunately, it has also been occasionally stressed in the literature that the forms of the above ”relativistic effects” are coordinate-dependent, their forms depend in turn on the kind of synchronization procedure adopted.
+To quote Leubner, K.
+Auflinger, and P.
+Krumm [11]: ”… there is a widespread belief among students that the familiar form of coordinate-dependent quantities like the measured velocity of light, the Lorentz transformation between two observers, ”addition of velocities”, ”time dilation”, ”length contraction” which they assume under the standard clock synchronization, is relativity proper.
+To demonstrate that this is by no means so, this paper studies the consequences of a non-standard synchronization, and it is shown that drastic changes in the appearance of all these quantities are thus induced.” .
 In the case of Lorentz coordinatization, one will experience e.g.
 the time dilation phenomenon.
-In contrast to this, in the case of absolute time coordinatization there are no relativistic kinematics effects and no time dilation will be found.
+In contrast to this, in the case of absolute time coordinatization, there are no relativistic kinematics effects, and no time dilation will be found.
 However, all coordinate-independent quantities like the particle path $\vec{x}(l)$ and momentum $|\vec{p}|$ remain independent of such a change in clock synchronization.
 2.8 Relativistic Particle Dynamics
-2.8.1 Covariant Approach
 The accelerated motion is described by a covariant equation of motion for a relativistic charged particle under the action of the four-force in the Lorentz lab frame.
 The trajectory of a particle $\vec{x}_{cov}(t)$ is viewed from the Lorentz lab frame as a result of successive infinitesimal Lorentz transformations.
-The lab frame time $t$ in the equation of motion cannot be independent from the space variables.
+The lab frame time $t$ in the equation of motion cannot be independent of the space variables.
 This is because Lorentz transformations lead to a mixture of positions and time, and the relativistic kinematics effects are considered to be a manifestation of the relativity of simultaneity.
-2.8.2 Noncovariant (”Single Frame”) Approach
 Let us consider the conventional particle tracking approach.
-It is generally accepted that in order to describe dynamics of relativistic particles in the lab reference frame, which we assume inertial, one only needs to take into account the relativistic dependence of the particle momentum on the velocity.
+It is generally accepted that in order to describe the dynamics of relativistic particles in the lab reference frame, which we assume inertial, one only needs to take into account the relativistic dependence of the particle momentum on the velocity.
 The treatment of relativistic particle dynamics involves only corrected Newton’s second law.
@@
 Conventional particle tracking treats the space-time continuum in a non-relativistic format, as a (3+1) manifold.
-In other words, in this approach, introducing as only modification to the classical case the relativistic mass, time differ from space.
-In fact, we have no mixture of positions and time (9).
+In other words, in this approach, introducing as an only modification to the classical case the relativistic mass, the time differs from space.
+We have no mixture of positions and time  [^9] Clarification of the true content of the non-covariant theory can be found in various advanced textbooks.
+To quote e.g.
+Ferrarese and Bini [12]: ”… within a single inertial frame, the time is an absolute quantity in special relativity also.
+As a consequence, if no more than one frame is involved, one would not expect differences between classical and relativistic kinematics.
+But in the relativistic context, there are differences in the transformation laws of the various relative quantities (of kinematics or dynamics), when passing from one reference frame to another.” We see that authors give a special role to the concept of a ”single inertial frame”.
+The name ”single inertial frame” tends to suggest that a distinctive trait of non-covariant theory is the absence of relativistic kinematics in the description of particle motion.
+This point was expressed by Friedman [15]: ”Within any single inertial frame, things look precisely the same as in Newtonian kinematics: there is an enduring Euclidean three-space, a global time $t$, and the law of motion.”.
 Most of the interesting phenomena in which charges move under the action of electromagnetic fields occur in very complicated situations.
@@
 mass) of a particle.
-This study of relativistic particle motion in a constant magnetic field, usual for accelerator engineering, looks precisely the same as in nonrelativistic Newtonian dynamics and kinematics.
-The trajectory of a particle $\vec{x}(t)$, which follows from the solution of the corrected Newton’s second law, does not include relativistic kinematics effects as relativity of simultaneity and the Galilean vectorial law of addition of velocities is actually used.
+This study of relativistic particle motion in a constant magnetic field, usual for accelerator engineering, looks precisely the same as in non-relativistic Newtonian dynamics and kinematics.
+The trajectory of a particle $\vec{x}(t)$, which follows from the solution of the corrected Newton’s second law, does not include relativistic kinematics effects as relativity of simultaneity and the Galilean vectorial law of addition of velocities is used.
 Let us discuss the important problem of the addition of velocities in relativity.
-Suppose that in the case of accelerated motion one introduces an infinite sequence of co-moving frames.
-At each instant, the rest frame is a Lorentz frame centered on the particle and moving with it.
-Suppose that in inertial frame where particle is at rest at a given time, the traveler was observing light itself.
-In other words measured speed of light $v=c$, and yet the frame is moving relative the lab frame.
-How will it look to the observer in the lab frame?
-According to relativistic law of addition of velocities the answer will be $c$.
-Maxwell’s equations remain in the same form when Lorentz transformations are applied to them, but Lorentz transformations give rise to non-Galilean transformation rules for velocities, and therefore the theory of relativity shows that, if Maxwell’s equations is to be valid in the lab frame, the trajectories of the particles must include relativistic kinematics effects.
-In other words, Maxwell,s equations can be applied in the lab frame only in the case when particle trajectories are viewed, from the lab frame as the result of successive infinitesimal Lorentz transformations.
-The absence of relativistic kinematics effects is the prediction of conventional non-covariant theory and is obviously absurd from the viewpoint of Maxwell’s electrodynamics.
+It is clear that the light velocity transforms as a particle velocity.
+We cannot take one kinematics for a particle’s velocity and the other kinematics for the velocity of light.
+We consider a case when the light source in the lab frame is accelerated from rest up to velocity $u$.
+Suppose that in the Lorentz inertial frame where the light source is at rest, the traveler was observing light itself.
+In other words, $v=c$, and yet the frame is moving.
+How will it look to the observer in the Lorentz lab frame?
+According to the relativistic law of addition of velocities, the answer will be $v=c$.
+Maxwell’s equations remain in the same form when Lorentz transformations are applied to them, but Lorentz transformations give rise to non-Galilean transformation rules for velocities, and therefore the theory of relativity shows that, if Maxwell’s equations are to be valid in the lab frame, the trajectories of the particles must include relativistic kinematics effects.
+In other words, Maxwell’s equations can be applied in the lab frame only in the case when particle trajectories are viewed, from the lab frame as the result of successive infinitesimal Lorentz transformations.
+The absence of relativistic kinematics effects is the prediction of conventional (3+1) non-covariant particle tracking and is obviously absurd from the viewpoint of Maxwell’s electrodynamics.
+The relativity of simultaneity (i.e.
+mixture of positions and time) does not have a place in this description.
 Therefore, something is fundamentally, powerfully, and absolutely wrong in coupling fields and particles within a ”single inertial frame”.
@@
 It is generally believed that the electrodynamics problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations.
-In all standard derivations it is assumed that usual Maxwell’s equations and corrected Newton’s second law can explain all experiments that are performed in a single inertial frame, for instance the lab reference frame.
-Going to electrodynamics problem, the differential form of Maxwell’s equations describing electromagnetic phenomena in the same inertial lab frame (in cgs units) is given by the following expressions:
+In all standard derivations, it is assumed that Maxwell’s equations and corrected Newton’s second law can explain all experiments that are performed in a single inertial frame, for instance, the lab reference frame.
+Going to the electrodynamics problem, the differential form of Maxwell’s equations describing electromagnetic phenomena in the same inertial lab frame (in cgs units) is given by the following expressions:
 $\vec{\nabla}\cdot\vec{E}=4\pi\rho~{},$ (5)
@@
 where $\delta(\vec{x}-\vec{x}_{n}(t))$ is three-dimensional delta function, $m_{n},e_{n},\vec{x}_{n}(t)$, and $\vec{v}_{n}=d\vec{x}_{n}(t)/dt$ denote the mass, charge, position, and the velocity of the $n$th particle, respectively.
-To evaluate radiation fields arising from an external sources Eq.
+To evaluate radiation fields arising from external sources Eq.
 (8) we need the velocity $\vec{v}_{n}$ and the position $\vec{x}_{n}$ as a function of lab frame time $t$.
-It is generally accepted by physics community that equation of motion, which describes how coordinates of the particle carrying the charge change with time $t$, is described by corrected Newton’s second law Eq.
+It is generally accepted by the physics community that the equation of motion, which describes how coordinates of the particle carrying the charge change with time $t$, is described by corrected Newton’s second law Eq.
 (4).
-This coupling of Maxwell’s equations and corrected Newton’s equation is commonly accepted as useful method in accelerator physics and, in particular, in analytical and numerical calculations of radiation properties.
-Such approach to relativistic dynamics and electrodynamics usually forces the accelerator physicist to believe that the design of particle accelerators possible without detailed knowledge of the theory of relativity.
+This coupling of Maxwell’s equations and corrected Newton’s equation is commonly accepted as a useful method in accelerator physics and, in particular, in analytical and numerical calculations of radiation properties.
+Such an approach to relativistic dynamics and electrodynamics usually forces the accelerator physicist to believe that the design of particle accelerators is possible without detailed knowledge of the theory of relativity.
 However, there is a common mistake made in accelerator physics connected with the difference between $\vec{x}(t)$ and $\vec{x}_{cov}(t)$ trajectories.
@@
 To evaluate fields arising from external sources we need to know their velocity and positions as a function of the lab frame time $t$.
-Suppose one wants to calculate properties of radiation.
-Given our previous discussion the question arises, whether one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particle moving along non-covariant trajectories like $\vec{x}(t)$.
+Suppose one wants to calculate the properties of radiation.
+Given our previous discussion, the question arises, whether one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by a particle moving along non-covariant trajectories like $\vec{x}(t)$.
 We claim that the answer to this question is negative.
-This algorithm for solving usual Maxwell’s equations in the lab frame, which is considered in all standard treatments as relativistically correct, is at odds with the theory of special relativity.
+This algorithm for solving Maxwell’s equations in the lab frame, which is considered in all standard treatments as relativistically correct, is at odds with the theory of special relativity.
 This essential point has never received attention in the physical community.
-Only the solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle trajectories in the lab frame.
-2.10 Clarification of the True Content of the ”Single Frame” Theory
-Let us now examine the logical content of the concept of a ”single inertial frame”.
-If a traveler in a co moving frame, similar to an observer in the lab frame, introduces a definite coordinate-time grid, there is always a definite transformation between these two four-dimensional coordinate systems.
-Thus, particle trajectories are always viewed from the lab frame as a result of successive transformations, and the form of these transformations depends on the choice of coordinate systems in the comoving and the lab frame.
-One might well wonder why it is necessary to discuss how different inertial frames are related to one another.
-The point is that all natural phenomena follow the principle of relativity, which is a restrictive principle: it says that the laws of nature are the same (or take the same form) in all inertial frames.
-In agreement with this principle, usual Maxwell’s equations can always be exploited in any inertial frame where electromagnetic sources are at rest using Einstein synchronization procedure in the rest frame of the source.
-The fact that one can deduce electromagnetic field equations for arbitrary moving sources by studying the form taken by Maxwell’s equations under the transformation between rest frame of the source and the frame where the source is moving is a practical application of the principle of relativity.
-The question now arises how to assign a time coordinate to the lab frame.
-Coordinates serve the purpose of labeling events in an unambiguous way, and this can be done in infinitely many different ways.
-The principle of relativity dictates that Maxwell’s equations can be applied in the lab frame only in the case when Lorentz coordinates are assigned and particle trajectories are viewing from the lab frame as a result of successive infinitesimal Lorentz transformations between the lab and comoving inertial frames.
-A ”single frame” (non-covariant 3+1) approach to relativistic particle dynamics has been used in particle tracking calculations for about seventy years.
-However, the type of clock synchronization which provides the time coordinate $t$ in the corrected Newton’s equation has never been discussed in literature.
-It is clear that without an answer to the question about the method of synchronization used, not only the concept of velocity, but also the dynamics law has no physical meaning.
-A ”single frame” approach to relativistic particle dynamics is forcefully based on a definite synchronization assumption but this is actually a hidden assumption.
-According to conventional particle tracking, the dynamical evolution in the lab frame is based on the use of the lab frame time $t$ as an independent variable.
-Such approach to relativistic particle dynamics is actually based on the use of a not standard (not Einstein) clock synchronization assumption in the lab frame.
-In fact, the usual for accelerator engineering study of relativistic particle motion in a constant magnetic field looks precisely the same as in nonrelativistic Newtonian mechanics and the trajectories of the electrons does not include relativistic kinematics effects.
-According to textbooks, this is no problem.
-If no more than one frame is involved, one does not need to use (and does not need to know) the theory of relativity.
-Only when one passes from one reference frame to another the relativistic context is important.
-Conventional particle tracking in a constant magnetic field is actually based on classical Newton mechanics.
-It is generally believed that the electrodynamics problem, similar to conventional particle tracking, can be treated within a description involving a single inertial frame and one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particles moving along the non-covariant (single frame) trajectories.
-This is misconception.
-The situation when only one frame is involved and the relativistic context is unimportant cannot be realized.
-The lab observer may argue, ”I don’t care about other frames.” Perhaps the lab observer doesn’t, but nature knows that, according to the principle of relativity, Maxwell’s equations are always valid in the Lorentz comoving frame.
-Electrodynamics equations can be written down in the lab frame only when a space-time coordinate system has been specified.
-An observer in the lab frame has only one freedom.
-This is the choice of a coordinate system (i.e.
-the choice of clock synchronization convention) in the lab frame.
-After this, the theory of relativity states that the electrodynamics equations in the lab frame are the result of transformation of Maxwell’s equations from the Lorentz comoving frame to the lab frame.
-2.11 Bibliography and Notes
-1.
-Many physicists tend to think of Galilean transformations as pre-relativistic transformations between spatial coordinates and time that are not compatible with the special theory of relativity.
-To quote e.g.
-Bohm [1] ”… the Galilean law of addition of velocities implies that the speed of light should vary with the speed of the observing equipment.
-Since this predicted variation is contrary to the fact, the Galilean transformations evidently cannot be the correct one.”.
-Similar statements can also be found in recently published pedagogical papers.
-To quote e.g.
-Drake and Purvis [2] ”One of the great insights to come relativity theory was the realization that Galilean transformations are wrong.
-The correct way to translate the space-time measure of events between inertial frames is with the Lorentz transformations” However, this is not true.
-Galilean transformations are simply transformations relating a given coordinate set to another coordinate set.
-The space-time continuum can be described in arbitrary coordinates, and choice of this set of coordinates cannot change the geometry of space-time.
-2.
-The mathematical argument that in the process of transition to arbitrary coordinates the geometry of the space-time does not change, is considered in textbooks as erroneous.
-To quote L.
-Landau and E.
-Lifshitz [3]: ”This formula is called the Galileo transformation.
-It is easily to verify that this transformation, as was to be expected, does not satisfy the requirements of the theory of relativity; it does not leave the interval between events invariant.”.
-This fact is ascribed to a lack of understanding of the difference between convention-dependent and convention-invariant parts of the theory.
-In pseudo-Euclidean geometry the interval between events is an invariant in arbitrary coordinates.
-A comparison with three-dimensional Euclidean space might help here.
-In the usual 3D Euclidean space, one can consider a Cartesian coordinate system ($x,y,z)$, a cylindrical coordinate system ($r,\phi,z$), a spherical coordinate system ($\rho,\theta,\phi$), or any other.
-Depending on the choice of the coordinate system one respectively has $ds^{2}=dx^{2}+dy^{2}+dz^{2}$, $ds^{2}=dr^{2}+r^{2}d\phi^{2}+dz^{2}$, $ds^{2}=d\rho^{2}+\rho^{2}d\theta^{2}+\rho^{2}\sin\theta^{2}d\phi^{2}$.
-The metric actually does not change, but the components of the metric do, depending only on the choice of coordinates.
-In general, in fact, we write $ds^{2}=g_{ik}dx^{i}dx^{k}$.
-Considering Cartesian coordinates we will always have $g_{ij}=\mathrm{diag}(1,1,1)$.
-Similarly, Lorentz transformations between inertial frames with Einstein coordinates leave the components of the metric tensor unvaried.
-3.
-There is a widespread belief that a Lorentz transformation reduces to a Galilean transformation in the non-relativistic limit.
-To quote e.g.
-French [4]: ”The reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to Galilean relation $t^{\prime}=t$ requires $x\ll ct$ as well as $v/c\ll 1$”.
-Similar statements can also be found in recently published textbooks.
-To quote Rafelski [5]: ”A wealth of daily experience shows that the Galilean coordinate transformation (GT) is correct in the nonrelativistic limit in which the speed of light is so large that it plays no physical role.
-Any coordinate transformation replacing the GT must also agree with this experience, and thus must contain the GT in the nonrelativistic limit.” This statement is absurd conclusion even from a mathematical standpoint.
-This has been recognized by some expects, perhaps most explicitly by Baierlein who states [6] that ”If the Lorentz transformation for infinitesimal $v/c$ were to reduce to the Galilean transformation, then iterative process could never generate a finite Lorentz transformation that is radically different from the Galilean transformation.
-But the finite transformations are radically different, and so -however subtly-the infinitesimal Lorentz transformation must differ significantly from the Galilean transformation.”
-4.
-Many authors of textbooks still attribute a measurable status to the conventional quantities.
-To quote e.g.
-Cristodoulides [7] ”The fact that Galilean transformation does not leave Maxwell’s equations has already been mentioned […] On the other hand, experiments show that the speed of light in vacuum is independent of the source or observer.”.
-5.
-Because we have empirical access only to the round-trip average speed of light, statements about the magnitude and isotrophy of the one-way speed of light must reflect the assumptions made in the choice of time coordinatization, and such entities change as the theory is re-synchronized.
-To quote C.
-Anderson, I.
-Vetharaniam, and G.
-Stedman [8]: ”No experiment, then, is a ”one-way” experiment.
-An empirical test of any property of the one-way speed of light is not possible.
-Such quantities as the one-way speed of light are irreducibly conventional in nature, and recognizing this aspect is to recognize a profound feature of nature”.
-6.
-The peculiarity of the kinematic consequence of using Galilean transformations is that the speed of light emitted by a moving source depends on the relative velocity between source and observer.
-A widespread theoretical argument used to support the incorrectness of Galilean transformations is the conclusion that a Galilean transformation of the velocity of light is not consistent with the explanation of reflection and refraction.
-This idea is a part of the material in well-known books.
-To quote Pauli [9] ”[…] it is essential that the spherical waves emitted by the dipoles in the body should interfere with the incident wave.
-If we now think of the body as at rest, and the light source moving relative to it, then […] the wave emitted by the dipoles will have velocity different from that of the incident wave.
-Interference is therefore not possible.” This conclusion is incorrect.
-It is clear that the incident and scattered wave at any given point have the same frequency and can interfere.
-7. For a general discussion of the difference between path and trajectory we suggest reading the book [10].
-8.
-The standard textbooks erroneously identify the properties of space-time with the familiar form that certain synchronization-dependent quantities assume under the Einstein’s clock synchronization.
-These quantities are, for example, the formulae for ”time dilation” and ”length contraction”.
-Fortunately, it has also been occasionally stressed in the literature that the forms of the above ”relativistic effects” are coordinate-dependent, their forms depend in turn on the kind of synchronization procedure adopted.
-To quote Leubner, K.
-Auflinger, and P.
-Krumm [11]: ”… there is a widespread belief among students that the familiar form of coordinate-dependent quantities like the measured velocity of light, the Lorentz transformation between two observers, ”addition of velocities”, ”time dilation”, ”length contraction” which they assume under the standard clock synchronization, is relativity proper.
-In order to demonstrate that this is by no means so, this paper studies the consequences of a non-standard synchronization, and it is shown that drastic changes in the appearance of all these quantities are thus induced.”
-9.
-Clarification of the true content of the non-covariant theory can be found in various advanced textbooks.
-To quote e.g.
-Ferrarese and Bini [12]: ”… within a single inertial frame, the time is an absolute quantity in special relativity also.
-As a consequence, if no more than one frame is involved, one would not expect differences between classical and relativistic kinematics.
-But in the relativistic context there are differences in the transformation laws of the various relative quantities (of kinematics or dynamics), when passing from one reference frame to another.” We see that authors give a special role to concept of a ”single inertial frame”.
-The name ”single inertial frame” tends to suggest that a distinctive trait of non-covariant theory is the absence of relativistic kinematics in the description of particle motion.
-This point was expressed by Friedman [15]: ”Within any single inertial frame, things looks precisely the same as in Newtonian kinematics: there is an enduring Euclidean three-space, a global (i.e.
-absolute) time $t$, and law of motion.
-But different inertial frames are related to one another in a non-Newtonian fashion.” According to conventional particle tracking, within the ”single” frame there is no relativistic kinematics effects.
-This, as we already mentioned, contradicts the Maxwell’s electrodynamics.
+Only the solution of the dynamics equations in covariant form gives the correct coupling between Maxwell’s equations and particle trajectories in the lab frame.
 3 Space-Time and Its Coordinatization
-3.1 Introductory Remarks
+3.1 Choice of Coordinates System in an Inertial Frame
 Let us discuss an ”operational interpretation” of the Lorentz and absolute time coordinatizations.
-We should underline that we claim the non covariant approach to relativistic particle dynamics is actually based on the use of a not standard and unusual clock synchronization assumption within the theory of relativity.
-It is important to know how to operationally interpret the absolute time convention i.e.
-how one should perform the clock synchronization in the lab frame.
-The result is very interesting, since it tell us about difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view.
-3.2 Choice of Coordinates System in an Inertial Frame
 Each physical phenomenon occurs in space and time.
 A concrete method for representing space and time is a frame of reference.
-One-and-the same space and time can be represented by various coordinate-time grids, i.e., by various frames of references.
-Even the simplest space-time coordinate systems require carefully description.
+One-and-the-same space and time can be represented by various coordinate-time grids, i.e., by various frames of reference.
+Even the simplest space-time coordinate systems require careful description.
 Clocks reveal the motion of a particle through the coordinate-time grid.
-The general approach to the determination of the motion of a particle is the following: at any instant a particle has a well-defined velocity $\vec{v}$ as measured in a laboratory frame of reference.
-How is a velocity of a particle found?
-The velocity is determined once the coordinates in the lab frame are chosen, and is then measured at appropriate time intervals along the particle’s trajectory.
+The general approach to the determination of the motion of a particle is the following: at any instant, a particle has a well-defined velocity $\vec{v}$ as measured in a laboratory frame of reference.
+How is the velocity of a particle found?
+The velocity is determined once the coordinates in the lab frame are chosen and is then measured at appropriate time intervals along the particle’s trajectory.
 But how to measure a time interval between events occurring at different points in space?
@@
 The concept of synchronization is a key concept in the understanding of special relativity.
-It is possible to think of various methods to synchronize distant clocks (1).
+It is possible to think of various methods to synchronize distant clocks  [^10] According to the thesis of the conventionality of simultaneity [13, 14, 15, 16, 17], the simultaneity of distant events is a conventional matter, as it can be legitimately fixed in different manners in any given inertial reference frame.
+To quote e.g.
+Moeller [13]: ”All methods for the regulation of clocks meet with the same fundamental difficulty.
+The concept of simultaneity between two events in different places has no exact objective meaning at all since we cannot specify any experimental method by which this simultaneity could be ascertained.
+The same is therefore true also for the concept of velocity.”.
 The choice of a convention on clock synchronization is nothing more than a definite choice of coordinates system in an inertial frame of reference of the Minkowski space-time.
 The space-time continuum can be described in arbitrary coordinates.
-By changing these arbitrary coordinates, the geometry of the four-dimensional space-time obviously does not change, and in the special theory of relativity we are not limited in any way in the choice of a coordinates system.
+By changing these arbitrary coordinates, the geometry of the four-dimensional space-time does not change, and in the special theory of relativity, we are not limited in any way in the choice of a coordinates system.
 Relying on the geometric structure of Minkowski space-time, one defines the class of inertial frames and adopts a Lorentz frame with orthonormal basis vectors.
-Within the chosen Lorentz frame, Einstein’s synchronization procedure of distant clocks (which based on the constancy of the speed of light in all inertial framers) and Cartesian space coordinates are enforced.
+Within the chosen Lorentz frame, Einstein’s synchronization procedure of distant clocks (which is based on the constancy of the speed of light in all inertial frames) and Cartesian space coordinates are enforced.
 3.2 Inertial Frame where a Source of Light is at Rest
@@
 This is because Maxwell’s theory has no intrinsic anisotropy.
-It has been stated that in their original form Maxwell’s equations are only valid in inertial frames.
+It has been stated that in their original form, Maxwell’s equations are only valid in inertial frames.
 However, Maxwell’s equations can be written down in coordinate representation only if the space-time coordinate system has already been specified.
-The problem of assigning Lorentz coordinates to the lab frame in the case of accelerated motion is complicated.
+The problem of assigning Lorentz coordinates to the lab frame in the case of light source motion is complicated.
 We would like to start with the simpler question of how to assign space-time coordinates to an inertial frame, where a source of light is at rest.
-We need to give a ”practical”, ”operational” answer to this question.
+We need to give a ”practical”, and ”operational” answer to this question.
 The most natural method of synchronization consists in putting all the ideal clocks together at the same point in space, where they can be synchronized.
-Then, they can be transported slowly to their original places (slow clock transport) (2).
+Then, they can be transported slowly to their original places (slow clock transport)  [^11] In the text first published in 1923, Eddington discussed, apparently for the first time, a procedure for synchronization using slow transport of clocks [18].
+The details can be found in review [8] ( see also [19]).
+.
 The usual Maxwell’s equations are valid in any inertial frame where sources are at rest and the procedure of slow clock transport is used to assign values to the time coordinate.
-The same considerations apply when charged particles are moving in non-relativistic manner.
-In particular, when oscillating, charged particles emit radiation, and in the non-relativistic case, when charges oscillate with velocities much smaller than $c$, dipole radiation is generated and described with the help of the Maxwell’s equations in their usual form.
-Let’s examine in a more detail how the dipole radiation term comes about.
-The retardation time in the integrands of the expression for the radiation field amplitude, can be neglected in the cases where the trajectory of the charge changes little during this time.
+The same considerations apply when charged particles are moving in a non-relativistic manner.
+In particular, when oscillating, charged particles emit radiation, and in the non-relativistic case, when charges oscillate with velocities much smaller than $c$, dipole radiation is generated and described with the help of Maxwell’s equations in their usual form.
+Let’s examine in more detail how the dipole radiation term comes about.
+The retardation time in the integrands of the expression for the radiation field amplitude can be neglected in the cases where the trajectory of the charge changes little during this time.
 It is easy to find the conditions for satisfying this requirement.
@@
 Then the retardation time $\sim a/c$.
-In order to ensure that the distribution of the charges in the system does not undergo a significant change during this time, it is necessary that $a\ll\lambda$, where $\lambda$ is the radiation wavelength.
-Thus, the dimensions of the system must be small compared to radiation wavelength.
+To ensure that the distribution of the charges in the system does not undergo a significant change during this time, it is necessary that $a\ll\lambda$, where $\lambda$ is the radiation wavelength.
+Thus, the dimensions of the system must be small compared to the radiation wavelength.
 This condition can be written in still another form $v\ll c$, where $v$ is of the order of magnitude of the velocities of the charges.
@@
 Therefore, one should not be surprised to find that dipole radiation theory gives fields very much like the instantaneous theory.
-The theory of relativity offers an alternative procedure of clocks synchronization based on the constancy of the speed of light in all inertial frames.
-This is usually considered a postulate but, as we have seen, it is just a convention.
+The theory of relativity offers an alternative procedure of clock synchronization based on the constancy of the speed of light in all inertial frames.
+This is usually considered as a postulate but, as we have seen, it is just a convention.
 The synchronization procedure that follows is the usual Einstein synchronization procedure.
@@
 Indeed, in dipole radiation theory we consider the small expansion parameter $v/c\ll 1$ neglecting terms of order $v/c$.
-In other words, in dipole radiation theory we use zero order non relativistic approximation.
-Einstein synchronization is defined in terms of light signals emitted by the dipole source at rest, assuming that light propagate with the same velocity $c$ in all directions.
-Using Einstein synchronization procedure in the rest frame of the dipole source, we actually select the Lorentz coordinate system.
-Slow transport synchronization is equivalent to Einstein synchronization in inertial system where the dipole light source is at rest (3).
+In other words, in dipole radiation theory we use zero-order nonrelativistic approximation.
+Einstein synchronization is defined in terms of light signals emitted by the dipole source at rest, assuming that light propagates with the same velocity $c$ in all directions.
+Using the Einstein synchronization procedure in the rest frame of the dipole source, we select the Lorentz coordinate system.
+Slow transport synchronization is equivalent to Einstein’s synchronization in an inertial system where the dipole light source is at rest  [^12] Many authors of textbooks still attribute a reality status of the one-way speed of light.
+To quote Hrasko [20]: ”It is sometimes claimed that Einstein’s synchronization of distant clocks $A$ and $B$ is circular.
+… This argument is, however, fallacious.
+It is true that one-way measurement of light velocity can indeed be performed only if clocks at the endpoints are synchronized correctly.
+But since they need not show the correct coordinate time, they can be synchronized without light signals by transporting them from the common site in a symmetrical manner.
+….
+As we see, the thought experiment described is capable of proving the constancy of light speed if it is true or to disprove it if it is false.
+It provided, therefore, solid logical foundation for Einstein’s synchronization prescription.” This logical argument is incorrect.
+Slow clock transport synchronization is equivalent to Einstein’s synchronization only in an inertial system where the light source is at rest..
 In other words, suppose we have two sets of synchronized clocks spaced along the $x$ axis.
@@
 If we would ride together with any clock in either set, we could see that it has the same time as the adjacent clocks, with which its reading is compared.
-This is because in our case of interest, when light source is at rest, field equations are the usual Maxwell’s equations and Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions.
-Using any of these synchronization procedures in the rest frame we actually select a Lorentz coordinate system.
-In this coordinate system the metric of the light source has Minkowski form $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
+This is because in our case of interest when the light source is at rest, field equations are the usual Maxwell’s equations and Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions.
+Using any of these synchronization procedures in the rest frame we select a Lorentz coordinate system.
+In this coordinate system, the metric of the light source has Minkowski form $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
 We want now to turn to the subject of electromagnetic waves.
@@
 We now consider the case when the light source in the lab frame is accelerated from rest up to velocity $v$ along the $x$-axis.
-In other words, we consider so-called active (physical) boost.
+In other words, we consider a so-called active (physical) boost.
 When we state that the source undergoes an acceleration, and the inertial observer (with measuring devices) does not, the acceleration means acceleration relative to the fixed stars.
@@
 any active boost of velocity) has an absolute meaning.
-In the case of an active boost we are dealing with motion of the same physical system, evolving in time and treated from the point of view of the same reference system.
+In the case of an active boost, we are dealing with the motion of the same physical system, evolving in time and treated from the point of view of the same reference system.
 A fundamental question to ask is whether our lab clock synchronization method depends on the state of motion of the light source or not.
 The answer simply fixes a convention.
-The simplest method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the light source was at rest, i.e.
+The simplest method of synchronization consists of keeping, without changes, the same set of uniformly synchronized clocks used in the case when the light source was at rest, i.e.
 we still enforce the clock transport synchronization ( or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest).
-This choice is usually the most convenient one from the viewpoint of connection to laboratory reality.
-This synchronization convention preserves simultaneity and is actually based on the absolute time (or absolute simultaneity) convention.
-Absolute simultaneity can be introduced in special relativity without affecting neither the logical structure, no the (convention-independent) predictions of the theory.
-We begin with the metric as true measure of space-time intervals for an inertial observer with coordinates $(t,x,y,z)$.
+This choice is the most convenient one from the viewpoint of connection to laboratory reality.
+This synchronization convention preserves simultaneity and is based on the absolute time (or absolute simultaneity) convention.
+Absolute simultaneity can be introduced in special relativity without affecting neither the logical structure nor the (convention-independent) predictions of the theory.
+We begin with the metric as the true measure of space-time intervals for an inertial observer with coordinates $(t,x,y,z)$.
 Using the Galilean boost: we substitute $x\to x-vt$, while leaving time unchanged into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain
 $ds^{2}=c^{2}(1-v^{2}/c^{2})dt^{2}+2vdxdt-dx^{2}-dy^{2}-dz^{2}~{}.$ (11)
-This metric describes the electrodynamics of the moving light source with viewpoint of the inertial observer measurements.
-From the formula itself it is easy to see that the result of measurement of the electromagnetic field configuration in the moving system in terms of coordinate $x$ and time $t$ of the measured device at rest in the inertial frame is the same as the measured result of this configuration at point $x-vt$ and time $t=0$.
+This metric describes the electrodynamics of the moving light source from the viewpoint of the inertial observer measurements.
+From the formula itself, it is easy to see that the result of measurement of the electromagnetic field configuration in the moving system in terms of coordinate $x$ and time $t$ of the measured device at rest in the inertial frame is the same as the measured result of this configuration at point $x-vt$ and time $t=0$.
 To understand the physical interpretation of the Eq.
@@
 It is always possible to create a new frame of reference by relabeling coordinates, and then discussing physical phenomena in terms of the new coordinate labels - a passive (coordinate) transformation.
-For example, it is always possible to create a so-called comoving coordinate system in the lab frame, and then discussing radiation from the moving source in therm of the new (comoving) coordinate labels.
+For example, it is always possible to create a so-called comoving coordinate system in the lab frame, and then discuss radiation from the moving source in therm of the new (comoving) coordinate labels.
 In the comoving coordinate system, fields are expressed as a function of the independent variables $x^{\prime},y^{\prime},z^{\prime}$, and $t^{\prime}$.
-The variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$ can be expressed in terms of the independent variables $x,y,z,t$ by means of a passive Galilean transformation, so that fields can be written in terms of $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$.
+The variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$ can be expressed in terms of the independent variables $x,y,z,t$ using a passive Galilean transformation so that fields can be written in terms of $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$.
 After the passive transformation, the Cartesian coordinates of the source transform as $x^{\prime}=x-vt,~{}y^{\prime}=y,~{}z^{\prime}=z$.
@@
 One must take into account that $(t,x,y,z)$ are coordinates of the rule clock structure at rest in the inertial frame.
-The equivalence of the active and passive pictures is due to the fact that moving system one way is equivalent to moving the coordinate system the other way by an equal amount.
+The equivalence of the active and passive pictures is because moving system one way is equivalent to moving the coordinate system the other way by an equal amount.
 Obviously, there is no a priory guarantee that the world behave like this.
-For instance, Newton mechanics is not invariant under passive rotation as demonstrated Newton’s bucket experiment.
-The empirical principle implies that within a single inertial frame every passive Galilean boost is postulated to corresponds to an active Galilean boost.
-According to this equivalence of passive and active boosts, in our case of interest the Maxwell’s equations always valid in the comoving frame.
+For instance, Newton’s mechanics is not invariant under passive rotation as demonstrated Newton’s bucket experiment.
+The empirical principle implies that within a single inertial frame, every passive Galilean boost is postulated to correspond to an active Galilean boost.
+According to this equivalence of passive and active boosts, in our case of interest, Maxwell’s equations are always valid in the comoving frame.
 In the comoving frame, fields are expressed as a function of the independent variables $x^{\prime},y^{\prime},z^{\prime}$, and $t^{\prime}$.
 The electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$.
-However, the variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$ can be expressed in terms of the independent variables $x,y,z,t$ by means of a Galilean transformation, so that fields can be written in terms of $x,y,z,t$.
+However, the variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$ can be expressed in terms of the independent variables $x,y,z,t$ using a Galilean transformation, so that fields can be written in terms of $x,y,z,t$.
 From the Galilean transformation $x^{\prime}=x-vt,~{}y^{\prime}=y,~{}z^{\prime}=z,~{}t^{\prime}=t$, after partial differentiation, one obtains $\partial/{\partial t}=\partial/{\partial t^{\prime}}-v\partial/{\partial x^{\prime}}$, $\partial/{\partial x}=\partial/{\partial x^{\prime}}$.
@@
 The solution of this equation $F[x-(c+v)t]+G[x+(-c+v)t]$ is the sum of two arbitrary functions, one of argument $x-(c+v)t$ and the other of argument $x+(-c+v)t$.
-Here we obtained the solution for waves which move in the $x$ direction by supposing that the field does not depend on $y$ and $z$.
-The first term represents a wave traveling forward in the positive $x$ direction, and the second term a wave traveling backwards in the negative direction.
-We conclude that the speed of light emitted by a moving source measured in the lab frame $(t,x)$ depends on the relative velocity of source and observer, in our example $v$.
+Here we obtained the solution for waves that move in the $x$ direction by supposing that the field does not depend on $y$ and $z$.
+The first term represents a wave traveling forward in the positive $x$ direction, and the second term represents a wave traveling backward in the negative direction.
+We conclude that the speed of light emitted by a moving source measured in the lab frame $(t,x,y,z)$ depends on the relative velocity of the source and observer, in our example $v$.
 In other words, the speed of light is compatible with the Galilean law of addition of velocities.
@@
 One fixed, such convention should be applied and kept in a consistent way for both dynamics and electrodynamics.
-In fact, the light velocity transforms as a particle velocity in the absolute time coordinatization as it mast be.
+In fact, the light velocity transforms as a particle velocity in the absolute time coordinatization as it must be.
 3.4 Discussion
@@
 It can be said that the particle has some inner variables - that we do not yet know about.
-Any phenomenological law, which is valid in the Lorentz rest frame, can be embedded in the four dimensional space-time only by using Lorentz coordinatization.
+Any phenomenological law, that is valid in the Lorentz rest frame, can be embedded in the four-dimensional space-time only by using Lorentz coordinatization.
 Now we wish to consider the electrodynamics.
@@
 inner variables) behind the laws of electrodynamics.
-It is obvious that electromagnetic fields are fundamental and the electrodynamics theory meets all requirements of the theory of relativity.
+Electromagnetic fields are fundamental and the electrodynamics theory meets all requirements of the theory of relativity.
 Consequently, one can use any synchronization to describe the electromagnetic field dynamics.
-The variety of methods of description physical systems makes it possible to select for each problem the representation most suitable for its solution.
+The variety of methods of description of physical systems makes it possible to select for each problem the representation most suitable for its solution.
 It is always possible to create a new frame of reference by relabeling coordinates, and then discussing physical phenomena in terms of the new coordinate labels - a passive transformation.
 An active transformation of a physical system is its motion, i.e., a variation in its characteristics under the effect of some internal or external interactions.
-We are dealing with motion of the same physical system, evolving in time and treated from the point of view of the same reference system.
-Now let us discuss correspondence between passive and active transformations.
-A comparison between the special relativity and the Newton theory might help here.
+We are dealing with the motion of the same physical system, evolving in time and treated from the point of view of the same reference system.
+Now let us discuss the correspondence between passive and active transformations.
+A comparison between special relativity and Newton’s theory might help here.
 The formalism of non-relativistic Newton’s theory is associated with absolute space structure.
-However, there is relativity of the uniform motions and we cannot determine the particle absolute velocity.
+However, there is relativity of the uniform motions and we cannot determine the particle’s absolute velocity.
 Passive Galilean boosts are dynamical symmetries of Newtonian mechanics.
-The equivalence of the active and passive Galilean boost is due to the fact that moving system one way is equivalent to moving the coordinate system the other way by an equal amount as described by the famous (Galileo) ship experiment .
-The reason of this (Galilean group) symmetry is simply that equations of motion do not explicitly depend on velocity.
+The equivalence of the active and passive Galilean boost is because that moving system one way is equivalent to moving the coordinate system the other way by an equal amount as described by the famous (Galileo) ship experiment.
+The reason for this (Galilean group) symmetry is simply that equations of motion do not explicitly depend on velocity.
 The internal dynamics of the system does not change under the Galilean boost.
-In the same time, Newtonian mechanics is not invariant under passive rotation.
-In fact, the form of equations of motion change due to pseudo-force terms (4).
-The Newton bucket argument by Newton was based on comparison of two different situations: a vessel of water at rest, vs the vessel after it has gradually been rotated from rest to a state of uniform angular velocity.
-The argument against the equivalence of the active and passive rotation is based on notion that in both cases there is no relative motion between the water and the bucket, while the shape of the water turns out to be different.
+At the same time, Newtonian mechanics is not invariant under passive rotation.
+The form of equations of motion change due to pseudo-force terms  [^13] For a general discussion of correspondence between passive and active transformations we suggest reading the paper [21]..
+The Newton bucket argument by Newton was based on a comparison of two different situations: a vessel of water at rest, vs the vessel after it has gradually been rotated from rest to a state of uniform angular velocity.
+The argument against the equivalence of the active and passive rotation is based on the notion that in both cases there is no relative motion between the water and the bucket, while the shape of the water turns out to be different.
 The passive rotation transformation is no more than a change of representation.
-There is no relativity of accelerations and we can determine the particle absolute acceleration in the inertial frame.
+There is no relativity of accelerations and we can determine the particle’s absolute acceleration in the inertial frame.
 The absolute acceleration means acceleration relative to the fixed stars.
-Let us now move to consider correspondence between active and passive transformations in special relativity.
-Within in single inertial frame the kinematical symmetry is corresponds to dynamical symmetry and the symmetry principle (i.e.
+Let us now move to consider the correspondence between active and passive transformations in special relativity.
+Within a single inertial frame, the kinematical symmetry corresponds to dynamical symmetry and the symmetry principle (i.e.
 equivalence of the active and passive Galilean boost) is satisfied.
-It should be note that passive Galilean boost within a single inertial frame is simply another parametrization of the observations of the inertial observer.
-According to passive transformation ($x\to x-vt$, $t\to t$), the motion of fixed stars with respect to the observer and his devices not changes.
+It should be noted that passive Galilean boost within a single inertial frame is simply another parametrization of the observations of the inertial observer.
+According to passive transformation ($x\to x-vt$, $t\to t$), the motion of fixed stars with respect to the observer and his devices does not change.
 It is important to note that $(t,x,y,z)$ are coordinates of the rule clock structure at rest in the inertial frame.
-In the Chapter 5 we will consider the problem of aberration of light in accelerated systems on the basis of the theory of special relativity.
-This problem can be adequately treated only by an approach which uses the absolute time coordinatization.
-In order to write electrodynamics laws in an accelerated frame, we define the metric of the accelerated frame using the inverse Galilean boost.
-We want to emphasize that in the case of an inverse Galilean boost of velocity we consider the effect of interaction which defined in terms of acceleration motion relative to the fixed stars.
-Any acceleration relative to the fixed stars has an absolute meaning.
-This means that the passive Galilean transformation within a single inertial frame is quite distinct from the (passive) inverse Galilean transformation, $x\to x_{n}+vt$, $t\to t_{n}$, i.e acceleration with respect to fixed stars of an observer (with his measuring instruments) from inertial frame to the accelerate frame.
+In Section 5 we will consider the problem of aberration of light in accelerated systems based on the theory of special relativity.
+This problem can be adequately treated only by an approach that uses absolute time coordinatization.
+To write electrodynamics laws in an accelerated frame, we define the metric of the accelerated frame using the inverse Galilean boost.
+We want to emphasize that in the case of an inverse Galilean boost of velocity, we consider the effect of interaction which is defined in terms of acceleration motion relative to the fixed stars.
+This means that the passive Galilean transformation within a single inertial frame is quite distinct from the (passive) inverse Galilean transformation, $x\to x_{n}+vt$, $t\to t_{n}$, i.e acceleration with respect to fixed stars of an observer (with his measuring instruments).
 One must take into account that $(t_{n},x_{n},y_{n},z_{n})$ are coordinates of the rule clock structure at rest in the accelerated frame.
-Such transition of the coordinates of the light source does not stand for a change of representation, but for a change in the absolute motion of the light source with respect to the fixed stars.
-The fact that in the real process of transmission to an accelerated (with respect to the fixed stars) frame the observer will experience the inertial force, is not accounted in the standard treatments.
-We argue that reciprocity of the inertial frames, which is considered in textbooks as relativistically correct, is at odds with special relativity.
+Such transition of the coordinates of the light source does not stand for a change of representation but for a change in the absolute motion of the light source (and rule-clock structure) with respect to the fixed stars.
+The fact that in the real process of transmission to an accelerated (with respect to the fixed stars) frame the observer will experience the inertial force, is not accounted for in the standard treatments.
+We argue that the reciprocity of the inertial frames, which is considered in textbooks as relativistically correct, is at odds with special relativity.
 The arguments concerning the relativity of motion in our case of interest cannot be applied, since the inertial and accelerated reference systems are not equitable.
-The formalism of relativistic physics is associated with absolute space-time structure.
-There is no relativity of the uniform motions and we can determine the system absolute velocity in the initial inertial frame.
-The reason of this asymmetry is simply that electrodynamics equations are fundamental and explicitly depend on velocity.
-According to the special relativity, there is a real difference between accelerated (with respect to the fixed stars) inertial frame and the inertial frame without accelerational history and this difference is closely associated with the inertial forces.
-It should be note that the duration of the acceleration period has a negligible effect on the anisotropy in the accelerated frame.
+According to special relativity, there is a real difference between the accelerated (with respect to the fixed stars) inertial frame and the inertial frame without accelerational history, and this difference is closely associated with the inertial forces.
+It should be noted that the duration of the acceleration period has a negligible effect on the anisotropy in the accelerated frame.
 Nevertheless, the integral effect of inertial forces completely determined the problem.
 A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer.
-We can treat a uniformly accelerating frame as if it were an in inertial frame with the addition of a uniform pseudo-gravitational field.
-This is basic content of equivalence principle.
+We can treat a uniformly accelerating frame as if it were an inertial frame with the addition of a uniform pseudo-gravitational field.
+This is the basic content of the equivalence principle.
 The principle of equivalence can be applied to solve non-inertial kinematics problems with dynamics methods.
-For instance, when the accelerated system starts moving with constance velocity the pseudo-gravitational field is zero.
-For the gravitational potential difference between the two frames we get (in the simplified second order approximation) $\phi=-v^{2}/2$.
-The fact that the inertial observer always turns out to have aged more that accelerated one, in the equivalence principle analysis is elementary consequence of the slowly down of the rate of the processes in the pseudo-gravitational potential.
-In fact, for clock at rest in the accelerated frame we have $dt_{prop}=dt_{n}(1+\phi/c^{2})=dt_{n}[1-v^{2}/(2c^{2})]$.
-In other words, the time dilation effect is absolute, but not relative.
 There is another way to solve the problem.
-From a space-time geometrical viewpoint, the explanation consists in using the fact that metric tensor must be continuous quantity.
-When the accelerated system starts moving with constant velocity the procedure of smooth tailoring of metric tensor can be performed.
+The formalism of relativistic physics is associated with absolute space-time structure.
+There is no relativity of the uniform motions and we can determine the system’s absolute velocity in the initial inertial frame.
+From a space-time geometrical viewpoint, the explanation consists in using the fact that the metric tensor must be a continuous quantity.
+When the accelerated system starts moving with constant velocity the procedure of smooth tailoring of the metric tensor can be performed.
 The (non-diagonal) Langevin metric is found by matching the accelerated frame and initial inertial frame metric tensors.
 Many experts who learned the theory of relativity using textbooks will find this statement disturbing.
-Indeed, according to textbooks, the problem is complete symmetric and both frames are equivalent.
-The popular belief remains that coordinates of these frames should be related by the Lorentz transformations and the metric in both frames will be diagonal, although it is an exercise in elementary algebra to prove discontinuity of metric tensor.
-With mathematical viewpoint, we must conclude that when authors of textbooks are dealing with transformation and this transformation does not preserve continuity of the metric tensor there are no symmetry between inertial frames.
-The initial inertial frame is not physically identical to the Newton’s absolute space.
-The special relativity taught us to think in terms of pseudo-Euclidean space-time geometry.
-In special relativity the absolute space and the absolute time of Newton are fused into absolute space-time.
-We will discuss this subject further in the Chapter 5, Chapter 9.
+Indeed, according to textbooks, the problem is completely symmetric and both frames are equivalent.
+The popular belief remains that the coordinates of these frames should be related by the Lorentz transformations and the metric in both frames will be diagonal.
+However, after diagonalization, the metric tensor in the accelerated frame must abruptly change from the Langevin value to the Minkowski value.
+Therefore, contrary to the textbook view, Lorentz transformation does not preserve the continuity of the metric tensor.
+We will discuss this subject further in Chapter 5.
 3.5 A Clock Resynchronization Procedure
-After the Galilean boost in the case of the absolute time coordinatization we can see that the homogeneous wave equation for the field in the lab frame has nearly but not quite the usual, standard form that takes when there is no acceleration up to velocity $v$.
+After the Galilean boost in the case of the absolute time coordinatization we can see that the homogeneous wave equation for the field in the lab frame has nearly but not quite the usual, standard form that takes when there is no acceleration.
 The main difference consists in the crossed term $\partial^{2}/\partial t\partial x$, which complicates the solution of the equation.
@@
 In this approximation, it is always possible to introduce a time transformation $t\to t+xv/c^{2}$.
-After this time shift we obtain the usual wave equation.
+After this time shift we obtain the usual wave equation  [^14] Rotation of the $x$-axis without changing the orientation of the $t$-axis is described by a Galilean transformation $x^{\prime}=x-vt$, $t^{\prime}=t$.
+Rotation of the $t$-axis without changing the orientation of the $x$-axis is described by the transformation $x^{\prime}=x$, $t^{\prime}=t+xv/c^{2}$.
+In the first-order approximation, this transformation mathematically is similar to the the Galilean transformation $x^{\prime}=x+vt$, $t^{\prime}=t$.
 Let us now consider the case of an arbitrary velocity along the $x$-axis.
-We have, then, a general method for finding solution of electrodynamics problem in the case of the absolute time coordinatization.
+We have, then, a general method for finding a solution of the electrodynamics problem in the case of absolute time coordinatization.
 The new independent variables $(x_{L},t_{L})$ can be expressed in terms of the old independent variables $(x,t)$:
 $ct_{L}=t\sqrt{1-v^{2}/c^{2}}+vx/[c^{2}\sqrt{1-v^{2}/c^{2}}]~{},~{}x_{L}=x/\sqrt{1-v^{2}/c^{2}}~{},$ (13)
-Since the change variables completed by the Galilean transformation is mathematically equivalent to the Lorentz transformation, it obviously follows that transforming to new variables $x_{L},t_{L}$ leads to the usual Maxwell’s equations.
+Since the change variables completed by the Galilean transformation is mathematically equivalent to the Lorentz transformation, it follows that transforming to new variables $x_{L},t_{L}$ leads to the usual Maxwell’s equations.
 In particular, the wave equation Eq.
@@
 $\Box_{L}^{2}E=\nabla_{L}^{2}E-\partial^{2}E/\partial(ct_{L})^{2}=0~{}.$ (14)
-In the new variables the velocity of light emitted by a moving source is constant in all directions, and equal to the electrodynamics constant $c$.
-The overall combination of the Galilean boost and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the lab frame, but in this context this transformation are only to be understood as useful mathematical device, which allow one to solve the electrodynamics problem in the choice of absolute time synchronization with minimal effort.
-We can now rise an interesting question: do we need to transform the results of the electrodynamics problem solution into the original variables?
+In the new variables, the velocity of light emitted by a moving source is constant in all directions, and equal to the electrodynamics constant $c$.
+The overall combination of the Galilean boost and variable changes yields the Lorentz transformation in the case of absolute time coordinatization in the lab frame, but in this context, this transformation is only to be understood as a useful mathematical device, which allows one to solve the electrodynamics problem in the choice of absolute time synchronization with minimal effort.
+We can now raise an interesting question: do we need to transform the results of the electrodynamics problem solution into the original variables?
 We state that the variable changes performed above have no intrinsic meaning - their meaning only being assigned by a convention.
@@
 Suppose that in the lab frame the velocities of ”1” and ”2” are $\vec{v}_{1}$, $\vec{v}_{2}$ and $\vec{v}_{1}\neq\vec{v}_{2}$.
-The question now arises how to assign a time coordinate to the lab reference frame.
+The question now arises of how to assign a time coordinate to the lab reference frame.
 We have a choice between an absolute time coordinate and a Lorentz time coordinate.
-The most natural choice, from the point of view of connecting to the laboratory reality, is the absolute time synchronization.
-In this case simultaneity is absolute, and for this we should prepare, for two sources, only one set of synchronized clocks in the lab frame.
-On the other hand, Maxwell’s equations are not form-invariant under Galilean transformations, that is, their form is different on the lab frame.
-In fact, the use of the absolute time convention, implies the use of much more complicated field equations, and these equations are different for each source.
+The most natural choice, from the point of view of connecting to the laboratory reality, is absolute time synchronization.
+In this case, simultaneity is absolute, and for this, we should prepare, for two sources, only one set of synchronized clocks in the lab frame.
+On the other hand, Maxwell’s equations are not form-invariant under Galilean transformations, that is, their form is different in the lab frame.
+The use of the absolute time convention implies the use of much more complicated field equations, and these equations are different for each source.
 Now we are in the position to assign Lorentz coordinates.
-The only possibility to introduce Lorentz coordinates in this situation consists in introducing individual coordinate systems (i.e.
-individual set of clocks) for each source.
-It is clear that if operational methods are at hand to fix the coordinates for the first source, the same methods can be used to assign values to the coordinates for the second source and these will be two different Lorentz coordinate systems.
-Suppose now that the second secondary source (e.g.
-mirror or pupil detection) interacts with the light beam radiated from the first source.
-The peculiarity of this problem with the viewpoint of the special relativity is that here we cannot prepare for sources a common Lorentz coordinatization in the single inertial frame.
+The only possibility to introduce Lorentz coordinates in this situation consists of introducing individual coordinate systems (i.e.
+individual sets of clocks) for each source.
+Suppose now that the second secondary source, for instance, a tangentially moving mirror, interacts with the light beam radiated from the first source at rest.
+The peculiarity of this problem with the viewpoint of special relativity is that here we cannot prepare for sources a common Lorentz coordinatization in the single inertial frame  [^15] It is clear that the electrodynamics equations must be identical for incoming and reflected waves.
+The time under Einstein’s synchronization for the moving mirror is readily obtained by introducing the time shift $xv/c^{2}$ in the first-order approximation.
+In this synchronization, we describe the reflection using Maxwell’s equations.
+This time shift has the effect of rotation of the radiation wavefront on the angle $v/c$ for both sources.
+Let us discuss the electrodynamics of the source at rest.
+In the absolute time coordinatization, the observer describes the light beam from the emitter at rest using the diagonal wave equation.
+The new crossed term that has to be put into the wave equation due to time shift leads to the non-diagonal wave equation (see footnote in Section 3.5).(see the next section for more details).
 3.6 Radiation by a Moving Source. Peculiarity of the Collinear Geometry
-The utilization of the theory of relativity becomes indispensable when we consider optical phenomena associated with a motion of light sources.
 So far we have considered the covariant and non covariant ways to solve the problem of radiation by a moving source in an inertial frame.
 Let us consider the acceleration of a dipole source in the lab inertial frame up to velocity $v$ along the $x$-axis.
-The question arises how to assign synchronization in the lab frame after the source acceleration.
-Before acceleration we picked a Lorentz coordinate system.
-Without changing synchronization in the lab frame after the source acceleration we have a complicated situation as concerns electrodynamics of moving charges.
-As a result of such boost, the transformation of time and spatial coordinates has no form of a Lorentz transformation.
-The Maxwell’s equations can be applied in the lab inertial frame only in the case when Lorentz coordinates are assigned.
-In order to keep the Lorentz coordinatiyation after acceleration, one needs to change the space-time coordinate system by introducing new variables $(x_{L},y_{L},z_{L},t_{L})$, Eq.
+The question arises of how to assign synchronization in the lab frame after the source acceleration.
+Before acceleration, we picked a Lorentz coordinate system.
+Without changing synchronization in the lab frame after the source acceleration we have a complicated situation concerning the electrodynamics of moving charges.
+As a result of such a boost, the transformation of time and spatial coordinates has no form of a Lorentz transformation.
+Maxwell’s equations can be applied in the lab inertial frame only in the case when Lorentz coordinates are assigned.
+To keep the Lorentz coordinatization after acceleration, one needs to change the space-time coordinate system by introducing new variables $(x_{L},y_{L},z_{L},t_{L})$, Eq.
 (13).
-In the new variables the velocity of light emitted by a moving source is equal to the electrodynamics constant $c$ in all directions.
-The aberration of light and Doppler effect are practical cases of study for illustration the difference between covariant and non-covariant approaches.
-The present treatment is designed to try and bring out the role of the Lorentz time transformation.
+In the new variables, the velocity of light emitted by a moving source is equal to the electrodynamics constant $c$ in all directions.
+The aberration of light and the Doppler effect are practical cases of study to illustrate the difference between covariant and non-covariant approaches.
 The most convenient way for calculating radiation is to make use of the formulae for Lorentz transition from one inertial frame of reference to another.
-It is convenient to introduce the reference frame $S^{\prime}$ fixed to the source.
-Then the problem solved as follows.
-In the frame $S^{\prime}$ the 4-vector of the light beam is specified, i.e.
-the frequency and the propagation direction of light are known.
-The frequency of light in the frame $S$ is easily to find using the Lorentz transformation formulae.
-Specifically, let the radiation have a dipole character and frequency $\omega_{0}$ in the frame of reference in which the source is at rest.
+Specifically, let the radiation have a dipole character and frequency $\omega_{0}$ in the frame of reference $S^{\prime}$ in which the source is at rest.
 Then in the laboratory frame of reference $S$ in which the source as a whole moves with velocity $v$ and the radiation comes along the velocity direction we observe the so-called radial Doppler effect.
 It is described by the well-known formula $\omega=\omega_{0}\sqrt{1-v^{2}/c^{2}}/(1-v/c)$.
-The effect of the factor $\sqrt{1-v^{2}/c^{2}}$ can be summarized in the following statement: on the moving object time is flowing slower than expected (time dilation).
-It is interesting to note that, the use of the conventional coupling of Maxwell’s equations and corrected Newton’s equation for the calculation of the radiation from a moving source does not necessarily leads to mistake.
-For rectilinear motion of the source and the emitted light beam, the non-covariant and covariant approaches produce the same trajectories, and Maxwell’ equations are compatible with the result of conventional particle tracking.
+The effect of the factor $\sqrt{1-v^{2}/c^{2}}$ can be summarized in the following statement: on the moving object, time is flowing slower than expected (time dilation).
+It is interesting to note that, the use of the conventional coupling of Maxwell’s equations and corrected Newton’s equation for the calculation of the radiation from a moving source does not necessarily lead to a mistake.
+For rectilinear motion of the source and the emitted light beam, the non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking.
 This method was incorrectly extended to the case of non-collinear geometry.
-At this point a reasonable question arises: why the same method gives incorrect result in the case of radiation of light from a source moving perpendicular to its radiated light beam propagation.
-It is not difficult to see that the peculiarity of the collinear geometry where there is a source moving along the same line as the radiated beam is that here the velocity is perpendicular to the plane of radiation wavefront (i.e plane of simultaneity).
+At this point, a reasonable question arises: why the same method gives incorrect results in the case of radiation of light from a source moving perpendicular to its radiated light beam propagation?
+It is not difficult to see that the peculiarity of the collinear geometry where there is a source moving along the same line as the radiated beam is that here the velocity is perpendicular to the plane of radiation wavefront (i.e.
+plane of simultaneity).
 Thus, for collinear motion, the plane of simultaneity in the absolute time coordinatization will have the same orientation for the Lorentz coordinatization.
-So far we have dealt with the longitudinal effect when the motion takes place along the straight line through the source and the receiver.
-In the transverse case the motion is perpendicular to this direction and the source is accelerated along the radiation wavefront (i.e.
+In the transverse case, the motion is perpendicular to this direction, and the source is accelerated along the radiation wavefront (i.e.
 initial plane of phase simultaneity).
@@
 Let us first consider a Lorentz transformation.
-The explanation consists in using a Lorentz boost to describe the uniform translation motion of the light source in the lab frame.
-It is clear, that wavefront phases which are simultaneous in $S^{\prime}$, but occur at different $x^{\prime}$ locations, are not simultaneous in $S$.
-If make a Lorentz boost, we automatically introduce a time transformation $t^{\prime}=t-xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation wavefront in the lab frame on the angle $v/c$ in the first order approximation.
+The explanation consists of using a Lorentz boost to describe the uniform translation motion of the light source in the lab frame.
+It is clear, that wavefront phases which are simultaneous in $S^{\prime}$ but occur at different $x^{\prime}$ locations, are not simultaneous in $S$.
+If make a Lorentz boost, we automatically introduce a time transformation $t^{\prime}=t-xv/c^{2}$, and the effect of this transformation is just a rotation of the radiation wavefront in the lab frame on the angle $v/c$ in the first-order approximation.
 According to Maxwell’s electrodynamics, radiation is always emitted in the direction normal to the radiation wavefront.
 Then, the radiated light beam is propagated at the angle $v/c$, yielding the phenomenon of the aberration of light.
-This remarkable phenomenon is what we shall discuss in the next chapter.
+This remarkable phenomenon is what we shall discuss in the next section.
 According to the conventional (3+1) approach, the simultaneity is absolute and there is no mixture of positions and time when sources change their velocities in the inertial frame.
 Thus it seems as if the conventional (3+1) approach is unable to account for the geometric phenomenon of aberration.
-In making use of the Maxwell’s equations and the absolute time transformation we obtain no deflection at all of the wavefront plane and the energy transport direction.
+In making use of Maxwell’s equations and the absolute time transformation we obtain no deflection at all of the wavefront plane and the energy transport direction.
 Let us now return to our consideration of the conventional coupling between Maxwell’s equations and particle trajectories in the ”single lab frame”.
-According to the theory of special relativity, there is no objection to the standard description of the optical phenomena associated with a motion of light source in the case when the radiation comes along the velocity direction.
-For example, for collinear geometry, the conventional coupling fields and particles may be a useful approach for analysis of the Doppler effect.
-In order to understand the relativistic red shifting of the light source in a moving system, we have to watch the machinery of the source and see what happens when it is moving.
-Since that is rather difficult, we shall take a very simple a kind of source, but it will work in principle.
+According to the theory of special relativity, there is no objection to the standard description of the optical phenomena associated with a motion of a light source in the case when the radiation comes along the velocity direction.
+For example, for collinear geometry, the conventional coupling fields and particles may be a useful approach for the analysis of the Doppler effect.
 […]
-According to covariant approach, the various relativistic kinematics effects concerning to the dipole radiation setup, turn up in successive orders of approximation.
+The relativistic motion of an electron is described by $d\vec{p}/dt=e\vec{v}\times\vec{H}/c$ with $\vec{p}=m\vec{v}/\sqrt{1-v^{2}/c^{2}}$.
+This equation can be written as $d\vec{v}/dt=\vec{v}\times\vec{\omega}_{L}$ where $\vec{\omega}_{L}=e\vec{H}\sqrt{1-v^{2}/c^{2}}/(mc)=\vec{\omega}_{L0}\sqrt{1-v^{2}/c^{2}}$ is the relativistic Larmor frequency.
+When the source is accelerated, the speed of electrons is increased, and therefore the mass is also increased and the electron is heavier.
+The emissivity presents a spectrum in which the frequency is given by $\omega=\omega_{L0}\sqrt{1-v^{2}/c^{2}}/(1-v/c)$.
+The frequency is multiplied by the relativistic Larmor frequency accounting at the same time through the denominator for the Doppler effect caused by the motion parallel to the magnetic field.
+We can therefore conclude that the frequency of oscillations of an electron is slowed down if the source is made to move with a velocity $v$.
+At first sight, the machinery of the source described does not involve the effect of relativistic time dilation at all.
+But it is there in the assumption that the mass of a moving object is equal to its relativistic mass $m_{r}=m/\sqrt{1-v^{2}/c^{2}}$.
+According to the covariant approach, the various relativistic kinematics effects concerning to the dipole radiation setup, turn up in successive orders of approximation.
 In the first order $(v/c)$.
@@
 - time dilation.
-The first order kinematics term $(v/c)$ plays an essential role only in the description of the dipole radiation in the perpendicular geometry.
+The first-order kinematics term $(v/c)$ plays an essential role only in the description of the dipole radiation in the perpendicular geometry.
 In the case of collinear geometry, a motion of the dipole source, according to the theory of relativity, influences the kinematics terms of the second order $(v/c)^{2}$ only.
-The ignorance of this distinction between collinear and non collinear boosts is the source of much confusion in the literature.
 In the non-covariant approach, a solution of the dynamics problem in the lab frame makes no reference to Lorentz transformations.
 This means that, for instance, within the lab frame the motion of particles looks precisely the same as predicted by classical mechanics, with its absolute time.
-The relativity of simultaneously (i.e.
-mixture of positions and time) do not have a place in this description.
 Such (the absolute time) method is suitable to account for the outcome of the experiments in collinear geometry.
-We argue that this algorithm for solving usual Maxwell’s equations in the lab frame is not applicable in the transverse case.
-However, the Maxwell’s equations in the lab frame are compatible only with covariant trajectories calculated by using Lorentz coordinates, therefore including such relativistic features as relativity of simultaneity.
-3.8 Absolute Time Synchronization in a Laboratory Frame
-The utilization of the electrodynamics in the absolute time coordinatization becomes indispensable when we use non-covariant (3+1) dimensional approach (i.e.
+We argue that this algorithm for solving Maxwell’s equations in the lab frame is not applicable in the transverse case.
+The utilization of the electrodynamics in absolute time coordinatization becomes indispensable when we use a non-covariant (3+1) dimensional approach (i.e.
 old kinematics) to relativistic particle dynamics.
-According to conventional particle tracking, the dynamical evolution in the lab frame is based on the use of the lab time $t$ as independent variable.
+According to conventional particle tracking, the dynamical evolution in the lab frame is based on the use of the lab time $t$ as the independent variable.
 The relativistic kinematics effects do not have a place in this description.
-Such approach to relativistic particle dynamics is actually based on the use of the absolute time synchronization assumption in the lab frame.
-If one wants to use the corrected Newton equations (in other words, using time as independent variable), only solution of the electrodynamics equations in non-covariant form (i.e.
+Such an approach to relativistic particle dynamics is based on the use of the absolute time synchronization assumption in the lab frame.
+If one wants to use the corrected Newton equations (in other words, using time as the independent variable), only the solution of the electrodynamics equations in non-covariant form (i.e.
 in the absolute time coordinatization) gives the correct coupling between electrodynamics equations and particle trajectories in the ”single frame”.
-That is the reason why the absolute time synchronization convention is extremely important from the applied point of view.
-3.9 Experimental Test of Coupling between the Electrodynamics and Dynamics
-In the Chapter 16 we shall discuss one of the consequences of our aberration of light theory, the radiation from an X-ray Free Electron Laser.
-Using XFEL example we can demonstrate that the coupling fields and particles in the conventional approach is intrinsically incorrect.
-The usual XFEL theory based on the use of old Newtonian kinematics for particle dynamics and the Einstein’s kinematics for the electrodynamics.
-In fact, the usual theoretical treatment of relativistic particle dynamics involves only a corrected Newton’s second law and is based on the use Galilean edition of velocities.
-This coupling of Maxwell’s equations and corrected Newton’s equation is commonly accepted as useful method in accelerator physics and, in particular, in numerical calculations of radiation properties.
-For rectilinear motion of the (ultra-relativistic) modulated electron beam, non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking.
-Let us now move on to consider the predictions of the existing XFEL theory in the case of non-collinear electron beam motion.
-As well-known result of conventional particle tracking states that after an electron beam is kicked by a weak dipole magnet there is a change in the trajectory of the electron beam, while the orientation of the modulated wavefront remains as before.
-In the absolute time coordinatization the simultaneity of ivents have absolute character.
-The absolute character of temporal coincidence of events is a consequence of the (hidden) absolute time synchronization convention.
-According to this old kinematics, the modulation wavefront remains unvaried.
-In other words, the kick results in a difference between the directions of the electron motion and the normal to the modulation wavefront (i.e.
-the kick results in a wavefront tilt).
-In existing XFEL theory the wavefront tilt is considered as real.
-However, in the ultra-relativistic limit, the wavefront tilt has no exact objective meaning at all since, due to finiteness of the speed of light, we cannot specify any experimental method by which this tilt could be ascertained.
-According to conventional (3+1) space-time treatment, a transverse kick does not change the orientation of a modulation wavefront, and hence suppresses the radiation emitted in the direction of the electron beam motion.
-In fact, according to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the modulation wavefront.
-This is because Maxwell’s theory has no intrinsic anisotropy.
-The electrons lying on the plane of simultaneity (i.e.
-on the modulation wavefront) gives rise to spherical radiation wavelets, and these combine according to Huygens’ principle to form what what is effective a radiated wave.
-If the plane of simultaneyty is $xy$-plane (i.e.
-beam modulation wavefront is perpendicular to the $z$-axis), then the Huygens’s construction shows that plane radiation wavefront will be emitted along $z$-direction.
-Only solution of the electrodynamics equations in non-covariant form gives correct coupling between the usual particle trajectories and the fields in the lab frame.
-The non-covariant approach within the framework of both mechanics and electrodynamics predicts an effect in complete contrast to the conventional treatment.
-In the Chapter 16 we demonstrate that our Galilean transformed electrodynamics within the lab frame (see Eq.11) says that by making measurement on the coherent radiation, one can observe only radiation in the kicked direction.
-Experiments (involving specially kicked modulated electron beam) show that this prediction is, in fact, true.
-The results of XFEL experiments demonstrated that even the direction of emission of coherent undulator radiation is beyond the predictive power of the conventional theory.
-A starting point in our study of relativity was this complete disagreement between XFEL theory and experiment.
-At this point, a reasonable question arises: since XFELs are designed without accounting for the special relativity, how can they actually operate?
-All existing XFEL codes are based on a model in which the modulated electron beam moves only along the undulator axis.
-In a typical configuration for an XFEL, the orbit of a modulated electron beam is controlled to avoid large excursions from the undulator axis.
-However, random errors in the focusing system can cause angular trajectory errors (or ”kicks”).
-Analysis of the trajectory errors on the XFEL amplification process showed that any XFEL undulator magnetic field must satisfy stringent requirements.
-However, studies of this critical aspect in the design of a XFEL sources are based on an incorrect coupling of fields and particles.
-The pleasant surprise is that the tolerances predicted are more stringent than they need be according to our corrected XFEL theory.
-This can be considered as one of the reason for the exceptional progress in XFEL developments over last decades.
-It is generally assumed that the usual Maxwell’s equations and corrected Newton’s second law can explain all experiments that are performed in a single inertial (lab) frame.
-According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity.
-There are many physicists who have already received knowledge about special relativity from textbooks and who would say, ”For those who want to learn just enough about special relativity so they can solve problems - it just changes Newton’s laws by introducing a correction factor to the mass.” This is misconception.
-We cannot take old kinematics for mechanics and Einstein’s kinematics for electrodynamics.
-3.10 Bibliography and Notes
-1.
-According to the thesis of conventionality of simultaneity [13, 14, 15, 16, 17], simultaneity of distant events is a conventional matter, as it can be legitimately fixed in different manners in any given inertial reference frame.
-To quote e.g.
-Moeller [13]: ”All methods for the regulation of clocks meet with the same fundamental difficulty.
-The concept of simultaneity between two events in different places obviously has no exact objective meaning at all, since we cannot specify any experimental method by which this simultaneity could be ascertained.
-The same is therefore true also for concept of velocity.” Let us illustrate a particular special example of the simultaneity convention.
-Let the synchronization of clocks in different spatial points be provided by light signals having, respectively, velocity $c_{1}$ in the direction parallel to the positive axis $x$, and velocity $c_{2}$ in the opposite direction.
-Then, a light signal sent from point $A$ at time $t_{A}$ will arrive at point $B$ at time $t_{B}=t_{A}+x_{AB}/c_{1}$.
-The reflected signal will arrive back at point $A$ at time $t^{\prime}_{A}=t_{B}+x_{AB}/c_{2}$.
-From these two expressions we get $t^{\prime}_{A}-t_{A}=x_{AB}(c_{2}+c_{1})/(c_{1}c_{2})$.
-Summing up we have $t_{B}=t_{A}+c_{2}(t^{\prime}_{A}-t_{A})/(c_{1}+c_{2})$.
-So we come to the synchronization first proposed in [14]: $t_{B}=t_{A}+\epsilon(t^{\prime}_{A}-t_{A})$, where $\epsilon=c_{2}/(c_{2}-c_{1})$.
-After substituting in this expression Einstein choice $c_{1}=-c_{2}=c$ we get $\epsilon=1/2$.
-This time order, as fixed by the standard synchronization, is frame dependent.
-This is the well-known thesis of relativity of simultaneity.
-The conventional nature of distant simultaneity in special relativity is not to be confused with the relativity of simultaneity.
-Clearly, the conventionality of simultaneity within a single inertial frame is quite distinct from the relativity of simultaneity in Einstein’s synchronization.
-If $\epsilon\neq 1/2$, then speed of light from $A$ to $B$ differs from the speed of light from $B$ to $A$.
-Although the Einstein synchronization choice, is preferred by physicists, it is nothing more ”physical” than any other.
-A nice pedagogical (but artificial) example of a non-standard synchronization is clock synchronization [11] in which $\epsilon=0$.
-To any convention on the simultaneity there will correspond a definite choice of the coordinate system in an inertial coordinate system of reference of the Minkowski space-time.
-A particular very unusual choice of coordinates, the absolute time coordinate choice, is exploited in this book.
-In our case, when time is absolute, we have $c_{1}=c+v$, $c_{2}=-c+v$.
-Substituting into the expression for $\epsilon$ we obtain $\epsilon=(1/2)(1-v/c)$.
-2.
-In the text first published in 1923, Eddington discussed, apparently for the first time, a procedure for synchronization using slow transport of clocks [18].
-The details can be found in review [8] ( see also [19]).
-3.
-We already pointed that we have empirical access only to the round-trip average speed of light.
-An empirical test of any property of the one-way speed of light is not possible.
-Many authors of textbooks still attribute a reality status of the one-way speed of light.
-To quote Hrasko [20]: ”It is sometimes claimed that Einstein synchronization of distant clocks $A$ ans $B$ is circular.
-The argument is very simple: Einstein synchronization is based on the equality of light velocity on the path from $A$ to $B$ and back from $B$ to $A$, but measurement of light velocity in one direction between to distant points is impossible unless the clocks at these points have already been synchronized.
-This argument is, however, fallacious.
-It is true that one-way measurement of light velocity can be performed only if clocks at the endpoints are synchronized correctly.
-But since they need not show the correct coordinate time, they can be synchronized without light signals by transporting them from common site in a symmetrical manner.
-The procedure consists of following steps: … .
-As we see, the thought experiment described is capable to prove constancy of light speed if it is true, or to disprove it if it is false.
-It provided, therefore, solid logical foundation for Einstein’s synchronization prescription.” This logical argument is incorrect.
-Slow clock transport synchronization is equivalent to Einstein’s synchronization in inertial system where the light source is at rest.
-4. For a general discussion of correspondence between passive and active transformations we suggest reading the paper [21].
 4 Aberration of Light Phenomenon. Inertial Frame of Reference
-4.1 Introductory Remarks
 The effect of light aberration in an inertial frame of reference is usually understood as a change in the direction of light propagation ascribed to boosted light sources.
-The phenomenon of aberration of light in an inertial frame of reference by no means simple to describe, even in the first order in $v/c$: a large number of incorrect results can be found in the textbooks.
-Thus, in order to keep the mathematical complexity of the discussion to a minimum, we will describe the effect of aberration by working only up to the first order.
-The appearance of relativistic effects in optical phenomena does not depend on a large speed of the radiation sources.
-Light is always a relativistic object, no matter how small the ratio $v/c$ may be.
-The subject of this chapter is the physical influence of the optical instrument on the aberration of light measurement.
-First, we will demonstrate that when one has finite-aperture mirror moving transversely and the plane wave of light is falling normally on the mirror, there is the aberration (deviation of the energy transport) for light reflected from the mirror.
-The next problem to be discussed in this chapter is of more practical importance.
-According to the Babinet’s principle, the problem of reflection from a transversely moving (finite-aperture) mirror complementary to the problem of transmission through a hole in the moving opaque screen.
+The phenomenon of aberration of light in an inertial frame of reference is by no means simple to describe, even in the first order in $v/c$: a large number of incorrect results can be found in the textbooks.
+Thus, to keep the mathematical complexity of the discussion to a minimum, we will describe the effect of aberration by working only up to the first order.
+The subject of this section is the physical influence of the optical instrument on the aberration of light measurement.
+First, we will demonstrate that when one has a finite-aperture mirror moving parallel to itself and the plane wave of light falls normally on the mirror, there is an aberration (deviation of the energy transport) for light reflected from the mirror.
+The next problem to be discussed in this section is of more practical importance.
+According to Babinet’s principle, the problem of reflection from a tangentially moving (finite-aperture) mirror is complementary to the problem of transmission through a hole in the moving opaque screen.
 One case of rather great interest is that which corresponds to transmission through a moving end of the telescope barrel.
-It is shown that the (electromagnetic) wave theory approach to the aberration of light in an inertial frame leads to a substantially different result for energy transport than the predicted by conventional theory.
 4.1 The ”Plane Wave” Emitter
 We shall try to understand the effect in a very simple case.
-The explanation of the effect of aberration is actually based on the use of the model of single plane-wave emitter.
+The explanation of the effect of aberration is based on the use of the model of a single plane-wave emitter.
 As a simple model of a plane-wave emitter, we use a two-dimensional array of identical coherent elementary sources (dipoles), uniformly distributed on a given ($x-y$) plane $P$.
@@
 It is always hiddenly assumed that the detector for the direction of the radiation is an energy propagation detector and the size of the detector aperture is sufficiently large compared with the radiation beam size.
-Indeed, what is usually considered as an aberration is, in fact, an apparent deviation of the energy transport direction.
+Indeed, what is usually considered an aberration is, in fact, a deviation of the energy transport direction.
 4.2 A Moving Emitter. The ”Single Frame” Description
@@
 Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the energy transport measurement.
-There could be two approaches to the analysis of the aberration o light radiated by a single moving emitter.
+There could be two approaches to the analysis of the aberration of light radiated by a single moving emitter.
 The first one is the covariant approach.
-The explanation of the effect of aberration of light in the case of a single moving emitter presented in the literature is actually based on the use of a Lorentz boost to describe how the direction of a beam of light depends on the velocity of the light source relative to the inertial frame of reference.
-Another non-covariant approach consists in using a ”single frame” description without reference to Lorentz transformations.
+The explanation of the effect of aberration of light in the case of a single moving emitter presented in the literature is based on the use of a Lorentz boost to describe how the direction of a beam of light depends on the velocity of the light source relative to the inertial frame of reference.
+Another non-covariant approach consists of using a ”single frame” description without reference to Lorentz transformations.
 The two approaches, treated according to Einstein’s or absolute time synchronization conventions give the same result in the case of a single moving emitter.
 The choice between these two different approaches in this case is a matter of pragmatics.
-In this section we demonstrate both approaches.
-Let us start with non-covariant approach.
-We must emphasize that there is no principle difficulty with the a non-covariant (3+1) approach in relativistic electrodynamics.
+In this section, we demonstrate both approaches.
+Let us start with the non-covariant approach.
+We must emphasize that there is no principle difficulty with the non-covariant (3+1) approach in relativistic electrodynamics.
 It is perfectly satisfactory.
@@
 It means that simultaneity is absolute and there is no mixture of positions and time when sources change their velocities in the inertial frame.
-A distinctive trait of our non-covariant theory is the absence of relativistic kinematics in the description of aberration of light phenomena.
-When one has a transversely moving emitter there is the deviation of the energy transport for radiated light.
+A distinctive trait of our non-covariant theory is the absence of relativistic kinematics in the description of the aberration of light phenomena.
+When one has a tangentially (to its surface) moving emitter there is the deviation of the energy transport for radiated light.
 According to the ”single frame” approach, this effect is a consequence of the fact that the Doppler effect is responsible for the angular frequency dispersion of the radiated light waves (Fig.
 1 left).
-Figure 1: The case when the $(x-y)$ plane emitter in an inertial lab frame is accelerated from rest up to velocity $v$ along the $x$-axis.
-It is assumed that there is no physical influence of the detector for the direction of the radiation on the measurement.
+Figure 1: The case when the plane wave emitter in an inertial lab frame is accelerated from rest up to velocity $v$ along the $x$-axis.
 The aberration increment is connected with the physical parameters by the relation: $\theta_{a}=v/c$.
-The explanation of the effect of aberration is based on the use of a Galilean (left) and Lorentz (right) boost to describe how the direction of a light beam depends on the velocity of light source relative to the lab frame.
-If make a Lorentz boost, we automatically introduce a time transformation $t\to t+xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation phase front in the lab frame.
+The explanation of the effect of aberration is based on the use of a Galilean (left) and Lorentz (right) boost.
 Let us consider the electrodynamics of the moving source.
-The explanation of the phenomenon of radiation in our case of interest consists in using a (active) Galileo boost to describe the uniform translational motion of the source in the inertial lab frame.
+The explanation of the phenomenon of radiation in our case of interest consists of using a (active) Galileo boost to describe the uniform translational motion of the source in the inertial lab frame.
 Maxwell’s equations are not preserved in form by the Galilean transformation, i.e.
 Maxwell’s equations are not invariant under Galilean transformation.
-The new terms that have to be put into the field equations due to use of Galilean transformation lead to the prediction of the Doppler effect.
-One should not to be surprised to find that electrodynamics problem of moving emitter has intrinsic anisotropy.
-If fact, anisontropy results directly from the time-dependence of the transverse position of the moving emitter with finite aperture.
-What must be recognized is that in the time-dependent emitter problem, the results will depend on the direction of the velocity vector.
-The present approach to moving emitter problem uses the Fourier transform methods.
+The new terms that have to be put into the field equations due to the use of Galilean transformation lead to the prediction of the Doppler effect.
+The present approach to the moving emitter problem uses the Fourier transform method.
 When we are dealing with linear systems it is useful to decompose a complicated input into a number of more simple inputs, to calculate the response of the system to each of these elementary functions, and to superimpose the individual responses to find the total response.
-Fourier analysis provides the basic means for performing such a decomposition (1).
+Fourier analysis provides the basic means for performing such a decomposition  [^16] For a general discussion of the Fourier transform methods of spatial filtering theory or Abbe diffraction theory we suggest reading the book [22]..
 […]
-The emitter with finite aperture is a kind of active medium which breaks up the radiated beam into a number of diffracted beams of plane waves.
+The emitter with a finite aperture is a kind of active medium that breaks up the radiated beam into a number of diffracted beams of plane waves.
 Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved.
@@
 […]
-There is a different physical viewpoint on the Doppler effect of the light beam radiated from a moving emitter that is equivalent to the presented above.
+There is a different physical viewpoint on the Doppler effect of the light beam radiated from a moving emitter that is equivalent to the one presented above.
 The phase of the wave at the world point $(\vec{r},t)$ is invariant at a change of the reference system.
@@
 This frequency change coincides with the result derived directly from the dispersion equation, as must be.
-Let us first remind the reader of the fact that the usual velocity of waves is defined as given the phase difference between the oscillations observed at two different points in a free plane wave.
-It is primary used for computing interference fringes that makes phase differences visible.
-In a plane wave we observe the phase velocity $\omega/k$.
+Let us first remind the reader of the fact that the usual velocity of waves is defined as the phase difference between the oscillations observed at two different points in a free plane wave.
+It is primarily used for computing interference fringes that make phase differences visible.
+In a plane wave, we observe the phase velocity $\omega/k$.
 Another (group, or energy propagation) velocity can be defined, if we consider the propagation of a peculiarity, that is change in amplitude impressed on a train of waves.
-A simple combination of groups obtains when two waves $\omega_{1}=\omega+\Delta\omega,~{}k_{1}=k+\Delta k$ and $\omega_{2}=\omega-\Delta\omega,~{}k_{2}=k-\Delta k$ are superimposed.
+A simple combination of groups is obtained when two waves $\omega_{1}=\omega+\Delta\omega,~{}k_{1}=k+\Delta k$ and $\omega_{2}=\omega-\Delta\omega,~{}k_{2}=k-\Delta k$ are superimposed.
 This represents a carrier with frequency $\omega$ and a modulation with frequency $\Delta\omega$.
@@
 Many textbooks on electromagnetic theory discuss the aberration of light phenomena in the context of plane wave.
-However, in dealing with plane wave one will have an incorrect model of the aberration of light.
+However, in dealing with plane waves one will have an incorrect model of the aberration of light.
 When an infinite sinusoidal wave travels, there is a uniform average energy density throughout the space.
@@
 It is impossible to know this.
-All experimental methods for measuring the aberration of light operate with light signals, and hence do not measure the phase velocity but the signal velocity and this velocity coincides with group velocity.
-In our example, the plane waves with different wavenumber vectors propagate out from the moving emitter with the different frequencies.
-Then equation $\Delta\omega/\Delta k_{x}=v$ holds for each scattered waves independently on the sign and the magnitude of the radiated angle.
+All experimental methods for measuring the aberration of light operate with light signals, and hence do not measure the phase velocity but the signal velocity and this velocity (in our case of interest) coincides with the group velocity.
+In our example, the plane waves with different wavenumber vectors propagate out from the moving emitter with different frequencies.
+Then equation $\Delta\omega/\Delta k_{x}=v$ holds for each scattered wave independently on the sign and the magnitude of the radiated angle.
 In fact, the $\Delta\omega$ in our case of interest is the Doppler shift $\Delta\omega=\vec{K}_{\perp}\cdot\vec{v}$ and the $\Delta k_{x}$ is simply the transverse component of the radiated wavenumber vector $\Delta k_{x}=K_{\perp}$.
-The last equations state that radiated light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=v$.
+The last equations state that a radiated light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=v$.
 4.3 A Moving Emitter. Explanation on the Basis of the Lorentz Transformations
-It should be note, however, that there is another satisfactory way of explaining the effect of aberration of light from the moving source.
-The explanation consists in using a Lorentz boost to describe the uniform translation motion of the light source in the lab frame.
-On the one hand, the Maxwell’s equations remain invariant with respect to Lorentz transformations.
-On the other hand, if make a Lorentz boost, we automatically introduce a time transformation $t^{\prime}=t-xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation phase front in the inertial frame.
-This is because the effect of this time transformation is just a dislocation in the timing of processes, which has the effect of rotating the plane of simultaneity on the angle $v/c$ in the first order approximation.
+It should be noted, however, that there is another satisfactory way of explaining the effect of aberration of light from the moving source.
+The explanation consists of using a Lorentz boost to describe the uniform translation motion of the light source in the lab frame.
+On the one hand, Maxwell’s equations remain invariant with respect to Lorentz transformations.
+On the other hand, if make a Lorentz boost, we automatically introduce a time transformation $t^{\prime}=t-xv/c^{2}$, and the effect of this transformation is just a rotation of the radiation phase front in the inertial frame.
+This is because the effect of this time transformation is just a dislocation in the timing of processes, which has the effect of rotating the plane of simultaneity on the angle $v/c$ in the first-order approximation.
 According to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the radiation phase front.
@@
 In other words, when a uniform translational motion of the source is treated according to Lorentz transformations, the aberration of light effect is described in the language of relativistic kinematics.
-According to the relativistic kinematics, the extra phase chirp $d\phi/dx=k_{x}=v\omega_{i}/c^{2}$ is introduced and the array of identical elementary sources of the moving emitter now have different phases.
+According to the relativistic kinematics, the extra phase chirp $d\phi/dx=k_{x}=v\omega_{i}/c^{2}$ is introduced and the array of identical elementary sources of the moving emitter now has different phases.
 As a consequence of this, the plane wave wavefront rotates after the Lorentz transformation.
@@
 From the initial conditions and the Lorentz transformation we find that $\omega=(\omega_{i}+vK_{\perp})$, $k_{z}=\sqrt{\omega^{2}_{i}/c^{2}-K_{\perp}^{2}}$, $k_{x}=(v\omega_{i}/c^{2}+K_{\perp})$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated dipole density, $\omega_{i}$ is the frequency of the emitter radiation before the acceleration.
-Substituting these expressions in dispersion equation we find that the latter is satisfied, as must be.
-As one of the consequences of the Doppler effect in the Lorentz coordinatization, we find an angular frequency dispersion of the light waves radiated from the moving emitter with finite aperture.
-The Doppler shift, $\Delta\omega$, of radiated light wave (in the first order approximation) is given by $\Delta\omega=\vec{K}_{\perp}\cdot\vec{v}$, where $K_{\perp}$ is the transverse component of the radiated wavenumber vector.
-The last equation state that radiated light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=v$.
+Substituting these expressions in the dispersion equation we find that the latter is satisfied, as must be.
+As one of the consequences of the Doppler effect in the Lorentz coordinatization, we find an angular frequency dispersion of the light waves radiated from the moving emitter with a finite aperture.
+The Doppler shift, $\Delta\omega$, of a radiated light wave (in the first order approximation) is given by $\Delta\omega=\vec{K}_{\perp}\cdot\vec{v}$, where $K_{\perp}$ is the transverse component of the radiated wavenumber vector.
+The last equation states that a radiated light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=v$.
 It is interesting to discuss what it means that there are two different (covariant and non-covariant) approaches that produce the same group velocity.
-The point is that both approaches describe correctly the same physical reality and since the group velocity has obviously an objective meaning (i.e.
+The point is that both approaches describe correctly the same physical reality and since the group velocity has an objective meaning (i.e.
 convention-invariant), both approaches yield the same physical results.
 The difference between the absolute time coordinatization and the Lorentz coordinatization is very interesting.
-In the Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization.
-We can interpret manipulations with rule-clock structure in the lab frame simply as a change of the time variable according to the transformation $t\to t+xv/c^{2}$.
+In Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization.
+We can interpret manipulations with a rule-clock structure in the lab frame simply as a change of the time variable according to the transformation $t\to t+xv/c^{2}$.
 The overall combination of Galileo transformation and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the inertial frame.
@@
 As a consequence, we should not care to transform the results of the electrodynamics problem solution into the original (3+1) variables.
-An idea of studying the relativistic electrodynamics using technique involving a change of variables is useful from a pedagogical point of view.
-It is worth remarking that the absent of a dynamical explanation for wavefront rotation in the Lorentz coordinatization has disturbed some physicists.
-It should be clear from the discussion in the Chapter 3 that a good way to think of the wavefront rotation is to regard it as a result of transformation to a new time variable in the framework of the Galilean (”single frame”) electrodynamics.
-4.5 Reflection from a Mirror Moving Transversely. Mistake in Existing Theory
+The idea of studying relativistic electrodynamics using a technique involving a change of variables is useful from a pedagogical point of view.
+It is worth remarking that the absence of a dynamical explanation for wavefront rotation in the Lorentz coordinatization has disturbed some physicists.
+It should be clear from the discussion in Chapter 3 that a good way to think of the wavefront rotation is to regard it as a result of the transformation to a new time variable in the framework of the Galilean (”single frame”) electrodynamics.
+4.4 Reflection from a Mirror Moving Parallel to Itself. Mistake in Existing Theory
 It is generally believed that for a mirror moving tangentially to its surface the law of reflection which holds for the stationary mirror is preserved, as shown in Fig.2.
@@
 This statement, presented in most textbooks, is incorrect.
-First, we examine the reasoning presented in textbooks (2).
+First, we examine the reasoning presented in textbooks  [^17] It is generally believed that there is no aberration for light reflected from mirrors moving parallel to itself.
+To quote e.g.
+Sommerfeld [23]: ”Thus, for a mirror moving tangentially to its surface the law of reflection which holds for the stationary mirror is preserved.” Similar statements can also be found in other textbooks.
+To quote e.g.
+Ugarov [24]: ” Hence, when the mirror moves parallel to itself the frequency of incident light is equal to that of reflected light, and the angle of incidence is equal to the angle of reflection.”.
 The reflection from the mirror is analyzed in two Lorentz reference frames.
@@
 In this frame, the mirror is at rest.
-In both frames, we use a Cartesian coordinate system in which $x-y$ plane is tangent to the reflection surface.
+In both frames, we use a Cartesian coordinate system in which the $x-y$ plane is tangent to the reflection surface.
 The $x$ direction coincides with the direction of $v$.
 For simplicity, we consider the case in which light is incident from the $z$ direction in the lab frame.
-Incident light is described by its four-dimensional wave vector, whose time like component is the angular frequency $\omega$ and whose space like components define the direction of propagation.
+Incident light is described by its four-dimensional wave vector, whose time-like component is the angular frequency $\omega$ and whose space-like components define the direction of propagation.
 In the lab frame, $(t,x,y,z)$, this vector has components $k_{1}=(\omega,0,0,-\omega/c)$, where the negative sign indicates propagation towards the mirror (Fig.
@@
 Our task is to determine the wave vector for the reflected beam.
-Figure 2: Transversely moving mirror with small aperture at normal incidence.
+Figure 2: A mirror moves parallel to itself.
 According to textbooks, there is no deviation of the energy transport for the reflected light beam.
-A monochromatic plane wave of light is falling normally on the small moving aperture mirror, and generates a reflected oblique beam.
-Figure 3: The effect of aberration of light is described in the language of relativistic kinematics, in terms of the wavenumber four vector.
+A monochromatic plane wave of light falls normally on the small moving aperture mirror and generates a reflected oblique beam.
+Figure 3: The effect of aberration of light is described in the language of relativistic kinematics, in terms of the wavenumber four-vector.
 Geometry of the reflection as seen from (a) lab frame, (b) inertial frame moving with the same velocity as the mirror.
 According to textbooks, there is no aberration for light reflected from a tangentially moving mirror.
-The argument that there is no aberration for light reflected from a transversely moving mirror runs something like this.
+The argument that there is no aberration for light reflected from a tangentially moving mirror runs something like this.
 It is easiest to consider the reflection in the moving frame (Fig.
@@
 The effect of reflection is to reverse the sign of the $z^{\prime}$ component of the wave vector, $k^{\prime}_{2}=(\omega,-v\omega/c^{2},0,\omega/c)$.
-We now obtain the reflected wave vector in the lab frame by applying the inverse Lorentz transformation: $k_{2}=(\omega,0,0,\omega/c)$.This vector represents a light beam traveling away from the mirror, having the same frequency as the incoming beam.
+We now obtain the reflected wave vector in the lab frame by applying the inverse Lorentz transformation: $k_{2}=(\omega,0,0,\omega/c)$.
+This vector represents a light beam traveling away from the mirror, having the same frequency as the incoming beam.
 This shows that the beam is reflected according to the usual geometrical optics laws, and the beam suffers no aberration.
-There are two errors in solving problems concerning reflection from a mirror moving transversely.
-The first error in the last argument follows from the fact that the concept of a plane wave and an infinite plane mirror is used in the case of the tangential motion.
+There are two errors in solving problems concerning reflection from a mirror moving tangentially to its surface.
+The first error in the last argument follows from the fact that the concept of a plane wave and an infinite plane mirror is used in the case of tangential motion.
 The absurdity of this, i.e.
-the impossibility of such a motion ever been detected is put forward as an obvious argument against the reasoning presented in textbooks.
-Indeed, if the mirror is infinite the problem is not time dependent (we must conclude that when we are dealing with hidden assumption that problem is not time dependent we have no aberration and that is not too surprising).
-It is obvious that only the motion of a finite mirror is meaningful.
-In view of the electrodynamics, only a velocity of the finite mirror has a physical meaning.
+the impossibility of such a motion ever being detected is put forward as an obvious argument against the reasoning presented in textbooks.
+Indeed, if the mirror is infinite the problem is not time-dependent (we must conclude that when we are dealing with the hidden assumption that the problem is not time-dependent we have no aberration and that is not too surprising).
+Only the motion of a finite mirror is meaningful.
+In view of the electrodynamics, only the velocity of the finite mirror has a physical meaning.
 The reasoning in textbooks ignores completely the interaction of the light and the moving mirror edges.
-We would also noticed that the concept of plane wave is used in textbooks.
-There is a common mistake made in the electrodynamics connected with the energy transport direction in the case of an (infinite) plane wave.
-When an infinite plane wave travels, there is a uniform average energy density throughout the space.
-It is impossible to know the energy transport direction when one has deal with a plane wave.
-All experimental methods for measuring the aberration increment operate with light signals, and hence do not measure the phase velocity (i.e.
-frequency and wavenumber vector) but the group velocity.
-It can be defined only if we consider the propagation of a peculiarity, that is change in amplitude impressed on a train of waves.
 What authors of textbooks generally overlooked is the fact that the energy transport problem is well-defined only if the source and mirror apertures have already been specified.
@@
 Suppose that an observer, which is at rest with respect to the inertial frame performs the direction of the energy transport measurement.
-The observer describes the light beam from the emitter at rest using diagonal (Minkowski) metric.
+The observer describes the light beam from the emitter at rest using the diagonal (Minkowski) metric.
 According to textbooks, the metric of the moving mirror is the same (i.e.
-Minkowski metric) as metric of the emitter at rest.
-The Maxwell’s equations can be applied in the lab inertial frame only in the case when Lorentz coordinates are assigned.
-The authors of textbooks incorrectly belief that the common Lorentz time coordinate axis for emitter and mirror can be assigned.
-This is misconception.
-The question arises how to assign synchronization in the lab frame after the mirror acceleration.
+Minkowski metric) as the metric of the emitter at rest.
+Maxwell’s equations can be applied in the lab inertial frame only in the case when Lorentz coordinates are assigned.
+The authors of textbooks incorrectly believe that the common Lorentz time coordinate axis for the emitter and mirror can be assigned.
+This is a misconception.
+The question arises of how to assign synchronization in the lab frame after the mirror acceleration.
 The peculiarity of this problem with the viewpoint of relativistic kinematics is that here the emitter (and observer with his measuring devices) is at rest in the lab inertial frame and the mirror is moving with the constant speed with respect to the lab frame and interacts with the radiated light beam.
@@
 But this will be the absolute time coordinatization for the boosted mirror and this boost will be described in such coordinatization by the Galilean transformation.
-Without changing synchronization in the lab frame we can prepare, for mirror and emitter, a common set of synchronized clocks only in the case of absolute time coordinatization i.e in the case when simultaneity is absolute.
-Suppose that we re-synchronize clocks in the lab frame in order to assign the Lorentz time coordinate for the boosted mirror.
-In this coordinatization, we describe the reflection of light using the usual Maxwell’s equations.
+Without changing synchronization in the lab frame we can prepare, for mirror and emitter, a common set of synchronized clocks only in the case of absolute time coordinatization i.e.
+in the case when simultaneity is absolute.
+Suppose we re-synchronize clocks in the lab frame to assign the Lorentz time coordinate for the boosted mirror.
+In this coordinatization, we describe the reflection of light using Maxwell’s equations.
 On the other hand, the effect of this time transformation is just a rotation of the radiation phase front of the incoming plane wave on the angle $v/c$.
 As a result, this new time coordinate in the lab frame is interpreted by saying that Maxwell’s equations are not applicable to the emitter radiation description.
-It is not difficult to see that the peculiarity of this coordinatization is that here the energy transport velocity is not equal to the phase velocity for incoming light beam.
-This problem in special relativity can be adequately treated by an approach which uses the absolute time coordinatization.
-We will demonstrate that when one has a transversely moving mirror with finite aperture and a plane wave of light is falling normally onto the mirror, there is a deviation of the energy transport for the reflected light.
+It is not difficult to see that the peculiarity of this coordinatization is that here the energy transport velocity is not equal to the phase velocity for the incoming light beam (see footnote in Section 3.6).
+This problem in special relativity can be adequately treated by an approach that uses absolute time coordinatization.
+We will demonstrate that when one has a tangentially moving mirror with a finite aperture and a plane wave of light is falling normally onto the mirror, there is a deviation of the energy transport for the reflected light.
 In this case, the effect of aberration of light results directly from the time-dependence of the position of the moving mirror, since a finite aperture has to be considered.
-What must be recognized is that in the time-dependent emitter-mirror problem, the solution involves light beams with different frequencies.
-It comes out quite natural that the result will depend on the velocity vector: in other words, we account to the influence of the Doppler effect.
-For a transversely moving mirror with finite aperture we cannot neglect the angular frequency dispersion, which is an effect of the first order in $v/c$.
-We already pointed that Chapter 3 is probably the most important chapter in this book.
-To our knowledge, neither operational interpretation of the absolute time coordinatization nor the difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view, are given elsewhere in the literature.
-Using the Galilean boost: we substitute $x\to x-vt$, while leaving time unchanged into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain metric Eq.
+What must be recognized is that in the absolute time coordinatization, the wave equation will depend on the velocity vector.
+It comes out quite naturally that the solution involves light beams with different frequencies.
+In other words, we account for the influence of the Doppler effect.
+For a tangentially moving mirror with finite aperture, we cannot neglect the angular frequency dispersion, which is an effect of the first order in $v/c$.
+We already pointed out that Section 3 is probably the most important section in this article.
+To our knowledge, neither the operational interpretation of absolute time coordinatization nor the difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view, are given elsewhere in the literature.
+Using the Galilean boost: we substitute $x\to x-vt$ while leaving time unchanged into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain metric Eq.
 11.
-This metric describes the electrodynamics of the moving mirror with viewpoint of the inertial observer measurements.
-There is an intuitively plausible way to understanding why authors of textbooks conclude that there is no aberration for light reflected from a transversely moving mirror.
+This metric describes the electrodynamics of the moving mirror from the viewpoint of the inertial observer measurements.
+There is an intuitively plausible way to understand why authors of textbooks conclude that there is no aberration for light reflected from a tangentially moving mirror.
 A comparison between the conventional aberration of light theory and the sound theory might help here.
-Bellow in the Section 4.15 we shall consider the sound emitter at rest in the atmosphere frame.
+In Section 4.13 we shall consider the sound emitter at rest in the atmosphere frame.
 We shall consider the case of a mirror, in the atmosphere frame, moving with velocity $v$ along its surface.
-According to the theory of sound, there is no deviation of the energy transport for reflected beam.
+According to the theory of sound, there is no deviation of the energy transport for the reflected beam.
 The group velocity is equal to the phase velocity.
-A plane wave is falling normally on the mirror, and generates an oblique sound beam.
-The (diagonal) wave equation of the moving mirror in the atmosphere frame is the same as wave equation of the emitter at rest.
-According to textbooks, there are important parallels between aberration of sound and light in the first order approximation.
-This is misconception.
-There is a fundamental difference between velocity of light and velocity of sound.
+A plane wave falls normally on the mirror and generates an oblique sound beam.
+The (diagonal) wave equation of the moving mirror in the atmosphere frame is the same as the wave equation of the emitter at rest.
+According to textbooks, there are important parallels between the aberration of sound and light in the first-order approximation  [^18] We should make one further remark about two errors in solving problems concerning reflection from a mirror moving tangentially to its surface.
+These errors are not independent.
+According to textbooks, the metric of the moving mirror is diagonal as the metric of the ”plane wave” emitter at rest, and the group velocity is equal to the phase velocity.
+So independently of the mirror size, there is no deviation of the energy transport for the reflected light beam..
+This is a misconception.
+There is a fundamental difference between the velocity of light and the velocity of sound.
 The satisfactory treatment of the aberration of light phenomena should be based on the pseudo-Euclidean space-time geometry.
@@
 It uses Fourier transforms of spatial filtering theory that is the Abbe diffraction theory.
-The essence of Abbe’s approach, in our case of interest, is that one regards the mirror as a kind of a diffraction grating which breaks up the incident beam of the plane wave into a number of diffracted beams constituted by plane waves.
+The essence of Abbe’s approach, in our case of interest, is that one regards the mirror as a kind of diffraction grating that breaks up the incident beam of the plane wave into a number of diffracted beams constituted by plane waves.
 Each of these beams corresponds to one of the Fourier components into which the reflected power of the mirror can be resolved.
@@
 The Bragg condition $\vec{k}_{s}=\vec{k}_{i}\pm\vec{K}_{\perp}$ shows how the direction of the incident and scattered wave are related.
-The first-order maxima dominate due to the fact that light is being scattered from a sinusoidal grating, rather than a set of discrete planes (grooves).
+The first-order maxima dominates because light is being scattered from a sinusoidal grating, rather than a set of discrete planes (grooves).
 We assume that the $\vec{K}_{\perp}$ vector is directed parallel to the side of the diffraction grating with the incident wave impinging on the grating perpendicularly, as shown in Fig.4.
@@
 This fact is interpreted as a Doppler effect, since the reflection is from a moving, rather than a stationary, set of waves.
-In the case of a transversely moving grating, light in the diffraction maxima undergoes a Doppler shift resulting from the fact that it has been reflected from moving waves with wavenumber vectors $\vec{K}_{\perp}$.
-Let us demonstrate that the new terms that have to put into the field equations due to use of absolute time coordinatization lead to the prediction of the Doppler effect.
+In the case of a tangentially moving grating, light in the diffraction maxima undergoes a Doppler shift resulting from the fact that it has been reflected from moving waves with wavenumber vectors $\vec{K}_{\perp}$.
+Let us demonstrate that the new terms that have to be put into the field equations due to the use of absolute time coordinatization lead to the prediction of the Doppler effect.
 We recall that with a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ dispersion equation in the absolute time coordinatization becomes: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c+k_{z}^{2}-\omega^{2}/c^{2}=0$.
@@
 In fact, $k_{z}=\sqrt{\omega^{2}_{i}/c^{2}-k_{x}^{2}}$, $k_{x}=K_{\perp}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated reflectance.
-From this dispersion equation, we find the requirement that the wavenumber $K_{\perp}$ and the frequency change $\Delta\omega$ are related by $\Delta\omega=K_{\perp}v$.
-The following important detail of such ”single inertial frame” description can hardly be emphasize enough.
+From this dispersion equation, we find the requirement that the wavenumber $K_{\perp}$ and the frequency change $\Delta\omega$ are related by $\Delta\omega=K_{\perp}v$  [^19] We shall now give one more derivation of this same interesting and important result.
+We note the very close correspondence between angular and frequency response of moving diffraction gratings and the Raman scattering [25].
+Since the theory of Bragg diffraction applies to light scattering by sound waves in liquids and solids, it is not surprising that we can obtain our results from the quantum theory of light scattering by phonons.
+This scattering process is also known as Brillouin scattering.
+The quantum theory leads to equation $\vec{k}_{s}=\vec{k}_{i}\pm\vec{K}_{\perp}$ by the requirement that the momentum is conserved between interacting photon and phonon.
+The momentum balance for scattering involving the emission or absorption of phonon leads to the Bragg condition.
+The frequency shift that occurs for light scattering from sound waves comes about as a result of the Doppler effect.
+Quantum mechanically, it is a consequence of the conservation of energy between the participating particles.
+In contrast, authors of special relativity textbooks conclude that there is no Doppler effect for light reflected from a tangentially moving grating.
+They discuss about a tangentially moving mirror but it is also true for grating.
+The difficulty we speak of is associated with the generally accepted assumption that the metric of the moving grating (mirror) is the same (i.e.
+Minkowski metric) as the metric of the emitter at rest.
+.
+The following important detail of such a ”single inertial frame” description can hardly be emphasized enough.
 If the source of light is at rest and the mirror is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves.
@@
 4).
-In our previous discussion of absolute time coordinatization we learned that the emitter at rest must be in the same time described by Maxwell’s electrodynamics.
+In our previous discussion of absolute time coordinatization, we learned that the emitter at rest must be at the same time described by Maxwell’s electrodynamics.
 A dispersion equation in the case of Maxwell’s electrodynamics is reduced to $k_{i}^{2}-\omega_{i}^{2}=0$.
 From the initial conditions we find that $\vec{k}_{i}=\vec{e}_{z}k_{z}$, $\omega_{i}=ck_{z}$.
-The contradiction, however disappears if we perform geometrical analysis of light reflection.
+The contradiction, however, disappears if we perform a geometrical analysis of light reflection.
 The peculiarity of the discussed geometry is that even after the Galilean transformation along the $x$-axis the dispersion equation in the absolute time coordinatization will have the same (diagonal) form $k_{z}^{2}-\omega^{2}=0$ for the incident wave.
-We worked out a very simple case in order to understand all the physical principle very clearly.
-In Section 5.11 we shall analyze the aberration of light with incoming wave which has rate of change of phase with respect to $x$-axis.
+We worked out a very simple case to understand all the physical principles very clearly.
+In Section 5.10 we shall analyze the aberration of light with an incoming wave which has a rate of change of phase with respect to the $x$-axis.
 […]
-Figure 5: Transversely moving mirror with a small aperture at normal incidence.
-When a plane wave of light is falling normally on the mirror, there is a deviation of the energy transport for the reflected light beam.
-This effect is a consequence of the fact that the Doppler effect is responsible for angular frequency dispersion of the light waves reflected from the mirror.
-As a result, the velocity of the energy transport is not equal to the phase velocity.
-One of the most important conclusions of the foregoing discussion is a remarkable prediction on the theory of the aberration of light, concerning the deviation of the energy transport for light reflected from a mirror moving transversely.
-Namely, when a plane wave of light is falling normally on the mirror, there is a deviation of the energy transport for reflected light beam (see Fig.
+Figure 5: Tangentially moving mirror with a small aperture at normal incidence.
+When a plane wave of light falls normally on the mirror, there is a deviation of the energy transport for the reflected light beam.
+The explanation of the effect of aberration is based on the use of absolute time coordinatization.
+One of the most important conclusions of the foregoing discussion is a remarkable prediction on the theory of the aberration of light, concerning the deviation of the energy transport for light reflected from a mirror moving tangentially to its surface.
+Namely, when a plane wave of light (in the absolute time coordinatization) falls normally on the moving mirror, there is a deviation of the energy transport for the reflected light beam (see Fig.
 5).
-This phenomenon can be regarded as a simple consequence of the Doppler effect.
-The argument that in the process of reflection from a transversely moving mirror the direction of propagation is not given by the normal to the wavefront is considered erroneous in literature (4).
+This phenomenon can be regarded as a simple consequence of the Doppler effect [^20] Doppler effect is just a mathematical description in an intermediate step: when a complete calculation is made, only the geometrical (aberration) effect takes place..
+The argument is that in the process of reflection from a tangentially moving mirror the direction of propagation is not given by the normal to the wavefront is considered erroneous in literature  [^21] There is a common misconception that the radiation wavefront orientation has objective meaning.
+To quote Norton [26]: ”One might try to escape the problem by supposing that the direction of propagation is not always given by the normal to the wavefront.
+We might identify the direction of propagation with the direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation.
+Whatever may be the merits of such proposals, they are unavailable to some trying to implement a principle of relativity.
+If the direction of propagation of a plane wave is normal to the wavefronts in one inertial frame then that must be true in all inertial frame.”.
 This fact is ascribed to a lack of understanding of the difference between convention-dependent and convention-invariant parts of the theory.
@@
 However, the phase front orientation (i.e.
-the plane of simultaneity in the judgement of an observer) has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this orientation could be ascertained.
+the plane of simultaneity in the judgment of an observer) has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this orientation could be ascertained.
 Since the phase front orientation does not exist as physical reality within the angular range $v/c$, a question arises: why do we need to account for the exact phase front orientation in our electrodynamics calculations?
@@
 A comparison with a gauge transformation in Maxwell’s electrodynamics might help here.
-Even if the phenomena are quite different, the common mathematical formulation permits us to draw this analogue.
 4.6 Discussion
-Physicists who try to understand the situation related to the use of the textbook approach in the aberration of light phenomena, are often troubled by the fact that in the situation where there is a mirror moving normally to its surface the reasoning presented in textbooks is correct.
+Physicists who try to understand the situation related to the use of the textbook approach in the aberration of light phenomena are often troubled by the fact that in the situation where there is an infinite mirror moving normally to its surface the reasoning presented in textbooks is correct.
 The reflection from the mirror is analyzed in two Lorentz reference frames.
-It is interesting to note that, for this case, relativistic kinematics correctly predicts the light frequency variation on reflection from a moving mirror.
-At this point a reasonable question arises: why the same method gives incorrect result in the case of reflection of light from a mirror moving tangentially to its surface?
-It is not difficult to see that the peculiarity of the situation where there is an infinite mirror moving normally to its surface is that here the ”mirror motion” is a real observable effect.
+In this case, relativistic kinematics correctly predicts the light frequency variation on reflection from a moving mirror.
+At this point, a reasonable question arises: why does the same method give incorrect results in the case of the reflection of light from a mirror moving tangentially to its surface?
+It is not difficult to see that the peculiarity of the normally moving (mirror) source is that here the velocity is perpendicular to the plane of oscillating elementary sources (dipoles).
+Thus, the plane of simultaneity in the absolute time coordinatization will have the same orientation for the Lorentz coordinatization.
+In the case of collinear geometry, a motion of the mirror, according to the theory of relativity, influences the relativistic kinematics terms of the second order only.
+Let us now consider a frequency measurement in the context of special relativity.
+Suppose that we have a Fabry-Perot interferometer (or a grating spectrometer).
+The frequency measured in this case is (standing wave) wavelength.
+The same statement is also true for all cases of light frequency measurements.
+Frequency measurements automatically obey the relationship between light frequency and interference.
+In contrast, the deviation of the energy transport direction is a geometrical effect.
+There is a fact governing all interference phenomena.
+It states that the interference pattern is independent of the chosen metric.
+In other words, according to the space-time geometry, phase is 4D invariant.
+For collinear geometry, Minkowski metric and metric Eq.
+11 give the same prediction (see Chapter 13 for more details).
+The next thing that we have to discuss is the mirror size.
+The second peculiarity of the situation where there is an infinite mirror moving normally to its surface is that here the ”mirror motion” is a real observable effect.
 Indeed, even if the mirror is infinite the problem is time dependent.
-In this situation, a velocity of the infinite mirror has physical meaning.
-Thus, for normal motion, the infinite mirror may be a useful concept for analysis of the Doppler effect.
-This method was incorrectly extended to the case of tangential motion.
-The second peculiarity of the collinear geometry where there is a mirror moving normally along the same line as the radiated beam is that here the velocity is perpendicular to the plane of radiation wavefront (i.e plane of simultaneity).
-Thus, for collinear motion, the plane of simultaneity in the absolute time coordinatization will have the same orientation for the Lorentz coordinatization.
-We shall discuss this subject further in the Chapter 10.
+In this situation, the velocity of the infinite mirror has physical meaning.
+Thus, for normal motion, the infinite mirror may be a useful concept for the analysis of the Doppler effect.
 4.7 Large Aperture Mirror
-In this chapter our discussion is limited to the region of problem parameters, in which we forget about the emitter edges.
-Although this aberration of light theory is just an approximation, it is a very great importance practically.
+In this section, our discussion is limited to the region of problem parameters, in which we forget about the emitter edges.
+Although this aberration of light theory is just an approximation, it is of very great importance practically.
 We shall also discuss the situation where the transverse size of the emitter is very small relative to the transverse size of the moving mirror.
-It should be note that this situation is not realized in the stellar aberration measurements.
-We shall work out this case in order to understand all the physical principles very clearly.
+It should be noted that this situation is not realized in the stellar aberration measurements.
+We shall work out this case to understand all the physical principles very clearly.
 It is easy to show that the deviation of the energy transport is absent in this case, Fig.
 6.
-The direct approach to moving mirror problem uses the Fourier transform methods.
-We now consider another method of calculating the aberration of light effect - we want to illustrate the great variety of possibilities.
-The way of thinking that made the law about the behavior of light reflected from a large aperture mirror evident is based on Babinet’s principle.
-It is well known that, when light comes through a hole of a given shape, made in an opaque screen, the distribution of intensity after the hole (i.e.
-the diffraction pattern) is the same as in the case when the hole is replaced by sources (dipoles) uniformly distributed over the hole.
-In other words, the diffracted plane wave from a hole, or from a source with the same shape of the hole are the same.
-Figure 6: Transversely moving mirror with a large aperture at normal incident.
-When a beam of light is falling normally on the mirror, there is no deviation of the energy transport for the reflected light beam.
-The velocity of the energy transport is equal to the phase velocity.
-This is a particular case of Babinet’s principle, which states that the sum of diffraction fields behind two complementary opaque screens is the incident wave.
-We know from this principle, that the solution we have found using Abbe’s approach also corresponds to that for large aperture hole in a moving opaque screen.
-We see clearly that there is no electromagnetic interaction of a light beam with screen.
-Indeed light beam is not scattered by the hole edges.
-Does this discussion about large aperture hole have any meaning?
-To see whether it does, we should remember about the Babinet’s principle.
-Here we only wish to show how easy the law of reflection from a large aperture mirror can be found with the help of the Babinet’s principle.
+The direct approach to the moving mirror problem uses the Fourier transform method.
+Figure 6: Tangentially moving mirror with a large aperture at normal incident.
+When a beam of light falls normally on the mirror, there is no deviation of the energy transport for the reflected light beam.
 4.8 Analysis of Transmission through a Hole in a Opaque Screen
-Above we demonstrated that when one has a small aperture mirror moving transversely and the plane wave of light is falling normally on the mirror, there is the aberration (deviation of the energy transport) for light reflected from the mirror.
+Above we demonstrated that when one has a finite aperture mirror moving tangentially to its surface and the plane wave of light falls normally on the mirror, there is an aberration (deviation of the energy transport) for light reflected from the mirror.
 The problem to be considered in this section is of more practical importance.
@@
 7.
-However, there is a common mistake made in relativistic optics, connected with aberration effects from a transversely moving screen containing a hole.
+However, there is a common mistake made in relativistic optics, connected with aberration effects from a tangentially moving screen containing a hole.
 We describe the system using, again, a Fourier transform method similar to that considered above.
-The screen containing a hole is a kind of diffraction grating which breaks up the incident beam of the plane wave into a number of diffracted beams of plane waves.
+The screen containing a hole is a kind of diffraction grating that breaks up the incident beam of the plane wave into a number of diffracted beams of plane waves.
 Each of these beams corresponds to one of the Fourier components into which a transmitted light beam can be resolved.
@@
 Figure 7: Aberration of light in an inertial frame of reference.
-Transversely moving screen which has a hole in it.
-According to textbooks, a monochromatic plane wave of light is falling normally on the screen and generates a transmitted oblique beam.
-There is no deviation of the energy transport for the transmitted oblique light beam.
-The velocity of the energy transport is equal to the phase velocity.
+Tangentially moving screen which has a hole in it.
+According to textbooks, a monochromatic plane wave of light falls normally on the screen and generates a transmitted oblique beam.
 Figure 8: The Bragg diffraction grating at the normal incident. The transmittance is sinusoidally space-modulated.
 Figure 9: Aberration of light in an inertial frame of reference.
-Transversely moving screen which has a hole in it.
-A monochromatic plane wave of light is falling normally on the screen and generates a transmitted light beam.
-The Doppler effect is responsible for the angular frequency dispersion of the light waves transmitted through the hole.
-As a result, the velocity of the energy transport is not equal to the phase velocity.
+Tangentially moving screen which has a hole in it.
+A monochromatic plane wave of light falls normally on the screen and generates a transmitted light beam.
+The explanation of the effect of aberration is based on the use of absolute time coordinatization.
 According to our approach, there is a remarkable prediction of the theory of aberration of light concerning the deviation of the energy transport for light transmitted through a hole in a moving screen.
-Namely, when one has a transversely moving screen with a hole in it and a plane wave of light is falling normally on the screen, there is a deviation of the energy transport for light transmitted through the hole (see Fig.
+Namely, when one has a tangentially moving screen with a hole in it and a plane wave of light is falling normally on the screen, there is a deviation of the energy transport for light transmitted through the hole (see Fig.
 9).
 4.9 Spatiotemporal Transformation of the Transmitted Light Beam
-Let us suppose that transmitted light pulse propagates in the $x-z$ plane.
+Let us suppose that the transmitted light pulse propagates in the $x-z$ plane.
 Now we are interested in the space-time intensity distribution in this plane.
-Spatiotemporal coupling arises naturally in transmitted radiation behind the screen, because the transmission process involves the introduction of an angular-frequency dispersion of the transmitted radiation.
+Spatiotemporal coupling arises naturally in transmitted radiation behind the screen because the transmission process involves the introduction of an angular-frequency dispersion of the transmitted radiation.
 The emitted light beam is represented with sufficient accuracy as the product of factors separately depending on space and time.
-However, when the manipulation of the emitted light requires the transmission through a hole in a moving opaque screen, such assumption fails.
+However, when the manipulation of the emitted light requires the transmission through a hole in a moving opaque screen, such an assumption fails.
 […]
 4.10 Applicability of the Ray Optics
-Let us move on to consider the predictions of the existing aberration of light theory in the case of a transversely moving screen which has a hole in it (Fig.
-7).
-According to conventional theory, the spot of the transmitted light beam on the observer screen also moves with the same velocity as the moving screen.
-It is important at this point to emphasize that the electrodynamics dictates that this would also lead to a consequent introduction of an angular-frequency dispersion.
-However, such angular spectrum change would mean a correction to a deviation of the energy transport direction of the transmitted light beam so that there is a glaring conflict with the prediction of the energy transport direction according to conventional aberration of light theory.
-The absence of the group velocity along the moving direction is the prediction of conventional aberration of light theory and is obviously absurd from the viewpoint of electrodynamics.
-Figure 10: Transverse moving screen which has a hole in it. Light corpuscles are falling normally on the screen and, according to literature, generates an oblique light beam.
-This incorrect statement is a straightforward consequence of the generally accepted way of looking at the aberration of light phenomena of most authors of the texbooks.
-Today one is told that the phenomenon of aberration of light could be interpreted, using corpuscular model of light.
-Light corpuscles are falling normally on the moving screen and, according to literature, generate oblique light beam as Fig.
+Today one is told that the phenomenon of aberration of light could be interpreted, using a corpuscular model of light.
+Light corpuscles falls normally on the moving screen and, according to literature, generate oblique light beam as Fig.
 10 shows.
@@
 If the optical system is spatially coherently illuminated, then a satisfactory treatment of the aberration of light should be based on the electromagnetic wave theory.
-Some experts believe that the applicability of corpuscular model in the theory of light should be reinterpreted as the applicability of ray optics.
+Figure 10: Tangentially moving screen which has a hole in it. Light corpuscles fall normally on the screen and, according to literature, generate an oblique light beam.
+Some experts believe that the applicability of the corpuscular model in the theory of light should be reinterpreted as the applicability of ray optics.
 Let us see what happens, according to the ray optics, in our case of interest.
-Light rays are falling normally on the moving screen and generate oblique ray beam as Fig.
-10 shows (6).
+Light rays fall normally on the moving screen and generate oblique ray beam as Fig.
+10 shows  [^23] To quote Brillouin [27]: ” Fig.
+5 explain the situation assuming a simplified device consisting of the parallel plate moving with uniform velocity $v$ in the horizontal direction.
+Monochromatic light is falling normally on the plate and generates an oblique ray.” This oblique effect is demonstrated in Fig.
+10..
 In this situation, we just treat light rays like little particles and the effect is entirely familiar.
 We would like now to discuss the region of applicability of ray optics.
-The situation relating to use the ray optics in the theory of aberration of light is complicated.
-One could naively expect that the region of applicability of ray optics, following from the textbooks reasoning, should be identified with any spatially incoherent radiation.
+The situation relating to the use of ray optics in the theory of aberration of light is complicated.
+One could naively expect that the region of applicability of ray optics, following the textbook’s reasoning, should be identified with any spatially incoherent radiation.
 However, incorrect results are obtained by doing so.
 In particular, a spatially completely incoherent source (e.g.
-an incandescent lamp or a star) is actually a system of elementary (statistically independent) point sources with different offsets.
-Radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individual elementary point sources.
+an incandescent lamp or a star) is a system of elementary (statistically independent) point sources with different offsets.
+A radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individual elementary point sources.
 An elementary source produces in front of a hole aperture effectively a plane wave.
-In other words, the transmission process involves the introduction of an angular frequency dispersion of the transmitted radiation.
-It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single point source characteristics like a deviation of the energy transport direction.
-This argument gives reason why ray optics is not applicable in the theory of aberration of light from (spatially) completely incoherent sources.
+In other words, the transmission process involves the unavoidable diffraction of a plane wave by an aperture.
+It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single-point source characteristics like a deviation of the energy transport direction.
+This argument gives the reason why ray optics is not applicable to the theory of aberration of light from (spatially) completely incoherent sources.
 We will illustrate the applicability of ray optics in the theory of aberration of light for a particular class of spatially incoherent light beams.
-In fact, the ray beam shown in Fig.
-10 can be realized as follows.
-Such beam may be produced by many commonly used lasers with a random spread of phases.
-One of the very useful properties of laser sources is their ability to produce fields that are highly directional.The intensity of such fields is concentrated in a very narrow solid angle.
-A method can be proposed for generating ray beam from primary sources by the use of array of randomly phased lasers.
-Such planar source generates rays which are falling normally on the moving screen (within the laser Rayleigh range) and generate oblique transmitted ray beam, Fig.
+A method can be proposed for generating ray beam from primary sources by the use of an array of randomly phased lasers.
+Such a planar source generates rays which are falling normally on the moving screen (within the laser Rayleigh range) and generate an oblique transmitted ray beam, Fig.
 10.
-Intuitively, a transversely moving screen containing a hole acts like a switcher for lasers.
-Surely, a luminous spot moving at velocity $v$ can be realized simpler, so to speak, ”manually”.
-We can arrange the laser-like sources along the $x$ axis and switch them on one after another (independently) from left to right with a given time lag.
+Intuitively, a tangentially moving screen containing a hole acts like a switcher for lasers.
+Surely, a luminous spot moving at velocity $v$ can be realized more simply, so to speak, ”manually”.
+We can arrange the laser-like sources along the $x$ axis and switch them one after another (independently) from left to right with a given time lag.
 Naturally, we can get a luminous spot moving at any velocity (even at $v>c$).
-From this example it is seen that in this process no information can be transmitted (along the $x$ axis) since each source radiates independently.
-4.12 Measure of Aberration
-[…]
+From this example, it is seen that in this process no information can be transmitted (along the $x$ axis) since each source radiates independently.
 4.11 Moving Large Aperture Emitter
 Let us now consider the case when a ”plane-wave” emitter in the lab inertial frame is accelerated from rest up to velocity $v$ along the $x$ axis.
-An emitter with finite aperture is a kind of active medium which breaks up the radiated beam into a number of diffracted beams of plane waves.
+An emitter with a finite aperture is a kind of active medium that breaks up the radiated beam into a number of diffracted beams of plane waves.
 Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved.
-We already know from our discussion from very beginning of this chapter that there is a deviation of the energy transport for the coherent light radiated by the transversely moving emitter, which is nothing else but well-known result: there is the aberration of light from the transversely moving emitter in the inertial frame of reference (Fig.
+We already know from our discussion at the very beginning of this section that there is a deviation of the energy transport for the coherent light radiated by the tangentially moving emitter, which is nothing else but a well-known result: there is the aberration of light from the tangentially moving emitter in the inertial frame of reference (Fig.
 1).
-The specific of our case of interest with the viewpoint of kinematics is that here the screen is at rest in the lab frame and the emitter is moving tangentially to its surface with constant speed with respect to the lab frame.
-For simplicity we shall assume that the transverse size of the moving ”plane-wave” emitter is very large relative to the transverse size of the hole in the screen.
+The specific of our case of interest from the viewpoint of kinematics is that here the screen is at rest in the lab frame and the emitter is moving tangentially to its surface with constant speed with respect to the lab frame.
+For simplicity, we shall assume that the size of the moving ”plane-wave” emitter is very large relative to the size of the hole in the screen.
 Suppose that an observer, which is at rest with respect to the screen performs the direction of the energy transport measurement, Fig.11.
 Figure 11: Aberration of light in an inertial frame of reference.
-The large aperture ”plane-wave” emitter moving tangentially to its surface.
-As one of the consequences of the Doppler effect, we find group velocity of the light waves radiated off a large aperture moving emitter.
-The screen is at rest and we have actually the problem of steady-state transmission.
-The transmitted light beam is going vertically because it has lost its horizontal (group) velocity component.
+The large aperture ”plane-wave” emitter moves tangentially to its surface.
+The screen is at rest and the aberration of light phenomenon is absent in this situation.
 If the screen is at rest and the source of light is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves.
-In our previous discussion we learned that the screen at rest must be described by Maxwell’s electrodynamics.
+In our previous discussion, we learned that the screen at rest must be described by Maxwell’s electrodynamics.
 In other words, the dispersion equation in the Lorentz coordinatization should be applied and kept in a consistent way for both incoming and scattered waves.
@@
 The peculiarity of the discussed geometry is that even after the Galilean transformation along the $x$ axis the dispersion equation in the absolute time coordinatization will have the same diagonal form $k^{2}_{z}-\omega^{2}=0$ for the (normal) incident wave.
-At close look at the physics of this subject shows that in the inertial lab frame, where the screen is at rest, we have actually the problem of steady-state transmission.
+A close look at the physics of this subject shows that in the inertial lab frame, where the screen is at rest, we have the problem of steady-state transmission.
 The Doppler effect is absent and the transmitted beam is going vertically because it has lost its horizontal (group velocity) component.
@@
 We only wish to emphasize here the following point.
-When the light passes through the small aperture hole we have a light beam whose fields have been perturbed by diffraction, and now not include information about emitter motion.
+When the light passes through the small aperture hole we have a light beam whose fields have been perturbed by diffraction, and now do not include information about emitter motion.
 4.12 A Point Source in an Inertial Frame of Reference
@@
 (b) A point-like ( or, more generally, spatially completely incoherent) source
-Source field diffraction can be divided into categories - the Fresnel (near-zone) diffraction and Fraunhofer (far-zone) diffraction.
+Source field diffraction can be divided into categories - the Fresnel (near-zone) diffraction and the Fraunhofer (far-zone) diffraction.
 In Fraunhofer diffraction, the phase of the wave is assumed to vary linearly across the detector aperture.
 This would occur if, for example, a plane wave were incident on the aperture at an angle with respect to the optical axis.
-In the Fresnel diffraction, we replace the assumption of a linear phase variation with quadratic phase variation.
+In the Fresnel diffraction, we replace the assumption of a linear phase variation with a quadratic phase variation.
 In the far zone, both types of sources produce in front of pupil detection effectively a plane wave.
@@
 Figure 12: Aberration of light in an inertial frame of reference.
-A point source produces in front of a hole aperture effectively a plane wave.
-If the motion of the point source is parallel to the screen, transmitted beam is going vertically.
+In Fraunhofer diffraction region, a point source produces in front of a hole aperture effectively a plane wave.
 The aberration of light phenomenon is absent in this situation.
@@
 One of the important conclusions of the discussion presented above (see Fig.
-11 ) is that the aberration of point source is absent in an inertial frame, Fig.
+11 ) is that the aberration of a point source is absent in an inertial frame, Fig.
 12.
-There are a number of remarkable effects which are a consequence of the fact that the information about a point source motion is not included into the light beam transmitted through the hole.
-In fact, this is the key to the binary star paradox discussed in the Chapter 6.
+There are a number of remarkable effects that are a consequence of the fact that the information about a point source motion is not included in the light beam transmitted through the hole.
+In fact, this is the key to the binary star paradox discussed in Chapter 6.
 For the rest of our discussion of the aberration of light phenomena in an inertial frame, it will be more convenient if we consider a somewhat modified source of the ”plane wave” emitter type.
-We have chosen the model described in the Section 4.2 because it is relatively simple.
-It is sufficiently complicated that it can stand as a prototype which can be generalized for the description of the aberration of light phenomena.
-No one has ever done all of ”thought experiments” we described in just this way, but we know what would happen from the laws of special relativity which are, of course, based on other experiments.
+We have chosen the model described in Section 4.2 because it is relatively simple.
+It is sufficiently complicated that it can stand as a prototype that can be generalized for the description of the aberration of light phenomena.
+No one has ever done all of the ”thought experiments” we described in just this way, but we know what would happen from the laws of special relativity which are, of course, based on other experiments.
 Fig.
@@
 It is assumed that the order of magnitude of the dimensions of the ”point” source is about $\lambda$, where $\lambda$ is the optical radiation wavelength.
-At this point, a reasonable question arises: why we are not discuss about the aberration of light radiated by a single laser.
-At first glance, the output light beam shown in Fig.
-13 may be produced by a laser.
-Indeed, one of the specific properties of a laser is ability to produce coherent light beams that are highly directional.
-We will not be able actually to get into the subject until later.
-We will analyze the aberration of light transmitted from the laser resonator which is accelerated in an inertial frame in the Section 8.6 and 10.4.
 4.13 The Difference between the Aberration of Light and Sound
@@
 Sound is a periodic motion of the air caused by vibrations of parts of the source.
-The explanation of the effect of sound aberration is actually based on the use of the model of single plane-wave emitter.
+The explanation of the effect of sound aberration is based on the use of the model of a single plane-wave emitter.
 As a model, we use a two-dimensional array of identical coherent elementary sources uniformly distributed on a given $(x-y)$ plane $P$.
-We take the elementary sources to start radiation waves simultaneous.
+We take the elementary sources to start radiation waves simultaneously.
 We shall discuss the situation where there is a finite aperture mirror moving tangentially to its surface.
-For simplicity, we shall assume that the transverse size of the moving mirror is very small relative to the transverse size of the ”plane-wave” emitter.
-The observer describes the sound beam from the emitter at rest (in the atmosphere frame) using usual wave equation $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$, where $v_{s}$ is the sound speed.
+For simplicity, we shall assume that the size of the moving mirror is very small relative to the size of the ”plane-wave” emitter.
+The observer describes the sound beam from the emitter at rest (in the atmosphere frame) using the diagonal wave equation $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$, where $v_{s}$ is the sound speed.
 This emitter will radiate plane wavefront in a vertical $z$ direction.
-The wave equation of the moving mirror in the atmosphere frame is the same as wave equation of the emitter at rest.
-We may consider the amplitude of the beam radiated by plane of oscillating sources as a whole to be resultant of radiated spherical waves.
-This is because diagonal wave equation has no intrinsic anisotropy.
-The elementary sources lying on the plane of simultaneity gives rise to spherical radiated wavelets, and these combine according to Huygens principle to form what is effectively a radiated wave.
-The wavefront emitted by the plane-wave emitter at given instant time is constructed as an envelope of all the spherical wavelets from the point sources on the emitter that have radiated until this instant.
+The wave equation of the moving mirror in the atmosphere frame is the same as that of the emitter at rest.
+We may consider the amplitude of the beam radiated by the plane of oscillating sources as a whole to be the result of radiated spherical waves.
+This is because the diagonal wave equation has no intrinsic anisotropy.
+The elementary sources lying on the plane of simultaneity give rise to spherical radiated wavelets, and these combine according to Huygens’s principle to form what is effectively a radiated wave.
+The wavefront emitted by the plane-wave emitter at a given instant time is constructed as an envelope of all the spherical wavelets from the point sources on the emitter that have radiated until this instant.
 According to the theory of sound, there is no deviation of the energy transport for the reflected sound beam.
-A plane wave is falling normally on the finite aperture mirror, and generates a reflected oblique beam.
+A plane wave falls normally on the finite aperture mirror and generates a reflected oblique beam.
 Note that the phase and group velocities are always equal in the atmosphere frame.
@@
 7.
-In the next chapter (see section 5.14) we shall continue our discussion of the aberration of sound theory.
-The textbook approach to the aberration of light usually forces the physicist to believe that there are important parallels between the aberration of sound and light in the first order approximation.
-In fact, the wave equation of the moving mirror (screen) is the same (i.e.
-diagonal wave equation) as wave equation of the emitter at rest.
-On the other hand, in the sound theory we talk about Newton space and time rather than about space-time with pseudo-Euclidean geometry.
-We come to the conclusion that the standard analysis of the aberration of light in the single inertial frame does not take into account the fundamental difference between velocity of light and velocity of sound.
+In the next section, we shall continue our discussion of the aberration of sound theory.
+The textbook approach to the aberration of light usually forces the physicist to believe that there are important parallels between the aberration of sound and light in the first-order approximation.
+The wave equation of the moving mirror (screen) is the same (i.e.
+diagonal wave equation) as the wave equation of the emitter at rest.
+On the other hand, in the sound theory, we talk about Newton’s space and time rather than about space-time with pseudo-Euclidean geometry.
+We conclude that the standard analysis of the aberration of light in the single inertial frame does not take into account the fundamental difference between the velocity of light and the velocity of sound.
 The satisfactory treatment of the relativistic modification should be based on the space-time geometric approach.
-The usual Minkowski metric is valid in inertial frame where observer is at rest and the procedure of slow clock transport (or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest) is used to assign values to the time coordinate.
-In this (Lorentz) coordinate system Minkowski metric describes the electrodynamics of the light source at rest with viewpoint of the inertial observer measurements.
-Now we consider the case when mirror in the inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis.
+The usual Minkowski metric is valid in the inertial frame where the observer is at rest and the procedure of slow clock transport (or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest) is used to assign values to the time coordinate.
+In this (Lorentz) coordinate system Minkowski metric describes the electrodynamics of the light source at rest with the viewpoint of the inertial observer measurements.
+Now we consider the case when the mirror in the inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis.
 The method of synchronization consists in keeping the same set of synchronized clocks used.
 This synchronization convention preserves the Minkowski metric of the inertial observer and the light source at rest.
-In this case we are dealing with the Galilean boost of the mirror.
+In this case, we are dealing with the Galilean boost of the mirror.
 We substitute $x\to x-vt$, while leaving time unchanged into the Minkowski metric of the inertial observer to obtain Eq.
 11.
-This metric describes the electrodynamics of the moving mirror with viewpoint of the inertial observer measurements.
+This metric describes the electrodynamics of the moving mirror from the viewpoint of the inertial observer measurements.
 In contrast, the accepted in the previous literature incorrect assumption that an inertial observer (i.e.
-observer at rest in the inertial frame) and the accelerated ( in the same inertial frame) mirror have common Minkowski metric has always been considered obvious.
-So that the commonly accepted approach to special relativity in the first order approximation and sound theory are really just variants of the same classical theory.
-4.16 Bibliography and Notes
-1. For a general discussion of the Fourier transform methods of spatial filtering theory or Abbe diffraction theory we suggest reading the book [22].
-2.
-It is generally believed that there is no aberration for light reflected from mirrors moving transversely.
-To quote e.g.
-Sommerfeld [23]: ”Thus, for a mirror moving tangentially to its surface the law of reflection which holds for the stationary mirror is preserved.” Similar statements can also be found in other textbooks.
-To quote e.g.
-Ugarov [24]: ” Hence, when the mirror moves parallel to itself the frequency of incident light is equal to that of reflected light, and the angle of incidence is equal to the angle of reflection.”
-3.
-There is another interpretation that we can give to the scattering process.
-We note the very close correspondence between angular and frequency response of moving diffraction gratings and the Raman scattering [25].
-Since the theory of Bragg diffraction applies to light scattering by sound waves in liquids and solids, it is not surprising that we can obtain our results from the quantum theory of light scattering by phonon.
-This scattering process is also known as Brillouin scattering.
-The quantum theory leads to equation $\vec{k}_{s}=\vec{k}_{i}\pm\vec{K}_{\perp}$ by the requirement that the momentum is conserved between interacting photon and phonon.
-The process of Brillouin scattering is a special case of Raman scattering.
-The momentum balance for scattering involving the emission or absorption of phonon leads to the Bragg condition.
-The requirement of conservation of energy leads to the equation $\omega_{s}=\omega_{i}\pm\omega_{phon}$ with $\omega_{s}$, $\omega_{i}$, and $\omega_{phon}$ being the radian frequencies, of the scattered photon, the incident photon, and the phonon.
-The frequency shift that occurs for light scattering from sound waves comes about as a result of the Doppler effect.
-Quantum mechanically, it is a consequence of the conservation of energy between the participating particles.
-We come to the conclusion that, according to the hidden choice of absolute time coordinatization in non-relativistic quantum mechanics, the lab observer actually sees the radiation after reflection as a result of a Galilean boost rather than a Lorentz boost.
-4.
-There is a common misconception that the radiation wavefront orientation has objective meaning.
-To quote Norton [26]: ”One might try to escape the problem by supposing that the direction of propagation is not always given by the normal to the wavefront.
-We might identify the direction of propagation with the direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation.
-Whatever may be the merits of such proposals, they are unavailable to some trying to implement a principle of relativity.
-If the direction of propagation of a plane wave is normal to the wavefronts in one inertial frame then that must be true in all inertial frame.” This incorrect statement is a straightforward consequence of the generally accepted way of looking at special relativity of most physicists.
-Accepting the postulate on the constancy of the speed of light one also automatically assumes Lorentz coordinates.
-According to such limiting understanding of the theory of relativity, it is assumed that only Lorentz coordinatization must be used to map the coordinates of events.
-5.
-Note that spatiotemporal coupling is discussed in literature usually in relation with ultrashort laser pulse propagation through a grating monochromator.
-Ultrashort laser pulses are usually represented as a products of electric field factors separately dependent on space and time.
-However, when the manipulation of ultrashort laser pulses requires the propagation through a grating monochromator, such assumption fails.
-In this situation one has to consider in addition (to phase fronts), planes of constant intensity, that is pulse fronts.
-A pulse-front tilt can be present in the beam due to propagation through an optical setup incorporating dispersive optical element.
-In the grating monochromator case, the different spectral components of the out-coming pulse travel in different directions.
-The electric field of a pulse including angular dispersion can be expressed in the Fourier domain $[k_{x},\omega]$ as $\bar{E}(k_{x}-p\omega,\omega)$, while the inverse Fourier transform from the $[k_{x},\omega]$ to the space-time domain $[x,t]$ can be expressed as $E(x,t+px)$, which is the electric field of pulse with a pulse-front tilt.
-The tilt angle $\theta_{tilt}$ is given by $\tan\theta_{tilt}=cp$.
-More specifically $p=dk_{x}/d\omega=kd\theta_{D}/d\omega=\lambda/(c\theta_{D}d)$, where $\lambda=2\pi c/\omega$, $\theta_{D}$ is the diffracted angle, and $d$ is the groove spacing.
-The diffracted angle $\theta_{D}$ is a function of frequency, according to the well-known plane grating equation.
-Assuming diffraction into the first order, one has $\lambda=(\cos\theta_{i}-\cos\theta_{D})d$, where $\theta_{i}$ is the incident angle.
-By differentiating this equation one obtains $d\theta_{D}/d\lambda=1/(\theta_{D}d)$, where we assume for simplisity grazing incidence geometry, $\theta_{i}\ll 1$ and $\theta_{D}\ll 1$.
-The physical meaning of the former equation is that different spectral components of the out-coming pulse travel in different directions.
-Therefore one concludes that the pulse-front tilt is invariably accompanied by angular dispersion.
-It follows that any device like a grating monochromator, that produces an angular dispersion, also introduces significant pulse-front tilt.
-In our case of interest we have deal with light diffracted by a (transversely) moving set of gratings and the Doppler effect is responsible for frequency dispersion $d\omega/dk_{x}=v$.
-Thus the spatitemporal coupling due to the light beam transmission through a hole in a moving screen and the usual pulse-front tilt distortion are quite different.
-6.
-To quote Brillouin [27]: ” Fig.
-5 explain the situation assuming a simplified device consisting of parallel plate moving with uniform velocity $v$ in the horizontal direction.
-Monochromatic light is falling normally on the plate and generates an oblique ray.” This oblique effect is demonstrated in our Fig.
-10.
+observer at rest in the inertial frame) and the accelerated ( in the same inertial frame) mirror have a common Minkowski metric has always been considered obvious.
+So the commonly accepted approach to special relativity in the first-order approximation and sound theory are just variants of the same classical theory.
 5 Aberration of Light Phenomenon. Noninertial Frame of Reference
-5.1 Introductory Remarks
-In this section we reexamine the issue of light transmission in non-inertial frames with particular reference to the aberration of light phenomena.
-We derive the aberration for a pulse of light traveling in the accelerated systems using the Langevin metric in general relativity (1).
-It comes out naturally if one writes the equation of the time transfer, from the inertial frame to the accelerated frame, in a generally covariant context.
+In this section, we reexamine the issue of light transmission in non-inertial frames with particular reference to the aberration of light phenomena.
+We derive the aberration for a pulse of light traveling in the accelerated systems using the Langevin metric in general relativity  [^24] The aberration of light problem in the accelerated systems is solved with the discovery of the essential asymmetry between the non-inertial and the inertial observers.
+This asymmetry is of the same nature as that of the well-known Sagnac effect [28, 29, 30].
+For instance, Langevin’s 1921 explanation of the Sagnac effect rested upon the assertion that ”any change of velocity, or any acceleration has an absolute meaning.” [31]..
 We only wish to emphasize here the following point.
 From a mathematical standpoint, there is no difference between calculations in the framework of the general theory of relativity and of the special theory of relativity in the absence of space-time curvature.
-The aberration of light problem is solved with discovery of the essential asymmetry between the non-inertial and the inertial observers.
-Actually, in resent years it seems to be almost normally accepted in scientific community that the ”theory of relativity” is just a name, not to be taken literally.
-One can conclude that not all is relative in relativity, because this theory also contains some features which are absolute.
-Has all our talk about asymmetry violated the relativity principle?
-At a first glance it might seem so, since the relativity principle is often interpreted as implying perfect symmetry among moving frames.
-The principle of relativity denied the possibility for an observer partaking in a uniform motion relative to an inertial frame of discovering by any measurement such a motion, of course, that one does not look outside.
+The aberration of light problem is solved with the discovery of the essential asymmetry between the non-inertial and the inertial observers.
 The arguments concerning the relativity of motion in our case of interest cannot be applied, since the inertial and non-inertial reference systems are not equitable.
 A typical resolution of the asymmetry paradox identifies acceleration as the agency of asymmetry.
-It should be note that the duration of the acceleration period has a negligible effect on the anisotropy in the accelerated frame, so the acceleration need not be considered explicitly in working the problem.
+It should be noted that the duration of the acceleration period has a negligible effect on the anisotropy in the accelerated frame, so the acceleration need not be considered explicitly in working the problem.
 Nevertheless, the acceleration completely determines the problem.
-It is worth remarking that the absent of a dynamical explanation for the asymmetry in special relativity has disturbed some physicists.
+It is worth remarking that the absence of a dynamical explanation for the asymmetry in special relativity has disturbed some physicists.
 A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer.
-An idea of treating the accelerated frame using the equivalence principle is useful from a pedagogical point of view.
-See section 9.1 for more ideas along this line.
-However, there remains an intriguing puzzle to solve: how can the (e.g.
+The idea of treating the accelerated frame using the equivalence principle is useful from a pedagogical point of view.
+We will discuss this subject further in Chapter 12.
+At this point, a reasonable question arises: how can the (e.g.
 earth-based and sun-based) observers tell which observer took the acceleration?
 The surprising fact is, the determination of which observer took the acceleration can be made only by observation of the ”fixed stars”.
-The acceleration is in principle defined in terms of motion relative to the fixed stars, and they must be consulted in order to determine whether an acceleration occurred.
-Thus when we state that the earth-based observer undergoes an acceleration, and the sun-based observer does not, there is a hidden assumption concerning the distribution of mass in the universe.
+The acceleration is in principle defined in terms of motion relative to the fixed stars.
 The implicit ”absolute” acceleration means acceleration relative to the fixed stars.
 5.1 Absolute Time Coordinatization in the Accelerated Systems
-We will consider the problem of aberration of light in the accelerated systems on the basis of the theory of special relativity.
-Let us demonstrate that for explanation of the optical effects in the rotating frame of reference one does not need neither modify the special theory of relativity, nor apply the general theory of relativity.
+We will consider the problem of aberration of light in accelerated systems on the basis of the theory of special relativity.
+Let us demonstrate that for an explanation of the optical effects in the rotating frame of reference, one does not need to modify the special theory of relativity nor apply the general theory of relativity.
 It is only necessary to strictly follow the special theory of relativity.
-In order to write electrodynamics laws in the accelerated frame we have to go just one step further, and define the metric of the non-inertial frame.
-One must take into account that metric tensor must be a continuous quantity.
+To write electrodynamics laws in the accelerated frame we have to go just one step further and define the metric of the non-inertial frame.
+One must take into account that the metric tensor must be a continuous quantity.
 The first mathematical idea is the smooth tailoring of the metric tensor.
-This problem in special relativity can be adequately treated only by an approach which uses the absolute time coordinatization.
-We describe the mathematical expansion of special theory of relativity onto accelerations in the Chapter 9.
-Also this can be done directly, it is some times possible to save time by getting the answer with some physical arguments.
-For example, in our case of interest one can take advantage of the following considerations, which are in agreement with rigorous mathematical derivation.
+This problem in special relativity can be adequately treated by an approach that uses absolute time coordinatization.
+We describe the mathematical expansion of the special theory of relativity onto accelerations at the end of this subsection.
+Also, this can be done directly, it is sometimes possible to save time by getting the answer with some physical arguments.
+For example, in our case of interest, one can take advantage of the following considerations, which are in agreement with rigorous mathematical derivation.
 Suppose that the initial inertial system $S$ is at rest with respect to the fixed stars and a system $S_{n}$ (and an observer with his measuring instruments) in the frame $S$ is accelerated from the rest with respect to the fixed stars up to velocity $v$ along the $x$-axis.
-In accelerated systems, only the theory maintaining an absolute simultaneity is logically consistent with the natural behavior of clocks.
+In accelerated systems, the theory of maintaining absolute simultaneity is logically consistent with the smooth tailoring of the metric tensor.
 The method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the system $S_{n}$ was at rest.
 It is well known that during the motion with acceleration (with respect to the fixed stars) the procedure of Einstein’s clock synchronization cannot be performed and the interval in the accelerated reference system $S_{n}$ will, by the moment when the system $S_{n}$ starts moving with constant velocity, have the non-diagonal form.
-Absolute simultaneity can be introduced in special relativity without affecting neither the logical structure, no the (convention-independent) predictions of the theory.
-We begin with the metric as the true measure of space-time intervals for an non-inertial observer $S_{n}$ with coordinates $(t_{n},x_{n},y_{n},z_{n})$.
-We transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$, using the inverse Galilean transformation: we substitute $x=x_{n}+vt$, while leaving time unchanged $t_{n}=t$ into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain (1)
+We begin with the metric as the true measure of space-time intervals for an accelerated observer $S_{n}$ with coordinates $(t_{n},x_{n},y_{n},z_{n})$.
+We transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$, using the inverse Galilean transformation: we substitute $x=x_{n}+vt$, while leaving time unchanged $t_{n}=t$ into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain
 $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}~{}.$ (15)
-This metric describes the electrodynamics of the accelerated light source with viewpoint of of the accelerated observer measurements as viewing this of the inertial observer.
-Inspecting Eq.
-(15) we can find the components of the metric tensor $g_{\mu\nu}$ in the coordinate system of $S_{n}$.
+This metric describes the electrodynamics of the accelerated light source with the viewpoint of the accelerated observer measurements as viewing this of the inertial observer.
+Inspecting Eq.(15) we can find the components of the metric tensor $g_{\mu\nu}$ in the coordinate system of $S_{n}$.
 We obtain $g_{00}=1-v^{2}/c^{2}$, $g_{01}=-v/c$, $g_{11}=-1$.
@@
 $dx_{n}=dx$, $dt_{n}=dt$.
-(see Chapter 9 for more details).
-The velocity of light emitted by a source at rest in the coordinate system $(t,x)$ for $S$ is $c$.
-In that case the Minkowski metric Eq.(1) associated with inertial frame $S$ predict a symmetry in the one-way speed of light.
-In the coordinate system $(t_{n},x_{n})$, however, the speed of light emitted by the accelerated source (i.e.
-source which is at rest with respect to the accelerated frame $S_{n}$) cannot be equal $c$ anymore because $(t_{n},x_{n})$ is related to $(t,x)$ via a Galilean transformation.
+(see Chapter 12 for more details).
+Now let us return to the mathematical expansion of the special theory of relativity onto accelerations.
+We will keep the mathematical complexity to a minimum and assume that the frame $S_{n}$ begins at $t=0$ to move with constant acceleration $a>0$ and does so till $t=T$.
+Now we would like to discuss an expression for the velocity from the viewpoint of the inertial (non-accelerated) observer.
+We know that uniformly accelerated motion must obey the equation $d(\gamma\vec{v})/dt=\vec{f}/m=\vec{a}$.
+This is readily integrated $\gamma\vec{v}=\vec{a}t$.
+We will then have $dx/dt=at/\sqrt{1+a^{2}t^{2}/c^{2}}$.
+We emphasize that $dx(t)/dt$ is the velocity of the noninertial frame as seen by the inertial observer.
+Under the inverse Galilean transformation $dx=dx_{n}+atdt/\sqrt{1+a^{2}t^{2}/c^{2}}$, $dt_{n}=dt$ the metric of the uniformly accelerating observer becomes
+$ds^{2}=c^{2}dt_{n}^{2}(1+a^{2}t_{n}^{2}/c^{2})^{-1}-2at_{n}dt_{n}dx_{n}(1+a^{2}t_{n}^{2}/c^{2})^{-1/2}-dx_{n}^{2}~{}.$ (16)
+Since the metric tensor of space-time must be a continuous quantity, we must match the coordinates in the accelerated and inertial stages.
+At $t=0$ this metric naturally coincides with the $S$ frame metric and we have $g_{00}=1$.
+During acceleration, we picked the absolute time coordinatization.
+Therefore, for the metric tensor to be continuous when the reference frame passes from non-inertial motion to inertial one we will make use of the metric Eq.
+16.
+We will therefore require that at $t=T$ the metric of a noninertial frame would pass continuously into the Langevin metric.
+For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer $(x_{n},t_{n})$ to the coordinates $(x,t)$ of the inertial observer by the Galilean boost $x_{n}=x-vt$, $t_{n}=t$, where $v=aT/\sqrt{1+a^{2}T^{2}/c^{2}}$.
+In this case, in the interval $T<t_{n}$ of coordinate time $t_{n}$ the reference frame of the accelerated system has the metric Eq.
+(15).
+We emphasize that $v$ is the velocity of the frame $S_{n}$ as seen by the inertial observer at the second (inertial) stage of $S_{n}$ motion.
+It becomes obvious that at $t=t_{n}=T$ the metric of the noninertial frame goes into the metric of the (accelerated) inertial frame in a continuous manner [^25] For a general discussion of the mathematical expansion of special relativity onto acceleration we suggest reading the book [16]..
+According to textbooks, inertial frames are characterized by motion at constant velocity relative to the fixed stars, and different inertial frames are related by Lorentz boosts.
+Therefore, it might appear that the coordinates and time in frames $S$ and $S_{n}$ (in the inertial stage) should be related by the Lorentz transformation.
+But here at the inertial part of the trajectory, the metric of both frames will be diagonal in the Lorentz coordinatization and so at $t_{n}=T$, the metric tensor of the accelerated frame must abruptly change from the value $g_{00}=(1+a^{2}T^{2}/c^{2})^{-1}$ to the value $g_{00}=1$.
+5.2 The Asymmetry between the Inertial and Accelerated Frames
+The velocity of light emitted by a source at rest in the coordinate system $(t,x,y,z)$ for $S$ is $c$.
+In that case the Minkowski metric Eq.(1) associated with inertial frame $S$ predicts a symmetry in the one-way speed of light.
+In the coordinate system $(t_{n},x_{n},y_{n},z_{n})$, however, the speed of light emitted by the accelerated source (i.e.
+source which is at rest with respect to the accelerated frame $S_{n}$) cannot be equal $c$ anymore because $(t_{n},x_{n},y_{n},z_{n})$ is related to $(t,x,y,z)$ via a Galilean transformation.
 This is readily verified if one recalls that the velocity of light in the reference system $S$ is equal to $c$.
@@
 This means that in the accelerated reference system of coordinates $(ct_{n},x_{n})$ the velocity of light parallel to the x-axis, is $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction as stated above.
-The reason why it is different from the electrodynamics constant $c$ is due to the fact that the clocks are synchronized following the absolute time convention, which is fixed because $(t_{n},x_{n})$ is related to $(t,x)$ via a Galilean transformation.
-5.3 The Asymmetry between the Inertial and Accelerated Frames
-We discovery of the essential asymmetry between the inertial and accelerated frames, namely, the Maxwell’s equations are not applicable from the viewpoint of an observer at rest with respect to accelerated frame $S_{n}$.
-In fact, the metric Eq.(15) associated with accelerated reference frame $S_{n}$ predicts an asymmetry in the one-way speed of light in the relative velocity direction.
-Accelerations (with respect to the fixed stars) have an effect on the propagation of light.
-On accelerated system $S_{n}$, the velocity of light emitted by a source at rest must be added to (or subtracted from) the speed due to acceleration and the velocity of light is different in opposite directions.
-In contrast, the Maxwell’s equations continue to hold from the viewpoint of an observer at rest with respect to inertial frame $S$.
-On inertial system $S$, the velocity of light emitted by a source at rest is $c$.
-The following important detail of such acceleration frame description can hardly be emphasize enough.
+We discovered the essential asymmetry between the inertial and accelerated frames, namely, Maxwell’s equations are not applicable from the viewpoint of an observer at rest with respect to accelerated frame $S_{n}$.
+The metric Eq.(15) associated with accelerated reference frame $S_{n}$ predicts an asymmetry in the one-way speed of light in the relative velocity direction.
+Accelerations (with respect to the fixed stars) affect on the propagation of light.
+On accelerated system $S_{n}$, the velocity of light emitted by a source at rest must be added to (or subtracted from) the speed due to acceleration, and the velocity of light is different in opposite directions.
+In contrast, on an inertial system, $S$, the velocity of light emitted by a source at rest is $c$.
+The following important detail of such acceleration frame description can hardly be emphasized enough.
 Metric applies to physical laws, not to physical facts.
-Indeed, it is always possible to chose such variables, in which metric of the accelerated source will be diagonal.
+Indeed, it is always possible to choose such variables, in which the metric of the accelerated source will be diagonal.
 We interpret the Langevin metric to mean that the law of electrodynamics is expressed by anisotropic field equations in the accelerated frame.
-By a physical facts in this context we mean the aberration of light radiated by a single ”plane wave” emitter in the accelerated frame.
-The electrodynamics equation needs to be integrated with initial condition for the radiation wavefront.
-After the boost we can see that acceleration has no effect on the wavefront orientation.
+By physical facts in this context we mean the aberration of light radiated by a single ”plane wave” emitter in the accelerated frame.
+The electrodynamics equation needs to be integrated with the initial condition for the radiation wavefront.
+After the boost we can see that acceleration does not affect the wavefront orientation.
 In fact, the variables $(t,x,y,z)$ can be expressed in terms of the variables $(t_{n},x_{n},y_{n},z_{n})$ by means of Galilean transformation $x_{n}=x-vt$, $t_{n}=t$, so that the wavefront of the emitted light beam is perpendicular to the vertical direction $z_{n}$ after the acceleration, Fig.
 14.
-We shall discuss this subject further in the Section 5.5.
-It would be well to begin with bird’s view of some main results.
+We shall discuss this subject further in Section 5.5.
+Figure 14: Aberration of light in an accelerated frame of reference $S_{n}$.
+Radiation wavefront orientation and the anisotropy of the speed of light are presented in the absolute time coordinatization ($t_{n}=t$) for the screen.
+Figure 15: Aberration of light in an accelerated frame of reference $S_{n}$.
+The screen is at rest in the accelerated frame.
+The aberration of the transmitted light is considered independent of the source speed.
+It would be well to begin with a bird’s view of some main results.
 Suppose that an observer in the accelerated frame $S_{n}$ performs an aberration measurement.
-How shall we describe the aberration of light from the ”plane wave” emitter which is at rest in the inertial frame $S$?
-In order to predict the result of the aberration measurement the accelerated observer should use the non-diagonal metric Eq.(15).
-Figure 14: Aberration of light in an accelerated frame of reference $S_{n}$.
-Radiation wavefront orientation and the anisotropy of speed of light presented in the absolute time coordinatization ($t_{n}=t$) for the screen.
-The crossed term in metric Eq.(15) generates anisotropy in the accelerated frame that is responsible for the change of radiation direction (aberration).
-Figure 15: Aberration of light in an accelerated frame of reference $S_{n}$.
-Screen is at rest in the accelerated frame.
-Radiation wavefront orientation and the anisotropy of speed of light presented in the absolute time coordinatization ($t_{n}=t$) for the screen.
-The aberration of the transmitted light is considered independent of the source speed and to have just a local origin exclusively based on the observer acceleration relative to the fixed stars.
+To predict the result of the aberration measurement the accelerated observer should use the non-diagonal metric Eq.(15).
 According to the asymmetry between the inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of light.
-Namely, if the opaque screen with hole was at rest relative to the fixed stars and the screen started from rest to motion, then the apparent angular position of the ”plane-wave” emitter seen in the accelerated frame through the aperture would jump by angle $-v/c$.
+Namely, if the opaque screen with a hole was at rest relative to the fixed stars and the screen started from rest to motion, then the apparent angular position of the ”plane-wave” emitter seen in the accelerated frame through the aperture would jump by angle $-v/c$.
 That is the transmission through the hole in the opaque screen in the frame $S_{n}$ appears as shown in Fig 14.
@@
 It is found that there is no aberration proceeding from the emitter, independently of their motion in the case of the small aperture hole.
-It is important to emphasize that the aberration of the light beam transmitted through the small aperture hole also does not dependent on the emitter motion in the accelerated system $S_{n}$, Fig.
+It is important to emphasize that the aberration of the light beam transmitted through the small aperture hole also is not dependent on the emitter motion in the accelerated system $S_{n}$, Fig.
 15.
@@
 5.3 Explanation of the Aberration on the Basis of Electrodynamics
-It is well known that the electrodynamics theory meets all requirements of the theory of relativity and therefore must accurately describe the properties of such a relativistic object as light.
-For example, we are able to demonstrate that velocity of light in the accelerated frame can be derived independently of a technique involving metric Eq.
+It is well known that the electrodynamics theory meets all requirements of the theory of relativity.
+For example, we can demonstrate that the velocity of light in the accelerated frame can be derived independently of a technique involving metric Eq.
 (15).
 In the inertial frame, fields are expressed as a function of the independent variables $x,y,z$, and $t$.
-According to the special relativity, the Maxwell’s equations always valid in the Lorentz reference frame.
+According to the special relativity, Maxwell’s equations are always valid in the Lorentz reference frame.
 The electric field $\vec{E}$ of an electromagnetic wave radiated by the emitter at rest in the inertial frame satisfies the equation $\Box^{2}\vec{E}=\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$.
-However, the variables $x,y,z,t$ can be expressed in terms of the variables $x_{n},y_{n},z_{n},t_{n}$ by means of a Galilean transformation, so that fields can be written in terms of $x_{n},y_{n},z_{n},t_{n}$.
+However, the variables $x,y,z,t$ can be expressed in terms of the variables $x_{n},y_{n},z_{n},t_{n}$ using a Galilean transformation, so that fields of an electromagnetic wave radiated by the emitter at rest in the accelerated frame can be written in terms of $x_{n},y_{n},z_{n},t_{n}$.
 Hence the wave equation transforms into
@@
 The solution of this equation $F[x_{n}-(c-v)t_{n}]+G[x_{n}+(-c-v)t_{n}]$ is the sum of two arbitrary functions, one of argument $x-(c-v)t$ and the other of argument $x+(-c-v)t$.
-The first term represents a wave traveling forward in the positive $x_{n}$ direction, and the second term a wave traveling backwards in the negative $x_{n}$ direction.
-Acceleration has an effect on the propagation of light.
-In fact, the coordinate velocity of light parallel to the $x$-axis is given by $dx/dt=c-v$ in the positive direction, and $dx/dt=-c-v$ in the negative direction.
+The first term represents a wave traveling forward in the positive $x_{n}$ direction, and the second term a wave traveling backward in the negative $x_{n}$ direction.
+Acceleration affects the propagation of light.
+The coordinate velocity of light parallel to the $x$-axis is given by $dx/dt=c-v$ in the positive direction, and $dx/dt=-c-v$ in the negative direction.
 We now consider the transmission through the hole in the opaque screen in the accelerated frame $S_{n}$, Fig.
@@
 We describe the aberration of light based on the observations made by an observer in the same accelerated frame as the screen.
-It is assumed that the detector for the direction of the radiation is an energy propagation detector and the size of the detector aperture is sufficiently large compared with the radiation beam size.
-In other words, it is assumed that there is no physical influence of the detector (e.g.
-aperture) on the measurement.
-We have already discussed the explanation of the aberration on the basis of electrodynamics in the previous chapter.
+We have already discussed the explanation of the aberration based on electrodynamics.
 Our earlier discussion is really about as far as anyone would normally need to go with the subject, but we are going to do it again.
-One reason is that one should know how to deal with what happens to waves in the accelerated frame from the viewpoint electrodynamics equations.
-First, we remind about the convenient way of describing diffraction by opaque screen we used in the previous chapter.
-The screen containing a hole is a kind of diffraction grating which breaks up the radiated beam into a number of diffracted beams of plane waves.
-Each of these beams corresponds to one of the Fourier components into which an transmittance can be resolved.
+One reason is that one should know how to deal with what happens to waves in the accelerated frame from the viewpoint of electrodynamics equations.
+The screen containing a hole is a kind of diffraction grating that breaks up the radiated beam into a number of diffracted beams of plane waves.
+Each of these beams corresponds to one of the Fourier components into which a transmittance can be resolved.
 Let us assume that the transmittance of the grating varies according to the law $T=g(K_{\perp})\cos{({K}_{\perp}x)}$.
@@
 In our example, the plane waves with different wavenumber vectors propagate out from the screen with different frequencies.
-Then equation $\Delta\omega/\Delta k_{x}=-v$ holds for each transmitted waves independently on the sign and the magnitude of the transmitted angle.
-The last equations state that transmitted light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=-v$, Fig.
+Then equation $\Delta\omega/\Delta k_{x}=-v$ holds for each transmitted wave independently on the sign and the magnitude of the transmitted angle.
+The last equations state that the transmitted light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=-v$, Fig.
 15.
 We would now like to discuss an apparent paradox.
-We obtained the electrodynamics equations in accelerated frame using the Galilean transformation of the Maxwell’s equations.
+We obtained the electrodynamics equations in an accelerated frame using the Galilean transformation of Maxwell’s equations.
 Intuition would seem to tell us that everything is at rest in the accelerated frame.
 But special relativity says that there is a derivative $\partial/\partial t_{n}=v\partial/\partial x_{n}$ because there is $\partial/\partial x_{n}$ that is not zero.
-Here we have a new kind of situation which is quite different from inertial frame electrodynamics.
-We will discuss this subject further in the Chapter 10.
+Here we have a new kind of situation that is quite different from inertial frame electrodynamics.
+We will discuss this subject further in Chapter 9.
+5.4 Simple Explanation of the Aberration of Light Effect in Accelerated Frame
+Above we considered the aberration of light problem in the accelerated frame.
+The aberration of light effect is not easy to calculate in the frame of reference attending acceleration.
+In this case, we used a metric tensor to obtain the electrodynamics equations in the accelerated reference frame.
+The presented approach uses the Fourier transform method.
+As one of the consequences of the crossed term in the wave equation, we found the group velocity that is responsible for the aberration.
+Now let us present a very simple explanation of the aberration of light in the accelerated frame based on electrodynamics.
+To find the aberration increment, we simply use Galilean transformation law for the electric and magnetic fields.
+We have an expression for the energy flow vector of the electromagnetic field.
+This vector, $\vec{S}=c\vec{E}\times\vec{B}/(4\pi)$, is called ”Poynting’s vector”.
+It tells us the rate at which the field energy moves around in space.
+To detect the aberration of light effect inside the accelerated frame, it is obvious that some coordinate system with reference direction is needed.
+We determine the reference directions perpendicular to the line motion in the inertial frame and in the accelerated frame.
+The motion of the aberrated light beam is assumed for simplicity, to lie in the same plane and the angular position of the aberrated beam is described by one angle.
+The analysis is much simplified if we treat separately the case of a radiated beam with its $\vec{E}$-vector parallel to the ”plane of incidence” (that is $xz$-plane) and the case of a radiated beam with the $\vec{E}$-vector perpendicular to the $xz$-plane.
+We will carry through the analysis for an incoming beam polarized perpendicular to the plane of incidence, but the principle is the same for both.
+So we take that $\vec{E}$, has only $y$- component.
+For a magnetic field, we get $\vec{B}=\vec{e}_{z}\times\vec{E}$, where $\vec{e}_{z}$ is the direction of Poynting vector $\vec{S}$ in the inertial frame.
+Above we demonstrated that we must match the (4D) coordinates in the accelerational and inertial parts of the trajectory.
+For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the Galilean boost.
+We can obtain the electrodynamics equations in an accelerated frame using the Galilean transformation (with velocity $-v$) of Maxwell’s equations.
+The fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame where the emitter is at rest are not the same as those $\vec{E}$ and $\vec{B}$ in the inertial frame before the active Galilean boost.
+With the Galilean transformation, the field transformation is $\vec{E}_{n}=\vec{E}-\vec{v}\times\vec{B}/c$, $\vec{B}_{n}=\vec{B}-\vec{v}\times\vec{E}/c$.
+This expression describes the first-order (in $v/c$) effects only.
+Here $\vec{v}=v\vec{e}_{x}$ is the velocity vector of the accelerated frame in the inertial frame.
+If the $\vec{E}$-vector perpendicular to the $xz$-plane, the magnetic field in the inertial frame has only a $x$-component and we have immediately $\vec{E}_{n}=\vec{E}=\vec{e}_{y}E$, $\vec{B}_{n}=(\vec{e}_{z}\times\vec{e}_{y})E-(\vec{e}_{x}\times\vec{e}_{y})vE/c=-\vec{e}_{x}E-\vec{e}_{z}vE/c$.
+Therefore there is a Poynting vector, $c\vec{E}_{n}\times\vec{B}_{n}/(4\pi)=[-(\vec{e}_{y}\times\vec{e}_{x})-(\vec{e}_{y}\times\vec{e}_{z})v/c]cE^{2}/(4\pi)=[\vec{e}_{z}-\vec{e}_{x}(v/c)]cE^{2}/(4\pi)$.
+Then, the radiated beam is propagated at the angle $-v/c$ with respect to the $z_{n}$-axis yielding the phenomenon of the aberration of light in the accelerated frame.
+We have found that we get the same result whether we analyze the energy transport direction using electrodynamics equations in the accelerated frame or using Galilean transformation law for electromagnetic fields.
+In the first instance, the equations state that the transmitted light beam with finite transverse size moves along $x$-direction with group velocity $d\omega/dk_{x}=-v$, in the second, the effect of light aberration is understood as a change in the direction of Poynting vector.
+The two approaches give the same result.
+In this section, we demonstrate that the direction of group velocity and the direction of the Poynting vector are two sides of the same coin.
+Acceleration affects the field equations.
+When the observer is at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be modified.
+One can easily show this by transforming coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer moving with velocity $-v$ with respect to the accelerated observer.
+From the Galilean transformation $x_{n}=x-vt,~{}y_{n}=y,~{}z_{n}=z,~{}t_{n}=t$, after partial differentiation, one obtains $\partial/{\partial t_{n}}=\partial/{\partial t}+v\partial/{\partial x}$, $\partial/{\partial x_{n}}=\partial/{\partial x}$.
+Our problem now is to work out the magnetic field of the transmitted light beam in terms of those of the incident plane wave.
+How can we do that?
+The magnetic field $\vec{B}_{n}$ satisfies electrodynamics equation ×$\vec{\nabla}\times\vec{B}_{n}=(\partial/{\partial t}+v\partial/{\partial x})\vec{E}_{n}/c$, in the accelerated frame.
+We are very specific and wright out explicitly the new component: $-\partial B_{z}/{\partial x}=(v/c)\partial E_{y}/{\partial x}$.
+In any measurement of the angular displacement of transmitted (through aperture) light beam, one will have to consider a finite radiation beam size.
+If we integrate this equation with respect to $x$ across transmitted light beam, we conclude that $\vec{B}_{n}=-\vec{e}_{x}E-\vec{e}_{z}vE/c$.
+This last result is just what we got by a field transformation argument.
+This is a good point to make a general remark about the physical reality of the energy flow (Poynting) vector.
+In any physically realizable aberration measurement, one will have to consider a finite radiation beam size.
+The direction of the energy transport (and, hence, the direction of the Poynting vector) has an exact objective meaning i.e.
+convention-invariant.
+However, there is a common mistake made in electrodynamics connected with the energy transport direction in the case of a plane wave.
+It is impossible to know the energy transport direction when one has to deal with a plane wave.
 5.5 A Resynchronization of the Accelerated Clocks
 It should be noted, however, that there is another satisfactory way of explaining the effect of aberration of light in the accelerated frame $S_{n}$.
-The explanation consists in using a clock re-synchronization procedure.
-Well known that in their original form Maxwell’s equations are valid in the inertial frames.
-But Maxwell’s equations can be written down only if the Lorentz coordinates has already been specified.
+The explanation consists of using a clock re-synchronization procedure.
 When the system $S_{n}$ starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed.
-The Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all direction.
-Using such synchronization procedure we actually select a Lorentz coordinate system for the screen.
-In this synchronization, we describe the transmission through the aperture using usual Maxwell’s equations.
+The Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions.
+Using such a synchronization procedure we select a Lorentz coordinate system for the screen.
+In this synchronization, we describe the transmission through the aperture using Maxwell’s equations.
 The interval in the accelerated reference system $S_{n}$ will have the diagonal form Eq.(1) for the transmitted light beam.
@@
 The plane wavefront rotates in the accelerated frame after the accelerated clock re-synchronization.
-The Maxwell’s equations can now be used to describe the transmitted (through the aperture) light beam.
-According to the Maxwell’s electrodynamics, the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light.
-The time $t^{\prime}_{n}$ under the Einstein’s synchronization in the $S_{n}$ frame is readily obtained by introducing the offset factor $x_{n}v/c^{2}$ and substituting $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first order approximation.
-This time shift has the effect of rotation the plane of simultaneity (that is emitter radiation wavefront) on the angle $-v/c$ in the first order approximation.
+According to Maxwell’s electrodynamics, the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light.
+The time $t^{\prime}_{n}$ under Einstein’s synchronization in the $S_{n}$ frame is readily obtained by introducing the offset factor $x_{n}v/c^{2}$ and substituting $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first-order approximation.
+This time shift has the effect of rotation of the plane of simultaneity (that is emitter radiation wavefront) on the angle $-v/c$.
 As a consequence of this, the plane wavefront rotates in the accelerated frame after re-synchronization, Fig.
 16.
-The new time coordinate in the accelerated frame is interpreted by saying that Maxwell’s equations are applicable to the light transmission (through the aperture) description.
-Then, the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light: the two approaches give the same result.
+The new time coordinate in the accelerated frame is interpreted by saying that Maxwell’s equations are applicable to the light transmission description.
+Then, the transmitted (through the aperture) light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light: the two approaches give the same result.
 The choice between these two different clock synchronizations is a matter of pragmatics.
@@
 Physical time $t^{\prime}_{n}$ determines the flow of time in a physical process in the accelerated frame $S_{n}$.
-The $dx^{\prime}_{n}$ is the physical distance between to points of 3-space.
+The $dx^{\prime}_{n}$ is the physical distance between two points of 3-space.
 5.6 Composition of Motions
@@
 Now we shall be more general and study the case where the motion is at an arbitrary velocity.
-We already know from our discussion in this chapter that metric tensor must be a continuous quantity.
+We already know from our discussion in this section that metric tensor must be a continuous quantity.
 The Langevin metric is found by matching the accelerated frame and inertial frame metric tensors.
 For smooth tailoring of the metric tensor, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the inverse Galilean boost.
-In this case the metric of inertial frame would pass into the Langevin metric of the accelerated frame.
+In this case, the metric of the inertial frame would pass into the Langevin metric of the accelerated frame.
 Let us consider the transformation of the direction of light propagation.
 The light (group) velocity transforms as a particle velocity.
-In this case we use the Galileo addition theorem of velocities for transforming the speed of propagation of light to the accelerated frame.
+In this case, we use the Galileo addition theorem of velocities for transforming the speed of propagation of light to the accelerated frame.
 Let us assume that the accelerated frame moves with velocity $v$, then the angle of aberration $\theta_{a}$ is such that   $\tan\theta_{a}=-v/c$.
 If $v/c\ll 1$ the angle of aberration is small, being approximated by $\theta_{a}=-v/c$.
-It is always possible to chose such variables, in which metric of the accelerated source will be diagonal.
+It is always possible to choose such variables, in which the metric of the accelerated source will be diagonal.
 In the Lorentz coordinatization, light propagates with constant velocity $c$ independently of the direction of propagation, and of the velocity of its source.
-If we look at transformation $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$, we see that velocity along the $x_{n}$-axis is the same in the Lorentz coordinatization too: $dx_{n}^{\prime}/dt_{n}^{\prime}=(dx_{n}/dt_{n})/[1-v^{2}/c^{2}-v(dx_{n}/dt_{n})]$.
+If we look at the transformation $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$, we see that velocity along the $x_{n}$-axis is the same in the Lorentz coordinatization too: $dx_{n}^{\prime}/dt_{n}^{\prime}=(dx_{n}/dt_{n})/[1-v^{2}/c^{2}-v(dx_{n}/dt_{n})]$.
 In fact, substituting $dx_{n}/dt_{n}=-v$ into this equation gives $dx_{n}^{\prime}/dt_{n}^{\prime}=-v$.
-Group velocity of light beam along the fixed stars motion is still given by $-v$, but when we use the new velocity of light this becomes $\sin\theta_{a}=-v/c$.
-This is modification of the aberration of light law in the Lorentz coordinatization.
-Suppose that accelerated observer performs a measurement of the light aberration.
-Since in the Lorentz coordinatization the time and distance have direct physical meaning, the accelerated observer finds that the angle of aberration is such that $\sin\theta_{a}=-v/c$.
+The group velocity of the light beam along the fixed stars motion is still given by $-v$, but when we use the new velocity of light this becomes $\sin\theta_{a}=-v/c$.
+This is a modification of the aberration of light law in the Lorentz coordinatization.
+Suppose that an accelerated observer measures the light aberration.
+Since in the Lorentz coordinatization, the time and distance have direct physical meaning, the accelerated observer finds that the angle of aberration is such that $\sin\theta_{a}=-v/c$.
 5.7 Discussion
-Now let us return to metric diagonalization and try to get better understanding of the relationship between the $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$ coordinate systems.
-We have already noted that the old coordinates $(x_{n},y_{n},z_{n},t_{n})$ are found by matching the accelerated frame and inertial frame metrics; what does mismatching of the coordinates $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$ mean, in terms of measurements made by accelerated observer?
-Here the metric in both reference frames will be diagonal, and, according to textbooks, the coordinates in this frames should be related by the Lorentz transformation.
-At first site the problem is complete symmetrical and both frames are equivalent.
+Now let us return to metric diagonalization and try to get a better understanding of the relationship between the $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$ coordinate systems.
+We have already noted that the old coordinates $(x_{n},y_{n},z_{n},t_{n})$ are found by matching the accelerated frame and inertial frame metrics; what does mismatching of the coordinates $(x,y,z,t)$ and $(x^{\prime}_{n},y^{\prime}_{n},z^{\prime}_{n},t^{\prime}_{n})$ mean, in terms of measurements made by the accelerated observer?
+Here the metric in both reference frames will be diagonal, and, according to textbooks, the coordinates in these frames should be related by the Lorentz transformation.
+At first sight, the problem is completely symmetrical and both frames are equivalent.
 Above we demonstrated that there is a real difference between the accelerated inertial frame and the inertial frame without accelerational history.
-From a mathematical viewpoint , we suggest the following explanation.
-After diagonalization, the metric tensor in the accelerated frame must abruptly change from the value $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the value $g_{00}=1,g_{01}=0$.
-We must conclude that when we are dealing with (Lorentz ) transformation and this transformation does not preserve continuity of the metric tensor we have no symmetry between inertial frames.
 Where does the asymmetry come from?
@@
 In the inertial frame, the wavefront of the emitted light beam is perpendicular to the vertical direction $z$.
-On the other hand, we know that when we are dealing with Lorentz coordinatization we have effect on the wavefront orientation in the accelerated frame $S_{n}$.
+On the other hand, we know that when we are dealing with Lorentz coordinatization we have an effect on the wavefront orientation in the accelerated frame $S_{n}$.
 According to relativistic kinematics, the extra phase chirp $d\phi/dx^{\prime}_{n}=k_{x}=-v\omega_{i}/c^{2}$ is introduced as a consequence of this, the plane wavefront rotates after the acceleration.
 Then, the radiated light beam is propagated at the angle $-v/c$ with respect to the $z_{n}$-axis, yielding the phenomenon of the aberration of light in the accelerated frame $S_{n}$.
-It is generally accepted that this is purely kinematic effect, involving no forces.
+This is purely a kinematic effect, involving no forces.
 Many people will, we think, find this disturbing.
-They would like to think that since in the Lorentz coordinatization the time and distance have direct physical meaning, there should be some physical (dynamical) reason for this wavefront rotation.
-We discuss the key of this paradox in great detail in the Section 9.1.
+We discuss the key to this paradox in great detail in Section 12.
 It is a dynamical line of arguments that explains this paradoxical situation with the asymmetry between inertial and accelerated reference frames.
@@
 The rate of a clock depends on the pseudo-gravitational potential at the place where the clock is situated.
-The time-offset relation can be interpreted as the accomulated time difference between spatially separated clocks.
-Suppose (for simplicity) that there is a constant acceleration $g$ along $x_{n}$-axis during the time $T=v/g$.
+The time-offset relation can be interpreted as the accumulated time difference between spatially separated clocks.
+Suppose (for simplicity) that there is a constant acceleration $g$ along the $x_{n}$-axis during the time $T=v/g$.
 The pseudo-gravitational acceleration is simply equal to the gradient of scalar potential: $g=-\partial\phi/\partial x_{n}$.
 The clock at higher gravitational potential (placed along the direction of acceleration) runs faster.
-For the pseudo-gravitational potential difference between two points along the $x_{n}$-axis we get $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$.
+For the pseudo-gravitational potential difference between two points along the $x_{n}$-axis, we get $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$.
 When the system $S$ starts moving with constant velocity the gradient of potential in the system $S_{n}$ is zero.
-For accumulated time difference between two spatially separated clocks we have $t_{n}(1)-t_{n}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$.
+For accumulated time difference between two spatially separated clocks, we have $t_{n}(1)-t_{n}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$.
 This time shift has the effect of rotation the plane of simultaneity on the angle $-v/c$.
-5.8 A Single Moving Emitter in a Noninertial Frame of Reference
-Above in the sections 4.3 - 4.4 we considered a single moving ”plane wave” emitter in the initial inertial frame of reference.
-Let us now analyze the aberration of light radiated by a single ”plane wave” emitter moving in the accelerated system.
-Before we go on to analyze an observations of a non-inertial observer, we should make one more remark about observations of an inertial observer.
-In the Chapter 4 we already emphasized that in the description of the aberration of light in the initial inertial frame of reference there are two choices of (four-dimensional) coordinatizations useful to consider:
-(a) Non-standard (absolute time) coordinatization for a moving source
-(b) Standard Lorentz coordinatization for a moving source
-When the light source in an inertial frame is accelerated from rest up to velocity $v$ along the $x$ axis, the simplest ( absolute time) method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the light source was at rest, i.e.
-we still enforce the clock transport synchronization ( or Einstein synchronization which is defined in terms of light signals emitted by the light source at rest).
-This choice is usually the most convenient one from the viewpoint of connection to laboratory reality.
-In fact, conventional particle tracking treats the space-time continuum in a non-relativistic format, as a (3 + 1) manifold.
-In other words, in this standard approach, introducing as only modification to the classical case the relativistic mass, time is absolute and we have no mixture of positions and time.
-Let us, then, discuss (according to the Chapter 3) how to assign Lorentz coordinates in the case when a light source in an inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis.
-In order to assign a Lorentz coordinate system in the inertial frame after the Galilean boost, one needs to perform the distant clock resynchronization $t^{\prime}=t-xv/c^{2}$.
-This new space-time coordinates in the lab frame are interpreted, mathematically, by saying that the metric is now diagonal and the speed of light from the moving source is isotropic and equal to $c$.
-Now let us return to observations of a non-inertial observer.
-In the description of the aberration of light radiated by a single moving emitter in an accelerated frame of reference there are also two choices of coordinatizations:
-(a) Non-standard (absolute time) coordinatization for an accelerated source
-(b) Standard Lorentz coordinatization for an accelerated source
+5.8 A Single Moving Emitter in an Accelerated Frame of Reference
+Above in Section 4, we considered the single emitter problem in the initial inertial frame of reference $S$.
+It is found that there is a deviation of the energy transport for the light radiated by the accelerated emitter, which is nothing else but a well-known result: there is the aberration of light from the tangentially moving emitter in the inertial frame of reference, Fig.
+1.
+We demonstrated that there could be two (covariant and no-covariant) approaches to the analysis of the aberration of light radiated by a single moving emitter.
+Let us now analyze the aberration of light radiated by a single ”plane wave” emitter moving in the accelerated system. In this case, there are also two choices of coordinatizations:
 (a) When the system $S_{n}$ in the initial inertial lab frame $S$ is accelerated from the rest up to velocity $v$ along the $x$-axis, the simplest ( absolute time) method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the system $S_{n}$ was at rest.
-Acceleration with respect to the fixed stars have an effect on the propagation of light.
+Acceleration with respect to the fixed stars affects the propagation of light.
 The velocity of light from the source which is at rest in the accelerated frame $S_{n}$ will, by the moment when the system $S_{n}$ starts moving with constant velocity, have anisotropy along the $x_{n}$ axis (see Eq.
@@
 Figure 17: Aberration of light in an accelerated frame of reference $S_{n}$.
-Emitter is at rest with respect to the fixed stars.
-Wave fronts orientation and the group velocity presented in the Einstein synchronization.
-The Einstein synchronization is defined in terms of light signals emitted by a source at rest (in the accelerated frame) assuming that light propagates with the same velocity $c$ in all direction.
-The aberration increment $\theta_{a}$ is connected with the physical parameters by the relation: $\theta_{a}=-2v/c$, where $v$ is the velocity of the fixed stars in the accelerated frame $S_{n}$.
+The emitter is at rest with respect to the fixed stars.
+Wavefront orientation and the group velocity are presented in the Einstein synchronization.
 (b) When the system $S_{n}$ starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed.
-The Einstein synchronization is defined in terms of light signals emitted by a source at rest (in the accelerated frame) assuming that light propagates with the same velocity $c$ in all direction.
-Using such (Einstein time) synchronization procedure we actually describe the light from the source at rest (in the accelerated frame) using usual Maxwell’s equations.
-Above we considered the single emitter problem in the initial inertial frame of reference $S$.
-It is found that there is a deviation of the energy transport for the light radiated by the transversely moving emitter, which is nothing else but well-known (from textbooks) result: there is the aberration of light from the transversely moving emitter in the inertial frame of reference (Fig.
-1).
+The Einstein synchronization is defined in terms of light signals emitted by a source at rest (in the accelerated frame) assuming that light propagates with the same velocity $c$ in all directions.
+Using such (Einstein time) synchronization procedure we describe the light from the source at rest (in the accelerated frame) using usual Maxwell’s equations.
 We now consider the case when the emitter is at rest in the initial inertial frame $S$ (i.e.
 at rest with respect to the fixed stars) and the observer, which is at rest with respect to the accelerated frame of reference $S_{n}$ performs the direction of the energy transport measurement.
-It is important to emphasize that the aberration of light radiated by the single emitter is also dependent of the emitter motion in the accelerated system $S_{n}$.
+It is important to emphasize that the aberration of light radiated by the single emitter is also dependent on the emitter motion in the accelerated system $S_{n}$.
 According to the asymmetry between the inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of light.
@@
 That is the aberration of light in the frame $S_{n}$ appears as shown in Fig 17.
-We are going to discuss the asymmetry between inertial and accelerated frames.
-Now let us assume the absolute time coordinatization.
-The Langevin metric, together with initial condition, describes the aberration of light effect for the emitter at rest in the accelerated frame.
-The essential asymmetry is rooted in the electrodynamics equations.
-The laws of special relativity say that we can describe radiation of the moving with velocity $-v$ in the accelerated frame emitter in terms of the accelerated observer if we describe the electrodynamics by using metric
+Let us describe the radiation of the moving with velocity $-v$ in the accelerated frame emitter in terms of the accelerated observer.
+Acceleration with respect to the accelerated frame affects the propagation of light.
+Let us assume the absolute time coordinatization.
+Using the Galilean boost: we substitute $x_{n}\to x_{n}+vt_{n}$, while leaving time unchanged into the Langevin metric to obtain
 $ds^{2}=c^{2}(1-4v^{2}/c^{2})dt_{n}^{2}-4vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}~{}.$ (18)
-This metric describes the electrodynamics of the moving light source with viewpoint of the accelerated observer measurements.
-From the formula itself it is easy to see that the result of measurement of the electromagnetic field configuration in the moving system in terms of coordinate $x_{n}$ and time $t_{n}$ of the measured device at rest in the accelerated frame is the same as the measured result of this configuration at point $x_{n}+vt_{n}$ and time $t_{n}=0$.
-In the inertial frame we have a similar situation.
-In the Chapter 3 we described the electrodynamics of the moving emitter by using metric Eq.
+This metric describes the electrodynamics of the light source at rest in the inertial frame with the viewpoint of the accelerated observer measurements.
+From the formula itself, it is easy to see that the result of measurement of the electromagnetic field configuration in the moving system in terms of coordinate $x_{n}$ and time $t_{n}$ of the measured device at rest in the accelerated frame is the same as the measured result of this configuration at point $x_{n}+vt_{n}$ and time $t_{n}=0$.
+In the inertial frame, we have a similar situation.
+In Chapter 3 we described the electrodynamics of the moving emitter by using metric Eq.
 (11).
-This is the same way we used in the accelerated frame.
-So we begin to understand the basic machinery of the asymmetry between (initial) inertial and accelerated frames in the absolute time coordinatization: inertial observer describes the electrodynamics of the light source at rest using Minkowski metric.
-We substitute $x=x_{n}+vt,t_{n}=t$ into Minkowski metric to obtain Langevin metric Eq.
+Within a single accelerated frame, in the inertial stage, the symmetry principle (i.e.
+equivalence of the active and passive Galilean boost) is satisfied.
+Let us now discuss the basic machinery of the asymmetry between (initial) inertial and accelerated frames in absolute time coordinatization: inertial observer describes the electrodynamics of the light source at rest using the Minkowski metric.
+We substitute $x=x_{n}+vt,t_{n}=t$ into the Minkowski metric to obtain the Langevin metric Eq.
 15 in the comoving frame.
-We may ask ourselves, what is the metric of the inertial observer with viewpoint of the accelerated observer?
-We transform coordinates $(t_{n},x_{n},y_{n},z_{n})$ that would be coordinates of an accelerated observer $S_{n}$ using the inverse Galilean transformation: we substitute $x_{n}=x-vt$, while leaving time unchanged $t_{n}=t$ into the Langevin metric Eq.
+We may ask ourselves, what is the metric of the inertial observer with the viewpoint of the accelerated observer?
+We transform coordinates $(t_{n},x_{n},y_{n},z_{n})$ that would be coordinates of an accelerated observer $S_{n}$ using the inverse Galilean transformation: we substitute $x_{n}=x-vt$ while leaving time unchanged $t_{n}=t$ into the Langevin metric Eq.
 15 to obtain the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ in the initial inertial frame, as it must be.
-This metric describes the electrodynamics of the light source at rest in the inertial frame with viewpoint of the inertial observer measurements as viewing this of the accelerated observer.
-In particular this demonstrates the time dilation effect of a physical clock at rest in the accelerated frame.
-The slowly down of the accelerated clock does not depends on the choice of reference system, in which this effect is measured (see Chapter 9 for more details).
+This metric describes the electrodynamics of the light source at rest in the inertial frame with the viewpoint of the inertial observer measurements as viewing this of the accelerated observer.
+In particular, this demonstrates the time dilation effect of a physical clock at rest in the accelerated frame.
+The slowdown of the accelerated clock does not depend on the choice of the reference system, in which this effect is measured (see Chapter 12 for more details).
 There is still another way of solving the problem.
-The explanation consists in using the Langevin metric diagonalization procedure.
+The explanation consists of using the Langevin metric diagonalization procedure.
 The new coordinatization in the accelerated frame is interpreted by saying that the accelerated observer should obtain diagonal metric $ds^{2}=c^{2}(dt^{\prime}_{n})^{2}-(dx_{n}^{\prime})^{2}-(dy_{n}^{\prime})^{2}-(dz_{n}^{\prime})^{2}$ for the emitter at rest in the accelerated frame and the metric $ds^{2}=c^{2}(1-v^{2}/c^{2})(dt^{\prime}_{n})^{2}-2vdx_{n}^{\prime}dt_{n}^{\prime}-(dx^{\prime}_{n})^{2}-(dy^{\prime}_{n})^{2}-(dz^{\prime}_{n})^{2}$ for the moving emitter.
 Now we see that there is a symmetry (reciprocity) in metrics (and consequently in electrodynamics equations) between the accelerated and inertial frames.
-This is consequence of the pseudo-Euclidean geometry of space-time.
-In fact, the difference between this metric and metric Eq.
+This is a consequence of the pseudo-Euclidean geometry of space-time.
+The difference between this metric and metric Eq.
 (11) is only the sign of velocity.
 Clearly, after the diagonalization, the asymmetry is rooted in the initial conditions.
-The plane radiation wavefront of moving emitter (independently of its velocity) rotates in the accelerated frame on the angle $-v/c$ after metric diagonalization.
-Of course in order to predict the result of aberration measurement the analysis can be done using directly the metric of the moving emitter in the absolute time coordinatization, Eq.
+The plane radiation wavefront of the moving emitter (independently of its velocity) rotates in the accelerated frame on the angle $-v/c$ after metric diagonalization.
+Of course, to predict the result of aberration measurement the analysis can be done using directly the metric of the moving emitter in the absolute time coordinatization, Eq.
 (18).
-According to electrodynamics (see Section 5.4), the extra group velocity $-2v$ is introduced as a consequence of the crossed term $-4vdxdt$, the radiated beam of the moving emitter is propagated at the angle $-2v/c$.
+According to electrodynamics, the group velocity $-2v$ is introduced as a consequence of the crossed term $-4vdxdt$.
 Let us now consider the most general case when the system $S_{n}$ in the frame $S$ accelerated from the rest up to velocity $v$ along the $x$ axis and emitter in the accelerated frame $S_{n}$ is accelerated from rest up to velocity $u$ along the $x_{n}$ axis.
 Suppose that an observer in the accelerated frame $S_{n}$ performs an aberration measurement.
-We can describe radiation of the moving with velocity $u$ in the accelerated frame emitter in therms of the accelerated observer if we describe the electrodynamics by using metric $ds^{2}=c^{2}[1-(v-u)^{2}/c^{2}]dt_{n}^{2}-2(v-u)dx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$.
-At close look at the physics of this subject shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-v/c+u/c$.
+We can describe radiation of the moving with velocity $u$ in the accelerated frame emitter in terms of the accelerated observer if we describe the electrodynamics by using metric $ds^{2}=c^{2}[1-(v-u)^{2}/c^{2}]dt_{n}^{2}-2(v-u)dx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$.
+A close look at the physics of this subject shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-v/c+u/c$.
 The point is that the crossed term in metric Eq.(15), which generates the anisotropy in the accelerated frame, depends only on the velocity $v$ of the accelerated frame relative to the initial inertial frame.
-Figure 18: Aberration of light in an accelerated frame of reference $S_{n}$.
-Emitter is at rest with respect to the fixed stars.
-The aberration of light might be calculated from classical theory, in which the light beam may be treated, say, as a corpuscular (raindrops) beam.
-On the basis of classical theory one gets a light aberration angle that is only one-half as big as that predicted by special relativity.
-Figure 19: The deflection of starlight by the field of the sun.
-The light deflection might be calculated from classical theory, in which the light beam may be treated, say, as a comet approaches with the velocity of light.
-On the basis of classical theory one gets a deflection of light that is only one-half as big as that predicted by general relativity.
-5.9 A Useful Analogy
-We have presented theoretical evidence of an aberration increment $2v/c$ concerning the single emitter problem in a non-inertial frame of reference.
-First, we notice that there is an extra factor 2.
-On the basis of classical theory one gets aberration increment that is only one-half as big as that predicted by special relativity, Fig.
-18.
-According to classical approach, there is no principle difference between the aberration of light and the aberration of raindrops.
-Now let us consider a completely different situation which, in the classical case, gives deflection of light that is only one-half as big as that predicted by special relativity.
-There is a methodological analogy between the aberration of light effect and the deflection of light in general relativity.
-A ray of light which, coming from fixed star, passes close by the sun will thus be attracted to it and will describe a somewhat concave orbit with respect to the sun.
-This deflection might be calculated from Newton’s theory, in which the ray of light may be treated, say, as an asteroid which approaches with the velocity of light.
-We then get a formula similar to that of Einstein, but giving only half the value of the deflection, Fig.
+Figure 18: Aberration of light in an accelerated frame of reference.
+A point source is at rest in the accelerated frame.
+Wavefront orientation is presented in the absolute time coordinatization for the screen.
+Figure 19: Aberration of light in an accelerated frame of reference. A point source is at rest with respect to the fixed stars.
+5.9 A Point Source in an Accelerated Frame of Reference
+Now let us return to observations of an accelerated observer.
+The peculiarity of point-like sources is that radiation emitted at one instant forms a sphere around the source and the measuring instrument always influences the measured radiation, Fig.
+18 - Fig.
 19.
-Errors inherent in the classical method applied to the computation of the deflection of light are due to the use of the velocity of light simply as a classical velocity.
-According to classical approach, there is no principal difference between the deflection of light and the deflection of asteroid.
-A methodological analogy with the calculation of the aberration increment from classical (raindrops) theory emerges.
-Does this analogy have any physical meaning?
-To see whether it does, we should compare the motion of relativistic particles in the gravitation field with the aberration of particle effect.
-We shall do this in the Section 7.5.
-Figure 20: Aberration of light in an accelerated frame of reference.
-Point source is at rest in the accelerated frame.
-Wavefront orientation presented in the absolute time coordinatization for the screen.
-Figure 21: Aberration of light in an accelerated frame of reference. Point source is at rest with respect to the fixed stars.
-5.10 A Point Source in a Non-inertial Frame of Reference
-Now let us return to observations of a non-inertial observer.
-The peculiarity of point-like sources is that radiation emitted at one instant form a sphere around the source and the measuring instrument always influences the measured radiation, Fig.
-20-Fig.
-21 .
-The aberration of point source is considered independent of the source speed and to have just a local origin exclusively based on the observer (with his measuring instruments) speed relative to the fixed stars.
-The crossed term in metric Eq.(15) generates anisotropy in the accelerated frame that is responsible for the change of transmitted radiation direction (2).
+The aberration of a point source is considered independent of the source speed and to have just a local origin exclusively based on the observer’s speed relative to the fixed stars.
+The crossed term in metric Eq.(15) generates anisotropy in the accelerated frame that is responsible for the change of transmitted radiation direction.
 The point source follows the same pattern as fixed stars under the same aberration angle i.e.
 its apparent position changes with an angular displacement common to fixed stars.
-In the framework of the conventional theory of the aberration of light, there is an outstanding puzzle concerning the stellar aberration.
+In the framework of the conventional theory of the aberration of light, there is an outstanding puzzle concerning stellar aberration.
 There are double star systems the components of which change their velocity on a time scale ranging from days to years.
-The components of such binary systems at some times can have velocities relative to the earth very different from one another; nevertheless it is well known that these components exhibits always the same aberration angle.
-Rotating binary systems follow the same pattern as all fixed stars and are observed within a period of a year under the same universal aberration angle, i.e.
+The components of such binary systems at some times can have velocities relative to the earth very different from one another; nevertheless, it is well known that these components always exhibit the same aberration angle.
+Rotating binary systems follow the same pattern as all fixed stars and are observed within a year under the same universal aberration angle, i.e.
 their apparent position changes with an annual period common to all distant stars.
-This argument suggests that results of the astronomical observations confirm our prediction for aberration of light from a point source in an non-inertial frame.
-In the next chapter we discuss this experimental test in more detail.
-In regard to light aberration one should differentiate between that from the point source and that from the laser source.
-Let us analyze the aberration of light transmitted from the laser resonator which is at rest in the accelerated frame.
-A close examination of of all experiments inside the accelerated frame shows that all optical phenomena in an optical resonator appeared to be independent of the acceleration relative to the fixed stars.
-It should be note that there is no influence of the difference between Langevin and Minkowski metrics on the parameter of a laser source.
-It is not hard to understand this result.
-It is clear that the crossed term in the Langevin metric, which generates aberration, cancels during the (round-trip) evolution of the radiation in the laser resonator.
-We will analyze the aberration of light transmitted from the laser resonator in great detail in the Section 8.6
+This argument suggests that the results of the astronomical observations confirm our prediction for the aberration of light from a point source in a rotating frame.
+In the next section, we discuss this experimental test in more detail.
 5.10 Tilted Incoming Plane Wave
-We go back now to the situation shown in Fig.20 - 21.
+We go back now to the situation shown in Fig.18 - Fig.
+19.
 A point source in the far zone produces in front of a hole aperture effectively a plane wave.
-In this chapter we considered only the case in which the point source on the optical axis (which is parallel to the $z_{n}$ axis) generates normally incident plane wave.
-It is obvious that the electrodynamics equations must be identical for all electromagnetic waves.
-In other words, the dispersion equation in the accelerated frame should be applied and kept in a consistent way for both incoming and scattered waves.
+In this section, we considered only the case in which the point source on the optical axis (which is parallel to the $z_{n}$ axis) generates a normally incident plane wave.
+The electrodynamics equations must be identical for all electromagnetic waves.
+In other words, the dispersion equation in the accelerated frame should be applied and kept consistently for both incoming and scattered waves.
 A dispersion equation in the case of Maxwell’s electrodynamics is reduced to $k_{x}^{2}+k_{y}^{2}+k_{z}^{2}-\omega^{2}=0$.
-The simplicity of considered case is that even after the Galilean transformation along the $x$-axis the dispersion equation in the accelerated frame will have the same (diagonal) form $k_{z}^{2}-\omega_{i}^{2}=0$ for the normal incident wave.
-Now let us see what happens with incoming wave which has rate of change of phase with respect to $x_{n}$.
+The simplicity of the considered case is that even after the Galilean transformation along the $x$-axis the dispersion equation in the accelerated frame will have the same (diagonal) form $k_{z}^{2}-\omega_{i}^{2}=0$ for the normal incident wave.
+Now let us see what happens with an incoming wave which has a rate of change of phase with respect to $x_{n}$.
 In other words, there is the angle $\alpha$ between the wave number vector $\vec{k}$ and the $z_{n}$-axis.
@@
 Let us assume that the wavefront of the emitted light beam is tilted by an angle $\alpha$ before the acceleration.
-Then we have $k_{x}=\alpha k$, $k_{z}=\sqrt{\omega^{2}_{i}/c^{2}-k_{x}^{2}}$.
-Acceleration has an effect on the field equations.
-With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ equation Eq.(16) becomes: $k_{x}^{2}-2vk_{x}\omega_{i}/c^{2}+k_{z}^{2}-(\omega_{i}-vk_{x})^{2}/c^{2}=0$ in the first order approximation with respect to a small parameter $v/c$.
-The screen containing a hole is a kind of diffraction grating which breaks up the incident plane wave into a number of diffracted plane waves.
-So the wave-number vector of the transmitted plane wave along the $x_{n}$ direction is $k_{x}+K_{\perp}$ where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated transmittance.
-From the dispersion equation, we find the requirement that the wavenumber $K_{\perp}$ and the frequency change $\Delta\omega$ are related by $\omega_{i}\Delta\omega=-\omega_{i}K_{\perp}v+c^{2}k_{x}K_{\perp}$.
-The last equations state that transmitted light beam moves along the $x_{n}$ direction with group velocity $d\omega/dk_{x}=-v+c^{2}k_{x}/\omega_{i}$.
-Then the apparent angular position of the point source is connected with the problem parameters by the relation $\theta=-v/c+\alpha$.
-Let us now look again at our dispersion equation.
-At first site the Doppler effect problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers.
-The question arises whether it is possible to determine experimentally the state of motion of the accelerated frame $S_{n}$ in the initial inertial frame $S$ by means of the Doppler shift $\Delta\omega=vk_{x}$ for accelerated source.
-In the Chapter 10 we demonstrate that there exist a specific feature of the Doppler shift measurements which prevents us to determine the velocity of the accelerated frame with respect to the initial inertial frame.
+Then we have $k_{x}=\alpha k=\alpha\omega_{i}/c$, $k_{z}=\sqrt{\omega_{i}^{2}/c^{2}-k_{x}^{2}}$.
+Acceleration affects the field equations.
+With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ equation Eq.(17) becomes: $(1-v^{2}/c^{2})k_{x}^{2}-2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$.
+The screen containing a hole is a kind of diffraction grating that breaks up the incident plane wave into a number of diffracted plane waves.
+So the wave-number vector of the transmitted plane wave along the $x_{n}$ direction is $k_{x}=\alpha k+K_{\perp}$ where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated transmittance.
+In keeping with our approximation, we should leave out the terms $v^{2}/c^{2}$, $\alpha^{2}$, and $\alpha v/c$.
+Our dispersion equation now looks like $\omega_{i}\Delta\omega=-\omega_{i}K_{\perp}v+c^{2}\alpha kK_{\perp}$.
+The last equations state that the transmitted light beam moves along the $x_{n}$ direction with group velocity $d\omega/dk_{x}=-v+c^{2}\alpha k/\omega_{i}$.
+Then the apparent angular position of the point source is connected with the problem parameters by the relation $\theta=-v/c+\alpha$ [^26] A similar situation would occur if we consider the case of a screen, in the inertial frame, moving with velocity $v$ along the $x$-axis.
+In Chapter 4, we have taken a simple example where a plane wave falls normally on the screen.
+But of course, we can generalize it.
+Let us assume that the wavefront of the incoming plane wave is tilted by an angle $\alpha$.
+According to the presented approach, there is a deviation of the energy transport for light transmitted through a hole in the moving screen.
+The aberration increment is connected with problem parameters by the relation $\theta=v/c+\alpha$..
 5.11 Absence of Reciprocity in the Aberration of Light Theory
-In the present section we shall continue our discussion of the aberration of light radiated by a single moving emitter.
-Imaging that there are two identical emitters.
+In the present section, we shall continue our discussion of the aberration of light radiated by a single moving emitter.
+Imagine that there are two identical emitters.
 Let us consider the case when the first emitter is at rest in an inertial frame and the second emitter is accelerated from rest up to velocity $v$ along the $x$ axis.
 Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the energy transport measurement.
-Now we must be careful about initial phasing of these emitters.
-As example, we consider the case in which initially the velocity component of the light beam along the $x$-axis is equal to zero.
-Then how does the light beam from the moving emitter looks?
+Now we must be careful about the initial phasing of these emitters.
+For example, we consider the case in which initially the velocity component of the light beam along the $x$-axis is equal to zero.
+Then how does the light beam from the moving emitter look?
 The inertial observer would find that angular displacement is equal to $\theta_{a}=v/c$.
 That is the radiation appears as shown in Fig.
-22.
-This is example what called the phenomenon of aberration of light and it is well known.
-Now let us return to observation of an accelerated observer.
-When the accelerated system starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed.
-In this synchronization, the accelerated observer describes the light beam from an emitter at rest using usual Maxwell’s equations.
-One may ask: ”Where is the information about the observer acceleration recorded in the case of the Lorentz clock synchronization?” When the accelerated observer performed the standard procedure of clock synchronization, the time shift has the effect of rotation the plane of simultaneity, that is source radiation wavefront, on the angle $-v/c$.
-It is easy to show that information about the observer acceleration is recorded in the phase front orientation (with respect to the coordinate axes of the accelerated frame) of the light beam radiated from the source at rest in the accelerated frame.
+20.
+This is an example of what is called the phenomenon of aberration of light and it is well known.
 Figure 20: The aberration of light from stationary and moving sources.
-According to the asymmetry between the initial inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of light.
-Namely, if the emitter is at rest relative to the fixed stars and the observer with measuring devices started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the emitter seen in the accelerated frame would jump by angle $\theta_{a}=-2v/c$.
-This situation is not symmetrical with respect to the change of the reference frames.
-If the observer is at rest relative to the fixed stars and the emitter started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the emitter seen in the inertial frame would jump by angle $\theta_{a}=v/c$.
-Figure 23: Reciprocity in a theory of the aberration of light.
-One must use the same source phasing to demonstrate the reciprocity.
-After the redirection procedure the inertial observer would find that angular displacement is $\theta_{a}=2v/c$.
-Let us describe what happens when accelerated observer performed the redirection of the accelerated emitter.
+Figure 21: Aberration of light when the accelerated observer redirected the accelerated emitter.
+After the redirection procedure, the inertial observer would find that angular displacement is $\theta_{a}=2v/c$.
+Let us describe what happens when the accelerated observer performed the redirection of the accelerated emitter.
 We consider the case in which finally the (group) velocity component of the light beam along the $x_{n}$-axis is equal to zero.
-After this redirection procedure the inertial observer would find that angular displacement is $\theta_{a}=2v/c$.
+After this redirection procedure, the inertial observer would find that angular displacement is $\theta_{a}=2v/c$.
 Suppose that the accelerated observer also performs an aberration measurement.
@@
 21 shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-2v/c$.
-The situation can be described quite naturally in the following way.
-In this case we define reference directions using light beams.
-We demonstrated that symmetry is a correct concept in the measurements of the angular displacement between the inertial $z$ and accelerated $z_{n}$ reference directions as it must be.
-Figure 24: Procedure of the accelerated source redirection.
+Figure 22: The procedure of the accelerated source redirection.
 The information about the source acceleration is recorded in the point source offset with respect to the optical axis (which is parallel to the $z_{n}$ axis).
 The wavefront orientation has no exact objective meaning.
-Because of this, the explanation of the effect of aberration (for sketch simplification) is based on the use of the absolute time (left) and Lorentz (right) coordinatization to describe how the direction of a light beam depends on the source offset.
+Because of this, the explanation of the effect of aberration (for sketch simplification) is based on the use of absolute time (left) and Lorentz (right) coordinatization to describe how the direction of a light beam depends on the source offset.
 Let us return to the redirection procedure of the accelerated source.
-Now the question is, where is the information about the source acceleration recorded in the case of redirecting light of the accelerated emitter.
-In order to understand the redirection of the light source in an accelerated system, we have to watch the machinery of the source and see what happens when the redirection takes place.
-Fig.
-13 shows a drawing of the ”plane wave” emitter type.
-Electrodynamics offers a procedure of redirection based on the offset of the point source with respect to the optical axis (which is parallel to the $z_{n}$ axis), Fig.
-24.
+Now the question is, where is the information about the source acceleration recorded in the case of redirecting light of the accelerated emitter?
+To understand the redirection of the light source in an accelerated system, we have to watch the machinery of the source and see what happens when the redirection takes place.
+Fig.13 shows a drawing of the ”plane wave” emitter type.
+Electrodynamics offers a procedure of redirection based on the offset of the point source with respect to the optical axis (which is parallel to the $z_{n}$ axis), Fig.22.
 The information about the source acceleration is recorded in the point source offset with respect to the optical axis.
-It is generally believed that the special relativity is reciprocal theory.
-The aberration of light problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers.
-In fact, without looking at anything external to the accelerated frame, one could determine the speed of the accelerated frame with respect to the initial inertial frame by means of aberration of point source measurements.
-No one has ever done such experiment, but we know what would happen from the astronomical observations.
-The motion of the stars with respect to the earth is never followed by any aberration.
-The aberration shift exist even in the case when star moves with the same velocity as the earth.
-There is an optical similarity between aberration of light from distant star (which is moving with the earth velocity) and from the earth-based point source.
-We will discuss this subject further in the Chapter 8.
-One finds many books which say that we cannot distinguish by any experiment which observer remains ”at rest” during the acceleration, because according to the principle of relativity only relative motion has any physical meaning.
-It follows that any observed asymmetry would lead to a contradiction with the principle of relativity.
-This argument is wrong.
-We discuss the apparent conflict between aberration of light and the principle of relativity in the Chapter 12.
-It is demonstrated that there is no conflict between the fundamental structure of special relativity on the one hand, and the aberration of light phenomena.
-Principle of special relativity is irrelevancy of velocity with regard to physical laws, not with regard to anything.
 5.12 Two Metrics
-There is an important theorem in the special relativity which is this: as a result of the Galilean boost, we obtain metric of the accelerated emitter in the initial inertial frame Eq.
+There is an important theorem in special relativity which is this: as a result of the Galilean boost, we obtain a metric of the accelerated emitter in the initial inertial frame Eq.
 11.
-In this coordinatization, we describe the non-accelerated light source using diagonal (Minkowski) metric.
-We transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer using passive Galilean transformation: we substitute $x\to x-vt,t\to t$ into Minkowski metric of the inertial observer to obtain Eq.11.
-This metric describes the electrodynamics of the moving emitter with viewpoint of the inertial observer measurements.
-The Langevin metric describes the electrodynamics of the accelerated emitter with viewpoint of the accelerated observer measurements as viewing this of the inertial observer.
-We transform coordinates of the inertial observer using the inverse Galilean transformation: we substitute $x\to x_{n}+vt,t\to t_{n}$ into Minkowski metric of the inertial observer to obtain Langevin metric Eq.15.
+In this coordinatization, we describe the non-accelerated light source using the diagonal (Minkowski) metric.
+We transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer using the equivalence of the active and passive pictures: we substitute $x\to x-vt,t\to t$ into Minkowski metric of the inertial observer to obtain Eq.11.
+This metric describes the electrodynamics of the moving emitter with the viewpoint of the inertial observer measurements.
+The Langevin metric describes the electrodynamics of the accelerated emitter with the viewpoint of the accelerated observer measurements as viewing this of the inertial observer.
+One must take into account that the metric tensor must be a continuous quantity.
+This problem in special relativity can be adequately treated only by an approach that uses absolute time coordinatization.
+For this purpose, it is sufficient to relate the coordinates and time to the Galilean boosts.
+We transform the coordinates of the inertial observer using the inverse Galilean transformation: we substitute $x\to x_{n}+vt,t\to t_{n}$ into Minkowski metric of the inertial observer to obtain Langevin metric Eq.15.
 Now we see that the Langevin metric Eq.15 of the accelerated emitter in the accelerated frame is the same as metric Eq.11 of the accelerated emitter in the inertial frame.
-In fact the difference is only that the sign of velocity $v$.
-These two metrics are actually two sides of the same coin.
+The difference is only that the sign of velocity $v$.
+These two metrics are two sides of the same coin.
 The point is that we describe the electrodynamics of the same accelerated (with respect to the fixed stars) emitter from the point of view of both an inertial observer and an accelerated observer.
-One must take into account that metric tensor must be continuous quantity.
-This problem in special relativity can be adequately treated only by an approach uses absolute time coordinatization.
-For this purpose, it is sufficient to relate the coordinates and time by the Galilean boosts.
 The commonly accepted approach to special relativity does not account for this theorem.
 The metric Eq.11 is represented here for the first time.
-It is difficult to see why this should be, since Langevin metric (which is found by matching the accelerated frame and inertial frame metric tensors by the Galilean boost) is widely used in the applications to the optical phenomena in a rotation frame.
+It is difficult to see why this should be since the Langevin metric (which is found by matching the accelerated frame and inertial frame metric tensors by the Galilean boost) is widely used in applications to the optical phenomena in a rotating frame.
 In the problem discussed we have still a puzzle.
-We create a so-called comoving coordinate system in the initial inertial frame, and then discussing radiation from the moving source in terms of the new (comoving) coordinate labels.
-After the passive transformation, the coordinates of the source transform as $x^{\prime}=x-vt,t^{\prime}=t$.
-According to the equivalence of the active and passive pictures within a single inertial frame, in our case of interest the Maxwell’s equations are always valid in the comoving coordinatization.
+We create a so-called comoving coordinate system in the initial inertial frame and then discuss radiation from the moving source in terms of the new (comoving) coordinate labels.
+After the passive transformation, the source coordinates transform as $x^{\prime}=x-vt,t^{\prime}=t$.
+According to the equivalence of the active and passive pictures within a single inertial frame, in our case of interest, Maxwell’s equations are always valid in the comoving coordinatization.
 The electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$.
 In other words, Minkowski metric $ds^{2}=c^{2}(dt^{\prime})^{2}-(dx^{\prime})^{2}-(dy^{\prime})^{2}-(dz^{\prime})^{2}$ describes the electrodynamics of the accelerated light source in comoving coordinatization.
-We will then ask how it is possible to take Minkowski metric for comoving coordinatization in the inertial frame and the Langevin metric in the accelerated frame.
+We will then ask how it is possible to take the Minkowski metric for comoving coordinatization in the inertial frame and the Langevin metric in the accelerated frame.
 This approach needs an explanation.
 The difference between measurements made in both reference frames is very interesting.
-First, we must now discuss a certain feature of the correspondence between passive and active transformations.
-It should be note that passive Galilean boost within a single inertial frame, $x^{\prime}=x-vt,t^{\prime}=t$, is simply another parametrization of the observations of the inertial observer.
+First, we must discuss a certain feature of the correspondence between passive and active transformations.
+It should be noted that passive Galilean boost within a single inertial frame, $x^{\prime}=x-vt,t^{\prime}=t$, is simply another parametrization of the observations of the inertial observer.
 The Eq.11 describes the inertial observer measurements.
-According to active transformations, the motion of fixed stars with respect to the inertial observe and his devices not changes and there is equivalence of the active and passive Galilean boost.
+According to active transformations within a single inertial frame, the motion of fixed stars with respect to the inertial observer and his devices does not change and there is the equivalence of the active and passive Galilean boost.
 In contrast, the Langevin metric describes the accelerated observer measurements.
@@
 On the other hand, the standard textbook presentation of the relativistic transformation is based on the hidden assumption that $x_{n}y_{n}z_{n}$-axes of the accelerated observer and $xyz$-axes of the inertial observer are parallel.
-In other words, it is assumed that observers have common 3-space.
+In other words, it is assumed that observers have a common 3-space.
 It is generally believed that simultaneous acceleration of the (rigid) reference space grid has direct physical meaning.
-This is misconception.
+This is a misconception.
 The commonly accepted derivation of the aberration of light effect does not account for the conventionality of distant simultaneity.
 There is an uncertainty of the relative position along the $z_{n}$-axis in the $x$-direction of amount $z_{n}v/c$ due to the uncertainty in the moments of acceleration.
-In fact in this case we have deal with space-like separated events.
-The $z_{n}$ axis orientation has no exact objective meaning, because the relativity of simultaneity takes place.
-The statement that the $z_{n}$-axis orientation with respect to the $x$-axis of the inertial observer has objective meaning to within a certain accuracy can be visualized by the picture of $z_{n}$ axis in the vertical orientation with approximate angle extension (blurring) given by $\Delta\theta=v/c$.
-This inclination angle between the inertial $z$ and accelerated $z_{n}$ reference directions has no exact objective meaning at all since, due to finiteness of the speed of light, we cannot specify any experimental method by which this displacement could be ascertained.
-We shall discuss this subject further in the Chapter 7.
+In this case we have to deal with space-like separated events.
+The $z_{n}$ axis orientation has no exact objective meaning because the relativity of simultaneity takes place.
+The statement that the $z_{n}$-axis orientation with respect to the $z$-axis of the inertial observer has objective meaning to within a certain accuracy can be visualized by the picture of $z_{n}$ axis in the vertical orientation with approximate angle extension (blurring) given by $\Delta\theta=v/c$.
+This inclination angle between the inertial $z$ and accelerated $z_{n}$ reference directions has no exact objective meaning at all since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this displacement could be ascertained.
+We shall discuss this subject further in Chapter 7.
 5.13 Ether Theory and the Aberration of Light Effect
-5.14.1 Introductory Remarks
 Now we wish to continue in our analysis a little further.
-We will look for a different way of calculating the aberration effect.
 Another solution to the accelerated emitter problem will be discussed going back to ether theory.
-We accept the ether theory in the original form: there is ether, which rules the speed of light.
-Because the ether is immovable, it is causes anisotropy in every frames moving relative to the ether.
+We accept the ether theory in its original form: there is ether, which rules the speed of light.
+Because the ether is immovable, it causes anisotropy in every frame moving relative to the ether.
 Because the non-inertial frame is accelerated with respect to the fixed stars; consequently we must feel an ”ether wind” when measuring light propagation.
 Firstly we discuss the effect of the ether wind on the light speed.
-According to the hypothesis of an ether at rest in the initial inertia frame, the velocity of light judged from an accelerated reference system would be $c+u$ for the beam propagating in the same direction as the accelerated system and $-c+u$ for the beam propagating in the opposite direction, where $u=-v$ is the ether velocity in the accelerated reference system.
+According to the hypothesis of an ether, the velocity of light judged from an accelerated reference system would be $c+u$ for the beam propagating in the same direction as the accelerated system and $-c+u$ for the beam propagating in the opposite direction, where $u=-v$ is the ether velocity in the accelerated reference system.
 The aberration effect is an effect of the ether-wind on the speed of light (Fig.
-25).
+23).
 Figure 23: Aberration of light in an accelerated frame.
-According to the ether theory, the ether wind generates anisotropy in the accelerated ( relative to the ether) frame that is responsible for change of radiation direction.
+According to the ether theory, the ether wind generates anisotropy in the accelerated ( relative to the ether) frame that is responsible for the change of radiation direction.
 5.13.1 Lorentz’s Theory of Corresponding States
-In 1895, Lorentz derived a first-order version of the so-called ”theorem of the corresponding states” and used it to give a very general explanation of the null results of the first order optical interference experiments.
-Lorentz’s theorem connects pairs of electromagnetic field configurations, one in a system at rest in the ether and one in a system in uniform motion through the ether, and say that, if one member of the pair is allowed by the laws of electrodynamics, so is the other.
+In 1895, Lorentz derived a first-order version of the so-called ”theorem of the corresponding states” and used it to give a very general explanation of the null results of the first-order optical interference experiments.
+Lorentz’s theorem connects pairs of electromagnetic field configurations, one in a system at rest in the ether and one in a system in uniform motion through the ether, and says that, if one member of the pair is allowed by the laws of electrodynamics, so is the other.
 The theorem works as follows.
-The mathematical description of the configuration in the moving system in therms of ”local time” (instead of ”real time” ) is the same as the description of its corresponding state in the system at rest in terms of the real time.
+The mathematical description of the configuration in the moving system in terms of ”local time” (instead of ”real time” ) is the same as the description of its corresponding state in the system at rest in terms of real time.
 Lorentz showed that two field configurations related to one another in many of their observable properties, in particular, they give rise to identical interference patterns.
 Thus, it is not possible to detect the orbital velocity using interference (e.g.
-Michelson-Morley) experiments (2).
+Michelson-Morley) experiments  [^27] For discussion of the relationship between the generally accepted way of looking at special relativity and Lorentz’s ether theory we suggest reading the paper [32]..
 We now discuss the application of Lorentz’s pre-relativistic theory to the aberration of light effect in a non-inertial frame of reference.
-Above we considered the problem of aberration of light in the accelerated systems on the basis of theory of special relativity.
+Above we considered the problem of aberration of light in the accelerated systems based on the theory of special relativity.
 We described the effect of aberration by working only up to the first order.
-In fact, the appearance of relativistic effects in optical phenomena does not depend on a large speed of the radiation sources.
-Light is always a relativistic object.
+We already know that light is always a relativistic object.
 We have already discussed the asymmetry between initial inertial frame and accelerated frames in this chapter.
-Here’s what we found out there: We pointed that metric tensor must be a continuous quantity.
-This problem in special relativity can be adequately treated only by an approach which uses the absolute time coordinatization.
-The usual Maxwell’s equations are valid in the initial inertial frame where optical system (e.g.
+Here’s what we found out there: We pointed out that the metric tensor must be a continuous quantity.
+This problem in special relativity can be adequately treated only by an approach that uses absolute time coordinatization.
+The usual Maxwell’s equations are valid in the initial inertial frame where the optical system (e.g.
 configuration of light source and lens, Fig.
-13) is at rest and procedure of slow clock transport is used to assign values of time coordinate.
-We require that metric of a non inertial frame associated with accelerated (with respect to the initial inertial frame) optical system would pass continuosly into Langevin metric.
+13) is at rest and the procedure of slow clock transport is used to assign values of time coordinate.
+We require that the metric of a noninertial frame associated with an accelerated (with respect to the initial inertial frame) optical system pass continuously into the Langevin metric.
 For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer $(x_{n},y_{n},z_{n},t_{n})$ to the coordinates $(x,y,z,t)$ of the inertial observer by the inverse Galilean boost $x_{n}=x-vt,y_{n}=y,z_{n}=z$, $t_{n}=t$.
@@
 17 where coordinates and time are transformed according to the Galilean transformation.
-The Langevin metric, together with initial condition, describes optical effects in the accelerated frame.
-Above in this chapter we considered the aberration of light problem in the accelerated frame.
+The Langevin metric, together with the initial condition, describes optical effects in the accelerated frame.
+Above in this section, we considered the aberration of light problem in the accelerated frame.
 The presented approach uses the Fourier transform method.
 The essential asymmetry is rooted in the electrodynamics equation Eq.
-16.
-As one of the consequences of the crossed term in the wave equation we found the group velocity that is responsible for the aberration.
+17.
+As one of the consequences of the crossed term in the wave equation, we found the group velocity that is responsible for the aberration.
 However, there is another satisfactory way of explaining the effect of aberration of light in the accelerated frame.
-The explanation consists in using a procedure of clock resynchronization.
-Using Einstein’s clock synchronization we actually select a Lorentz coordinate system for the optical setup at rest in the accelerated frame.
-In this synchronization we describe the field configuration using usual Maxwell’s equations.
-The time $t^{\prime}_{n}$ under the Einstein’s synchronization is obtained by substituting $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first order approximation.
+The explanation consists of using a procedure of clock resynchronization.
+Using Einstein’s clock synchronization, we select a Lorentz coordinate system for the optical setup at rest in the accelerated frame.
+In this synchronization, we describe the field configuration using Maxwell’s equations.
+The time $t^{\prime}_{n}$ under Einstein’s synchronization is obtained by substituting $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first-order approximation.
 Clearly, after the resynchronization, the asymmetry is rooted in the initial condition.
 The plane radiation wavefront of the emitter rotates in the accelerated frame on the angle $-v/c$ after resynchronization.
-Lorentz’s treatment of optical setup in an inertial frame moving through the ether is similar to our treatment of such setup in the accelerated (with respect to the fixed stars) frame.
-Now we must analyze this similarities.
+At first sight, Lorentz’s treatment of optical setup in an inertial frame moving through the ether is similar to our treatment of such setup in the accelerated (with respect to the fixed stars) frame.
+Now we must analyze this similarity.
 One thing is clear: Our initial inertial frame (i.e.
 inertial frame without accelerational history with respect to the fixed stars) can be interpreted as what Lorentz thought of as a frame at rest in the ether.
-In the Lorentz’s theory, and more generally in Physics before the development of special relativity, the connection between an observer at rest who used $(x,y,z,t)$ coordinates and an observer bound to the moving body who used $(x_{n},y_{n},z_{n},t)$ coordinates was given by the Galileo relation: $x_{n}=x-vt,y_{n}=y,z_{n}=z,t_{n}=t$.
+In Lorentz’s theory, and more generally in Physics before the development of special relativity, the connection between an observer at rest who used $(x,y,z,t)$ coordinates and an observer bound to the moving body who used $(x_{n},y_{n},z_{n},t)$ coordinates was given by the Galileo relation: $x_{n}=x-vt,y_{n}=y,z_{n}=z,t_{n}=t$.
 Galilean boost in Lorentz’s theory is empirically equivalent to the smooth tailoring of the metric tensor.
-In this case the metric of initial inertial frame would pass into the Langevin metric of the accelerated frame.
+In this case, the metric of the initial inertial frame would pass into the Langevin metric of the accelerated frame.
 Lorentz of course would not have thought of his theory in this way, since he did not have the resources of Minkowski space-time to do so.
-Because of this Galilean boost, the homogeneous wave equation for field in the moving frame, Eq.
-16, has not standard form.
+Because of this Galilean boost, the homogeneous wave equation for the field in the moving frame, Eq.
+17, has no standard form.
 The main difference consists in the crossed term, which complicates the solution of the equation.
 To get around of this difficulty, Lorentz observed that simplification is always possible.
-In solving wave equation, Lorentz found it convenient to introduce new variable $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$.
+In solving wave equation, Lorentz found it convenient to introduce a new variable $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$.
 It is called ”local time”.
 The substitution of this time shift in Lorentz theory is a mathematical device used to change the electrodynamics equations into Maxwell’s form.
-At first cite, it seems that this Lorentz’s theorem of corresponding states for first order phenomena provides all the guidance we need to exact treatment of aberration of light in the accelerated frame.
-One could naively expect that Lorentz’s theory would predict in full generality that observation of aberration of light from the earth-based incoherent source would reveal the earth orbital motion, Fig.
-24.
-At this point, a reasonable question arises: why Lorentz concluded from his theorem of corresponding states that no first order effects of the orbital motion of the earth could be detected?
+At first sight, it seems that Lorentz’s theorem of corresponding states for first-order phenomena provides all the guidance we need to exact treatment of the aberration of light in the accelerated frame.
+One could naively expect that Lorentz’s theory would predict in full generality that observation of aberration of light from the earth-based incoherent source would reveal the earth’s orbital motion, Fig.
+22.
+At this point, a reasonable question arises: why did Lorentz conclude from his theorem of corresponding states that no first-order effects of the orbital motion of the earth could be detected?
 5.13.2 Analogy with the Sound Theory
 A comparison between the Lorentz ether theory and the sound theory might help here.
-Above in the Section 4.15 we considered the sound emitter problem in the atmosphere frame.
+Above in Section 4.13, we considered the sound emitter problem in the atmosphere frame.
 We considered the case of a screen, in the atmosphere frame, moving with velocity $v$ along its surface.
 According to the theory of sound, there is no deviation of the energy transport for sound transmitted through a hole in the moving screen.
-A plane wave is falling normally on the screen, and generates a transmitted oblique sound beam, Fig.
+A plane wave falls normally on the screen and generates a transmitted oblique sound beam, Fig.
 7.
 The group velocity is equal to the phase velocity.
-The wave equation $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$ always valid in the atmosphere frame.
+The diagonal wave equation $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$ always valid in the atmosphere frame.
 This equation is not applicable from the viewpoint of an observer at rest with respect to the accelerated frame.
-Accelerations with respect to the atmosphere frame have an effect on the propagation of sound.
+Accelerations with respect to the atmosphere frame affect the propagation of sound.
 In particular, the wave equation transforms into
@@
 where coordinates and time are transformed according to the Galilean transformation $x_{n}=x-vt$, $t_{n}=t$.
-The wave equation needs to be integrated with initial condition for the sound wavefront.
-Acceleration has no effect on the wavefront orientation.
-So that wavefront of the emitted sound beam is perpendicular to the $z_{n}$ axis.
-Because the air is immovable, it is causes anisotropy in every frames moving relative to the atmosphere frame.
-Consequently we must feel an ”air wind” when measuring sound propagation.
+The wave equation needs to be integrated with the initial condition for the sound wavefront.
+Acceleration does not affect the wavefront orientation.
+So wavefront of the emitted sound beam is perpendicular to the $z_{n}$ axis.
+Because the air is immovable, it causes anisotropy in every frame moving relative to the atmosphere frame.
+Consequently, we must feel an ”air wind” when measuring sound propagation.
 According to the wave equation Eq.
-18, the velocity of sound judged from an accelerated reference system would be $v_{s}-v$ for the beam propagating in the same direction as the accelerated system and $-v_{s}-v$ for the beam propagating in the opposite direction, where $-v$ is the wind velocity in the accelerated reference system.
+19, the velocity of sound judged from an accelerated reference system would be $v_{s}-v$ for the beam propagating in the same direction as the accelerated system and $-v_{s}-v$ for the beam propagating in the opposite direction, where $-v$ is the wind velocity in the accelerated reference system.
 The aberration effect is an effect of the air wind on the speed of sound (Fig.
-25).
+23).
 Let us now analyze the aberration of sound radiated by a single ”plane-wave” emitter in the atmosphere frame.
-The wavefront radiated by the plane-wave emitter at given instant time is constructed as an envelope of all the spherical wavelets from the point sources on the emitter that have radiated until this instant.
+The wavefront radiated by the plane-wave emitter at a given instant time is constructed as an envelope of all the spherical wavelets from the point sources on the emitter that have radiated until this instant.
 It is not difficult to see that when the emitter is moving at a constant velocity $v$ along the $x$ axis, the wavefront will propagate at the speed $v_{s}$ at the angle $\theta=v/v_{s}$ from the vertical.
@@
 Now let us return to observations of an accelerated observer.
-The wave equation needs to be integrated with initial condition for the sound wavefront.
-Galilean boost has no effect on the wavefront orientation.
-So that wavefront of the emitted sound beam is tiled at the angle $v/v_{s}$ with respect to the $z_{n}$ axis.
-The wave number vector of the radiated plane wave is fixed by initial condition.
+The wave equation needs to be integrated with the initial condition for the sound wavefront.
+Galilean boost does not affect the wavefront orientation.
+So the wavefront of the emitted sound beam is tiled at the angle $v/v_{s}$ with respect to the $z_{n}$ axis.
+The wave number vector of the radiated plane wave is fixed by the initial condition.
 Then we have $k_{x}=vk/v_{s}=v\omega_{s}/v_{s}^{2}$.
 The non-diagonal wave equation Eq.
-18 states that radiated sound beam moves along $x_{n}$ axis with group velocity $d\omega/dk_{x}=-v+v_{s}^{2}k_{x}/\omega_{s}=0$ (see Section 5.11 for more detail).
-Thus we obtain the same addition theorem for sound beam velocities as for the particle velocities in the classical physics.
-So that the two (ether and sound) theories were really just variants of the same classical theory.
-According to Lorentz, the wave equation of the moving light (sound) emitter in the ether (atmosphere) frame is the same as wave equation of the emitter at rest.
-As a consequence of this, after the boost we can see that screen acceleration in the ether (atmosphere frame) has no effect on the wavefront orientation and the wave equation.
-Thus Lorentz concluded that one could not determine the orbital motion of the earth by means of aberration of earth-based source measurements.
-We come to the conclusion that the Lorentz analysis of the aberration of light does not take into account the fundamental difference between velocity of light and velocity of sound.
-When Lorentz formulated his pre-relativistic theory he assumed a space-time background with a well defined geometrical structure: Newtonian space-time.
-At close look at the physics of this subject shows that there is a real difference between the inertial frame without accelerational history with respect to the fixed stars and the Lorentz classical ether at rest.
-The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of unified space-time model.
-In special relativity the absolute space and the absolute time of Newton are fused into absolute space-time.
-5.15 Bibliography and Notes
-1.
-The aberration of light problem in the accelerated systems is solved with discovery of the essential asymmetry between the non-inertial and the inertial observers.
-This asymmetry is of the same nature as that of the well-known Sagnac effect [28, 29, 30].
-For instance Langevin’s 1921 explanation of the Sagnac effect rested upon the assertion that ”any change of velocity, or any acceleration has an absolute meaning.” [31].
-2. For discussion of the relationship between the general accepted way of looking at special relativity and the Lorentz’s ether theory we suggest reading the paper [32].
+19 states that the radiated sound beam moves along $x_{n}$ axis with group velocity $d\omega/dk_{x}=-v+v_{s}^{2}k_{x}/\omega_{s}=0$ (see Section 5.10 for more detail).
+Thus we obtain the same addition theorem for sound beam velocities as for the particle velocities in classical physics.
+According to Lorentz, the wave equation of the moving light (sound) emitter in the ether (atmosphere) frame is the same as the wave equation of the emitter at rest.
+Thus Lorentz concluded that one could not determine the orbital motion of the earth using aberration of earth-based source measurements.
+So the two (ether and sound) theories were just variants of the same classical theory.
+We conclude that the Lorentz analysis of the aberration of light does not take into account the fundamental difference between the velocity of light and the velocity of sound.
+When Lorentz formulated his pre-relativistic theory he assumed a space-time background with a well-defined geometrical structure: Newtonian space-time.
+A close look at the physics of this subject shows that there is a real difference between the inertial frame without an accelerational history with respect to the fixed stars and the Lorentz classical ether at rest.
+The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of a unified space-time model.
+In special relativity, the absolute space and the absolute time of Newton are fused into absolute space-time.
 6 Stellar Aberration
 6.1 The Corpuscular Model of Light and the Stellar Aberration
-It is generally believed that the phenomenon of aberration of light could be interpreted, using the corpuscular model of light, as being analogous of the observation of the oblique fall of raindrops by a moving observer.
-This is a classical kinematics method to the computation of the stellar aberration used in astronomy for about three hundred years (1).
+It is generally believed that the aberration of light could be interpreted, using the corpuscular model of light, as being analogous to observing the oblique fall of raindrops by a moving observer.
+This is a classical kinematics method to the computation of the stellar aberration used in astronomy for about three hundred years  [^28] The phenomenon of the annual apparent motion of celestial objects about their locations, named stellar aberration, was discovered by Bradley in 1727, who also explained it employing the corpuscular model of light [33]..
 What had to be added in the 20th century, was that the dynamical laws of Newton for light were found to be all wrong, and electromagnetic wave theory had to be introduced to correct them.
 As well-known, according to textbooks the physical basis of the stellar aberration is the fact that the velocity of light is finite and changes its direction when seen from another reference frame.
-It is a consequence of the formula for addition of velocities applied to a light beam when the observer is changing its reference frame.
-For an observer on the earth, it is, with respect to the solar referential frame, of about $v=$ 30 km/s, corresponding to the earth motion around the sun.
+It is a consequence of the formula for the addition of velocities applied to a light beam when the observer is changing its reference frame.
+For an observer on the earth, it is, with respect to the solar referential frame, of about $v=$ 30 km/s, corresponding to the earth’s motion around the sun.
 It is sufficient to describe the effect of stellar aberration by working only up to the first order $v/c=10^{-4}$.
-According to the conventional approach, the study of stellar aberration is intimately connected with the old (Newtonian) kinematics: the Galilean vectorial law of addition of velocities is actually used.
+According to the conventional approach, the study of stellar aberration is intimately connected with the old (Newtonian) kinematics: the incorrect assumption that an earth-based observer and sun-based observer have a common 3-space is used.
 6.2 Wave Theory of the Stellar Aberration
-Most authors treat light propagating through the telescope barrel as rain of photons, and not as a plane wave (2) .
-Questioning the validity of standard reasoning we argue that a satisfactory treatment of stellar aberration should be based on the coherent wave optics.
+Most authors treat light propagating through the telescope barrel as rain of photons, and not as a wave  [^29] There is a widespread belief that the rays of light coming from the star are falling on the telescope tube and not interact with its sides.
+To quote French [4]: ”Regarding light as being composed of rain of photons, we can easily calculate the change in apparent direction of a distant object such a star.” Similar statements can also be found in recently published textbooks.
+To quote Rafelski [5]: ”We consider a light ray originating in a distant star.
+… The following discussion is addressing the observation of well-focusing light rays and not (spherical) plane wave light.
+… This picture of the experimental situation is accurate since the light emitted by a star consists of an incoherent flux of photons produced in independent atomic processes.” This conclusion is incorrect.
+We emphasize that the design of a telescope optical system is based on the classical theory of diffraction (originated directly from the classical electromagnetic theory).
+The resolution of a telescope is limited by the diffraction of light waves and the telescope always influences the measured radiation (due to unavoidable diffraction of a starlight by the telescope aperture).
+In a well-corrected optical system and a circular pupil, the size of the Airy disc of diffraction pattern (i.e.
+the size of the image of a point source) is inversely proportional to the pupil’s diameter.
+.
+Questioning the validity of standard reasoning we argue that a satisfactory treatment of stellar aberration should be based on coherent wave optics.
 It can be easily demonstrated that the light produced by a distant star is approximately coherent over a circular area whose diameter, in all practical cases, is much larger than the telescope diameter.
-Hence, we sample such a tiny portion of the coherent area of the starlight with our telescopes that the waveforms are effectively (flat) plane waves (3).
+Hence, we sample such a tiny portion of the coherent area of the starlight with our telescopes that the waveforms are effectively (flat) plane waves  [^30] A star is a completely incoherent source.
+The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34].
+Any star can be considered as very far away from the sun.
+In all cases of practical interest, telescopes are situated in the far zone of the (distant star) source.
+Consider a star like Sirius, which is one of the nearest stars.
+The coherent area of light observed from Sirius has a diameter of about 6 m.
+This correlation was observed by Brown and Twiss in 1956 [35]..
 Let us verify that this assertion is correct.
@@
 In its elementary sense, the degree of coherence between the two points simply describes the contrast of the interference fringes that are obtained when the two points are taken as secondary sources.
-Let a source $S$ illuminates the two pinholes $S_{1}$ and $S_{2}$, as shown in Fig.26.
+Let a source $S$ illuminate the two pinholes $S_{1}$ and $S_{2}$, as shown in Fig.24.
 The source is perfectly non-coherent.
@@
 It is sometimes said that the spatial coherence in light beams increases with distance ”by mere propagation”.
-It would be nice to find an explanation which is elementary in the sense that we can see what is happened physically.
+It would be nice to find an explanation that is elementary in the sense that we can see what is happening physically.
 Suppose that a quasi-monochromatic wave is an incident on an aperture in an opaque screen, as illustrated in Fig.
-26.
+24.
 In general, this wave may be partially coherent.
-The detailed structure of an optical wave undergoes changes as the wave propagates through space.
-In a similar fashion, the detailed structure of the spatial coherence undergoes changes, and in this sense, the transverse coherence function is said to propagate.
-Knowing the spatial coherence on the aperture, we wish to find the spatial coherence on the observing screen at distance $z$ beyond the aperture.
-The stellar radiation is a stochastic object and for any starlight beam there exist some characteristic linear dimension, $\Delta r$, which determines the scale of spatially random fluctuations.
+The detailed structure of an optical wave changes as the wave propagates through space.
+In a similar fashion, the detailed structure of the spatial coherence changes, and in this sense, the transverse coherence function is said to propagate.
+Knowing the spatial coherence on the aperture, we wish to find the spatial coherence on the observing screen at a distance $z$ beyond the aperture.
+The stellar radiation is a stochastic object and for any starlight beam, there exists some characteristic linear dimension, $\Delta r$, which determines the scale of spatially random fluctuations.
 Fig.25 illustrates the type of spiky pattern on an aperture in an opaque screen.
@@
 Here $\Delta r$ may be estimated as the typical linear dimension of spikes.
-[…]
 Figure 24: Young’s interferometer.
@@
 Figure 26: Reciprocal width relation of Fourier transform pair.
-Free space basically act as a Fourier transformation.
+Free space basically acts as a Fourier transformation.
 This means that the field in the far zone is, phase factor and proportionality factor aside, the spatial Fourier transform of the field at the source plane.
+[…]
 A star is an incoherent source with the scale of spatially random fluctuations $\lambdabar\Delta r\sim\lambdabar$.
@@
 In the case of astronomical observations by telescope, each spike on the earth’s surface has a finite thickness which is the same order as the spatial coherence length of the starlight.
-It can be said that the coherent area of the observed starlight would be of the order of $z/(kd)$, where $z$ is the distance between the star and earth.
-The star Bradley chose was Draconis, which is one of the nearest stars.
-In this case the diameter of the coherent area on the earth surface can be estimated to be about 100 m.
+It can be said that the coherent area of the observed starlight would be of the order of $z/(kd)$, where $z$ is the distance between the star and the earth.
+The star Bradley chose was Draconis, one of the nearest stars.
+In this case, the diameter of the coherent area on the earth’s surface can be estimated to be about 100 m.
 6.3 Examination of Stellar aberration in the Context of Special Relativity
-It is generally believed that the theory of relativity appears to conform to the phenomenon of stellar aberration discovered by Bradley by claiming it is a consequence of the motion of observers relative to light sources(4).
-The lack of symmetry between the cases when either the star or telescope on the earth is moving is shown clearly on the basis of the separation of binary stars.
-Spectroscopic binaries have velocities exceeding the earth’s velocity round the sun.
+It is generally believed that the theory of relativity appears to conform to the phenomenon of stellar aberration discovered by Bradley by claiming it is a consequence of the motion of observers relative to light sources 313131 In the paper on the theory of relativity, Einstein deduced the aberration formula from the idea that the velocity of $v$ is the relative velocity of the star-earth system.
+The idea was represented by many authors of textbooks.
+To quote Moeller [13]: ”This phenomenon, which is called aberration, was observed … by Bradley who noticed that the stars seem to perform a collective annual motion in the sky.
+This apparent motion is simply because the observed direction of a light ray coming from a star depends on the velocity of the earth relative to the star.”.
+The lack of symmetry between the cases when either the star or telescope on the earth is moving is shown clearly based on the separation of binary stars.
+Spectroscopic binaries have velocities exceeding the earth’s velocity around the sun.
 They revolve around their common center of gravity within days, a period during which the motion of the earth is practically constant.
-The components of the binary system should be easily separable, when their changing velocities are comparable to the earth’s velocity round the sun.
-This is, however, not observed (Fig.29-Fig.30) (5).
+The components of the binary system should be easily separable when their changing velocities are comparable to the earth’s velocity around the sun.
+This is, however, not observed (Fig.27-Fig.28)  [^32] In 1950 Ives [36] stressed for the first time that the presence of binaries in the sky gave rise to an important difficulty for the theory of relativity.
+It is stated that the idea that aberration may be described in terms of relative motions of the bodies concerned is immediately refuted by the existence of spectroscopic binaries with velocities comparable with that of the Earth in its orbit.
+Still, this exhibit’s aberrations are not different from other stars.
+For example, a spectroscopic binary, Mizar A, has well-known orbital parameters, from which can be calculated an observable angular separation of 1’10” if aberration were due to relative velocity.
+The empirical value is less than 0.01”, clearly incompatible with authors of textbooks point of view [37, 38]..
 Figure 27: Aberration shift as inferred from astronomical observations.
-The plane wave fronts of starlight entering the telescope are imaged by the lens to a diffraction spot which lies in the focal plane.
-Two cases are selected with different velocities of a earth-based telescope and a star.
+The plane wavefronts of starlight entering the telescope are imaged by the lens to a diffraction spot that lies in the focal plane.
+Two cases are selected with different velocities of an earth-based telescope and a star.
 In the first case (left) the star is at rest with respect to the fixed stars.
@@
 Figure 28: Aberration shift as predicted by the conventional theory of the stellar aberration.
-Two cases are selected with different velocities of a earth-based telescope and a star.
-In the first case (left) the star is at rest with respect to the fixed stars.
-In the second case (right) the star moves with the same velocity as the earth.
-The conventional theory predicts no aberration of the image in this case.
-In Chaper 4, we presented a critical reexamination of the textbook statement that wavefronts and raindrops are to have the same aberration.
+In Chapter 4, we presented a critical reexamination of the textbook statement that wavefronts and raindrops are to have the same aberration.
 We used the theory of relativity to show that when one has a transversely moving mirror and a plane wave of light is falling normally on the mirror, there is a deviation of the energy transport for light reflected from the mirror.
-This effect is a consequence of the fact that the Doppler effect is responsible for angular frequency dispersion of light waves reflected from the moving mirror with finite aperture.
+This effect is a consequence of the fact that in the absolute time coordinatization, the wave equation will depend on the velocity vector.
 As a result, the velocity of the energy transport is not equal to the phase velocity.
-According to the Babinet’s principle, this remarkable prediction of our theory should be correct also for light transmitted through a hole punched in a moving opaque screen or, consequently, through the moving open end of a telescope barrel.
-In principle the binary star paradox is resolved by noting that when light passes through the end of a telescope barrel we have a light beam whose fields have been perturbed by diffraction, and, therefore, do not include information about star motion relative to the fixed stars.
+According to Babinet’s principle, this remarkable prediction of our theory should be correct also for light transmitted through a hole punched in a moving opaque screen or, consequently, through the moving open end of a telescope barrel.
+In principle, the binary star paradox is resolved by noting that when light passes through the end of a telescope barrel we have a light beam whose fields have been perturbed by diffraction, and, therefore, do not include information about star motion relative to the fixed stars.
 If the telescope were at rest relative to the fixed stars and the star started to move from rest, then the apparent position of the star seen in the telescope would never jump by any angle.
-However, other difficulties arise in the explanation of the earth-based observations, i.e the change in an apparent positions of the fixed stars, which happens when the earth-based telescope changes of its motion relative to the fixed stars.
+However, other difficulties arise in the explanation of the earth-based observations, i.e.
+the change in apparent positions of the fixed stars, which happens when the earth-based telescope changes its motion relative to the fixed stars.
 It should be stressed that it is the telescope and not the star that must change its velocity (relative to the fixed stars) to cause aberration.
-An objective of this chapter is to consider non-relativistic interpretations of the stellar aberration, and clearly demonstrate that they are incorrect.
-Classical kinematics effects leads to serious mistakes if applied to the computation of stellar aberration as seen on the earth rotating around the sun (i.e.
-in a non-inertial geocentric frame of reference).
-The problem of the earth-based measurements is solved with the discovery of the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars (6).
+The problem of the earth-based measurements is solved with the discovery of the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars  [^33] This has been recognized by some experts, perhaps most explicitly by Selleri who states [38] that ”Thus a complete explanation of the aberration effect is given in terms of variations of the earth absolute velocity due to orbital motion, while the star/earth relative velocity is irrelevant.
+Thus acceleration (of the planet, this time) enters once more into a game.”.
 We derive the aberration for a pulse of light traveling on the surface of the earth using the Langevin metric.
-Though stellar aberration behaves asymmetrically, it does not contradict special relativity, because the heliocentric (i.e.
+Even though stellar aberration behaves asymmetrically, it does not contradict special relativity, because the heliocentric (i.e.
 sun-based) reference system is inertial and the geocentric reference system is non-inertial.
 6.4 Heliocentric Inertial Frame of Reference
-Above, we demonstrated that when one has some hole in the opaque screen at rest with respect to the fixed stars and the transversely moving point source, there is no aberration (deviation of the energy transport) for light transmitted through the hole.
+Above, we demonstrated that when one has some hole in the opaque screen at rest with respect to the fixed stars and the (tangentially to the screen) moving point source, there is no aberration (deviation of the energy transport) for light transmitted through the hole.
+The transmission appears as shown in Fig 12.
 The absence of the effects of moving (relative to the fixed stars) source in this setup automatically implies the same problem for stellar aberration theory in the heliocentric frame of reference.
@@
 Suppose that an observer, which is at rest relative to the telescope, performs the direction of the energy transport measurement.
-At close look at the physics of this subject shows that in the heliocentric frame of reference, where the telescope is at rest, we have actually the problem of steady-state transmission.
-Then how does the transmitted light beam looks?
+A close look at the physics of this subject shows that in the heliocentric frame of reference, where the telescope is at rest, we have the problem of steady-state transmission.
+Then how does the transmitted light beam look?
 It looks as though the transmitted beam is going along the telescope axis because it has lost its horizontal group velocity component.
-That is the transmission appears as shown in Fig 12.
 It takes the case of a telescope positioned perpendicular to the plane phase front.
 In other words, the telescope pointed directly at the star.
-If the motion of the star is parallel to the phase front (i.e.
+If star’s motion is parallel to the phase front (i.e.
 perpendicular to the telescope axis), starlight entering the end of the telescope would be able to pass its full length.
-One of the most important conclusions of the discussion presented above is that the aberration of starlight phenomenon is absent in this situation.
-In particular, the binary components remains unresolved which means that their velocity has no influence on aberration.
+One of the most important conclusions of the discussion presented above is that the aberration of the starlight phenomenon is absent in this situation.
+In particular, the binary components remain unresolved which means that their velocity does not influence aberration.
 6.5 Earth-Based Non-Inertial Frame of Reference
 The analogy between the obliquity of raindrops and the stellar aberration is incorrect.
-Only on the basis of the theory of relativity and the wave optics, we are able to describe all earth-based experimental observations of stellar aberration.
-Our theory predicts an effect of stellar aberration in complete agreement to the Bradley’s results, Fig.
-31.
+Only based on the theory of relativity and wave optics, we are able to describe all earth-based experimental observations of stellar aberration.
+Our theory predicts an effect of stellar aberration in complete agreement with Bradley’s results, Fig.
+29.
 According to the asymmetry between the inertial and rotating frames, there is a remarkable prediction on the theory of the aberration of light.
-Namely, if the telescope is at rest relative to the earth and the earth rotating relative to the fixed stars, then the direction of a star as seen from the earth is not the same as the direction when viewed by a hypothetical sun-based observer.
-Apparent angle is less than the actual angle.
+Namely, if the telescope is at rest relative to the earth and the earth is rotating relative to the fixed stars, then the direction of a star as seen from the earth is not the same as the direction when viewed by a hypothetical sun-based observer.
+The apparent angle is less than the actual angle.
 The difference between the actual angle and apparent angle $\theta_{a}$ is connected with the physical parameters by the relation: $\theta_{a}=v/c$, where $v$ is the velocity of the earth in its orbit around the sun.
 It could be said that the crossed term in metric Eq.(15) generates anisotropy in the rotating frame that is responsible for the change of radiation direction (aberration).
-That is the transmission through the telescope aperture in the rotating frame appears as shown in Fig.
-20.
-Figure 31: The direction of a star as seen from the earth is not the same as the direction when viewed by a hypothetical observer at the sun center.
-Apparent angle $\phi$ is less than the actual angle $\theta$.
-The difference between the actual angle and apparent angle is connected with the physical parameters by the relation: $\theta-\phi=v/c$, where $v$ is the velocity of the earth in its orbit around the sun.
-The stellar aberration in the geocentric frame of reference is considered independent of the star speed and to have just a local origin exclusively based on the observer speed with respect to the sun.
-The main facts which a theory of stellar aberration in the earth-based frame of reference must explain are (1) the annual apparent motion of the fixed stars about their locations and (2) the null apparent aberration of rotating binary systems.
-We presented here a theory which accounts for all these, and in addition gives new results.
-We demonstrated that the aberration of light is a complex phenomenon which must be branched out into a number of varieties according to their origin.
-These branching out takes place depending on what is the cause of aberration - whether it is a jump in the velocity (relative to the fixed stars) of the observer or of the light source.
-Aberration could undergo further splitting - depending on the physical influence of the optical instrument on the measurement.
+The transmission through the telescope aperture in the rotating frame appears as shown in Fig.18 - Fig.
+19.
+Figure 29: The direction of a star as seen from the earth.
+The stellar aberration in the geocentric frame of reference is considered independent of the star’s speed and to have just a local origin exclusively based on the observer’s speed with respect to the sun.
 According to our interpretation, there are many kinds of aberration and the stellar aberration in the earth-based frame is only one of these.
-We demonstrated that all earth-based experiments can be explained on the basis of the effect of the measuring instrument (i.e.
+We demonstrated that all earth-based experiments can be explained based on the effect of the measuring instrument (i.e.
 physical influence of the telescope on the measurement) and the acceleration of traveling earth-based observer relative to the fixed stars.
 6.6 Physical Coordinate System in Space
-The aberration of light is the geometric phenomenon.
-In order to detect the aberration effect inside the earth-based frame, it is obvious that some coordinate system with reference direction is needed.
-A conventional approach to the aberration of light effect is forcefully based on a definite assumption of reference direction, but this is actually a hidden assumption.
+The aberration of light is a geometric phenomenon.
+To detect the aberration effect inside the earth-based frame, it is obvious that some coordinate system with reference direction is needed.
+A conventional approach to the aberration of light effect is forcefully based on a definite assumption of reference direction, but this is a hidden assumption.
 Traditionally the physical interpretation of the aberration of light effect in terms of measurements performed with rods that are at rest in the observer’s frame.
@@
 Fortunately, it is possible to find the space description on a more fundamental level than that of measuring rods.
-The reference axis in the earth-based frame is formed by gravitation field vector.
+The reference axis in the earth-based frame is formed by a gravitation field vector.
 The plumb-line direction is known as the nadir, leading to the earth’s center.
@@
 The region of applicability of our stellar aberration theory is more wider than one might think.
-Above we assumed for simplicity that a star image is actually a point spread function in the image plane of a telescope.
+Above we assumed for simplicity that a star image is a point spread function in the image plane of a telescope.
 In other words, it is assumed that the input signal is effectively a plane wave.
 Suppose that a telescope is able to distinguish detail in the star image.
-A complete understanding of the relation between object and image can be obtained if the effect of diffraction are included.
+A complete understanding of the relation between object and image can be obtained if the effect of diffraction is included.
 The effect of diffraction is to convolve that ideal image with the Fraunhofer diffraction pattern of the telescope pupil.
-The star is a spatially completely incoherent source (7).
-This means that such source is actually a system of elementary (statistically independent) point sources with different offsets.
+The star is a spatially completely incoherent source  [^34] In the statistical optics for star emission, the fields are described classically at the level of Maxwell’s equations and the emitting medium is treated (as an ensemble of atoms) by quantum mechanics.
+The star is a system of elementary point sources.
+The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$, where $\lambda$ is the (visible) radiation wavelength.
+Within an elementary source volume ($\lambda^{3}$) there is an enormous number of atoms.
+The semi-classical theory is treating these atoms as coherent, interacting radiating dipoles.
+The induced macroscopic dipole moment in an elementary source leads to classical electromagnetic radiation.
+The distinguishing characteristic of statistical optics is the fact that electromagnetic fields are treated in a completely classical manner until they interact with the atoms of the photosensitive material on which they are incident.
+Thus there is no necessity to deal with quantization of the electromagnetic field; only the interaction of the classical field and matter is quantized.
+The photodetector has the effect of converting the continuous cycle-averaged classical intensity of the light beam into a succession of discrete photocounts..
+This means that such a source is a system of elementary (statistically independent) point sources with different offsets.
 An elementary source with a given offset produces in front of a telescope pupil effectively a plane wave.
 An elementary source offset tilts the far zone field.
-Radiation field generated by an completely incoherent source can be seen as a linear superposition of fields of individual elementary point sources.
+A radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individual elementary point sources.
 The image of an elementary point source is a point spread function.
 In other words, there is always the physical influence of the telescope on the measurement of a completely incoherent source.
-It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single point source characteristics like independence on source motion.
-This argument gives reason why our theory of stellar aberration is correct also for imaging of arbitrary completely incoherent sources.
-6.8 Bibliography and Notes
-1.
-The phenomenon of the annual apparent motion of celestial objects about their locations, named stellar aberration, was discovered by Bradley in 1727, who also explained it employing the corpuscular model of light [33].
-2.
-There is a widespread belief that the rays of light coming from the star are falling on the telescope tube and not interact with its sides.
-To quote French [4]: ”Regarding light as being composed of rain of photons, we can easily calculate the change in apparent direction of a distant object such a star.” Similar statements can also be found in recently published textbooks.
-To quote Rafelski [5]: ”We consider a light ray originating in a distant star.
-… The following discussion is addressing the observation of well focusing light-rays and not (spherical) plane wave light.
-… This picture of the experimental situation is accurate since the light emitted by a star consists of an incoherent flux of photons produced in independent atomic processes.” This conclusion is incorrect.
-We emphasize that the design of a telescope optical system is based on the classical theory of diffraction (originated directly from the classical electromagnetic theory).
-The resolution of a telescope is limited by the diffraction of light waves and the telescope always influences the measured radiation (due to unavoidable diffraction of a star light by the telescope aperture).
-In a well corrected optical system and in a circular pupil, the size Airy disc of diffraction pattern (i.e.
-size of the image of a point source) is inversely proportional to the diameter of the pupil.
-3.
-A star is a completely incoherent source.
-The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34].
-Any star can be considered as very far away from the sun.
-In all cases of practical interest, telescopes are situated in the far-zone of the (distant star) source.
-Consider a star like Sirius, which is one of the the nearest stars.
-The coherent area of light observed from Sirius has a diameter of about 6 m.
-This correlation was observed by Brown and Twiss in 1956 [35].
-4.
-It is widely believed that stellar aberration depends on the relative velocity of the source (star) and observer.
-In the paper on the theory of relativity Einstein deduced the aberration formula from the idea that the velocity of $v$ is the relative velocity of the star-earth system.
-The idea was represented by many authors of textbooks.
-To quote Moeller [13]: ”This phenomenon, which is called aberration, was observed … by Bradley who noticed that the stars seem to perform a collective annual motion in the sky.
-This apparent motion is simply due to the fact that the observed direction of a light ray coming from a star depends on the velocity of the earth relative to the star.”
-5.
-In 1950 Ives [36] stressed for the first time that the presence of binaries in the sky gave rise to an important difficulty for the theory of relativity.
-It is stated that the idea that aberration may be described in terms of relative motions of the bodies concerned is immediately refuted by the existence of spectroscopic binaries with velocities comparable with that of the Earth in its orbit.
-Still this exhibit aberrations not different from other stars.
-For example, a spectroscopic binary, Mizar A, has well-known orbital parameters, from which can be calculated an observable angular separation of 1’10” if aberration were due to relative velocity.
-The empirical value is less than 0.01”, clearly incompatible with authors of textbooks point of view [37, 38].
-There is no available explanation for the fact that, while the observational data on stellar aberration are compatible with moving earth, the symmetric description, when the star possesses the relative transverse motion, does not apparently lead to observations compatible with predictions.
-6.
-Stellar aberration exist as observable phenomenon only in the presence of changing states of motion (i.e.
-acceleration).
-The problem of the earth-based measurements is related with the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars.
-This has been recognized by some expects, perhaps most explicitly by Selleri who states [38] that ”Thus a complete explanation of the aberration effect is given in terms of variations of the earth absolute velocity due to orbital motion, while the star/earth relative velocity is irrelevant.
-Thus acceleration (of planet, this time) enters once more into a game.”
-7.
-From everything that authors of textbooks have written about the star light (”rain of photons”) it clearly follows that they did not understand the several aspects of the statistical optics.
-In the statistical optics for star emission, the fields are described classically at the level of Maxwell’s equations and the emitting medium is treated (as a ensemble of atoms) by quantum mechanics.
-The atomic polarization induced by the radiation field appears as a driving term in Maxwell’s equations and sustains the oscillations.
-The star is actually a system of elementary (statistically independent) point sources with different offsets.
-The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$, where $\lambda$ is the (visible) radiation wavelength.
-Within an elementary source volume ($\lambda^{3}$) there is an enormous number of atoms.
-Semi-classical theory is treating these atoms as coherent, interacting radiating dipoles.
-The induced macroscopic dipole moment in an elementary source leads to the classical electromagnetic radiation.
-The distinguishing characteristic of the statistical optics is the fact that electromagnetic fields are treated in a completely classical manner until they interact with the atoms of the photosensitive material on which they are incident.
-Thus there is no necessity to deal with quantization of the electromagnetic field; only the interaction of the classical field and matter is quantized.
-Thus in the semiclassical theory, where the atom is treated quantum mechanically but the electromagnetic field is treated classically, the photodetector has the effect of converting the continuous cycle-averaged classical intensity of the light beam into a succession of discrete photocounts.
-Light interact with matter in a fundamentally random way.
-As a consequence, any measurement of (star) light will be accompanied by certain unavoidable fluctuations in the detector.
-Stochastic fluctuations of the classical intensity can influence the statistical properties of the photoevents that are observed.
-Note in particular that the variance $\sigma^{2}_{K}$ of photoevents $K$ consists of two distinct terms, each of which has a physical interpretation: $\sigma^{2}_{K}=[K^{2}-<K>^{2}]/<K>^{2}=1/<K>+\sigma^{2}_{W}$.
-The first term is simply the variation of the counts that would be observed if classical intensities were constant and the photocounts were purely Poisson.
-We refer to this contribution to the count fluctuations as ”quantum noise”.
-The second therm, $\sigma^{2}_{W}$, is clearly zero if there are no fluctuations of the classical intensity.
-For instance, in the case of laser light, this component would be identically zero, and the count variance would be simply that arising from the photon ”shot noise”.
-This is intrinsic fluctuations associated with quantum fluctuations of vacuum and can give information only on the mean rate ot photons production.
-When star (thermal) light is incident on the photodetector, the classical fluctuations are nonzero, and the variance of the photocounts is larger than that expected for a Poisson distribution.
-The classical thermal fluctuations contains rich information about radiation source.
-In particular, thermal noise provide us with information for analyzing (using Van Cittert-Zernike theorem) the spatial source profile.
-The ratio of the classical variance to the ”photon shot noise” variance is equal to $\delta_{c}=<K>\sigma^{2}_{W}$.
-Parameter $\delta_{c}$ named as the photocount degeneracy parameter and can be interpreted as the average number of counts that occur in a single coherence interval of the incident radiation.
-When $\delta_{c}\ll 1$, it is highly probable that will be no more than one count per coherence interval of the wave, with the result that shot noise predominates over classically induced noise.
-On the other hand, when $\delta_{c}\gg 1$, there are many photoevents present in each coherence interval of the wave.
-The result is ”bunching” of the photoevents by the classical intensity fluctuations and an increase of the variance of the counts to the point where the classically induced fluctuations are far stronger than the shot noise variations.
+It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single-point source characteristics like independence on source motion.
+This argument gives the reason why our theory of stellar aberration is correct also for imaging of arbitrary completely incoherent sources.
 7 Notion of Ordinary Space in Special Relativity
-We introduce new approach to the aberration of light theory in non-inertial frames of reference, finding another way in which this complicated problem can be solved.
-For example, when viewed from the inertial frame, the aberration of light effect in the accelerated frame (Fig.
-16) is easy to calculate in the framework of the special theory of relativity taking advantage of the relativity of simultaneity.
-In contrast, in order to compute the aberration of light effect in non-inertial systems we used a metric tensor.
 7.1 Inertial Frame View of Observations of the Noninertial Observer
-First we want to rise the following interesting and important point.
+We introduce a new approach to the aberration of light theory in non-inertial frames of reference, finding another way in which this complicated problem can be solved.
+First, we want to raise the following interesting and important points.
 The laws of physics in any inertial reference frame should be able to account for all physical phenomena, including the observations made by non-inertial observers.
 It is widely believed that all phenomena in non-inertial (e.g.
-rotating) reference system should be considered only in the framework of the space-time geometric approach using Langevin metric.
-However, the analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard (Einstein) special relativity taking advantage of well known relativistic kinematic effects.
-For example, when viewed from the inertial lab system, the interpretation of the Sagnac effect is simple and the phase difference (attributed to the Sagnac effect) between counter-propagating waves may be derived from the relativistic law of velocity composition.
-In other words, Sagnac effect in the rotating frame, as viewed from the inertial lab frame, presents a kinematic effect (Einstein’s velocity addition) of special theory of relativity.
-It is implicitly assumed that the speed of light is $c$ in the lab frame which, in accordance with Einstein special relativity, is independent of motion of the source.
+rotating) reference systems should be considered only in the framework of the space-time geometric approach using the Langevin metric.
+However, the analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard (Einstein) special relativity taking advantage of well-known relativistic kinematic effects.
+For example, when viewed from the inertial lab system, the interpretation of the Sagnac effect is simple, and the phase difference (attributed to the Sagnac effect) between counter-propagating waves may be derived from the relativistic law of velocity composition.
+In other words, the Sagnac effect in the rotating frame, as viewed from the inertial lab frame, presents a kinematic effect (Einstein’s velocity addition) of the special theory of relativity.
+It is implicitly assumed that the speed of light is $c$ in the lab frame which, in accordance with Einstein’s special relativity, is independent of the motion of the source.
 In contrast, the Sagnac effect is not easy to calculate in a frame of reference attending rotation.
-In this case, authors of textbook used a metric tensor (Langevin metric) in a plane four-dimensional Minkowski space-time to calculate the propagation time difference between counter-running waves.
+In this case, authors of textbooks used a metric tensor (Langevin metric) in a plane four-dimensional Minkowski space-time to calculate the propagation time difference between counter-running waves.
 Can we not look at the aberration of light effect in the same way?
-Here also we used the Langevin metric in a non-inertial frame of reference.
+Here we also used the Langevin metric in a non-inertial frame of reference.
 Let us first consider the application of the classical kinematic method to the computation of the aberration of light effect in a non-inertial frame, as viewed from an inertial frame.
-Authors of textbooks obtain the same addition theorem for light beam velocities as for particle velocities in the classical physics.
+Authors of textbooks obtain the same addition theorem for light beam velocities as for particle velocities in classical physics.
 This result is particularly remarkable because the study of light aberration is intimately connected with old kinematics.
-If we neglect $v^{2}/c^{2}$ comparison with unity, the Galilean vectorial law of addition of velocities is actually used.
+If we neglect $v^{2}/c^{2}$ comparison with unity, the Galilean vectorial law of addition of velocities is used.
 However, the classical kinematics method leads to a serious mistake if applied to the computation of the aberration of light effect.
-Let us consider the aberration of light in the inertial frame.
-According to our approach, there is a remarkable prediction concerning the deviation of the energy transport for light transmitted through a hole on a moving screen, Fig.
-9.
-In order to predict the result of the aberration measurement the accelerated observer should use the Langevin metric.
-The crossed term in this metric generates anisotropy in the accelerated frame that is responsible for the change of radiation direction, Fig.
-14.
-In contrast, based only on the (Galilean or Einstein’s) velocity addition, we arrive at the conclusion that there is no aberration of the transmitted light beam in the accelerated frame.
-The error in the last argument follows from the fact that incorrect assumption that an inertial observer and accelerated observer have common 3-space is used.
+According to our approach, there is a remarkable prediction concerning the aberration of light phenomena in the accelerated frame, Fig.
+20.
+To predict the result of the aberration measurement the accelerated observer should use the Langevin metric.
+The crossed term in this metric generates anisotropy in the accelerated frame that is responsible for the change of radiation direction.
+In contrast, based only on the (Galilean or Einstein’s) velocity addition, we conclude that acceleration does not spoil the symmetry between the accelerated reference frame and the initial inertial frame.
+The error in the last argument follows from the fact that the incorrect assumption that an inertial observer and accelerated observer have a common 3-space is used.
 According to the special relativity, inertial observer and accelerated (with respect to the fixed stars) observer have different 3-spaces.
@@
 Relativistic effects do not have a place in this description.
-According to classical approach, there is no principal difference between the aberration of light and the aberration of raindrops.
-We first notice that one of the postulate of special relativity states that if two distinct events cannot be connected by a causal signal that travels no faster than a light signal, then they cannot be connected by a causal signal at all.
-We will show that the aberration of light effect is a corollary to the relativistic kinematic effects.
-The appearance of relativistic effects in radiation phenomena does not depends on a large speed of the radiation sources.
-Light is always a relativistic object.
-In particular the relativity of simultaneity is responsible for aberrations to the first order in $v/c$.
+According to the classical approach, there is no principal difference between the aberration of light and the aberration of raindrops.
 Now, it is very interesting to show that the geometric effects in our ordinary space world are closely associated with the relativity of simultaneity.
@@
 16 the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light.
-The question cannot be avoided relative to what a light beam propagated in the accelerated frame with angular displacement $-v/c$?
-Suppose that an observer in the accelerated frame performs the direction of the light beam measurement and the plane wavefront of transmitted light beam is imaged by a lens to a diffraction spot which lies in the focal plane on the optical axis.
-Measurement of the direction of the optical axis with respect to the frame axes is equivalent to the determination the angular displacement.
-In order to detect the aberration of light effect inside the accelerated frame, it is obvious that some coordinate system with reference direction is needed.
+The question cannot be avoided relative to what a light beam propagated in the accelerated frame with angular displacement $-v/c$.
+Suppose that an observer in the accelerated frame performs the direction of the light beam measurement and the plane wavefront of the transmitted light beam is imaged by a lens to a diffraction spot that lies in the focal plane on the optical axis.
+Measurement of the direction of the optical axis with respect to the frame axes is equivalent to the determination of the angular displacement.
+To detect the aberration of light effect inside the accelerated frame, it is obvious that some coordinate system with reference direction is needed.
 We must inquire in detail by what method we assign coordinates.
-This method involves some sort of physical procedure; eventually it must be such that it will be give us coordinates in both (inertial and accelerated) frame of reference.
-Figure 32: Aberration of light in the accelerated frame of reference.
-In the Einstein time synchronization for the screen, the accelerated observer describes the transmitted light beam and the reference beam from an emitter at rest using usual Maxwell’s equations.
-The $x_{n}$-axis is parallel to the wavefront of the reference beam.
-The aberration increment may be defined as an angle between the reference and transmitted light beams.
-Figure 33: Aberration of light in the inertial frame of reference.
-In the absolute time synchronization for the screen, the inertial observer describes the reference beam from an emitter at rest using usual Maxwell’s equations.
-The $x$-axis is parallel to the wavefront of the reference beam.
-The aberration increment may be defined as an angle between the reference and transmitted light beams.
-In ordinary space we find that the accelerated frame moves with respect to the inertial frame along the line motion and the inertial frame moves with respect to the accelerated frame along the line motion.
+This method involves some sort of physical procedure; eventually, it must be such that it will give us coordinates in both (inertial and accelerated) frames of reference.
+In ordinary space, we find that the accelerated frame moves with respect to the inertial frame along the line motion and the inertial frame moves with respect to the accelerated frame along the line motion.
 The angle between the axis of the observer’s coordinate system and the line motion is a simple ordinary space geometric parameter.
 Using the line motion as a reference $x_{n}$-axis, the accelerated observer can then define the second reference axis.
-We need to give a ”practical”, ”operational” answer to the question of how to assign an axis perpendicular to the $x_{n}$-axis.
-Clearly, it is possible to define a reference direction using a light beam.
+We need to give a ”practical”, and ”operational” answer to the question of how to assign an axis perpendicular to the $x_{n}$-axis.
+It is possible to define a reference direction using a light beam.
 We will define the second reference direction in the following way.
-Let us suppose that the aberration direction inside the accelerated frame is determined with reference to the fixed direction of light beam from a ”plane-wave” emitter which is at rest in the accelerated frame.
+Let us suppose that the aberration direction inside the accelerated frame is determined concerning the fixed direction of the light beam from a ”plane-wave” emitter which is at rest in the accelerated frame.
 In other words, a coordinate system is formed here by electromagnetic axis and line motion.
-The motion of the aberrated light beam are assumed for simplicity, to lie in the same plane and the angular position of the aberrated beam is described by one angle.
+The motion of the aberrated light beam is assumed for simplicity, to lie in the same plane and the angular position of the aberrated beam is described by one angle.
 When the accelerated system starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed.
-The Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all direction.
+The Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions.
 In this synchronization, the accelerated observer describes the light beam from an emitter at rest using usual Maxwell’s equations.
@@
 We consider the case in which the velocity component of the (reference) light beam along the $x_{n}$-axis is equal to zero.
-In other words, the $x_{n}$-axis is parallel to the wavefront of the reference ”plane wave”, Fig.
-32.
+In other words, the $x_{n}$-axis is parallel to the wavefront of the reference ”plane wave”.
 The number that specifies the aberration increment of the light beam transmitted through the hole may be defined as an angle between the reference and transmitted light beams.
 Figure 30: Inertial frame view of observations of the accelerated observer.
-In the Lorentz coordinatization, the reference light beam in the accelerated frame is emitted in the direction normal to the wavefront.
-If make a Lorentz boost, we introduce a time transformation $t\to t+xv/c^{2}$ and the effect of this transformation is just an inclination of the electromagnetic axis in the accelerated frame.
 Let us try to get an understanding of the relationship between the reference directions inside the inertial frame and the reference directions inside the accelerated frame.
-Such approach in our case of (aberration of light) interest is based on the use of local (reference) light sources.
+Such an approach in our case of (aberration of light) interest is based on the use of local (reference) light sources.
 We determine the reference directions perpendicular to the line motion in the inertial frame and the accelerated frame using two local light sources.
-In effect, one source should be stay at rest in the accelerated frame (Fig.
-32) while the second should be at rest in the inertial frame (Fig.
-33).
+In effect, one source should stay at rest in the accelerated frame while the second should be at rest in the inertial frame.
 In other words, the reference electromagnetic axis in each frame is formed by an individual light beam.
@@
 The equivalence of all local physical frames of reference underlies the theory of relativity, so when the aberration angle emerges, it emerges in the same manner in all local physical reference systems.
-A comparison with a light clock might help here.
-When inertial observer looks at the light clock inside the accelerated frame, he sees that clock run slowly.
-Not only does this particular kind of clock run more slowly, but if the theory of relativity is correct, any other clock, operating on any principle whatsoever, would also appear to run slower.
 Now let us return to observations of an accelerated observer, as viewed from an inertial frame.
-Let us start with non-covariant approach.
+Let us start with the non-covariant approach.
 The aberration of light problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations.
@@
 The explanation consists in using Lorentz coordinatization to describe the uniform translation motion of the light source in the inertial frame.
-In the Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization.
+In Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization.
 We can interpret manipulations with rule clock structure in the inertial frame simply as a change of the time variable according to the transformation $t\to t+xv/c^{2}$.
-The overall combination of Galileo transformation and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the inertial frame.
-Now speed of light is independent of motion of the source in the inertial frame.
-This means that the velocity of light relative the moving system is $c-v$ in the positive $x$-direction, and $-c-v$ in the negative direction.
-It is immediately understood that, as a result of motion of accelerated system along the $x$-axis (i.e.
-along the plane of simultaneity of the reference beam in the accelerated frame) with velocity $v$, there is relativity of simultaneity.
+The overall combination of Galileo transformation and variable changes yields the Lorentz transformation in the case of absolute time coordinatization in the inertial frame.
+Now speed of light is independent of the motion of the source in the inertial frame.
+This means that the velocity of light relative to the moving system is $c-v$ in the positive $x$-direction, and $-c-v$ in the negative direction.
+It is immediately understood that, as a result of the motion of the accelerated system along the $x$-axis (i.e.
+along the plane of simultaneity of the reference beam in the accelerated frame) with velocity $v$, there is the relativity of simultaneity.
 This reasoning is in analogy with Einstein’s train-embankment experiment.
-On the one hand, the wave equation remains invariant with respect to Lorentz transformations.On the other hand, if use the Lorentz coordinatization , we automatically introduce a time transformation $t\to t+xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation phase front in the accelerated frame on the angle $v/c$.
+On the one hand, the wave equation remains invariant with respect to Lorentz transformations.
+On the other hand, if use the Lorentz coordinatization, we automatically introduce a time transformation $t\to t+xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation phase front in the accelerated frame on the angle $v/c$.
 From what has preceded it is easy to see that in the accelerated system, from the point of view of the inertial observer, the electromagnetic axis, assigned by the accelerated observer, will not parallel to the $z$-axis of the inertial observer, Fig.
-34.
-We arrive at the conclusion that in the accelerated frame transmitted light beam propagates along the $z$-axis of the inertial frame of reference with angular displacement $-v/c$ with respect to the reference beam (i.e.
+30.
+We conclude that in the accelerated frame transmitted light beam propagates along the $z$-axis of the inertial frame of reference with angular displacement $-v/c$ with respect to the reference beam (i.e.
 with respect to the $z_{n}$-axis) in the accelerated frame.
-To continue our discussion of inertial frame view of observations of the non-inertial observer, let us consider the opposite situation.
 We determine the reference directions in the inertial frame and the accelerated frame using two local light sources.
 This has interesting consequences.
-For example, suppose that an observer in the inertial frame performs the direction of transmitted light measurement relative to the reference light beam from the emitter which is at rest with respect to the inertial frame (Fig.
-33).
+First of all, the accelerated observer and the inertial observer can directly measure the angular displacement of the transmitted light beam with respect to the reference light beam of the inertial frame.
+Suppose that an observer in the inertial frame performs the direction of transmitted light measurement relative to the reference light beam from the emitter which is at rest with respect to the inertial frame.
 The inertial observer would find that the angular displacement is positive and equal to $\theta_{a}=v/c$.
 Now we ask about the angular displacement of transmitted light with respect to the reference beam inside the accelerated frame.
-On the one hand, we know that angular displacement of the transmitted light with respect to the reference light beam from the emitter which is at rest with respect to the accelerated frame is $-v/c$.
+On the one hand, we know that the angular displacement of transmitted light with respect to the reference light beam from the emitter which is at rest with respect to the accelerated frame is $-v/c$.
 On the other hand, the angular displacement of the inertial reference light beam is $-2v/c$, Fig.
 17.
-We come to the conclusion that the accelerated observer can directly measure the angular displacement of the transmitted light with respect to the reference light beam of the inertial frame.
+We conclude that the accelerated observer can directly measure the angular displacement of the transmitted light with respect to the reference light beam of the inertial frame.
 In fact, the relative angular displacement is $\theta_{a}(\mathrm{transmitted~{}beam})-\theta_{a}(\mathrm{inertial~{}reference~{}beam})=-v/c+2v/c=v/c$.
-This result is consistent with the inertial observer measurements (Fig.
-33), as must be.
+This result is consistent with the inertial observer measurements, as must be 353535 We have been talking about the electromagnetic axis.
+Suppose that we determine the reference directions in the inertial frame and the accelerated frame using two relativistic electron sources.
+We emphasize an important difference between a reference light beam and a reference relativistic electron beam.
+In our experimental arrangement, it would be impossible to measure the angular displacement of transmitted light exactly using the reference electron beam axis..
 7.2 The Non-Existence of an Instantaneous 3-Space
@@
 Let us assume, according to textbooks, that the line motion of observers is the same in the accelerated frame as in the inertial frame.
-At first site, if reference axes were orthogonal to the line motion they must be parallel to each other.
-Since there exist the angular displacement $v/c$, the situation seems paradoxical.
-Figure 35: The lab observer sees the rotation of the moving axes with respect to the lab frame axes, which is closely associated with the Wigner rotation.
-First of all, as noted above, the key idea in resolution of this paradox is the realization of the fact that there is no absolute notion of simultaneity in the theory of special relativity.
-In the case of the relativity of simultaneity we have a mixture - of positions and time.
-In other words, in the space measurement of one observer there is mixed a little bit of the time as seen by the other.
-The point, then, is not the existence of a common ordinary space.
-The fact that two observers with different trajectories have different 3-spaces is not considered in textbooks.
-If we look at the situation carefully we see that the accepted in textbooks assumption of common (x) line motion of observers is based on the belief that observers have common 3-space.
-This is misconception.
-Right way would be to use the light beam as the common reference direction.
+At first sight, if reference axes were orthogonal to the line motion they must be parallel to each other.
+Since there exists the angular displacement $v/c$, the situation seems paradoxical.
+Figure 31: The inertial observer sees the rotation of the moving axes with respect to the lab frame axes, which is closely associated with the relativity of simultaneity.
+First of all, as noted above, the key idea in the resolution of this paradox is the realization of the fact that there is no absolute notion of simultaneity in the theory of special relativity.
+In the case of the relativity of simultaneity, we have a mixture - of positions and time.
+In other words, in the space measurement of one observer, there is mixed a little bit of the time as seen by the other.
+If we look at the situation carefully we see that the generally accepted assumption of common (x) line motion of observers is based on the belief that observers have a common 3-space.
+This is a misconception.
+The right way would be to use the light beam as a reference direction.
 We have already seen that the inertial observer can directly measure the angular displacement between the inertial ($z$) and accelerated ($z_{n}$) electromagnetic reference directions and this result is consistent with the accelerated observer measurement.
@@
 Let us accept that the electromagnetic $z_{n}$ axis in the inertial frame is the reference light beam in the accelerated frame.
-Now the lab observer sees the rotation of the moving frame axes with respect to the lab frame axes, Fig.
-35.
+Now the inertial observer sees the rotation of the moving frame axes with respect to the inertial frame axes, Fig.
+31.
 This rotation is a relativistic kinematic effect.
-One can verify directly that the axes of accelerated frame are actually rotated with respect to the lab frame axes by the angle equal to $\theta_{w}=v/c$ in the first order in $v/c$.
-Let us clear up the meaning of the parameter $\theta_{w}$.
-It can be associated with the angle of the so-called Wigner rotation (see the Section 7.5 for more details).
-Our results show that we cannot remain with the framework of parallel axes $z_{n},z$ when considering the inertial frame view of light beam observations of the non-inertial observer in perpendicular geometry.
-This inclination angle $v/c$ can be interpreted as a rotation angle $\theta_{w}=v/c$ of the moving axes with respect to the lab frame axes.
+One can verify directly that the axes of the accelerated frame are rotated with respect to the inertial frame axes by the angle equal to $\theta_{w}=v/c$ in the first order in $v/c$ (see the next section for more details).
 7.3 Acceleration of a Rigid Body and Special Relativity
@@
 The concept of rigid motion, as known from Newtonian mechanics, cannot be taken over unchanged to the special theory of relativity.
-This follows from the fact that in the space measurement of one observer there is mixing of space and time, as seen by other.
-First, let us study the relativistic kinematics of rotation of the moving frame axes with respect to the inertial frame.
-Consider an emitter-detector setup in the lab inertial frame, Fig.
-36.
-Suppose that the setup is accelerated from the rest with respect to the lab frame up to velocity $v$ along the $x$-axis.
-The main difference between the relativistic kinematics and the kinematics of Newton is that in the lab frame, the accelerated emitter-detector setup exhibits a shearing motion.
+First, let us study the relativistic kinematics of rotation of the moving frame axes with respect to the inertial frame axes.
+Consider two emitter-detector setups in the initial inertial frame.
+Suppose one setup is accelerated from the rest with respect to the inertial frame up to velocity $v$ along the $x$-axis, Fig.
+32.
+The explanation of the effect of aberration is based on the use of absolute time coordinatization  [^36] Another explanation consists of using a clock re-synchronization procedure in the initial inertial frame.
+The time under Einstein’s synchronization for the moving source is readily obtained by introducing the offset factor $xv/c^{2}$ in the first-order approximation.
+This time shift has the effect of rotation of the radiation wavefront on the angle $v/c$ for both sources.
+In this synchronization, we describe the radiation of the moving source using Maxwell’s equations.
+It is interesting to discuss the electrodynamics of the source at rest.
+The new crossed term that has to be put into the wave equation due to time transformation leads to the prediction of the group velocity $d\omega/dk_{x}=-v$ (see footnote in Section 3.5).
+On the other hand, there is the angle $v/c$ between the wave number vector $\vec{k}$ and the $z$-axis.
+Then, accounting for the initial condition, the radiated light beam of the source at rest moves along the $x$-axis with group velocity $d\omega/dk_{x}=0$ as it must be..
+The main difference between the relativistic kinematics and the kinematics of Newton is that in the inertial frame, the accelerated emitter-detector setup exhibits a shearing motion.
 This counterintuitive fact is not a paradox.
-In fact, this shearing contribution adds up to the Wigner rotation.
 The shearing (3-space geometric) effect is closely associated with the relativity of simultaneity.
-Above we demonstrated that aberration of light effect in the accelerated frame, as viewed from the inertial lab frame, may be derived from the Lorentz transformations.
-In other words, aberration of light effect in the accelerated frame, as viewed from the inertial lab frame, presents a kinematics effect of special theory of relativity, Fig.
-37.
+Above we demonstrated that the aberration of light effect in the accelerated frame, as viewed from the inertial frame, may be derived from the Lorentz transformations.
+In other words, aberration of light effect in the accelerated frame, as viewed from the initial inertial frame, presents a kinematics effect of the special theory of relativity, Fig.
+33.
 It is shown that the relativity of simultaneity is responsible for aberrations in the first order of $v/c$.
-We discovered of the essential asymmetry between the inertial and accelerated frame, namely, an acceleration (with respect to the inertial frame) have an effect on the propagation of light in the accelerated frame.
-On accelerated system, the accelerated detector does not receive light radiated by the accelerated emitter, Fig.
-37.
-In contrast, the inertial detector at rest continues to receive light radiated by the inertial emitter at rest, Fig.
+We discovered the essential asymmetry between the inertial and accelerated frame, namely, an acceleration (with respect to the inertial frame) has an effect on the propagation of light in the accelerated frame.
+In an accelerated system, the accelerated detector does not receive light radiated by the accelerated emitter, Fig.
+33.
+In contrast, the detector at rest in the inertial frame continues to receive light, Fig.
 32 left.
 Figure 32: Aberration of light in an inertial frame of reference.
-The aberration of light effect can be described within the standard special relativity taking advantage of the Wigner rotation theory.
-When one has a transversely moving emitter there is the deviation of the energy transport for radiated light.
-This effect is a consequence of the fact that the Doppler effect is responsible for the angular frequency dispersion of the radiated light waves (see Fig.
-1).
-As viewed from the Lorentz lab frame, the coordinate axes of the accelerated frame is rotated through the angle $v/c$ with respect to the coordinate axes of the lab frame.
-In the first order approximation in $v/c$, the accelerated emitter-detector setup exhibits a shearing motion (during acceleration).
-This shearing contribution adds up to the Wigner rotation.
 The wavefront orientation has no exact objective meaning.
 Because of this, only the wavefront orientation in the absolute time coordinatization for the moving source is presented here for sketch simplification.
-Figure 37: Aberration of light in an accelerated frame of reference.
-In order to predict result of the aberration measurement, the accelerated observer should use the Langevin metric.
-On accelerated system, the accelerated detector does not receive light radiated by the accelerated emitter.
-Figure 38: Aberration of light in an inertial frame of reference as predicted by the conventional theory (see Fig. 1).
+Figure 33: Aberration of light in an accelerated frame of reference. In an accelerated system, the accelerated detector does not receive light radiated by the accelerated emitter.
+Figure 34: Aberration of light in an inertial frame of reference as predicted by the conventional theory.
 Figure 35: Aberration of light in an accelerated frame of reference as predicted by the conventional theory.
-An acceleration does not spoil the motional symmetry between the accelerated reference frame and inertial reference frame.
-Common textbook presentations of the special relativity use the (3+1) approach which deals with hidden assumption that distant clocks are synchronized according to the absolute simultaneity convention.
-In fact, according to textbooks, when a reference frame at rest is put into motion, then all points in the 3D reference grid have a start to move at same time.
-Consequently, rigidity in the sense introduced by textbooks cannot be considered as a relativistic kinematics property.
+Acceleration does not spoil the motional symmetry between the accelerated reference frame and the inertial reference frame.
+Common textbook presentations of the special relativity use the (3+1) approach which deals with the hidden assumption that distant clocks are synchronized according to the absolute simultaneity convention.
+According to textbooks, when a reference frame at rest is put into motion, then all points in the 3D reference grid start to move at the same time.
 The standard presentation of the aberration of light effect is based on the hidden assumption that $(x_{x},y_{n},z_{n})$ axes of the accelerated frame and $(x,y,z)$ axes of the inertial frame are parallel.
 In other words, it is incorrectly assumed that accelerated and inertial observers have common 3-space, Fig.
-38,39.
-According to the conventional theory, an acceleration does not spoil the motional symmetry between the accelerated reference frame and inertial reference frame.
+34 - Fig.35.
+According to the conventional theory, acceleration does not spoil the motional symmetry between the accelerated reference frame and the inertial reference frame.
 Our result (Fig.
-36,37 ) is at odds with the prediction from textbooks.
-The commonly accepted derivation of the aberration of light effect does not account for the Wigner rotation, which in our case closely associated with the relativity of simultaneity.
+32 - Fig.33 ) is at odds with the prediction from textbooks.
+The commonly accepted derivation of the aberration of light effect does not account for the Wigner rotation, which in our case is closely associated with the relativity of simultaneity.
+In the next chapter we shall consider the Wigner rotation theory in great detail.
 7.4 Discussion
 Let us return to observations of an inertial observer.
-Special theory of relativity says that the orthogonal emitter-detector setup exhibits a shearing motion during acceleration, Fig.
-36.
+The special theory of relativity says that the orthogonal emitter-detector setup exhibits a shearing motion during acceleration, Fig.
+32.
 This is a purely kinematic effect, involving no forces.
@@
 That’s not an easy question, but here is one way of thinking about this.
-It is clear that our ordinary intuition are quite wrong when we have deal with two space-like separated events.
-We see that the accepted in textbooks assumption of rigidity of accelerated frame is based on the belief that simultaneous acceleration of the reference space grid has direct physical meaning.
-This is misconception.
-In fact, when we have two distant events, we have deal with the conventionality of distant simultaneity within the time interval $L/c$, where $L$ is the space separation of these two events.
+It is clear that our ordinary intuition is quite wrong when we have to deal with two space-like separated events.
+We see that the generally accepted assumption of rigidity of the accelerated frame is based on the belief that simultaneous acceleration of the reference space grid has direct physical meaning.
+This is a misconception.
+When we have two distant events, we have to deal with the conventionality of distant simultaneity within the time interval $L/c$, where $L$ is the space separation of these two events.
 The position along the $x$-axis of the accelerated emitter with respect to the accelerated detector has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this position could be ascertained.
-There is an uncertainty (blurring) of the relative position in the $x$-direction of amount $Lv/c$ (due to the uncertainty in the moments of the acceleration).
+There is an uncertainty (blurring) of the relative position in the $x$-direction of amount $Lv/c$ (due to the uncertainty in the moments of the acceleration)  [^37] Our next example has to do with a light clock.
+A clock which we call a ”light clock” is a rod with a mirror at each end.
+According to textbooks, when the inertial observer looks at the orthogonal light clock going by, he sees that light, in going from mirror to mirror is taking a zigzag path.
+This standard explanation is based on the hidden assumption that $(x_{n},y_{n},z_{n})$ axes of the moving observer and $(x,y,z)$ axes of the inertial observer are parallel.
+This is a misconception.
+Let us consider the phase (and initialization) of light clocks.
+The word ”phase” means the zigzag shape of the path followed by the light beam.
+The shape of the zigzag pattern cannot be measured in the lab frame.
+The position of the accelerated bottom mirror with respect to the accelerated top mirror has no exact objective meaning.
+There is uncertainty of amount $lv/c$, where $l$ is the rod length..
 There are several points to be made about the above results.
-Above we discussed that the aberration of light effect in accelerated frames can be described in an inertial frame taking advantage of the Wigner rotation theory.
-To be complete we must add an analysis of the vertical displacement of the detector.
+We already know that the aberration of light effect in accelerated frames can be described in an inertial frame taking advantage of the Wigner rotation theory.
+One of the important conclusions of the discussion presented above is that inertial and accelerated observers have no common line of motion.
+To be complete we must add an analysis of the angular displacement of the $x_{n}$-axis.
 We have considered only the case in which motion is non-relativistic in speed.
-In other words, we use a first order relativistic approximation.
-It is easy to see that for the vertical displacement we have $\Delta x\simeq\theta_{w}Lv/c\simeq Lv^{2}/c^{2}$.
-So special relativity says that there is no difference between the (kinematic) shearing motion and the (Wigner) rotation motion (during acceleration) as $v^{2}/c^{2}\to 0$.
-7.5 Aberration of Relativistic Particles
+In other words, we use a first-order relativistic approximation.
+It is easy to see that for the displacement of the detector along the $x$-axis in the $z$ direction we have $\Delta z\simeq\theta_{w}Lv/c\simeq Lv^{2}/c^{2}$.
+So special relativity says that there is no difference between the (kinematic) shearing motion (during acceleration) and the (Wigner) rotation motion as $v^{2}/c^{2}\to 0$.
+8 Kinematics of the Wigner Rotation
+The subject of this section is the Wigner rotation  [^38] As known, a composition of noncollinear Lorentz boosts does not result in a different boost but in a Lorentz transformation involving a boost and a spatial rotation, the latter being known as Wigner rotation [39, 40].
+Wigner rotation is sometimes called Thomas rotation (see e.g.
+[13, 41])..
+We have already discussed the Wigner rotation in the previous section.
+Our earlier discussion is really about the aberration of light.
+The Wigner rotation is also related to the aberration of particles.
+The Wigner rotation is a relativistic kinematic effect, which consists in that the coordinate axes of a reference frame, moves along a curvilinear trajectory rotating about the axes of a Lorentz inertial frame.
+To deduce this expression is the main purpose of this chapter.
+We have tried to keep the mathematical complexity of the discussion to a minimum.
+8.1 The Commutativity of Collinear Lorentz Boosts
+Let us now consider a relativistic particle, accelerating in the initial inertial frame, and let us analyze its evolution within Lorentz coordinate systems.
+The permanent rest frame of the particle is not inertial.
+To get around that difficulty one introduces an infinite sequence of comoving frames.
+At each instant, the rest frame is a Lorentz frame centered on the particle and moving with it.
+As the particle velocity changes to its new value at an infinitesimally later instant, a new Lorentz frame centered on the particle and moving with it at the new velocity is used.
+Let us denote the three inertial frames by $K,K^{\prime},K"$.
+The initial inertial frame is $K$, $K^{\prime}$ is the particle’s rest frame with velocity $\vec{v}=\vec{v}(\tau)$ relative to $K$, and $K"$ is the particle’s rest frame at the next instant of proper time $\tau+d\tau$, which moves relative to $K^{\prime}$ with infinitesimal velocity $d\vec{v^{\prime}}$.
+All inertial reference frames are assumed to be Lorentz reference frames.
+To have this, we impose that $K^{\prime}$ is connected to $K$ by the Lorentz boost $L(\vec{v})$, with $\vec{v}$, which transforms a given four vector event $X$ in space-time into $X^{\prime}=L(\vec{v})X$.
+The relation $X"=L(d\vec{v^{\prime}})L(\vec{v})X$ presents a step-by-step change from $K$ to $K^{\prime}$ and then to $K"$.
+There is another composition of reference-frame transformations that describes the same particle evolution in the Minkowski space-time.
+Let $R^{\prime}$ be an inertial frame with velocity $d\vec{v}$ relative to the initial inertial frame $R$.
+We impose that $R^{\prime}$ is connected to $R$ by the Lorentz boost $L(d\vec{v})$.
+The Lorentz rest frame $R"$ is supposed to move relative to the Lorentz frame $R^{\prime}$ with velocity $\vec{v}$.
+The relation $X"=L(\vec{v})L(d\vec{v})X$ presents a step-by-step change from $R$ to $R^{\prime}$ and then to the rest frame $R"$.
+Let us examine the transformation of the three-velocity in the theory of relativity.
+For a rectilinear motion along the $z$ axis, it is performed in accordance with the following equation: $v_{z}(\tau+d\tau)=(dv_{z}+v_{z})/(1+v_{z}dv_{z}/c^{2})$.
+As it happens with the composition of Galilean boosts, collinear Lorentz boosts commute: $L(dv_{z})L(v_{z})=L(v_{z})L(dv_{z})$.
+This means that the resultant of successive collinear Lorentz boosts is independent of which transformation applies first.
+8.2 The Noncommutativity of Two Lorentz Boosts in Nonparallel Directions
+In contrast with the case of Lorentz boosts in collinear directions, Lorentz boosts in different directions do not commute.
+While the successive application of two Galilean boosts is a Galilean boost and the successive application of two rotations is a rotation, the successive application of two non-collinear Lorentz boosts is not a Lorentz boost.
+Let us compare the succession $K\to K^{\prime}\to K"$ with the succession $R\to R^{\prime}\to R"$ in the case when the particle’s acceleration in the frame $K^{\prime}$ is perpendicular to the line of flight of the initial inertial frame.
+The frame $K"$ is supposed to move relative to $K^{\prime}$ with velocity $d\vec{v^{\prime}}_{x}$.
+Now, since we can write the result in terms of succession $L(\vec{v}_{z})L(d\vec{v}_{x})$ as well as in terms of succession $L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})$, there is a need to clarify a number of questions associated with these compositions of Lorentz frames.
+We can easily understand that the operational interpretation of the succession $L(\vec{v}_{z})L(d\vec{v}_{x})$ is particularly simple, involving physical operation used in the measurement of the particle’s velocity increment $d\vec{v}_{x}$ in the initial inertial frame.
+We should be able to understand the operational interpretation of the succession $L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})$.
+The inertial observer sees the time dilation in the Lorentz frame which moves with respect to the initial inertial frame with velocity $\vec{v}_{z}$: $dt/\gamma=d\tau$.
+Because of time dilation in the moving frame, the velocity increment in the initial inertial frame $dv_{x}$ corresponds to a velocity $\gamma dv_{x}$ and $-\gamma dv_{x}$ in the frames $K^{\prime}$ and $K"$ respectively.
+The resulting boost composition can be represented as $X"=L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})X=L(\vec{v}_{z})L(d\vec{v}_{x})X$.
+In other words, Lorentz boosts in different directions do not commute: $L(\vec{v}_{z})L(d\vec{v}_{x})\neq L(d\vec{v}_{x})L(\vec{v}_{z})$.
+8.3 Wigner Rotation
+A large number of incorrect expressions for the Wigner rotation can be found in the literature.
+Therefore, there is a need to consider the Wigner rotation in detail.
+Rather than working out all the transformation matrices for four-vector components, we would like to show a geometrical approach that is very useful for this problem.
+[…]
+There are several points to be made about the Wigner rotation theory.
+We wish to remark that the expression for the Wigner rotation angle in the proper frame (as viewed from the Lorentz $K$ frame), $d\vec{\Phi}_{K"}=(\gamma-1)\vec{v}\times d\vec{v}/v^{2}$, can be presented in the form $d\vec{\Phi}_{K"}=(1-1/\gamma)\vec{v}_{K"}\times d\vec{v}_{K"}/v_{K"}^{2}$.
+Thus, the Wigner rotation angle in the proper frame is expressed in terms of the $K$ frame velocity $\vec{v}_{K"}=-\vec{v}$ and its increment $d\vec{v}_{K"}=-\gamma d\vec{v}$ in the proper frame.
+Transformed back to the lab Lorentz frame this becomes $d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$.
+This is the Wigner rotation angle of the proper frame axes in the initial inertial (lab) frame $K$  [^39] The correct expression for the Thomas (Wigner) rotation was first obtained by V.
+Ritus [42].
+In deriving expressions for the Thomas (Wigner) rotation, the majority of authors (see e.g.
+[41]) was supposedly guided by the incorrect expression for Thomas (Wigner) rotation from Moeller’s monograph [13].
+The expression obtained by Moeller is given by $\vec{\delta\Phi}=(1-\gamma)\vec{v}\times d\vec{v}/v^{2}=(1-\gamma)\vec{\delta\theta}$ (and subsequently expression for Thomas precession $\Omega_{\mathrm{T}}=(1-\gamma)\omega_{0}$).
+It should be noted that, in his monograph, Moeller stated several times that this expression is valid in the lab Lorentz frame.
+Clearly, this expression and correct result differ both in sign and magnitude.
+An analysis of why Moeller obtained an incorrect expression for the Thomas (Wigner) rotation in the lab frame is the focus of Ritus paper [43].
+As shown in [43], the Moeller’s mistake is not computational, but conceptual in nature.
+In review [44] it is shown that the correct result was obtained in the works of several authors, which were published more than a half century ago but remained unnoticed against the background of numerous incorrect works..
+We note that owing to the relativistic effect of time dilation in the reference frame that moves to the lab frame, the Wigner rotation angle in the proper frame is always $\gamma$ time higher than in the lab frame  [^40] We note that in 1986, M.
+Stranberg obtained an expression for the Thomas (Wigner) rotation correct both in the lab inertial frame and the comoving reference frame [45].
+It is noteworthy that [45] is one of the few papers that explicitly states that the angle of the Thomas (Wigner) rotation in the comoving reference frame is $\gamma$ times higher than in the lab frame..
+We derived the exact relation $d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$ using only rudimentary knowledge of special relativity.
+In textbooks on the theory of relativity, the spatial rotation associated with the composition of two Lorentz boosts in non-parallel directions is often introduced using the algebraic approach.
+This is one of the reasons why authors of textbooks obtained an incorrect expression for the Wigner rotation.
+They describe the rotation of a moving object without the geometrical meaning of such rotation and encounter serious difficulties in the interpretation of the applied calculations and of the results.
+Figure 36: The interpretation of the Wigner rotation about the proper frame axes.
+8.4 How to Measure a Wigner rotation?
+The Wigner rotation is a relativistic kinematic effect, which consists in that the coordinate axes of a reference frame, moving along a curvilinear trajectory rotating about the axes of a Lorentz lab frame.
+In the Lorentz lab frame, for infinitely small transformations (due to acceleration) we obtained the formula
+$d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}=\left(1-\frac{1}{\gamma}\right)\vec{\delta\theta}~{},$ (20)
+where $d\vec{v}$ is the vector of small velocity change due to acceleration, $\Phi$ is the Wigner rotation angle of the spatial coordinate axes of the proper system relative to Lorentz lab frame, and $\theta$ is the orbital angle.
+But how to measure this orientation?
+A moving coordinate system changes its position in time.
+The question arises whether it is possible to give an experimental interpretation of the rotation of a moving coordinate system.
+We illustrate the problem of how to represent the orientation of the moving coordinate system with a simple example.
+The execution of successive transformations from $R$ to $R^{\prime}$ at velocity $d\vec{v_{x}}$ and from $R^{\prime}$ to $R"$ at velocity $\vec{v_{z}}$ equivalent to the composition of boost $L(\vec{v}_{z}+d\vec{v}_{x})$ and rotation.
+The Wigner rotation Eq.(20) is performed additionally to the Lorentz boost at velocity $\vec{v_{z}}+d\vec{v_{x}}$.
+The interpretation of this rotation about the initial inertial frame of reference is closely associated with the length contraction  [^41] Length of moving object is determined not only by the properties of space-time but also by our synchronization convention.
+Therefore, similar to the radiation phase front orientation, it has no exact objective meaning.
+Such quantities (under the standard clock synchronization) usually are called ”relativistic kinematics effects”.
+The Wigner rotation is also a coordinate effect with no exact objective meaning..
+Figure 37: The interpretation of the Wigner rotation about the lab frame.
+Suppose that the lab observer after the Lorentz boost at velocity $v_{z}$ rotates the coordinate system on the angle $\vec{\delta\theta}=\vec{v_{z}}\times d\vec{v_{x}}/v_{z}^{2}$ and now $x_{r}$ locates orthogonally to the vector $\vec{v_{z}}+d\vec{v_{x}}$.
+Similarly, the axis $z_{r}$ is parallel to the vector $\vec{v_{z}}+d\vec{v_{x}}$.
+Consider a rod directed along $x$-axis in the comoving frame.
+The motion takes place in the plane $(x,z)$ and the rod is located perpendicularly to the velocity $\vec{v_{z}}$.
+After the rotation of the lab frame axes, the projection of the rod on the $z_{r}$ axes will be simply $l\delta\theta$, where $l$ is the rod length in the $R$ frame and also in the lab frame after the first boost along the $z$-axis.
+After the second Lorentz boost at velocity $d\vec{v_{x}}$ this projection will be contracted down to $l\delta\theta/\gamma$ (Fig.
+37).
+Let the observer in the lab frame fix the position of the axes of the comoving frame.
+In ultrarelativistic limit $\gamma\to\infty$ these axes will be parallel to the rotated lab frame axes $(x_{r},z_{r})$.
+The projection of the rod on the $z_{r}$ axis will be zero.
+In the case of an arbitrary velocity, the axes of the comoving frame are not parallel to the rotated lab frame axes $(x_{r},z_{r})$.
+According to contracted projection, the angle will be $-\vec{\delta\theta}/\gamma$.
+And one can verify directly that the axes of the comoving frame are rotated with respect to the initial lab frame axes $(x,z)$ by the angle equal to $\delta\theta-\delta\theta/\gamma$, which is just the Wigner rotation angle in accordance with equation Eq.(20).
+Here we only wished to show how naturally Lorentz transformations lead to the Wigner rotation phenomenon.
+8.5 Discussion
+A large number of incorrect expressions for the Wigner (Thomas) rotation can be found in the literature.
+Let us discuss the reasons that underlie the derivation of the incorrect expressions.
+The Wigner rotation also relates the angular rotation velocity (Thomas precession) of the spin of elementary particles following a curvilinear orbit.
+In 1959, Bargman, Michel, and Telegdi (BMT) proposed a consistent relativistic theory for the dynamics of the spin as observed in the lab frame, which was successfully tested in experiments [46].
+It is commonly believed that the BMT equation contains the standard (and incorrect) result for the Wigner rotation in the Lorentz lab frame.
+There are many physicists who have already received knowledge about the Wigner (Thomas) rotation from well-known textbooks and who would say, ”The extremely precise measurements of the magnetic-moment anomaly of the electron made on highly relativistic electrons are based on the BMT equation, of which the Wigner rotation is an integral part, and can be taken as experimental confirmations of the standard expression for the Wigner rotation”.
+This misconception about the experimental test of the incorrect expression for the Wigner rotation in the lab frame is common and pernicious.
+We should remark that the results in the Bargmann-Michel-Telegdi paper were obtained by the method of semi-classical approximation of the Dirac equation.
+The Wigner rotation was not considered at all, because the Dirac equation allows obtaining the solution for the total particle’s spin motion without explicitly splitting it into the Larmor and Wigner parts.
+In should be noted that the situation relating to the use of the Wigner (Thomas) rotation theory in the relativistic spin dynamics is complex.
+The starting point of Bargman, Michel, and Telegdi was the particle rest frame and the equation of motion for particle angular momentum, which they generalized to the Lorentz lab frame and then transformed back to the rest frame.
+This brings up an interesting question: Why it is convenient to transform this equation to the rest frame as of that instant?
+The approach in which we deal with the proper spin is much preferred in the experimental practice due to the clear physical meaning of the (3D) spin vector $\vec{s}$.
+Unlike the (3D) momentum $\vec{p}$ , which has definite components in each reference frame, angular momentum is defined only in one particular reference frame.
+It does not transform.
+Any statement about it refers to the rest frame as of that instant.
+If we say that in the lab frame, the spin of a particle makes the angle $\phi$ with its velocity, we mean that in the particle’s rest frame, the spin makes this angle with the line motion of the lab frame.
+The analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard (Einstein) special relativity taking advantage of well-known relativistic kinematic effects.
+In the case of spin dynamics, one has no occasion to make measurements with non-inertial devices.
+The conventional method used to explain the spin dynamics in the lab frame is very unusual.
+First, we already pointed out that the laws of physics in the lab frame are able to account for all physical phenomena, including the observations made by non-inertial observers.
+In particular, it is possible to determine the result of spin orientation measurements with respect to the lab frame axes in the proper frame.
+In this case, when viewed from the lab frame, the interpretation of the spin rotation experiments in the lab frame is simple and may be derived from the prediction of the proper observer as concerns the measurement made by the lab observer.
+Therefore, the spin rotation measurement by a polarimeter in the lab frame is interpreted from the viewpoint of the proper observer as viewing of the lab observer.
+This complexity is one of the reasons why authors of textbooks obtained an incorrect expression for the Wigner rotation.
+We will discuss this subject further in Chapter 17.
+It is also important to elucidate the reason why did the error in the Wigner rotation theory remained so long undetected.
+Wigner (Thomas) rotation is generally presented only in the context of spin kinematics as a peculiar effect of special relativity.
+This is supposedly the reason why until recently no researches have noticed so serious a discrepancy between the results of different works.
+We emphasize that the Wigner rotation is a very basic relativistic phenomenon.
+It is as basic as the time dilation and length contraction.
+The theory of special relativity shows that the aberration effect in the rotating earth-based frame, as viewed from the inertial sun-based frame (in the Lorentz coordinatization), presents a kinematic effect of the special theory of relativity.
+In this case, when viewed from the sun-based frame, the aberration image shift in the earth-based frame is regulated by the Wigner rotation.
+According to the Wigner rotation theory, two observers with different trajectories have different 3-spaces.
+In contrast, the accepted in previous literature incorrect assumption that an earth-based observer and an inertial (e.g.
+sun-based) observer have a common 3-space has always been considered obvious.
+9 Aberration of Relativistic Particles
 We want now to discuss the phenomenon of the aberration of relativistic particles.
-We begin by recalling a very interesting (methodological) analogy between the aberration of light effect and the deflection of light in general relativity.
-Consider a ”point” mass $m$ moving with a velocity $V$ in the gravitation field of a point mass $M$.
-Both the motion of particles and the path of light in the gravitation field are described by the same equation of motion.
-The deflection angle for a particle is $\theta_{d}=(1+V^{2}/c^{2})\theta_{c}$, where $\theta_{c}$ is the particle deflection angle according to classical approach.
-The factor $(1+V^{2}/c^{2})$ is new from relativistic (Schwarzschild) metric, but reduced to familiar Newtonian limit as $V^{2}/c^{2}\to 0$.
-For light $V/c=1$ and there is an extra factor 2.
-On the basis of classical theory one gets a deflection of light that is only one-half as big as predicted by general relativity.
-Suppose we have a particle source.
+Suppose we have a particle source at rest in the initial inertial frame.
 All particles which come out of the source will have the same velocity $V$.
 Particles trace paths along the $z$-axis.
-We consider the case when the source is at rest in the initial inertial frame and the observer (with his measuring devices) is accelerated from rest up to velocity $v$ along the $x$ axis.
+Let us consider the case when the observer (with his measuring devices) and the particle source are accelerated from rest up to velocity $v$ along the $x$ axis.
 It is assumed that $v\ll V$.
 Suppose that an observer in the accelerated frame performs an aberration measurement.
-How shall we describe the aberration of particle beam from the source which is at rest in the inertial frame?
-In order to predict the result of the aberration measurement the accelerated observer should use the Langevin metric Eq.(15).
-The aberration increment is connected with the problem parameters by the relation $\theta_{a}=-[1+(1-1/\gamma_{p})]\theta_{c}$, where $\theta_{c}=v/V$ is the particle deflection angle according to classical approach.
-Here $\gamma_{p}=1/\sqrt{1-V^{2}/c^{2}}$ is the relativistic factor.
-The term $[1+(1-1/\gamma_{p})]$ is new from relativistic (Langevin) metric, but reduced to familiar classical limit as $V^{2}/c^{2}\to 0$.
-For light $V/c=1$ and there is an extra factor 2.
-A correct solution of this problem in the rotating frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of a relativistic particle propagation.
-In the Chapter 8 (see Section 8.2) we will find that it is very easy to treat the problem by the use of dynamical line of arguments.
-The electromagnetic forces which govern the properties of an emitted electron beam must be affected by source acceleration in such a way that they lead to a deviation of the electron transport direction.
-We can obtain the electromagnetic fields in accelerated frame using Galilean transformation.
-We want now to find the aberration of ultrarelativistic particles from stationary and moving sources.
-Imaging that there are two identical sources.
-For the sake of simplicity we assume that the relativistic factor of the beam $\gamma_{p}\gg 1$.
-Let us consider the case when the first source is at rest in an inertial frame and the second source is accelerated from rest up to velocity $v$ along the $x$-axis.
-Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the particle transport measurement.
-Now we must be careful about the initial direction of the velocity of a particle beam.
-As example, we consider the case in which particles are initially directed along the $z$-axis.
-Then how does the particle beam from the moving source looks?
-The inertial observer would find that angular displacement is equal to $\theta_{a}=v/c$.
-This is example what called aberration of relativistic particles and it is well known.
-Figure 40: The aberration of ultrarelativistic ($\gamma_{p}\gg 1)$ particles from stationary and moving sources.
-If the particle source is at rest relative to the fixed stars and the observer with measuring devices started from rest to uniform motion with velocity $v\ll c$, then the apparent angular position of the source seen in the accelerated frame would jump by angle $\theta_{a}=-2v/c$.
-This situation is not symmetrical with respect to the change of the reference frames.
-If observer is at rest relative to the fixed stars and the source started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the source seen in the inertial frame would jump by angle $\theta_{a}=v/c$
-Figure 41: Reciprocity in the theory of the aberration of particles.
-One must use the same beam directing to demonstrate the reciprocity.
-After the redirection procedure of the accelerated source the inertial observer would find that the angular displacement is $\theta_{a}=2v/c$ (for the sake of simplicity it is assumed that $\gamma_{p}\gg 1$).
-The difference between measurements made in both reference frames is very interesting.
-According to the asymmetry between the inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of particles.
-Namely, if the source is at rest relative to the fixed stars and the observer with measuring devices started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the emitter seen in the accelerated frame would jump by angle $\theta_{a}=-2v/c$.
-This situation is not symmetrical with respect to the change of the reference frames, Fig.
-40.
-It could be said that the crossed term in metric Eq.(15) generates anisotropy in the accelerated frame that is responsible for the difference between aberration increments.
-When the accelerated system starts moving with constant velocity the procedure of beam redirection can be performed.
-Let us describe what happens when accelerated observer performed the redirection of the accelerated source.
-We consider the case in which finally the particles trace paths along the $z_{n}$-axis.
-After this redirection procedure the inertial observer would find that angular displacement is $\theta_{a}=2v/c$.
-Suppose that the accelerated observer also performs an aberration measurement.
-Fig.
-41 shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-2v/c$.
-Above in this chapter we demonstrated that aberration of light effect in the accelerated frame, as viewed from the inertial lab frame, presents a well known kinematic effect of special theory of relativity.
-What about particles?
+How shall we describe the aberration of particle beam from the source which is at rest in the accelerated frame?
+To predict the result of the aberration measurement the accelerated observer should use the Langevin metric Eq.(15) (see Section 9.2).
+The aberration increment is connected with the problem parameters by the relation $\theta_{a}=-(1-\sqrt{1-V^{2}/c^{2}})v/V$.
+We should point out that this result is valid for an arbitrary Lorentz factor $\gamma=1/\sqrt{1-V^{2}/c^{2}}$.
+The term $(1-1/\gamma)v/V$ is new from the relativistic (Langevin) metric but reduced to the familiar classical limit as $V^{2}/c^{2}\to 0$.
+A correct solution to this problem in the accelerated frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of relativistic particle propagation.
+9.1 Inertial Frame View of Observations of the Noninertial Observer
+We can, of course, analyze this situation in various other ways.
+Above we demonstrated that the aberration of light effect in the accelerated frame, as viewed from the initial inertial frame, presents a well-known kinematic effect of the special theory of relativity.
 Can we not look at the aberration of particles in the accelerated frame in the same way?
 As we shall see, the aberration of particles effect in non-inertial frames of reference can be described in an inertial frame within standard special relativity taking advantage of the Wigner rotation theory.
-The Wigner rotation is a relativistic kinematic effect, which consists in that a coordinate axes of a non-inertial reference frame is rotating about the axes of a (Lorentz) initial inertial frame.
-What is the formula for Wigner rotation?
-It turns out that this is complicated, and it takes a great deal of study and sophistication to appreciate it.
-(We shall do this in the Chapter 17).
-We only write it down to give the impression.
-The expression for the rotation angle in the inertial frame can be presented in vector form $\vec{\theta_{w}}=(1-1/\gamma_{p})\vec{V}\times\vec{v}/V^{2}$, where $\vec{V}$ is the vector of particle velocity in the initial Lorentz frame before acceleration, $\vec{v}$ is the vector of small velocity change ($v\ll V$) due to acceleration, $\theta_{w}$ is the Wigner rotation angle of the spatial coordinate axes of the system comoving with a particle relative to Lorentz lab frame.
-We shall discuss this expression further in the Chapter 17.
-To find the rotation magnitude in the stated problem, we introduce a composition of boosts.
-Let $S$ be the initial inertial frame of reference, $S_{n}$ an accelerated frame with velocity $\vec{v}$ relative to the inertial frame, and $S^{\prime}$ is a particle comoving frame which moves (after the redirection of the accelerated source) relative to the $S_{n}$ with velocity $\vec{V}$ along the $z_{n}$-axis.
+The Wigner rotation is a relativistic kinematic effect, which consists in that the coordinate axes of a non-inertial reference frame are rotating about the axes of a (Lorentz) initial inertial frame.
+The expression for the rotation angle in the inertial frame can be presented in vector form $\vec{\theta_{w}}=(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$, where $\vec{V}$ is the vector of particle velocity in the initial Lorentz frame before acceleration, $\vec{v}$ is the vector of small velocity change ($v\ll V$) due to acceleration, $\theta_{w}$ is the Wigner rotation angle of the proper frame axes in the (Lorentz) inertial frame.
+We want now to see how this formula for the Wigner rotation works in the case when the observer and particle source are accelerated up to velocity $v$ in the inertial frame.
+To find the rotation magnitude in the stated problem, we introduce a composition of Lorentz boosts.
+Let $S$ be the initial inertial frame of reference, $S_{n}$ an accelerated frame with velocity $\vec{v}=\vec{e}_{x}v$ relative to the inertial frame, and $S^{\prime}$ is a frame which moves relative to the $S_{n}$ with velocity $\vec{V}$ along the $z_{n}$-axis.
 Two sequential boosts from the inertial frame $S$ to $S_{n}$ and then to $S^{\prime}$ are equivalent to the boost from $S$ to $S^{\prime}$ and the subsequent rotation.
-The comoving frame $S^{\prime}$ is rotated through the angle $\theta_{w}$ with respect to the inertial frame $S$.
-Now we must be careful about direction of rotation.
+The frame $S^{\prime}$ is rotated through the angle $\theta_{w}$ with respect to the inertial frame $S$.
+Now we must be careful about the direction of rotation.
 There is a good mnemonic rule.
 The rule says that the direction of the Wigner rotation in the Lorentz lab frame is the same as the direction of the velocity rotation in the Lorentz lab frame.
-We can easily understand that the coordinate axes of the particle comoving frame $(x^{\prime},y^{\prime},z^{\prime})$ will be parallel to the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$.
-As viewed from the inertial frame, the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$ is also rotated through the angle $\theta_{w}$ with respect to the coordinate axes of the inertial frame $(x,y,z)$, Fig.
-41.
+We can easily understand that the coordinate axes of the frame $(x^{\prime},y^{\prime},z^{\prime})$ will be parallel to the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$.
+As viewed from the inertial frame, the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$ are also rotated through the angle $\theta_{w}$ with respect to the coordinate axes of the inertial frame $(x,y,z)$.
 Our results show that we cannot remain with the framework of parallel axes $z_{n},z$ when considering the inertial frame view of particle beam observations of the accelerated (up to the velocity $v$) observer in perpendicular geometry.
-For light $V/c\to 1$ and the axes of accelerated frame are then rotated with respect to the inertial frame axes by angle $\theta_{w}=v/c$.
+So we have expression for the aberration increment in the accelerated frame: $\theta_{a}=-\theta_{w}=-(1-1/\gamma)v/V$.
+For light $V/c\to 1$ and the aberration increment becomes $\theta_{a}=-v/c$.
+9.2 Explanation of the Aberration of Electrons on the Basis of Electrodynamics
+It is very easy to treat the aberration of particle problems by the use of a dynamical line of arguments.
+The electromagnetic forces that govern the properties of an emitted electron beam must be affected by source acceleration in such a way that they lead to a deviation of the electron transport direction.
+We can obtain the electromagnetic fields in an accelerated frame using Galilean transformation.
+For those who have already received knowledge about special relativity from well-known textbooks, it may be expedient to begin its aberration of particle (theoretical and experimental) development with the microscopic approach.
+The point is that one can compute any relativistic quantity directly from the underlying theories of matter without involving relativistic kinematics effects at all.
+The aberration of particles associated with the transformation from the inertial frame to the accelerated frame may be regarded as a result of the action of a certain force.
+This problem has no relation to the problem of the ether and is already formulated within the framework of special relativity.
+Let us consider what happens if the electron gun, an electron source for a relativistic electron beam, in the initial inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis.
+The electromagnetic forces that govern the properties of an emitted electron beam must be affected by the source acceleration in such a way that they lead to a deviation of the electron transport direction.
+There is a machinery behind the electron momentum $\vec{p}$ changes when the source and observer (with his measuring instruments) are at rest in the accelerated frame.
+Its origin is explained in the framework of the electrodynamics theory.
+We can obtain the electrodynamics equations in an accelerated reference frame using the Galilean transformation.
+When we consider the electrodynamics equations under Galilean transformation, there is a question: what is the transformation law for the electromagnetic fields $\vec{E}$ and $\vec{B}$?
+The fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame where the particle source is at rest are not the same as those $\vec{E}$ and $\vec{B}$ in the inertial frame before the active Galilean boost.
+We need to find the relation between $\vec{E}_{n}$ and $\vec{B}_{n}$ and $\vec{E}$ and $\vec{B}$.
+To do this, we transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$, using a Galilean transformation: we substitute $x_{n}=x-vt$, while leaving time unchanged $t_{n}=t$ into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain Langevin metric $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$.
+We can obtain the electrodynamics equations in an accelerated frame using the Galilean transformation (with velocity $-v$) of Maxwell’s equations.
+With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$.
+This expression describes first-order (in $v/c$) effects only.
+Here $\vec{v}$ is the velocity vector of the source in the initial inertial frame.
+Figure 38: Aberration of electrons when the electron source and observer are at rest in the accelerated frame.
+[…]
+We see that the aberration of particles associated with the transformation from the inertial coordinate system to the accelerated coordinate system may be regarded as a result of the action of Lorentz force.
+The electromagnetic forces that govern the properties of an emitted electron beam must be affected by acceleration with respect to the fixed stars in such a way that they lead to the deviation of the electron beam transport direction, Fig.
+38.
+The formula $\vec{\theta}_{a}=-(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$ correspond to the result we demonstrated for observation of the non-inertial observer taking advantage of the Wigner rotation theory.
+The theory of special relativity shows that the aberration of particles effect in the accelerated frame, as viewed from the initial inertial frame (in the Lorentz coordinatization), presents a kinematic effect of the special theory of relativity.
+In this case, when viewed from the initial inertial frame, the aberration image shift in the accelerated frame is regulated by the Wigner rotation.
+According to the Wigner rotation theory, two observers with different trajectories have different 3-spaces.
+In contrast, the accepted in previous literature incorrect assumption that a non-inertial (e.g.
+earth-based) observer and an inertial (e.g.
+sun-based) observer have a common 3-space has always been considered obvious.
+It is important to stress at this point that the dynamical line of arguments discussed here explains what the Wigner rotation physically means in the rotating frame.
+The relativistic kinematics effects in the rotating frame, as viewed from the inertial observer, are only an interpretation of the behavior of the electromagnetic fields.
+9.3 Magnetic Field Measurements in the Accelerated Electron Source
+We would now like to describe an apparent paradox.
+In special relativity, we are confronted with a few things that might seem counterintuitive.
+But this is still a mathematically based theory.
+It tells us a peculiar thing: when we are accelerated electron source, the fields in the accelerated frame are not static; the problem is time dependent.
+That is what special relativity says!
+When the electron source and observer are at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be nonzero.
+Intuition would seem to tell us that everything is at rest, so the electric field is not changing with time.
+But special relativity says that there is a derivative $\partial\vec{E}/\partial t_{n}=v\partial\vec{E}/\partial x_{n}$ because there is $\partial\vec{E}/\partial x_{n}$ that is not zero.
+This is just a feature of the 4D formalism with no special consequence.
+A comparison with a gauge transformation in classical electromagnetism might help here.
+In electrodynamics, it is often convenient to introduce the Coulomb gauge.
+In this case field propagates instantaneously.
+But this instantaneous propagation is just a mathematical description in an intermediate step: when a complete calculation is made, proper cancellations of the instantaneous propagation take place.
+We obtained the electrodynamics equations in an accelerated frame using the inverse Galilean transformation (with velocity $-v$) of Maxwell’s equations.
+With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$.
+The electric $\vec{E}_{n}$ and magnetic $\vec{B}_{n}$ fields are not changing with time - all of the fields are static so that the problem is not time-dependent.
+There are many physicists who have already received knowledge about electrodynamics from textbooks and who would say, ”There must be currents to get a static magnetic field at all - and currents can only be from moving charges”.
+Since there are no moving charges , the situation seems indeed paradoxical.
+Here we have a new kind of situation that is quite different from (initial) inertial frame electrodynamics.
+How can that be?
+When the electron source and observer are at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be nonzero.
+One can easily show this by transforming coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$.
+The variables $x,y,z,t$ can be expressed in terms of the independent variables $x_{n},y_{n},z_{n},t_{n}$ using a Galilean transformation so that fields can be written in terms of $x_{n},y_{n},z_{n},t_{n}$.
+From the Galilean transformation $x_{n}=x-vt,~{}y_{n}=y,~{}z_{n}=z,~{}t_{n}=t$, after partial differentiation, one obtains $\partial/{\partial t_{n}}=\partial/{\partial t}+v\partial/{\partial x}$, $\partial/{\partial x_{n}}=\partial/{\partial x}$.
+So $\partial/\partial t_{n}=v\partial/\partial x_{n}$ at $\partial/\partial t=0$.
+A natural assumption is to consider that time derivative $\partial/\partial t_{n}$ arises from the Galilean boost.
+Now we shall be more general and study the case of the Lorentz coordinatization.
+When the system $S_{n}$ starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed.
+The time $t^{(L)}_{n}$ under the Einstein’s synchronization in the $S_{n}$ frame is readily obtained by introducing the offset factor $x_{n}v/c^{2}$ and substituting $t^{(L)}_{n}=t_{n}-x_{n}v/c^{2}$ in the first order approximation.
+The new time coordinate in the accelerated frame is interpreted by saying that Maxwell’s equations are applicable to the aberration of electrons description.
+After partial differentiation, one obtains $\partial/{\partial t^{(L)}_{n}}=v\partial/{\partial x_{n}}$, $\partial/{\partial x^{(L)}_{n}}=\partial/{\partial x_{n}}+(v/c^{2})\partial/{\partial t_{n}}$.
+So $\partial/\partial t^{(L)}_{n}=v\partial/\partial x^{(L)}_{n}$ neglecting terms of order of $v^{2}/c^{2}$.
+By changing the (four-dimensional) coordinate system, one cannot obtain a physics in which new physical phenomena appear.
+When the situation is described as we have done here, there doesn’t seem to be any paradox at all; it comes out quite natural that separation between space and time is no longer possible in general.
+The argument that the result is paradoxical runs something like this: If we could measure magnetic field $\vec{B}_{n}$ by the usual magnetometer there would indeed be a paradox.
+But we could not do that.
+We already discussed above that magnetic field $\vec{B}_{n}$ leads to change in the transverse momentum of test particles given by $\theta_{a}=p_{\perp}/p=(1-1/\gamma)v/V$.
+If a particle motion velocity is non-relativistic, the binomial expansion yields $\vec{\theta}_{a}=-[V^{2}/(2c^{2})]\vec{V}\times\vec{v}/V^{2}$ and the momentum perturbation presents a relativistic effect.
+So special relativity says that there are no changes in the (test) particle momentum as $V^{2}/c^{2}\to 0$.
+If we look at the usual (i.e.
+classical) magnetometer with test particles, we see no evidence of the magnetic field $B_{n}$ effect.
+Not only does this particular kind of classical magnetometer can not measure the magnetic field $B_{n}$, but if the theory of relativity is correct, any other classical magnetometer, operating on any principle whatsoever, would also appear to demonstrate no evidence of the magnetic field $B_{n}$ effect.
+To understand magnetic field measurements in the accelerated electron source, we have to watch the machinery of the classical magnetometer and see what happens when it is at rest in the accelerated frame.
+Since it is rather difficult, we shall take a simple kind of usual magnetometer.
+It is relatively easy to observe nuclear magnetism through the phenomena of ”nuclear magnetic resonance”.
+A proton resonance apparatus can be used as a proton resonance magnetometer.
+We give a general analysis of this type of classical magnetometer independently of the experimental accuracy.
+In general, a current loop, which has a magnetic moment $\vec{m}$ in the inertial lab frame in which it is at rest, has an electric dipole moment equal to $\vec{d}=\vec{v}\times\vec{m}/c$ when it is moving with velocity $\vec{v}$ relative to the lab frame.
+The magnetic properties of materials are attributed to atomic current loops.
+Then a magnetized body that is accelerated relative to the lab frame should have an electric polarization relative to the lab frame.
+Thus, in the lab frame, the additional potential energy of a moving particle with the proper magnetic moment $\vec{m}$ in the magnetic field $\vec{B}$ and the electrical field $\vec{E}$ becomes $U_{p}=-\vec{m}\cdot\vec{B}-\vec{d}\cdot\vec{E}$.
+We obtained the electrodynamics equations in an accelerated frame using the Galilean transformation (with velocity $-\vec{v}$) of Maxwell’s equations.
+With the Galilean transformation, the magnetic dipole transformation for the case $\vec{d}=0$, is $\vec{m}_{n}=\vec{m}$, $\vec{d}_{n}=-\vec{v}\times\vec{m}/c$.
+As we know from vector algebra, $(\vec{a}\times\vec{b})\cdot\vec{c}=\vec{a}\cdot(\vec{b}\times\vec{c})$; so our terms are also the same as $(-\vec{v}\times\vec{E})\cdot\vec{m}=-\vec{v}\cdot(\vec{E}\times\vec{m})$, $(-\vec{v}\times\vec{m})\cdot\vec{E}=-\vec{v}\cdot(\vec{m}\times\vec{E})$, as evaluated in the accelerated frame.
+Hence we obtain the additional potential energy equal to zero just as we expected.
+We could not measure magnetic field $\vec{B}_{n}$ by usual magnetometer.
+9.4 Aberration of Particles from a Moving Source
+Let us consider what happens if the electron gun, an electron source for a relativistic electron beam, in the initial inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis, Fig.
+39.
+In other words, we consider a so-called active (physical) boost.
+In the case of an active boost, we are dealing with the motion of the same physical system, evolving in time and treated from the point of view of the same reference system.
+The simplest method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the particle source was at rest, i.e.
+we still enforce the clock transport synchronization ( or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest).
+This choice is usually the most convenient one from the viewpoint of the smooth tailoring of the metric tensor.
+This synchronization convention preserves simultaneity and is based on the absolute time (or absolute simultaneity) convention.
+Figure 39: Aberration of electrons from a moving source in an inertial frame. The motion of the electron gun is perpendicular to the optical axis.
+We can obtain the electric field $\vec{E}$ and the magnetic $\vec{B}$ observed in the inertial frame where the particle source is moving with velocity $v$ using the Galilean transformation.
+When we consider the electrodynamics equations under Galilean transformation, there is a question: what is the transformation law for the electromagnetic fields?
+The equivalence of the active and passive pictures is due to the fact that moving system one way is equivalent to moving the coordinate system the other way by an equal amount.
+It is for this reason that fields in the inertial frame before the active Galilean boost are the same as those $\vec{E^{\prime}}$ and $\vec{B^{\prime}}$ in the comoving coordinate system.
+With the Galilean transformation, the field transformation, for the case $\vec{B^{\prime}}=0$, is $\vec{E}=\vec{E^{\prime}}$, $\vec{B}=\vec{v}\times\vec{E^{\prime}}/c$.
+This expression describes first-order (in $v/c$) effects only.
+Here $\vec{v}$ is the velocity vector of the source in the lab frame.
+[…]
+We described the effect of aberration of particles by working only up to the first order $v/c$.
+An elementary explanation of this effect is well-known.
+The aberration of particles in the inertial lab frame can be easily explained in terms of the Galilean transformation of velocities between different reference frames.
+It is important to stress at this point that the dynamical line of arguments discussed here explains what the aberration of particles physically means in the inertial frame.
+The kinematics effects are only an interpretation of the behavior of the electromagnetic fields.
+Now let us return to the magnetic field measurements in the accelerated electron source, Fig.
+39.
+In the lab frame, we are dealing with motion of the electron source, evolving in time and treated from the point of view of the inertial system.
+It is not difficult to see that the peculiarity of the situation is that here the ”source motion” is a real observable effect.
+Indeed, the problem is time-dependent.
+In this situation, the velocity of the source has a physical meaning.
+In the case of a moving electron source, special relativity says that magnetic field $\vec{B}=\vec{v}\times\vec{E}/c$ leads to a change in the transverse momentum of test particles given by $\theta_{a}=p_{\perp}/p=v/V$ and this result is valid also in the non-relativistic limit $V^{2}/c^{2}\to 0$.
+Now we can raise an interesting question.
+Suppose that an inertial observer is measuring the magnetic field with an usual magnetometer.
+Now we know that there are moving charges, so that means that when there is a current in the moving source, the source generates a magnetic field $\vec{B}=\vec{v}\times\vec{E}/c$ which can be measured with a usual magnetometer.
+In order to understand this, we have to watch the machinery of the NMR magnetometer.
+Now that our magnetometer is at rest in the lab frame, that means that the additional potential energy of a proton with the proper magnetic moment $\vec{m}$ in the magnetic field $\vec{B}$ is equal to $-\vec{m}\cdot\vec{B}\neq 0$.
+The moving electron source in the lab frame generates a magnetic field which can be measured with NMR magnetometer.
+We come to the conclusion that the magnetic field generated by a moving source in an inertial frame is a real field in the sense we have defined it.
 10 Earth-Based Setups to Detect the Aberration Phenomena
 10.1 Potential of Earth-Based Electron Microscopes
-Above we discussed the new relativistic particle kinematics for rotating frame of references.
-Since we live on rotating earth frame, difference in relativistic kinematics between rotating and non-rotating frames of references is of practical as well as theoretical significance.
+Above we discussed the new relativistic particle kinematics for rotating frames of reference.
+Since we live on a rotating earth frame, the difference in relativistic kinematics between rotating and non-rotating frames of reference is of practical as well as theoretical significance.
 We propose an aberration of particle experiment using an earth-based particle source.
-Experiments can be explained on the basis of special relativity.
-A correct solution of this problem in the earth-based frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of a relativistic particle beam propagation.
+Experiments can be explained based on special relativity.
+A correct solution to this problem in the earth-based frame requires the use of metric tensor even in first-order experiments.
 We propose an experiment using commercial 200-kV scanning transmission electron microscope (STEM) as a particle source.
 One could use a vertically oriented optical setup.
-We suppose that the electron beam is imaged by a lens to a spot which lies in the image plane.
+We suppose that the electron beam is imaged by a lens to a spot that lies in the image plane.
 Measurement of the spot shift with respect to the optical axis (which is parallel to the zenith-nadir axis) is equivalent to the determination of the angular displacement.
-Due to the high stability, the aberration increment could be observed (1).
+Due to the high stability, the aberration increment could be observed  [^42] The dedicated STEM with high stability was made [47].
+It was developed based on a 200-kV commercial instrument.
+The electron source was installed in the anti-seismic room.
+Temperature fluctuations in the room were within 0.2 ${}^{o}C$.
+The magnetic field in the room was less than $0.2mG$ near the source column.
+To prevent thermal drift, a shield box was placed over the high-voltage generator.
+The column was covered with rubber sheets to reduce the temperature fluctuations.
+The image drift measured during a period of 61 minutes was $\simeq 6~{}\mathrm{nm}$.
+The drift speed and direction during this period were approximately constant.
+The average drift speed was 0.12 nm/min.
+It is substantially smaller than that conventional commercial microscope, in which the drift speed is about 2 nm/min..
 Consider the inertial sun-based reference system.
-In this system there is the earth which rotates around the sun with orbital velocity $v$.
-Developing into powers of $v/c$ we can classify effects for velocities $v\ll c$ as of the first order, of the second order or of higher orders.
-The essentially relativistic effects are of the second (or higher) orders.
+In this system, there is the earth which rotates around the sun with orbital velocity $v$.
+Developing into powers of $v/c$ we can classify effects for velocities $v\ll c$ as of the first order, of the second order, or higher orders.
 Clearly, in the case of the rotating (around the sun) earth-based frame, we consider the small expansion parameter $v/c\simeq 10^{-4}$ neglecting terms of order of $v^{2}/c^{2}$.
 Suppose that an earth-based observer performs an aberration of particle measurement.
-The aberration angle varies with one-year temporal period.
-An approximate formula to express the aberration angle can be found by using Langevin metric in the earth-based frame or by using Wigner rotation theory in the sun-based frame.
+The aberration angle varies with a one-year temporal period.
+An approximate formula to express the aberration angle can be found by using the Langevin metric in the earth-based frame or by using the Wigner rotation theory in the sun-based frame.
 In the aberration of particles, the aberration angle is the apparent angular deviation of the position of the particle source relative to the real location of the source.
-The reference axis in the earth-based frame can be formed by plumb line, which is the most fundamental local earth-based coordinate system.
-When measured with an earth-based particle source, the aberration angle is $\vec{\theta}_{a}=-\vec{\theta}_{w}=-(1-1/\gamma_{p})\vec{V}\times\vec{v}/V^{2}$, where $v$ is the orbital velocity of the earth relative to the sun, $V$ is the particle velocity, and $\gamma_{p}=1/\sqrt{1-V^{2}/c^{2}}$ is the particle relativistic factor.
+The reference axis in the earth-based frame can be formed by a plumb line, which is the most fundamental local earth-based coordinate system.
+When measured with an earth-based particle source, the aberration angle is $\vec{\theta}_{a}=-\vec{\theta}_{w}=-(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$, where $\vec{v}$ is the orbital velocity of the earth relative to the sun, $\vec{V}$ is the velocity of the accelerated electron, and $\gamma=1/\sqrt{1-V^{2}/c^{2}}$ is the particle relativistic factor.
 This means that the angle of Wigner rotation depends on $V^{2}/c^{2}$.
-It should be note that presented formula can be considered only as a first order approximation in $v/c$.
-In the theory of (infinitesimal) Wigner rotation we must also consider the relativistic parameter $V/c$.
+It should be noted that the presented formula can be considered only as a first-order approximation in $v/c$.
+In the theory of (infinitesimal) Wigner rotation, we must also consider the relativistic parameter $V/c$.
 If a particle motion velocity is non-relativistic, the binomial expansion yields $\vec{\theta}_{a}=-[V^{2}/(2c^{2})]\vec{V}\times\vec{v}/V^{2}$.
 Let us calculate the parameters for the STEM setup.
-The Wigner rotation angle parameter (constant of aberration) defines the scale of the aberration increment, and is equal to $\theta_{w}=(1-1/1/\gamma_{p})v/V\simeq 40\mu\mathrm{rad}$.
-Here $\gamma_{p}\simeq 1.39$, $V/c=0.7$, $v/c\simeq 100\mu\mathrm{rad}$.
+The Wigner rotation angle parameter (constant of aberration) defines the scale of the aberration increment and is equal to $\theta_{w}=(1-1/1/\gamma)v/V\simeq 40$ $\mu\mathrm{rad}$.
+Here $\gamma\simeq 1.39$, $V/c=0.7$, $v/c\simeq 100$ $\mu\mathrm{rad}$.
 Earth moves around the sun, with a consequent change of the direction of $\vec{v}$.
@@
 Figure 40: Definitions related observer’s position on the earth.
-A specific aspect of our case study needs further investigations.
-The proposed method of measuring the angle of aberration involves the use of earth-based sources and have a big advantage.
+A specific aspect of our case study needs further investigation.
+The proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage.
 The rotation of the earth on its axis should produce a corresponding shift of the image.
@@
 In principle, records could be taken over a period of one day.
-The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth rotation axis (2).
+The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth’s rotation axis 434343 The constant of diurnal aberration defines the scale of the aberration increment and is equal to $\theta_{d}=(1-1/1/\gamma)v_{e}/V\simeq 0.6$ $\mu\mathrm{rad}$.
+Here $\gamma\simeq 1.39$, $V/c=0.7$, $v_{e}/c\simeq 1,55$ $\mu\mathrm{rad}$ where $v_{e}=0.45$ km/s is the linear velocity of the earth rotation at equator..
 In practice, the measurement procedure is complicated by a number of factors, Fig.
-42.
-A corrections has to be made for the observer position on the earth’s surface and the obliquity of the ecliptic $\delta=23^{o}$.
+40.
+A correction has to be made for the observer’s position on the earth’s surface and the obliquity of the ecliptic $\delta=23^{o}$.
 The image appears to move in an ellipse.
-The shape of the aberration ellipse obviously depends on the observer’s latitude $\phi$.
-For an observer on the equator the ellipse degenerates into a line segment and for an observer at the pole of the earth, the ellipse is a circle.
-At the value $l=2$ mm of the focal length (3), the major axis of the aberration ellipse of the order of 80 ($\cos\delta)$ nm.
-The value of minor axis depends on the observer’s position and is given by 80 ($\cos\delta\sin\phi))$ nm.
-When the latitude is $\phi=35^{o}$, the value of the scale of image shift per hour depends on the local time and various from a value about 10 nm/60 min to about 20 nm/60 min through the day (4).
-8.2 Explanation of the Aberration of Electrons on the Basis of Electrodynamics
-A large number of incorrect expression for the Wigner rotation (it is often called the ”Thomas rotation”) can be found in the literature.
-For those who have already received knowledge about the Wigner rotation from the well-known textbooks, it may be expedient to begin its aberration of particle (theoretical and experimental) development with the microscopic approach.
-The point is that one can compute any relativistic quantity directly from the underlying theories of matter without involving relativity at all.
-Wigner rotation associated with the transformation from the inertial (sun-based) coordinate system to the rotation (earth-base) coordinate system may be regarded as a result of the action of certain force.
-This problem has no relation to the problem of the ether and is already formulated within the framework of special relativity.
-The electromagnetic forces which govern the properties of an emitted electron beam must be affected by (earth-based) source acceleration around the sun in a such a way that they lead to a deviation of the electron transport direction.
-In fact, there is a machinery behind the electron momentum $\vec{p}$ changes when the source and observer (with his measuring instruments) are at rest in the rotating frame.
-Its origin is explained in the framework of the electrodynamics theory.
-We can obtain the electrodynamics equations in rotating reference frame using the Galilean transformation.
-When we consider the electrodynamics equations under Galilean transformation, there is a question: what is the transformation law for the electromagnetic fields $\vec{E}$ and $\vec{B}$?
-The fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame where the particle source is at rest are not the same as those $\vec{E}$ and $\vec{B}$ in the inertial frame before the active Galilean boost.
-We need to find the relation between $\vec{E}_{n}$ and $\vec{B}_{n}$ and $\vec{E}$ and $\vec{B}$.
-To do this, we transform coordinates $(t,x)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$, using a Galilean transformation: we substitute $x_{n}=x-vt$, while leaving time unchanged $t_{n}=t$ into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain Langevin metric $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$.
-We can obtain the electrodynamics equations in rotating frame using the Galilean transformation (with velocity $-v$) of the Maxwell’s equations.
-With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$.
-This expression describe first order (in $v/c$) effects only.
-Here $\vec{v}$ is the velocity vector of the earth in the sun-based frame.
-[…]
-This formula correspond to the result we demonstrated for observation of the non-inertial observer taking advantage of the Wigner rotation theory.
-The theory of special relativity shows that the aberration of particles effect in the rotating earth-based frame, as viewed from the inertial sun-based frame (in the Lorentz coordinatization), presents a kinematic effect of special theory of relativity.
-In this case, when viewed from the sun-based frame, the aberration image shift in the earth-based frame is regulated by the Wigner rotation.
-According to the Wigner rotation theory, two observers with different trajectories have different 3-spaces.
-In contrast, the accepted in previous literature incorrect assumption that an earth-based observer and an inertial (e.g.
-sun-based) observer have common 3-space has always been considered obvious.
-It is important to stress at this point that the dynamical line of arguments discussed here explains what the Wigner rotation physically means in the rotating frame.
-The relativistic kinematics effects in the rotating frame, as viewed from the inertial observer, is only an interpretation of the behavior of the electromagnetic fields.
+The shape of the aberration ellipse depends on the observer’s latitude $\phi$.
+For an observer on the equator, the ellipse degenerates into a line segment, and for an observer at the pole of the earth, the ellipse is a circle.
+At the value $l=2$ mm of the focal length  [^44] Electron-optical system of the STEM contains an electron gun and several magnetic lenses, stacked vertically to form a lens column.
+The illumination system of the instrument comprises the electron gun, together with two condenser lenses that focus the electrons onto the spacemen.
+Electrons entering the lens column appear to come from a virtual source.
+The first condenser lens (C1) is a strong magnetic lens, with a focal length $f_{1}$ of typically 2 mm.
+Using a virtual electron source as its object, C1 produces a real image.
+Because the lens is located $D_{1}=15$ cm below the object (virtual source), its object distance is 15 cm $\gg f_{1}$ and the image distance $\simeq f_{1}$.
+The second condenser lens (C2) is a weak magnetic lens ($f_{2}\simeq 10$ cm) that provides little magnification but allows the diameter of illumination at the specimen to vary continuously over a wide range.
+The distance between lens centers is typically 10 cm.
+Spacemen is located $D_{3}=25$ cm below the C2., the major axis of the aberration ellipse of the order of 80 ($\cos\delta)$ nm.
+The value of the minor axis depends on the observer’s position and is given by 80 ($\cos\delta\sin\phi))$ nm.
+When the latitude is $\phi=35^{o}$, the value of the scale of image shift per hour depends on the local time and varies from a value of about 10 nm/60 min to about 20 nm/60 min through the day  [^45] Experimental results (see Fig.
+44 p.167 in [47]) shows that the image shift is quite close to the theoretical prediction $\simeq 10~{}\mathrm{nm}$.
+It is important to note that we used the typical focal length (2 mm) in our estimations.
+The details of the optical setup may influence the image shift.
+.
 10.2 Thought Experiment using Point-Like Source at Optical Frequencies
-It is generally believed that the theory of relativity appears to conform to the phenomenon of stellar aberration claiming it is a consequence of the motion of observers relative to light sources.
-In particular, the conventional method used to explain the aberration phenomena is based on the belief that there is no aberration for the starlight radiated from the star moving with the same velocity as the earth.
-However, aberration shift, as inferred from astronomical observations, behaves asymmetrically.
-It should be stressed that it is the telescope and not the star that must change its velocity (relative to the fixed stars) to cause aberration.
 We want now to discuss a simple scaling model for the stellar aberration, Fig.
-43.
+41.
 Let us consider the scale transformation of the source linear dimension and the distance between the source and observer.
 We will obtain a condition for optical similarity between the aberration of light from a distant star and from the earth-based source.
-This scale transformation may help in the initial design of the earth-based setup using incandescent lamp as a point-like source instead of a star.
+This scale transformation may help in the initial design of the earth-based setup using the incandescent lamp as a source instead of a star.
 Figure 41: Scaling model for the stellar aberration.
@@
 We know (without using our knowledge of special relativity) what would happen from the astronomical observations.
-Without looking at anything external to the earth, one could determine the speed of the earth around the sun by means of aberration of light measurements.
-The peculiarity of point-like sources is that radiation emitted at one instant form a sphere around the source and the measuring instrument always influences the measured radiation.
-Source field diffraction can be divided into categories - the Fresnel (near-zone) diffraction and Fraunhofer ( or telescopic) diffraction.
-If the Fraunhofer (or equivalently, far zone) approximation $z>D^{2}/\lambda$ is satisfied, then quadratic phase factor is approximately unity over the detection aperture.
-Here $D$ is the aperture width, $\lambda$ is the wavelength of the radiation, and $z$ is the distance between source and observer.
-In the far zone a point source produces in front of pupil detection effectively a plane wave.
-The plane wave fronts of starlight entering the telescope are imaged by the lens to a diffraction spot which lies in the focal plane.
-It worth noting that we consider the divergence of the transmitted radiation, $\lambda/D$, that is relatively small compared to the aberration angle, $\theta_{a}=v/c$.
+Without looking at anything external to the earth, one could determine the speed of the earth around the sun using aberration of light measurements.
+A star ( or an incandescent lamp) is a system of elementary statistically independent point sources.
+Point-source field diffraction can be divided into categories - the Fresnel (near-zone) diffraction and Fraunhofer ( or telescopic) diffraction.
+If the Fraunhofer (or equivalently, far zone) approximation $z>D^{2}/\lambda$ is satisfied, then the quadratic phase factor is approximately unity over the detection aperture.
+Here $D$ is the aperture width, $\lambda$ is the wavelength of the radiation, and $z$ is the distance between the source and observer.
+In the far zone, a point source produces in front of pupil detection effectively a plane wave.
+The plane wavefronts of starlight entering the telescope are imaged by the lens to a diffraction spot that lies in the focal plane.
+It is worth noting that we consider the divergence of the transmitted radiation, $\lambda/D$, that is relatively small compared to the aberration angle, $\theta_{a}=v/c$.
 An incandescent lamp is a completely incoherent source.
-The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem (5).
-In our case of interest, imaging system is situated in the far-zone of the source (6).
-Under these circumstances, the Van Cittert-Zernike theorem takes its simplest form.
+The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34].
+In our case of interest, imaging system is situated in the far zone $D^{2}/(\lambda z)<1$.
 It is the source linear dimension $d$ (source diameter) that determines the coherent area of the observed radiation $(z\lambda/d)^{2}$.
-The condition for neglecting the transverse size of the source, $d$, formulated as the requirement that the change of the correlation function along the aperture be less than unity, that is $d\ll D$.
-Let us estimate the parameters for the earth-based setup using incandescent lamp.
-We want to describe some thought - but possible - experiment.
-At optical frequencies, the conditions required for validity of the Fraunhofer approximation can be sever ones.
+The condition for neglecting the transverse size of the source, $d$, formulated as the requirement that the change of the correlation function along the aperture be less than unity, that is $d<D$ 464646 In all cases of practical interest, telescopes are situated in the far zone of the distant stars.
+If the far zone approximation $z>d$ is satisfied, then the Van Cittert-Zernike theorem takes its simplest form.
+For example, the coherence area of the light emitted by the circular incoherent source is $A_{c}=4\lambda^{2}z^{2}/(\pi d^{2})$.
+The minimal coherent area of light observed on the earth’s surface from the nearest star has a diameter of about $\lambda z/d\simeq$ 10 m.
+This diameter is larger than a typical telescope aperture $D$.
+In this situation, a star can be considered as a point source and a star image is a point spread function in the image plane of a telescope.
+The star consists of an extended collection of independent ”point” sources.
+The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$ and far zone condition $z>D^{2}/\lambda$ for the elementary ”point” source is automatically satisfied.
+Our scaling model is based on the far zone approximation.
+We have only to change the distance and the source diameter.
+With the limitation of the small parameters assumption ($Dd/(z\lambda)<1$, $D^{2}/(z\lambda)<1$) the proposed scaling method gives $d<D$..
+Let us estimate the parameters for the earth-based setup using the incandescent lamp.
+We want to describe some thought - but possible - experiments.
+At optical frequencies, the conditions required for the validity of the Fraunhofer approximation can be severe ones.
 For example, at a wavelength of $\lambda=0.3$ $\mu$m and an aperture width of $D=3$ cm, the observation distance must satisfy $z>3000$ m.
 Obviously, the required conditions are met in a number of important problems.
-No one has ever done such experiment, but we know (without using our knowledge of special relativity) what would happen from the astronomical observations.
-The purpose of this thought experiment is to suggest that without looking at anything external to the earth, one could determine the speed of the earth around the sun by means of aberration of light measurements.
-The aberration of light problem in the accelerated frame demonstrates the essential asymmetry between the non-inertial and inertial observers.
-There is a certain degree of analogy between the aberration of light and the Sagnac effect.
-Clearly, it is possible to detect a rotating velocity without looking outside using Sagnac phenomenon.
-A close look at the physics of these two subjects shows things which are common to these phenomena: In both situations these are experiments in the first order in $v/c$.
-Indeed, the aberration of light is the geometric effect in our ordinary space which is (similar to the Sagnac phenomenon) closely associated with the first order (crossed) term in the Langevin metric.
-8.4 Simple Explanation of the Aberration of Light Effect in Rotating Frame
-Above (see Chapter 5) we considered the aberration of light problem in the accelerated frame.
-The aberration of light effect is not easy to calculate in frame of reference attending rotation.
-In this case, we used a metric tensor to obtain the electrodynamics equations in rotating reference frame.
-The presented approach uses the Fourier transform method.
-As one of the consequences of the crossed term in the wave equation, we found the group velocity that is responsible for the aberration.
-Indeed, what is usually considered as an aberration is, in fact, a deviation of the energy transport direction.
-In the Section 8.2 we presented a very simple explanation of the aberration of relativistic particles in the rotating frame on the basis of electrodynamics.
-To find the aberration increment, we simply used Galilean transformation law for the electric and magnetic fields.
-Can we not look at the aberration of light effect in the same way?
-We have an expression for the energy flow vector of the electromagnetic field.
-This vector, $\vec{S}=c\vec{E}\times\vec{B}/(4\pi)$, is called ”Poynting’s vector”.
-It tell us the rate at which the field energy moves around in space.
-In order to detect the aberration of light effect inside the accelerated frame, it is obvious that some coordinate system with reference direction is needed.
-We determine the reference directions perpendicular to the line motion in the inertial frame and in the accelerated frame.
-The motion of the aberrated light beam are assumed for simplicity, to lie in the same plane and the angular position of the aberrated beam is described by one angle (as illustrated in Fig.
-32, Fig.
-33 ).
-The analysis is much simplified if we treat separately the case of a radiated beam with its $\vec{E}$-vector parallel to the ”plane of incidence” (that is $xz$-plane) and the case of a radiated beam with the $\vec{E}$-vector perpendicular to the $xz$-plane.
-We will carry through the analysis for an incoming beam polarized perpendicular to the plane of incidence, but the principle is the same for both.
-So we take that $\vec{E}$, has only $y$- component.
-For a magnetic field we get $\vec{B}=\vec{e}_{z}\times\vec{E}$, where $\vec{e}_{z}$ is the direction of Poynting vector $\vec{S}$ in the inertial frame.
-In the previous chapters we demonstrated that we must match the (4D) coordinates in the accelerational and inertial parts of the trajectory.
-For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the Galilean boost.
-We can obtain the electrodynamics equations in rotating frame using the Galilean transformation (with velocity $-v$) of the Maxwell’s equations.
-The fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame where the emitter is at rest are not the same as those $\vec{E}$ and $\vec{B}$ in the inertial frame before the active Galilean boost.
-With the Galilean transformation, the field transformation is $\vec{E}_{n}=\vec{E}-\vec{v}\times\vec{B}/c$, $\vec{B}_{n}=\vec{B}-\vec{v}\times\vec{E}/c$.
-This expression describe first order (in $v/c$) effects only.
-Here $\vec{v}=v\vec{e}_{x}$ is the velocity vector of the earth in the sun-based frame.
-If the $\vec{E}$-vector perpendicular to the $xz$-plane, the magnetic field in the inertial frame has only a $x$-component and we have immediately $\vec{E}_{n}=\vec{E}=\vec{e}_{y}E$, $\vec{B}_{n}=(\vec{e}_{z}\times\vec{e}_{y})E-(\vec{e}_{x}\times\vec{e}_{y})vE/c=-\vec{e}_{x}E-\vec{e}_{z}vE/c$.
-Therefore there is a Poynting vector, $c\vec{E}_{n}\times\vec{B}_{n}/(4\pi)=[-(\vec{e}_{y}\times\vec{e}_{x})-(\vec{e}_{y}\times\vec{e}_{z})v/c]cE^{2}/(4\pi)=[\vec{e}_{z}-\vec{e}_{x}(v/c)]cE^{2}/(4\pi)$.
-Then, the radiated beam is propagated at the angle $-v/c$ with respect to the $z_{n}$-axis yielding the phenomenon of the aberration of light in the accelerated frame.
-We have found that we get the same result whether we analyze the energy transport direction using electrodynamics equations in the rotating frame, or using Galilean transformation law for electromagnetic fields.
-In the first instance, the equations state that transmitted light beam with finite transverse size moves along $x$-direction with group velocity $d\omega/dk_{x}=-v$, in the second, the effect of light aberration is understood as a change in the direction of Poynting vector.
-The two approaches give the same result.
-In this section we demonstrate that the direction of group velocity and the direction of Poynting vector are actually two sides of the same coin.
-When the observer are at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be modified.
-One can easily show this by transforming coordinates $(t,x,y,z)$ that would be coordinates of an inertial (sun-based) observer moving with velocity $-v$ with respect to the earth-based observer.
-From the Galilean transformation $x_{n}=x-vt,~{}y_{n}=y,~{}z_{n}=z,~{}t_{n}=t$, after partial differentiation, one obtains $\partial/{\partial t_{n}}=\partial/{\partial t}+v\partial/{\partial x}$, $\partial/{\partial x_{n}}=\partial/{\partial x}$.
-Acceleration has an effect on the field equations.
-In fact, the wave equation transforms into Eq.
-(16).
-In any measurement of the angular displacement of transmitted (through aperture) light beam, one will have to consider a finite radiation beam size.
-We have, then, a general method for finding the group velocity of transmitted light beam.
-We can solve the Eq.
-(16) by using Fourier transform method.
-The new terms that have to be put into the wave equation due to use Galilean transformation leads to prediction of the group velocity, $-v$, along the $x$-axis.
-Our problem now is to work out the magnetic field of the transmitted light beam in terms of those of the incident plane wave.
-How can we do that?
-The magnetic field $\vec{B}_{n}$ satisfies Maxwell’s equation ×$\vec{\nabla}\times\vec{B}_{n}=(\partial/{\partial t}+v\partial/{\partial x})\vec{E}_{n}/c$, in the rotating frame.
-We are very specific and wright out explicitly the new component: $-\partial B_{z}/{\partial x}=(v/c)\partial E_{y}/{\partial x}$.
-If we integrate this equation with respect to $x$ across transmitted light beam, we conclude that $\vec{B}_{n}=-\vec{e}_{x}E-\vec{e}_{z}vE/c$.
-This last result is just what we got by a field transformation argument.
-This is a good point to make a general remark about the physical reality of the energy flow (Poynting) vector.
-In any physically realizable aberration measurement, one will have to consider a finite radiation beam size.
-It is always hiddenly assumed that the detector for the direction of the radiation is an energy propagation detector and the size of the detector aperture is sufficiently large compared with the radiation beam size.
-The direction of the energy transport (and, hence, the direction of the Poynting vector) has an exact objective meaning i.e.
-convention-invariant.
-However, there is a common mistake made in electrodynamics connected with the energy transport direction in the case of a plane wave.
-It is impossible to know the energy transport direction when one has deal with plane wave.
+No one has ever done such an experiment, but we know (without using our knowledge of special relativity) what would happen from the astronomical observations.
+The purpose of this thought experiment is to suggest that without looking at anything external to the earth, one could determine the speed of the earth around the sun using aberration of light measurements.
 10.3 Working Way of the Setup Using Incoherent Source at Optical Frequencies
-It should be note that Fraunhofer diffraction patterns can be observed at distances much closer than implied by equation $z>D^{2}/\lambda$ provided the aperture is illuminated by a spherical wave converging towards the observer, or if a positive lens is properly situated between the observer and the aperture.
+It should be noted that Fraunhofer diffraction patterns can be observed at distances much closer than implied by equation $z>D^{2}/\lambda$ provided the aperture is illuminated by a spherical wave converging towards the observer, or if a positive lens is properly situated between the observer and the aperture.
 A working way of the earth-based optical setup to detect the aberration of light is schematically shown in Fig.
-44.
+42.
 This is the well-known two-lens image formation scheme.
-This scheme allows for magnification by changing the focal distance of the second lens but for simplicity, in the following we will assume that the two focal distances are the same (i.e.
+This scheme allows for magnification by changing the focal distance of the second lens but for simplicity, in the following, we will assume that the two focal distances are the same (i.e.
 we consider 1:1 imaging).
-This two-lens setup is usually employed for image-processing purposes, as it can be better used for image modification compared to the single-lens system.
+This two-lens setup is usually employed for image-processing purposes, as it can be better used for image modification than the single-lens system.
 Given the two-lens setup discussed above, we consider the relatively simple problem of characterization of point source radiation in the image plane.
-We assume that the two lenses in Fig.
+We assume the two lenses in Fig.
 42 are identical.
@@
 Within an elementary source volume ($\lambda^{3}$) there is an enormous number of atoms.
-Semi-classical theory is treating these atoms as coherent, interacting radiating dipoles.
-The induced macroscopic dipole moment in an elementary source leads to the classical electromagnetic radiation.
-In the optics for such ”point” source emission, the fields are described classically at the level of Maxwell’s equations and the emitting medium is treated (as a ensemble of atoms) by quantum mechanics.
-The atomic polarization induced by the radiation field appears as a driving term in Maxwell’s equations and sustains the oscillations.
-With the help of modern lithography it is not difficult to produce an unresolved point source at optical wavelengths and let sufficiently bright radiation from it (7).
+In the optics for such ”point” source emission, the fields are described classically at the level of Maxwell’s equations and the emitting medium is treated (as an ensemble of atoms) by quantum mechanics.
+With the help of modern lithography, it is not difficult to produce an unresolved point source at optical wavelengths and let sufficiently bright radiation from it  [^47] It may be possible to use the masked spatially incoherent UV light source.
+One could use the light source onto a (micron diameter) hole of a specially designed mask.
+The mask should be placed at the distance $l<<d$, where $d$ is the transverse size of the spatially incoherent UV light source..
 Consider a diffracting aperture that is circular, and let the radius of the aperture be $a$.
 The physical properties of the lens can be combined in a single number $f$ called the focal length.
-We assume that the image produced by a diffraction-limited optical system (i.e.
+We assume that the image is produced by a diffraction-limited optical system (i.e.
 a system that is free from aberrations).
@@
 One can take advantage of well-known resolution criteria like Rayleigh criteria i.e.
-$0.6\lambda/NA=2\mu$m.
+$20.6\lambda/NA=$ 2 $\mu$m.
 One could use a vertically oriented optical setup.
-Constant of aberration defines the scale of the aberration increment, and is equal to $v/c=100\mu$rad.
+The constant of aberration defines the scale of the aberration increment, and is equal to $v/c=100$ $\mu$rad.
 Measurement of the spot shift with respect to the optical axis (which is parallel to the nadir direction) is equivalent to the determination of the angular displacement.
-As already pointed out, the proposed method of measuring the angle of aberration involves the use of earth-based sources and have a big advantage.
+As already pointed out, the proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage.
 The rotation of the earth on its axis should produce a corresponding shift of the image.
@@
 The major axis of the ellipse is about 14 $\mu$m.
-The value of minor axis depends on the observer’s position and is given by $14(\sin\phi)$ $\mu$m (8).
+The value of the minor axis depends on the observer’s position and is given by $14(\sin\phi)$ $\mu$m  [^48] The important question is that of light detection.
+The small image size form rather sever restriction on the use of the CCD as a detector.
+In fact, for the case considered numerically image size is near 2 $\mu$m.
+One possibility to overcome this practical difficulty is photon detection with an MCP (microchannel plate photomultiplier) detector.
+It may be possible to use a (commercially available) precision 2D translation stage with a resolution of 0.5 $\mu$m to control the motion of a hole in the opaque screen.
+The aperture diameter in this example may be set to 1 micron or smaller.
+The screen is fixed horizontally on the translation stage which is placed in the image plane.
+The MCP-PMT may realize the measurement of (transmitted through a hole) photons.
+The proposed design has the advantage of device compactness.
+The height of our proposed optical configuration is about one meter..
 Figure 42: A working way of the earth-based optical setup to detect the aberration of light. Two-lens image formation setup using the earth-based incoherent source at optical frequency.
 We derived our results under the assumption that the order of magnitude of the dimensions of the source is about $\lambda$.
-It should be note that the resolution of proposed optical techniques is not limited by size of incoherent source.
-Radiation field generated by an incoherent source can be seen as a linear superposition of fields of individual elementary point sources.
+It should be noted that the resolution of proposed optical techniques is not limited by the size of the incoherent source.
+The radiation field generated by an incoherent source can be seen as a linear superposition of fields of individual elementary point sources.
 The image of an elementary point source is a point spread function.
-Optical system using incoherent light could be completely described in therm of convolution integral relating the object distribution to the image distribution.
+An optical system using incoherent light could be completely described in therm of convolution integral relating the object distribution to the image distribution.
 In such a description, the optical system is characterized by its impulse response, i.e.
@@
 The rounding or softening of the source’s edges is characteristic of the convolution (”smoothing”) integral.
-It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single point source characteristics like oscillatory image shift with a period of one day.
-This argument gives reason why our technique can be extended to allow for the reconstruction of the angle of aberration from a measurement of image shift of incoherent source.
+It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single-point source characteristics like oscillatory image shift within a period of one day.
+This argument gives the reason why our technique can be extended to allow for the reconstruction of the angle of aberration from a measurement of image shift of an incoherent source.
 The edge object has become useful in aberration shift detection.
@@
 The edge width of the source is strictly related to the practical resolution achievable.
-For our a particular example, the edge width should be about $1\mu$m .
-8.6 Absent of the Aberration of Light from the Earth-based Laser Source
-Let us make some comments on the aberration of light experiments using an earth-based source.
-In regard to light aberration one should differentiate between that from the incoherent source and that from the laser source.
-One could naively expect that the rotation of the earth around the sun produces aberration in an amount large enough to be taken into account in precise observation work using laser as a earth-based light source.
-Let us stress that the aberration of light cannot be measured using laser.
-In order to understand this phenomenon, we consider the simple case when a laser resonator is equipped with plane mirrors.
+For our particular example, the edge width should be about $1\mu$m.
+11 The Aberration of Light from a Laser Source
+11.1 Theory of a Laser with Optical Resonator
+Suppose that an earth-based observer performs an aberration of light measurement using a laser as a light source.
+Let us stress that the earth’s orbital motion cannot be measured using a laser.
+To understand this phenomenon, we consider the simple case when a laser resonator is equipped with plane mirrors.
 For oscillation to occur, the total loss in power due to diffraction and reflection loss must be less than the power gained by travel through the active medium.
-Thus diffraction loss is expected to be an important factor, both in determining the start-oscillation condition, and in determining the distribution of energy in the interferometer during oscillation.
+Thus diffraction loss is expected to be an important factor, both in determining the start-oscillation condition and in determining the distribution of energy in the interferometer during oscillation.
 The purpose of our study is to investigate the effect of diffraction on the electromagnetic field in a Fabry-Perot interferometer in free space.
-The conclusion can be applied equally well to gaseous lasers provided the interferometer immersed in the active medium.
+The conclusion can be applied equally well to gaseous lasers provided the interferometer is immersed in the active medium.
 The electric field $\vec{E}$ of an electromagnetic wave in a passive resonator satisfies the equation $\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$.
-For simplicity, we shall assume that the field in the resonator to be circularly polarized.
+For simplicity, we shall assume that the field in the resonator is circularly polarized.
 Such an assumption does not violate the generality of the analysis because the polarization degeneracy takes place in the system under study.
 The electric field vector in the plane resonator is presented as an oscillation superposition with different longitudinal wave numbers, $E_{x}+iE_{y}=\sum_{m}\widetilde{E}_{m}(x,y,t)\exp(-i\omega_{m}t)\sin(K_{m}z+\delta)$, where $K_{m}=m\pi/l-i/(nl)$, $\omega_{m}=m\pi c/l$, $\delta=i/n$, $n$ is the (complex) refractive index of the mirror matter ($|n|\gg 1)$, $l$ is the distance between the mirrors, and $m$ is an integer number ($m\gg 1$).
-It is evident that this expression satisfies Leontovich’s boundary conduction (9) on the mirror surface when $z=0$ and $z=l$.
+It is evident that this expression satisfies Leontovich’s boundary conduction  [^49] We have presented an example of introducing approximate boundary conditions.
+These approximate boundary conditions were derived by Leontovich.
+They are valid on the surface of a material having a large value of the modulus of the refractive index.
+An extended theoretical treatment of Leontovich’s boundary conditions appears in the monograph [48].
+on the mirror surface when $z=0$ and $z=l$.
 We assume the field change per one resonator pass to be small i.e.
@@
 Substituting the ansatz in the wave equation we get $\nabla_{\perp}\Phi+2i\omega_{m}[\Lambda/c^{2}+2/(nl)]\Phi=0$.
-In a plane resonator the diffraction effects at the edges of the mirrors can be taken into account if one considers the space between the mirrors as a waveguide, and uses diffraction theory at the open end of waveguide.
-If almost an integer number of half waves is laid along the length between the mirrors then at open resonator end we may put boundary conditions on the amplitude function $\Phi(x,y)$, equivalent by its effect to the open waveguide end action (so-called impedance boundary conditions of resonance type (10)): $\Phi+(1+i)\beta_{0}\sqrt{cl/(4\omega_{m})}\partial\Phi/\partial\zeta=0$, where $\zeta$ is the direction of the normal to the imaginary side surface of the resonator, and the parameter $\beta_{0}=0.824$ (11).
-The boundary condition do not depend on the polarization which means that polarization degeneracy takes place in the system under study.
-Therefore, the problem concerning the open plane resonator excitation by means of equivalent boundary condition can be reduced to the more usual classical formulation of the problem of closed resonator excitation.
 […]
-The rigorous results of the three-dimensional theory of a plane Fabry-Perot resonator have shown that a radiation in the resonator may be presented by a set of modes.
+[…]
+The rigorous results of the three-dimensional theory of a plane Fabry-Perot resonator have shown that radiation in the resonator may be presented by a set of modes.
 Each mode is characterized by the decrement and an eigenfunction of the field amplitude distribution along the mirror surface.
@@
 We designate a normal mode of a circular plane mirrors as a $TEM_{\nu j}$ mode, with $\nu$ denoting the order of angular variations ( the angular variation are sinusoidal in form) and $j$ denoting the order of radial variation.
-The lowest order, of $TEM_{00}$ mode, we designate as the dominate mode for circular plane mirrors.
-Suppose that an earth-based observer performs an aberration of light measurement using laser as a light source.
-One could use a vertically oriented optical setup.
-Above, we demonstrate that acceleration has an effect on the field equations.
+The lowest order, of $TEM_{00}$ mode, we designate as the dominant mode for circular plane mirrors.
+11.2 Absent of the Aberration of Light from the Earth-Based Laser Source
+Above, we demonstrate that acceleration affects the field equations.
 The electric field $\vec{E}$ of an electromagnetic wave into the resonator satisfies the wave equation transformed into Eq.
 (17), where coordinates and time are transformed according to a Galilean transformation.
-At first glance, measurement of the diffraction losses of the dominate mode is equivalent to the determination of the angular displacement.
-Figure 45: Absent ot the aberration of light from the earth-based laser source.
-The electromagnetic wave travels in the resonator forward and back reflecting from the mirrors.
-The wavefront orientation in the Lorentz coordinatization for the laser source is presented here for sketch simplification.
-It should be note, however, that there is no influence of the difference between the Langevin and Minkowski metrics on the parameters of a laser with optical resonator.
-In our previous discussion of the Langevin metric, we learned that after the Galilean transformation of the wave equation we obtain the complicated anisotropic equation Eq.
+At first glance, the measurement of the diffraction losses of the dominate mode is equivalent to the determination of the angular displacement.
+It should be noted, however, that there is no influence of the difference between the Langevin and Minkowski metrics on the parameters of a laser with an optical resonator.
+In our previous discussion of the Langevin metric, we learned that after the Galilean transformation of the wave equation, we obtain the complicated anisotropic equation Eq.
 (17).
@@
 It is enough to assume the Lorentz covariance of the electrodynamics theory involved in the optical resonator calculations.
-We can make a mathematical trick for the solution of the electrodynamics field equations with anisotropic terms: in order to eliminate these terms in the transformed field equations, we make a change variables.
-The overall combination of the Langevin metric and variable changes actually yields the Minkowski metric in the accelerated frame, but in this context this transformation are only to be understood as useful mathematical device, which allow one to solve the electrodynamics problem in the case of the Langevin metric with minimal effort.
+We can make a mathematical trick for the solution of the electrodynamics field equations with anisotropic terms: to eliminate these terms in the transformed field equations, we make a change of variables.
+The overall combination of the Langevin metric and variable changes yields the Minkowski metric in the accelerated frame, but in this context, this transformation is only to be understood as a useful mathematical device, which allows one to solve the electrodynamics problem in the case of the Langevin metric with minimal effort.
 We state that the variable changes performed above have no intrinsic meaning - their meaning only being assigned by a convention.
-It is clear that the convention-independent results of calculations are precisely the same in the new variables.
+The convention-independent results of calculations are precisely the same in the new variables.
 By changing the (four-dimensional) coordinate system, one cannot obtain a physics in which new physical phenomena appear.
@@
 round-trip or two beams) measurements.
-Let us analyze the aberration of light transmitted from the laser resonator which is at rest in the accelerated frame.
-When the measuring data is analyzed, the accelerated observer finds that the deviation of the energy transport direction is absent.
+Let us now analyze the aberration of light transmitted from the laser resonator which is at rest in the accelerated frame.
+When the measuring data is analyzed, the accelerated observer finds that the deviation of the energy transport direction is absent, Fig.
+43.
 It could be said that the crossed term in the Langevin metric, which generates aberration, cancels during the (round-trip-to-round-trip) evolution of the radiation in the optical resonator.
-We would like to emphasize a very important difference between laser and incoherent source.
-The radiation of incoherent source clearly depends on an initial conditions and there is influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer.
-In contrast, when we are dealing with optical resonator the (3-space) boundary conditions define the energy propagation direction.
-Now let us consider another problem in optical resonator operation that we should briefly discuss.
-This has to do with the optical experiments of second order.
-It is important to remark that a close examination of all experiments inside the uniformly moving accelerated $S_{n}$ frame, however, shows that all optical phenomena in the optical resonator appeared to be independent of the uniform motion relative to the fixed stars.
-Imaging that there are two identical laser sources.
-We set up the lasers such that light is travels between mirrors parallel to the $z$-axis.
-Let us consider the case when the first source is at rest in an inertial frame and the second source is accelerated from rest up to velocity $v$ along the $x$-axis (i.e.
-transverse to the direction of optical beam).
-Suppose that an observer in the inertial frame performs a measurement of the frequency $\omega_{m}=m\pi c/l$ of the $TEM_{00}$ mode.
-In order to measure the frequency, the observer has to specify frequency standard.
-Suppose he took an ordinary atom, which had natural frequency $\omega_{o}$ at rest.
-Let us consider the opposite situation.
-Suppose that an observer in the accelerated frame performs a measurement of the $TEM_{00}$ mode.
-It is also assumed that the same atom at rest in the accelerated frame is used as frequency standard.
-In the case of Langevin metric Eq.(15), the speed of light propagated between the mirrors in the transverse $z_{n}$-axis direction cannot be equal $c$ anymore.
-Using Eq.(15) we obtain $(dz_{n}/dt_{n})^{2}=c^{2}(1-v^{2}/c^{2})$.
-Therefore, the frequency of the $TEM_{00}$ mode is $\omega_{m}=m\pi c\sqrt{1-v^{2}/c^{2}}/l$.
-Because $t_{n}=t$ (note that coordinate clock in the accelerated frame , although fixed in this frame , read the same time as the clocks in the inertial frame) and $\omega_{m}=m\pi c/l$ has the physical meaning of the frequency of the dominate mode at rest in the inertial frame, this implies that the frequency of the $TEM_{00}$ mode at rest in the accelerated frame is red shifted compared to $\omega_{m}$ in the original inertial frame.
-Taking into account the Langevin metric we obtain the time dilation of a physical clock at rest in the accelerated frame as compared to physical clock in the inertial frame.
-In the next chapter we shall discuss the asymmetry in the experiments involved the measurement of quantities of the second order in $v/c$.
-An observer on the accelerated frame cannot observe the time dilation effect using optical resonator.
-In order to detect the red shift of the $TEM_{00}$ mode at rest in the accelerated frame, it is obvious that some reference frequency is needed.
-The point is that the natural frequency of the atom observed in the accelerated frame is $\omega_{0}\sqrt{1-v^{2}/c^{2}}$.
-We have found that the accelerated observer will measure the same mode frequency.
-It is not hard to understand this result.
-It is clear that an observer performs the measurement of standing wave mode of the optical resonator.
-If we analyze the geometry of the situation, we find that from the standing wave measurements in the optical resonator we can only extract information about two-way speed of light.
-In other words, the measurement of the two-way speed of light is universal.
-Although the optical cavity we have been describing seems to be quite different from a simple electromagnetic model for the light-clock of special relativity, the two systems are, of course closely related.
-The discussion of the behavior of an accelerated light clock is presented in the next chapter.
-8.7 Bibliography and Notes
-1.
-The dedicated STEM with high stability was made [39].
-It was developed based on a 200-kV commercial instrument.
-The electron source was installed in the anti-seismic room.
-Temperature fluctuations in the room were within 0.2 ${}^{o}C$.
-The magnetic field in the room was less than $0.2mG$ near the source column.
-To prevent thermal drift, a shield box was placed over the high-voltage generator.
-The column was covered with rubber sheets to reduce the temperature fluctuations.
-The image drift measured during a period of 61 minutes was $\simeq 6~{}\mathrm{nm}$.
-The drift speed and direction during this period were approximately constant.
-The average drift speed was 0.12 nm/min.
-It is substantially smaller than that conventional commercial microscope, in which the drift speed is about 2 nm/min.
-2.
-The rotation of the earth on its axis produces aberration in an amount large enough to be taken into account in precise observation work using electron microscope as a particle source.
-The Wigner rotation angle parameter (constant of diurnal aberration) defines the scale of the aberration increment, and is equal to $\theta_{w}=(1-1/1/\gamma_{p})v_{e}/V\simeq 0.6\mu\mathrm{rad}$.
-Here $\gamma_{p}\simeq 1.39$, $V/c=0.7$, $v_{e}/c\simeq 1,55\mu\mathrm{rad}$ where $v_{e}=0.45$km/s is the linear velocity of the earth rotation at equator.
-3.
-Electron-optical system of the STEM contains an electron gun and several magnetic lenses, stacked vertically to form a lens column.
-The illumination system of the instrument comprises the electron gun, together with two condenser lenses that focus the electrons onto the spacemen.
-Electrons entering the lens column appear to come from a virtual source.
-The first condenser lens (C1) is a strong magnetic lens, with a focal length $f_{1}$ of typically 2 mm.
-Using a virtual electron source as its object, C1 produces a real image.
-Because the lens is located $D_{1}=15$ cm bellow the object (virtual source), its object distance is 15 cm $\gg f_{1}$ and the image distance $\simeq f_{1}$.
-The second condenser lens (C2) is a weak magnetic lens ($f_{2}\simeq 10$ cm) that provides little magnification but allows the diameter of illumination at the specimen to varied continuously over a wide range.
-Distance between lens centers typically 10 cm.
-Spacemen is located $D_{3}=25$ cm bellow the C2.
-4.
-Experimental results (see Fig.
-44 p.167 in [39]) shows that the image shift is quite close to the theoretical prediction $\simeq 10~{}\mathrm{nm}$.
-It is important to note that we used the typical focal length (2 mm) in our estimations.
-It is also assumed that the second condenser lens provides no magnification ($M\simeq 1$).
-The details of the optical setup may influence the image shift.
-5.
-The original optical source (such as a star or an incandescent lamp) consists of an extended collection of independent radiators.
-Accordingly, it is of some special interest to know precisely how mutual intensity propagates away from an incoherent source.
-The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34].
-6.
-In all cases of practical interest, telescopes are situated in the far zone of the distant stars.
-If the far zone approximation $z>2d^{2}/\lambda$ is satisfied, then the Van Cittert-Zernike theorem takes its simplest form.
-For example, the coherence area of the light emitted by the circular incoherent source of diameter $d$ at distance $z$ is $A_{c}=4\lambda^{2}z^{2}/(\pi d^{2})$.
-The maximal coherent area of light observed on the earth’s surface from a nearest star has diameter of about 10 m.
-In other words, any star can be considered as a point source and a star image is actually a point spread function in the image plane of a telescope.
-The star consists of an extended collection of independent ”point” sources.
-The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$.
-It is obvious that the star diameter $d$ is much larger than a telescope aperture $D$ and that far zone condition $z>D^{2}/\lambda$ for elementary ”point” source is automatically satisfied.
-Our scaling model is based on the far zone approximation.
-We have only to change the distance and the source diameter.
-With the limitation of the small far zone parameters ($d^{2}/(z\lambda)<1$, $D^{2}/(z\lambda)<1$) assumption, the proposed scaling method gives $d\ll D$.
-7.
-It may be possible to use the masked spatially incoherent UV light source.
-One could use the light source onto (micron diameter) hole of specially designed mask.
-The mask should be placed at the distance $l<<d$, where $d$ is the transverse size of the spatially incoherent UV light source.
-8.
-The important question is that of light detection.
-The small image size form rather sever restriction on the use of the CCD as a detector.
-In fact, for the case considered numerically image size is near 2 $\mu$m.
-One possibility to overcome this practical difficulty is the photon detection with an MCP (micro channel plate photo multiplier) detector.
-It may be possible to use a (commercially available) precision 2D translation stage with a resolution of 0.5 $\mu$m to controlling the motion of a hole in the opaque screen.
-The aperture diameter in this example may be set to 1 micron or smaller.
-The screen is fixed horizontally on the translation stage which is placed in the image plane.
-The MCP-PMT may realize the measurement of (transmitted through a hole) photons.
-The proposed design has the advantage of device compactness.
-The height of our proposed optical configuration is about one meter.
-9.
-We have presented an example of introducing an approximate boundary conditions.
-This approximate boundary conditions were derived by Leontovich.
-They are valid on the surface of a material having a large value of the modulus of the refractive index.
-An extended theoretical treatment of Leontovich’s boundary conditions appears in the monograph [40].
-10.
-Analytical method is based on the introduction of particular, complex boundary condition, which are called impedance boundary conditions [41].
-Once these boundary conditions for the field are formulated on the virtual side surface of the open plane resonator, the original problem, which is an open resonator excitation, is simplified to the that of a closed resonator.
-11. $\beta_{0}$ turns out to be related with one of the most famous mathematical functions, the Riemann zeta function $\zeta$. In fact, it is given by $-\zeta(1/2)/\sqrt{\pi}$, see [41].
+Figure 43: Aberration of light in an accelerated frame of reference. It is assumed that the accelerated laser operates in the steady-state regime.
+Figure 44: Aberration of light from a ”plane wave” emitter in an accelerated frame of reference.
+According to the difference between the ”plane wave” emitter and the laser source, there is a remarkable prediction in the theory of the aberration of light.
+Namely, if the ”plane wave” emitter was at rest relative to the fixed stars and this source started from rest to motion relative to the fixed stars, then the apparent angular position of the ”plane wave” emitter seen in the accelerated frame would jump by angle $-v/c$, Fig.
+44.
+To understand the aberration of light, we will have to use the special theory of relativity and not be satisfied with classical theory.
+How can such a difference between a ”plane wave” emitter and a laser source come about?
+Let us first ask: ”What happens to accelerated optical resonator?” When we were dealing with optical resonators we used the electrodynamics boundary conditions in 3-space, and they defined the energy propagation direction along the optical axis (i.e.
+perpendicular to the mirror surface).
+According to special relativity, an acceleration does not change the direction of the optical axis (i.e.
+mirrors orientation) with respect to the accelerated frame (3-space) axes.
+An observer in the rest frame of a laser source finds that the direction of the energy transport is independent of the motion with respect to the fixed stars, Fig.
+43.
+According to special relativity, accelerated and inertial observers have common absolute space-time but have different metrics and, consequently, different 3-spaces.
+When we are dealing with optical resonator the (3-space) boundary conditions define the energy propagation direction.
+In contrast, the radiation of an incoherent source depends on initial conditions and there is an influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer.
+The solution is presented in terms of the absolute time coordinate, hence the wavefront of the emitted light beam is perpendicular to the direction $z_{n}$ after the acceleration.
+Using this initial condition and the Langevin metric in a frame of reference attending acceleration enables us to explain the aberration of light from a ”plane wave” emitter, Fig.
+44.
+11.3 A Moving Laser Source in an Inertial Frame of Reference
+Let us consider the radiation emitted by the accelerated laser source in the inertial frame.
+The aberration of light effect can be described within the standard special relativity taking advantage of the relativity of simultaneity.
+When one has a transversely (i.e.
+parallel to the mirrors) moving laser there is a deviation of the energy transport for light transmitted from the optical resonator.
+In the first order approximation in $v/c$, the inertial observer would find that angular displacement is $\theta_{a}=2v/c$, Fig.
+45.
+In regard to light aberration one should differentiate between that from the ”plane wave” emitter and that from the laser source.
+Suppose we have a ”plane wave” emitter.
+We consider the case when the ”plane wave” emitter is accelerated from rest up to velocity $v$ in the direction perpendicular to its optical axis.
+Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the light transport measurement.
+Then how does the light beam from the moving ”plane wave” emitter look?
+The inertial observer would find that angular displacement is equal to $\theta_{a}=v/c$, Fig.
+46.
+Figure 45: Aberration of light from a moving laser source in the initial inertial frame of reference.
+The motion of the laser resonator is parallel to the mirrors.
+It is assumed that the accelerated laser operates in the steady-state regime.
+Figure 46: Aberration of light from a moving ”plane wave” emitter in the initial inertial frame of reference.
+The first thing we would say about inertial frame measurements is that there is an intuitively plausible way to understand the aberration of light from a moving ”plane wave” emitter.
+An elementary explanation of this effect is well-known.
+This phenomenon is fully understandable in terms of the transformation of velocities between different reference frames.
+The aberration of light in the initial inertial frame can also be easily explained on the basis of the corpuscular theory of light.
+This is plausible if one keeps in mind that a light signal represents a certain amount of electromagnetic energy.
+Energy, like mass, is a quantity that is conserved, so that a light signal resembles, in many aspects, material particles.
+Therefore, we should expect that group velocities of light signals obey the same addition theorem for particle velocities.
+A closer treatment based on the wave theory of light confirms this expectation.
+In the case of the moving laser source, intuition would seem to tell us that the aberration increment would be equal to $\theta_{a}=v/c$.
+But the special relativity says that there is an extra factor 2.
+The commonly accepted derivation of the aberration of light effect does not account for the mixture of positions and time in the accelerated frame.
+It is incorrectly assumed that accelerated and inertial observers have a common 3-space.
+In contrast to this, as viewed from the initial inertial frame, the coordinate axes of the accelerated frame are rotated through the angle $v/c$ with respect to the coordinate axes of the inertial frame.
 12 Special Relativity and the Reciprocal Symmetry
-In this chapter we will collect a number of useful facts concerning the space-time measurements made by different observers moving with respect to each other.
-The situation with measurements in special relativity is not symmetrical with respect to the change of the reference frames.
-The fact that in the real process of transmission to an accelerated reference frame (i.e in the process of an observer accelerating with respect to the fixed stars) the observer will experience the pseudo-gravitational force, is not considered in textbooks.
 12.1 Pseudo-gravitation Field and the Langevin Metric
-This section concentrates on the nature of special relativistic effects.
-Dynamics, based on the field equations, is usually hidden in the language of the relativistic kinematics.
+This chapter concentrates on the nature of special relativistic effects.
+Dynamics, based on the field equations, is usually hidden in the language of relativistic kinematics.
 The metric structure of space-time is only an interpretation of the behavior of the dynamical matter fields in the view of different observers.
-It is worth remarking that the absent of a dynamical explanation for the Langevin metric in special relativity has disturbed some physicists.
+It is worth remarking that the absence of a dynamical explanation for the Langevin metric in special relativity has disturbed some physicists.
 A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer.
-Langevin metric associated with the transformation from the inertial coordinate system to the accelerated (with respect to the fixed stars) coordinate system may be regarded as a result of the action of certain force.
-This problem formulated within the framework of general relativity.
-Note that it is not correct to say that this is explanation of the effects discussed in the last four chapters.
-One of the aims in this section is to show that one can describe the non-inertial frames in the pseudo-Euclidean space-time using the language of general relativity.
+Langevin metric associated with the transformation from the inertial coordinate system to the accelerated (with respect to the fixed stars) coordinate system may be regarded as a result of the action of a certain force.
+This problem is formulated within the framework of general relativity.
+Note that it is not correct to say that this is an explanation of the effects discussed in the last chapters.
+One of the aims of this chapter is to show that one can describe the non-inertial frames in the pseudo-Euclidean space-time using the language of general relativity.
 Suppose that the reference frame $S$ is at rest with respect to the fixed stars.
@@
 After the acceleration, the system of inertia $S$ is (reciprocally) moving in the direction of $x_{n}$-axis with the velocity $-v$ relative to $S_{n}$.
-When the system $S_{n}$ starts moving with constant velocity it will be inertial frame of reference.
-At first glance, we have the symmetry (reciprocity) between the inertial frame $S$ and inertial frame $S_{n}$.
-According to principle of covariance (i.e.
-principle of reciprocity), accelerated (with respect to the fixed stars) observer should obtain Minkowski metric $ds^{2}=c^{2}dt_{n}^{2}-dx_{n}^{2}$.
-This metric is reciprocal to the metric of the inertial observer $ds^{2}=c^{2}dt^{2}-dx^{2}$.
-Where does the asymmetry comes from?
+When the system $S_{n}$ starts moving with constant velocity it will be the inertial frame of reference.
+At first glance, we have the symmetry (reciprocity) between the inertial frame $S$ and the inertial frame $S_{n}$.
+According to the principle of covariance (i.e.
+principle of reciprocity), accelerated (with respect to the fixed stars) observer should obtain Minkowski metric $ds^{2}=c^{2}dt_{n}^{2}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$.
+This metric is reciprocal to the metric of the inertial observer $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
+Where does the asymmetry come from?
 Equivalence principle analysis does not need any tricks to account for the asymmetry between an inertial frame and an accelerated (with respect to the fixed stars) frame.
@@
 By a ”pseudo-gravitational field”, we mean an apparent field (not a real gravitational field) that acts on all objects proportionally to their mass; by ”uniform” we mean that the force felt by each object is independent of its position.
-This is basic content of equivalence principle.
+This is the basic content of the equivalence principle.
 The principle of equivalence can be applied to solve non-inertial kinematic problems with dynamics methods.
@@
 When the system $S$ starts moving with constant velocity the pseudo-gravitational field is zero.
-For the pseudo-gravitational potential difference between the two frames we get after some calculation $\phi=(\partial\phi/\partial x_{n})gT^{2}/2=-v^{2}/2$.
+For the pseudo-gravitational potential difference between the two frames, we get after some calculation $\phi=(\partial\phi/\partial x_{n})gT^{2}/2=-v^{2}/2$.
 It is assumed that $\phi=0$ at $x_{n}=x$.
@@
 Nevertheless, what general relativity does say about real gravitational fields does hold in the restrictive sense for pseudo-gravitational fields.
-One thing we need here is that time runs slower as we descend into the potential well of a pseudo force field.
+One thing we need here is that time runs slower as we descend into the potential well of a pseudo-force field.
 We can use that fact to our advantage when analyzing the metric in the non-inertial frame.
 The equivalence principle implies that gravity can shift the frequency of an electromagnetic wave, and cause clocks to run slow.
-The light frequency increases with increase absolute value of the potential of the gravitational field.
+The light frequency increases with the increase absolute value of the potential of the gravitational field.
 If a pulse of light emitted at a point where the gravitational potential is $\phi_{1}$, has at that point the frequency $\omega_{1}$, then upon arriving at a point where the potential is $\phi_{2}$, it will have a frequency measured in units of the proper time at that point equal to $\omega_{2}=\omega_{1}[1+(\phi_{1}-\phi_{2})/c^{2}]$.
@@
 Thus physical time elapses the more slowly the smaller the gravitation potential at a given point in a space, i.e.
-the lager the absolute value.
-If of two identical clocks is placed in a gravitation field for some time, the clock which has been in the field will thereafter appear to be slow.
+the larger the absolute value.
+If two identical clocks are placed in a gravitation field for some time, the clock that has been in the field will thereafter appear to be slow.
 For clock at rest in the accelerated frame we have $dt^{(p)}=dt_{n}(1+\phi/c^{2})=dt_{n}[1-v^{2}/(2c^{2})]$.
-The fact that the inertial observer always turns out to have aged more than the non-inertial one, in the equivalence principle analysis is elementary consequence of the slowly down of the rate of the processes in the pseudo-gravitational potential, which can be shown by means of simple algebraic calculations.
-Furthermore it is important to point out that the phenomenon of the relativity o simultaneity can be understood in terms of dynamical consideration.
+The fact that the inertial observer always turns out to have aged more than the non-inertial one, in the equivalence principle analysis is an elementary consequence of the slowdown of the rate of the processes in the pseudo-gravitational potential, which can be shown using simple algebraic calculations.
+Furthermore, it is important to point out that the phenomenon of the relativity of simultaneity can be understood in terms of dynamical consideration.
 The time-offset relation can be interpreted as the accumulated time difference between two spatially separated clocks because of pseudo-gravity experienced by the accelerated observer.
 This suggests that the clock at higher gravitational potential (placed along the direction of acceleration) runs faster.
-For the pseudo-gravitational potential difference between two points along the $x_{n}$-axis we get $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$.
+For the pseudo-gravitational potential difference between two points along the $x_{n}$-axis, we get $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$.
 When the system $S$ starts moving with constant velocity the gradient of potential in the system $S_{n}$ is zero.
-For accumulated time difference between two spatially separated clocks we have $t^{(p)}(1)-t^{(p)}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$.
+For accumulated time difference between two spatially separated clocks, we have $t^{(p)}(1)-t^{(p)}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$.
 The next thing we must investigate is the origin of the Lorentz deformation.
@@
 This is a purely kinematic effect, involving no forces.
-We come to the conclusion that the various relativistic effects are not independent of each other, but that these effects can all be understood through a general equivalence principle.
-We have now all quantities we wanted.
+We conclude that the various relativistic effects are not independent of each other, but that these effects can all be understood through a general equivalence principle.
+We have now all the quantities we wanted.
 Let us put them all together into the expression for the new space-time variables.
@@
 The $|dx_{n}^{\mathrm{(p)}}|$ is nothing, but the physical distance between two points of 3-space.
-In new physical coordinates we obtain interval in the following (Langevin) form $ds^{2}=(cdt_{n}^{\mathrm{(p)}})^{2}$ $(dx_{n}^{\mathrm{(p)}})^{2}$ $=(1-v^{2}/c^{2})c^{2}dt^{2}_{n}-2vdx_{n}dt_{n}-dx_{n}^{2}$.
-The previous derivation of the Langevin metric in the accelerated frame from pseudo-gravitation field includes one delicate point.
+The interval in physically measurable differentials is a diagonal quantity.
+The physically measurable differentials $dt_{n}^{\mathrm{(p)}}$ and $dx_{n}^{\mathrm{(p)}}$ are linear combinations of the coordinate time $dt_{n}$ and the coordinate differential $dx_{n}$, the coefficients of these combinations are built up of the components of the metric tensor.
+In new physical coordinates we obtain interval in the following form $ds^{2}=(cdt_{n}^{\mathrm{(p)}})^{2}$ $(dx_{n}^{\mathrm{(p)}})^{2}$ $=(1-v^{2}/c^{2})c^{2}dt^{2}_{n}-2vdx_{n}dt_{n}-dx_{n}^{2}$.
+The previous derivation of the Langevin metric in the accelerated frame from the pseudo-gravitation field includes one delicate point.
 Metric applies to physical laws, not to physical facts.
-We interpret the Langevin metric to mean that the law of electrodynamics is expressed by equation that have the form Eq.(16) in the accelerated frame.
-By a physical facts in this context we mean, for example, the aberration of light radiated by a single ”plane wave” emitter in the accelerated frame.
-The electrodynamics equation needs to be integrated with initial condition for the radiation wavefront.
+We interpret the Langevin metric to mean that the law of electrodynamics is expressed by an equation that has the form Eq.(17) in the accelerated frame.
+By physical facts in this context we mean, for example, the aberration of light radiated by a single ”plane wave” emitter in the accelerated frame.
+The electrodynamics equation needs to be integrated with the initial condition for the radiation wavefront.
 After the inverse Galilean boost, the coordinates transform as $x_{n}=x-vt,t_{n}=t$, so that the wavefront of the emitted light beam is perpendicular to the vertical direction $z_{n}$ after the acceleration, Fig.
@@
 The Langevin metric, together with this initial condition, describes the aberration of light effect in the accelerated frame.
-Although both observers consider themselves as at rest, they both predict that the Maxwell’s equations continue to hold in the inertial frame and the Minkowski metric is not applicable in the accelerated frame.
-But observers have totally different explanation for this difference of metrics.
-Inertial observer says that physical processes in the accelerated frame describes by the Langevin metric due to the active Galilean boost.
-In other words, the analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard pseudo-Euclidean geometry of space-time.
-Accelerated observer, on the other hand, says that Langevin metric in his frame can be understood in terms of dynamical consideration.
 12.2 Equivalence Principle and the Gravitational Physics
-We should make one further remark about the equivalence principle (1).
-The widely held opinion that such a principle is the basis of the general relativity is not entirely correct.
+We should make one further remark about the equivalence principle  [^52] For a general discussion of the equivalence principle we suggest reading the paper [50]..
+The widely held opinion that such a principle is the basis of general relativity is not entirely correct.
 The gravitational field is not simply a force that is universal and has the same action on all bodies.
@@
 The first thing we must understand is the Riemann tensor, since it is this tensor that is the gravitational field: if it vanishes (and only in this case) a field does not exist.
-Actually this argument is already fully sufficient to show that the equivalence principle in the usual formulation does not hold.
+Actually, this argument is already fully sufficient to show that the equivalence principle in the usual formulation does not hold.
 The equivalence principle is valid for mechanical processes and does not hold for electrodynamics processes.
-This means that by means of measurements within the a system it is possible to establish whether the system is inertial or falls freely in a homogeneous gravitational field.
+This means that by means of measurements within the system it is possible to establish whether the system is inertial or falls freely in a homogeneous gravitational field.
 In a homogeneous gravity field, a free charge is uniformly accelerated with respect to an inertial frame, and therefore the charge will radiate, and radiation will occur in both strong and weak fields.
 In the uniformly accelerated frame of reference, the free charge is not accelerated with respect to the inertial frame of reference.
-But in then follows in accordance with electrodynamics that it can never radiate, and therefore the equivalence principle is indeed violated .
-With the special and general relativity we can discuss the case of earth-based experiments where pseudo-gravity and the real gravity are inextricably mixed.
-With this understanding, we next need to describe the gravitational fields near the sun due to the sun mass itself.
+But in accordance with electrodynamics, it can never radiate, and therefore the equivalence principle is indeed violated.
+With special and general relativity we can discuss the case of earth-based experiments where pseudo-gravity and real gravity are inextricably mixed.
+With this understanding, we next need to describe the gravitational fields near the sun due to the sun’s mass itself.
 Let us consider the simplest instance of a transformation from an inertial frame, in which the space-time is Minkowskian, to the earth-based frame.
-Newton gravity is known to be valid when gravitation fields are to weak to produce velocities near the speed of light.
+Newton’s gravity is known to be valid when gravitation fields are too weak to produce velocities near the speed of light.
 In other words, the gravitational potential energy of a particle is much less than its rest-mass energy.
-In such situation, general relativity must make the same predictions as Newton gravity.
-For weak gravitational field the ordinary Newtonian potential $\phi$ completely determines the metric, which has the form $ds^{2}=(1+2\phi/c^{2})c^{2}dt^{2}-(1-2\phi/c^{2})(dx^{2}+dy^{2}+dz^{2})$.
-The sign of $\phi$ is chosen negative, so that far from the source of Mass $M$, we have $\phi=-GM/r$.
+In such a situation, general relativity must make the same predictions as Newton gravity.
+For weak gravitational field, the ordinary Newtonian potential $\phi$ completely determines the metric, which has the form $ds^{2}=(1+2\phi/c^{2})c^{2}dt^{2}-(1-2\phi/c^{2})(dx^{2}+dy^{2}+dz^{2})$.
+The sign of $\phi$ is chosen negatively so that far from the source of Mass $M$, we have $\phi=-GM/r$.
 Now the condition above that the field be weak means that $|m\phi|<<mc^{2}$, so that $|\phi/c^{2}|<<1$.
-This metric is really only correct to first order in $\phi/c^{2}$.
+This metric is only correct to the first order in $\phi/c^{2}$.
 We say that space-time becomes flat asymptotically at large distances from the source.
@@
 Where $\vec{v}$ is the velocity of the earth, and $(d\vec{r})^{2}$ is the square of the coordinate distance.
-The absence of equivalence of the equations in an accelerated reference system and in an inertial frame with a static gravitational field is obvious, since in the first case the non-diagonal metric coefficients will be functions of the time, whereas in the second case the metric exhibits diagonal properties and does not depend on the time.
-Therefore, in the first case the charge at rest will radiate, whereas in the second case the charge at rest will not radiate.
+The absence of equivalence of the equations in an accelerated reference system and in an inertial frame with a static gravitational field is obvious since in the first case the non-diagonal metric coefficients will be functions of the time, whereas in the second case the metric exhibits diagonal properties and does not depend on the time.
+Therefore, in the first case, the charge at rest will radiate, whereas in the second case, the charge at rest will not radiate.
 It is for this reason that one cannot deduce from the observations the validity of the equivalence principle as a physical principle.
-The relation between gravity and inertia is a subject that needs to be understood more deeply.
 Now let us return to observations of an earth-based observer.
 One additional expression for the invariant interval is needed: $ds^{2}=(1+2\phi/c^{2}-v^{2}/c^{2})d(ct)^{2}-2(\vec{v}\cdot d\vec{r})dt-(1-2\phi/c^{2})(d\vec{r})^{2}$.
-To the order of calculations, this result is a simple superposition of the Langevin metric with the metric in the sun centered inertial frame for the weak external gravitational field of the sun.
+To the order of calculations, this result is a simple superposition of the Langevin metric with the metric in the sun-centered inertial frame for the weak external gravitational field of the sun.
 Both orbit speed and gravitational field dilate time and contract space.
@@
 An excellent illustration of the absence of reciprocity is the time dilation effect.
-Consider the course of time in two inertial reference frames, one of which, $S$, will be considered to be at rest, while another one, $S_{n}$, will accelerate with respect to the first one up to velocity $v$ along the $x$- axis.
-As we already discussed, the transformation from the inertial frame to the accelerated frame apparently first used by Langevin.
-The mathematical idea that are used in the literature of time dilation effect is the smooth tailoring of the metric tensor.
+Consider the course of time in two inertial reference frames, one of which, $S$, will be considered to be at rest with respect to the fixed stars, while another one, $S_{n}$, will accelerate with respect to the first one up to velocity $v$ along the $x$-axis.
+As we already discussed, the transformation from the inertial frame to the accelerated frame was apparently first used by Langevin.
+The mathematical idea used in the literature on the time dilation effect is the smooth tailoring of the metric tensor.
 The transformation to the accelerated frame, with coordinates denoted by ”n”, has the Galilean form, $t_{n}=t$, $x_{n}=x-vt$.
-After the Galilean transformation, the metric of non-inertial frame goes to into the Langevin metric of the accelerated inertial frame $S_{n}$ in a continuous manner (see Section 9.6 for more detail).
+After the Galilean transformation, the metric of the non-inertial frame goes to the Langevin metric of the accelerated inertial frame $S_{n}$ in a continuous manner (see Chapter 5 for more detail).
 In terms of these new coordinates the invariant interval given by Eq.
@@
 With the cross-term, $dt_{n}dx_{n}$, clocks in the accelerated frame are not synchronized in the standard way (Einstein synchronization) of sending a light signal back and forth between the clocks.
-Suppose we have two identical physical clocks at one and the same point of the inertial reference system $S$.
+Suppose we have two identical physical clocks at the same point of the inertial reference system $S$.
 Consider their readings to coincide at the initial moment $t=0$.
@@
 the time, shown by the first physical clock will coincide with the time $t=t_{n}$ of coordinate clock.
-Since the second physical clock is at rest, its reading of its proper time is $d\tau_{n}=\sqrt{1-v^{2}/c^{2}}dt_{n}$.
+Since the second physical clock is at rest, the reading of its proper time is $d\tau_{n}=\sqrt{1-v^{2}/c^{2}}dt_{n}$.
 This demonstrates the time dilation of a physical clock at rest in the accelerated frame as compared to physical clocks in the inertial frame.
-We have found that the slowly down of the accelerated physical clock does not depend on the choice of reference system, in which this effect is measured as it must be.
+We have found that the slowdown of the accelerated physical clock does not depend on the choice of reference system, in which this effect is measured.
 Now we go on to consider another situation, a very practical one.
@@
 Now frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ accelerate with respect to the inertial frame $S$, up to the same speed $v$ but move in opposite directions.
-The problem is complete symmetrical in the inertial frame $S$ and no difference in the time dilation of a physical clock at rest in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ (as compared to physical clock in the inertial frame $S$) can exist.
+The problem is completely symmetrical in the inertial frame $S$ and no difference in the time dilation of a physical clock at rest in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ (as compared to physical clock in the inertial frame $S$) can exist.
 The following simple analysis confirms this conclusion.
@@
 The situation seems paradoxical.
-Now the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ move relative to each other with a non zero speed.
-We see an interesting thing: the presence of some non zero relative velocity cannot be a cause for the effect of time dilation.
+Now the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ move relative to each other with a non-zero speed.
+We see an interesting thing: the presence of some non-zero relative velocity cannot be a cause for the effect of time dilation.
 Above we discussed that one can describe the non-inertial frames using the language of general relativity.
-Time dilation effect associated with the transformation from inertial frame to the accelerated frame may be regarded as a result of the action of certain force.
-A resolution of the symmetry paradox identifies a pseudo-gravitational field within the systems $S_{n}^{(1)}$ and $S_{n}^{(2)}$ as the agency of time dilation effect.
-Proper time clocks having the same accelerational histories really do run at same rates and yield identical measurement results when at rest in different inertial systems.
+The time dilation effect associated with the transformation from the inertial frame to the accelerated frame may be regarded as a result of the action of a certain force.
+A resolution of the symmetry paradox identifies a pseudo-gravitational field within the systems $S_{n}^{(1)}$ and $S_{n}^{(2)}$ as the agency of the time dilation effect.
+Proper time clocks having the same accelerational histories do run at the same rates and yield identical measurement results when at rest in different inertial systems.
 Now we go on to consider another situation.
@@
 Taking into account the Minkowski metric we obtain $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt$ as it must be.
-In our calculations we find that mathematics for relating proper times is extremely simple.
+In our calculations, we find that mathematics for relating proper times is extremely simple.
 That is not too surprising.
 We used Galilean boosts and the Galilean addition of velocities.
-We shall discuss the time dilation effect for moving reference (atomic) clock further in the Chapter 11.
 12.4 Light-Clock. Observations of the Noninertial Observer
-9.4.1 Introductory remarks
-In order to understand the slowing of the clock in an accelerated system, we have to watch the machinery of the clock and see what happens when it is moving.
+To understand the slowing of the clock in an accelerated system, we have to watch the machinery of the clock and see what happens when it is moving.
 We shall try to understand the effect in a very simple case.
-A clock which we shall call ”light-clock” is a rod (meter stick) with a mirror at each end.
+A clock which we shall call ”light clock” is a rod (meter stick) with a mirror at each end.
 The tick of time is the reflection of light in the mirror.
-Let us now consider the parallel and orthogonal to motion orientations of the light clock.
+Let us now consider the parallel and orthogonal motion orientations of the light clock.
 The equivalence of all clocks underlines the theory of relativity and we shall show that irrespective of the orientation of the light path with respect to the motion of the light-clock, the clock scores the same time.
-Two light-clocks are shown in Fig.
-46
+Two light clocks are shown in Fig.
+47
 Figure 47: Parallel and orthogonal light-clock setups.
-9.4.2 Parallel Light-Clock. Langevin Metric
-Now we have to watch the machinery of the physical clock and see what happens to the accelerated light-clock.
+Now we have to watch the machinery of the physical clock and see what happens to the accelerated light clock.
 We describe the clock based on the observations made by an observer in the same accelerated frame as the clock.
-In a first step, we set up the light clock such that light travels between mirrors parallel to the direction of fixed stars motion.
-In the case of Langevin metric, the speed of light radiated by the accelerated emitter cannot be equal $c$ anymore.
-If $ds$ is the infinitesimal displacement along the world line of a beam of light, then $ds^{2}=0$ and using Eq.(15) we obtain $c^{2}=(dx_{n}/dt_{n}+v)^{2}$.
+In the first step, we set up the light clock such that light travels between mirrors parallel to the direction of fixed stars motion.
+In the case of the Langevin metric, the speed of light radiated by the accelerated emitter cannot be equal to $c$ anymore.
+If $ds$ is the infinitesimal displacement along the world line of a beam of light, then $ds^{2}=0$, and using Eq.(15) we obtain $c^{2}=(dx_{n}/dt_{n}+v)^{2}$.
 This means that in the accelerated reference system of coordinates $(ct_{n},x_{n})$ the velocity of light radiated by the accelerated emitter parallel to the $x_{n}$-axis, is $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction.
@@
 Therefore, the time interval between the sending and reception of the light signal is $\Delta t=t_{1}+t_{2}=2l_{n}/[c(1-v^{2}/c^{2})]$.
-One must first determine the length of physical rods.
+One must first determine the length of the physical rods.
 For the measurement of distances between fixed points on the accelerated frame we shall use physical measuring rods of the same kind as those used in the inertial frame, but now at rest relative to the accelerated frame.
-In connection with the process of measuring distances in accelerated frames a problem arises which did not occur in inertial frames.
-For according to the special relativity, no rigid bodies can exist, since they would provide a means of transmitting signals with velocity greater than velocity of light.
 With regard to spatial distances, the interpretive principle is that $\sqrt{-ds^{2}}$ gives the length of an infinitesimal rod whose endpoints are simultaneous according to standard simultaneity in the rod’s rest frame.
 Taking into account the Langevin metric Eq.(15), we obtain at $dt_{n}=0$ that $dx_{n}=dx$, i.e.
-the length of physical rod in the inertial frame will coincide with the coordinate distance in the accelerated frame.
+the length of the physical rod in the inertial frame will coincide with the coordinate distance in the accelerated frame.
 When we apply this rule to physical rods that are at rest in the accelerated frame, we encounter the complication that $dt_{n}=0$ does not automatically correspond to standard simultaneity in the accelerated frame.
@@
 The spatial geometry in the accelerated frame is characterized by the differential spatial line element on it.
-It should be note that spatial geometry has a conventional character.
+It should be noted that spatial geometry has a conventional character.
 However, the proper spatial line element $dl_{n}$ is of special significance, since it defines a coordinate-independent spatial geometry in the accelerated frame.
@@
 With the help of Eq.(15), we find $dl_{n}^{2}=dx_{n}^{2}/(1-v^{2}/c^{2})$, which determines the spatial geometry in the accelerated frame.
-This equation is asymmetric with respect to the lengths $dx_{n}=dl$ and $dl_{n}$, since it relates the physical length $dl$ of the rod in the inertial frame $S$ to the physical length $dl_{n}$ in the accelerated frame $S_{n}$.
-Our analysis based on using standard measuring rods in an accelerated frame to measure it geometrical properties shows that coordinate distance $dx_{n}$ has physical length $dx_{n}/\sqrt{1-v^{2}/c^{2}}$.
-With reference to the accelerated frame it is interpreted as an effect of acceleration with respect to the fixed stars, i.e.
+This equation is asymmetric with respect to the lengths $dx_{n}=dl$ and $dl_{n}$ since it relates the physical length $dl$ of the rod in the inertial frame $S$ to the physical length $dl_{n}$ in the accelerated frame $S_{n}$.
+Our analysis based on using standard measuring rods in an accelerated frame to measure its geometrical properties shows that coordinate distance $dx_{n}$ has physical length $dx_{n}/\sqrt{1-v^{2}/c^{2}}$.
+It is interpreted as an effect of acceleration with respect to the fixed stars, i.e.
 the geometry of space is non-Euclidean in an accelerated frame.
-This shows that the physical meter stick on which the mirrors are mounted undergoes compress on in the direction of fixed stars motion.
-Thus we have ascertained the asymmetry of meter stick’s length directly from the metric.
-Therefore, the time interval between the sending and reception of the light signal in the parallel light-clock in the accelerated frame is $\Delta t_{n}=(2l/c)/\sqrt{1-v^{2}/c^{2}}$.
-Because $t_{n}=t$ and $\Delta t=2l/c$ has the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame.
-9.4.3 Orthogonal Light-Clock. Langevin Metric
-Let us consider a light-clock accelerated up to velocity $v$ transverse to the direction of optical pulse.
-In other words, we set up light-clock such that the effect of a meter stick contraction is not present.
-We describe the orthogonal light-clock based on the observations made by an observer in the same accelerated frame as the light-clock.
-If $ds$ is the infinitesimal displacement along the world line of a beam of light, then $ds^{2}=0$ and using Eq.(15) we obtain $(dz_{n}/dt_{n})^{2}=c^{2}(1-v^{2}/c^{2})$.
+This shows that the physical meter stick on which the mirrors are mounted undergoes compression in the direction of fixed stars motion.
+Thus we have ascertained the asymmetry of the meter stick’s length directly from the metric.
+Therefore, the time interval between the sending and reception of the light signal in the parallel light clock in the accelerated frame is $\Delta t_{n}=(2l/c)/\sqrt{1-v^{2}/c^{2}}$.
+Because $t_{n}=t$ and $\Delta t=2l/c$ have the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame.
+Let us consider a light-clock accelerated up to velocity $v$ transverse to the direction of the optical pulse.
+In other words, we set up light clock such that the effect of a meter stick contraction is not present.
+We describe the orthogonal light clock based on the observations made by an observer in the same accelerated frame as the light clock.
+If $ds$ is the infinitesimal displacement along the world line of a beam of light, then $ds^{2}=0$, and using Eq.(15) we obtain $(dz_{n}/dt_{n})^{2}=c^{2}(1-v^{2}/c^{2})$.
 Therefore, the time interval between the sending and reception of the light signal is $\Delta t_{n}=2l/[c\sqrt{1-v^{2}/c^{2}}]$.
-Because $t_{n}=t$ and $\Delta t=2l/c$ has the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame.
-The theory of relativity shows us that irrespective of the orientation of the light path with respect to the motion of the light-clock, the clock scores the same time.
+Because $t_{n}=t$ and $\Delta t=2l/c$ have the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame.
 12.5 Metrics in the Accelerated Frame
-Let us follow out the consequences of assuming that the inertial frame has no accelerational history and the (observer) space-time metric at rest has diagonal form.
+Let us follow the consequences of assuming that the inertial frame has no accelerational history and the (observer) space-time metric at rest has a diagonal form.
 We start with the two emitters in the inertial frame.
 The first emitter is at rest and the second is moving with velocity $v$.
-As a result of Galilean boost, we obtained the metric of the accelerated emitter in the inertial frame Eq.
+As a result of the Galilean boost, we obtained the metric of the accelerated emitter in the inertial frame Eq.
 (11).
-In this coordinatization, we describe the non-accelerated light source using diagonal metric, $ds^{2}=c^{2}dt^{2}-dx^{2}$.
+In this coordinatization, we describe the non-accelerated light source using diagonal metric, $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
 Let us now return to our consideration of the accelerated frame.
 In the accelerated frame we have the reciprocal velocity.
-Second emitter is at rest and the first is moving with velocity $-v$.
+The second emitter is at rest and the first is moving with velocity $-v$.
 Now what about reciprocity in physics?
-According to the conventional theory, an acceleration does not spoil the motional symmetry between the accelerated frame and the inertial frame and the accelerated observer should obtain diagonal metric for the second emitter at rest in the accelerated frame.
+According to the conventional theory, an acceleration does not spoil the motional symmetry between the accelerated frame and the inertial frame and the accelerated observer should obtain a diagonal metric for the second emitter at rest in the accelerated frame.
 The principle of reciprocity implies that the metric Eq.
@@
 (11) of the moving emitter in the inertial frame.
-According to principle of reciprocity, the difference is only that the sign of velocity $v$.
+According to the principle of reciprocity, the difference is only that the sign of velocity $v$.
 The principle of reciprocity states that all inertial frames are equivalent.
-However, this relativity principle does not hold and the Lorentz covariance of electrodynamics equations does not dictate the Lorentz covariance of the solutions.
-The commonly accepted approach to special relativity does not account for the inertial (pseudo-gravitational) force within the accelerated frame as the agency of asymmetry.
-Space-time geometrical analysis of the accelerated frame (see Section 9.1) leads to the asymmetry (non-reciprocity) between the inertial frame and the accelerated frame.
+However, this relativity principle does not hold and the Lorentz covariance of all physical equations does not dictate the reciprocal symmetry of nature.
+We are going to discuss the asymmetry between inertial and accelerated frames.
+Space-time geometrical analysis of the accelerated frame leads to the asymmetry (non-reciprocity) between the inertial frame and the accelerated frame.
 We obtain in the accelerated frame the Langevin metric Eq.
 (15) of the second emitter.
-This metric describe the electrodynamics equations in the accelerated frame.
-We found of the essential asymmetry between inertial and accelerated frames, namely, the Maxwell’s equations are not applicable from the viewpoint of an observer at rest with respect to the accelerated frame.
+This metric describes the electrodynamics equations in the accelerated frame.
+We found the essential asymmetry between inertial and accelerated frames, namely, Maxwell’s equations are not applicable from the viewpoint of an observer at rest with respect to the accelerated frame.
 Because we are using absolute time coordinatization, the initial conditions (i.e.
-radiation wavefront orientations) are identical in both frames.
-We are going to discuss the asymmetry between inertial and accelerated frames.
-In the discussion up to this point, we assumed absolute time coordinatization.
-The Langevin metric, together with initial condition, describes the aberration of light effect for the second emitter at rest in the accelerated frame.
-The essential asymmetry is rooted in the electrodynamics equations.
-However, there is another satisfactory way of explaining the effect of aberration of light in the accelerated frame.
-The explanation consists in using a metric diagonalization procedure.
-The new coordinatization in the accelerated frame is interpreted by saying that the accelerated observer should obtain diagonal metric for the second emitter at rest and the (Langevin) metric Eq.
-(15) for the first (moving) emitter.
-Now we see that there is a symmetry (reciprocity) in electrodynamics equations between the accelerated and the inertial frame.
-This is a consequence of the pseudo-Euclidean geometry of space-time.
-Our impossibility to detect orbital velocity using second order optical (interference) experiments is a result of this metric reciprocity, see Chapter 11 for more details.
-Now the question is, where asymmetry comes from?
-Clearly, after the diagonalization, the asymmetry is rooted in the initial conditions.
-The plane radiation wavefront of both emitters rotates in the accelerated frame on the angle $-v/c$ after metric diagonalization.
-The previous derivation of the metrics in the accelerated frame includes one delicate point.
-According to a reciprocity in metrics, time dilation is symmetric.
-This contradiction is resolved by noting that metric diagonalization procedure does not preserve continuity of the metric tensor.
-In the previous chapters we learned that the accelerated frame is connected with the initial inertial frame by the Galilean transformation, but not the Lorentz one.
-In fact, the Langevin metric is found by matching the accelerated frame and inertial frame metric tensors.
-We shall discuss the smooth tailoring of the metric tensor further in the next section.
-9.6 Mathematical Expansion of Special Relativity onto Acceleration
-Before we begin the next topic, we would like to describe a number of mathematical ideas that are used in the literature of time dilation effect.
-Suppose we have two clocks in the inertial frame $S$ without accelerational history.
-Let first of these clocks always be at rest in the frame $S$ and the second clock accelerates and start to move with constant velocity $v$ along the $x$ axis.
-The proper time of the first clock coincides with the coordinate time $d\tau=dt$.
-Since the motion of both clocks is uniform, frames $S$ and $S_{n}$ will be inertial.
-According to textbooks, the coordinates and time in this frames should be related by the Lorentz transformation and the metrics of both frames will be diagonal.
-At first site the problem is complete symmetrical and there is no difference in the metrics in the frames $S$ and $S_{n}$.
-The situation seems paradoxical.
-We know that independent of the chosen coordinatization the proper time of the second accelerated clock is always less than that of the first clock.
-The time dilation of the accelerated clock is an absolute effect but not relative.
-It is independent of the time coordinate definition at the inertial part of the trajectory.
-This is an intrinsic property of space-time geometry.
-A more careful analysis shows that one must take into account that the metric tensor must be a continuous quantity.
-In the previous chapters we demonstrated that many problems in special relativity can be adequately treated only by an approach which uses the non-standard absolute time synchronization (i.e non diagonal metrics).
-We will keep the mathematical complexity to a minimum and assume that the second clock begins at $t=0$ to move with constant acceleration $a>0$ and does so till $t=T$.
-Now we would like to discuss an expression for the velocity from the viewpoint of inertial (non accelerated) observer.
-We know that uniformly accelerated motion must obey the equation $d(\gamma\vec{v})/dt=\vec{f}/m=\vec{a}$.
-This is readily integrated $\gamma\vec{v}=\vec{a}t$.
-We will then have $dx/dt=at/\sqrt{1+a^{2}t^{2}/c^{2}}$.
-We emphasize that $dx(t)/dt$ is the velocity of the second clock as seen by the inertial observer.
-Under the inverse Galilean transformation $dx=dx_{n}+atdt/\sqrt{1+a^{2}t^{2}/c^{2}}$, $dt_{n}=dt$ the metric of the uniformly accelerating observer becomes (3)
-$ds^{2}=c^{2}dt_{n}^{2}(1+a^{2}t_{n}^{2}/c^{2})^{-1}-2at_{n}dt_{n}dx_{n}(1+a^{2}t_{n}^{2}/c^{2})^{-1/2}-dx_{n}^{2}~{}.$ (19)
-At $t=0$ this metric naturally coincides with the $S$ frame metric and we have $g_{00}=1$.
-But here at the inertial part of the trajectory the metric of both frames will be diagonal in the Lorentz coordinatization and so at $t=T$ the metric tensor of the accelerated frame must abruptly change from the value $g_{00}=(1+a^{2}t_{n}^{2}/c^{2})^{-1}$ to the value $g_{00}=1$.
-This shows that we must match the coordinates in the accelerational and inertial parts of the (second clock) trajectory.
-During acceleration we picked the absolute time coordinatization.
-Therefore, for the metric tensor to be continuous when the reference frame passes from non-inertial motion to inertial one we will make use of the metric Eq.
-19.
-We will therefore require that at $t=T$ the metric of a non inertial frame associated with the second clock would pass continuously into the Langevin metric.
-For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer $(x_{n},t_{n})$ to the coordinates $(x,t)$ of the inertial observer by the Galilean boost $x_{n}=x-vt$, $t_{n}=t$, where $v=aT/\sqrt{1+a^{2}T^{2}/c^{2}}$.
-In this case, in the interval of coordinate time $T<t_{n}$ the reference frame of the second clock has the metric $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$.
-We emphasize that $v$ is the velocity of the second clock as seen by the inertial observer at the second (inertial) stage of motion.
-It becomes obvious that at $t=t_{n}=T$ the metric of non inertial frame goes into the metric of the (accelerated) inertial frame in a continuous manner.
-9.7 Bibliography and Notes
-1. For a general discussion of the equivalence principle we suggest reading the paper [42].
-2.
-In the general case, the problem to assigning Lorentz coordinates in an inertial frame is complicated.
-Let us consider a source, arbitrary accelerating in the inertial frame, and let us analyze its evolution with a Lorentz coordinate system.
-The permanent rest frame of the source is obviously non inertial.
-To get around that difficulty, one introduces an infinite sequence of comoving coordinate systems.
-At each instant, the comoving coordinate system is a Lorentz coordinate system centered on the source and moving with it.
-As the source velocity changes to its new value at an infinitesimally latter instant, a new Lorentz coordinate system centered on the source and moving with it at the new velocity is used to observe the source.
-We should make one further remark about this covariant algorithm.
-An opinion is sometimes expressed that described above algorithm includes a hidden postulate.
-It seems necessary a dynamical assumption to justify attributing to an accelerated clock the same rate as a clock in inertial motion in relation to which it is momentary at rest.
-This is, in view of some authors, an extra condition that a clock must satisfy.
-It is assumed that the effect of the motion on the clock depends only on it instantaneous speed, not its acceleration.
-This condition often refereed to as a ”clock hypothesis”.
-To quote Brown [17]: ”If the accelerating forces are small in relation to the internal restorative forces of the clock, then the clock’s proper time will be proportional to the Minkowski distance along its world line.
-… This condition is often referred to as the clock hypothesis, and its justification, as we have seen, rests on accelerative forces being small in the appropriate sense.” We state that the clock hypothesis does not have status of independent hypothesis is not needed as an independent postulate in the theory of relativity.
-As discussed above, the transformations within a single inertial frame are simply another parametrization of the observations of the inertial observer.
-According to such transformations, the observer (and his clocks) motion relative to the fixed stars does not changes.
-In other words, when we discuss only the observations of an inertial observer the problem of accelerated clock does not exist at all.
-3. For a general discussion of the mathematical expansion of special relativity onto acceleration we suggest reading the book [16].
-10 Relativistic Measurement
-In our approach to physics in general and of the theory of relativity in particular we think it very important always to remember that we are dealing with objective physical quantities measured by the measurements.
-We illustrate the problem of measurement in the special relativity by typical examples.
+radiation wavefront orientations) are identical in both frames  [^53] Many physicists tend to think that the validity of Maxwell’s laws in all inertial frames implies the elimination of the privileged frame.
+To quote e.g.
+Dieks [51]: ” As we already mentioned, it is a basic principle of the special theory of relativity that the line element $ds$ supplies all information about the physics of the situation, as described in the given coordinates.”.
+We would now like to describe an apparent paradox.
+The argument that all inertial frames are equivalent runs something like this.
+We know that to keep Lorentz coordinates in the accelerated frame we need to perform a metric diagonalization.
+Here the metric in both reference frames will be diagonal, and, according to textbooks, the coordinates in these frames should be related by the Lorentz transformation.
+At first sight, the problem is completely symmetrical and both frames are equivalent.
+However, there is a real difference between the accelerated inertial frame and the inertial frame without accelerational history.
+Indeed, we should expect the metric tensor to be changed.
+The difference between the inertial and accelerated frames is understandable.
+After diagonalization, the metric tensor in the accelerated frame must abruptly change from the value $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the value $g_{00}=1,g_{01}=0$.
+We must conclude that when we are dealing with (Lorentz ) transformation and this transformation does not preserve the continuity of the metric tensor we have no symmetry between inertial frames.
+13 Relativistic Measurements
 13.1 The Notion of Time in Special Relativity Theory
-There are many changes in what concepts are important when we go from classical to relativistic theory.
-We have already discussed some of them in previous chapters.
-We will show the connection between the classical notion of time and the relativistic time.
-The theory of relativity show us that the notion of time is not what we would have expected on the basis of our intuitive ideas.
+The theory of relativity shows us that the notion of time is not what we would have expected on the basis of our intuitive ideas.
 We attempt to give a definition of time on the basis of a principle ”Each physical magnitude is defined by the method it is measured.” Let us see how this definition works.
-In the case of length we know that there exist a length standard and we know that any body has its particular length.
-It should be note, however, that we are dealing with objective physical length only in the case of a stationary object.
+In the case of length, we know that there exists a length standard and we know that anybody has a particular length.
+It should be noted, however, that we are dealing with objective physical length only in the case of a stationary object.
 In contrast to this, the length of a moving object (e.g.
@@
 Now we are ready to consider the time problem in special relativity theory.
-Here is one example which shows in a circumstance that is easy to understand.
+Here is one example which shows a circumstance that is easy to understand.
 Suppose we know the law of muon decay in the Lorentz rest frame.
 When a Lorentz transformation of the decay law is tried, one obtains the prediction that, in the lab frame, the characteristic lifetime of a particle has increased from $\tau_{0}$ to $\gamma\tau_{0}$.
-We can interpreted this result by saying that after the travel distance $\gamma v\tau_{0}$, the population in the lab frame would reduced to 1/2 of the origin population.
+We can interpret this result by saying that after the travel distance $\gamma v\tau_{0}$, the population in the lab frame would reduced to 1/2 of the origin population.
 The time measured in this case is ”length” which is metrized as described above.
-That leads us to interesting question.
+That leads us to an interesting question.
 Is the same statement perhaps also true for all cases of time measurements in the framework of special relativity?
@@
 The frequency measured in this case is (standing wave) ”length”.
-Frequency measurements automatically obey the relationship between the frequency and interference.
+Frequency measurements automatically obey the relationship between frequency and interference.
 Another example of the relation between interference and frequency measurements is a grating spectrometer.
-The frequency measurement in this case is a position of the light spot along the dispersion direction.
+The frequency measurement in this case is the position of the light spot along the dispersion direction.
 There is no specific, important physical difference between diffraction grating and interferometer.
-Grating spectrometer is also based on the interference of incident and reflected light waves.
+The grating spectrometer is also based on the interference of incident and reflected light waves.
+Now let us see how we can compare the clock rates in the initial inertial frame $S$ and accelerated frame $S_{n}$.
+It is generally believed that observing clock rates in the two frames $S$ and $S_{n}$ moving relative to each other, one can only compare readings of one clock from one frame with readings of several clocks from another frame because two clocks from different frames get together at the same point in space only once.
+In one of the frames, there must be at least two clocks that are supposed to be synchronized.
+We have already pointed out that in all cases of time measurements in the framework of special relativity the time measured in this case is actually ”length”.
+In the case of length measurement of resting objects, we are dealing with objective (i.e.
+convention independent) physical quantity.
+Because we have empirical access only to the length measurements, it is impossible to agree with the textbook statement that one can only compare readings of one clock from one frame with readings of several synchronized clocks from another frame.
+Let us try out our operational interpretation of time measurements on the light clock example, to see how it works.
+Let us consider a specific thought experiment.
+We shall discuss the situation where time marks can be imprinted on the moving object (screen) by a clock.
+First, we describe the orthogonal light-clock based on the observations made by an observer in the same accelerated frame as the light-clock.
+A screen is at rest in the initial inertial frame.
+Suppose that the top mirror is semitransparent for marking the (moving) screen to measure the time between clock clicks.
+In the case of Langevin metric Eq.(15), the speed of light radiated by the accelerated emitter in the transverse direction is $c\sqrt{1-v^{2}/c^{2}}$.
+Therefore, the time interval between the sending and reception of the light signal is $\Delta t_{n}=2l/[c\sqrt{1-v^{2}/c^{2}}]$.
+The spacing of the time marks, i.e the distance $\Delta x_{n}$ through which the object (screen in the initial inertial frame) falls in the time is $\Delta x_{n}=\Delta x=2lv/[c\sqrt{1-v^{2}/c^{2}}]$.
+Taking into account the Langevin metric Eq.(15), we obtain at $dt_{n}=dt=0$ that $dx_{n}=dx$, i.e.
+the length of the physical rod in the inertial frame coincidence with the coordinate distance in the accelerated frame.
+Because $t_{n}=t$ and $\Delta t=2l/c$ have the physical meaning of the time indicated by a physical clock at rest in the initial inertial frame, this implies that, as physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame.
+The accelerated clock rate measured by the inertial observer is actually spacing of the time marks on the screen in the initial inertial frame.
+Second, we wise versa describe the orthogonal light-clock based on the observations made by an observer in the same initial inertial frame as the light-clock.
+Therefore, the time interval between the sending and reception of the light signal is $\Delta t=2l/c$.
+The spacing of the time marks, i.e.
+the distance $\Delta x$ through which the screen in the accelerated frame falls in the time is $\Delta x=\Delta x_{n}=2lv/c$.
+Analysis based on using standard measuring rods in an accelerated frame to measure its geometrical properties shows that coordinate distance $dx_{n}$ has physical length $dx_{n}/\sqrt{1-v^{2}/c^{2}}$.
+We have ascertained the asymmetry of the meter stick’s length directly from the Langevin metric.
+This shows that the physical spacing of the time marks measured in the accelerated frame undergoes compress on in the direction of fixed stars motion.
+Because $\Delta x_{n}=\Delta x$ and $\Delta x=2vl/c$ has the physical meaning of the time rate indicated by a physical clock at rest in the inertial frame, this implies that the accelerated observer can observe the shortness of mark spacing $\Delta l_{n}=2lv\sqrt{1-v^{2}/c^{2}}]/c$.
+Here $\Delta l_{n}$ is the physical mark spacing measured in the accelerated frame.
+This thought experiment illustrates that there is no reciprocity in time dilation.
 13.2 Invariance of the Two-Way Speed of Light and the Interference Phenomena
 We must now discuss a certain feature of the phenomenon of interference.
-Because of our usage of Galilean transformations within electrodynamics we have some apparent paradoxes.
-One finds many books which say that a Galilean transformation of the velocity of light is not consistent with the electron-theoretical explanation of refraction and reflection.
-Let us now consider a reflection and refraction in the context of special relativity.
-We are now in a position to understand what is happening when light radiated by a moving emitter goes through a stationary material like glass.
+Let us consider a reflection and refraction in the context of special relativity.
+We are now in a position to understand what happens when light radiated by a moving emitter goes through a stationary material like glass.
 We see that what we have to do is to calculate the interference between the incident radiation and the radiation due to the material.
@@
 So the possibility of calculating interference effects is not clear.
-Figure 47: Light radiated by a moving emitter reflects from a stationary mirror in the collinear geometry.
-There is interference between the incident radiation and the radiation due to the material.
-Galilean transformation of the velocity of light is consistent with the electron-theoretical explanation of refraction and reflection.
-The Minkowski metric and the metric Eq.
-(11) give the same prediction for the wavelength of the standing wave.
+Figure 48: Light radiated by a moving emitter reflects from a stationary mirror in the collinear geometry. There is interference between the incident radiation and the radiation due to the material.
 If the mirror is at rest and the emitter is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves.
 In other words, the same metric should be applied and kept in a consistent way for both incident and scattered waves.
-The requirement that the electrodynamics equations must be identical for both incident and reflected waves appear to be a paradox here.
-In fact, we have two different metrics.
+The requirement that the electrodynamics equations must be identical for both incident and reflected waves appears to be a paradox here.
+Indeed, we have two different metrics.
 We have already made a few remarks about two different metrics in the aberration of light problem.
-In the Chapter 4, we considered only the first order approximation.
-We pointed out that the paradox is avoided because the peculiarity of the aberration of light (perpendicular) geometry is that even after the Galilean transformation the electrodynamics equations in the absolute time coordinatization will have Maxwell’s form for the incident plane wave.
-Let us now see what happens in the collinear geometry, Fig.
-47.
-Here we just summarize the general idea.
+In the Section 4, we considered only the first-order approximation.
+We pointed out that the paradox is avoided because the peculiarity of the aberration of light geometry is that even after the Galilean transformation the electrodynamics equations in the absolute time coordinatization will have Maxwell’s form for the normally incident plane wave.
+Let us see what happens in the collinear geometry, Fig.
+48.
+Here we summarize the general idea.
 We will show that the paradox is avoided because Minkowski metric and metric Eq.
@@
 Because we have empirical access only to the round-trip average speed of light the value of the one-way speed of light is just a matter of convention without physical meaning.
-In contrast to this, the two-way speed of light, directly measurable along a round-trip, has physical meaning.
+In contrast, the two-way speed of light, directly measurable along a round-trip, has physical meaning.
 We must now discuss a certain feature of the phenomenon of interference.
-One finds many books which say that a Galilean transformation of the velocity of light is not consistent with the electron-theoretical explanation of refraction and reflection.
-It is widely accepted that if we consider a moving source and a stationary glass, the incident light wave and wave scattered by the dipoles of the glass cannot interfere as required by the electron theory of dispersion since their velocity are different.
+It is widely accepted that if we consider a moving source and a stationary glass, the incident light wave and wave scattered by the dipoles of the glass cannot interfere as required by the electron theory of dispersion since their velocity are different in the absolute time coordinatization.
 This is simply not true in physics.
@@
 Thus, the incident and scattered wave at any given point have the same frequency and can interfere.
-The effect of the different velocities is to produce a relative phase which varies with position in space.
+The effect of the different velocities is to produce a relative phase that varies with position in space.
 This affects the velocity and amplitude envelope of the single wave which results from the superposition of the two separate waves.
 The following simple analysis confirms these ideas.
-We will describe effect by assuming that the metric is non-diagonal, Eq.
+We will describe the effect by assuming that the metric is non-diagonal, Eq.
 11.
@@
 The superposition of these two waves is represented by $\exp i(\omega t-kx)+\exp i(\omega t+k^{\prime}x+\phi)=2[\cos[(k+k^{\prime})x/2+\phi/2]]\exp i[\omega t-(k-k^{\prime})x/2+\phi/2]$.
-There is the cosine factor representing an amplitude envelope which is stationary in space and whose periodicity is inversely proportional to the difference in the propagating constants $k$ and $k^{\prime}$ of the two component waves.
+There is the cosine factor representing an amplitude envelope that is stationary in space and whose periodicity is inversely proportional to the difference in the propagating constants $k$ and $k^{\prime}$ of the two component waves.
 This can be written in a simpler form $2[\cos[\omega x/[c(1-v^{2}/c^{2})]+\phi/2]]\exp i[\omega t-xv\omega/[c^{2}(1-v^{2}/c^{2})]+\phi/2]$.
 Suppose that the source at rest is emitting waves at frequency $\omega_{0}$.
-In the lab frame after the Galilean transformation the velocity of incoming wave is $c+v$.
+In the lab frame after the Galilean transformation the velocity of the incoming wave is $c+v$.
 Thus if $\omega_{0}$ is the natural frequency, the observed in the lab frame frequency would be $\omega=\omega_{0}/[1-v/(c+v)]$.
 Therefore, the observed in the lab frame frequency is $\omega=\omega_{0}(1+v/c)$.
-The shift in frequency observed in the above situation is the well known Doppler effect.
-Our equation for superposition of two waves now looks like $2[\cos[\omega_{0}x/[c(1-v/c)]+\phi/2]]\exp i[\omega_{0}(1+v/c)t-xv\omega_{0}/[c^{2}(1-v/c)]+\phi/2]$.
+The shift in frequency observed in the above situation is the well-known Doppler effect.
+Our equation for the superposition of two waves now looks like $2[\cos[\omega_{0}x/[c(1-v/c)]+\phi/2]]\exp i[\omega_{0}(1+v/c)t-xv\omega_{0}/[c^{2}(1-v/c)]+\phi/2]$.
 Suppose that an observer in the laboratory performs the standing wave measurement.
-We should examine what parts of the measured data depends on the choice of synchronization convention and what parts do not.
+We should examine what parts of the measured data depend on the choice of synchronization convention and what parts do not.
 We state that time oscillation has no intrinsic meaning - its meaning only being assigned by a convention.
 In particular, one can see the connection between the time shift $xv\omega_{0}(1+v/c)/c^{2}$ in $\exp i[\omega_{0}(1+v/c)t-xv\omega_{0}/[c^{2}(1-v/c)]+\phi/2]$ and the issue of distant clock synchrony.
-Note that the scale of time (frequency) is also unrecognizable from physical viewpoint.
-Suppose that the laboratory observer performs a measurement of the wavelength of the standing wave.
+Note that the scale of time (frequency) is also unrecognizable from a physical viewpoint.
+Suppose that the laboratory observer measures the wavelength of the standing wave.
 The relation $dl_{i}^{2}=(-g_{11}+g_{01}g_{01}/g_{00})dx^{2}$ gives the connection between the spatial line element and the metric.
-With the help of Eq.(11), we find $dl_{i}^{2}=dx^{2}/(1-v^{2}/c^{2})$, which determines the spatial geometry in the inertial frame.
-This is explained as due to Lorentz contraction of the measuring rods in the inertial frame.
+With the help of Eq.(11), we find $dl_{i}^{2}=dx^{2}/(1-v^{2}/c^{2})$, which determines the spatial geometry in the case of the Langevin metric.
+This is explained as due to Lorentz contraction of the measuring rods.
 Then, when the measured data is analyzed, the laboratory observer finds that the wave number is equal to $\sqrt{1-v^{2}/c^{2}}\omega_{0}/[c(1-v/c)]=\omega_{0}\sqrt{1+v/c}/\sqrt{1-v/c}$.
-The observer finds that the wavelength of radiation from moving source (the source moves towards the observer) is decreased by the factor $\sqrt{1-v/c}/\sqrt{1+v/c}$.
+The observer finds that the wavelength of radiation from the moving source (the source moves towards the observer) is decreased by the factor $\sqrt{1-v/c}/\sqrt{1+v/c}$.
 We see that it is the same factor that we can obtain by assuming that the metric is diagonal.
 In other words, the laboratory observer will measure the same two-way speed of light, irrespective of the choice of metric.
-The theory of relativity shows us that there is a certain quantity, which we call two-way speed of light, that does not change under any reference frame transformations.
-Another way in which our complicated problem can be solved is the following interesting one.
-The reason why the laboratory observer will measure the same two-way speed of light is that the method is based on the phenomenon of interference.
+The reason why the laboratory observer will measure the same wavelength is that the method is based on the phenomenon of interference.
 There is a fact governing all interference phenomena.
@@
 It states that there is a certain quantity, which we call phase, that does not change under any reference frame transformations.
-The relativistic invariance of phase is enormously useful in making analysis (as we will illustrate in a few examples in the Chapter 11) without knowing all the formulas.
+The relativistic invariance of phase is enormously useful in making an analysis (as we will illustrate in a few examples in the next section) without knowing all the formulas.
 In the problem discussed we have still a puzzle.
 Let us consider the acceleration of a dipole source in the lab inertial frame up to velocity $v$ along the $x$-axis.
-Without changing synchronization in the lab frame after the source acceleration we have a complicated situation as concerns dynamics and electrodynamics of moving charges.
-Conventional particle tracking in the single inertial frame is actually based on classical Newton mechanics.
+Without changing synchronization in the lab frame after the source acceleration we have a complicated situation concerning the dynamics and electrodynamics of moving charges.
+The dynamical evolution of the particle according to the conventional (non-covariant) particle tracking may be considered as the result of successive infinitesimal Galilean transformations.
+The corrected Newton’s law is valid for each step.
 It is not difficult to see that the peculiarity of this situation is that here a solution of the dynamics problem in the inertial frame makes no reference to Lorentz transformations.
@@
 The relativity of simultaneously (i.e.
-mixture of positions and time) do not have a place in this description.
-The dynamical evolution of the particle according to the non-covariant particle tracking may be considered as the result of successive infinitesimal Galilean transformations.
-The corrected Newton law is valid for each step.
-On the other hand, the Maxwell’s equations can be applied in the inertial frame only in the case when Lorentz coordinates are assigned.
+mixture of positions and time) does not have a place in this description.
+On the other hand, Maxwell’s equations can be applied in the inertial frame only in the case when Lorentz coordinates are assigned.
 We will then ask how it is possible to take old kinematics for mechanics and Einstein’s kinematics for electrodynamics.
@@
 How can that be?
-We already know that the Galilean boosts commute.
+We have already discussed the explanation of this puzzle in Section 3.6.
+The Galilean boosts commute.
 The collinear Lorentz boosts also commute.
-Now let us see what happens to the moving source.
-The peculiarity of the collinear geometry where there is a source moving along the same line as the radiated beam is that here the velocity is perpendicular to the plane of radiation wavefront (i.e plane of simultaneity).
-Thus, for collinear motion, the plane of simultaneity in the absolute time coordinatization will have the same orientation for the Lorentz coordinatization.
-In the case of collinear geometry, a motion of the source, according to the special relativity, influences the kinematic terms of the higher (than $v/c$) order only.
-At first site, the Newton’s particle dynamics (in the case of collinear geometry) does not involve the relativistic effects at all.
-But the higher order effects are there in the assumption that the mass of the moving particle is equal to its relativistic mass.
-We shall discuss this subject further in the Chapter 13.
-10.3 Field Measurements in the Accelerated Frame
-In special relativity, we are confronted with a few things that might seem counter intuitive.
-But this is still a mathematically based theory.
-It tell us a peculiar thing: that when we are accelerated electron source, the fields in the accelerated frame is not static; the problem is time dependent.
-That what special relativity says!
-When the electron source and observer are at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be nonzero.
-Intuition would seem to tell us that everything is at rest, so the electric field is not changing with time.
-But special relativity says that there is a derivative $\partial\vec{E}/\partial t_{n}=v\partial\vec{E}/\partial x_{n}$ because there is $\partial\vec{E}/\partial x_{n}$ that is not zero.
-This is just a feature of the 4D formalism with no special consequence.
-A comparison with a gauge transformation in classical electromagnetism might help here.
-In electrodynamics, it is often convenient to introduce the Coulomb gauge.
-In this case field propagates instantaneously.
-But this instantaneous propagation is just mathematical description in an intermediate step: when a complete calculation is made, proper cancellations of the instantaneous propagation take place.
-10.3.1 The Aberration of Particles from a Moving Source
-Let us consider what happens if the electron gun, an electron source for a relativistic electron beam, in the initial inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis, Fig.
-48.
-In other words, we consider so-called active (physical) boost.
-In the case of an active boost we are dealing with motion of the same physical system, evolving in time and treated from the point of view of the same reference system.
-The simplest method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the particle source was at rest, i.e.
-we still enforce the clock transport synchronization ( or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest).
-This choice is usually the most convenient one from the viewpoint of the smooth tailoring of the metric tensor.
-This synchronization convention preserves simultaneity and is actually based on the absolute time (or absolute simultaneity) convention.
-Figure 48: Aberration of electrons from a moving source in an inertial frame.
-The motion of the electron gun is perpendicular to the optical axis.
-If an observer is at rest relative to the fixed stars and the electron gun started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the source seen in the inertial frame would jump by angle $\theta_{a}=v/V$, where $V$ is the velocity of the accelerated electrons.
-We can obtain the electric field $\vec{E}$ and the magnetic $\vec{B}$ observed in the inertial frame where the particle source is moving with velocity $v$ using the Galilean transformation.
-When we consider the electrodynamics equations under Galilean transformation, there is a question: what is the transformation law for the electromagnetic fields?
-The equivalence of the active and passive pictures is due to the fact that moving system one way is equivalent to moving the coordinate system the other way by an equal amount.
-It is for this reason that fields in the inertial frame before the active Galilean boost are the same as those $\vec{E^{\prime}}$ and $\vec{B^{\prime}}$ in the comoving coordinate system.
-With the Galilean transformation, the field transformation, for the case $\vec{B^{\prime}}=0$, is $\vec{E}=\vec{E^{\prime}}$, $\vec{B}=\vec{v}\times\vec{E^{\prime}}/c$.
-This expression describe first order (in $v/c$) effects only.
-Here $\vec{v}$ is the velocity vector of the source in the lab frame.
-[…]
-Figure 49: Aberration of electrons when the electron source and observer are at rest in the accelerated frame.
-The aberration increment is connected with the problem parameters by the relation $\theta_{a}=(1-1/\gamma)v/V$.
-Here $\gamma=1/\sqrt{1-V^{2}/c^{2}}$ is the relativistic factor.
-We described the effect of aberration of particles by working only up to the first order $v/c$.
-An elementary explanation of this effect is well-known.
-The aberration of particles in the inertial lab frame can be easily explained in terms of Galilean transformation of velocities between different reference frames.
-It is important to stress at this point that the dynamical line of arguments discussed here explains what the aberration of particles physically means in the inertial frame.
-The kinematics effects is only an interpretation of the behavior of the electromagnetic fields.
-In Chapter 8 we already discussed the aberration of electrons.
-We had to restrict ourselves to discussing the aberration only for the case when the electron source and observer are at rest in the accelerated frame.
-The physical principles that produces the aberration increment were, however, made clear.
-We have seen in Chapter 8 that the Wigner rotation associated with the transformation from inertial coordinate system to the accelerated coordinate system may be regarded as a result of the action of certain electromagnetic (Lorentz) force.
-In fact, the electromagnetic forces which govern the properties of an emitted electron beam must be affected by acceleration with respect to the fixed stars in such a way that they lead to a deviation of the electron transport direction, Fig.
-49.
-The microscopic approach for the motion of electrons in the accelerated frame gives a similar (for the motion of electrons in the inertial frame) answer for the total transverse momentum, $p_{\perp}$, of the accelerated electron beam except with the following modification.
-Because the electron source is at rest in the accelerated frame, the initial condition is modified.
-We can rewrite our formula for the transverse momentum in the inertial frame as $p_{\perp}=p_{0}+\Delta p_{\perp}$, where $\Delta p_{\perp}=vT/c^{2}=mv/\sqrt{1-V^{2}/c^{2}}-mv$.
-In the accelerated frame, the initial momentum of the emitted electron $p_{0}$ is equal to zero.
-Therefore, when the electron source is at rest in the accelerated frame, there is an aberration increment given by $\theta_{a}=p_{\perp}/p=\Delta p_{\perp}/p=(1-1/\gamma)v/V$.
-So we see that we could analyze this complicated situation either by the idea that there is an electromagnetic forces and that they lead to a deviation of the energy transport, or else by the Lorentz coordinatization and that the aberration of particles effect in the accelerated frame (as viewed from the inertial frame) presents a kinematics effect (Wigner rotation) of special theory of relativity.
-10.3.2 Magnetic Field Measurements in the Accelerated Electron Source
-Let us return to the case when the electron source and observer are at rest in the accelerated frame, Fig.
-49.
-We would now like to describe an apparent paradox.
-We obtained the electrodynamics equations in accelerated frame using the inverse Galilean transformation (with velocity $-v$) of the Maxwell’s equations.
-With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$.
-The electric $\vec{E}_{n}$ and magnetic $\vec{B}_{n}$ fields are not changing with time - all of the fields are static, so that the problem is not time dependent.
-There are many physicists who have already received knowledge about the electrodynamics from textbooks and who would say, ”There must be currents in order to get a static magnetic field at all - and currents can only from moving charges”.
-Since there are no moving charges , the situation seems indeed paradoxical.
-Here we have a new kind of situation which is quite different from (initial) inertial frame electrodynamics.
-How can that be?
-When the electron source and observer are at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be nonzero.
-One can easily show this by transforming coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$.
-The variables $x,y,z,t$ can be expressed in terms of the independent variables $x_{n},y_{n},z_{n},t_{n}$ by means of a Galilean transformation, so that fields can be written in terms of $x_{n},y_{n},z_{n},t_{n}$.
-From the Galilean transformation $x_{n}=x-vt,~{}y_{n}=y,~{}z_{n}=z,~{}t_{n}=t$, after partial differentiation, one obtains $\partial/{\partial t_{n}}=\partial/{\partial t}+v\partial/{\partial x}$, $\partial/{\partial x_{n}}=\partial/{\partial x}$.
-So $\partial/\partial t_{n}=v\partial/\partial x_{n}$ at $\partial/\partial t=0$.
-A natural assumption is to consider that time derivative $\partial/\partial t_{n}$ arises from the Galilean boost.
-Now we shall be more general and study the case of the Lorentz coordinatization.
-When the system $S_{n}$ starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed.
-The time $t^{(L)}_{n}$ under the Einstein’s synchronization in the $S_{n}$ frame is readily obtained by introducing the offset factor $x_{n}v/c^{2}$ and substituting $t^{(L)}_{n}=t_{n}-x_{n}v/c^{2}$ in the first order approximation.
-The new time coordinate in the accelerated frame is interpreted by saying that Maxwell’s equations are applicable to the aberration of electrons description.
-After partial differentiation, one obtains $\partial/{\partial t^{(L)}_{n}}=v\partial/{\partial x_{n}}$, $\partial/{\partial x^{(L)}_{n}}=\partial/{\partial x_{n}}+(v/c^{2})\partial/{\partial t_{n}}$.
-So $\partial/\partial t^{(L)}_{n}=v\partial/\partial x^{(L)}_{n}$ neglecting terms of order of $v^{2}/c^{2}$.
-By changing the (four-dimensional) coordinate system, one cannot obtain a physics in which new physical phenomena appear.
-When the situation is described as we have done it here, there doesn’t seem to be any paradox at all; it comes out quite natural that separation between space and time is no longer possible in general.
-The argument that the result is paradoxical runs some thing like this: If we could measuring magnetic field $\vec{B}_{n}$ by usual magnetometer there would indeed be a paradox.
-But we could not do that.
-We already discussed above that magnetic field $\vec{B}_{n}$ leads to change in the transverse momentum of test particles given by $\theta_{a}=p_{\perp}/p=(1-1/\gamma)v/V$.
-If a particle motion velocity is non-relativistic, the binomial expansion yields $\vec{\theta}_{a}=-[V^{2}/(2c^{2})]\vec{V}\times\vec{v}/V^{2}$ and the momentum perturbation presents a relativistic effect.
-So special relativity says that there are no changes in the (test) particle momentum as $V^{2}/c^{2}\to 0$.
-If we look at the usual (i.e.
-classical) magnetometer with test particles, we see no evidence of the magnetic field $B_{n}$ effect.
-Not only does this particular kind of classical magnetometer can not measure the magnetic field $B_{n}$, but if the theory of relativity is correct, any other classical magnetometer, operating on any principle whatsoever, would also appear to demonstrate no evidence of the magnetic field $B_{n}$ effect (1).
-Now let us return to observations of an inertial observer, Fig.
-48.
-In this case we are dealing with motion of the electron source, evolving in time and treated from point of view of the lab system.
-It is not difficult to see that the peculiarity of the situation is that here the ”source motion” is a real observable effect.
-Indeed, the problem is time dependent.
-In this situation, a velocity of the source has physical meaning.
-In the case of moving electron source the special relativity says that magnetic field $\vec{B}=\vec{v}\times\vec{E}$ leads to change in the transverse momentum of test particles given by $\theta_{a}=p_{\perp}/p=v/V$ and this result is valid also in the non-relativistic limit $V^{2}/c^{2}\to 0$.
-Now we can raise an interesting question.
-Suppose that lab observer is measuring the magnetic field with an usual magnetometer.
-Now we know that there are moving charges, so that means that when there is a current in the moving source, the source generates a magnetic field which can be measured with an usual magnetometer.
-We come to the conclusion that the magnetic field generated by a moving source in an inertial frame is in fact a real field in the sense we have defined it.
-10.4 Inertial Frame View of Accelerated Light Clock
-In the Chapter 9, we described the orthogonal light-clock based on observations made by an observer in the same accelerated frame as the light-clock.
-Now we want to extend the description in an important way - in a way that says something in detail about how light is propagated in the inertial frame.
-We begin by recalling the aberration of light from a moving laser source in an inertial frame of reference.
-Although the laser seems to be quite different from a simple electromagnetic model for the light-clock of special relativity, the two systems are, of course closely related.
-10.4.1 The Aberration of Light from a Moving Laser Source
-The aberration of light effect can be described within the standard special relativity taking advantage of the Wigner rotation theory.
-When one has a transversely (i.e.
-parallel to the mirrors) moving laser there is the deviation of the energy transport for light transmitted from the optical resonator.
-In the first order approximation in $v/c$, the inertial observer would find that angular displacement is $\theta_{a}=2v/c$, Fig.
-50.
-How shall we describe the aberration of light beam from the laser source which is accelerated from rest up to velocity $v$ in the inertial frame?
-First, we notice that there is an extra factor 2.
-On the basis of conventional theory one gets aberration increment only one-half as big as that predicted by special relativity.
-One-half of this increment is a consequence of the fact that the Doppler effect is responsible for the angular frequency dispersion of the radiated light waves.
-Another source of the aberration of light is the Wigner rotation.
-As viewed from the (Lorentz) initial inertial frame, the coordinate axes of the accelerated frame is rotated through the angle $v/c$ with respect to the coordinate axes of the inertial frame.
-In regard to light aberration one should differentiate between that from the ”plane wave” emitter and that from the laser source.
-Suppose we have a ”plane wave” emitter.
-We consider the case when the ”plane wave” emitter is accelerated from rest up to velocity $v$ in the direction perpendicular to its optical axis.
-Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the light transport measurement.
-Then how does the light beam from the moving ”plane wave” emitter looks?
-The inertial observer would find that angular displacement is equal to $\theta_{a}=v/c$, Fig.
-51.
-The first thing we would say about inertial frame measurements is that there is an intuitively plausible way to understanding the aberration of light from a moving ”plane wave” emitter.
-An elementary explanation of this effect is well-known.
-This phenomenon is fully understandable in terms of transformation of velocities between different reference frames.
-The aberration of light in the initial inertial frame can also be easily explained on the basis of corpuscular theory of light.
-This is plausible if one keeps in mind that a light signal represents a certain amount of electromagnetic energy.
-Energy, like mass, is a quantity that is conserved, so that a light signal resembles, in many aspects, a material particles.
-Therefore, we should expect that group velocities of light signals obey the same addition theorem for particle velocities.
-A closer treatment based on wave theory of light confirm this expectation.
-In the case of the moving laser source, intuition would seem to tell us that aberration increment would be the same.
-But the special relativity says that there is an extra factor 2.
-The commonly accepted derivation of the aberration of light effect does not account for the Wigner rotation.
-It is incorrectly assumed that accelerated and inertial observers have common 3-space.
-Figure 50: Aberration of light from a moving laser source in an inertial frame of reference.
-The motion of the laser resonator is parallel to the mirrors.
-If an observer is at rest relative to the fixed stars and the laser source started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the laser source seen in the inertial frame would jump by angle $\theta_{a}=2v/c$.
-It is assumed that the accelerated laser operates in the steady-state regime.
-Figure 51: Aberration of light from a moving ”plane wave” emitter in an inertial frame of reference.
-The motion of the source is perpendicular to the optical axis.
-If an observer is at rest relative to the fixed stars and the ”plane wave” emitter started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the source seen in the inertial frame would jump by angle $\theta_{a}=v/c$.
-Figure 52: Aberration of light in an accelerated frame of reference.
-It is assumed that the accelerated laser operates in the steady-state regime.
-The aberration of light from the accelerated laser source is absent.
-The electromagnetic wave travels in the resonator forward and back reflecting from mirrors.
-The crossed term in the Langevin metric, which generates aberration, cancels during evolution of the radiation in the optical resonator.
-Figure 53: Aberration of light from a ”plane wave” emitter in an accelerated frame of reference.
-If the ”plane wave” emitter is at rest in the accelerated frame, then the apparent angular position of the source seen in the accelerated frame would jump by angle $\theta_{a}=-v/c$.
-Now we ask about the angular displacement of light transmitted from the laser resonator inside the accelerated frame.
-The accelerated observer would find that the deviation of the energy transport direction is absent, Fig.
-52.
-According to asymmetry between the ”plane wave” emitter and the laser source, there is a remarkable prediction on the theory of the aberration of light.
-Namely, if the ”plane wave” emitter was at rest relative to the fixed stars and this source started from rest to motion relative to the fixed stars, then the apparent angular position of the ”plane wave” emitter seen in the accelerated frame would jump by angle $-v/c$, Fig.
-53.
-Obviously, in order to understand the aberration of light, we will have to use the special theory of relativity and not be satisfied with classical theory.
-How can such difference between a ”plane wave” emitter and a laser source come about?
-Let us first ask: ”What happens to accelerated optical resonator?” When we were dealing with optical resonators we used the electrodynamics boundary conditions in 3-space, and they defined the energy propagation direction along the optical axis (i.e.
-perpendicular to the mirror surface).
-According to special relativity, an acceleration does not change the direction of the optical axis (i.e mirrors orientation) with respect to the accelerated frame (3-space) axes.
-An observer in the rest frame of a laser source finds that the direction of the energy transport is independent of the motion with respect to the fixed stars, Fig.
-52.
-We turn now to the radiation emitted by the accelerated laser source in the inertial frame.
-As viewed from the inertial frame, the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$ is rotated through the Wigner angle $\theta_{w}$ with respect to the axes of the inertial frame $(x,y,z)$.
-For light $V/c\to 1$ and the $z_{n}$-axis of accelerated frame is then rotated with respect to the $z$-axis by angle $\theta_{w}=v/c$.
-As a result, an extra factor of 2 appears in the aberration increment, in Fig.
-50.
-Let us now discuss more of the difference between laser and incoherent source.
-Suppose we have a ”plane wave” emitter which is at rest in an accelerated frame.
-How shall we find the general wave solution in the case of a point source?
-The answer is that electrodynamics equation needs to be integrated with initial condition for the radiation wavefront.
-According to the special relativity, physical process of light creation takes place in a metric space-time geometry, so the metric tensor must be used to describe the light creation and propagation in empty space.
-We only wish to emphasize here the following point.
-According to special relativity, accelerated and inertial observers have common absolute space-time but have different metrics and, consequently, different 3-spaces.
-When we are dealing with optical resonator the (3-space) boundary conditions define the energy propagation direction.
-In contrast, when we are dealing with a point source the combination of initial condition (i.e orientation of the radiation wavefront ) and metric tensor is important.
-We would like to emphasize a very important difference between laser and incoherent source.
-The radiation of incoherent source clearly depends on an initial conditions and there is influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer.
-The solution presented in terms of the absolute time coordinate, hence the wavefront of the emitted light beam is perpendicular to the direction $z_{n}$ after the acceleration.
-Using this initial condition and the Langevin metric in a frame of reference attending acceleration enables us to explain the aberration of light from a ”plane wave” emitter, Fig.
-53.
-10.4.2 Orthogonal Light Clock Operation
-Slowing of the clock in a moving system is a very peculiar phenomenon.
-We have already discussed accelerated clocks in the previous chapter.
-We described the parallel and orthogonal light clocks based on the observations made by an observer in the same accelerated frame as the clock.
-Now let us return to observations of an inertial observer.
-We must say immediately that there is no objection to the standard description of the parallel light clock in the inertial lab frame.
-What must be recognized is that the concept of Wigner rotation is only introduced in the description of an orthogonal light clock.
-Let us consider a light clock accelerated up to velocity $v$ transverse to the direction of optical pulse.
-First, we examine the reasoning presented in textbooks.
-Suppose that an inertial system $S$ (and an observer with his measuring instruments) is at rest with respect to the fixed stars and a system $S_{n}$ in the lab frame $S$ is accelerated from the rest with respect to the fixed stars up to velocity $v$ along the $x$-axis.
-A clock which we call ”light-clock” is a rod with mirror at each end.
-According to textbooks, when inertial observer looks at the orthogonal light clock going by, he sees that the light, in going from mirror to mirror is taking a zigzag path.
-If in a giving time the clock moves forward a distance proportional to $v$ in Fig.
-54, the distance the light travels in the same time is proportional to $c$, and the vertical distance is therefore proportional to $\sqrt{c^{2}-v^{2}}$.
-That is it takes a longer time for light to go from end to end in the moving clock than in stationary clock.
-Therefore the apparent time between clicks is longer for the moving clock in the proportion $1/\sqrt{1-v^{2}/c^{2}}$.
-Figure 54: The ”light clock” thought experiment.
-The calculation of the time interval between the sending and reception of the light signal in the reference frame $S$ relative to which a light clock moves.
-According to textbooks, it is assumed that $(x_{n},y_{n},z_{n})$ axes of the moving frame and $(x,y,z)$ axes of the lab observer are parallel.
-Nevertheless, there is argument against this commonly accepted derivation of the slowing down of an orthogonal light clock in a moving system.
-The standard explanation is based on the hidden assumption that $(x_{n},y_{n},z_{n})$ axes of the moving observer and $(x,y,z)$ axes of the lab observer are parallel.
-In other words, it is assumed that observers have common 3-space.
-This is misconception.
-In fact, two observers with different trajectories in the Lorentz coordinatization have different 3-spaces.
-The accepted in textbooks assumption of rigidity of accelerated frame is based on the belief that simultaneous acceleration of the reference space grid $(x_{n},y_{n},z_{n})$ has direct physical meaning.
-However, the position along the $x$-axis of the accelerated first mirror with respect to the accelerated second mirror has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this position could be ascertained.
-There is an uncertainty (blurring) of the relative position in the $x$-direction of amount $lv/c$, where $l$ is the rod length.
-This uncertainty in the relative position exist due to the uncertainty in the moments of the acceleration.
-The theory of relativity shows us that the relationship of positions and times are not what we would expect intuitively.
-When we have two distant events, we have deal with the conventionality of distant simultaneity within the time interval $l/c$, where $l$ is the space separation of these two events.
-Now let us see how we can compare the clock rates in the frame $S$ and $S_{n}$.
-It is generally believed that observing clock rates in the two frames $S$ and $S_{n}$ moving relative to each other, one can only compare readings of one clock from one frame with readings of several clocks from another frame, because two clocks from different frames get together at the same point in space only once.
-In one of the frames there must be at least two clocks which are supposed to be synchronized.
-We have already pointed out that in all cases of time measurements in the framework of special relativity the time measured in this case is actually ”length”.
-In the case of length measurement of resting objects we are dealing with objective (i.e.
-convention independent) physical quantity.
-Because we have empirical access only to the length measurements, it is impossible to agree with textbook statement that one can only compare readings of one clock from one frame with readings of several synchronized clocks from another frame.
-Let us try out our operational interpretation of time measurements on the light clock example, to see how it works.
-Let us consider a specific thought experiment.
-We shall discuss the situation where time marks can be imprinted on the moving object (screen) by a clock, as sketched in Fig.
-55.
-First, we describe the orthogonal light-clock based on the observations made by an observer in the same accelerated frame as the light-clock.
-In the case of Langevin metric Eq.(15), the speed of light radiated by the accelerated emitter in the transverse direction is $c\sqrt{1-v^{2}/c^{2}}$.
-Therefore, the time interval between the sending and reception of the light signal is $\Delta t_{n}=2l/[c\sqrt{1-v^{2}/c^{2}}]$.
-The spacing of the time marks, i.e the distance $\Delta x_{n}$ through which the object (screen in the lab frame) falls in the time is $\Delta x_{n}=\Delta x=2lv/[c\sqrt{1-v^{2}/c^{2}}]$.
-Taking into account the Langevin metric Eq.(15), we obtain at $dt_{n}=dt=0$ that $dx_{n}=dx$, i.e.
-the length of physical rod in the inertial frame coincidence with the coordinate distance in the accelerated frame.
-Because $t_{n}=t$ and $\Delta t=2l/c$ has the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that, as physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame.
-The accelerated clock rate measured by the lab observer is actually spacing of the time marks on the lab screen.
-Figure 55: The ”light clock” thought experiment.
-Time marks can be imprinted on the moving screen by clock.
-Observations made by an observer in the same accelerated frame as the light clock.
-The anisotropy of speed of light presented in the absolute time coordinatization.
-The distance $\Delta x_{n}$ through which the screen in the lab frame falls in the time (i.e.
-spacing of the time marks) is $\Delta x_{n}=\Delta x=2vl\gamma/c$.
-Second, we describe the orthogonal light-clock based on the observations made by an observer in the same inertial frame as the light-clock, Fig.
-56.
-Therefore, the time interval between the sending and reception of the light signal is $\Delta t=2l/c$.
-The spacing of the time marks, i.e the distance $\Delta x$ through which the screen in the accelerated frame falls in the time is $\Delta x=\Delta x_{n}=2lv/c$.
-Analysis based on using standard measuring rods in an accelerated frame to measure it geometrical properties shows that coordinate distance $dx_{n}$ has physical length $dx_{n}/\sqrt{1-v^{2}/c^{2}}$.
-We have ascertained the asymmetry of meter stick’s length directly from the Langevin metric.
-This shows that physical spacing of the time marks measured in the accelerated frame undergoes compress on in the direction of fixed stars motion.
-Because $\Delta x_{n}=\Delta x$ and $\Delta x=2vl/c$ has the physical meaning of the time rate indicated by a physical clock at rest in the inertial frame, this implies that the accelerated observer can observe the shortness of mark spacing $\Delta l_{n}=2lv\sqrt{1-v^{2}/c^{2}}]/c$.
-Here $\Delta l_{n}$ is the physical mark spacing measured in the accelerated frame.
-In fact, we deal with the same time dilation effect.
-This thought experiment illustrates that there is no reciprocity in time dilation.
-Figure 56: The ”light clock” thought experiment.
-Time marks can be imprinted on the moving screen by clock.
-Observations made by an observer in the same inertial frame as the light clock.
-The distance $\Delta x_{n}$ through which the screen in the accelerated frame falls in the time (i.e.
-spacing of the time marks) is $\Delta x_{n}=\Delta x=2vl/c$.
-Figure 57: The ”light clock” thought experiment.
-Phase of the orthogonal light clock in the accelerated frame.
-Observation made by an observer in the inertial lab frame.
-The phase of the accelerated light clock cannot be measured.
-Figure 58: The ”light clock” thought experiment.
-Observation made by an observer in the inertial lab frame.
-The position along $x$- axis of the accelerated first mirror with respect to the second mirror has no exact objective meaning.
-As viewed from the Lorentz lab frame, the orientation of coordinate axes of the accelerated frame with respect to coordinate axes of the lab frame is regulated by the Wigner rotation.
-The accepted in textbooks assumption of rigidity of accelerated orthogonal light clock is based on the incorrect belief that axes of the accelerated frame and axes of the lab frame are parallel.
-In order to keep the complexity of the discussion to a minimum, we describe the Wigner rotation effect by working only up to the first order in $v/c$.
-Figure 59: The ”light clock” thought experiment.
-Phase of the orthogonal light clock in the accelerated frame.
-Observation made by an observer in the inertial lab frame.
-According to textbooks, it is incorrectly assumed that accelerated and inertial observer have common 3-space.
-The possibility of measurement of the phase is prediction of conventional light clock theory and is obviously absurd from the viewpoint of special relativity.
-We demonstrated that apparent time between clicks is longer for accelerated clock.
-Now, continuing our discussion of the light clock operation in the frame $S$ and $S_{n}$, let us consider the phase of light clocks.
-The word ”phase” in this case has meaning the zigzag shape of the path followed by the light beam.
-We have already (Fig.
-55 - 56) discussed how a clock rate can be measured when the clock is moving.
-Above we found that the rate of moving clock (i.e.
-the time interval between the sending and reception of the light signal) measured by the lab observer is actually spacing of the time marks on the lab screen.
-Is the same statement also true for phase measurements?
-Let us try to demonstrate that it is.
-To see the idea, suppose that the lab observer has two screens and we shall discuss the up-and-down imprinting method (i.e.
-the situation where time marks can be imprinting on the both screens by a moving clock).
-We learned in the Chapter 7 that the orientation of coordinate axes of the accelerated frame with respect to coordinate axes of the Lorentz lab frame is regulated by the Wigner rotation.
-We can show that, as consequence of Wigner rotation, the shape of the zigzag pattern cannot be measured in the lab frame, Fig.
-57.
-The standard presentation of the orthogonal light clock operation is based on the hidden assumption that accelerated and inertial observers have common 3-space.
-According to the conventional theory, when a reference frame at rest is put into motion, then all points in the 3D reference grid have a start to move at same time.
-Consequently, rigidity of the orthogonal light clock in the sense introduced by textbooks cannot be considered as relativistic kinematics property, Fig.
-58.
-Our result (Fig.
-57) is at odds with the prediction from textbooks, Fig.
-59.
-The commonly accepted description of the orthogonal light clock operation does not account for the Wigner rotation, which in our case closely associated with the relativity of simultaneity.
-The possibility of measurement of the phase is prediction of conventional light clock theory and is obviously absurd from the viewpoint of special relativity.
-We want now to describe briefly the mechanism of initialization of the orthogonal light clock in the accelerated frame.
-In order to understand how light is going from mirror to mirror, we have to watch the machinery of the light clock and see what happens when it is moving in the inertial frame.
-We shall take a very simple a kind of light clock initialization, but it will work in principle.
-Although the Fabry-Perot interferometer we have been describing in the Chapter 8 seems to be quite different from the light clock of special relativity , the systems are closely related.
-We consider the simple case when an optical cavity is equipped with plane mirrors.
-Suppose that both mirrors are semitransparent for marking the both screens to measure the time between clock clicks.
-The light clock can be initialized using a short laser pulse (from an external laser in the accelerated frame) coupled in through the top (down) semitransparent mirror.
-Suppose that the initialization laser pulse is impinging onto the top mirror of the light clock at point $A$, Fig.57.
-Special relativity tells us a peculiar thing: that light signal travels from the top mirror to the bottom mirror at the location denoted $B$.
-The position of the top mirror when it meets the returned signal is at $C$.
-This motion ($A\to B\to C$) constitutes the first tick of the clock, and subsequent ticks ($C\to D\to F$) are due to additional reflections of the pulse from the top and the bottom mirrors.
-That’s what special relativity theory says.
-How can that be?
-It is clear that our ordinary intuition are quite wrong when we have deal with different 3D-spaces for different observers in the 4D space-time.
-In the space measurement of one observer there is mixing of space and time, as seen by other.
-Most paradoxes of the moving light clock operation are closely linked to this mixing.
-However, these paradoxes are only apparent, and can be removed by employing the pseudo-Euclidean space-time geometry.
-The position along the $x$-axis of the accelerated bottom mirror with respect to the accelerated top mirror has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this position could be ascertained .
-There is an uncertainty (blurring) of the relative position in the $x$-direction of amount $lv/c$ (due to the uncertainty in the moments of the acceleration), Fig.58.
-10.5 Bibliography and Notes
-1.
-In order to understand magnetic field measurements in the accelerated electron source, we have to watch the machinery of the classical magnetometer and see what happens when it is at rest in the accelerated frame.
-Since it is rather difficult, we shall take a simple a kind of usual magnetometer.
-It is relatively easy to observe the nuclear magnetism by the phenomena of ”nuclear magnetic resonance”.
-A proton resonance apparatus can be used as proton resonance magnetometer.
-We give a general analysis of this type of classical magnetometer independently of the experimental accuracy.
-In general a current loop, which has a magnetic moment $\vec{m}$ in the inertial lab frame in which it is at rest, has an electric dipole moment equal to $\vec{p}=\vec{v}\times\vec{m}/c$ when it is moving with velocity $\vec{v}$ relative to the lab frame.
-The magnetic properties of materials are attributed to atomic current loops.
-Then a magnetized body which is accelerated relative to the lab frame should have an electric polarization relative to the lab frame.
-Thus, in the lab frame, the additional potential energy of a moving particle with the proper magnetic moment $\vec{m}$ in the magnetic field $\vec{B}$ and the electrical field $\vec{E}$ becomes $U_{p}=-\vec{m}\cdot\vec{B}-\vec{p}\cdot\vec{E}$.
-We obtained the electrodynamics equations in accelerated frame using the Galilean transformation (with velocity $-\vec{v}$) of the Maxwell’s equations.
-With the Galilean transformation, the magnetic dipole transformation for the case $\vec{p}=0$, is $\vec{m}_{n}=\vec{m}$, $\vec{p}_{n}=-\vec{v}\times\vec{m}/c$.
-As we know from vector algebra, $(\vec{a}\times\vec{b})\cdot\vec{c}=\vec{a}\cdot(\vec{b}\times\vec{c})$; so our terms are also the same as $(-\vec{v}\times\vec{E})\cdot\vec{m}=-\vec{v}\cdot(\vec{E}\times\vec{m})$, $(-\vec{v}\times\vec{m})\cdot\vec{E}=-\vec{v}\cdot(\vec{m}\times\vec{E})$, as evaluated in the accelerated frame.
-Hence we obtain the additional potential energy equal to zero just as we expected.
-We could not measuring magnetic field $\vec{B}_{n}$ by usual magnetometer.
-Now let us return to observations of an inertial observer.
-In the case of moving electron source the special theory of relativity says that magnetic field $\vec{B}=\vec{v}\times\vec{E}$ can be measured with the usual magnetometer.
-Now that our magnetometer is at rest in the lab frame, so that means that the additional potential energy of a proton with the proper magnetic moment $\vec{m}$ in the magnetic field $\vec{B}$ is equal to $-\vec{m}\cdot\vec{B}\neq 0$.
-The moving electron source in the lab frame generates a magnetic field which can be measured with an usual magnetometer.
+The peculiarity of the collinear geometry where there is a source moving along the same line as the radiated beam is that here the velocity is perpendicular to the plane of radiation wavefront (i.e.
+plane of simultaneity).
+In the case of collinear geometry, a motion of the source, according to special relativity, influences the kinematic terms of the higher (than $v/c$) order only.
+At first sight, Newton’s particle dynamics (in the case of collinear geometry) does not involve relativistic effects at all.
+However, the higher-order effects are there in the assumption that the mass of the moving particle is equal to its relativistic mass.
 14 A Preferred Inertial Frame and the Second Order Experiments
-It should be note that all well known methods to test the special relativity are second order optical measurements.
-The cardinal example is given by the Michelson-Morley interference experiment.
-A close examination of all second order optical experiments inside an inertial frame shows that phenomena appear to be independent of the uniform motion relative to the fixed stars.
-It is not possible to detect the orbital velocity of the earth using second order interference experiments.
-This is consequence of the geometry of space-time.
-A number of other experiments were curried out, which all attempted to measure the velocity of the earth relative to the sun.
-They are not connected with light interference phenomena.
-We shall discuss experiments that compare the rates of different atomic clocks that move together relative to the earth-based frame.
-The question arises whether it is possible to determine experimentally the state of motion of the initial inertial frame $S$ in the accelerated frame $S_{n}$ by means of the time dilation effect for moving reference (atomic) clock.
-Until now, all test of time dilation with atomic clocks in the earth-based frame have involved circle path geometry.
-According to the special relativity, in any clock experiment which probe the influence of the orbital velocity one must use the linear (one-way) path geometry.
-These are experiments of decays of rapidly moving unstable particles.
-However, the modern experimental techniques are not sensitive enough to register the effect of orbital velocity on the average distance traveled by a disintegrated particle.
 14.1 Second Order Optical Experiments and Special Relativity
-The example of the time dilation effect is the resonance absorption of gamma rays in a Mossbauer rotating disk experiments.The measurements of the resonance absorption of gamma rays for the same relative velocity between the source and the observer give a blue shift when the source is at the center, and a red shift when it is at the tip.
-This experiment demonstrates that there is asymmetry in the transverse Doppler effect (1).
-These measurements also show that there is no a frequency shift when the source and observer are at the tip.
-The pulse from the source of one accelerated observer on the circle, which has red shift in the center of a circle due to time dilation, is perceived by diametrically opposite observer without frequency alteration due to his time dilation.
-This will take place despite the fact that each of the observer, having exchanged the pulse, is moving in the lab reference frame $S$ with velocity $v$, in opposing motion.
-At first site the transverse Doppler effect problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers.
-The question arises whether it is possible to determine experimentally the state of motion of the inertial frame $S$ in the accelerated frame $S_{n}$ by means of the Doppler red-blue shift of second order for moving sources.
+An example of the time dilation effect is the resonance absorption of gamma rays in Mossbauer rotating disk experiments.
+The measurements of the resonance absorption of gamma rays for the same relative velocity between the source and the observer give a blue shift when the source is at the center and a red shift at the tip.
+This experiment demonstrates that there is an asymmetry in the transverse Doppler effect  [^54] The first experimental confirmation of the time dilation in the resonance absorption of gamma rays was reported by Champeney, Isaak, and Khan in 1965 [52]..
+These measurements also show no frequency shift when the source and observer are at the tip.
+The pulse from the source of one accelerated observer on the circle, which has a red shift in the center of a circle due to time dilation is perceived by a diametrically opposite observer without frequency alteration due to his time dilation.
+This will occur despite the fact that each of the observers, having exchanged the pulse, is moving in the lab reference frame $S$ with velocity $v$, in opposing motion.
+At first sight, the transverse Doppler effect problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers.
+The question arises whether it is possible to determine experimentally the state of motion of the inertial frame $S$ in the accelerated frame $S_{n}$ using the Doppler red-blue shift of second order for moving sources.
 In particular, the question arises whether it is possible to ascertain the state of motion of the earth relative to the sun-based frame.
-Let us consider the measurements of the resonance absorption of gamma rays in a Mossbauer rotating disk experiments.
+Let us consider the measurements of the resonance absorption of gamma rays in a Mossbauer rotating disk experiment.
 Suppose that an absorber is placed in the center of a disk (which is at rest to the earth-based frame) and an emitter is placed on the edge of the disk.
-The emitter is moving round absorber along a circle with constant velocity $w$.
-In order to keep the mathematical complexity of the discussion to a minimum, we will describe the effects by working only up to the second order in $w/c$ and $v/c$, where $v$ is the earth orbital velocity.
-The frequency of the emitter changes with the velocity $\vec{w}$ in a manner as it suggest itself from the transverse Doppler effect.
+The emitter is moving around the absorber along a circle with constant velocity $w$.
+To keep the mathematical complexity of the discussion to a minimum, we will describe the effects by working only up to the second order in $w/c$ and $v/c$, where $v$ is the earth’s orbital velocity.
+The frequency of the emitter changes with the velocity $\vec{w}$ in a manner as it suggests itself from the transverse Doppler effect.
 Indeed, the inner frequency of the emitter depends on its velocity $\vec{w}$, so that (using Langevin metric in the earth-based frame) $\omega(t_{n})=\omega_{0}\sqrt{1-|\vec{v}+\vec{w}|^{2}}$, then neglecting small terms $\omega(t_{n})=\omega_{0}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(t_{n})/c^{2}]$.
-Since the physical clock is at rest in the earth-based frame, its reading of its proper time is $d\tau=(1-v^{2}/2)dt_{n}$.
+Since the physical clock is at rest in the earth-based frame, the reading of its proper time is $d\tau=(1-v^{2}/2)dt_{n}$.
 We expect therefore the radiation emitted with the constant frequency $\omega_{0}$ from the emitter to arrive in the absorber with a varying frequency $\omega(\tau)=\omega_{0}[1-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(\tau)/c^{2}]$.
@@
 Then the measure of the absorption would be expected to vary with the position of the emitter relative to the disk center.
-In the actual experiment no effect on the absorption was found, when the disk was made to rotate.
+In the actual experiment, no effect on the absorption was found when the disk was made to rotate.
 The negative result of the above experiment can be understood supposing that only a part of the whole frequency fluctuation is considered.
 A more careful analysis shows that one must take into account the change in frequency due to the so-called radial Doppler effect.
-We discuss a setup based on a perpendicular geometry and at first site the radial Doppler effect should be absent.
-We already know from our discussion in previous chapters that there is the aberration of light in the earth-based frame and the radial component of the emitter velocity is equal to $\vec{w}\cdot\vec{v}/c$.
+We discuss a setup based on a perpendicular geometry and at first sight, the radial Doppler effect should be absent.
+We already know from our discussion in previous sections that there is the aberration of light in the earth-based frame and the radial component of the emitter velocity is equal to $\vec{w}\cdot\vec{v}/c$.
 Thus we find that the emitted radiation falling on the disk center with periodically changing frequency $\delta\omega=\omega_{0}\vec{v}\cdot\vec{w}(\tau)/c^{2}$ will be exactly in resonance with the changing inner frequency.
@@
 The frequencies of the light emitted parallel and antiparallel to the beam direction were measured by a spectrograph (at rest in the laboratory).
-Now let us to analyze this experiment.
+Now let us analyze this experiment.
 We discuss a setup based on a parallel geometry.
 The radiation from the moving atoms, giving the Doppler shifted lines, was observed together with the radiation from the resting atoms existing in the same working volume, and giving an anshifted line.
-In that way Ives and Stilwell replaced the difficult problem of the determination of the asymmetry of shift of the ”red” and blue shifted lines with respect to the unshifted line.
-At first site, the orbital motion of the earth relative to the sun can be observed through this type of experiment (2).
-Indeed, if the inner frequency of the atom depends on its velocity $\vec{w}$ so that (using Langevin metric in the earth-based frame) $\omega(t_{n})=\omega_{0}\sqrt{1-|\vec{v}-\vec{w}|^{2}}$ and the frequency of radiation received spectrograph should depend on the orbital velocity.
+In that way, Ives and Stilwell replaced the difficult problem of the determination of the asymmetry of the shift of the ”red” and blue-shifted lines with respect to the unshifted line.
+At first sight, the orbital motion of the earth relative to the sun can be observed through this type of experiment  [^55] In Ives and Stilwell experiment [53] the effect which occurred on the ground of the orbital motion of the earth relative to the sun was within the experimental limitations.
+This situation pertains because of the high value of the emitter’s speed $w\gg v$ involved.
+Presently we give a general analysis of this type of experiment independently of the experimental accuracy..
+Indeed, if the inner frequency of the atom depends on its velocity $\vec{w}$ so that (using Langevin metric in the earth-based frame) $\omega(t_{n})=\omega_{0}\sqrt{1-|\vec{v}+\vec{w}|^{2}}$ and the frequency of radiation received spectrograph should depend on the orbital velocity.
 In rotating disk experiments the effect which occurred on the ground of the transverse Doppler effect is compensated by the other effect.
-In the case of perpendicular geometry we explained of the results of the experiments taking into account the aberration of light together with the radial (i.e.
+In the case of perpendicular geometry, we explained the results of the experiments taking into account the aberration of light together with the radial (i.e.
 classical) Doppler effect.
-We know that in the collinear geometry the aberration of light is absent and one must take into account only the radial Doppler effect.
-If we analyze the geometry of the situation, we find that the observed frequency (i.e frequency of the radiation falling on the detector) is $\omega=\omega_{0}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(t_{n})/c^{2}]/[1-w/(c-\vec{w}\cdot\vec{v}/w)]$, where (according to Langevin metric) $c^{*}=c-\vec{w}\cdot\vec{v}/w$ is the coordinate velocity of light in the earth-based frame.
+We know that in the collinear geometry, the aberration of light is absent and one must take into account only the radial Doppler effect.
+If we analyze the geometry of the situation, we find that the observed frequency (i.e frequency of the radiation falling on the detector) is $\omega=\omega_{0}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})+\vec{v}\cdot\vec{w}(t_{n})/c^{2}]/[1-w/(c+\vec{w}\cdot\vec{v}/w)]$, where (according to Langevin metric) $c^{*}=c+\vec{w}\cdot\vec{v}/w$ is the coordinate velocity of light in the earth-based frame.
 Thus we find just as in the case of the rotating disk experiment that also in measurements of the Doppler red shift of second order for moving atoms the two types of effects caused by the orbital motion of the earth compensate each other and no observable effect remains.
 Usually, such a cancellation is found to stem from a deep underlying principle.
-From the series of negative results like that obtained by Michelson and Morley it seems reasonable to suppose that it is not accidental that various optical experiments trying to determine orbital velocity $v$ remained unsuccessful.
-And we should accept the view that there exist a general law of nature which prevents us to determine $v$ by laboratory experiments.
-Nevertheless, in this case there does not appear to be any such profound implication.
+From the series of negative results like that obtained by Michelson and Morley, it seems reasonable to suppose that it is not accidental that various optical experiments trying to determine orbital velocity $v$ remained unsuccessful.
+And we should accept the view that there exists a general law of nature that prevents us from determining $v$ by laboratory experiments.
+Nevertheless, in this case, there does not appear to be any such profound implication.
 This is a specific feature of the optical experiments of second order.
-Indeed, we already know from our discussion in previous chapters that it is possible to detect the orbital velocity without looking at anything external to the earth using aberration of light phenomenon.
-Now many people who learn special relativity in the usual way find this disturbing.
-The argument that aberration shift must be symmetrical runs some thing like this (3): It is always possible to chose such variables, in which metric of the accelerated frame will be diagonal.
-The validity of Maxwell’s laws in all inertial frames implies that light propagates with the same velocity $c$ in all inertial frames.
-It implies the elimination of the privilege frame, so meaning that the notion of absolute motion does not have physical content at all.
-What is wrong with a principle that the metric supplies all information about the physics of the situation, as described in the given coordinates?
-We already discussed this subject in the Chapter 5.
-At first glance, after the diagonalization of the Langevin metric, we have the symmetry between inertial frames.
-Where does the asymmetry comes from?
-The electrodynamics equations are not supplies all information about the physics of the situation.
-To solve the electrodynamics equations it is necessary to determine the initial conditions.
-We already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization.The time under the Einstein synchronization is readily obtained by introducing the time offset factor.
-This time shift has the effect of rotation of the plane of simultaneity.
-As a consequence of this, the plane wavefront rotates in the accelerated frame after metric diagonalization.
-Now the information about the direction of observer acceleration is recorded in the radiation wavefront orientation.
-The electrodynamics equations are Lorentz covariant.
-This is a consequence of the (pseudo-Euclidean) geometry of space-time.
-However, from the mathematical viewpoint, the Lorentz covariance of electrodynamics equations is insufficient to guarantee the covariant solutions.
-In contrast, the accepted in previous literature incorrect assumption that covariant equations must have covariant solutions has always been considered obvious.
-We will discuss this subject further in the Chapter 12.
+Indeed, we already know from our discussion in previous sections that it is possible to detect the orbital velocity without looking at anything external to the earth using aberration of light phenomenon.
 14.2 A Specific Feature of the Second Order Optical Experiments
-We have seen that the principle of relativity is experimentally confirmed only in restricted sense.
-Presently we find that in the second order experiments the two types of effects caused by the orbital motion of the earth compensate each other and no observable effect remains.
-And we may accept the view that there exist a specific feature of the second order experiments which prevents us to determine the velocity of the earth relative to the sun by laboratory experiments.
+We have seen that the principle of relativity is experimentally confirmed only in a restricted sense.
+Presently we find that in the second order experiments, the two types of effects caused by the orbital motion of the earth compensate each other and no observable effect remains.
+We may accept the view that there exists a specific feature of the second-order experiments that prevents us from determining the velocity of the earth relative to the sun by laboratory experiments.
 Let us verify that this assertion is correct.
@@
 The light from a source is split by a suitable apparatus into two beams and the two overlapping beams produce an interference pattern.
-The two beam originate in the same source.
-Thus, it is not possible to detect the orbital velocity using second order (interference) experiments.
-Phase is 4D invariant (i.e.
-it is simply number).
-It is independent on the chosen inertial frame and on the chosen coordinatization.
+The two beams originate from the same source.
+Thus, it is not possible to detect the orbital velocity using second-order (interference) experiments.
+Phase is 4D invariant.
+It is independent of the chosen inertial frame and the chosen coordinatization.
 This is a consequence of the geometry of space-time.
-In the interference experiments neither the Doppler effect nor the aberration of light exist separately as well defined physical phenomena and only phase is meaningful quantity.
-This discussion clearly show that the special relativity calculations yields the observed null fringe shift in the earth-based frame.
-Until now we have considered only the case in which we have deal with interference pattern as in Michelson interferometer.
-Now we shall more general and study the case of the Doppler shift of second order for moving atoms.
+This discussion clearly shows that the special relativity calculations yield the observed null fringe shift in the earth-based frame.
+The most famous of these interference experiments is one performed by Michelson and Morley.
+Now we shall study the case of the Doppler shift of second order for moving atoms.
 The frequencies of the emitted light were measured by a grating spectrograph.
 There is no specific, important physical difference between diffraction grating and interferometer.
-Grating spectrometer is also based on the interference of incident and reflected light waves.
-A more interesting example is the measurements of the resonance absorption of gamma rays in a Mossbauer rotating disk experiments.
+A more interesting example is the measurements of the resonance absorption of gamma rays in a Mossbauer rotating disk experiment.
 We shall try to understand the effect in a very simple case.
@@
 We inquire about the field at a large distance on the opposite side of the absorber.
-According to the electrodynamics, an electric field is the vector sum of the fields produced by the external source and the fields produced by each charges in the plane of absorber.
-We know that absorber consists of atoms which contains nuclear.
+According to electrodynamics, an electric field is the vector sum of the fields produced by the external source and the fields produced by each charge in the plane of the absorber.
+We know that absorber consists of atoms that contain nuclear.
 When the electric field of the source acts on the nuclear it drives the nuclear charge.
 And moving charges generate a field - they constitute new radiators.
-This means that the two plane waves which interfere are propagate in the same direction.
+This means that the two plane waves which interfere are propagating in the same direction.
 We have in this case interference of the incident and (coherently) forward scattered waves.
 Our model of the nuclear oscillator has some damping force.
-Because we put into take account of damping, the index of refraction is now a complex number and we have destructive interference.
-In this case we deal also with interference phenomenon.
-11.3 Time Dilation Effect for Moving Reference Atomic Clock
-The next thing that we have to discuss is the time dilation effect. We have already discussed this subject in Section 9.3. Here’s what we found there:
-1.
-We have found that the slowly down of the accelerated (with respect to the fixed stars) physical clock does not depend on the choice of reference system, in which this effect is measured.
-In other words, the time dilation effect is absolute, but not relative.
-2.
-According to the asymmetry between (initial) inertial and accelerated frames, there is a remarkable prediction on the theory of special relativity.
-Namely, suppose that frame $S$ is at rest with respect to the fixed stars and frame $S_{n}$ is accelerated with respect to the inertial frame $S$ up to the speed $v$.
-Now two clocks symmetrically accelerated in the frame $S_{n}$ up to the the same speed $v$ but moving in opposite direction.
-However, there is a difference in time dilation of accelerated clocks in the frame $S_{n}$.
-It is clear that clocks have different accelerational histories with respect to the fixed stars.
-We would now like to describe an apparent paradox.
-The argument that all inertial frames are equivalent runs something like this.
-We know that in order to keep a Lorentz coordinates in the accelerated frame we need to perform a metric diagonalization.
-Here the metric in both reference frames will be diagonal, and, according to textbooks, the coordinates in this frames should be related by the Lorentz transformation.
-At first site the problem is complete symmetrical and both frames are equivalent.
-However, there is a real difference between the accelerated inertial frame and the inertial frame without accelerational history.
-Indeed, we should expect the metric tensor to be changed.
-The difference between the inertial and accelerated frames is understandable.
-After diagonalization, the metric tensor in the accelerated frame must abruptly change from the value $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the value $g_{00}=1,g_{01}=0$.
-We must conclude that when we are dealing with (Lorentz ) transformation and this transformation does not preserve continuity of the metric tensor we have no symmetry between inertial frames.
-The question arises whether it is possible to determine experimentally the state of motion of the inertial frame $S$ in the accelerated frame $S_{n}$ by means of the time dilation effect for moving reference (atomic) clock.
-At first site it may appear as if orbital motion of the earth could be observed by time dilation experiments.
-Although in principle this can be done it is necessary to analyze the question in greater detail.
-The proper time interval may be identified with particle lifetime, atomic transition periods, nuclear half-lives, but the crucial point here is that this interval is completely determined by his movement in the initial inertial frame.
-In 1971 Hafele and Keating synchronized four atomic clocks with reference clock at the laboratory, and then flew them around the earth in commercial jets (4) .
-The clocks were flown first from east to west and then from west to east.
-When the clocks returned to the laboratory they were compared with the reference clock.
-The clocks flown in the plane ran slower than the clocks that remainder on the earth.
-The agreement between the observed and predicted discrepancies confirms the predictions of special relativity for the clock acceleration with respect to the earth.
-In the actual experiment no effect of earth orbital velocity on the clock dilation was found.
-[…]
+Because we put into take account damping, the index of refraction is now a complex number and we have destructive interference.
+In this case, we deal also with the interference phenomenon.
 14.3 Wigner Rotation and the Trouton-Noble Experiment
-Now we shall discuss other second order experiments.
+Now we shall discuss other second-order experiments.
 They are not connected with the propagation of light and can be understood in terms of general dynamical consideration.
-The most important of these experiments was the experiment of Trouton and Noble.
+The most important of these experiments was the experiment of Trouton and Noble [54].
 The Trouton-Noble experiment looks for the turning motion of a charged parallel plate capacitor suspended at rest in the frame of the earth in order to measure the earth’s motion relative to the sun (i.e.
@@
 Denoting the angle between $\vec{v}$ and $\vec{r}$ by $\theta$, we find for the absolute value of moment of force $M=e^{2}v^{2}\sin{\theta}\cos{\theta}/(c^{2}r)$.
-In the actual experiment a charged condense was suspended on an elastic string.
+In the actual experiment, a charged condense was suspended on an elastic string.
 The condenser was placed so that $\theta=\pi/4$, i.e.
 so that the line perpendicular to the surface of the condenser plates subtended an angle $\pi/4$ with the supposed direction of the orbital velocity of the earth, the direction of $\vec{v}$.
-This experiment led to negative result.
+This experiment led to negative results.
 i.e.
@@
 The fact that capacitor in the earth-based frame produces no torque is in complete agreement with the special relativity and could be explained in a rather trivial way.
-We can obtain the electrodynamics equations in the earth-based frame using the Galilean transformation (with velocity $-v$) of the Maxwell’s equations.
+We can obtain the electrodynamics equations in the earth-based frame using the Galilean transformation (with velocity $-v$) of Maxwell’s equations.
 With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$.
 Here $\vec{v}$ is the velocity vector of the earth in the sun-based frame.
-Thus we must conclude that this magnetic field produces no torque in the earth-based frame where capacitor is at rest.
-In the experiment discussed we have still a puzzle.
+Thus we must conclude that this magnetic field produces no torque in the earth-based frame where the capacitor is at rest.
 The existence of the magnetic field is responsible for the existence of the magnetic force and this force provides a torque on the charge in the sun-based frame.
-It is important to point out that experiment can be fully accounted for in terms of two inertial frames.
 We have the apparent paradox of different mechanical equations for force and torque governing the motion of a charged particle in different inertial frames; there is a torque and so a time rate of change of (3D) angular momentum in one inertial frame, but not in another relatively moving inertial frame.
-We will show that the real cause of the paradox is - the incorrect assumption that an earth-based observer and an sun-based observer have common 3-space.
+We will show that the real cause of the paradox is - the incorrect assumption that an earth-based observer and a sun-based observer have a common 3-space.
 According to the special relativity, two observers with different trajectories have different 3-spaces.
 The standard presentation of the Trouton-Noble experiment is based on the hidden assumption that $(x_{n},y_{n},z_{n})$ axes of the moving earth-based observer and $(x,y,z)$ axes of the sun-based observer are parallel.
-We demonstrate that only solution of the dynamics equations in covariant form (accounting for the special relativity in the transformation of coordinate axes) gives the explanation of Trouton-Noble experiment in a complete agreement with the special relativity.
-The key idea in resolution of the paradox arises in all usual explanation of the Trouton-Noble experiment is realization of the fact that in the space measurement of the earth-based observer there is mixing of space and time, as seen by the sun-based observer.
+We demonstrate that only the solution of the dynamics equations in covariant form (accounting for the special relativity in the transformation of coordinate axes) gives the explanation of the Trouton-Noble experiment in complete agreement with the special relativity.
+The key idea in the resolution of the paradox that arises in all usual explanations of the Trouton-Noble experiment is the realization of the fact that in the space measurement of the earth-based observer there is the mixing of space and time, as seen by the sun-based observer.
 While naively one should expect the angle between $\vec{v}$ and $\vec{r}_{n}$ in the earth-based frame to be the same angle between $\vec{v}$ and $\vec{r}$ in the sun-based frame it is not the case.
@@
 Thus we find that the moment of force produced by a pair of charges in the earth-based frame is reduced to $M_{n}=e^{2}v^{2}\sin\theta\cos\theta/(2c^{2}r)$.
-[…]
-11.5 Bibliography and Notes
+But we are still not finished!
+We must make a further correction to the moment of force in the earth-based frame.
+A more careful analysis shows that we must take into account the change in $v_{\perp}=v\sin\theta$ due to the Wigner rotation.
+As the force acts in the direction $\vec{r}$ there is a component of the charge acceleration perpendicular to the velocity $\vec{v}$.
+This acceleration component produces a rate of rotation of the proper (earth-base) frame axes in the sun-based frame.
+The rate at which the Wigner rotation occur is given $d\vec{\Phi}/dt=[v^{2}/(2c^{2})](\vec{v}/v^{2})\times d\vec{v}/dt$.
+The direction of the earth-based frame rotation in the sun-based frame is the same as the direction of the velocity rotation in the sun-based frame.
+We already have all the formulas that we need.
+For $d\theta/dt$ we get $d\theta/dt=[v^{2}/(2c^{2})]e^{2}\sin\theta/(mvr^{2})$, where $m$ is the mass of charge.
+So we have that the second correction to the moment of force is $mrdv_{\perp}/dt=v\cos\theta d\theta/dt=e^{2}v^{2}\sin\theta\cos\theta/(2c^{2}r)$.
+The two types of effects compensate the influence of the magnetic field in the sun-base frame and the moment of force produced by a pair of charges in the earth-based frame vanishes.
+14.4 Time Dilation Effect for Moving Reference Atomic Clock
+The next thing that we have to discuss is the time dilation effect. We have already discussed this subject in the Chapter 12. Here’s what we found there:
 1.
-The measurements of the resonance absorption of gamma rays in a Mossbouer rotating disk experiment [43] for the same relative velocity between the source and the observer give a blue shift when the source is at the center, and a red shift when it is at the tip.
+We have found that the slow down of the accelerated (with respect to the fixed stars) physical clock does not depend on the choice of the reference system, in which this effect is measured.
+In other words, the time dilation effect is absolute, but not relative.
 2.
-In Ives and Stilwell experiment [44] the effect which occurred on the ground of the orbital motion of the earth relative to the sun was within the experimental limitations.
-This situation pertains because the high value of the emitters speed $w\gg v$ involved.
-Presently we give a general analysis of this type of experiments independently of the experimental accuracy.
-3.
-Many physicists tend to think that the validity of Maxwell’s laws in all inertial frames implies the elimination of the privilege frame.
-To quote e.g.
-Dieks [45]: ” As we already mentioned, it is a basic principle of the special theory of relativity that the line element $ds$ supplies all information about the physics of the situation, as described in the given coordinates.”
-4.
-The first experimental confirmation of the time dilation predicted in the circle path geometry with atomic clocks and aircraft was reported by Hafele and Keating in 1971 [46].
-A very interesting result have been obtained by GPS team.
+According to the asymmetry between (initial) inertial and accelerated frames, there is a remarkable prediction in the theory of special relativity.
+Namely, suppose that frame $S$ is at rest with respect to the fixed stars and frame $S_{n}$ is accelerated with respect to the inertial frame $S$ up to the speed $v$.
+Now two clocks symmetrically accelerated in the frame $S_{n}$ up to the the same speed $v$ but moving in opposite direction.
+However, there is a difference in time dilation of accelerated clocks in the frame $S_{n}$.
+It is clear that clocks have different accelerational histories with respect to the fixed stars.
+The question arises whether it is possible to determine experimentally the state of motion of the inertial frame $S$ in the accelerated frame $S_{n}$ using the time dilation effect for moving reference (atomic) clock.
+At first sight, it may appear as if orbital motion of the earth could be observed by time dilation experiments.
+Although in principle this can be done it is necessary to analyze the question in greater detail.
+The proper time interval may be identified with particle lifetime, atomic transition periods, and nuclear half-lives.
+Still, the crucial point here is that this interval is completely determined by clock movement in the initial inertial frame.
+In 1971 Hafele and Keating synchronized four atomic clocks with reference clock at the laboratory and then flew them around the earth in commercial jets [^56] The first experimental confirmation of the time dilation predicted in the circle path geometry with atomic clocks and aircraft was reported by Hafele and Keating in 1971 [55].
+A very interesting result has been obtained by the GPS team.
 The first example of the different inertial frames having the same accelerational histories is the global positioning system (GPS).
@@
 There are no asymmetries among the GPS satellites.
-That is, times are equal in every GPS satellites [47].
-We see that the presence of some non zero relative velocity cannot be a cause for the effect of time dilation.
-In this experiment no effect of earth orbital velocity on the clock dilation was found.
-The negative result of the GPS experiment can be understood remembering that in this situation the clocks also are moving around the earth (i.e.
-around the reference clock) along the circular path and no observable effect of orbital velocity remains.
+That is, times are equal in every GPS satellite [56].
+We see that the presence of some non-zero relative velocity cannot be a cause for the effect of time dilation.
+In this experiment, no effect of earth orbital velocity on the clock dilation was found.
+The negative result of the GPS experiment can be understood by remembering that in this situation the clocks also are moving around the earth (i.e.
+around the reference clock) along the circular path and no observable effect of orbital velocity remains..
+The clocks were flown first from east to west and then from west to east.
+When the clocks returned to the laboratory they were compared with the reference clock.
+The clocks flown in the plane ran slower than the clocks that remained on the earth.
+The agreement between the observed and predicted discrepancies confirms the predictions of special relativity for the clock acceleration with respect to the earth.
+In the actual experiment, no effect of earth’s orbital velocity on the clock dilation was found.
+[…]
 15 The Principle of Relativity and the Modern Cosmology
-12.1 The Relativity Principle of Special Relativity
-In the last few chapters we have treated the essential ideas necessary for an understanding of the aberration of light phenomena.
-We have not had to worry yet about the fact that the source-observer asymmetry associated with the stellar aberration phenomenon at first glance contradicts the principle of relativity.
-The principle of relativity which we owe to Poincare, who first coined the term, denied the possibility for an observer partaking in a uniform motion relative to the fixed stars of discovering, by any measurement, such a motion, under the assumption, of course, that one does not look outside.
-Now let us to analyze the stellar aberration experiment.
-Perhaps we should say: ” In the case of stellar aberration, the peculiarity of the aberration of light measurements is that the accelerated (earth-based) observer looks outside to the stars.” But no!
-The motion of the stars with respect to the earth is never followed by any aberration.
-The aberration shift (as inferred from astronomical observations) exists even in the case when star moves with the same velocity as the earth.
-We cannot remain with the framework of ”looking outside to the star” when considering the earth-based observations of the stellar aberration.
-In the Section 8.3 we discussed a simple scaling model for the stellar aberration.
+Above we have treated the essential ideas necessary for an understanding of the aberration of light phenomena.
+At first glance, the source-observer asymmetry associated with the stellar aberration phenomenon contradicts the principle of relativity.
+This principle which we owe to Poincare, who first coined the term, denied the possibility for an observer partaking in a uniform motion relative to the fixed stars of discovering, by any measurement, such a motion, under the assumption, of course, that one does not look outside.
+Now let us analyze the stellar aberration experiment.
+Perhaps we should say: ” In the case of stellar aberration, the peculiarity of the aberration of light measurements is that the accelerated (earth-based) observer looks outside to the stars.” However, the motion of the stars with respect to the earth is never followed by any aberration.
+The aberration shift (as inferred from astronomical observations) exists even in the case when a star moves with the same velocity as the earth.
+We cannot remain within the framework of ”looking outside to the star” when considering the earth-based observations of the stellar aberration.
+In Chapter 10 we discussed a simple scaling model for stellar aberration.
 We obtained a condition for optical similarity between the aberration of light from a distant star and from the earth-based incoherent source.
-We find that without looking outside, it is possible to determine the speed of the earth around the sun by means of aberration of light measurements.
-No one has ever done such experiment with earth-based incoherent source, but we know would happen from the astronomical observations.
-It should be note that all well known methods to test the special relativity are second order measurements .
-The cardinal example is given by the Michelson-Morley experiment.
-A close examination of all second order experiments inside an inertial frame shows that phenomena appear to be independent of the uniform motion relative to the fixed stars.
-From the series of negative results like that obtained by Michelson and Morley it seems reasonable to suppose that it is not accidental that various optical experiments trying to determine orbital velocity remained unsuccessful.
-And we should accept the view that there is exist a general law of of nature which prevents us to determine orbital velocity by laboratory experiments.
-Nevertheless, in this case there does not appear to be any such profound implication.
-This is specific feature of the optical experiments of second order.
-All these methods are based on phenomenon of interference.
-There is a fact governing all interference phenomena.
-The law is called the relativistic invariance of phase (i.e.
-it is simply number).
-It is not change under any reference frame transformations.
-It is not possible to detect the orbital velocity using second order interference experiments.
-This is consequence of the geometry of space-time.
-We conclude the following: The usual formulation of the theory of relativity deals with formulated above principle of irrelevancy of velocity, but it should be understood only in a limiting sense when we are dealing with second order optical measurements.
-There is no conflict between the fundamental structure of special theory of relativity on the one hand, and the aberration of light phenomena.
+We find that without looking outside, it is possible to determine the speed of the earth around the sun using aberration of light measurements.
+There is no conflict between the fundamental structure of the special theory of relativity on the one hand, and the aberration of light phenomena.
 First, from what is said in the previous chapters it should be clear that special relativity does not require the formulated above principle of irrelevancy of velocity.
-Principle of special relativity applies to physical laws, not to physical facts.
-Not only in Einstein writings, but in every textbook on special theory of relativity can be found the formulation of principle of special relativity as follows: the laws of physics are the same in all inertial frames (1).
+The principle of special relativity applies to physical laws, not to physical facts.
 The special principle of relativity says that the same laws should hold in all inertial frames.
-We interpret this to mean that the laws can be expressed by equations that have the same form in all inertial frames.
-According to special principle of relativity all inertial frames are equivalent with regard to physical laws, not with regard to physical facts.
-By a physical fact in this context we mean properties of existing objects - in our case of interest the properties of incoherent source.
-By mean a physical law we mean for example electrodynamics equations.
-Principle of special relativity is irrelevancy of velocity with regard to physical laws, not with regard to anything.
-We point that the essence of the special theory of relativity consists in the following postulate: all physical process proceed in four-dimentional space-time, the geometry of which is pseudo-Euclidean.
-But give pseudo-Euclidean geometry of space-time does not dictate the irrelevancy of velocity with regard to anything.
-From a mathematical viewpoint, the argument looks something like the following.
-The inertial frames are characterized by motion at constant velocity relative to the fixed stars and different inertial frames are related by Lorentz boosts.
-Relativistic symmetry is usually identified with Lorentz symmetry, i.e.
-with symmetry under Lorentz boosts.
-At the same time it is common knowledge that Lorentz boosts alone do not make up a group.
-Lorentz boosts do not form a group due presence of the Wigner rotation.
-In general, two successive Lorentz boosts are not equivalent to boost.
-The product of two Lorentz boosts in different directions is equal to the product of a pure boost and a spatial rotation, the Wigner rotation.
-We must conclude that when we are dealing with transformations and set of these transformations does not form a group we have no symmetry between inertial frames.
-There is a real difference between the accelerated inertial frame and the inertial frame without accelerational history.
-It is comes out quite natural that this difference is closely associated with the Wigner rotation phenomenon.
-The existence in nature of something corresponding to the concept of absolute acceleration entails the existence in nature of something corresponding to the concept of absolute velocity.
-If we look at the situation carefully we see that the time dilation effect problem in the accelerated frame demonstrate the essential asymmetry between accelerated and initial ( i.e.
+We interpret this to mean that the laws can be expressed by equations that have the same form in all inertial frames  [^57] In the special theory of relativity the form-invariant transformations between inertial frames are Lorentz transformations [57].
+For a general discussion about the Principle of Relativity we suggest reading the paper [58].
+.
+According to the special principle of relativity, all inertial frames are equivalent with regard to physical laws, not with regard to physical facts.
+The existence in nature of the concept of absolute acceleration entails the existence in nature of the concept of absolute velocity.
+If we look at the situation carefully we see that the time dilation effect problem in the accelerated frame demonstrates the essential asymmetry between accelerated and initial ( i.e.
 without acceleration history) inertial frame.
-The proper time of any particle moving arbitrary in the initial inertial frame is always slower than the physical time in this initial inertial frame.
+The proper time of any particle moving arbitrarily in the initial inertial frame is always slower than the physical time in this initial inertial frame.
 It is a fundamental property of the pseudo-Euclidean space-time.
-There is no relativity of the uniform motions and we can determine the particle absolute velocity in the initial inertial frame.
+There is no relativity of the uniform motions and we can determine the particle’s absolute velocity in the initial inertial frame.
 The formalism of classical physics is associated with absolute space structure.
-This structure must be associated with some ”substrance”, the ether, which incidentally provides an absolute rest frame.
+This structure must be associated with some ”substance”, the ether, which incidentally provides an absolute rest frame.
 So is space-time the new ether?
@@
 It clearly does provide an ”absolute rest reference frame”.
-The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of unified space-time model, with no a priori splitting of space from time.
+The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of a unified space-time model, with no a priori splitting of space from time.
 The initial inertial frame is not physically identical to the absolute space.
@@
 Newtonian mechanics is invariant under Galilean transformations.
-However, from the mathematical viewpoint, the relativistic Lorentz transformation is qualitatively different from the non-relativistic Galilean one.
+However, from the mathematical viewpoint, the Lorentz transformation is qualitatively different from the Galilean one.
+It is common knowledge that Lorentz boosts alone do not make up a group.
 The set of Galileo boosts form a group.
-So we see that principle of Galilean relativity states that all inertial frames are equivalent.
-Equivalence between two frames means that none of them preferred to the other.
-Galileo’s version of relativity principle was incorrectly extended to the whole of physics.
-The relativity principle does not hold for the whole range of validity of the Lorentz covariant physical laws and the Lorentz covariance is not a fundamental symmetry of physics (because the accelerated inertial frame is connected with the initial inertial frame by the Galilean transformation, but not the Lorentz one).
-12.2 Aberration and the Modern Cosmology
-Now we want to discuss a serious trouble - the failure of the aberration of light theory presented above.
-There are difficulties associated with the ideas of modern cosmology.
-The difficulty we speak of is associated with the concept of source-observer asymmetry, when applied to the motion relative to the cosmic microwave background (CMB).
-One may say that perhaps there is no use worrying about these difficulties since there are so many things about universe that we still don’t understand.
-It is obvious that we live in a universe ruled by the unknown.
-Indeed, dark energy and dark matter together contribute to almost 96 percents of the total energy-mass of the universe and we have no knowledge of what dark matter (or dark energy) is.
+So we see that the principle of Galilean relativity states that all inertial frames are equivalent.
+Equivalence between two frames means that none of them is preferred to the other.
+The reason for this (Galilean group) symmetry is simply that equations of motion do not explicitly depend on velocity.
+The internal dynamics of the system do not change under the Galilean boost.
+Galileo’s version of the relativity principle was incorrectly extended to the whole of physics.
+The relativity principle does not hold for the whole range of validity of the Lorentz covariant physical laws and the Lorentz covariance is not a fundamental symmetry of physics.
+In fact, since the metric tensor of space-time must be a continuous quantity, the accelerated inertial frame is connected with the initial inertial frame by the Galilean transformation, but not the Lorentz one.
+The Galilean transformations do not preserve the form-invariance of Maxwell’s equations under a change of inertial frame.
+The reason for this absence of Galileo’s relativity is simply that electrodynamics equations explicitly depend on velocity.
+Therefore, according to special relativity, there is a real difference between the accelerated (with respect to the fixed stars) inertial frame and the inertial frame without accelerational history.
+It should be noted that this difference is closely related to the inertial forces.
+This is the basic content of the equivalence principle.
+The relation between inertia and gravity is a subject that needs to be understood more deeply.
+Now we want to discuss a serious difficulty associated with the ideas of modern cosmology.
+The difficulty we speak of is associated with the concept of source-observer asymmetry when applied to the motion relative to the cosmic microwave background (CMB).
+One may say that perhaps there is no use worrying about these difficulties since there are so many things about the universe that we still don’t understand.
+We live in a universe ruled by the unknown.
+Indeed, dark energy and dark matter together contribute to almost 96 percent of the total energy-mass of the universe and we have no knowledge of what dark matter (or dark energy) is.
 It is important to realize that when we try to talk about working assumptions in modern cosmology, they are approximate and will change.
-Therefore what we discuss in this section will not be accurate in a certain sense.
 We cannot go into details of the modern cosmology principles at this time.
@@
 According to the cosmological principle, the universe should appear isotropic, without any preferred direction, to a comoving observer, having no peculiar motion relative to the cosmic fluid of the expanding universe.
-However, a peculiar motion of such an observer might introduce a dipole anisotropy in the observed properties of a class of objects and which in turn, might be exploited to infer the peculiar velocity of the observer.
-For instance the CMB radiation shows an anisotropy distribution.
+However, a peculiar motion of such an observer might introduce a dipole anisotropy in the observed properties of a class of objects which in turn, might be exploited to infer the peculiar velocity of the observer.
+For instance, the CMB radiation shows an anisotropy distribution.
 The CMB dipole is almost ubiquitously assumed to be of kinematical origin, i.e.
@@
 It is widely believed that the dipole anisotropy is produced by the Doppler effect due to the relative motion between the earth, i.e.
-an observer, and the frame where cosmic microwave background looks nearly isotropic.
+an observer, and the frame where the cosmic microwave background looks nearly isotropic.
 By subtracting the dipole, the CMB is defined as the rest frame for the universe.
 The observed dipole indicates that the solar system is moving at 370 km/s relative to the observed universe in the direction of galactic longitude $l=264^{o}$ and latitude $b=48^{O}$.
-This is quite far from the galactic rotation direction 250 km/s towards $l=90^{o}$ and $b=0$.
-The motion relative to the cosmic microwave background results from the sum of many components of velocity due to gravitational attraction of various mass concentrations.
-The existence of clusters and super clusters of galaxies and our motion is a natural consequence of the large scale organization of matter.
-The peculiar velocity consists of the five vector contributions: the motion of the earth around the sun (30 km/s), the hypothetical circular motion around the our galaxy (250 km/s), the motion of our galaxy in the local group, and the motion of the local group with respect to the cosmic microwave background.
-The origin of the velocity of the local group is still uncertain and has been under discussion over past two decades.
-This peculiar velocity is believed to be generated by the spatial inhomogeneties of mass (mainly dark matter) distribution in nearly large scale structures.
+This is quite far from the galactic rotation direction of 250 km/s towards $l=90^{o}$ and $b=0$.
+The motion relative to the cosmic microwave background results from the sum of many components of velocity due to the gravitational attraction of various mass concentrations.
+The existence of clusters and super clusters of galaxies and our motion is a natural consequence of the large-scale organization of matter.
+The peculiar velocity consists of the five vector contributions: the motion of the earth around the sun (30 km/s), the hypothetical circular motion around our galaxy (250 km/s), the motion of our galaxy in the local group, and the motion of the local group with respect to the cosmic microwave background.
+The origin of the velocity of the local group is still uncertain and has been under discussion over the past two decades.
+This peculiar velocity is believed to be generated by the spatial inhomogeneities of mass (mainly dark matter) distribution in nearly large-scale structures.
 The velocity of the local group with respect to the cosmic microwave background is estimated to be 500 km/s.
-Now let us see how aberration of light various with the speed relative to the rest frame of the universe.
+Now let us see how aberration of light varies with the speed relative to the rest frame of the universe.
 Consider first the textbook explanation.
 It is very easy to understand how the aberration of light effect comes about from the point of view of the conventional aberration of light theory.
-The existence of an aberration of the line of sight by motion can be recognized by considering an observer in a car driving through rainstorm: raindrops falling vertically appear to be obliquely.
+The existence of an aberration of the line of sight by motion can be recognized by considering an observer in a car driving through a rainstorm: raindrops falling vertically appear to be oblique.
 An important application of the Galilean velocity transformation law is provided by stellar aberration, the change in the apparent direction of a star caused by the earth’s motion around the sun.
@@
 What about the motion relative to the CMB rest frame?
-According to textbooks, if the earth motion were uniform, the aberration effect would be undetectable since the ”true” direction of the star is unknown.
+According to textbooks, if the earth’s motion were uniform, the aberration effect would be undetectable since the ”true” direction of the star is unknown.
 Indeed, who can say where a given star should be?
@@
 But, according to conventional theory, the aberration of light phenomenon does not exist in an aberration of light experiments using an earth-based light source.
-We now in position to understand the nature of our difficulty with the motion relative to the CMB.
+We are now in a position to understand the nature of our difficulty with the motion relative to the CMB.
 It is believed that the dipole anisotropy is produced by the Doppler effect due to acceleration (during the billions of years) of the solar system with respect to the rest frame of the universe.
-This acceleration is believed to be generated by the spatial inhomogenetics of (Dark matter) mass distribution in nearly large scale structures.
+This acceleration is believed to be generated by the spatial inhomogeneities of (Dark matter) mass distribution in nearly large-scale structures.
 A correct solution of the aberration problem in the earth-based frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric plays a fundamental role in the non-inertial kinematics of a light ( or relativistic particle ) beam produced by the earth-based source.
-According to the relativistic theory of aberration presented in this book, the proper rotation of the earth on its axis should produce a corresponding shift of the image.
-The aberration shift will depend also upon the value of $v_{\perp}^{CMB}$, the component of the solar system velocity (relative to the CMB) perpendicular to the earth rotation axis.
-The rotation of the earth produces aberration in an amount larger enough to be taken into account in precise observation work using electron microscope as an earth-based particle source.
-Experimental results show (see the Chapter 8) that the image shift is quite close to the theoretical prediction for the (30 km/s) orbital velocity and clearly indicate that the signal associated with motion (370 km/s) relative to the CMB does not exist.
+According to the relativistic theory of aberration presented in this article, the proper rotation of the earth on its axis should produce a corresponding shift of the image.
+The aberration shift will depend also upon the value of $v_{\perp}^{CMB}$, the component of the solar system velocity (relative to the CMB) perpendicular to the earth’s rotation axis.
+The rotation of the earth produces aberration in an amount larger enough to be taken into account in precise observation work using an electron microscope as an earth-based particle source.
+Experimental results show (see Chapter 10) that the image shift is quite close to the theoretical prediction for the (30 km/s) orbital velocity and clearly indicate that the signal associated with motion (370 km/s) relative to the CMB does not exist.
 The simplest explanation is that the CMB dipole might be of non-kinematical origin.
-More recent observations of astronomers also cast doubt on the CMB dipole, being the ultimate representatives of the solar peculiar velocity.
+More recent observations by astronomers also cast doubt on the CMB dipole, being the ultimate representative of the solar peculiar velocity.
 In recent years observations have emerged hinting at an anisotropic universe.
 The discovery of the preferred direction in the universe was serendipitous.
-Until very recently the velocity of the solar system in the rest frame of universe is inferred from CMB temperature dipole anisotropy.
-Obviously, an independent measurements of this velocity is needed to fully establish the kinematical origin of the dipole.
+Until very recently the velocity of the solar system in the rest frame of the universe is inferred from CMB temperature dipole anisotropy.
+Obviously, an independent measurement of this velocity is needed to fully establish the kinematical origin of the dipole.
 Another such quantity that could be employed to look for departures from isotropy is the angular distribution of distant radio sources in the sky.
 This could provide an independent check on the interpretation of CMB dipole.
-The radio data clearly indicate that significantly larger dipole exist in the rest frame of the radio galaxies.
-While the velocity of the solar system inferred from the CMB temperature dipole anisotropy is 370 km/s, the radio dipole measurements finds the speed of motion to be around 1000 km/s (i.e.
-to be around three times larger that that of CMB).
-From the Hubble diagram of quasars, motion of the solar system is derived, which out to be the largest value ever found, 8000 km/s (2).
-On the other hand, a common direction for all these dipoles, determined from completely independent surveys by different groups employing different techniques, indicate that these dipoles are not resulting from some systematics in the observations or the data analysis, but could instead suggest an inherent anisotropy.
-This is totally unexpected in a standard model of universe.
+The radio data clearly indicate that significantly larger dipoles exist in the rest frame of the radio galaxies.
+While the velocity of the solar system inferred from the CMB temperature dipole anisotropy is 370 km/s, the radio dipole measurements find the speed of motion to be around 1000 km/s (i.e.
+to be around three times larger than that of CMB).
+From the Hubble diagram of quasars, the motion of the solar system is derived, which is out to be the largest value ever found, $8\cdot 10^{3}$ km/s  [^58] Recent observations of the cosmic dipole anisotropies is revolutionized our understanding of the universe.
+The various dipoles, including CMB dipole, all pointing along the same direction, suggest a preferred direction in the universe, raising thereby uncomfortable questions about the cosmological principle, the basis of the standard model in modern cosmology [59, 60, 61]..
+On the other hand, a common direction for all these dipoles, determined from completely independent surveys by different groups employing different techniques, indicates that these dipoles are not resulting from some systematics in the observations or the data analysis, but could instead suggest an inherent anisotropy.
+This is totally unexpected in a standard model of the universe.
 We have a preferred direction, aligned with the CMB dipole, in the universe.
@@
 There is a difficulty with such some sort of an ”axis” of the universe which, in turn, would be against the cosmological principle.
-Three independent dipole vectors pointing along the same particular direction could imply an anisotropic universe, violating the cosmological principle, a cornerstone of the modern cosmology(2).
-Let us be conservative and say that there are two kinds of the earth velocity that the total velocity with respect to the initial (i.e.
+Three independent dipole vectors pointing along the same particular direction could imply an anisotropic universe, violating the cosmological principle, a cornerstone of modern cosmology.
+Let us be conservative and say that there are two kinds of the earth velocity.
+The total velocity with respect to the initial (i.e.
 privilege ) inertial frame could be the sum of the orbital velocity and the velocity of the sun.
 The velocity of the sun consists the two contributions: the circular motion around our galaxy and the motion of our galaxy.
-In earth-based experiments where we measure the earth velocity with respect to the privilege frame by seeing aberration of light (or particles), we are measuring the recorded (in the accelerational history) velocity.
-So the velocity with respect to the privilege frame consists of the two contributions: a recorded motion plus an unrecorded motion.
-We know that there is definitely a recorded (orbital) motion, and we have a formula for it.
-It is therefore impossible to get all the earth velocity to be record in the accelerational history in the way we hoped.
-It is not a legal theory if we have nothing but modern cosmology.
-Something else has to be added.
+In earth-based experiments where we measure the earth’s velocity with respect to the privileged frame by seeing the aberration of light (or particles), we are measuring the recorded (in the accelerational history) velocity.
+So the velocity with respect to the privilege frame consists of two contributions: a recorded motion plus an unrecorded motion.
+We know that there is definitely a recorded (orbital) motion.
+It is therefore impossible to get all the earth’s velocity to be recorded in the accelerational history in the way we hoped.
+Clearly something else has to be added.
 We would like to think a little more about why the earth is moving with respect to the initial inertial frame.
@@
 So it may help our understanding if we look a little more closely at where the forces come from.
-We must say immediately that orbital motion is result of non-gravitational forces.
+We must say immediately that orbital motion is the result of non-gravitational forces.
 It is well known that a full 98 percent of all the angular momentum in the solar system is concentrated in the planets, yet a staggering 99.8 percent of all the mass in our solar system is in our sun.
-Perhaps the first scientifically respectable theory of the origin of the solar system was given by Hoyle (1960).
-He invoked the action of magnetic field to transfer angular momentum from the central body, the sun, to the ejected matter which eventually formed the planets.
-In contrast, the unrecorded motion around our galactic and the motion of our galactic is result of the action of gravity.
+Perhaps the first scientifically respectable theory of the origin of the solar system was given by Hoyle (1960) [62].
+He invoked the action of the magnetic field to transfer angular momentum from the central body, the sun, to the ejected matter which eventually formed the planets.
+In contrast, the unrecorded motion around our galactic and the motion of our galactic is the result of the action of gravity.
 We started by talking about the gravitational interaction of the earth with the sun.
 We believe the theory of gravity is so much that we allow it to tell us about the force of one galactic to another.
-Perhaps we are making to great an extrapolation of our limiting knowledge of gravity to galactic scales.
+Perhaps we are making too great an extrapolation of our limiting knowledge of gravity to galactic scales.
 Perhaps the entire difficulty is that a modification of gravity may be responsible for unrecorded velocities.
-We may someday find out one things we don’t understand today, for example dark matter, can, in fact be explained as modification of the gravitation theory.
-Today it doesn’t seem likely, but no one can say for sure.
-There are so many things about universe that we don’t understand.
-12.3 Bibliography and Notes
-1.
-By ”the same” is means form-invariance: the mathematical relations between the physical quantities remain identical after a coordinate transformation has been performed on these quantities.
-In special theory of relativity the (form-invariant) transformations between inertial frames are Lorentz-transformations [48].
-2.
-Recent observations of the cosmic dipole anisotropies is revolutionized our understanding of the universe.
-Attempts to recover the CMB dipole from counts of later universe sources such as radio galaxies and quasars, which are assumed to be in the CMB frame, largely agree that CMB dipole direction is recovered, but not the magnitude.
-The various dipoles, included CMB dipole, all pointing along the same direction, suggest a preferred direction in the universe, raising thereby uncomfortable questions about the cosmological principle, the basis of the standard model in modern cosmology [49, 50, 51].
-.
+There are so many things about the universe that we don’t understand.
 16 Relativistic Particle Dynamics
-In previous chapters we considered the kinematics of the theory of relativity, which concerns the study of the four vectors of positions, velocity and acceleration.
-Kinematics studies trajectories as geometrical objects, independently of their causes.
-This means that it is not possible to predict the trajectory of a particle evolving under a given dynamical field using just a kinematic treatment.
-In dynamics we consider the effect of interaction on motion.
 16.1 Manifestly Covariant Particle Dynamics
@@
 Relying on the geometric structure of Minkowski space-time, one can define the class of inertial frames and can adopt a Lorentz frame with orthonormal basis vectors for any given inertial frame.
-In any Lorentz coordinate system the law of motion becomes
+In any Lorentz coordinate system, the law of motion becomes
 $m\frac{d^{2}x_{\mu}}{d\tau^{2}}=eF^{\mu\nu}\frac{dx_{\nu}}{d\tau}~{},$ (21)
@@
 The electromagnetic field is described by a second-rank, antisymmetric tensor with components $F^{\mu\nu}$.
-The coordinate-independent proper time $\tau$ is a parameter describing the evolution of physical system under the relativistic laws of motion, Eq.
+The coordinate-independent proper time $\tau$ is a parameter describing the evolution of the physical system under the relativistic laws of motion, Eq.
 (21).
-The covariant equation of motion for a relativistic charged particle under the action of the four-force $K_{\mu}=eF^{\mu\nu}dx_{\nu}/d\tau$ in the Lorentz lab frame, Eq.(20), is a relativistic ”generalization” of the Newton’s second law.
-The three-dimensional Newton second law $md\vec{v}/dt=\vec{f}$ can always be used in the instantaneous Lorentz comoving frame.
-Relativistic ”generalization” means that the previous three independent equations expressing Newton second law are be embedded into the four-dimensional Minkowski space-time (1).
+The covariant equation of motion for a relativistic charged particle under the action of the four-force $K_{\mu}=eF^{\mu\nu}dx_{\nu}/d\tau$ in the Lorentz lab frame, Eq.(21), is a relativistic ”generalization” of Newton’s second law.
+The three-dimensional Newton’s second law $md\vec{v}/dt=\vec{f}$ can always be used in the instantaneous Lorentz comoving frame.
+Relativistic ”generalization” means that the previous three independent equations expressing Newton’s second law are embedded into the four-dimensional Minkowski space-time  [^59] The idea of embedding is based on the fact that the usual Newton’s second law can always be used in any Lorentz frame where the particle, whose motion we want to describe, is at rest [63].
+In other words, if an instantaneously comoving Lorentz frame is given at some instant, one can precisely predict the evolution of the particle in this frame during an infinitesimal time interval.
+In geometric language, Newton’s law is strict on a hyperplane perpendicular to the world line.
+However, the hyperplane tilts together with its normal $u_{\mu}$ as one moves along the world line.
+For the embedding, we need an operator $\hat{P}_{\perp}$ that continually projects vectors of Minkowski space on hyperplanes perpendicular to the world line.
+The desired operator is $(\hat{P}_{\perp})_{\mu\nu}=\eta_{\mu\nu}-u_{\mu}u_{\nu}/u^{2}$ [63].
+The presence of the projector operator $\hat{P}_{\perp}$ suggests that we have only three independent equations..
 The immediate generalization of $md\vec{v}/dt=\vec{f}$ to an arbitrary Lorentz frame is Eq.(21), as can be checked by reducing to the rest frame.
@@
 Having written down the motion equation in a 4-vector form, Eq.(21), and determined the components of the 4-force, we satisfied the principle of relativity for one thing, and, for another, we obtained the four components of the equation of particle motion.
-This is covariant relativistic generalization of the three dimensional Newton’s equation of motion which is based on particle proper time as the evolution parameter.
-We next wish to describe a particle motion in the Lorentz lab frame using the lab time $t$ as evolution parameter.
+This is a covariant relativistic generalization of the three-dimensional Newton’s equation of motion based on particle proper time as the evolution parameter.
+We next wish to describe a particle motion in the Lorentz lab frame using the lab time $t$ as an evolution parameter.
 First, we examine the reasoning presented in textbooks.
@@
 $\mathcal{E}^{2}/c^{2}-|\vec{p}|^{2}=mc^{2}~{}.$ (24)
-According to the non-covariant (3+1) approach we seek for the initial value solution to these equations.
-Using explicit expression for Lorentz force we find that the three equations Eq.(21) automatically imply the constraint Eq.(23), once this is satisfied initially at $t=0$.
+According to the non-covariant (3+1) approach we seek the initial value solution to these equations.
+Using the explicit expression for Lorentz force we find that the three equations Eq.(22) automatically imply the constraint Eq.(24), once this is satisfied initially at $t=0$.
 In the (3+1) approach, the four equations of motion ”split up” into (3+1) equations and we have no mixture of space and time parts of the dynamics equation Eq.(21).
@@
 One could expect that the particle’s trajectory in the lab frame, following from the previous reasoning $\vec{x}(t)$, should be identified with $\vec{x}_{cov}(t)$.
-However, paradoxical result are obtained by doing so.
+However, paradoxical results are obtained by doing so.
 In particular, the trajectory $\vec{x}(t)$ does not include relativistic kinematics effects.
 In the non-covariant (3+1) approach, the solution of the dynamics problem in the lab frame makes no reference to Lorentz transformations.
-This means that, for instance, within the lab frame the motion of particles in constant magnetic field looks precisely the same as predicted by Newtonian kinematics: relativistic effects do not have a place in this description.
+This means that, for instance, within the lab frame the motion of particles in a constant magnetic field looks precisely the same as predicted by Newtonian kinematics: relativistic effects do not have a place in this description.
 In conventional particle tracking a particle trajectory $\vec{x}(t)$ can be seen from the lab frame as the result of successive Galileo boosts that track the motion of the accelerated (in a constant magnetic field) particle.
-The usual Galileo rule for addition of velocities is used to determine the Galileo boosts tracking a particular particle, instant after instant, along its motion along the curved trajectory.
+The usual Galileo rule for the addition of velocities is used to determine the Galileo boosts tracking a particular particle, instant after instant, along its motion along the curved trajectory.
 The old kinematics is especially surprising, because we are based on the use of the covariant approach.
-Where does it comes from?
-The previous commonly accepted derivation of the equations for the particle motion in the three dimensional space from the covariant equation Eq.(20) includes one delicate point.
+Where does it come from?
+The previous commonly accepted derivation of the equations for the particle motion in the three-dimensional space from the covariant equation Eq.(21) includes one delicate point.
 In Eq.(22) and Eq.(23) the restriction $\vec{p}=m\vec{v}/\sqrt{1-v^{2}/c^{2}}$ has already been imposed.
-One might well wonder why, because in the accepted covariant approach, the solution of the dynamics problem for the momentum in the lab frame makes no reference to the three-dimensional velocity.
-In fact, equation Eq.(20) tells us that the force is the rate of change of the momentum $\vec{p}$, but does not tell us how momentum varies with speed.
+One might well wonder why because in the accepted covariant approach, the solution of the dynamics problem for the momentum in the lab frame makes no reference to the three-dimensional velocity.
+Equation Eq.(21) tells us that the force is the rate of change of the momentum $\vec{p}$, but does not tell us how momentum varies with speed.
 The four-velocity cannot be decomposed into $u=(c\gamma,\vec{v}\gamma)$ when we deal with a particle accelerating along a curved trajectory in the Lorentz lab frame.
@@
 In other words, the presentation of the time component as the relation $d\tau=dt/\gamma$ between the $\tau$ and lab time $t$ is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
-It is important to stress in this point that the situation when only one clock in the comoving frame is involved in dynamics cannot be realized.
+It is important to stress at this point that the situation when only one clock in the comoving frame is involved in dynamics cannot be realized.
 The Newton law can be written down in the proper frame only when a space-time coordinate system $(x^{\prime},y^{\prime},z^{\prime},\tau)$ has been specified.
 The type of clock synchronization which provides the time coordinate in the Newton equation has never been discussed in textbooks.
-It is clear that without an answer to the question about the method of distant clock synchronization used, not only the concept of acceleration, but also the dynamics law has no physical meaning.
+It is clear that without an answer to the question about the method of distant clock synchronization used, not only the concept of acceleration but also the dynamics law has no physical meaning.
 A proper frame approach to relativistic particle dynamics is forcefully based on a definite (Einstein) coordinatization assumption.
 After this, the dynamics theory states that the equation of motion in the proper frame is $md^{2}\vec{x^{\prime}}/d\tau^{2}=\vec{f}^{\prime}$.
-It should be stressed that in the case of velocity increment $\Delta\vec{v^{\prime}}=\Delta\vec{x^{\prime}}/\Delta\tau$ we also deal with the distant events.
-It can be said with some abuse of language that 3-velocity vector is always a ”spatially extended” object.
+It should be stressed that in the case of velocity increment $\Delta\vec{v^{\prime}}=\Delta\vec{x^{\prime}}/\Delta\tau$, we also deal with the distant events.
+It can be said with some abuse of language that a 3-velocity vector is always a ”spatially extended” object.
 Let us return to our consideration of the relation $d\tau=dt/\gamma$ between the $\tau$ and lab time $t$.
-The calculation carried out in the case of an spatially extended object shows that the temporal coincidence of two distant events has absolute character: $\Delta\tau=0$ implies $\Delta t=0$.
-It should be note that usually the notation ”$\tau$” arose from the proper time.
+The calculation carried out in the case of a spatially extended object shows that the temporal coincidence of two distant events has absolute character: $\Delta\tau=0$ implies $\Delta t=0$.
+It should be noted that usually the notation ”$\tau$” arose from the proper time.
 Now let us recall the standard concept of the object’s proper time.
@@
 As we already know, the proper time on moving object flows slower than the time $t$.
-This phenomena was called time dilation.
+This phenomenon was called time dilation.
 The proper time can also be introduced for a point particle moving with acceleration.
-This standard interpretation of the proper time $\tau$ has evidently nothing to do with dynamics and one may wonder where this contradiction existing between the name and the content of time in dynamic law come from.
-In order to avoid being to abstract for to long we have given one example.
-Let us discuss the case in which the velocity increment is not in the direction of the uniform motion.
-For example, there may be a particle which is just moving ”upward” with velocity increment $\Delta y^{\prime}/\Delta\tau$ and the lab system is moving ”horizontally”.
-Now, according to textbooks, we simply using relation $d\tau=dt/\gamma$ with the standard result $\Delta v_{y}=\Delta v^{\prime}_{y}/\gamma$.
-However, paradoxical result are obtained by doing so.
-In particular, the trajectory of the particle does not include mixture of positions and time.
-One of the important conclusions of the discussion presented above in the Chapter 7 is that there is no absolute notion of simultaneity in the theory of special relativity.
+This standard interpretation of the proper time $\tau$ has nothing to do with dynamics and one may wonder where this contradiction existing between the name and the content of time in dynamic law comes from.
+To avoid being to abstract for too long we have given one example.
+Let us discuss the case where the velocity increment is not in the direction of the uniform motion.
+For example, there may be a particle which is moving ”upward” with velocity increment $\Delta y^{\prime}/\Delta\tau$ in the frame $S_{n}$ and the $S_{n}$ system is moving ”horizontally”.
+Now, according to textbooks, we simply use relation $d\tau=dt/\gamma$ with the standard result $\Delta v_{y}=\Delta v^{\prime}_{y}/\gamma$.
+However, paradoxical results are obtained by doing so.
+In particular, the trajectory of the particle does not include a mixture of positions and time.
+One of the important conclusions of the discussion presented above in Section 6 is that there is no absolute notion of simultaneity in the theory of special relativity.
 Therefore there is no notion of an instantaneous 3-space.
@@
 The standard presentation of the velocity increment transformation as the relation $\Delta v_{y}=\Delta v^{\prime}_{y}/\gamma$ is based on the hidden assumption that $y^{\prime}$-axis and $y$-axis are parallel.
-On the other hand Lorentz transformation makes axes ($y^{\prime}$ and $y$) oblique-angled due to Wigner rotation.
+On the other hand, Lorentz transformation makes axes ($y^{\prime}$ and $y$) oblique-angled due to Wigner rotation.
 A difference between non-covariant and covariant point particle trajectories follows from this immediately (We will discuss the transformation of velocities in greater detail in Section 16.5).
@@
 Surprises can surely be expected when we return from the four-vectors language to the three-dimensional velocity vector $\vec{v}$, which can be represented in terms of the components of four-vector as $\vec{v}=d\vec{x}/dx_{0}$.
-In contrast with the pseudo-Euclidean four-velocity space, the relativistic three-velocity space is a three-dimensional space with constant negative curvature, i.e.
-three-dimensional space with Lobachevsky geometry.
 It is well known that for rectilinear accelerated motion the usual momentum-velocity relation holds.
 In fact, for the rectilinear motion the combination of the usual momentum-velocity relation and the covariant three-velocity transformation (according to Einstein’s law of velocity addition) is consistent with the covariant three-momentum transformation and both (non-covariant and covariant) approaches produce the same trajectory.
-We can see why by examine the transformation of the three velocity in the theory of relativity.
+We can see why by examining the transformation of the 3-velocity in the theory of relativity.
 For a rectilinear motion, this transformation is performed as $v=(v^{\prime}+V)/(1+v^{\prime}V/c^{2})$.
@@
 But we want to express the new momentum in terms of the primed momentum and energy, and we note that $p=(p^{\prime}+\mathcal{E}^{\prime}V/c^{2})/\sqrt{1-V^{2}/c^{2}}$.
-Thus, for a rectilinear motion, the combination of Einstein addition law for parallel velocities and the usual momentum-velocity relation is consistent with the covariant momentum transformation.
-The dynamical evolution of the particle according to the non-covariant particle tracking may be considered as the result of successive infinitesimal Lorentz transformations.
-The corrected Newton law is valid for each step.
-In other words, the two observers with rectilinear trajectories have common 3-space.
-This result was incorrectly extended to an arbitrary trajectory.
+Thus, for a rectilinear motion, the combination of Einstein’s addition law for parallel velocities and the usual momentum-velocity relation is consistent with the covariant momentum transformation.
 We already know that the collinear Lorentz boosts (similar to the collinear Galilean boosts) form a commutative group.
@@
 As well-known, the composition of non-collinear Lorentz boosts is equivalent to a Lorentz boost followed by a spatial rotation.
-This Wigner rotation is relativistic effect that does not have a non-covariant analogue.
-One of the consequences of non-commutativity of non-collinear Lorentz boosts is the unusual momentum-velocity relation $\vec{p}_{cov}\neq m\vec{v}_{cov}/\sqrt{1-v^{2}_{cov}/c^{2}}$, which also does not have any non-covariant analogue.
-The theory of relativity shows that the unusual momentum-velocity relation discussed above is related with the acceleration along curved trajectories.
-In this case there is no notion of a common ordinary space and there is a difference between covariant and non-covariant particle trajectories.
-Only the solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle trajectories in the lab frame.
-A closer analysis of the concept of velocity, i.e.
-a discussion of the methods by which a time coordinate can actually be assigned in the lab frame, opens up the possibility of a description of such physical phenomena as radiation from a relativistic electron accelerating along a curved trajectory in accordance with the theory of relativity.
+This Wigner rotation is a relativistic effect that does not have a non-covariant analogue.
+One of the consequences of non-commutativity of non-collinear Lorentz boosts is the unusual momentum-velocity relation $\vec{p}_{cov}\neq m\vec{v}_{cov}/\sqrt{1-v^{2}_{cov}/c^{2}}$, which also does not have any non-covariant analog.
+The theory of relativity shows that the unusual momentum-velocity relation discussed above is related to the acceleration along curved trajectories.
+In this case, there is no notion of a common ordinary space and there is a difference between covariant and non-covariant particle trajectories.
 16.4 The Usual Integration of the 4D Covariant Equation of Motion
 Attempts to solve the dynamics equation Eq.(21) in manifestly covariant form can be found in the literature.
-The trajectory which is found does not include relativistic kinematics effects.
+The trajectory that is found does not include relativistic kinematics effects.
 Therefore, it cannot be identified with $\vec{x}_{cov}(t)$ even if, at first glance, it appears to be derived following covariant prescription.
-First, we analyze textbooks presentation of the integration of the covariant equation of motion.
+First, we analyze the textbook’s presentation of the integration of the covariant equation of motion.
 Consider, for example, the motion of a particle in a given electromagnetic field.
-The simplest case, of great practical importance, is that of an uniform electromagnetic field meaning that $F^{\mu\nu}$ is constant on the whole space-time region of interest.
-In particular we consider the motion of a particle in a constant homogeneous magnetic field, specified by tensor components $F^{\mu\nu}=B(e^{\mu}_{2}e^{\nu}_{3}-e^{\nu}_{2}e^{\mu}_{3})$ where $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are orthonormal space like basis vectors $e^{2}_{2}=e^{2}_{3}=-1$, $e_{2}\cdot e_{3}=0$.
+The simplest case, of great practical importance, is that of a uniform electromagnetic field meaning that $F^{\mu\nu}$ is constant on the whole space-time region of interest.
+In particular, we consider the motion of a particle in a constant homogeneous magnetic field, specified by tensor components $F^{\mu\nu}=B(e^{\mu}_{2}e^{\nu}_{3}-e^{\nu}_{2}e^{\mu}_{3})$ where $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are orthonormal space like basis vectors $e^{2}_{2}=e^{2}_{3}=-1$, $e_{2}\cdot e_{3}=0$.
 In the lab frame of reference where $e^{\mu}_{0}$ is taken as the time axis, and $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are space vectors the field is indeed purely magnetic, of magnitude $B$ and parallel to the $e_{1}$ axis.
@@
 The components of the equation of motion are then $du^{(0)}/d\tau=0$, $du^{(1)}/d\tau=0$, $du^{(2)}/d\tau=-eBu^{(3)}/(mc)$, $du^{(3)}/d\tau=eBu^{(2)}/(mc)$.
-We seek for the initial value solution to these equations as done in the existing literature (2).
-A distinctive feature of the initial value problem in relativistic mechanics, is that the dynamics is always constrained.
-In fact, the evolution of the particle is subject to $mdu^{\mu}/d\tau=eF^{\mu\nu}u_{\nu}$, but also to the constraint $u^{2}=c^{2}$.
+We seek the initial value solution to these equations as done in the existing literature 606060 From the solution of the covariant equation of motion authors of textbooks [5, 63, 64] conclude that the covariant motion of a charged particle in a constant magnetic field is a uniform circular motion.
+The trajectory of the particle does not include relativistic kinematics effects and the Galilean vectorial law of addition of velocities is actually used.
+The old kinematics comes from the relation $d\tau=dt/\gamma$.
+It is only after the authors of textbooks have make those replacement for $d\tau$ that authors obtain the usual formula for non-covariant trajectory..
+A distinctive feature of the initial value problem in relativistic mechanics is that the dynamics is always constrained.
+The evolution of the particle is subject to $mdu^{\mu}/d\tau=eF^{\mu\nu}u_{\nu}$, but also to the constraint $u^{2}=c^{2}$.
 However, such a condition can be weakened requiring its validity at certain values of $\tau$ only, let us say initially, at $\tau=0$.
-Therefore, if $Y(\tau)$ vanishes initially, i.e.
+Therefore, if $Y(\tau)=u^{2}-c^{2}$ vanishes initially, i.e.
 $Y(0)=0$, then $Y(\tau)=0$ at any $\tau$.
@@
 One could expect that the particle’s trajectory in the lab frame, following from the previous reasoning $[0,X^{(2)}(t),X^{(3)}(t)]$, should be identified with $\vec{x}_{cov}(t)$.
-However, paradoxical result are obtained by doing so.
-In particular, the trajectory $[0,X^{(2)}(t),X^{(3)}(t)]$ does not include relativistic kinematics effects.
-We found that the usual integration of the four-dimensional covariant equation of motion Eq.(20) gives particle trajectory which looks precisely the same as in Newton dynamics and kinematics.
+However, the trajectory $[0,X^{(2)}(t),X^{(3)}(t)]$ does not include relativistic kinematics effects.
+We found that the usual integration of the four-dimensional covariant equation of motion Eq.(21) gives particle trajectory which looks precisely the same as in Newton’s kinematics and dynamics.
 The trajectory of the electron does not include relativistic effects and the Galilean vectorial law of addition of velocities is actually used.
 So we must have made a mistake.
-Notice that we are using textbooks to solve covariant equation of motion.
+Notice that we are using textbooks to solve covariant equations of motion.
 We did not make a computational mistake in our integrations, but rather a conceptual one.
@@
 What must be recognized is that the concept of velocity is only introduced in the second integration step.
-However, in accepted covariant approach, the solution of the dynamics problem for the momentum in the lab frame makes no reference to three-dimensional velocity.
-In fact, the initial condition which we used is $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$ and includes $\gamma c$ and $\gamma v$, which are actually notations for the time and space parts of the initial four-momentum.
-The three-dimensional trajectory and respectively velocity, which are convention-dependent, are only found after the second integration step.
-Then, where does the old kinematics comes from?
+However, in the accepted covariant approach, the solution of the dynamics problem for the momentum in the lab frame makes no reference to three-dimensional velocity.
+The initial condition that we used is $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$ and includes $\gamma c$ and $\gamma v$, which are actually notations for the time and space parts of the initial four-momentum.
+The three-dimensional trajectory and velocity, which are convention-dependent, are only found after the second integration step.
+Then, where does the old kinematics come from?
 The second integration was performed using the relation $d\tau=dt/\gamma$.
@@
 We should then expect to get results similar to those obtained in the case of the (3+1) non-covariant particle tracking.
-In fact, based on the structure of the four components of the equation of motion Eq.(20), we can arrive to another mathematically identical formulation of the dynamical problem.
+In fact, based on the structure of the four components of the equation of motion Eq.(21), we can arrive at another mathematically identical formulation of the dynamical problem.
 The fact that the evolution of the particle in the lab frame is subject to a constraint has already been mentioned.
@@
 In conventional particle tracking a particle trajectory $\vec{x}(t)$ can be seen from the lab frame as the result of successive Galileo boosts that track the motion of the accelerated (in a constant magnetic field) particle.
-The usual Galileo rule for addition of velocities is used to determine the Galileo boosts tracking a particular particle, instant after instant, along its motion along the curved trajectory.
+The usual Galileo rule for the addition of velocities is used to determine the Galileo boosts tracking a particular particle, instant after instant, along its motion along the curved trajectory.
 We cannot take old Newton’s kinematics for mechanics and Einstein’s kinematics for electrodynamics.
-Only solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle traectories in the lab frame.
+Only the solution of the dynamics equations in covariant form gives the correct coupling between the Maxwell’s equations and particle trajectories in the lab frame.
 The trajectory $\vec{x}_{cov}(t)$ is viewed from the lab frame as the result of successive Lorentz transformations.
-In this case relativistic kinematics effects arise.
-Let us try out our algorithm for reconstructing $\vec{x}_{cov}(t)$ on some example, to see how it works.
+In this case, relativistic kinematics effects arise.
+Let us try out our algorithm for reconstructing $\vec{x}_{cov}(t)$ on some examples, to see how it works.
 We will find formulas relating the velocity of a particle in one reference system to its velocity in a second reference system.
@@
 We begin with the simple case of relativistically combining perpendicular velocities.
-There is no length contraction for perpendicular distances, but time dilation still occur.
+There is no length contraction for perpendicular distances, but time dilation still occurs.
 Therefore the particle velocity in the frame $K$ is just $v_{x}=V$, $v_{y}=v_{y}^{\prime}/\gamma$.
@@
 It only makes sense to say two velocities are perpendicular if they are measured in the same reference frame.
-Hence when authors of textbooks say that $\vec{e}_{x}^{\prime}v_{x}^{\prime}$ and $\vec{e}_{x}V$ are perpendicular they actually mean that $(x^{\prime},y^{\prime},z^{\prime})$ axes and $(x,y,z)$ axes are parallel.
-In other words, it is (hidden) assumed that $K$ and $K^{\prime}$ observers have common 3-space.
-This is misconception.
+Hence when authors of textbooks say that $\vec{e}_{y^{\prime}}v_{y}^{\prime}$ and $\vec{e}_{x}V$ are perpendicular they actually mean that $(x^{\prime},y^{\prime},z^{\prime})$ axes and $(x,y,z)$ axes are parallel.
+In other words, it is (hidden) assumed that $K$ and $K^{\prime}$ observers have common a 3-space.
+This is a misconception.
 One of the consequences of non-commutativity of non-collinear Lorentz boosts is not the existence of a common ordinary space.
 Suppose that in the Lorentz frame $K^{\prime}$ the traveler was observing particle motion along the $y^{\prime}$ axis.
-In other words measured velocity of the particle $v^{\prime}_{y}$ , and yet the frame is moving relative to the frame $K$ along $x$ axis.
+In other words, the measured velocity of the particle $v^{\prime}_{y}$ , and yet the frame is moving relative to the frame $K$ along $x$ axis.
 Then to an observer at rest in the Lorentz frame $K$, the frame $K^{\prime}$ with the moving particle seems to rotate about the frame $K$.
@@
 This means that we take limit $\gamma\gg 1\gg v^{\prime}_{y}/c$.
-We consider the small parameter $v^{\prime}_{y}/c$ and use the second order approximation.
+We consider the small parameter $v^{\prime}_{y}/c$ and use the second-order approximation.
 For ultrarelativistic limit $\gamma\gg 1$ and the axes of $K^{\prime}$ frame are then rotated with respect to the $K$ frame axes by angle $\theta_{w}=v^{\prime}_{y}/(\gamma c)$.
 We want to look now at the consequences of this rotation.
-First we notice that there is a projection of the velocity $v^{\prime}_{y}$ on the $x$-axis, so the velocity component along the $x$-axis in the frame $K$ is going to be smaller than $V$.
+First, we notice that there is a projection of the velocity $v^{\prime}_{y}$ on the $x$-axis, so the velocity component along the $x$-axis in the frame $K$ is going to be smaller than $V$.
 Our problem now is to work out the decrement of the horizontal velocity in the $K$ frame.
@@
 Integrating with respect to infinitesimal value $dv^{\prime}_{y}$, we find that the horizontal velocity is $v_{x}=V-(v^{\prime}_{y})^{2}/(2\gamma^{2}c)$.
-In the ultrarelativistic approximation we find the simple result $v_{y}=v^{\prime}_{y}/\gamma$, $v_{x}=V-v_{y}^{2}/(2c)$, so that Lorentz transformations simply leads to a rotation of particle velocity $V$ of the angle $v_{y}/c=v_{y}/V$.
-In contrast, according to textbooks, it follows that the total particle speed in the Lorentz $K$ frame increase from $V$ to $V+v_{y}^{2}/(2c)$.
+In the ultrarelativistic approximation we find the simple result $v_{y}=v^{\prime}_{y}/\gamma$, $v_{x}=V-v_{y}^{2}/(2c)$, so that Lorentz transformations simply lead to a rotation of particle velocity $V$ of the angle $v_{y}/c=v_{y}/V$.
+In contrast, according to textbooks, it follows that the total particle speed in the Lorentz $K$ frame increases from $V$ to $V+v_{y}^{2}/(2c)$.
 Our result is at odds with the prediction from textbooks.
-The commonly accepted derivation of the composition of velocities based on the use the relation $dt^{\prime}=dt/\gamma$ and does not account for the mixture of (transverse) positions and time.
-Now we go on to consider another situation, a very practical one.
-We want to study the case of relativistically combining perpendicular velocities where velocity $v_{y}^{\prime}$ approaches to that of light, $v_{y}^{\prime}\to c$, and the velocity $V$ is small, $V/c\ll 1$.
-According to textbooks, the velocity addition formula gives $v_{x}\to V$, $v_{y}=v_{y}^{\prime}/\gamma\to v_{y}^{\prime}-V^{2}/(2c)$.
-So that two Lorentz boosts simply leads to a rotation of particle velocity $v_{y}^{\prime}$ of the angle $V/c$.
-We can believed that it is right; it is example what called aberration of (ultra) relativistic particles and it is well known.
-How can that be?
-Perhaps it is not puzzling, through, when we remember that what we called inertial frame $K^{\prime}$ is really accelerated frame with respect to the inertial lab frame $K$.
-It is important to stress in this point that we consider the motion of a particle in the lab frame and we want to describe how to determine the covariant trajectory $\vec{x}(t)$.
-Above in the Chapter 7 we demonstrated that the aberration of particles in the accelerated frame $K^{\prime}$ can be described in the lab frame within standard special relativity taking advantage of the Wigner rotation theory.
-As viewed from the lab frame, the coordinate axes of the accelerated frame is rotated through the angle $V/c$ with respect to the coordinate axes of the lab frame.
-We can easily understand that the particle motion in the $K^{\prime}$ frame will be parallel to the $y$ coordinate axis of the lab frame.
-It is clear that we should then expect to get results similar to those obtained in the textbooks.
-When we studied asymptotic $V\to c$, $v_{y}^{\prime}/c\ll 1$, we saw that when particle is moving nonrelativistically in the frame $K^{\prime}$ (i.e.
-when the aberration effect is negligible), a particle motion will be parallel to the coordinate axis $y^{\prime}$ and we have no problem of demonstrating that there is the connection between the velocity transformation and the Wigner rotation.
-It is hoped that presented example will draw wide attention to the central role that the Wigner rotation plays in the transformation of non-collinear velocities.
-To emphasize the physical concepts involved, rather than mathematical formalism, we consider not the most general case of the (corrected) velocity addition theorem.
-However, the special case considered already involves all necessary ingredients.
+The commonly accepted derivation of the composition of velocities is based on the use of the relation $dt^{\prime}=dt/\gamma$ and does not account for the mixture of (transverse) positions and time.
 16.6 On the Relativistic Velocity Composition and the Wigner Rotation
 There are two difficulties in solving problems concerning covariant particle tracking.
-First of all, conventional particle tracking in the single inertial frame is actually based on classical Newton mechanics without reference to Lorentz transformations.
+First of all, conventional particle tracking in the single inertial frame is based on classical Newton mechanics without reference to Lorentz transformations.
 According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity.
-In other words, in this standard approach, introducing as only modification to the classical case the relativistic mass, time differ from space.
-It fact, we have no mixture positions and time.
+In other words, in this standard approach, introducing as an only modification to the classical case of the relativistic mass, the time differs from space.
+We have no mixture of positions and time.
 This essential point has never received attention in the physical community.
 Another difficulty is the confusion surrounding the combination of non-parallel velocities in special relativity.
-Let us to consider composition of two velocities when they are not lying on a straight line.
-In conventional particle tracking (similar to the Newton mechanics), one composes them geometrically by the method of parallelogram with velocities regarded as the usual vectors.
+Let us consider the composition of two velocities when they are not lying on a straight line.
+In conventional particle tracking (similar to the Newton’s mechanics), one composes them geometrically by the method of a parallelogram with velocities regarded as the usual vectors.
 In relativistic kinematics, the composition of the velocities is studied analytically from the composition of Lorentz boosts.
 At the same time it is common knowledge that Lorentz boosts alone do not make a group.
-We must conclude that when we are dealing with transformations and set of these transformations does not form a group we have no legal theory.
+We must conclude that when we are dealing with transformations and a set of these transformations does not form a group we have no legal theory.
 Authors of textbooks believe that the relativity principle requires that Einstein’s velocity addition law gives the composition law of a group.
@@
 Many experts who learned the theory of relativity using textbooks will find this statement disturbing.
-The popular belief remains that Einstein’s law of addition of velocities is associative, although it is an exercise in elementary algebra to prove non-associativity (3).
+The popular belief remains that Einstein’s law of addition of velocities is associative, although it is an exercise in elementary algebra to prove non-associativity  [^61] The non-commutativity and non-associativity of the relativistic composition of non-collinear velocities was first recognized by Mocanu [65] and latter resolved by Ungar [66], by showing role played by the Wigner rotation in the transformation..
 It is therefore impossible to get covariant trajectory, $\vec{x}_{cov}(t)$, if we have nothing but Lorentz boosts.
 Something else has to be added.
-Lorentz boosts do not form a group due to presence of the Wigner rotation.
-It is comes out quite natural that breakdown of commutativity and associativity in Einstein addition, turn out to be Wigner rotation.
+Lorentz boosts do not form a group due to the presence of the Wigner rotation.
+It comes out quite natural that the breakdown of commutativity and associativity in Einstein’s addition turns out to be Wigner rotation.
 16.7 Convention-Invariant Particle Tracking
@@
 The path is rather a purely geometrical notion.
-Complete paths or their parts may consist of, e.
-g., line segments, arcs, circles, helical curves.
-If we take the origin of the (Cartesian) coordinate system and we connect the point to the point laying on the path and describing the motion of the particle, then the creating vector will be a position vector $\vec{x}(l)$.
+Complete paths or their parts may consist of, e.g., line segments, arcs, circles, and helical curves.
+If we take the origin of the (Cartesian) coordinate system and connect the point to the point laying on the path, then the creating vector will be a position vector $\vec{x}(l)$.
 The derivative of a vector is the vector tangent to the curve described by the radius vector $\vec{x}(l)$.
 The sense of the $d\vec{x}(l)/dl$ is determined by the sense of the curve arc $l$.
-We already know from our discussion in Introduction that the path $\vec{x}(l)$ has exact objective meaning i.e.
+We already know from our discussion in Chapter 2 that the path $\vec{x}(l)$ has exact objective meaning i.e.
 it is convention-invariant.
-The components of the momentum four vector $mu=(\mathcal{E}/c,\vec{p})$ have also exact objective meaning.
-In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle in the lab frame is convention dependent and has no exact objective meaning.
+The components of the momentum four-vector $mu=(\mathcal{E}/c,\vec{p})$ have also exact objective meaning.
+In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle in the lab frame is convention-dependent and has no exact objective meaning.
 We want now to describe how to determine the position vector $\vec{x}(l)_{cov}$ in covariant particle tracking.
@@
 $\frac{d^{2}\vec{x}_{cov}}{dl^{2}}=\frac{d\vec{x}_{cov}}{dl}\times\left(\frac{e\vec{B}}{pc}\right)~{}.$ (26)
-These three equations corresponds exactly to the equations for the components of the position vector that can be found using the non-covariant particle tracking approach, and $\vec{x}(l)_{cov}$ is exactly equal to $\vec{x}(l)$ as it must be.
-The point is that both approaches describe correctly the same physical reality and since the curvature radius of the path in the magnetic field, and consequently the three-momentum, has obviously an objective meaning (i.e.
+These three equations correspond exactly to the equations for the components of the position vector that can be found using the non-covariant particle tracking approach, and $\vec{x}(l)_{cov}$ is exactly equal to $\vec{x}(l)$ as it must be.
+The point is that both approaches describe correctly the same physical reality and since the curvature radius of the path in the magnetic field, and consequently, the three-momentum, has an objective meaning (i.e.
 is convention-invariant), both approaches yield the same physical results.
-13.8 The Nature of Mass
-This section concentrates on the nature of mass.
-The concept of mass is one of the most fundamental notions in physics, comparable in importance only to concept of space and time.
-With advent of the special theory of relativity, physicists and philosophers focused their attention on the concept of relativistic mass.
-The remarkable progress in experimental and theoretical physics made during the past few decades has considerably deepened our knowledge concerning the nature of mass.
-We conclude this section with a brief discussion of the physical meaning of the Minkowski geometry.
-13.8.1 Phenomenology and Relativistic Extensions
-In order fully to understand the meaning of the embedding of the Newton’s dynamics law in the Minkowski space-time, one must keep in mind that, above, we characterized Newton’s equation in the Lorentz comoving frame as a phenomenological law.
-The microscopic interpretation of the inertial mass of a particle is not given.
-In other words, it is generally accepted that Newton’s second law is a phenomenological law and the rest mass is introduced in an ad hoc manner.
-The system of coordinates in which the equations of Newton’s mechanics are valid can be defined as Lorentz rest frame.
-The relativistic generalization of the Newton’s second law to any Lorentz frame permits us to make correct predictions.
-We are in the position to formulate the following general statement: any phenomenological law, which is valid in the Lorentz rest frame, can be embedded in the four dimensional space-time only by using Lorentz coordinatization (i.e.
-Einstein synchronization convention).
-Suppose we do not know why a muon disintegrates, but we know the law of decay in the Lorentz rest frame.
-This law would then be a phenomenological law.
-The relativistic generalization of this law to any Lorentz frame allows us to make a prediction on the average distance traveled by the muon.
-In contrast, in the non covariant (3+1) space and time approach there is no time dilation effect, since for Galilean transformations the time scales do not change.
-Therefore, in the (3+1) non covariant approach, there is no kinematics correction factor $\gamma$ to the travel distance of relativistically moving muons.
-The two approaches give, in fact, a different result for the travel-distance, which must be, however, convention-invariant.
-This glaring conflict between results of covariant and non covariant approaches can be explained as follows: it is a dynamical line of arguments that explains this paradoxical situation with the relativistic $\gamma$ factor.
-In fact, there is a machinery behind the muon disintegration.
-Its origin is explained in the framework of the Lorentz-covariant quantum field theory.
-In the microscopic approach to muon disintegration, Einstein and absolute time synchronization conventions give the same result for such convention-invariant observables like the average travel distance, and it does not matter which transformation (Galilean or Lorentz) is used.
-13.8.2 The Relativistic Mass
-In the non covariant (3+1) space and time approach, there is no time dilation nor length contraction, because for Galilean transformations time and spatial coordinates scales do not change.
-Moreover, it can easily be verified that Newton’s second law keeps its form under Galilean transformations.
-Therefore, in the (3+1) non covariant approach, there is no kinematics correction factor $\gamma$ to the mass in Newton’s second law.
-However, in contrast to kinematics effects like time dilation and length contraction, the correction factor $\gamma$ to the mass in the Newton’s second law has direct objective meaning.
-In fact, if we assign space-time coordinates to the lab frame using the absolute time convention, the equation of motion is still given by Newton’s second law corrected for the relativistic dependence of momentum on velocity even though, as just stated, it has no kinematical origin.
-Understanding this result of the theory of relativity is similar to understanding previously discussed results: at first we use Lorentz coordinates and later the (3+1) non covariant approach in terms of a microscopic interpretation that must be consistent with the principle of relativity.
-It is well-known from classical electrodynamics that the electromagnetic field of an electron carries a momentum proportional to its velocity for $v\ll c$, while for an arbitrary velocity $v$, the momentum is altered by the relativistic $\gamma$ factor in the case when the absolute time convention is used.
-Many attempts have been made to explain the electron mass as fully originating from electromagnetic fields.
-However, these attempts have failed.
-In fact, it is impossible to have a stationary non-neutral charge distribution held together by purely electromagnetic forces.
-In other words, mass and momentum of an electron cannot be completely electromagnetic in origin and in order to grant stability there is a necessity for compensating electromagnetic forces with non electromagnetic fields.
-From this viewpoint, Newton’s second law is an empirical phenomenological law where the relativistic correction factor $\gamma$ to the mass is introduced in an ad hoc manner.
-From a microscopic viewpoint, today accepted explanation of how structureless particles like leptons and quarks acquire mass is based on the coupling to the Higgs field, the Higgs boson having been recently experimentally observed at the LHC.
-This mechanism can be invoked to explain Newton’s second law from a microscopic viewpoint even for structureless particles like electrons.
-However, at larger scales, an interesting and intuitive concept of the origin of physical inertia is illustrated, without recurring to the Higgs field, by results of Quantum Chromodynamics (QCD) for protons and neutrons, which are not elementary and are composed of quarks and gluon fields.
-If an initial, unperturbed nuclear configuration is disturbed, the gluon field generates forces that tend to restore this unperturbed configuration.
-It is the distortion of the nuclear field that gives rise to the force in opposition to the one producing it, in analogy to the electromagnetic case.
-But in contrast to the electromagnetic model of an electron, the QCD model of a nucleon is stable, and other compensation fields are not needed.
-Now, the gluon field mass can be computed from the total energy (or momentum) stored in the field, and it turns out that the QCD version in which quark masses are taken as zero provides a remarkably good approximation to reality.
-Since this version of QCD is a theory whose basic building blocks have zero mass, the most of the mass of ordinary matter (more than 90 percent) arises from pure field energy.
-In other words, the mass of a nucleon can be explained almost entirely from a microscopic viewpoint, which automatically provides a microscopic explanation of Newton’s second law of motion.
-In order to predict, on dynamical grounds, the inertial mass of a relativistically moving nucleon one does not need to have access to the detailed dynamics of strong interactions.
-It is enough to assume Lorentz covariance (i.e.
-Lorentz form-invariance of field equations) of the complete QCD dynamics involved in nucleon mass calculations.
-The previous discussion, results in a most general statement: it is enough to assume Lorentz covariance of the quantum field theory involved in micro-particle (elementary or not elementary) mass calculations in order to obtain the same result for the relativistic mass correction from the two synchronization conventions discussed here, and it does not matter which transformation (Galilean or Lorentz) is used.
-13.8.3 What does Space-Time Geometry Explain?
-It is important to stress at this point that the dynamical line of argument discussed here explains what the Minkowski geometry physically means.
-The pseudo-Euclidean geometric structure of space-time is only an interpretation of the behavior of the dynamical matter fields in the view of different observers, which is an observable, empirical fact.
-It should be clear that the relativistic properties of the dynamical matter fields are fundamental, while the geometric structure is not.
-Dynamics, based on the field equations, is actually hidden in the language of kinematics.
-The Lorentz covariance of the equations that govern the fundamental interactions of nature is an empirical fact, while the postulation of the pseudo-Euclidean geometry of space-time is a mathematical interpretation of it that yields the laws of relativistic kinematics: at a fundamental level this postulate is, however, based on the way fields behave dynamically.
-13.9 Bibliography and Notes
-1.
-Let us try to get a better understanding of the geometric restatement of Newton’s second law.
-To derive the covariant form of relativistic dynamics, we should embed the three-dimensional vector relation $md\vec{v}/dt=\vec{f}$ into the four-dimensional geometry of Minkowski space [52].
-The idea of embedding is based on the principle of relativity i.e.
-on the fact that the usual Newton’s second law can always be used in any Lorentz frame where the particle, whose motion we want to describe, is at rest.
-In other words, if an instantaneously comoving Lorentz frame is given at some instant, one can precisely predict the evolution of the particle in this frame during an infinitesimal time interval.
-In geometric language, the Newton law is strict on a hyperplane perpendicular to the world line.
-However, the hyperplane tilts together with its normal $u_{\mu}$ as one moves along the world line.
-For the embedding we need an operator $\hat{P}_{\perp}$ that continually projects vectors of Minkowski space on hyperplanes perpendicular to world line.
-The desired operator is $(\hat{P}_{\perp})_{\mu\nu}=\eta_{\mu\nu}-u_{\mu}u_{\nu}/u^{2}$ [52].
-In the instantaneously comoving frame one can unambiguously construct a four-force $K_{\mu}=[0,\vec{f}]$.
-Then, in an arbitrary Lorentz frame, the components $K_{\mu}$ can be found through the appropriate Lorentz transformation.
-In the rest frame obviously $u_{\mu}K^{\mu}=0$.
-It follows that, since $u_{\mu}K^{\mu}$ is an invariant, the four-force $K_{\mu}$ is perpendicular to the four-velocity $u_{\mu}$ in any Lorentz frame.
-The desired embedding of Newton’s second law in hyperplanes perpendicular to the world line is found by imposing $(\hat{P}_{\perp})_{\mu\nu}(mdu_{\nu}/d\tau-K_{\nu})=0$.
-This is a tensor equation in Minkowski space-time that relates geometric objects and does not need coordinates to be expressed.
-The evolution of a particle can be described in terms of world line $\sigma(\tau)$, and the 4-velocity by $u=d\sigma/d\tau$, having a meaning independently of any coordinate system.
-Similarly in geometric language, the electromagnetic field is described by the second-rank, antisymmetric tensor $F$, which also requires no coordinates for its definition.
-This tensor produces a 4-force on any charged particle given by $\hat{P}_{\perp}\cdot(mdu/d\tau-eF\cdot u)=0$ [52].
-This is the basic dynamics law for relativistic charged particles expressed in terms of geometric objects and automatically included the principle of relativity.
-The presence of the projector operator $\hat{P}_{\perp}$ suggests that we have only three independent equations.
-In the case of Maxwell’s equation we are able to rewrite the equations in the relativistic form without any change in the meaning at all, just with a change notations.
-It is important to noticed that the situation with dynamics equations is more complicated.
-In order fully to understand the meaning of the embedding of the dynamics law in the hyperplanes perpendicular to the world line, one must keep in mind that, above, we characterized the Newton’s equation in the Lorentz comoving frame as a phenomenological law.
-The system of coordinates in which the equations of Newton’s mechanics are valid can be defined as Lorentz rest frame.
-The relativistic generalization of the Newton’s second law to any Lorentz frame permits us to make correct predictions.
-The projector operator guarantees that this coordinate system restriction will be satisfied.
-2.
-In general, the covariant equation of motion can be solved only by numerical methods; however, it is always attractive to find instances where exact solutions can be obtained.
-The simplest case of great practical importance is that the motion of a particle in a constant homogeneous magnetic field.
-From the solution of covariant equation of motion authors of textbooks [5, 52, 53] conclude that the covariant motion of a charge particle in a constant magnetic field is a uniform circular motion.
-The trajectory of the particle does not include relativistic kinematics effects and the Galilean vectorial law of addition of velocities is actually used.
-Actually, the old kinematics comes from the relation $d\tau=dt/\gamma$.
-This relation between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate in the lab frame, is based on the use of the absolute time convention.
-It is only after the authors of textbooks have make those replacement for $d\tau$ that authors obtain the usual formula for non-covariant trajectory.
-3.
-The non-commutativity and non-associativity of the relativistic composition of non-collinear velocities was first recognized by Mocanu [54] and latter resolved by Ungar [55], by showing role played by the Wigner rotation in the transformation.
+17 Relativistic Spin Dynamics
+We would like to show that while learning relativistic particle dynamics we simultaneously learned about a large number of other subjects.
+Let us consider the example of the motions of spin in a given field.
+Spin dynamics equations can be expressed as tensor equations in Minkowski space-time.
+We shall limit ourselves to the case of a particle with a magnetic moment $\vec{\mu}$ in a microscopically homogeneous electromagnetic field.
+Evidently, the torque affects only the spin and the force affects only the momentum.
+It follows that the motion of the system as a whole in any frame is determined entirely by its charge, independent of magnetic dipole moment.
+This part of the motion has been treated in the previous chapter.
+We need to consider spin motion only now.
+17.1 Magnetic Dipole at Rest in an Electromagnetic Field
+Let us first consider the spin precession for a nonrelativistic charged particle.
+The proportionality of magnetic moment $\vec{\mu}$ and angular momentum $\vec{s}$ has been confirmed in many ”gyromagnetic” experiments on many different systems.
+The constant of proportionality is one of the parameters charactering a particular system.
+It is normally specified by giving the gyromagnetic ratio or $g$ factor, defined by $\vec{\mu}=ge\vec{s}/(2mc)$.
+This formula says that the magnetic moment is parallel to the angular momentum and can have any magnitude.
+For an electron $g$ is very nearly 2.
+Suppose that a particle is at rest in an external magnetic field $\vec{B}_{R}$.
+The equation of motion for the angular momentum in its rest frame is $d\vec{s}/d\tau=\vec{\mu}\times\vec{B}_{R}=eg\vec{s}\times\vec{B}_{R}/(2mc)=\vec{\omega}_{s}\times\vec{s}$.
+In other words, the spin precesses around the direction of the magnetic field with the frequency $\omega_{s}=-eg\vec{B}_{R}/(2mc)$.
+In the same nonrelativistic limit the velocity processes around the direction of $\vec{B}_{R}$ with the frequency $\omega_{p}=-(e/mc)\vec{B}_{R}$: $d\vec{v}/d\tau=(e/mc)\vec{v}\times\vec{B}_{R}$.
+Thus, for $g=2$ spin and velocity precess with the same frequency, so that the angle between them is conserved.
+17.2 Derivation of the Covariant (BMT) Equation of Motion of Spin
+In seeking the equation for the spin motion, we shall be guided by the known dynamics in the rest frame and the known relativistic transformation laws.
+We emphasize that spin is defined in a particular frame (the rest frame).
+Therefore, to form expressions with known transformation behavior, we need to introduce a four-quantity related to the spin.
+A convenient choice is a four- (pseudo)-vector $S$ defined by the requirement that in the rest frame, its space-like components are the spin components, while the time-like component is zero.
+We shall call $S$ four-spin; when normalized by dividing by its invariant length, it will be called polarization four-vector.
+It is space-like, and therefore in no frame does it space-like part vanish.
+Let the spin of the particle be represented in the rest frame by $\vec{s}$.
+The four-vector $S^{\alpha}$ is by definition required to be purely spatial at time $\tau$ in an instantaneous Lorentz rest frame $R(\tau)$ of the particle and to coincide at this time with the spin $\vec{s}(\tau)$ of the particle; that is $S_{R}^{\alpha}(\tau)=(0,\vec{S}_{R}(\tau))=(0,\vec{s}(\tau))$.
+At a later instant $\tau+\Delta\tau$ in an instantaneous inertial rest frame $R(\tau+\Delta\tau)$, we have similarly $S_{R}^{\alpha}(\tau+\Delta\tau)=(0,\vec{S}_{R}(\tau+\Delta\tau))=(0,\vec{s}(\tau+\Delta\tau))$.
+The BMT equation [46] is the manifestly covariant equation of motion for a four-vector spin $S^{\alpha}$ in an electromagnetic field $F^{\alpha\beta}$:
+$\frac{dS^{\alpha}}{d\tau}=\frac{ge}{2mc}\left[F^{\alpha\beta}S_{\beta}+\frac{1}{c^{2}}u^{\alpha}\left(S_{\lambda}F^{\lambda\mu}u_{\mu}\right)\right]-\frac{1}{c^{2}}u^{\alpha}\left(S_{\lambda}\frac{du^{\lambda}}{d\tau}\right)~{},$ (27)
+where $u_{\mu}=dx_{\mu}/d\tau$ is the four-dimensional particle velocity vector. With Eq.(21), one has
+$\frac{dS^{\alpha}}{d\tau}=\frac{e}{mc}\left[\frac{g}{2}F^{\alpha\beta}S_{\beta}+\frac{g-2}{2c^{2}}u^{\alpha}\left(S_{\lambda}F^{\lambda\mu}u_{\mu}\right)\right]~{},$ (28)
+The BMT equation is valid for any given inertial frame, and consistently describes, together with the covariant-force law, the motion of a charged particle with spin and magnetic moment.
+The covariant equation of spin motion for a relativistic particle under the action of the four-force $Q^{\mu}=eF^{\mu\nu}u_{\nu}$ in the Lorentz lab frame, Eq.(27), is a relativistic ”generalization” of the equation of motion for a particle angular momentum in its rest frame.
+Relativistic ”generalization” means that the three independent equations expressing the Larmor spin precession are embedded into the four-dimensional Minkowski space-time.
+In Lorentz coordinates there is a kinematics constraint $S^{\mu}u_{\mu}=0$, which is the orthogonality condition of four-spin and four-velocity.
+Because of this constraint, the four-dimensional dynamics law, Eq.(27), actually includes only three independent equations of motion.
+Using the explicit expression for Lorentz force we find that the four equations Eq.(27) automatically imply the constraint $S^{\mu}u_{\mu}=0$ as it must be.
+To prove this we may point out that one has in every Lorentz frame $S^{0}=\vec{S}\cdot\vec{v}$.
+While $S^{0}$ vanishes in the rest frame, $dS^{0}/d\tau$ need not.
+In fact $d(S^{\mu}u_{\mu})/d\tau=0$ implies $dS^{0}/d\tau=\vec{S}\cdot d\vec{v}/d\tau$.
+The immediate generalization of $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$ and $dS^{0}/d\tau=\vec{S}\cdot d\vec{v}/d\tau$ to arbitrary Lorentz frames is Eq.(27) as can be checked by reducing to the rest frame.
+A methodological analogy with the relativistic generalization of Newton’s second law emerges.
+To fully understand the meaning of embedding of the spin dynamics law in the Minkowski space-time, one must keep in mind that, above, we characterized the spin dynamics equation in the Lorentz comoving frame as a phenomenological law.
+The microscopic interpretation of the magnetic moment of a particle is not given.
+In other words, it is generally accepted that the spin dynamics law is a phenomenological law and the magnetic moment is introduced in an ad hoc manner.
+The system of coordinates in which the classical equations of motion for particle angular momentum are valid can be defined as Lorentz rest frame.
+The relativistic generalization of the three-dimensional equation $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$ to any Lorentz frame permits us to make correct predictions.
+17.3 Change Spin Variables
+When Bargman, Michel, and Telegdi first discovered the correct laws of spin dynamics, they wrote a manifestly covariant equation in Minkowski space-time, Eq.(28), which describes the motion of the four-spin $S^{\mu}$.
+The derivation of this equation was very similar to the four-tensor equations that were already known to relativistic particle dynamics.
+The equation Eq.(28) is more complex than one might think.
+It is composed of a set of coupled differential equations.
+To find a solution directly from the system seems quite difficult, even for a very symmetric, uniform magnetic field setup.
+How to solve this four-tensor equation is an interesting question.
+In relativistic spin dynamics, it is done in one particular way, which is very convenient.
+To apply Eq.(28) to specific problems it is convenient to introduce a three vector $\vec{s}~{}$ by the equation $\vec{s}=\vec{S}+S^{0}\vec{p}c/(\mathcal{E}+mc^{2})$.
+With the help of this relation, one can work out the equation of motion for $\vec{s}~{}$.
+Let us restrict our treatment of spinning particle dynamics to purely transverse magnetic fields.
+This means that the magnetic field vector $\vec{B}$ is oriented normally to the particle line motion.
+In this practically important case one has, after somewhat lengthy calculations [46]:
+$\frac{d\vec{s}}{d\tau}=\vec{\Omega}\times\vec{s}=-\frac{e}{2mc}\left[\left(g-2+\frac{2}{\gamma}\right)\gamma\vec{B}\right]\times\vec{s}~{},$ (29)
+What must be recognized is that in the accepted covariant approach (indeed, Eq.(28) is manifestly covariant), the solution of the dynamics problem for the spin in the lab frame makes no reference to the three-dimensional velocity.
+In fact, the Eq.(29) includes relativistic factor $\gamma$ and vector $\vec{v}/c$, which are actually notations: $\gamma=\mathcal{E}/(mc^{2})$, $\vec{v}/c=\vec{p}c/\mathcal{E}$.
+All quantities $\mathcal{E},\vec{p},\vec{B}$ are measured in the lab frame and have exact objective meaning i.e.
+they are convention-independent.
+The evolution parameter $\tau$ is also measured in the lab frame and has exact objective meaning.
+For instance, it is not hard to demonstrate that $d\tau=mdl/|\vec{p}|$, where $dl$ is the differential of the path length.
+Now we have an equation in the most convenient form to be solved.
+Suppose we let the charged spinning particle in the lab frame through a bending magnet with the length $dl$.
+We know that $d\theta=-eBdl/(|\vec{p}|c)$ is the orbital angle of the particle in the lab frame.
+Note that $d\tau=mdl/|\vec{p}|$.
+Then, Eq.(29) tells us that we may write the spin rotation angle with respect to the lab frame axes $\Omega d\tau$ as $\Omega d\tau=[(g/2-1)\gamma d\theta+d\theta]$.
+The spin vector $\vec{s}$ directly gives the spin as perceived in a comoving system.
+If we say that in the lab frame, the spin of a particle makes the angle $\phi$ with its velocity, we mean that in the particle’s rest frame, the spin makes this angle with the line motion of the lab frame.
+We must conclude that the conventional method used to explain the spin dynamics in the lab frame is very unusual.
+Indeed, the spin rotation measurement in the lab frame is interpreted from the viewpoint of the proper observer as viewing this of the lab observer.
+17.4 An Alternative Approach to the BMT Theory
+Above we described the BMT equation, Eq.(29), in the standard manner.
+It uses a spin quantity defined in the proper frame but observed with respect to the lab frame axes.
+That means that we know the orientation of the proper spin with respect to the lab coordinate system which is moving with velocity $-\vec{v}$ and acceleration $-\gamma d\vec{v}/d\tau$ in the proper frame.
+Let’s look at what the equation Eq.(29) says in a little more detail.
+It will be more convenient if we rewrite this equation as
+$d\vec{s}=\vec{\Omega}d\tau\times\vec{s}=-eg\gamma\vec{B}d\tau/(2mc)\times\vec{s}+e(\gamma-1)\vec{B}d\tau/(mc)\times\vec{s}~{}.$ (30)
+Now let’s see how we can write the right-hand side of Eq.(30).
+The first term is that we would expect the spin rotation due to a torque with respect to the proper frame axes $d\vec{\phi}_{L}=-eg\gamma\vec{B}d\tau/(2mc)=(g/2)\gamma d\vec{\theta}$.
+Here $d\vec{\theta}=-eBdl/(|\vec{p}|c)$ is the angle of the velocity rotation in the lab frame.
+It has also been made evident by our analysis in Section 8 that the angle of rotation $d\vec{\phi}_{W}=-e(\gamma-1)\vec{B}d\tau/(mc)=(\gamma-1)d\vec{\theta}$ corresponds to the Wigner rotation of the lab frame axes with respect to the proper frame axes.
+With these definitions, we have $d\vec{s}=\vec{\Omega}d\tau\times\vec{s}=d\vec{\phi}_{L}\times\vec{s}-d\vec{\phi}_{W}\times\vec{s}$.
+Now we introduce our new approach to the BMT theory, finding another way in which this complicated problem can be solved.
+We know that $d\vec{\phi}_{L}$ and $d\vec{\phi}_{W}$ are the rotations with respect to the proper frame axes.
+Actually, we only need to find the spin motion with respect to the lab frame axes.
+Now we must be careful about signs of rotations.
+There is a good mnemonic rule to learn the signs of different rotations.
+The rule says, first, that the direction of the velocity rotation in the proper frame is the same as the direction of the velocity rotation in the lab frame.
+Second, the direction of the lab frame rotation in the proper frame is the same as the direction of the velocity rotation in the proper frame.
+Third, the sign of the spin rotation due to a torque at $g>0$ (it is handy to remember that for an electron $g$ is positive and very nearly 2) means that the direction of the spin rotation in the proper frame is the same as the direction of the velocity rotation in the proper frame.
+We now ask about the proper spin rotation with respect to the lab frame axes.
+This is easy to find.
+The relative rotation angle is $d\vec{\phi}_{L}-d\vec{\phi}_{W}$.
+So we begin to understand the basic machinery behind spin dynamics.
+We see why the Wigner rotation of the lab frame axes in the proper frame must be taken into account if we need to know the proper spin dynamics with respect to the lab frame axes.
+Why the new derivation of the BMT equation is so simple?
+The reason is that the splitting of the particle spin motion with respect to the lab frame axes into the dynamic and kinematic parts can only be realized in the proper frame.
+In the proper frame, we do not need to know any more about a relativistic ”generalization” of the (phenomenological) classical equation of motion for the particle angular momentum.
+In this case, it is possible to separate the spin dynamics problem into the trivial (Larmor) dynamic problem and the kinematic problem of Wigner rotation of the lab frame axes in the proper frame.
+17.5 Spin Tracking
+Having written down the spin motion equation in a four-vector form, Eq.(28), and determined the components of the four-force, we satisfied the principle of relativity for one thing, and, for another, we obtained the four components of the equation for the spin motion.
+This is a covariant relativistic generalization of the usual three-dimensional equation of magnetic moment motion, which is based on the particle’s proper time as the evolution parameter.
+We next wish to describe the spin motion with respect to the Lorentz lab frame using the lab time $t$ as the evolution parameter.
+When going from the proper time $\tau$ to the lab time $t$, the frequency of spin precession with respect to the lab frame can be obtained using the well-known formula $d\tau=dt/\gamma$. We then find
+$\frac{d\vec{s}}{dt}=\vec{\varpi}\times\vec{s}=-\frac{e}{2mc}\left[\left(g-2+\frac{2}{\gamma}\right)\vec{B}\right]\times\vec{s}~{}.$ (31)
+The frequency of spin precession can be written in the form $\varpi=\omega_{0}[1+\gamma(g/2-1)]$, where $\omega_{0}$ is the particle revolution frequency.
+The old kinematics comes from the relation $d\tau=dt/\gamma$.
+The presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization that provides the time coordinate $t$ in the lab frame is based on the use of the absolute time convention.
+In the previous section, we saw that the particle path $\vec{x}(l)$ has an exact objective meaning i.e.
+it is convention-invariant.
+The spin orientation $\vec{s}$ at each point of the particle path $\vec{x}(l)$ has also exact objective meaning.
+In contrast to this, and consistently with the conventionality of the three-velocity, the function $\vec{s}(t)$ describing the spinning particle in the lab frame has no exact objective meaning.
+We now want to describe how to determine the spin orientation along the path $\vec{s}(l)$ in covariant spin tracking.
+Using the covariant equation Eq.(27) we obtain Eq.(29).
+If we use the relation $d\tau=mdl/|\vec{p}|$ our convention-invariant equation of spin motion reads
+$\frac{d\vec{s}}{dl}=-\frac{e\mathcal{E}}{m|\vec{p}|c^{3}}\left[\left(\frac{g}{2}-1+\frac{mc^{2}}{\mathcal{E}}\right)\vec{B}\right]\times\vec{s}=\left[\left(\frac{g}{2}-1\right)\frac{\mathcal{E}}{mc^{2}}+1\right]\frac{d\vec{\theta}}{dl}\times\vec{s}~{},$ (32)
+which is based on the path length $l$ as the evolution parameter.
+These three equations correspond exactly to the equations for components of the proper spin vector that can be found from the non-covariant spin tracking equation Eq.(31).
+We want to emphasize that there are two different (covariant and noncovariant) approaches that produce the same spin orientation $\vec{s}(l)$ along the path.
+The point is that both approaches describe correctly the same physical reality and the orientation of the proper spin $\vec{s}$ at any point of particle path in the magnetic field has an objective meaning, i.e.
+is convention-invariant.
 18 Relativity and Electrodynamics
@@
 Now let us use these equations to discuss the phenomena called radiation.
-To evaluate radiation fields arising from an external sources in Eq.
+To evaluate radiation fields arising from external sources in Eq.
 (8), we need to know the velocity $\vec{v}$ and the position $\vec{x}$ as a function of the lab frame time $t$.
-As discussed above, it is generally accepted that one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particles moving along non-covariant trajectory like $\vec{x}(t)$.
-The trajectory $\vec{x}(t)$, which follow from the solution of the corrected Newton’s second law Eq.
+As discussed above, it is generally accepted that one should solve the Maxwell’s equations in the lab frame with current and charge density created by particles moving along non-covariant trajectory like $\vec{x}(t)$.
+The trajectory $\vec{x}(t)$, which follows from the solution of the corrected Newton’s second law Eq.
 (4) under the absolute time convention, does not include, however, relativistic kinematics effects.
-We argue that this algorithm for solving usual Maxwell’s equations in the lab frame, which is considered in all standard treatments as relativistically correct, is at odds with the principle of relativity.
-However, the usual Maxwell’s equations in the lab frame, Eq.
-(8), are compatible only with covariant trajectories calculated by using Lorentz coordinates, therefore including such relativistic features as the relativity of simultaneity and the Wigner rotation.
-We now consider the particle motion in a giving magnetic field.
+We now consider the particle motion in a given magnetic field.
 In their usual form, Maxwell’s equations are valid only in Lorentz reference frames.
-According to correct coupling of fields and particles, from the Eq.
+According to the correct coupling of fields and particles, from the Eq.
 10 follows that
@@
 where $\vec{v}_{cov}=d\vec{x}_{cov}/dt$. The covariant trajectory of a particle is viewed from the lab frame as the result of successive Lorentz transformations.
-We find in this book that the Wigner rotation turns out to play an important role as the regulator of the velocity addition law.
-As we have already pointed, the four-velocity cannot be decomposed into $u=(c\gamma,\vec{v}\gamma)$ when we deal with a particle accelerating along a curved trajectory in the Lorentz lab frame (1).
-The presentation of the time component simply as the relation $d\tau=dt^{\prime}/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
-One of the consequences of non-commutativity of non-collinear Lorentz boosts is a difference between covariant and non covariant particle trajectories in a giving magnetic field.
-One can see that this essential point has never received attention by the physical community.
-As a result, a correction of the conventional radiation theory is required.
+[…]
 18.1 Why did the Error in Radiation Theory Remain So Long Undetected?
@@
 At this point, a reasonable question arises: why the error in radiation theory should have so long remained undetected?
-In order to answer this question, we shall discuss the subject more mathematically.
+To answer this question, we shall discuss the subject more mathematically.
 For an arbitrary parameter $v/c$ covariant calculations of the radiation process is very difficult.
 There are, however, circumstances in which calculations can be greatly simplified.
-As example of such circumstance is a non-relativistic radiation setup.
+An example of such a circumstance is a non-relativistic radiation setup.
 The non-relativistic asymptote provides the essential simplicity of the covariant calculation.
 The reason is that the non-relativistic assumption implies the dipole approximation which is of great practical significance.
-In accounting only for the dipole part of the radiation we neglect all information about the electron trajectory That means that the dipole radiation does not show any sensitivity to the difference between covariant and non-covariant particle trajectories.
-We want now to solve electrodynamics equations mathematically in a general way and consider the radiation associated with the succeeding terms in (multi-pole) expansion of the field in powers of the ratio $v/c$.
-Radiation theory is naturally developed in the space-frequency domain, as one is usually interested in radiation properties at a given position in space and at a certain frequency.
-In this book we define the relation between temporal and frequency domain via the following definition of Fourier transform pair:
+In accounting only for the dipole part of the radiation we neglect all information about the electron trajectory That means that the dipole radiation does not show any sensitivity to the difference between covariant and non-covariant particle trajectories  [^63] Similar to the non-relativistic asymptote, the ultrarelativistic asymptote also provides the essential simplicity of the covariant calculation.
+The reason is that the ultrarelativistic assumption implies the paraxial approximation.
+Note that the geometry of the electron motion in a bending magnet has a cylindrical symmetry.
+There are a number of remarkable effects which are a consequence of the cylindrical symmetry and the paraxial approximation.
+It can be demonstrated that there is no need to use covariant particle tracking for derivation of the bending magnet radiation.
+However, there is one situation where the conventional theory fails.
+The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion along the magnetic field (see the next chapter for more details)..
 […]
 Suppose we are interested in the radiation generated by an electron and observed far away from it.
-In this case it is possible to find a relatively simple expression for the electric field (see Appendix I).
+In this case, it is possible to find a relatively simple expression for the electric field.
 We indicate the electron velocity in units of $c$ with $\vec{\beta}$, the electron trajectory in three dimensions with $\vec{r}(t)$ and the observation position with $\vec{r}_{0}$.
@@
 In the far zone, by definition, the unit vector $\vec{n}$ is nearly constant in time.
-If the position of the observer is far away enough from the charge, one can make the expansion
-$\left|\vec{r}_{0}-\vec{r}(t)\right|=r_{0}-\vec{n}\cdot\vec{r}(t)~{}.$ (30)
+If the position of the observer is far away enough from the charge, one can make the expansion $\left|\vec{r}_{0}-\vec{r}(t)\right|=r_{0}-\vec{n}\cdot\vec{r}(t)$.
 We then obtain the following approximate expression for the the radiation field in the space-frequency domain:
@@
 where $\omega$ is the frequency, $(-e)$ is the negative electron charge and we make use of Gaussian units.
-A different constant of proportionality in Eq.(LABEL:revwied) and in the well known textbooks is to be ascribed to the use of different units and definition of the Fourier transform.
-First we will limit our consideration to the case of sources moving in a non-relativistic fashion.
-According to the principle of relativity, usual Maxwell’s equations can always be used in any Lorentz frame where sources are at rest.
-The same considerations apply where sources are moving in non-relativistic manner.
-In particular, when oscillating, charge particles emit radiation, and in the non-relativistic case, when the velocities of oscillating charges $v\ll c$, dipole radiation will be generated and described with the help of the Maxwell’s equations in their usual form, Eq.
-(8).
-Let’s examine in a more detail how the dipole radiation term comes about.
+A different constant of proportionality in Eq.(LABEL:revwied) and the well-known textbooks is to be ascribed to the use of different units and the definition of the Fourier transform.
+Let’s examine in more detail how the dipole radiation term comes about.
 The time $\vec{r}(t)\cdot(\vec{n}/c)$ in the integrands of the expression for the radiation field amplitude, Eq.
@@
 Then the time $\vec{r}(t)\cdot(\vec{n}/c)\sim a/c$.
-In order to ensure that the distribution of the charges in the system does not undergo a significant change during this time, it is necessary that $a\ll\lambda$.
+To ensure that the distribution of the charges in the system does not undergo a significant change during this time, it is necessary that $a\ll\lambda$.
 This condition can be written in still another form $v\ll c$, where $v$ is of the order of magnitude of the velocities of the charges.
-We consider the radiation associated with the first order term in the expansion of the Eq.
-(LABEL:revwied) in power of $\vec{r}(t)\cdot(\vec{n}/c)$.
+We consider the radiation associated with the zero order term in the expansion of the Eq.
+(LABEL:revwied) in the power of $\vec{r}(t)\cdot(\vec{n}/c)$.
 In doing so, we neglected all information about the electron trajectory $\vec{r}(t)$.
@@
 But that is only the first and most practically important term.
-The other terms tell us that there are higher order corrections to the dipole radiation approximation.
+The other terms tell us that there are higher-order corrections to the dipole radiation approximation.
 The calculation of this correction requires detailed information about the electron trajectory.
-Obviously, in order to calculate the correction to the dipole radiation, we will have to use the covariant trajectory and not be satisfied with the non-covariant approach.
+To calculate the correction to the dipole radiation, we will have to use the covariant trajectory and not be satisfied with the non-covariant approach.
 However, correction to the multipole radiation is expected to be small.
-For example, covariant (trajectory) correction for quadrupole radiation is treated as a correction of higher (than quadrupole) order.
+For example, covariant (trajectory) correction for quadrupole radiation is treated as a correction of a higher (than quadrupole) order.
 18.2 An Illustrative Example
-In the next chapter we present a critical reexamination of existing synchrotron radiation theory.
-But before the discussion of this topic it would be well to illustrate error in standard coupling fields and particles in accelerator and plasma physics by considering the relatively simple example, wherein the essential physical features are not obscured by unnecessary mathematical difficulties.
-This illustrative example is mainly addressed to readers with limiting knowledge of accelerator and synchrotron radiation physics.
+It would be well to illustrate errors in standard coupling fields and particles in accelerator and plasma physics by considering the relatively simple example, wherein the essential physical features are not obscured by unnecessary mathematical difficulties.
+This illustrative example is mainly addressed to readers with limited knowledge of accelerator and synchrotron radiation physics.
 Fortunately, the error in standard coupling fields and particles can be explained in a very simple way.
-14.2.1 The Velocity of an Electron Accelerated by the Kicker Field
 An electron kicker setup is a practical case of study for illustrating the difference between covariant and non-covariant trajectories.
-Let us consider the simple case when an ultrarelativistic electron moving with the velocity $v$ along $z$-axis in the lab frame is kicked by a weak dipole magnetic field directed along $x$-axis.
-We assume for simplicity that the kick angle is small compare with $1/\gamma$, where $\gamma=1/\sqrt{1-v^{2}/c^{2}}$ is the relativistic factor.
+Let us consider the simple case when an ultrarelativistic electron moving with the velocity $v$ along the $z$-axis in the inertial (lab) frame is kicked by a weak dipole magnetic field directed along the $x$-axis.
+We assume for simplicity that the kick angle is small compared with $1/\gamma$, where $\gamma=1/\sqrt{1-v^{2}/c^{2}}$ is the relativistic factor.
 This means that we take the limit $\gamma\gg 1\gg\gamma v_{y}/v$.
@@
 The trajectory of the electron, which follows from the solution of the corrected Newton’s second law under the absolute time convention, does not include relativistic effects.
-Therefore, as usual for Newtonian kinematics, Galilean vectorial law of addition of velocities is actually used.
+Therefore, as usual for Newtonian kinematics, the Galilean vectorial law of addition of velocities is used.
 Non-covariant particle dynamics shows that the electron direction changes after the kick, while the speed remains unvaried (Fig.
-60).
+49).
 According to non-covariant particle tracking, the magnetic field $B\vec{e}_{x}$ is only capable of altering the direction of motion, but not the speed of the electron.
@@
 Taking the ultrarelativistic limit $v\simeq c$ and using the second order approximation we get $v_{z}=v[1-v_{y}^{2}/(2v^{2})]=v[1-v_{y}^{2}/(2c^{2})]$.
-Figure 60: A setup for illustration the difference between covariant and non-covariant trajectories.
+Figure 49: A setup for illustrating the difference between covariant and non-covariant trajectories.
 The motion of a relativistic electron accelerated by a kicker field.
 It is assumed for simplicity that $\gamma\gg 1\gg\gamma v_{y}/v$.
-According to the non-covariant particle tracking, the magnetic field $B\vec{e}_{x}$ is only capable altering the direction of motion, but not the speed of the electron.
+According to the non-covariant particle tracking, the magnetic field $B\vec{e}_{x}$ is only capable of altering the direction of motion, but not the speed of the electron.
 In contrast, covariant particle tracking, which is based on the use of Lorentz coordinates, yields different results for the velocity of the electron.
@@
 Upstream of the kicker, the particle is at rest in the frame $S^{\prime}$.
-In order to have this, we impose that $S^{\prime}$ is connected to $S$ by the Lorentz boost $L(\vec{v})$, with $\vec{v}$ parallel to the $z$ axis, which transforms a given four vector event $X$ in a space-time into $X^{\prime}=L(\vec{v})X$.
-Let us analyze the particle evolution within $S^{\prime}$ frame.
+To have this, we impose that $S^{\prime}$ is connected to $S$ by the Lorentz boost $L(\vec{v})$, with $\vec{v}$ parallel to the $z$ axis, which transforms a given four vector event $X$ in space-time into $X^{\prime}=L(\vec{v})X$.
+Let us analyze the particle evolution within the $S^{\prime}$ frame.
 Our particle is at rest and the kicker is running towards it with velocity $-\vec{v}$.
@@
 If we neglect terms in $(\gamma v_{y}/c)^{3}$, the relativistic correction in the composition of velocities does not appear in this approximation.
-First, we analyze textbook presentation of a composition of motions.
-Let $S"$ be a frame fixed with respect to the particle downstream the kicker.
-As is known, the non collinear Lorentz boosts does not commute.
-In our second order approximation we can neglect the difference between the $\gamma v_{y}/c$ and $\gamma_{z}v_{y}/c$, where $\gamma_{z}=1/\sqrt{1-v_{z}^{2}/c^{2}}$.
+First, we analyze the textbook presentation of a composition of motions.
+Let $S"$ be a frame fixed with respect to the particle downstream of the kicker.
+As is known, the non-collinear Lorentz boosts do not commute.
+In our second-order approximation, we can neglect the difference between the $\gamma v_{y}/c$ and $\gamma_{z}v_{y}/c$, where $\gamma_{z}=1/\sqrt{1-v_{z}^{2}/c^{2}}$.
 Here $v_{z}=v(1-\theta_{k}^{2}/2)$ and $\theta_{k}=v_{y}/v=v_{y}/c$ in our (ultrarelativistic) case of interest.
@@
 For the simple case of parallel velocities, the addition law is $L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e}_{z}v)=L(\vec{e}_{z}v_{z})$.
-The resulting (passive) Lorentz boost composition can be represented as $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e}_{z}v_{z})X$ and textbooks state that the magnetic field, in the Lorentz coordinatization, is only capable altering the direction of motion, but not the speed of the electron.
+The resulting Lorentz boost composition can be represented as $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e}_{z}v_{z})X$ and textbooks state that the magnetic field, in the Lorentz coordinatization, is only capable altering the direction of motion, but not the speed of the electron.
 According to textbooks, in relativistic kinematics, the composition of the perpendicular velocities is studied analytically from the composition of Lorentz boosts.
@@
 In the ultrarelativistic approximation we find the simple result $v_{tot}=v_{z}=v[1-v_{y}^{2}/(2c^{2})]$, so that a Lorentz transformation leads to a rotation of the particle velocity $v_{z}$ of the angle $v_{y}/c$ (Fig.
-61).
+50).
 It becomes clear that the contradiction is related to the Wigner rotation.
-In fact, as we already discussed in the previous chapter, the relation $dt^{\prime}=dt/\gamma$ cannot be used when we deal with relativistic composition of perpendicular velocities.
-It should be note, however, that there is another satisfactory way of the covariant electron tracking.
+As we already discussed in the previous sections, the relation $dt^{\prime}=dt/\gamma$ cannot be used when we deal with the relativistic composition of perpendicular velocities in the Lorentz coordinatization.
+It should be noted, however, that there is another satisfactory way of covariant electron tracking.
 Let us now return to our consideration of the motion of a relativistic electron accelerated by the kicker field and let us discuss the observation of an inertial lab observer without passing from one reference frame to another.
@@
 In other words, we consider an active boost.
-In the case of an active boost we are dealing with motion of the same physical system, evolving in time and treated from the point of view of the same reference system.
-The simplest method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the electron was upstream the kicker.
-This synchronization convention preserves simultaneity and is actually based on the absolute time (or absolute simultaneity) convention.
+In the case of an active boost, we are dealing with the motion of the same physical system, evolving in time and treated from the point of view of the same reference system.
+The simplest method of synchronization consists of keeping, without changes, the same set of uniformly synchronized clocks used in the case when the electron was upstream the kicker.
+This synchronization convention preserves simultaneity and is based on the absolute time (or absolute simultaneity) convention.
 The usual Galileo rule of addition of velocities is used to determine the Galileo boosts tracking an electron, instant after instant, along its motion in a constant magnetic field.
-In the Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization.
-The overall combination of Galileo transformation and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the lab frame.
+In Section 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization.
+The overall combination of Galileo transformation and variable changes yields the Lorentz transformation in the case of absolute time coordinatization in the lab frame.
 Let us analyze the resynchronization process of the lab distant clocks during the acceleration of the electron.
@@
 Then, an instant after entering the magnetic field, the electron velocity changes of the infinitesimal value $d\vec{v}$ along the $y$-axis.
-At this first step, Eq.(21) allows us to express the differential $d\vec{v}$ through the differential $dt$ in the Lorentz coordinate system assigned upstream the kicker.
+At this first step, Eq.(22) allows us to express the differential $d\vec{v}$ through the differential $dt$ in the Lorentz coordinate system assigned upstream of the kicker.
 If clock synchronization is fixed, this is equivalent to the application of the absolute time convention.
-In order to keep Lorentz coordinates in the lab frame, as discussed before, we need to perform a clock resynchronization by introducing an infinitesimal time shift.
+To keep Lorentz coordinates in the lab frame, as discussed before, we need to perform a clock resynchronization by introducing an infinitesimal time shift.
 The simplest case is when the kick angle $\theta_{k}$ is very small, and we evaluate transformations, working only up to the order $(\theta_{k}\gamma)^{2}$.
@@
 Thus, Eq.(22) allows us to express the small velocity change $\Delta\vec{v}$ after the kick in the initial Lorentz coordinates system, and to perform clock resynchronization only downstream the kicker.
-Therefore, after the kick we can consider the composition of two Lorentz boosts along the perpendicular $x$ and $z$ directions.
+Therefore, after the kick we can consider the composition of two Lorentz boosts along the perpendicular $y$ and $z$ directions.
 The first boost imparts the velocity $v\theta_{k}\vec{e}_{y}$ to the electron along the $y$-axis and the second boost imparts the additional velocity $-(v\theta_{k}^{2}/2)\vec{e}_{z}$ along the $z$ axis, while the restriction to second order assures that the boosts commute.
 Figure 50: The motion of a relativistic electron accelerated by the kicker field.
-In order to describe the kicker influence on the electron trajectory as Lorentz transformation, one needs to perform a clock resynchronization by introducing a time shift and change the scale of time, that is the rhythm of all clocks, from $t$ to $\gamma_{y}t$, with $\gamma_{y}=1+\theta_{k}^{2}/2$.
-It follows that the total electron speed in the Lorentz lab frame downstream the kicker decreases from $v$ to $v(1-\theta^{2}_{k}/2)$.
-In order to keep a Lorentz coordinates system in the lab frame after the kick, that is equivalent to describe the kicker influence on the electron trajectory as Lorentz transformation, we need to perform a clock resynchronization by introducing a time shift and change the scale of time, that is the rhythm of all clocks, from $t$ to $\gamma_{y}t$, with $\gamma_{y}\simeq 1+\theta_{k}^{2}/2$.
-It is immediately understood that the speed of electron downstream the kicker is no longer independent of the electron motion in the magnetic field (Fig.
-61).
+To describe the kicker influence on the electron trajectory as Lorentz transformation, one needs to perform a clock resynchronization by introducing a time shift and change the scale of time.
+To keep a Lorentz coordinates system in the lab frame after the kick, which is equivalent to describing the kicker influence on the electron trajectory as Lorentz transformation, we need to perform a clock resynchronization by introducing a time shift and change the scale of time, that is the rhythm of all clocks, from $t$ to $\gamma_{y}t$, with $\gamma_{y}\simeq 1+\theta_{k}^{2}/2$.
+It is immediately understood that the speed of the electron downstream of the kicker is no longer independent of the electron motion in the magnetic field (Fig.
+50).
 No relativistic correction to the velocity component along the $y$-axis appears in the second order, but a correction of the longitudinal velocity component, changing $v_{z}$ to $v_{z}/\gamma_{y}$ with $v_{z}=v(1-\theta_{k}^{2}/2)$ and $v_{z}/\gamma_{y}=v(1-\theta_{k}^{2})$.
 It follows that the total electron speed in the lab frame, after clock resynchronization downstream the kicker, decreases from $v$ to $v(1-\theta_{k}^{2}/2)$.
-The time dilation does not come into the calculation of the velocity increment, but appears in the correction of the initial (relativistic) velocity $\vec{v}=v\vec{e}_{z}$.
-We have found that we get the same result whether we using passive Lorentz transformations or using active Lorentz transformations as it must be.
+The time dilation does not come into the calculation of the velocity increment but appears in the correction of the initial (relativistic) velocity $\vec{v}=v\vec{e}_{z}$.
 Note that we discuss particle tracking in the limit of a small kick angle $\gamma v_{y}/c\ll 1$.
 However, even in this simple case and for a single electron we are able to demonstrate the difference between non-covariant and covariant particle trajectories.
-The electron speed decreases from $v$ to $v(1-\theta_{k}^{2}/2)$.
-This result is at odds with the prediction from non-covariant particle tracking.
-14.2.2 Discussion
-We now want to point out that when we accelerate the electron in the lab frame by a kicker, the information about this acceleration is included into the covariant trajectory.
-The result of covariant approach clearly depends on the absolute value of the kick angle.
-The argument concerning the relativity motion in our case of interest cannot be applied.
-Not all is relative in special relativity, because any change of velocity has an absolute meaning.
-The ”absolute” acceleration means acceleration with respect to the fixed stars.
-There are several points to be made about the above results.
-One interesting question is, where physically is the information about the electron acceleration?
-We answer this question in great detail in the Section 15.6.
-It is a dynamical line of arguments that explains this paradoxical situation with the trajectory of the accelerated electron.
-Dynamics, based on the field equations, is actually hidden in the language of relativistic kinematics.
-Without proof, we may state the results.
-The distribution of the electromagnetic fields from a fast moving charge is described by Ginzburg-Frank formula (see Appendix IV).
-We remark first that for a rapidly moving electron we have nearly equal transverse and mutually perpendicular electric and magnetic fields.
-These are indistinguishable from the fields of a beam of radiation.
-From microscopic viewpoint, when the electron passes through a kicker, its fields are perturbed, and now include information about acceleration.
-Let us quickly look at some other cases of interest.
-Above we consider the simple case when kick angle is small.
-When the electron passes through the weak kicker there is no synchrotron radiation (to be more precise, in this case radiation is indistinguishable from the self-electromagnetic fields of the electron).
-We expect that an electron that passes through the weak kicker is still ”field-dressed”.
-Suppose the ultrarelativistic electron in the lab frame is kicked by dipole magnet field, but the kick angle is large compare with $1/\gamma$.
-In other words, we now consider an electron moving along a standard synchrotron radiation setup.
-At the exit of strong kicker we have a ”naked” (or ”field-free”) electron i.e.
-an electron that is not accompanied by virtual radiation fields.
-There is a process of formation of the ”field dressed” electron (i.e.
-formation of the fields from a fast moving charge) within the formation length from the very beginning of the straight section downstream of the strong kicker.
-The information about acceleration is included now into the synchrotron radiation.
-But the acceleration history (together with the self fields of the ultrarelativistic electron) is washed out during the kicking process, and now electron fields not include information about acceleration.
-That is to say, the information about acceleration is not included into the covariant trajectory downstream the strong kicker (we will discuss this subject further in the Section 14.5).
-14.2.3 The Motion of an Electron Accelerated by a Kicker in a Bending Magnet
+18.3 The Motion of an Electron Accelerated by a Kicker in a Bending Magnet
 In our relativistic but non-covariant study of electron motion in a given magnetic field, the electron has the same velocity and consequently the same relativistic factor $\gamma$ upstream and downstream of the kicker.
-Suppose we now put the electron through a bending magnet (i.e.
+Suppose we put the electron through a bending magnet (i.e.
 a uniform magnetic field directed along the $y$-axis ), Fig.
-62.
+51.
 The motion in the bending magnet we obtained is practically the same as in the case of non-relativistic dynamics, the only difference being the appearance of the relativistic factor $\gamma$ in the determination of cyclotron frequency $\omega_{c}=eB/(m\gamma)$.
@@
 Figure 51: Geometry for radiation production from a bending magnet.
-The motion of a relativistic electron accelerated by a kicker field.
 According to the non-covariant particle tracking, the magnetic field $B\vec{e}_{x}$ of the kicker is only capable altering the direction of motion, but not the speed of the electron.
-One could naively expect that according to covariant particle tracking, since the total speed of electron in the lab frame downstream of the kicker decreases from $v$ to $v(1-\theta_{k}^{2}/2)$, this would also lead to a consequent decrease of the three-momentum $|\vec{p}|$ from $m\gamma v$ to $m\gamma v(1-\gamma^{2}\theta_{k}^{2}/2)$ in our approximation.
-However, such a momentum change would mean a correction to the radius $R$ of order $\gamma^{2}\theta_{k}^{2}$ so that there is a glaring conflict with the calculation of the radius according to non covariant tracking.
-Since the curvature radius of the trajectory in the bending magnet has obviously an objective meaning, i.e.
+One could naively expect that according to covariant particle tracking since the total speed of an electron in the lab frame downstream of the kicker decreases from $v$ to $v(1-\theta_{k}^{2}/2)$, this would also lead to a consequent decrease of the three-momentum $|\vec{p}|$ from $m\gamma v$ to $m\gamma v(1-\gamma^{2}\theta_{k}^{2}/2)$ in our approximation.
+However, such a momentum change would mean a correction to the radius $R$ of order $\gamma^{2}\theta_{k}^{2}$ so that there is a glaring conflict with the radius calculation according to noncovariant tracking.
+Since the curvature radius of the trajectory in the bending magnet has an objective meaning, i.e.
 it is convention-invariant, this situation seems paradoxical.
-The paradox is solved taking into account the fact that in Lorentz coordinates the three-vector of momentum $\vec{p}$ is transformed, under Lorentz boosts, as the space part of the four vector $p_{\mu}$.
+The paradox is solved taking into account the fact that in Lorentz coordinates the three-vector of momentum $\vec{p}$ is transformed, under Lorentz boosts, as the space part of the four-vector $p_{\mu}$.
 Let us consider a composition of Lorentz boosts that track the motion of the relativistic electron accelerated by the kicker field.
@@
 We have $p_{\mu}=[\mathcal{E}/c,\vec{p}]$.
-We consider the Lorentz frame $S^{\prime}$ fixed with respect to the electron upstream the kicker, and in the special case when electron is at rest $p_{\mu}^{\prime}=[mc,\vec{0}]$.
+We consider the Lorentz frame $S^{\prime}$ fixed with respect to the electron upstream the kicker, and in the special case when an electron is at rest $p_{\mu}^{\prime}=[mc,\vec{0}]$.
 We turn focus on what happens in $S^{\prime}$.
@@
 to the energy of the electron.
-In fact, the energy increases from $mc^{2}$ to $mc^{2}+m(\gamma v_{y})^{2}/2$.
-We remind that $S^{\prime}$ is connected to the lab frame $S$ by a Lorentz boost.
+The energy increases from $mc^{2}$ to $mc^{2}+m(\gamma v_{y})^{2}/2$.
+We remember that $S^{\prime}$ is connected to the lab frame $S$ by a Lorentz boost.
 Now, with a boost to a frame moving at velocity $\vec{v}=-v\vec{e}_{z}$, the transformation of the longitudinal momentum component, normal to the magnetic field of the bend, is $p_{z}=\gamma(p_{z}^{\prime}+vp_{0}^{\prime}/c)=\gamma mv$.
 Therefore we can see that the momentum component along the $z$-axis remains unchanged in our approximation of the Lorentz transformation.
-We also have, from the transformation properties of four-vectors, that the time component $p_{0}=\gamma(p^{\prime}_{0}+vp^{\prime}_{z})=\gamma mc$
-Let us now return to our consideration on the covariant electron trajectory calculation in the Lorentz lab frame when a constant magnetic field is applied.
-We analyzed a very simple (but very practical) kicker setup and we noticed that, in fact, the three-momentum is not changed; so we have already verified that this transformation is the same as the non covariant transformation for the three-momentum, i.e.
-$\vec{p}_{cov}=\vec{p}$.
-We also found that there is a difference between covariant and non covariant output velocities, $v_{cov}<v$.
-In these transformations we therefore demonstrated that $\vec{p}_{cov}\neq m\vec{v}_{cov}/\sqrt{1-v^{2}_{cov}/c^{2}}$ for curved trajectory in ultrarelativistic asymptotic.
+We also have, from the transformation properties of four-vectors, that the time component $p_{0}=\gamma(p^{\prime}_{0}+vp^{\prime}_{z})=\gamma mc$.
 It is interesting to discuss what it means that there are two different (covariant and noncovariant) approaches that produce the same particle three-momentum.
-The point is that both approaches describe correctly the same physical reality and the curvature radius of the trajectory in the magnetic field (and consequently the three-momentum) has obviously an objective meaning, i.e.
+The point is that both approaches describe correctly the same physical reality and the curvature radius of the trajectory in the magnetic field (and consequently the three-momentum) has an objective meaning, i.e.
 is convention-invariant.
-In contrast to this, the velocity of the particle has objective meaning only up to a certain accuracy, because the finiteness of velocity of light takes place.
-14.2.4 Red Shift of the Critical Frequency of the Synchrotron Radiation
-Next we discuss the interesting problem of emission of synchrotron radiation in a bending magnet with and without kick.
+In contrast to this, the velocity of the particle has objective meaning only up to a certain accuracy, because the finiteness of the velocity of light takes place.
+18.4 Redshift of the Critical Frequency of the Synchrotron Radiation
+Next, we discuss the interesting problem of the emission of synchrotron radiation in a bending magnet with and without the kick (see the next chapter).
 Let us consider the setup pictured in Fig.
-62.
+51.
 Suppose that an ultrarelativistic electron moving along the $z$-axis in the lab frame is kicked by a weak dipole field directed along the $x$-axis before entering a uniform magnetic field directed along the $y$-axis, i.e.
@@
 Let us discuss the case when the source is heading towards the observer (Fig.
-63).
+52).
 An electromagnetic source propagates through the system as a function of time following a certain trajectory $\vec{x}(t^{\prime})$.
@@
 As a result, an observer sees the motion of the electromagnetic source as a function of $t$.
-The prime used here to indicate the retarded times should not be confused with the primes referring to a Lorentz transformed frame in the proceeding sections.
+The prime used here to indicate the retarded times should not be confused with the primes referring to a Lorentz-transformed frame in the proceeding sections.
 […]
 Figure 52: Geometry for synchrotron radiation from a bending magnet.
-Radiation from an electron passing through the setup is observed through a spectral filter by a fixed observer positioned on the tangent to the bend at point $P$, as shown in Fig.
-(a).
-Electromagnetic source propagates through the system, as a function of time, as shown in Fig.
-(b).
-However, electromagnetic signal emitted at time $t^{\prime}$ at a given position $x(t^{\prime})$ arrives at observer position at a different time $t$, due to finite speed of light.
-As a result, the observer in Fig.
-(a) sees the electromagnetic source motion as a function of $t$.
-The apparent motion is a hypocycloid, and not the real motion $x(t^{\prime})$.
-The observer sees a time-compressed motion of the sources, which go from point $A$ to point $B$ in an apparent time corresponding to an apparent distance $2R\theta dt/dt^{\prime}$.
-Figure 64: Formation length for bend.
-Formation length $L_{f}$ can also be considered as a longitudinal size of a single-electron source.
-It does not make sense at all to talk about the position where electromagnetic signals are emitted within $L_{f}$.
+Figure 53: Formation length for the bend.
 It is clear from the above that, according to conventional synchrotron radiation theory, if we consider radiation, the introduction of the kick only amounts to a rigid rotation of the angular distribution along the new direction of the electron motion.
@@
 We now see a second order correction $\theta_{k}^{2}$ that is, however, multiplied by a large factor $\gamma^{2}$.
-It should be note, however, that there is another satisfactory way of explaining the red shift.
+It should be noted, however, that there is another satisfactory way of explaining the redshift.
 We can reinterpret this result with the help of a non-covariant treatment, which deals with non-covariant particle trajectories, and with Galilean transformations of the electromagnetic field equations.
@@
 However, Maxwell’s equations do not remain invariant with respect to Galilean transformation, and the velocity of light has increased from $c$, without kick, to $c(1+\theta_{k}^{2}/2)$ with kick.
-The reason for the velocity of light being different from the electrodynamics constant $c$ is due to the fact that, according to the absolute time convention, the clocks after the kick are not resynchronized.
+The reason for the velocity of light being different from the electrodynamics constant $c$ is because, according to the absolute time convention, the clocks after the kick are not resynchronized.
 The ratio of the electron velocity to that of light is convention independent i.e.
-it does not depend on the distant clock synchronization or on the rhythm of the clocks.
-Our calculations show that covariant and non-covariant treatments (at the correct coupling fields and particles) give the same result for the red shift prediction, which is obviously convention-invariant and depends only on the (dimensionless) parameter $v_{coord}/c_{coord}=v(1-\theta^{2}_{k}/2)/c=v/[c(1+\theta^{2}_{k}/2)]$, where $v_{coord}$ is the coordinate electron velocity, and $c_{coord}$ is the coordinate velocity of light.
-In order to confirm the predictions of our synchrotron radiation theory, we propose an experimental test at third generation synchrotron radiation sources.
+it does not depend on the distant clock synchronization or the rhythm of the clocks.
+Our calculations show that covariant and non-covariant treatments (at the correct coupling fields and particles) give the same result for the redshift prediction, which is convention-invariant and depends only on the (dimensionless) parameter $v_{coord}/c_{coord}=v(1-\theta^{2}_{k}/2)/c=v/[c(1+\theta^{2}_{k}/2)]$, where $v_{coord}$ is the coordinate electron velocity, and $c_{coord}$ is the coordinate velocity of light.
+To confirm the predictions of our synchrotron radiation theory, we propose an experimental test at third-generation synchrotron radiation sources.
 Synchrotron radiation from bending magnets is emitted within a wide range of frequencies.
-The possibility of using narrow bandwidth sources in an experimental study on the red shift in synchrotron radiation spectrum looks more attractive.
+The possibility of using narrow bandwidth sources in an experimental study on the redshift in the synchrotron radiation spectrum looks more attractive.
 This allows one to increase the sensitivity of the output intensity on the redshift, and to relax the requirement on beam kicker strength and photon beam line aperture.
@@
 They cause the electron beam to follow a periodic undulating trajectory with the consequence that interference effects occur.
-Undulators have typically many periods.
+Undulators typically have many periods.
 The interference of radiation produced in different periods results in a bandwidth that scales as the inverse number of periods.
-Therefore, the use of insertion devices installed at third generation synchrotron radiation facilities would allow us to realize a straightforward increase in the sensitivity to the red shift at a relatively small kick angle, $\theta_{k}<1/\gamma$.
-In the next chapter we discuss this experimental test in more detail.
-14.3 Bibliography and Notes
+Therefore, the use of insertion devices installed at third-generation synchrotron radiation facilities would allow us to realize a straightforward increase in the sensitivity to the redshift at a relatively small kick angle, $\theta_{k}<1/\gamma$.
+In other words, we are stating that, despite the many measurements done over decades, synchrotron radiation theory is not an experimentally well-confirmed theory.
+18.5 Discussion
 […]
+Let us quickly look at some other cases of interest.
+Above we consider the simple case when the kick angle is small.
+When the electron passes through the weak kicker there is no synchrotron radiation (to be more precise, in this case, radiation is indistinguishable from the self-electromagnetic fields of the electron).
+We expect that an electron that passes through the weak kicker is still ”field-dressed”.
+Suppose the ultrarelativistic electron in the lab frame is kicked by a dipole magnet field, but the kick angle is large compared with $1/\gamma$.
+In other words, we now consider an electron moving along a standard synchrotron radiation setup.
+At the exit of the strong kicker, we have a ”naked” (or ”field-free”) electron i.e.
+an electron that is not accompanied by virtual radiation fields.
+There is a process of formation of the ”field dressed” electron (i.e.
+formation of the fields from a fast-moving charge) within the formation length from the very beginning of the straight section downstream of the strong kicker.
+The information about acceleration is included now in the synchrotron radiation.
+But the acceleration history (together with the self fields of the ultrarelativistic electron) is washed out during the kicking process, and now electron fields not include information about acceleration.
+That is to say, the information about acceleration do not included in the covariant trajectory downstream of the strong kicker (we will discuss this subject in the next chapter).
 19 Synchrotron Radiation
 19.1 Introductory Remarks
-Accelerator physics was always thought in terms of the old (Newtonian) kinematics that is not compatible with Maxwell’s equations.
+Accelerator physics was always thought of in terms of the old (Newtonian) kinematics that is not compatible with Maxwell’s equations.
 We would like now to use our ideas about dynamics and electrodynamics to consider in some greater detail the question: ”Why did the error in the synchrotron radiation theory remain so long undetected?”
 The phenomena of electromagnetic radiation which we want to study are relatively complicated.
-For an arbitrary setup covariant calculations of the radiation process is very difficult.
+For an arbitrary setup, covariant calculations of the radiation process are very difficult.
 There are, however, circumstances in which calculations can be greatly simplified.
-As example of such circumstance is a synchrotron radiation setup.
+An example of such a circumstance is a synchrotron radiation setup.
 Similar to the non-relativistic asymptote, the ultrarelativistic asymptote also provides the essential simplicity of the covariant calculation.
@@
 We assume that the transverse velocity is small compared to the velocity of light.
-In other words, we use a second order relativistic approximation for the transverse motion.
-Instead of small (total) velocity parameter $(v/c)$ in the non-relativistic case, we use a small transverse velocity parameter $(v_{\perp}/c)$.
+In other words, we use a second-order relativistic approximation for the transverse motion.
+Instead of a small (total) velocity parameter $(v/c)$ in the non-relativistic case, we use a small transverse velocity parameter $(v_{\perp}/c)$.
 The next step is to analyze the longitudinal motion, following the same method.
-We should remark that the analysis of the longitudinal motion in a synchrotron radiation setup is very simple.
+We should note that the analysis of the longitudinal motion in a synchrotron radiation setup is very simple.
 If we evaluate the transformations up to second order $(v_{\perp}/c)^{2}$, the relativistic correction in the longitudinal motion does not appear in this approximation.
-According to covariant approach, the various relativistic kinematics effects concerning to the synchrotron radiation setup, turn up in successive orders of approximation.
+According to the covariant approach, the various relativistic kinematics effects concerning to the synchrotron radiation setup, turn up in successive orders of approximation.
 In the first order $(v_{\perp}/c)$.
@@
 - time dilation.
-Relativistic correction in law of composition of velocities, which already appears in the second order, and results directly from the time dilation.
-The first order kinematics term $(v_{\perp}/c)$ plays an essential role only in the description of the coherent radiation from a modulated electron beam.
-In a storage ring the distribution of the longitudinal position of the electrons in a bunch is essentially uncorrelated.
+Relativistic correction in the law of composition of velocities, which already appears in the second order, results directly from the time dilation.
+The first-order kinematics term $(v_{\perp}/c)$ plays an essential role only in the description of the coherent radiation from a modulated electron beam.
+In a storage ring, the distribution of the longitudinal position of the electrons in a bunch is essentially uncorrelated.
 In this case, the radiated fields due to different electrons are also uncorrelated and the average power radiated is a simple sum of the radiated power from individual electrons; that is we sum intensities, not fields.
@@
 19.2 Paraxial Approximation for the Radiation Field
-The general method to derive the frequency spectrum is to transform the electric field from the time domain to the frequency domain by the use of Fourier transform.
+The general method to derive the frequency spectrum is to transform the electric field from the time domain to the frequency domain by the use of the Fourier transform.
 First, let us rewrite Eq.
@@
 […]
-Eq. (LABEL:revwied) and Eq. (LABEL:rrevwied) are equivalent but include different integrands. This is no mistake, as different integrands can lead to the same integral (see Appendix I).
+Eq. (LABEL:revwied) and Eq. (LABEL:rrevwied) are equivalent but include different integrands. This is no mistake, as different integrands can lead to the same integral (see Appendix A1).
 We call $z_{0}$ the observation distance along the optical axis of the system, while $(x_{0},y_{0})$ fixes the transverse position of the observer.
 Using the complex notation, in this and in the following sections we assume, in agreement with Eq.
-(28), that the temporal dependence of fields with a certain frequency is of the form:
+(LABEL:ftdef2), that the temporal dependence of fields with a certain frequency is of the form:
 $\vec{E}\sim\vec{\bar{E}}(z_{0},x_{0},y_{0},\omega)\exp(-i\omega t)~{}.$ (40)
@@
 $\vec{E}=\vec{E}_{a}\exp\left(\frac{i\omega}{c}z_{0}-i\omega t\right)~{}.$ (41)
-In the following we will always assume that the ultra-relativistic approximation is satisfied, which is the case for SR setups.
+In the following, we will always assume that the ultra-relativistic approximation is satisfied, which is the case for SR setups.
 As a consequence, the paraxial approximation applies too.
@@
 We consider the motion in a static magnetic field.
-According to the conventional particle tracking the magnitude of the velocity is constant and is equal $v=ds/dt=\mathrm{const.}$, where $s(z)$ is the longitudinal coordinate along the path.
+According to conventional particle tracking the magnitude of the velocity is constant and is equal $v=ds/dt=\mathrm{const.}$, where $s(z)$ is the longitudinal coordinate along the path.
 The transverse components of the envelope of the field in Eq.
-(LABEL:rrevwied) in the far zone and in paraxial approximation finally becomes
+(LABEL:rrevwied) in the far zone and paraxial approximation finally becomes
 […]
@@
 Traditionally, all courses in synchrotron radiation theory have begun in the same way, retracing the path followed in the historical development of the subject.
-One first learns about synchrotron radiation from bending magnet.
-We will begin in this chapter by dealing with the ”advanced” subject of the synchrotron radiation theory.
-For undulator setup covariant calculations of the radiation process is quite simple.
+One first learns about synchrotron radiation from bending magnets.
+We will begin this chapter by dealing with the ”advanced” subject of the synchrotron radiation theory.
+For undulator setup covariant calculations of the radiation process are quite simple.
 To generate specific synchrotron radiation characteristics, radiation is often produced from special insertion devices called undulators.
-The resonance approximation, that can always be applied in the case of undulator radiation setups, yields simplifications of the theory.
+The resonance approximation, which can always be applied in the case of undulator radiation setups, yields simplifications of the theory.
 This approximation does not replace the paraxial one, but it is used together with it.
@@
 The frequency emitted by a particle going through an undulator can be obtained by considering the interference between the parts of the radiation created at successive periods (see Fig.
-65).
+54).
 The frequency of this field is Doppler shifted.
-The shortest wavelength is observed on undulator axis.
-In resonance approximation, all undulator radiation at shortest wavelength is emitted within an angle much smaller than $1/\gamma$ (see Fig.
-66).
+The shortest wavelength is observed on the undulator axis.
+In resonance approximation, all undulator radiation at the shortest wavelength is emitted within an angle much smaller than $1/\gamma$ (see Fig.
+55).
 This automatically selects observation angles of interest.
-In fact, if we consider observation angles outside the diffraction angle, we obtain zero intensity with accuracy $1/N_{w}$.
+If we consider observation angles outside the diffraction angle, we obtain zero intensity with accuracy $1/N_{w}$.
 We now have to ask: Why did the error in insertion device theory remain so long undetected?
-We answer this question in great detail bellow.
+We answer this question in great detail below.
 In discussing this we will deal only with the radiation into the central cone (see Fig.
-66).
+55).
 In doing so, we neglected all information about the electron trajectory.
-In this approximation the electron oscillation amplitude is always (independently of how large the undulator strength parameter) much smaller than the radiation diffraction size at the undulator exit, because $N_{w}\gg 1$.
+In this approximation, the electron oscillation amplitude is always (independently of how large the undulator strength parameter is) much smaller than the radiation diffraction size at the undulator exit because $N_{w}\gg 1$.
 Thus, the undulator radiation theory gives fields very much like the (dipole-like) instantaneous theory.
@@
 (43) can be used to characterize the far field from an electron moving on any path.
-In this section we present a simple derivation of the frequency representation of the radiated field produced by an electron in the planar undulator.
+In this section, we present a simple derivation of the frequency representation of the radiated field produced by an electron in the planar undulator.
 The magnetic field on the undulator axis has the form
@@
 $K=\frac{e\lambda_{w}B_{w}}{2\pi mc^{2}}~{},$ (48)
-$m$ being the electron mass at rest and $B_{w}$ being the maximal magnetic field of the undulator on axis.
-In this case the electron path is given by
+$m$ being the electron mass at rest and $B_{w}$ being the maximal magnetic field of the undulator on the axis.
+In this case, the electron path is given by
 $x(z)=r_{w}\cos(k_{w}z)~{},$ (49)
@@
 (51) alone, without extra-terms (both interfering and not) from the other parts of the electron path.
-This means that one should have complete information about the electron path and calculate extra-terms to be added to Eq.
-(46) in order to have the total field from a given setup.
+This means that one should have complete information about the electron path and calculate extra terms to be added to Eq.
+(51) to have the total field from a given setup.
 Yet, we can find particular situations for which the contribution from Eq.
@@
 (51) exhibits simplifications, independently of the frequency of interest due to the long integration range with respect to the scale of the undulator period.
-A well known expression for the angular distribution of the first harmonic field in the far zone (see Appendix II for a detailed derivation) can be obtained from Eq.
+A well-known expression for the angular distribution of the first harmonic field in the far zone (see Appendix A2 for a detailed derivation) can be obtained from Eq.
 (51).
@@
 Eq.
-(55) describes a field with spherical wavefront centered in the middle of the undulator.
+(60) describes a field with a spherical wavefront centered in the middle of the undulator.
 19.3.2 Why did the Error in Insertion Device Theory Remain so Long Undetected?
@@
 (141).
-Within the resonance approximation ($N_{w}\gg 1$) for the frequencies around the first harmonic it can be simplified to the well-known expression Eq.
+Within the resonance approximation ($N_{w}\gg 1$) for the frequencies around the first harmonic, it can be simplified to the well-known expression Eq.
 (60) where the field is horizontally polarized and azimuthal symmetric.
@@
 Taking the first zero of the $\sin(\cdot)/(\cdot)$ function at $C=0$ we will be able to determine the natural angular width of the radiation for the first harmonic $\theta_{c}$.
-The cone with aperture $\theta_{c}$ is usually called central cone.
+The cone with aperture $\theta_{c}$ is usually called the central cone.
 It can be found that $\theta_{c}^{2}=1/(2N_{w}\bar{\gamma}_{z}^{2})\ll 1/\bar{\gamma}_{z}^{2}$.
-Now we would like to understand what is the characteristic transverse size of the field distribution at the exist of the undulator.
-The radiation from magnetic poles always interferes coherently at zero angle with respect to undulator axis.
+Now we would like to understand what is the characteristic transverse size of the field distribution at the exit of the undulator.
+The radiation from magnetic poles always interferes coherently at zero angle with respect to the undulator axis.
 This interference is constructive within an angle of about $\sqrt{c/(\omega L_{w})}$.
@@
 It follows that $r_{w}^{2}/(cL_{w}/\omega)=K^{2}\omega/(L_{w}k_{w}^{2}\gamma^{2})=K^{2}/[\pi N_{w}(1+K^{2}/2)]\ll 1$, where we use the fact that $\gamma^{2}=(1+K^{2}/2)\bar{\gamma}_{z}^{2}$.
-This inequality holds independently of the value of $K$, because $N_{w}\gg 1$.
+This inequality holds independently of the value of $K$ because $N_{w}\gg 1$.
 Thus, the electron oscillating amplitude is always much smaller than the radiation diffraction size at the undulator exit.
 We consider the radiation associated with the first order term in the expansion of the Eq.
-(158) in power of small parameter $[K\theta_{x}\omega/(\gamma k_{w}c)]$ (see Eq.
+(147) in the power of small parameter $[K\theta_{x}\omega/(\gamma k_{w}c)]$ (see Eq.
 (150)).
@@
 (53) since the term $K\theta_{x}\omega\cos(k_{w}z^{\prime})/(\gamma k_{w}c)$ is neglected.
-In this approximation the electron orbit scale is always much smaller than the radiation diffraction size and Eq.
+In this approximation, the electron orbit scale is always much smaller than the radiation diffraction size, and Eq.
 (60) gives fields very much in agreement with the dipole radiation theory.
@@
 We consider the small expansion parameter $v_{\perp}/c$, neglecting terms of order $(v_{\perp}/c)^{3}$, but not of order $(v_{\perp}/c)^{2}$.
-In other words we use a second order relativistic approximation for transverse motion.
+In other words, we use a second-order relativistic approximation for transverse motion.
 We should remark that the analysis of the longitudinal motion in the ultrarelativistic approximation is much simpler than in the case of transverse motion.
-It is easy to see that the acceleration in the constant magnetic field yields an transverse electron velocity $v_{\perp}$ and $\Delta v_{z}=-v(v_{\perp}/c)^{2}/2$ parallel to the $z$-axis.
+It is easy to see that the acceleration in the constant magnetic field yields a transverse electron velocity $v_{\perp}$ and $\Delta v_{z}=-v(v_{\perp}/c)^{2}/2$ parallel to the $z$-axis.
 If we evaluate the transformations up to the second order $(v_{\perp}/c)^{2}$, the relativistic correction in the longitudinal motion does not appear.
@@
 Eq.
-(55) can be generalized to the case of a particle with a given offset $\vec{l}$ and deflection angle $\vec{\eta}$ with respect to the longitudinal axis, assuming that the magnetic field in the undulator is independent of the transverse coordinate of the particle (see Appendix III).
+(60) can be generalized to the case of a particle with a given offset $\vec{l}$ and deflection angle $\vec{\eta}$ with respect to the longitudinal axis, assuming that the magnetic field in the undulator is independent of the transverse coordinate of the particle (see Appendix A3).
 Although this can be done using Eq.
 (161) directly, it is sometimes possible to save time by getting the answer with some trick.
-For example, in the undulator case one takes advantage of the following geometrical considerations, which are in agreement with rigorous mathematical derivation.
+For example, in the undulator case, one takes advantage of the following geometrical considerations, which are in agreement with rigorous mathematical derivation.
 First, we consider the effect of an offset $\vec{l}$ on the transverse plane, with respect to the longitudinal axis $z$.
@@
 This induces a relative red shift in the resonant wavelength $\Delta\lambda/\lambda\sim\eta^{2}/2$.
-In practical cases of interest we may estimate $\eta\sim 1/\gamma$.
+In practical cases of interest, we may estimate $\eta\sim 1/\gamma$.
 Then, $\Delta\lambda/\lambda\sim 1/\gamma^{2}$ should be compared with the relative bandwidth of the resonance, that is $\Delta\lambda/\lambda\sim 1/N_{w}$, $N_{w}$ being the number of undulator periods.
@@
 $\widetilde{{E}}\left(z_{0},\vec{\eta},\vec{l},\vec{\theta}\right)=-\frac{K\omega eLA_{JJ}}{2c^{2}z_{0}\gamma}\exp\left[i\frac{\omega z_{0}}{2c}\left|\vec{\theta}-\frac{\vec{l}}{z_{0}}\right|^{2}\right]\mathrm{sinc}\left[\frac{\omega L\left|\vec{\theta}-\left({\vec{l}}/{z_{0}}\right)-\vec{\eta}\right|^{2}}{4c}\right]$ (62)
-Finally, in the far-zone case, we can always work in the limit for $l/z_{0}\ll 1$, that allows one to neglect the term ${\vec{l}}/{z_{0}}$ in the argument of the $\mathrm{sinc}(\cdot)$ function, as well as the quadratic term in $\omega l^{2}/(2cz_{0})$ in the phase.
+Finally, in the far-zone case, we can always work in the limit for $l/z_{0}\ll 1$, which allows one to neglect the term ${\vec{l}}/{z_{0}}$ in the argument of the $\mathrm{sinc}(\cdot)$ function, as well as the quadratic term in $\omega l^{2}/(2cz_{0})$ in the phase.
 Thus Eq.
@@
 19.3.4 Influence of the Kick According to Correct Coupling of Fields and Particles
-According to the correct coupling of fields and particles, there is a remarkable prediction of undulator radiation theory concerning to the undulator radiation from the single electron with and without kick.
+According to the correct coupling of fields and particles, there is a remarkable prediction of undulator radiation theory concerning the undulator radiation from the single electron with and without kick.
 Namely, when a kick is introduced, there is a red shift in the resonance wavelength of the undulator radiation in the velocity direction.
@@
 (161) can be expressed by the formula $\Delta\Phi_{T}=\omega\theta_{k}^{2}z^{\prime}/(2c)$, where we account for that $v\simeq c$ in ultrarelativistic approximation.
-Suppose that without kick the electron moves along the constrained trajectory parallel to the undulator axis.
-The field which produces this electron in the far zone is given by Eq.
+Suppose that without a kick the electron moves along the constrained trajectory parallel to the undulator axis.
+The field that produces this electron in the far zone is given by Eq.
 (60).
-Referring back to the Eq.
+Referring back to Eq.
 (64), we see that the conventional undulator radiation theory gives the following expression for radiation field after the kick
@@
 (66), the difference consisting in the term $\omega\theta_{k}^{2}/(2c)$ in the argument of $\mathrm{sinc}$ function.
-Attention must be called to the difference in resonance frequency between the undulator radiation setup with and without the kick.
+Attention must be paid to the difference in resonance frequency between the undulator radiation setup with and without the kick.
 Remembering the definition of the detuning parameter $C=k_{w}\Delta\omega/\omega_{r}$, we can write the redshift in resonance frequency as $\Delta\omega/\omega_{r}=-\omega_{r}\theta_{k}^{2}/(2k_{w}c)$.
-With this we also pointed out that the red shift can be written as $\Delta\omega/\omega_{r}=-\gamma^{2}\theta_{k}^{2}/(1+K^{2}/2)$.
+With this, we also pointed out that the red shift can be written as $\Delta\omega/\omega_{r}=-\gamma^{2}\theta_{k}^{2}/(1+K^{2}/2)$.
 We now see a second order correction $\theta_{k}^{2}$ that is, however, multiplied by the large factor $\bar{\gamma}_{z}^{2}$.
 We are now ready to investigate, more generally, what form the field expression takes under the introduction of a kick.
-Suppose that, without kick, the electron moves along the trajectory with angle $\vec{\eta}$ with respect to the undulator axis.
+Suppose that, without the kick, the electron moves along the trajectory with angle $\vec{\eta}$ with respect to the undulator axis.
 The field produced by this electron is given by Eq.
@@
 Now this all leads to an interesting situation.
-According to the conventional theory, the resonance wavelength depends only on the observation angle with respect to the electron velocity direction.
+According to conventional theory, the resonance wavelength depends only on the observation angle with respect to the electron velocity direction.
 Equation (70) says that for any kick angle $\vec{\theta}_{k}$ and for any angle $\vec{\eta}$ between the undulator axis and the initial electron velocity direction, the radiation along the velocity direction has no redshift.
@@
 The result of the covariant approach Eq.
-(67) clearly depends on the absolute value of the kick angle $\theta_{k}$ and the radiation along the velocity direction has the red shift only when the kick angle has nonzero value.
-We must conclude that when we accelerate the electron in the lab frame upstream the undulator, the information about this acceleration is included into the covariant trajectory.
+(72) depends on the absolute value of the kick angle $\theta_{k}$ and the radiation along the velocity direction has the redshift only when the kick angle has a nonzero value.
+We must conclude that when we accelerate the electron in the lab frame upstream of the undulator, the information about this acceleration is included in the covariant trajectory.
 19.3.5 Experimental Test of SR Theory in 3rd Generation Light Source
 One way to demonstrate incompatibility between the standard approach to relativistic electrodynamics, which deals with the usual Maxwell’s equations, and particle trajectories calculated by using non-covariant particle tracking, is to make a direct laboratory test of synchrotron radiation theory.
-In other words, we are stating here that, despite the many measurements done during decades, synchrotron radiation theory is not an experimentally well-confirmed theory.
+In other words, we are stating here that, despite the many measurements done over decades, synchrotron radiation theory is not an experimentally well-confirmed theory.
 Let us analyze the potential for exploiting synchrotron radiation sources in order to confirm the predictions of corrected synchrotron radiation theory.
-The emittance of the electron beam in new generation synchrotron radiation sources is small enough, so that one can neglect finite electron beam size and angular divergence in the soft X-ray wavelength range, and such synchrotron radiation source can be examined under the approximation of a filament electron beam.
+The emittance of the electron beam in new-generation synchrotron radiation sources is small enough, so that one can neglect finite electron beam size and angular divergence in the soft X-ray wavelength range, and such synchrotron radiation source can be examined under the approximation of a filament electron beam.
 This allows us to take advantage of analytical presentations for single electron synchrotron radiation fields.
@@
 The basic setup for a test experiment is sketched in Fig 56.
-The soft X-ray undulator beam line should be tuned to a minimum photon energy (typically this limit is related with the so called ”water window” wavelength range).
+The soft X-ray undulator beamline should be tuned to a minimum photon energy (typically this limit is related to the so-called ”water window” wavelength range).
 The radiation pulse goes through a monochromator filter $F$ and its energy is subsequently measured by the detector.
 No precise monochromatization of the undulator radiation is required in this case: a monochromator line width $\Delta\omega/\omega\simeq 10^{-3}$ is sufficient.
-In order for proposed test experiment to be curried out, it is necessary to control the beam kicking e.g.
+For the proposed test experiment to be curried out, it is necessary to control the beam kicking e.g.
 by corrector magnet.
-In the case of no kick the maximum pulse energy registered by the detector will coincide with the monochromator line tuned to resonance.
+In the case of no kick, the maximum pulse energy registered by the detector will coincide with the monochromator line tuned to resonance.
 When the kick is introduced the conventional synchrotron radiation theory still predicts a zero red shift in the resonance wavelength.
 In contrast to this, one of the immediate consequences of the corrected theory is the occurrence of a non-zero redshift.
-The proposed experimental procedure is relatively simple, because is based on relative measurements in the (electron beam) velocity direction with and without transverse kick.
+The proposed experimental procedure is relatively simple because is based on relative measurements in the (electron beam) velocity direction with and without transverse kick.
 Such a measurement is critical, in the sense that the prediction of conventional theory is the absence of redshift and (to our knowledge) has never been performed at synchrotron radiation facilities.
-However, XFEL based experiment confirm our correction for spontaneous undulator emission (1).
+However, XFEL based experiment confirms our correction for spontaneous undulator emission, see the next chapter for more details .
 19.4 Synchrotron Radiation from Bending Magnets
@@
 The observer in the standard treatment is assumed to be located in a vertical plane tangent to the circular trajectory at the origin, at an angle $\theta$ above the level of the orbit.
-In other words, in this geometry the $z$ axis is not fixed, but depends on the observer’s position.
+In other words, in this geometry, the $z$ axis is not fixed but depends on the observer’s position.
 Note that the geometry of the electron motion has a cylindrical symmetry, with the vertical axis going through the center of the circular orbit.
@@
 There are a number of remarkable effects which are a consequence of the cylindrical symmetry and the paraxial approximation.
-In discussing this we will demonstrate that there is no needs to use covariant particle tracking for derivation of the bending magnet radiation.
+In discussing this we will demonstrate that there is no need to use covariant particle tracking for derivation of the bending magnet radiation.
 However, there is one situation where the conventional theory fails.
-The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion along the magnetic field (i.e with respect to nominal orbit).
-It would be well to begin with bird’s view of the main results.
+The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion along the magnetic field (i.e.
+with respect to nominal orbit).
+It would be well to begin with a bird’s view of the main results.
 We want to solve the electrodynamics problem based on Maxwell’s equations in their usual form.
-In this case we should analyze the particle evolution within the framework of special relativity, where the problem of assigning Lorentz coordinates to the lab frame in the case of accelerating motion is complicated.
+In this case, we should analyze the particle evolution within the framework of special relativity, where the problem of assigning Lorentz coordinates to the lab frame in the case of accelerating motion is complicated.
 The only possibility to introduce Lorentz coordinates in this situation consists in introducing individual coordinate systems (i.e.
 individual rule-clock structure) for each point of the path.
-In order to keep Lorentz coordinates in the lab frame, as discussed before, we need only to perform a clock resynchronization by introducing a time shift and change the scale of time, that is the rhythm of clocks.
+In order to keep Lorentz coordinates in the lab frame, as discussed before, we need only to perform a clock resynchronization by introducing a time shift and changing the scale of time, that is the rhythm of clocks.
 Because of cylindrical symmetry, the observer is assumed to be located in the vertical plane tangent to the circular trajectory at the origin.
@@
 In the ultrarelativistic (paraxial) approximation, the observer sees a uniform acceleration of the electron $a=v^{2}/R\simeq c^{2}/R$ in the transverse direction.
-We can, then, write velocity and offset of the electron as follows $v_{\perp}=at$, $r_{\perp}=at^{2}/2$.
+We can, then, write the velocity and offset of the electron as follows $v_{\perp}=at$, $r_{\perp}=at^{2}/2$.
 Relativity of simultaneity gives time shift $\Delta t=v_{\perp}r_{\perp}/c^{2}=v_{\perp}^{2}t/(2c^{2})$.
@@
 This is a coincidence.
-It is because we have deal with uniform acceleration in the transverse direction using a second order (paraxial) approximation when an electron is moving along an arc of a circle.
+It is because we have to deal with uniform acceleration in the transverse direction using a second-order (paraxial) approximation when an electron is moving along an arc of a circle.
 19.4.1 Conventional Theory
 Consider a single relativistic electron moving on a circular orbit and an observer.
-It is worth to underline the difference between the more general geometry which we use and the geometry used in most synchrotron radiation textbooks for the treatment of bending magnet radiation.
+It is worth underlining the difference between the more general geometry that we use and the geometry used in most synchrotron radiation textbooks for the treatment of bending magnet radiation.
 In standard treatments of bending magnet radiation, the horizontal observation angle $\theta_{x}$ is always equal to zero.
@@
 $\Phi_{0}=-{\omega R\theta_{x}\over{2c}}\left({1\over{\gamma^{2}}}+{\theta_{x}^{2}\over{3}}+\theta_{y}^{2}\right)~{}.$ (85)
-In standard treatments of bending magnet radiation, the phase term $\exp(i\Phi_{0})$ is absent. In fact, the horizontal observation angle $\theta_{x}$ is always equal to zero.
+In standard treatments of bending magnet radiation, the phase term $\exp(i\Phi_{0})$ is absent. The horizontal observation angle $\theta_{x}$ is always equal to zero.
 19.4.2 Why did the Error in Synchrotron Radiation Remain so Long Undetected?
@@
 As already remarked, in conventional (non-covariant) particle tracking the description of the dynamical evolution in the lab frame is based on the use of the absolute time convention.
-In this case simultaneity is absolute, and we only need one set of synchronized clocks in the lab frame, to be used for the description of the accelerated motion.
+In this case, simultaneity is absolute, and we only need one set of synchronized clocks in the lab frame, to be used for the description of the accelerated motion.
 However, the use of the absolute time convention automatically implies the use of much more complicated field equations, and these equations are different for each value of the particle velocity i.e.
@@
 We start by considering an electron moving along a circular trajectory that lies in the $(x,z)$-plane and tangent to the $z$ axis.
-Because of cylindrical symmetry, in order to calculate spectral and angular photon distributions, it is not necessary to consider an observer at general location.
+Because of cylindrical symmetry, in order to calculate spectral and angular photon distributions, it is not necessary to consider an observer at the general location.
 The observer is assumed to be located in the vertical plane tangent to the circular trajectory at the origin.
@@
 (13): $x_{L}=\gamma_{x}x$, $t_{L}=(t/\gamma_{x}+\gamma_{x}xv_{x}/c^{2})$.
-We are dealing with a second order approximation and $\gamma_{x}=1+v_{x}^{2}/(2c^{2})$.
+We are dealing with a second-order approximation and $\gamma_{x}=1+v_{x}^{2}/(2c^{2})$.
 The overall combination of Galilean transformation and variable changes actually yields to the transverse Lorentz transformation.
-Since the Galilean transformation, completed by the introduction of the new variables, is mathematically equivalent to a Lorentz transformation, it obviously follows that transforming to new variables leads to the usual Maxwell’s equations.
-In order to keep Lorentz coordinates in the lab frame, as discussed before, we need only to perform a clock resynchronization by introducing the time shift $\Delta t=t_{L}-t=-[v_{x}^{2}/(2c^{2})]t+xv_{x}/c^{2}$.
+Since the Galilean transformation, completed by the introduction of the new variables, is mathematically equivalent to a Lorentz transformation, it follows that transforming to new variables leads to the Maxwell’s equations.
+To keep Lorentz coordinates in the lab frame, as discussed before, we need only to perform a clock resynchronization by introducing the time shift $\Delta t=t_{L}-t=-[v_{x}^{2}/(2c^{2})]t+xv_{x}/c^{2}$.
 The relativistic correction to the particle’s offset ”$x$” does not appear in this expansion order, but only in order of $v_{x}^{3}/c^{3}$ and $x_{L}=x$ in our case of interest.
-Although we have only shown that time shift in one rather special case, the result is right for any offset and (transverse) velocity direction: $\Delta t=t_{L}-t=-[|\vec{v}_{\perp}|^{2}/(2c^{2})]z^{\prime}/c+\vec{r}_{\perp}\cdot\vec{v}_{\perp}/c^{2}$.
+Although we have only shown time shift in one rather special case, the result is right for any transverse velocity direction: $\Delta t=t_{L}-t=-[|\vec{v}_{\perp}|^{2}/(2c^{2})]z^{\prime}/c+\vec{r}_{\perp}\cdot\vec{v}_{\perp}/c^{2}$.
 To finish our analysis we need only find a relativistic correction to the longitudinal motion.
 We remark again that if we evaluate the transformations up to the second order $(v_{\perp}/c)^{2}$, the relativistic correction in the longitudinal motion does not appear in this approximation.
-We have demonstrated the covariant method that can be used for any trajectory - a general way of funding what happens directly in space-frequency domain and in paraxial approximation.
+We have demonstrated the covariant method that can be used for any trajectory - a general way of funding what happens directly in the space-frequency domain and in paraxial approximation.
 Let us now see how to apply this covariant method to a special situation.
@@
 In the ultrarelativistic approximation, we have a uniform acceleration of the electron $a=v^{2}/R\simeq c^{2}/R$ in the transverse direction.
-We can, then, write velocity and offset of the electron as follows $v_{x}=at=az^{\prime}/v=az^{\prime}/c$, $x=at^{2}/2=az^{\prime 2}/(2c^{2})$.
-We have now all quantities we wanted.
-Let us put them all together in relativistic time shift: $\Delta t=t_{L}-t=-a^{2}z^{\prime 3}/(2c^{5})+a^{2}z^{\prime 3}/(2c^{5})=0$.
+We can, then, write the velocity and offset of the electron as follows $v_{x}=at=az^{\prime}/v=az^{\prime}/c$, $x=at^{2}/2=az^{\prime 2}/(2c^{2})$.
+We have now all the quantities we wanted.
+Let us put them all together in a relativistic time shift: $\Delta t=t_{L}-t=-a^{2}z^{\prime 3}/(2c^{5})+a^{2}z^{\prime 3}/(2c^{5})=0$.
 There is no difference!
-We do not need to use covariant particle tracking for derivation of the bending magnet radiation.
+We do not need to use covariant particle tracking to derive the bending magnet radiation.
 Usually, such a beautiful cancellation is found to stem from a deep underlying principle.
-Nevertheless, in this case there does not appear to be any such profound implication.
+Nevertheless, in this case, there does not appear to be any such profound implication.
 This is a coincidence.
-This cancellation is not surprising, if one analyzes the general expression for the radiation field from bending magnet in the far zone Eq.(78).
+This cancellation is not surprising, if one analyzes the general expression for the radiation field from the bending magnet in the far zone Eq.(83).
 In our previous discussion of undulator radiation, we learned that the relativistic correction appears only when the transverse electron trajectory is included in the total phase $\Phi_{T}$ under the integral Eq.(43).
 Referring back to Eq.(45) for the phase factor $\Phi_{T}$, we see that the term which depends on the transverse position of the electron can be written as $\exp i(\omega/c)[\theta_{x}x(z^{\prime})+\theta_{y}y(z^{\prime})]$.
-We conclude that the observation angle in the total phase factor under the integral must be related with the contribution of the transverse electron trajectory.
+We conclude that the observation angle in the total phase factor under the integral must be related to the contribution of the transverse electron trajectory.
 Now look at Eq.(83).
@@
 This means that the transverse constraint motion of the electron in the bending magnet does not affect synchrotron radiation.
-So we are justified using a non-covariant approach for considering the constrained electron motion along the nominal orbit in $(x,z)$-plane.
-We point out that the cancellation in relativistic time shift and the independence of the Fraunhofer propagator (to be more precise, in space-frequency domain we are dealing with a paraxial approximation of Green’s function of nonhomogeneous Helmholtz equation) on the observation angle $\theta_{x}$ in the far zone can be regarded as the two sides of the same coin: they are manifestation of the cylindrical symmetry when an electron is moving along an arc of a circle.
+We are justified in using a non-covariant approach to considering the constrained electron motion along the nominal orbit in the $(x,z)$-plane.
+We point out that the cancellation in relativistic time shift and the independence of the Fraunhofer propagator (to be more precise, in the space-frequency domain we are dealing with a paraxial approximation of Green’s function of nonhomogeneous Helmholtz equation) on the observation angle $\theta_{x}$ in the far zone can be regarded as the two sides of the same coin: they are manifestation of the cylindrical symmetry when an electron is moving along an arc of a circle.
 Because of cylindrical symmetry, in order to calculate spectral and angular photon distributions in the far zone, it is not necessary to consider an observer at a general location.
@@
 In this case observation angle $\theta_{x}=0$ and the observation angle $\theta_{y}$ is above the level of the orbit.
-In other words, in this very special geometry the $z$-axis is not fixed, but depends on the observer position.
+In other words, in this very special geometry, the $z$-axis is not fixed but depends on the observer’s position.
 However, this way of proceeding can hardly help to obtain radiation fields in the near zone.
-Indeed, in the near zone we are dealing with the Fresnel propagator, which obviously depends on the constrained motion of the electron.
+Indeed, in the near zone, we are dealing with the Fresnel propagator, which depends on the constrained motion of the electron.
 We use far-zone arguments only to show that there is no influence of the difference between the non-covariant and covariant trajectories on the synchrotron radiation from bending magnets.
@@
 19.4.3 Influence of the Kick According to Conventional Theory
-Up to this point we considered an electron moving along a circular trajectory that lies in the $(x,z)$-plane and tangent to the $z$ axis.
+Up to this point, we considered an electron moving along a circular trajectory that lies in the $(x,z)$-plane and tangent to the $z$ axis.
 The phase difference in the fields will be determined by the position of the observer and by the electron trajectory.
@@
 (155), Eq.
-(171) in the Appendix III) can be used to characterize the field from an electron moving on any trajectory.
+(158) in Appendix A3) can be used to characterize the field from an electron moving on any trajectory.
 Using Eq.
@@
 It is clear from the above that the field distribution in the far zone depends only on the observation angle with respect to the electron velocity direction.
-According to the conventional (incorrect) coupling of fields and particles, there is a prediction of radiation theory concerning to the bending magnet radiation from a single electron with and without kick.
+According to the conventional (incorrect) coupling of fields and particles, there is a prediction of radiation theory concerning the bending magnet radiation from a single electron with and without the kick.
 Namely, when a kick is introduced, there is a rigid rotation of the angular distribution in the far zone.
@@
 So, accelerator physicists did not appreciate that there was a contribution to the synchrotron radiation from relativistic kinematics effects.
-The question now arises how can storage rings actually operate.
+The question now arises how can storage rings operate?
 The point is that this example deals with a situation where electron beam kinetics is determined by the emission of synchrotron radiation from bending magnets.
-However, because of the cylindrical symmetry, covariant and non-covariant solutions for the electron motion along an arc of a circle yield similar properties of synchrotron radiation except the following modification.
+However, because of the cylindrical symmetry, covariant and non-covariant solutions for the electron motion along an arc of a circle yield similar properties of synchrotron radiation except for the following modification.
 The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion in the vertical direction.
-But synchrotron radiation from bending magnets is emitted within a wide range of frequencies, and the output intensity is not sensitive on the red shift.
+However synchrotron radiation from bending magnets is emitted within a wide range of frequencies, and the output intensity is not sensitive to the redshift.
 19.5 How to Solve Problems Involving Many Trajectory Kicks
-We shall now discuss the situation where there are $n$ arbitrary spaced kickers, all different from one another in terms of the rotation angle introduced.
+We shall now discuss the situation where there are $n$ arbitrarily spaced kickers, all different from one another in terms of the rotation angle introduced.
 Let us consider how we may apply covariant particle tracking in this circumstance, and try to understand what is happening when we have for example an undulator downstream of the kicker setup.
@@
 However, this situation is not surprising, if one analyzes the general expression for the radiation field from a single electron Eq.(LABEL:revwied).
-In fact, we should note that, in general, one needs to know the entire history of the electron from $t^{\prime}=-\infty$ to $t^{\prime}=\infty$ since the integration in Eq.(LABEL:revwied) is performed (at least formally) between these limits.
+We should note that, in general, one needs to know the entire history of the electron from $t^{\prime}=-\infty$ to $t^{\prime}=\infty$ since the integration in Eq.(LABEL:revwied) is performed (at least formally) between these limits.
 However, this statement should be interpreted physically, depending on the situation under study: integration should in fact be performed from and up to times when the electron does not contribute to the field anymore.
-We should pointed out that it is the electrodynamics theory, which ultimately decides what part of the particle trajectory is important for calculating undulator radiation and what part can be neglected.
+We should point out that it is the electrodynamics theory, which ultimately decides what part of the particle trajectory is important for calculating undulator radiation and what part can be neglected.
 The most important, general statement concerning the relevant part of the particle trajectory, is that it must be calculated according to the covariant method (if one wants to use the usual Maxwell’s equations).
@@
 In other words, we now consider an electron moving along a standard synchrotron radiation setup.
-The electron enters the setup via a bending magnet, passes through a straight section, an undulator, and another straight section.
+The electron enters the setup via a bending magnet and passes through a straight section, an undulator, and another straight section.
 Finally, it leaves the setup via another bend.
-Note that, although the integration in Eq.(38) is performed from $\infty$- ∞ to $\infty$, the only (edge) part of the trajectory into the bending magnets contributing to the integral is of order of the radiation formation length $L_{f}$.
+Note that, although the integration in Eq.(43) is performed from $\infty$- ∞ to $\infty$, the only (edge) part of the trajectory into the bending magnets contributing to the integral is of the order of the radiation formation length $L_{f}$.
 Mathematically, it is reflected in the fact that $\Phi_{T}(z^{\prime})$ in Eq.(43) exhibits more and more rapid oscillations as $z^{\prime}$ becomes larger than the formation length.
-At the critical wavelength the formation length is simply of order of $R/\gamma$, $R$ being the radius of the bend.
-That simply corresponds to an orbiting angular interval $\Delta\theta\simeq 1/\gamma$.
-Typically, the critical wavelength of the radiation from a bending magnet in synchrotron radiation source is about 0.1 nm and the formation length in this case is only few millimeters.
+At the critical wavelength, the formation length is simply of the order of $R/\gamma$, $R$ is the radius of the bend.
+That corresponds to an orbiting angular interval $\Delta\theta\simeq 1/\gamma$.
+Typically, the critical wavelength of the radiation from a bending magnet in a synchrotron radiation source is about 0.1 nm.
+The formation length in this case is only a few millimeters.
 […]
@@
 Figure 57: Standard undulator radiation setup.
-When the electron passes through a bending magnet there is the synchrotron radiation, washing out the (Frank- Ginzburg) fields of the fast moving charge.
-At the exit of the bending magnet we have ”naked” electron.
-There is a process of formation of the field-dressed electron within the formation length from the very beginning of the straight section downstream the bending magnet.
-The field-dressed ultrarelativistic electron has a visible transverse size of order a few microns for third generation synchrotron radiation sources.
+When the electron passes through a bending magnet there is synchrotron radiation, washing out the (Frank- Ginzburg) fields of the fast-moving charge.
+At the exit of the bending magnet, we have a ”naked” electron.
+There is a process of formation of the field-dressed electron within the formation length from the very beginning of the straight section downstream of the bending magnet.
+The field-dressed ultrarelativistic electron has a visible transverse size of the order of a few microns for third-generation synchrotron radiation sources.
 […]
@@
 We start with the description of the field formation process along the straight section downstream of the switcher, based on the covariant approach.
-First of all we have to synchronize distant clocks within the lab frame.
+First of all, we have to synchronize distant clocks within the lab frame.
 The synchronization procedure that follows is the usual Einstein synchronization procedure.
 It is assumed that in the Lorentz lab frame the electron proceeds following a rectilinear trajectory with velocity $v$.
-This assumption is used as initial condition.
-Then we can analyze situation downstream the switcher by using the usual Maxwell’s equations.
+This assumption is used as an initial condition.
+Then we can analyze the situation downstream of the switcher by using the usual Maxwell’s equations.
 When one analyzes the process of ”field-dressed” electron formation from the viewpoint of the noncovariant approach, one assumes the same initial conditions (rectilinear trajectory with velocity $v$) for the electron motion.
-Then one solves the electrodynamics problem of fields formation by using the usual Maxwell’s equations.
-We already mentioned that the type of clock synchronization which results in time coordinate $t$ in an electron trajectory $\vec{x}(t)$ is never discussed in accelerator physics.
+Then one solves the electrodynamics problem of field formation by using the usual Maxwell’s equations.
+We already mentioned that the type of clock synchronization that results in the time coordinate $t$ in an electron trajectory $\vec{x}(t)$ is never discussed in accelerator physics.
 However, we know that the usual Maxwell’s equations are only valid in the Lorentz frame.
 The noncovariant approach is obviously based on a definite synchronization assumption, but this is actually a hidden assumption.
-In our case of interest, within the lab frame the Lorentz coordinates are then automatically enforced.
-So one should not be surprised to find that in this simple case of rectilinear motion there is no difference between covariant and non covariant calculations of the initial conditions at the undulator entrance.
+In our case of interest, within the lab frame, the Lorentz coordinates are then automatically enforced.
+So one should not be surprised to find that in this simple case of rectilinear motion, there is no difference between covariant and noncovariant calculations of the initial conditions at the undulator entrance.
 Because of the characteristics of undulator radiation, in order to calculate the radiation field within the central cone, we simply use the instantaneous (i.e.
@@
 What do we expect for the undulator radiation?
-At first glance the situation is similar to the switcher setup and the electron trajectory is again divided into two parts: before and after the kicker.
+At first glance, the situation is similar to the switcher setup, and the electron trajectory is again divided into two parts: before and after the kicker.
 The most important difference, however, is that electrodynamics now dictates that both trajectories are important for the calculation of the undulator radiation.
-When the electron passes through the kicker there is no synchrotron radiation (to be more precise, in this case radiation is indistinguishable from the self-electromagnetic fields of the electron), washing out the virtual radiation fields like in the switcher case.
-We expect that an electron that passes through a kicker is still ”field-dressed”, but we have an electron whose fields has been perturbed, and now include information about the acceleration with respect to an inertial frame.
-According to the conventional theory, as usual for Newtonian kinematics, the Galilean vectorial law of addition of velocities is actually used.
+When the electron passes through the kicker there is no synchrotron radiation (to be more precise, in this case, radiation is indistinguishable from the self-electromagnetic fields of the electron), washing out the virtual radiation fields like in the switcher case.
+We expect that an electron that passes through a kicker is still ”field-dressed”, but we have an electron whose field has been perturbed, and now include information about the acceleration with respect to an inertial frame.
+According to the conventional theory, as usual for Newtonian kinematics, the Galilean vectorial law of addition of velocities is used.
 Non-covariant particle dynamics shows that the direction of the electron trajectory changes after the kick, while its speed remains unvaried.
@@
 The electron speed decreases from $v$ to $v(1-\theta^{2}_{k}/2)$.
-This result is at odds with the prediction from non-covariant particle tracking, because the corrected relativistic addition law for non-parallel velocities is used to calculate the electron trajectory.
+This result is at odds with the prediction from noncovariant particle tracking because the corrected relativistic addition law for non-parallel velocities is used to calculate the electron trajectory.
 According to the conventional algorithm for solving electrodynamics field equations, which deals with the usual Maxwell’s equations, and particle trajectories calculated by using non-covariant particle tracking, the undulator radiation along the velocity direction has no redshift of resonance frequency for any kick angle $\theta_{k}$.
@@
 19.6 Synchrotron Radiation in the Case of Particle Motion on a Helix
-The presence of red shift in bending magnet radiation automatically implies the same problem for conventional cyclotron radiation theory.
+The presence of a redshift in bending magnet radiation automatically implies the same problem for conventional cyclotron radiation theory.
 In the ultrarelativistic limit, there are well-known analytical formulas that describe the spectral and angular distribution of cyclotron radiation emitted by an electron moving in a constant magnetic field having a non-relativistic velocity component parallel to the field, and an ultrarelativistic velocity component perpendicular to it.
@@
 Let us discuss in some detail the relativistic cyclotron radiation.
-Here we shall only give some final results and discuss their relation with the conventional synchrotron radiation theory from bending magnet.
-In the case of an uniform translation motion with non-relativistic velocity along the magnetic field direction (Fig.
-69), a widely accepted (in astrophysics) expression for the angular and spectral distributions of radiation from an ultra-relativistic electron on a helical orbit is given by (2)
+Here we shall only give some final results and discuss their relation with the conventional synchrotron radiation theory from bending magnets.
+In the case of a uniform translation motion with non-relativistic velocity along the magnetic field direction (Fig.
+58), a widely accepted (in astrophysics) expression for the angular and spectral distributions of radiation from an ultra-relativistic electron on a helical orbit is given by  [^66] A widely accepted expression for the angular and spectral distributions of radiation from an ultra-relativistic electron on a helical orbit were calculated in [69, 70].
+At present, relativistic cyclotron radiation results are textbook examples (see e.g.
+[71]) and do not require a detailed description.
 ${\vec{\widetilde{E}}}(\chi,\alpha)$$\sim\Bigg{\{}\vec{e}_{x}\left[(\xi^{2}+\psi^{2})K_{2/3}\left(\frac{\omega}{2\omega_{c}}\left(1+\frac{\psi^{2}}{\xi^{2}}\right)^{3/2}\right)\right]$ (105)
 $-i\vec{e}_{y}\left[(\xi^{2}+\psi^{2})^{1/2}\psi K_{1/3}\left(\frac{\omega}{2\omega_{c}}\left(1+\frac{\psi^{2}}{\xi^{2}}\right)^{3/2}\right)\right]\Bigg{\}}~{},$
-where $K_{1/3}$ and $K_{2/3}$ are the modified Bessel functions, $\xi=1/\gamma$, $\psi=\chi-\alpha$, ($\chi$ is the angle between $\vec{v}$ and $\vec{B}$ and $\alpha$ that between $\vec{n}$ and $\vec{B}$); the angle $\psi$ is clearly the angular distance between the direction of the electron velocity $\vec{v}$ and the direction of observation $\vec{n}$.
+where $K_{1/3}$ and $K_{2/3}$ are the modified Bessel functions, $\xi=1/\gamma$, $\psi=\chi-\alpha$, ($\chi$ is the angle between $\vec{v}$ and $\vec{B}$ and $\alpha$ that between $\vec{n}$ and $\vec{B}$); the angle $\psi$ is the angular distance between the direction of the electron velocity $\vec{v}$ and the direction of observation $\vec{n}$.
 Here the $\omega_{c}$ is defined by $3eB\gamma^{2}/(2mc)$.
 Figure 58: Geometry for radiation production from helical motion.
-Actually we have already discussed radiation from an ultrarelativistic electron on a helical orbit in the previous section.
+Actually, we have already discussed radiation from an ultrarelativistic electron on a helical orbit in the previous section.
 Equation Eq.
-(96) is the result we worked out above for the bending magnet radiation from a single electron with angular deflection with respect to nominal orbit.
+(101) is the result we worked out above for the bending magnet radiation from a single electron with angular deflection with respect to the nominal orbit.
 Eq.
-( 100) does not look the same as Eq.
+( 105) does not look the same as Eq.
 (101).
@@
 […]
-Then, making the necessary variable changes, the formula reduces to Eq.( 100).
+Then, making the necessary variable changes, the formula reduces to Eq.( 105).
 19.6.2 Methodology of Solving Problems Involving Boosts
 The calculation leading to Eq.
-( 100) is rather elaborate.
+( 105) is rather elaborate.
 It is therefore desirable to have an independent derivation.
@@
 Eq.
-( 100) holds, indeed, in the frame $S$ for a particle whose velocity is $(v_{x},v_{y},v_{z})=(v_{0}\sin\chi\sin\phi,v_{0}\cos\chi,v_{0}\sin\chi\cos\phi)$.
+( 105) holds, indeed, in the frame $S$ for a particle whose velocity is $(v_{x},v_{y},v_{z})=(v_{0}\sin\chi\sin\phi,v_{0}\cos\chi,v_{0}\sin\chi\cos\phi)$.
 The Lorentz transformation, which leads to the value $v_{y}=v_{0}\cos\chi$ for the $y$-component of the velocity yields $(v_{x},v_{y},v_{z})=(v^{\prime}\sin\phi^{\prime}/\gamma_{y},v_{y},v^{\prime}\cos\phi^{\prime}/\gamma_{y})$, where $\gamma_{y}=1/\sqrt{1-v^{2}_{y}/c^{2}}$, $v^{\prime}$ is the velocity of the electron in the frame $S^{\prime}$ and the phase angle $\phi^{\prime}=\phi$ is invariant.
@@
 If one transforms the radiation field for a particle in a circular motion in the system $S^{\prime}$, one obtains the result that the effect of a boost amounts to a rigid rotation of the angular-spectral distribution of the radiation emitted by the electron moving with velocity $v_{0}$ on a circle that is, once more, Eq.
-( 100) (3) .
+( 105)  [^67] The covariant way of analyzing the radiation for helical motion was considered in [72].
+It is generally believed that $\vec{x}(t)=\vec{x}(t)_{cov}$ and this is the reason why in the [72] there is no distinction between the two (noncovariant and covariant) approaches to describe the electron motion on a helix downstream of the kicker setup..
 The notation ”” used here to indicate the velocity in the transverse to the magnetic field direction should not be confused with the ”” referring to an acceleration of the electron in the transverse direction in the proceeding sections.
 It comes out quite naturally that the covariant way of analyzing the radiation for helical motion considered above is based on the Lorentz transformation.
-In other words, within the lab frame the Lorentz coordinates are automatically enforced.
+In other words, within the lab frame, the Lorentz coordinates are automatically enforced.
 It assumed that in the Lorentz lab frame the electron proceeds following a helical trajectory with velocity $v_{0}$.
-This is employed as initial condition.
+This is employed as an initial condition.
 In the ultrarelativistic approximation a Lorentz boost along the field direction with non relativistic velocity $v_{y}$ leads to the circular motion of the electron with the same velocity $v_{0}$.
@@
 The restriction to the second order provides an essential simplicity of the calculations.
-We can use a sequence of two commuting non-collinear Lorentz boosts in order to discuss the particle motion downstream the kicker.
-When the kick is introduced, the covariant particle tracking predicts a change of the initial condition at the entrance of the synchrotron radiation setup.
-If we will keep the Lorentz coordinate system in the lab frame downstream of the kicker, we will find that the covariant velocity on the helical orbit after the kick decreases from $v$ to $v-v\theta^{2}_{k}/2$ and the covariant way of analyzing the radiation for a helical motion with covariant velocity $v_{0}=v-v\theta^{2}_{k}/2$ considered above will leads to a red shift in the critical wavelength.
-It should be note that the single passive Lorentz boost to the reference system $S^{\prime}$ discussed above is only a mathematical trick.
+We can use a sequence of two commuting non-collinear Lorentz boosts in order to discuss the particle motion downstream of the kicker.
+When the kick is introduced, the covariant particle tracking predicts a change in the initial condition at the entrance of the synchrotron radiation setup.
+If we keep the Lorentz coordinate system in the lab frame downstream of the kicker, we will find that the covariant velocity on the helical orbit after the kick decreases from $v$ to $v-v\theta^{2}_{k}/2$ and the covariant way of analyzing the radiation for a helical motion with covariant velocity $v_{0}=v-v\theta^{2}_{k}/2$ considered above will lead to a red shift in the critical wavelength.
+It should be noted that the single passive Lorentz boost to the reference system $S^{\prime}$ discussed above is only a mathematical trick.
 A good way to think of this transformation is to regard it as a result of transformation to new variables.
-According to this kinematic transformation, the electron (and the observer) motion relative to the fixed stars does not changes.
+According to this kinematic transformation, the electron’s and the observer’s motion relative to the fixed stars does not change.
 An absolute (i.e.
-physically real) acceleration is always a dynamical process.
-We must conclude that when we accelerate with respect to the fixed stars an electron in the lab inertial frame upstream the uniform magnetic field, this acceleration is absolute and the information about this acceleration is included into the covariant trajectory.
+physically real) acceleration is always a dynamic process.
+We must conclude that when we accelerate with respect to the fixed stars an electron in the lab inertial frame upstream of the uniform magnetic field, this acceleration is absolute and the information about this acceleration is included in the covariant trajectory.
 19.6.3 On the Advanced ”Paradox” Related to the Coupling Fields and Particles
@@
 Let us see what happens with a kicker, which is installed in the straight section upstream of the bending magnet and is characterized by a kick angle $\theta_{k}=\theta_{0}$.
-When the kick in the $y$ direction is introduced, there is a red shift of the critical wavelength which arise because, according to corrected addition velocities law, the electron velocity decreases from $v$ to $v-v\theta_{0}^{2}/2$ after the kick.
+When the kick in the $y$ direction is introduced, there is a redshift of the critical wavelength which arises because, according to corrected addition velocities law, the electron velocity decreases from $v$ to $v-v\theta_{0}^{2}/2$ after the kick.
 The red shift of the critical frequency $\omega_{c}$ can be expressed by the formula $\Delta\omega_{c}/\omega_{c}=-(3/2)\gamma^{2}\theta_{0}^{2}$.
 We see a second order correction $\theta_{0}^{2}$ that is, however, multiplied by a large factor $\gamma^{2}$.
-The result of the covariant approach clearly depends on the absolute value of the kick angle $\theta_{0}$ and the radiation along the velocity direction has a red shift only when the kick angle has nonzero value.
+The result of the covariant approach depends on the absolute value of the kick angle $\theta_{0}$ and the radiation along the velocity direction has a redshift only when the kick angle has a nonzero value.
 The implicit ”absolute” acceleration means acceleration relative to the fixed stars.
@@
 It comes about as the result of the difference between two Lorentz coordinate systems in the lab frame.
-By trying to accelerate the electron upstream the bending magnet we have changed Lorentz coordinates for that particular source.
+By trying to accelerate the electron upstream of the bending magnet we have changed Lorentz coordinates for that particular source.
 We know that in order to keep a Lorentz coordinates system in the lab frame after the kick we need to perform a clock resynchronization.
@@
 When we do not perturb the electron motion (relative to the fixed stars) upstream of the bending magnet, no clock resynchronization takes place, while when we do perturb the motion, clock resynchronization is introduced.
-Figure 70: Two sets of initial conditions resulting in the same uniform motion along the magnetic field direction in the case of absolute time coordinatization in the lab frame.
+Figure 59: Two sets of initial conditions result in the same uniform motion along the magnetic field direction in the case of absolute time coordinatization in the lab frame.
 The magnitude of the electron velocity and the orientation of the velocity vector with respect to the magnetic field vector are identical in both setups.
 We would now like to describe an apparent paradox.
-The argument that the difference between these two situations, ending with a final uniform translation along the magnetic fields direction, is paradoxical can be summarized in the following way: in the case of absolute time coordinatization in the lab frame, the initial conditions at the bending magnet entrance are apparently identical.
+The argument that the difference between these two situations, ending with a final uniform translation along the magnetic field direction, is paradoxical can be summarized in the following way: in the case of absolute time coordinatization in the lab frame, the initial conditions at the bending magnet entrance are apparently identical.
 In fact, the magnitude of the electron velocity and the orientation of the velocity vector with respect to the magnetic field vector are identical in both setups.
-We must conclude that when we accelerate the electron in the lab frame upstream the bending magnet, the information about this acceleration is not included into the non-covariant trajectory.
+We must conclude that when we accelerate the electron in the lab frame upstream of the bending magnet, the information about this acceleration is not included in the noncovariant trajectory.
 Where is the information about the electron acceleration recorded in the case of absolute time coordinatization?
 Since an electron is a structureless particle, the situation seems indeed paradoxical.
-The above statement includes one delicate point.
-It is not true that an ultrarelativistic electron is a ”structureless particle”.
-Electrodynamics deals with observable quantities.
-Let us consider the measurement of the red shift in the bending magnet radiation from our kicked electron.
-We can measure the accurate value of the red shift using a spectrometer in the lab frame, and this leads to a description of the setup in the space-frequency domain.
 […]
@@
 According to the old kinematics, the orientation of the virtual radiation phase front is unvaried.
-However, Maxwell’s equations do not remain invariant with respect to Galilean transformations and, as discuss throughout this book, the choice of the old kinematics implies using anisotropic field equations.
+However, Maxwell’s equations do not remain invariant with respect to Galilean transformations and, as discussed throughout this book, the choice of the old kinematics implies using anisotropic field equations.
 As a result, the phase front remains plane but the direction of propagation is not perpendicular to the phase front.
@@
 Then, having chosen the absolute time synchronization, electrodynamics predicts that the virtual radiation beam propagates in the kicked direction with the phase front tilt $\theta_{k}$.
-This is the key to the ”paradox” discussed here.
 The information about the electron acceleration is recorded in the perturbation of the self-electromagnetic fields of the electron.
-Mathematically information is recorded in the phase front tilt of the virtual radiation beam.
-15.7 Bibliography and Notes
-1.
-It should be note that results of the beam splitting experiment at LCLS confirm our correction for spontaneous undulator emission [62].
-It apparently demonstrated that after a modulated electron beam is kicked on a large angle compared to the divergence of the XFEL radiation, the modulation wavefront is readjusted along the new direction of the motion of the kicked beam.
-Therefore, coherent radiation from the undulator placed after the kicker is emitted along the kicked direction practically without suppression (see the Chapter 16 for more detail).
-In the framework of the conventional theory, there is a second outstanding puzzle concerning the beam splitting experiment at the LCLS.
-In accordance with conventional undulator radiation theory, if the modulated electron beam is at perfect resonance without kick, then after the kick the same modulated beam must be at perfect resonance in the velocity direction.
-However, experimental results clearly show that when the kick is introduced there is a red shift in the resonance wavelength.
-The maximum power of the coherent radiation is reached when the undulator is detuned to be resonant to the lower longitudinal velocity after the kick [62].
-It should be remarked that any linear superposition of a given radiation field from single electrons conserves single-particle characteristics like parametric dependence on undulator parameters and polarization.
-Consider a modulated electron beam kicked by a weak dipole field before entering a downstream undulator.
-Radiation fields generated by this beam can be seen as a linear superposition of fields from individual electrons.
-Now experimental results clearly show that there is a red shift in the resonance wavelength for coherent undulator radiation when the kick is introduced.
-It follows that the undulator radiation from the single electron has red shift when the kick is introduced as well.
-This argument suggests that results of the beam splitting experiment in reference [62] confirm our correction for spontaneous undulator emission.
-2.
-A widely accepted expression for the angular and spectral distributions of radiation from an ultra-relativistic electron on a helical orbit were calculated in [57, 58].
-At present, relativistic cyclotron radiation results are textbook examples (see e.g.
-[59]) and do not require a detail description.
-3.
-The covariant way of analyzing the radiation for helical motion was considered in [60].
-It is generally believed that $\vec{x}(t)=\vec{x}(t)_{cov}$ and this is the reason why in the [60] there is no distinction between the two (non covariant and covariant) approaches to describe the electron motion on a helix downstream of the kicker setup.
 20 Relativity and X-Ray Free Electron Lasers
 20.1 Introductory Remarks
-In the previous chapter we attempted to answer the question of why the error in radiation theory should have so long remained undetected.
-According to covariant approach, the various relativistic kinematics effects concerning to the synchrotron radiation setup, turn up in successive orders of approximation.
+In the previous chapter, we attempted to answer the question of why the error in radiation theory should have so long remained undetected.
+According to the covariant approach, the various relativistic kinematics effects concerning the synchrotron radiation setup, turn up in successive orders of approximation.
 Instead of small (total) velocity parameter $(v/c)$ in the non-relativistic case, we use a small transverse velocity parameter $(v_{\perp}/c)$.
@@
 It is demonstrated that due to a combination of the ultrarelativistic (i.e.
-paraxial) approximation and a very special symmetry of the conventional synchrotron radiation setup there is a cancellation of the second order relativistic kinematics effects except the non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion in the (bending) magnetic field direction.
-But synchrotron radiation from a bending magnets is emitted within a wide range of frequencies, and the output intencity is not sensitive on the red shift.
+paraxial) approximation and a very special symmetry of the conventional synchrotron radiation setup, there is a cancellation of the second-order relativistic kinematics effects except the non-zero redshift of the critical frequency, which arises when there are perturbations of the electron motion in the (bending) magnetic field direction.
+However, synchrotron radiation from bending magnets is emitted within a wide range of frequencies, and the output intensity is not sensitive on the redshift.
 That means that the spontaneous synchrotron radiation does not show sensitivity to the difference between covariant and non-covariant particle trajectories.
-But in the 21st century with the operation XFELs this situation changes.
-An XFEL is an example where the first order kinematics term $(v_{\perp}/c)$ plays an essential role in the description of the XFEL radiation and, in this case, covariant coupling of fields and particles predicts an effect in complete contrast to the conventional treatment.
-In this chapter we present a critical reexamination of existing XFEL theory.
-The main emphasize of this chapter is on coherent undulator radiation from the modulated electron beam.
+But in the 21st century with the operation XFELs, this situation changes.
+An XFEL is an example where the first-order kinematics term $(v_{\perp}/c)$ plays an essential role in the description of the XFEL radiation and, in this case, covariant coupling of fields and particles predicts an effect in complete contrast to the conventional treatment.
+In this chapter, we present a critical reexamination of existing XFEL theory.
+The main emphasis of this chapter is on coherent undulator radiation from the modulated electron beam.
 This chapter mainly addressed to readers with limiting knowledge of accelerator and XFEL physics.
 Figure 60: A well-known result of conventional (non-covariant) particle tracking.
-A micro-bunching electron beam passing through a weak dipole magnet (kicker) and undergoes a kick of an angle $\theta_{k}$.
+A micro-bunching electron beam passes through a weak dipole magnet (kicker) and undergoes a kick of an angle $\theta_{k}$.
 The propagation axis of the electron beam is deflected, while the wavefront orientation is preserved.
-The usual XFEL theory based on the use of old Newtonian kinematics for particle dynamics and the Einstein’s kinematics for the electrodynamics.
+The usual XFEL theory is based on the use of old Newtonian kinematics for particle dynamics and Einstein’s kinematics for electrodynamics.
 In fact, the usual theoretical treatment of relativistic particle dynamics involves only a corrected Newton’s second law and is based on the use Galilean edition of velocities.
-For rectilinear motion of the modulated electron beam, non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking.
+For the rectilinear motion of the modulated electron beam, non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking.
 However, one of the consequences of the relativity of simultaneity (i.e.
 mixture of positions and time) is a difference between covariant and non-covariant kinematics of a modulated electron beam in a given magnetic field.
-The theory of relativity shows that discussed above difference related with the acceleration along curved trajectories.
-There are several cases where the first order relativistic effect can occur in XFELs, mainly through the introduction of an trajectory kick (1).
-The most elementary of the effect that represents a crucial test of the correct coupling fields and particles is a problem involves the production of coherent undulator radiation by modulated ultrarelativistic electron beam kicked by a weak dipole field before entering a downstream undulator.
-It would be well to begin with bird’s view of some of the main results.
+The theory of relativity shows that the discussed above difference is related to the acceleration along curved trajectories.
+There are several cases where the first-order relativistic effect can occur in XFELs, mainly through the introduction of a trajectory kick  [^68] An angular kick is often an essential part of many XFEL-related diagnostic or experimental procedures.
+The standard gain length measurement procedure in XFELs employs such kicks.
+Other applications include ”beam-splitting” schemes where different polarization components are separated by means of an angular kick to the modulated electron beam [73, 74]..
+The most elementary of the effects that represent a crucial test of the correct coupling fields and particles is a problem involving the production of coherent undulator radiation by modulated ultrarelativistic electron beam kicked by a weak dipole field before entering a downstream undulator.
+It would be well to begin with a bird’s view of some of the main results.
 Let us now move on to consider the predictions of the existing XFEL theory in the case of non-collinear electron beam motion.
-As well-known result of conventional particle tracking states that after an electron beam is kicked by a weak dipole magnet there is a change in the trajectory of the electron beam, while the orientation of the modulated wavefront remains as before (Fig.
-71).
+A well-known result of conventional particle tracking states that after an electron beam is kicked by a weak dipole magnet there is a change in the trajectory of the electron beam, while the orientation of the modulated wavefront remains as before (Fig.
+60).
 In other words, the kick results in a difference between the directions of the electron motion and the normal to the modulation wavefront (i.e.
@@
 In existing XFEL theory the wavefront tilt is considered as real.
-According to conventional treatment, a transverse kick does not change the orientation of a modulation wavefront, and hence suppresses the radiation emitted in the direction of the electron motion (2).
+According to conventional treatment, a transverse kick does not change the orientation of a modulation wavefront and hence suppresses the radiation emitted in the direction of the electron motion  [^69] In a typical configuration for an XFEL, the orbit of a modulated electron beam is controlled to avoid large excursions from the undulator axis.
+All existing XFEL codes are based on a model in which the modulated electron beam moves only along the undulator axis.
+However, random errors in the focusing system can cause angular trajectory errors (or ”kicks”).
+The discrepancy between the directions of the electron motion and wavefront normal after the kick has been discussed in the literature.
+It was pointed out that coherent radiation is emitted towards the modulation wavefront normal.
+Thus, according to the conventional coupling of fields and particles (which we claimed was incorrect), the discrepancy between the two directions decreases the radiation efficiency [75].
+Analysis of the trajectory errors on the XFEL amplification process showed that any XFEL undulator magnetic field must satisfy stringent requirements.
+The pleasant surprise is that the tolerances predicted are more stringent than they need to be according to the corrected (4D) XFEL theory.
+This can be considered as one of the reasons for the exceptional progress in XFEL developments over the last decades..
 Figure 61: A result of covariant particle tracking.
 In the ultrarelativistic limit, the orientation of the modulation wavefront, i.e.
-the orientation of plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the modulation electron beam is treated using Lorentz coordinates.
-The theory of relativity dictates that a modulated electron beam in the ultrarelativistic limit has the same kinematics, in Lorentz coordinates, as laser beam.
+the orientation of the plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the modulation electron beam is treated using Lorentz coordinates.
+The theory of relativity dictates that a modulated electron beam in the ultrarelativistic limit has the same kinematics, in Lorentz coordinates, as a laser beam.
 According to Maxwell’s equations, the wavefront of a laser beam is always orthogonal to the propagation direction.
 The covariant approach within the framework of both mechanics and electrodynamics predicts an effect in complete contrast to the conventional treatment.
-Namely, in the ultrarelativistic limit, the wavefront of modulation, that is a plane of simultaneity, is always perpendicular to the electron beam velocity (Fig.
-72).
-As a result, the Maxwell’s equations predict strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction.
-Experiments show that this prediction is, in fact, true (3).
-The results of XFEL experiments demonstrated that even the direction of emission of coherent undulator radiation is beyond the predictive power of the conventional theory (4).
-It is worth remarking that the absent of a dynamical explanation for the modulation wavefront readjusting in the Lorentz coordinatization has disturbed some XFEL experts.
-We only wish to emphasize that a good way to think of the modulation wavefront readjusting is to regard it as a result of transformation to a new time variable in the framework of the Galilean (”single frame”) electrodynamics.
+Namely, in the ultrarelativistic limit, the wavefront of modulation, a plane of simultaneity, is always perpendicular to the electron beam velocity (Fig.
+61).
+As a result, Maxwell’s equations predict a strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction.
+Experiments show that this prediction is, in fact, true.
+The results of XFEL experiments demonstrated that even the direction of emission of coherent undulator radiation is beyond the predictive power of the conventional theory, see Section 20.6 for more details.
+It is worth remarking that the absence of a dynamical explanation for the modulation wavefront readjusting in the Lorentz coordinatization has disturbed some XFEL experts.
+We only wish to emphasize that a good way to think of the modulation wavefront readjusting is to regard it as a result of the transformation to a new time variable in the framework of the Galilean (”single frame”) electrodynamics.
 20.2 Modulation Wavefront Orientation
@@
 If we have adopted a method for timing distant events (i.e.
-a synchronization convention), we can also specify a method for measuring the orientation of the modulation wavefront: if electrons located at the position with maximum density cross the $y$-axis simultaneously at certain position $z$, then the modulation wavefront is perpendicular to $z$-axis.
+a synchronization convention), we can also specify a method for measuring the orientation of the modulation wavefront: if electrons located at the position with maximum density cross the $y$-axis simultaneously at a certain position $z$, then the modulation wavefront is perpendicular to $z$-axis.
 In other words, the modulation wavefront is defined as a plane of simultaneous events (the events being the arrival of particles located at maximum density): in short, a ”plane of simultaneity”.
-Let us formulate the initial conditions in the lab frame in terms of orientation of the modulation wavefront and beam velocity.
-Suppose that $v_{z}$ is the velocity of the comoving frame $R(\tau)$ with respect to the lab frame $K(\tau)$ along the common $z$-axis in positive direction.
-In the lab frame we select a special type of coordinate system, a Lorentz coordinate system to be precise.
+Let us formulate the initial conditions in the lab frame in terms of the orientation of the modulation wavefront and beam velocity.
+Suppose that $v_{z}$ is the velocity of the comoving frame $R(\tau)$ with respect to the lab frame $K(\tau)$ along the common $z$-axis in a positive direction.
+In the lab frame, we select a special type of coordinate system, a Lorentz coordinate system to be precise.
 Within a Lorentz frame (i.e.
 inertial frame with Lorentz coordinates), Einstein’s synchronization of distant clocks and Cartesian space coordinates $(x,y,z)$ are enforced.
-In order to have this, we impose that $R$ is connected to $K$ by the Lorentz boost $L(\vec{v}_{z})$, with $\vec{v}_{z}$, which transforms a given four vector event $X$ in space-time into $X_{R}=L(\vec{v_{z}})X$.
+To have this, we impose that $R$ is connected to $K$ by the Lorentz boost $L(\vec{v}_{z})$, with $\vec{v}_{z}$, which transforms a given four vector event $X$ in space-time into $X_{R}=L(\vec{v_{z}})X$.
 We now consider the acceleration of the beam in the lab frame up to velocity $v_{y}$ along the $y$-axis.
-The question arises how to assign synchronization in the lab frame after the beam acceleration.
-Before acceleration we picked a Lorentz coordinate system.
-Then, after the acceleration, the beam velocity changes of an small value $v_{y}$ along the $y$-axis.
-Without changing synchronization in the lab frame after the particle acceleration we have a complicated situation as concerns electrodynamics of moving charges.
-As a result of such boost, the transformation of time and spatial coordinates has the form of a Galilean transformation.
-In order to keep a Lorentz coordinate system in the lab frame after acceleration, one needs to perform a clock resynchronization by introducing the time shift $t\to t+yv_{y}/c^{2}$.
-This form of time transformation is justified by the fact that we are dealing with a first order approximation.
+The question arises of how to assign synchronization in the lab frame after the beam acceleration.
+Before acceleration, we picked a Lorentz coordinate system.
+Then, after the acceleration, the beam velocity changes of a small value $v_{y}$ along the $y$-axis.
+Without changing synchronization in the lab frame after the particle acceleration we have a complicated situation concerning the electrodynamics of moving charges.
+As a result of such a boost, the transformation of time and spatial coordinates has the form of a Galilean transformation.
+To keep a Lorentz coordinate system in the lab frame after acceleration, one needs to perform a clock resynchronization by introducing the time shift $t\to t+yv_{y}/c^{2}$.
+This form of time transformation is justified by the fact that we are dealing with a first-order approximation.
 Therefore, $v_{y}/c$ is so small that $v_{y}^{2}/c^{2}$ can be neglected and one arrives at the coordinate transformation $y\to y+v_{y}t$, $t\to t+yv_{y}/c^{2}$.
-The Lorentz transformation just described differs from a Galilean transformation just by the inclusion of the relativity of simultaneity, which is only relativistic effect that appearing in the first order in $v_{y}/c$.
+The Lorentz transformation just described differs from a Galilean transformation just by the inclusion of the relativity of simultaneity, which is the only relativistic effect that appears in the first order in $v_{y}/c$.
 The relation $X_{R}=L(\vec{v_{z}})L(\vec{v_{y}})X$ presents a step-by-step change from the lab reference frame $K(\tau+d\tau)$ to $K(\tau)$ and then to the proper reference frame $R$.
-The shift in the time when electrons located at the position with maximum density cross the $y$-axis of the lab frame $\Delta t=yv_{y}/c^{2}$ has the effect of a rotation the modulation wavefront on the angle $v_{z}\Delta t/y=v_{z}v_{y}/c^{2}$ in the first order approximation.
+The shift in the time when electrons located at the position with maximum density cross the $y$-axis of the lab frame $\Delta t=yv_{y}/c^{2}$ has the effect of a rotation of the modulation wavefront on the angle $v_{z}\Delta t/y=v_{z}v_{y}/c^{2}$ in the first order approximation.
 In ultrarelativistic limits, $v_{z}\simeq c$, and the modulation wavefront rotates exactly as the velocity vector $\vec{v}$.
 What does this wavefront readjustment mean in terms of measurements?
-In the absolute time coordinatization the simultaneity of a pair of events has absolute character.
+In absolute time coordinatization the simultaneity of a pair of events has absolute character.
 The absolute character of the temporal coincidence of two events is a consequence of the absolute time synchronization convention.
 According to this old kinematics, the modulation wavefront remains unvaried.
-However, according to the covariant approach we establish a criterion for the simultaneity of events, which is based on the invariance of the speed of light.
+However, according to the covariant approach, we establish a criterion for the simultaneity of events, which is based on the invariance of the speed of light.
 It is immediately understood that, as a result of the motion of electrons along the $y$ axis (i.e.
@@
 This reasoning is in analogy with Einstein’s train-embankment thought experiment.
-The wavefront orientation has no exact objective meaning, because the relativity of simultaneity takes place.
-The statement that the wavefront orientation has objective meaning to within a certain accuracy can be visualized by the picture of wavefront in the proper orientation with approximate angle extension (blurring) given by $\Delta\theta\simeq v_{z}(v_{y}/c^{2})$.
+The wavefront orientation has no exact objective meaning because the relativity of simultaneity takes place.
+The statement that the wavefront orientation has objective meaning within a certain accuracy can be visualized by the picture of the wavefront in the proper orientation with approximate angle extension (blurring) given by $\Delta\theta\simeq v_{z}(v_{y}/c^{2})$.
 This relation specifies the limits within which the nonrelativistic theory can be applied.
-In fact, it follows that for a very non relativistic electron beam for which $v_{z}^{2}/c^{2}$ is very small, the angle ”blurring” becomes very small too.
+It follows that for a very nonrelativistic electron beam for which $v_{z}^{2}/c^{2}$ is very small, the angle ”blurring” becomes very small too.
 In this case angle of wavefront tilt $\theta=v_{y}/v_{z}$ is practically sharp $\Delta\theta/\theta\simeq v_{z}^{2}/c^{2}\ll 1$.
@@
 20.3 XFEL Radiation Setup
-The most elementary of the effect that represents a crucial test of the correct coupling fields and particles is a problem involves the production of coherent undulator radiation by modulated ultrarelativistic electron beam kicked by a weak dipole field before entering a downstream undulator.
-We want to study the process of emission of coherent undulator radiation from such setup.
-The key element of a XFEL source is the udulator, which forces the electrons to move along curved periodical trajectories.
+The most elementary of the effects that represent a crucial test of the correct coupling fields and particles is a problem involving the production of coherent undulator radiation by modulated ultrarelativistic electron beam kicked by a weak dipole field before entering a downstream undulator.
+We want to study the process of emission of coherent undulator radiation from such a setup.
+The key element of an XFEL source is the undulator, which forces the electrons to move along curved periodical trajectories.
 There are two popular undulator configurations: helical and planar.
-To understand the basic principles of undulator source operation, let us consider the helical undulator (it is interesting to note that the first XFEL experiment demonstrating the apparent wavefront readjusting used helical undulator (3)).
+To understand the basic principles of undulator source operation, let us consider the helical undulator.
 The magnetic field on the axis of the helical undulator is given by
@@
 We neglected the transverse variation of the magnetic field.
-It is necessary to mention that in XFEL engineering we deal with a very high quality of the undulator systems, which have a sufficiently wide good-field-region, so that our studies, which refer to a simple model of undulator field nevertheless yields a correct quantitative description in large variety of practical problems.
+It is necessary to mention that in XFEL engineering we deal with very high-quality undulator systems, which have a sufficiently wide good-field-region, so that our studies, which refer to a simple model of undulator field nevertheless yield a correct quantitative description in a large variety of practical problems.
 The Lorentz force $\vec{F}=-e\vec{v}\times\vec{B}_{w}/c$ is used to derive the equation of motion of electrons with charge $-e$ and mass $m$ in the presence of magnetic field
@@
 As a rule, the electron rotation angle $\theta_{w}$ is small and the longitudinal electron velocity $v_{z}$ is close to the velocity of light, $v_{z}=\sqrt{v^{2}-v^{2}_{\perp}}\simeq v(1-\theta_{w}^{2}/2)\simeq c$.
-Let us consider a modulated ultrarelativistic electron beam moving alone the $z$ axis in the field of the helical undulator.
-In the present study we introduce the following assumptions.
+Let us consider a modulated ultrarelativistic electron beam moving alone on the $z$ axis in the field of the helical undulator.
+In the present study, we introduce the following assumptions.
 First, without kick the electrons move along constrained helical trajectories in parallel with the $z$ axis.
@@
 $n=n_{0}(\vec{r}_{\perp})[1+a\cos\omega(z/v_{z}-t)]~{},$ (114)
-where $a=\mathrm{const.}$ In other words we consider the case in which there are no variation in amplitude and phase of the density modulation in the transverse plane.
-Under these assumptions the transverse current density may be written in the form
+where $a=\mathrm{const.}$ In other words we consider the case in which there are no variations in amplitude and phase of the density modulation in the transverse plane.
+Under these assumptions, the transverse current density may be written in the form
 $\vec{j}_{\perp}=-e\vec{v}_{\perp}(z)n_{0}(\vec{r}_{\perp})[1+a\cos\omega(z/v_{z}-t)]~{}.$ (115)
-Even through the measured quantities are real, it is generally more convenient to use complex representation, starting with real $\vec{j}_{\perp}$, one defines the complex transverse current density:
+Even though the measured quantities are real, it is generally more convenient to use complex representation, starting with real $\vec{j}_{\perp}$, one defines the complex transverse current density:
 $j_{x}+ij_{y}=-ec\theta_{w}n_{0}(\vec{r}_{\perp})\exp(-ik_{w}z)[1+a\cos\omega(z/v_{z}-t)]~{}.$ (116)
@@
 Maxwell equations can be manipulated mathematically in many ways in order to yield derived equations more suitable for certain applications.
-For example, from Maxwell equations Eq.(8) we can obtain an equation which depends only on the electric field vector $\vec{E}$ (in Gaussian units):
+For example, from Maxwell equations Eq.(8) we can obtain an equation that depends only on the electric field vector $\vec{E}$ (in Gaussian units):
 $c^{2}\vec{\nabla}\times(\vec{\nabla}\times{\vec{E}})=-\partial^{2}\vec{E}/\partial t^{2}-4\pi\partial\vec{j}/\partial t~{}.$ (117)
@@
 Thus, this nonhomogeneous wave equation is the complete and correct formula for radiation.
-However we want to apply it to still simpler circumstance in which second term (or, the current term) in the right-hand side provides the main contribution to the value of the radiation field.
+However, we want to apply it to still simpler circumstances in which the second term (or, the current term) on the right-hand side provides the main contribution to the value of the radiation field.
 It is relevant to remember that our case of interest is the coherent undulator radiation and the divergence of this radiation is much smaller compared to the angle $1/\gamma$.
-It can be shown that when this condition is fulfilled the gradient term, $4\pi c^{2}\vec{\nabla}\rho$, in the right-hand side of the nonhomoheneous wave equation can be neglected.
+It can be shown that when this condition is fulfilled the gradient term, $4\pi c^{2}\vec{\nabla}\rho$, in the right-hand side of the nonhomogeneous wave equation can be neglected.
 Thus we consider the wave equation
@@
 This is the condition for synchronism between the transverse electromagnetic wave and the fast transverse current wave with the propagation constant $\omega/v_{z}-k_{w}$.
-With the current wave traveling with the same phase speed as electromagnetic wave, we have the possibility of obtaining a spatial resonance between electromagnetic wave and electrons.
-If this the case, a cumulative interaction between modulated electron beam and transverse electromagnetic wave in empty space takes place.
+With the current wave traveling at the same phase speed as the electromagnetic wave, we have the possibility of obtaining a spatial resonance between the electromagnetic wave and electrons.
+If this is the case, a cumulative interaction between the modulated electron beam and a transverse electromagnetic wave in empty space takes place.
 We are therefore justified in considering the contributions of all the waves except the synchronous one to be negligible as long as the undulator is made of a large number of periods.
 Here follows an explanation of the resonance condition which is elementary in the sense that we can see what is happening physically.
-The field of electromagnetic wave has only transverse components, so the energy exchange between the electron and electromagnetic wave is due to transverse component of the electron velocity.
+The field of electromagnetic wave has only transverse components, so the energy exchange between the electron and electromagnetic wave is due to a transverse component of the electron velocity.
 For effective energy exchange between the electron and the wave, the scalar product $-e\vec{v}_{\perp}\cdot\vec{E}$ should be kept nearly constant along the whole undulator length.
@@
 Since we are interested in coherent emission around the modulation wavelength the theory of coherent undulator radiation is naturally developed in the space-frequency domain.
-In fact, in this case one is usually interested into radiation properties at fixed modulation frequency.
+In fact, in this case, one is usually interested in radiation properties at fixed modulation frequency.
 We first apply a temporal Fourier transformation to the inhomogeneous wave equation to obtain the inhomogeneous Helmholtz equation
@@
 […]
-This is just a mathematical description of the concept of Huygens’ secondary sources and waves, and is of course well-known, but we still recalled how it follows directly from the Maxwell’s equations.
-We may consider the amplitude of the beam radiated by plane of oscillating electrons as a whole to be the resultant of radiated spherical waves.
+This is just a mathematical description of the concept of Huygens’ secondary sources and waves and is of course well-known, but we still recall how it follows directly from Maxwell’s equations.
+We may consider the amplitude of the beam radiated by the plane of oscillating electrons as a whole to be the resultant of radiated spherical waves.
 This is because Maxwell’s theory has no intrinsic anisotropy.
-The electrons lying on the plane of simultaneity gives rise to spherical radiated wavelets, and these combine according to Huygens’ principle to form what is effectively a radiated wave.
+The electrons lying on the plane of simultaneity give rise to spherical radiated wavelets, and these combine according to Huygens’ principle to form what is effectively a radiated wave.
 If the plane of simultaneity is the $xy$-plane (i.e.
-beam modulation wavefront is perpendicular to the $z$- axis), then the Huygens’ construction shows that plane wavefronts will be emitted along the $z$-axis.
+beam modulation wavefront is perpendicular to the $z$- axis), then Huygens’ construction shows that plane wavefronts will be emitted along the $z$-axis.
 In summary: according to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the modulation wavefront.
@@
 when an inertial frame where the Einstein synchronization procedure is used to assign values to the time coordinates.
-Einstein’s time order should be applied and kept in consistent way in both dynamics and electrodynamics.
+Einstein’s time order should be applied and kept in a consistent way in both dynamics and electrodynamics.
 Our previous description implies quite naturally that Maxwell’s equations in the lab frame are compatible only with covariant trajectories $\vec{x}_{cov}(t)$, calculated by using Lorentz coordinates and, therefore, including relativistic kinematics effects.
@@
 In other words, the electron motion and the wavefront normal have different directions.
-Therefore, according to conventional coupling of fields and particles that we deem incorrect, the coherent undulator radiation in the kicked direction produced in a downstream undulator is expected to be dramatically suppressed as soon as the kick angle is larger than the divergence of the output coherent radiation.
-In order to estimate the loss in radiation efficiency in the kicked direction according to the conventional coupling of fields and particles, we make the assumption that the spatial profile of the modulation is close to that of the electron beam and has a Gaussian shape with standard deviation $\sigma$.
+Therefore, according to the conventional coupling of fields and particles that we deem incorrect, the coherent undulator radiation in the kicked direction produced in a downstream undulator is expected to be dramatically suppressed as soon as the kick angle is larger than the divergence of the output coherent radiation.
+To estimate the loss in radiation efficiency in the kicked direction according to the conventional coupling of fields and particles, we make the assumption that the spatial profile of the modulation is close to that of the electron beam and has a Gaussian shape with standard deviation $\sigma$.
 A modulated electron beam in an undulator can be considered as a sequence of periodically spaced oscillators.
@@
 When all the oscillators are in phase there is, therefore, strong emission in the direction $\theta=0$.
-If we have a triangle with a small altitude $r\simeq\theta z$ and long base $z$, than the diagonal $s$ is longer than the base.
+If we have a triangle with a small altitude $r\simeq\theta z$ and long base $z$, then the diagonal $s$ is longer than the base.
 The difference is $\Delta=s-z\simeq z\theta^{2}/2$.
@@
 This is because in this case the contributions of various oscillators are uniformly distributed in phase from 00 to $2\pi$.
-In the limit for a small size of the electron beam, $\sigma\to 0$, the interference will be constructive within an angle of about $\Delta\theta\backsimeq\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$, where $L_{w}=\lambda_{w}N_{w}$ is the undulator length.
-In the limit for a large size of the electron beam, the angle of coherence is about $\Delta\theta\backsimeq c/(\omega\sigma)$ instead.
+In the limit for small size of the electron beam, $\sigma\to 0$, the interference will be constructive within an angle of about $\Delta\theta\backsimeq\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$, where $L_{w}=\lambda_{w}N_{w}$ is the undulator length.
+In the limit for the large size of the electron beam, the angle of coherence is about $\Delta\theta\backsimeq c/(\omega\sigma)$ instead.
 The boundary between these two asymptotes is for sizes of about $\sigma_{\mathrm{dif}}\backsimeq\sqrt{cL_{w}/\omega}$.
 The parameter $\omega\sigma^{2}/(cL_{w})$ can be referred to as the electron beam Fresnel number.
-It is worth noting that, for XFELs, the transverse size of electron beam $\sigma$ is typically much larger than $\sigma_{\mathrm{dif}}$ (i.e electron beam Fresnel number is large).
+It is worth noting that, for XFELs, the transverse size of electron beam $\sigma$ is typically much larger than $\sigma_{\mathrm{dif}}$ (i.e.
+electron beam Fresnel number is large).
 Thus, we can conclude that the angular distribution of the radiation power in the far zone has a Gaussian shape with standard deviation $\sigma_{\mathrm{\theta}}\backsimeq c/(\sqrt{2}\omega\sigma)$.
-However, still according to the conventional treatment, after the electron beam is kicked we have the already-mentioned discrepancy between direction of the electron motion and wavefront normal.
+However, still according to the conventional treatment, after the electron beam is kicked we have the already-mentioned discrepancy between the direction of the electron motion and wavefront normal.
 Then, the radiation intensity along the new direction of the electron beam can be approximated as $I\backsimeq I_{0}\exp[-\theta_{k}^{2}/(2\sigma_{\mathrm{\theta}}^{2})]$, where $I_{0}$ is the on-axis intensity without kick and $\theta_{k}$ is the kick angle.
 The exponential suppression factor is due to the tilt of the modulation wavefront with respect to the direction of motion of the electrons.
-We presented a study of very idealized situation for illustrating the difference between conventional and covariant coupling of fields and particles.
-We solved the dynamics problem of the motion of a relativistic electrons in the prescribed force field of weak kicker magnet by working only up to the order of $\gamma\theta_{k}$.
+We presented a study of the very idealized situation to illustrate the difference between the conventional and covariant coupling of fields and particles.
+We solved the dynamics problem of the motion of relativistic electrons in the prescribed force field of a weak kicker magnet by working only up to the order of $\gamma\theta_{k}$.
 This approximation is of particular theoretical interest because it is relatively simple and at the same time forms the basis for understanding relativistic kinematic effects such as relativity of simultaneity.
 Let us discuss the region of validity of our small kick angle approximation $\theta_{k}\gamma\ll 1$.
-Since in XFELs the Fresnel number is rather large, we can always consider a kick angle which is relatively large compared to the divergence of the output coherent radiation, and, at the same time, it is relatively small compared to the angle $1/\gamma$.
+Since in XFELs, the Fresnel number is rather large, we can always consider a kick angle that is relatively large compared to the divergence of the output coherent radiation, and, at the same time, it is relatively small compared to the angle $1/\gamma$.
 In fact, from $\omega\sigma^{2}/(cL_{w})\gg 1$, with some rearranging, we obtain $\sigma_{\theta}^{2}\simeq c^{2}/(\omega^{2}\sigma^{2})\ll c/(\omega L_{w})$.
 Then we recall that $\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$.
-Therefore, the first order approximation used to investigate the kicker setup in this chapter is of practical interest in XFEL engineering.
+Therefore, the first-order approximation used to investigate the kicker setup in this chapter is of practical interest in XFEL engineering.
 It is one of the aims of this chapter is to demonstrate the kind of experimental predictions we are expecting from our corrected radiation theory.
 We worked out a very simple case in order to illustrate all the essential physical principles very clearly.
-Surprisingly, the first order approximation used to investigate the kicker setup in this section has also important practical applications.
+Surprisingly, the first-order approximation used to investigate the kicker setup in this section has also important practical applications.
 Above we have shown that our correct coupling of fields and particles predicts an effect in complete contrast to the conventional treatment.
 Namely, in the ultrarelativistic limit, the plane of simultaneity, that is wavefront orientation of the modulation, is always perpendicular to the electron beam velocity.
-As a result, we predict strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction, Fig.
-73.
+As a result, we predict a strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction, Fig.
+62.
 Figure 62: Basic setup for the experimental test of XFEL theory.
-The correct coupling fields and particles predicts an effect in complete contrast to the conventional treatment.
 According to the covariant approach, in the ultrarelativistic limit, the wavefront of modulation is always perpendicular to the electron beam velocity.
-As a result, the Maxwell’s equations predict strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction.
+As a result, Maxwell’s equations predict a strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction.
 Experiments show that this prediction is, in fact, true.
 XFEL experts actually witnessed an apparent wavefront readjusting due to the relativistic kinematics effect, but they never drew this conclusion.
-In this book, we are actually first in considering the idea that results of the conventional theory of radiation by relativistically moving charges are not consistent with the special relativity.
+In this book, we are actually first in considering the idea that the results of the conventional theory of radiation by relativistically moving charges are not consistent with special relativity.
 In previous literature, identification of the trajectories in the source part of the usual Maxwell’s equations with the trajectories calculated by conventional particle tracking in the (”single”) lab frame has always been considered obvious.
 20.4 Modulation Wavefront Tilt in Maxwell’s Theory. Logical Inconsistency
-In existing literature theoretical analysis is presented, of an XFEL driven by an electron beam with wavefront tilt, and this analysis is based on the exploitation of usual Maxwell’s equations and standard simulation codes.
+In existing literature, theoretical analysis is presented, of an XFEL driven by an electron beam with wavefront tilt, and this analysis is based on the exploitation of usual Maxwell’s equations and standard simulation codes.
 Using only a kicker setup (i.e.
-without undulator radiation setup) we can demonstrate that the coupling fields and particles in the conventional XFEL theory is intrinsically incorrect.
-The existing XFEL theory based on the use of the absolute time convention (i.e.
+without undulator radiation setup) we can demonstrate that the coupling fields and particles in the conventional XFEL theory are intrinsically incorrect.
+The existing XFEL theory is based on the use of the absolute time convention (i.e.
 old kinematics) for particle dynamics.
-Here we will give a simple proof of the conflict between conventional particle tracking and Maxwell’s electrodynamics.
-The purpose is to show how one can demonstrate in a simple way that the conventional XFEL theory is absolutely incapable of correctly describing the distribution of the electromagnetic fields from a fast moving modulated electron beam downstream the kicker.
-Under the Maxwell’s electrodynamics, the fields of a modulated electron beam moving with a constant velocity exhibit an interesting behavior when the velocity of charges approaches that of light: namely, in the space-time domain they resemble more and more close of a laser beam (see Appendix IV).
+Here we will give simple proof of the conflict between conventional particle tracking and Maxwell’s electrodynamics.
+The purpose is to show how one can demonstrate in a simple way that the conventional XFEL theory is incapable of correctly describing the distribution of the electromagnetic fields from a fast-moving modulated electron beam downstream the kicker.
+Under Maxwell’s electrodynamics, the fields of a modulated electron beam moving with a constant velocity exhibit an interesting behavior when the velocity of charges approaches that of light: namely, in the space-time domain they resemble more and more closer to a laser beam (see Appendix IV).
 In fact, for a rapidly moving modulated electron beam we have nearly equal transverse and mutually perpendicular electric and magnetic fields: in the limit $v\to c$ they become indistinguishable from the fields of a laser beam, and according to Maxwell’s equations, the wavefront of the laser beam is always perpendicular to the propagation direction.
@@
 Let us now consider the effect of the kick on the electron modulation wavefront.
-If we rely on the conventional particle tracking, the kick results in a difference between the directions of electron motion and the normal to the modulation wavefront, i.e.
+If we rely on conventional particle tracking, the kick results in a difference between the directions of electron motion and the normal to the modulation wavefront, i.e.
 in a tilt of the modulation wavefront.
-This is already a conflict result, because we now conclude that, according to the conventional ”single frame” approach, the direction of propagation after the kick is not perpendicular to the radiation beam wavefront.
+This is already a conflict result because we now conclude that, according to the conventional ”single frame” approach, the direction of propagation after the kick is not perpendicular to the radiation beam wavefront.
 In other words, the radiation beam motion and the radiation wavefront normal have different directions.
 The virtual radiation beam (which is indistinguishable from a real radiation beam in the ultrarelativistic asymptote) propagates in the kicked direction with a wavefront tilt.
-This is what we would get for the case when our analysis is based on the conventional particle tracking, and is obviously absurd from the viewpoint of Maxwell’s electrodynamics.
-In existing literature theoretical analysis is presented, of an XFEL driven by an electron beam with wavefront tilt, and this analysis is based on the exploitation of usual Maxwell’s equations and standard simulation codes.
+This is what we would get for the case when our analysis is based on conventional particle tracking, and is absurd from the viewpoint of Maxwell’s electrodynamics.
+In existing literature, theoretical analysis is presented, of an XFEL driven by an electron beam with wavefront tilt, and this analysis is based on the exploitation of usual Maxwell’s equations and standard simulation codes.
 Using only a kicker setup (i.e.
-without undulator radiation setup) we demonstrated that the coupling fields and particles in the conventional XFEL theory is intrinsically incorrect.
+without undulator radiation setup) we demonstrated that the coupling fields and particles in the conventional XFEL theory are intrinsically incorrect.
 The difficulty above is a part of the continual problem of XFEL physics, which started with coherent undulator radiation from an ultrarelativistic modulated electron beam in the kicked direction, and now has been focused on the wavefront tilt of the self-electromagnetic fields of the modulated electron beam.
+20.5 Discussion
+Finally, we make some remarks on another way in which our complicated problem can be solved.
+We know the result for Lorentz coordinatization.
+Now let us understand why the emission of coherent undulator radiation from the modulated electron beam in the kicked direction exists in the framework of absolute time synchronization.
+The physics is actually very simple, and if we treat the production of coherent undulator radiation by kicked electron beam as the aberration of light we shall see that we can understand almost everything that happens in the undulator.
+Our Galilean-transformed electrodynamics says that by making a measurement of the coherent radiation, one can observe only radiation in the kicked direction.
+This is exactly analogous to the aberration (deviation of the energy transport) for light radiated from the single moving emitter in an inertial frame, Fig.
+1.
+There is an intuitively plausible way to understand the aberration of light from a kicked electron beam.
+The aberration of light in an inertial frame can be easily explained on the basis of corpuscular theory, Fig.
+39.
+This phenomenon is fully understandable in terms of the transformation of velocities between different reference frames.
+In the case of the transversely moving electron beam in an undulator, intuition would seem to tell us that the aberration increment would be the same.
+This is plausible if one keeps in mind that a (undulator) radiation pulse represents a certain amount of electromagnetic energy.
+Energy, like mass, is a quantity that is conserved, so that a coherent X-ray pulse resembles, in many aspects, material particles.
+Therefore, we should expect that group velocities of undulator radiation pulses obey the same addition theorem for particle velocities in the inertial lab frame.
+A closer treatment based on the wave theory of light confirms this expectation.
+Now let us demonstrate that the Galilean transformed electrodynamics will give a deviation of the energy transport direction for radiated X-ray pulse.
+The modulated electron beam with a finite aperture in an undulator is a kind of active medium that breaks up the radiated beam into a number of diffracted beams of plane waves.
+Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved.
+Let us assume that the density of the elementary source varies in the transverse direction according to the law $\rho_{e}=g(K_{\perp})\cos{({K}_{\perp}y)}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated electron beam density.
+From the Galilean transformation, after partial differentiation, one obtains wave equation Eq.(12).
+The new terms that have to be put into the field equations due to the use of Galilean transformation lead to the prediction of the Doppler effect.
+As one of the consequences of the Doppler effect in the absolute time coordinatization, we find an angular frequency dispersion of the light waves radiated from the kicked (modulated) electron beam with a finite aperture.
+The Doppler shift, $\Delta\omega$, of a radiated light wave (in the first order approximation) is given by $\Delta\omega=K_{\perp}v_{y}$, where $K_{\perp}$ is the transverse component of the radiated wavenumber vector.
+The last equation states that a radiated coherent X-ray beam with finite transverse size moves along the $y$ direction with group velocity $d\omega/dk_{y}=v_{y}$.
+Thus, according to correct coupling conventional particle tracking and electrodynamics, we identify the direction of X-ray beam propagation with the direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation.
+20.6 Experimental Test of Coupling between the Electrodynamics and Dynamics
+The fact that our theory predicts reality in a satisfactory way is well-illustrated by comparing the prediction we just made with the results of an experiment involving ”X-ray beam splitting” of a circularly-polarized XFEL pulse from the linearly-polarized XFEL background pulse, a technique used in order to maximize the degree of circular polarization.
+The XFEL experiments demonstrated that after a modulated electron beam is kicked on a large angle compared to the divergence of the XFEL radiation, the modulation wavefront is readjusted along the new direction of motion of the kicked beam, Fig.
+62.
+This is the only way to justify coherent radiation emission from the short undulator placed after the kicker and along the kicked direction, see Fig.
+14 in [76].
+These results came unexpectedly, but from a practical standpoint, the ”apparent wavefront readjusting” immediately led to the realization that the unwanted, linearly polarized radiation background could be fully eliminated without extra hardware.
+In existing literature a theoretical analysis of XFELs driven by an electron beam with wavefront tilt was presented in [77, 78], based on the use the Maxwell’s equations.
+In fact, the Maxwell solver was used as a part of the standard simulation code.
+We state that this approach is fundamentally incorrect.
+In ultrarelativistic asymptote, a modulation wavefront tilt is absurd from the viewpoint of Maxwell’s electrodynamics.
+In the case of an XFEL we deal with an ultrarelativistic electron beam and within the Lorentz lab frame (i.e.
+within the validity of Maxwell’s equations) the tilted modulation wavefront is at odds with the special relativity.
+It should be noted that the results of the beam splitting experiment at LCLS confirm our correction for spontaneous undulator emission [74].
+It apparently demonstrated that after a modulated electron beam is kicked on a large angle compared to the divergence of the XFEL radiation, the modulation wavefront is readjusted along the new direction of the motion of the kicked beam.
+Therefore, coherent radiation from the undulator placed after the kicker is emitted along the kicked direction practically without suppression (see Chapter 19 for more detail).
+In the framework of the conventional theory, there is a second outstanding puzzle concerning the beam-splitting experiment at the LCLS.
+In accordance with conventional undulator radiation theory, if the modulated electron beam is at perfect resonance without the kick, then after the kick the same modulated beam must be at perfect resonance in the velocity direction.
+However, experimental results clearly show that when the kick is introduced there is a redshift in the resonance wavelength.
+The maximum power of the coherent radiation is reached when the undulator is detuned to be resonant to the lower longitudinal velocity after the kick [74].
+It should be remarked that any linear superposition of a given radiation field from single electrons conserves single-particle characteristics like parametric dependence on undulator parameters and polarization.
+Consider a modulated electron beam kicked by a weak dipole field before entering a downstream undulator.
+Radiation fields generated by this beam can be seen as a linear superposition of fields from individual electrons.
+Now experimental results clearly show that there is a redshift in the resonance wavelength for coherent undulator radiation when the kick is introduced.
+It follows that the undulator radiation from the single electron has a redshift when the kick is introduced as well.
+This argument suggests that the results of the beam splitting experiment in reference [74] confirm our correction for spontaneous undulator emission.
+20.7 Myth about the Reality of the Modulation Wavefront Tilt
 In existing XFEL theory the wavefront tilt is considered as real.
@@
 $\theta_{tilt}=0$) when the evolution of the modulated electron beam is treated using Lorentz coordinates.
-No physical effects may depends on an arbitrary constant or an arbitrary function (5).
-16.5 Final Remarks
-Finally, we make some remarks on another way in which our complicated problem can be solved.
-We know the result for Lorentz coordinatization.
-Now let us understand why the emission of coherent undulator radiation from the modulated electron beam in the kicked direction exist in the framework of absolute time synchronization.
-The physics is actually very simple, and if we treat the production of coherent undulator radiation by kicked electron beam as the aberration of light we shall see that we can understand almost everything that happens in the undulator.
-Our Galilean transformed electrodynamics says that by making a measurement on the coherent radiation, one can observe only radiation in the kicked direction.
-This is exactly analogous to the aberration (deviation of the energy transport) for light radiated from the single moving emitter in an inertial frame, Fig.
-1.
-There is an intuitively plausible way to understanding the aberration of light from a kicked electron beam.
-The aberration of light in an inertial frame can be easily explained on the basis of corpuscular theory, Fig.
-48.
-This phenomenon is fully understandable in terms of transformation of velocities between different reference frames.
-In the case of the transversely moving electron beam in an undulator, intuition would seem to tell us that aberration increment would be the same.
-This is plausible if one keeps in mind that a (undulator) radiation pulse represents a certain amount of electromagnetic energy.
-Energy, like mass, is a quantity that is conserved, so that a coherent X-ray pulse resembles, in many aspects, a material particles.
-Therefore, we should expect that group velocities of undulator radiation pulses obey the same addition theorem for particle velocities in the inertial lab frame.
-A closer treatment based on wave theory of light confirm this expectation.
-Now let us demonstrate that the Galilean transformed electrodynamics will give a deviation of the energy transport direction for radiated X-ray pulse.
-The modulated electron beam with finite aperture in an undulator is a kind of active medium which breaks up the radiated beam into a number of diffracted beams of plane waves.
-Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved.
-Let us assume that the density of the elementary source varies in the transverse direction according to the law $\rho_{e}=g(K_{\perp})\cos{({K}_{\perp}y)}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated electron beam density.
-From the Galilean transformation, after partial differentiation, one obtains wave equation Eq.(12).
-The new terms that have to be put into the field equations due to use of Galilean transformation lead to the prediction of the Doppler effect.
-As one of the consequences of the Doppler effect in the absolute time coordinatization, we find an angular frequency dispersion of the light waves radiated from the kicked (modulated) electron beam with finite aperture.
-The Doppler shift, $\Delta\omega$, of radiated light wave (in the first order approximation) is given by $\Delta\omega=K_{\perp}v_{y}$, where $K_{\perp}$ is the transverse component of the radiated wavenumber vector.
-The last equation states that radiated coherent X-ray beam with finite transverse size moves along the $y$ direction with group velocity $d\omega/dk_{y}=v_{y}$.
-Thus, according to correct coupling conventional particle tracking and electrodynamics, we identify the direction of X-ray beam propagation with direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation.
-16.6 Bibliography and Notes
-1.
-An angular kick is often an essential part of many XFEL related diagnostic or experimental procedures.
-The standard gain length measurement procedure in XFELs employs such kicks.
-Other applications include ”beam-splitting” schemes where different polarization components are separated by means of an angular kick to the modulated electron beam [61, 62].
-2.
-In a typical configuration for an XFEL, the orbit of a modulated electron beam is controlled to avoid large excursions from the undulator axis.
-All existing XFEL codes are based on a model in which the modulated electron beam moves only along the undulator axis.
-However, random errors in the focusing system can cause angular trajectory errors (or ”kicks”).
-The discrepancy between directions of the electron motion and wavefront normal after the kick have been discussed in the literature.
-One particular consequence that received attention following the [63] is the effect of the trajectory error (single kick error) on the XFEL amplification process.
-It was pointed out that coherent radiation is emitted towards the wavefront normal of the beam modulation.
-Thus, according to conventional coupling of fields and particles (which we claimed incorrect), the discrepancy between the two directions decreases the radiation efficiency [63].
-Analysis of the trajectory errors on the XFEL amplification process showed that any XFEL undulator magnetic field must satisfy stringent requirements.
-However, semi-analytical studies of this critical aspect in the design of a XFEL sources are based on an incorrect coupling of fields and particles.
-The pleasant surprise is that the tolerances predicted are more stringent than they need be according to the corrected XFEL theory.
-This can be considered as one of the reason for the exceptional progress in XFEL developments over last decades.
-3.
-The fact that our theory predicts reality in a satisfactory way is well-illustrated by comparing the prediction we just made with the results of an experiment involving ”X-ray beam splitting” of a circularly-polarized XFEL pulse from the linearly-polarized XFEL background pulse, a technique used in order to maximize the degree of circular polarization.
-The XFEL experiments apparently demonstrated that after a modulated electron beam is kicked on a large angle compared to the divergence of the XFEL radiation, the modulation wavefront is readjusted along the new direction of motion of the kicked beam, Fig.
-73.
-This is the only way to justify coherent radiation emission from the short undulator placed after the kicker and along the kicked direction, see Fig.
-14 in [64].
-These results came unexpectedly, but from a practical standpoint, the ”apparent wavefront readjusting” immediately led to the realization that the unwanted, linearly-polarized radiation background could be fully eliminated without extra-hardware.
-4.
-In existing literature a theoretical analysis of XFELs driven by an electron beam with wavefront tilt was presented in [65, 66], based on the use the usual Maxwell’s equations.
-In fact, the Maxwell solver was used as a part of the standard simulation code.
-We state that this approach is fundamentally incorrect.
-In ultrarelativistic asymptote a modulation wavefront tilt is absurd with the viewpoint of Maxwell’s electrodynamics.
-In the case of an XFEL we deal with an ultrarelativistic electron beam and within the Lorentz lab frame (i.e.
-within the validity of the Maxwell’s equations) the tilted modulation wavefront is at odds with the principle of relativity.
-5.
-In existing XFEL theory the wavefront tilt is considered as real.
+No physical effects may depend on an arbitrary constant or an arbitrary function.
 Let us consider one example.
@@
 To estimate the importance of the last effect, we should calculate the dispersion of the longitudinal velocities due to both effects.
-The deviation of the longitudinal velocity from nominal value is $\Delta v_{z}=v\Delta\gamma/\gamma^{3}-v\Delta\theta^{2}/2$.
+The deviation of the longitudinal velocity from the nominal value is $\Delta v_{z}=v\Delta\gamma/\gamma^{3}-v\Delta\theta^{2}/2$.
 The finite angular spread of the electron beam results in a difference in time when each electron arrives at the same longitudinal position.
-This is so called normal debunching effect.
+This is so-called normal debunching effect.
 From the viewpoint of the existing XFEL theory, the time difference is enhanced by the kick angle $\theta_{k}$.
 In this case, according to conventional (non-covariant) particle tracking, the angle of wavefront tilt is $\theta_{tilt}=\theta_{k}$.
-It is a widespread belief that the wavefront tilt has physical meaning, and that the deviation of the longitudinal velocity component (i.e.
+It is a widespread belief that the wavefront tilt has a physical meaning and that the deviation of the longitudinal velocity component (i.e.
 velocity component which is perpendicular to the modulation wavefront within the framework of Galilean kinematics) is now given by the expression $\Delta v_{z}=-v|\Delta\vec{\theta}+\vec{\theta}_{k}|^{2}/2$.
-If such picture is correct, the crossed term $v\vec{\theta}_{k}\cdot\Delta\vec{\theta}$ leads to a significant degradation of the modulation amplitude.
+If such a picture is correct, the crossed term $v\vec{\theta}_{k}\cdot\Delta\vec{\theta}$ leads to a significant degradation of the modulation amplitude.
 This mechanism is called smearing of modulation and should be distinguished from the normal debunching [75].
-Many experts would like to think that any debunching process obviously has objective meaning.
-The theory of relativity says, however, that normal debunching has objective meaning, but smearing effects not exist at all.
+Many experts would like to think that any debunching process has objective meaning.
+The theory of relativity says, however, that normal debunching has an objective meaning, but smearing effects do not exist at all.
 The explanation of the new debunching mechanism clearly demonstrates the essential dependence of the smearing effect on the choice of the coordinate system in the four-dimensional space, which from the physical point of view is meaningless.
-Now let us understand physically why the new debunching mechanism does not exist in framework of Galilean kinematics.
+Now let us understand physically why the new debunching mechanism does not exist in the framework of Galilean kinematics.
 In this old kinematics the crossed term $v\vec{\theta}_{k}\cdot\Delta\vec{\theta}$ leads to a degradation of modulation amplitude in the forward direction.
 Our Galilean transformed electrodynamics says, however, that by making a measurement on the coherent radiation, one can observe only radiation in the kicked direction.
-But the crossed term is absent in the expression for the deviation of the velocity component along the kicked direction.
+However, the crossed term is absent in the expression for the deviation of the velocity component along the kicked direction.
 It comes out quite naturally that the smearing effect is not a real phenomenon.
-17 Kinematics of the Wigner Rotation
-17.1 Introductory Remarks
-The subject of this chapter is the Wigner rotation (1).
-We have already discussed the Wigner rotation in Chapter 7.
-Our earlier discussion is really about aberration of light and particles.
-The Wigner rotation is also relates the angular rotation velocity of the spin of an elementary particle following a curvilinear orbit.
-The Wigner rotation is a relativistic kinematic effect, which consists in that a coordinate axes of a reference frame, moves along a curvilinear trajectory rotating about the axes of a Lorentz lab frame.
-The point is that two observers with different trajectories have different 3-spaces.
-In other words, in the space measurement of one observer there is mixing of space and time, as seen by the other.
-When we studied aberration of light and particles, we begun by writing down an expression for Wigner rotation angle.
-To deduce this expression is the main purpose of this chapter.
-We have tried to keep the mathematical complexity of the discussion to a minimum.
-17.2 The Commutativity of Collinear Lorentz Boosts
-Lorentz transformations are essential to the further mathematical development of the Wigner rotation theory, so the next two subsections detail the usual applications together with some physical discussion.
-Let us now consider a relativistic particle, accelerating in the lab frame, and let us analyze its evolution within Lorentz coordinate systems.
-The permanent rest frame of the particle is obviously not inertial.
-To get around that difficulty one introduces an infinite sequence of comoving frames.
-At each instant, the rest frame is a Lorentz frame centered on the particle and moving with it.
-As the particle velocity changes to its new value at an infinitesimally later instant, a new Lorentz frame centered on the particle and moving with it at the new velocity is used to observing the particle.
-Let us denote the three inertial frames by $K,R(\tau),R(\tau+d\tau)$.
-The lab frame is $K$, $R(\tau)$ is the rest frame with velocity $\vec{v}=\vec{v}(\tau)$ relative to $K$, and $R(\tau+d\tau)$ is the rest frame at the next instant of proper time $\tau+d\tau$, which moves relative to $R(\tau)$ with infinitesimal velocity $d\vec{v^{\prime}}$.
-All inertial reference frames are assumed to be Lorentz reference frames.
-In order to have this, we impose that $R(\tau)$ is connected to $K$ by the Lorentz boost $L(\vec{v})$, with $\vec{v}$, which transforms a given four vector event $X$ in a space-time into $X_{R}=L(\vec{v})X$.
-The relation $X_{R}=L(d\vec{v^{\prime}})L(\vec{v})X$ presents a step-by-step change from $K$ to $R(\tau)$ and then to $R(\tau+d\tau)$.
-There is another composition of reference-frame transformations which describes the same particle evolution in the Minkowski space-time.
-Let $K(\tau)$ be an inertial frame with velocity $d\vec{v}$ relative to the lab frame $K(\tau+d\tau)$.
-We impose that $K(\tau)$ is connected to $K(\tau+d\tau)$ by the Lorentz boost $L(d\vec{v})$.
-The Lorentz rest frame $R$ is supposed to move relative to the Lorentz frame $K(\tau)$ with velocity $\vec{v}$.
-The relation $X_{R}=L(\vec{v})L(d\vec{v})X$ presents a step-by-step change from $K(\tau+d\tau)$ to $K(\tau)$ and then to the rest frame $R$.
-Let us examine the transformation of the three-velocity in the theory of relativity.
-For a rectilinear motion along the $z$ axis it is performed in accordance with the following equation: $v_{z}(\tau+d\tau)=(dv_{z}+v_{z})/(1+v_{z}dv_{z}/c^{2})$.
-Like it happens with the composition of Galilean boosts, collinear Lorentz boosts commute: $L(dv_{z})L(v_{z})=L(v_{z})L(dv_{z})$.
-This means that the resultant of successive collinear Lorentz boosts is independent of which transformation applies first.
-17.3 The Noncommutativity of Two Lorentz Boosts in Nonparallel Directions
-In contrast with the case of Lorentz boosts in collinear directions, Lorentz boosts in different directions do not commute.
-While the successive application of two Galilean boosts is Galilean boost and the successive application of two rotations is a rotation, the successive application of two non-collinear Lorentz boosts is not a Lorentz boost.
-The composition of non-collinear boosts will results to be equivalent to a boost, followed by spatial rotation, the Wigner rotation.
-Let us compare the succession $K\to R(\tau)\to R(\tau+d\tau)$ with the succession $K(\tau+d\tau)\to K(\tau)\to R$ in the case when the acceleration in the rest frame is perpendicular to the line of flight of the lab frame in the rest frame.
-The frame $R(\tau+d\tau)$ is supposed to move relative to $R(\tau)$ with velocity $d\vec{v^{\prime}}_{x}$.
-Now, since we can write the result in terms of succession $L(\vec{v}_{z})L(d\vec{v}_{x})$ as well as in terms of succession $L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})$, there is a need to clarify a number of questions associated with these compositions of Lorentz frames.
-We can easily understand that the operational interpretation of the succession $L(\vec{v}_{z})L(d\vec{v}_{x})$ is particular simple, involving physical operation used in the measurement of the particle’s velocity increment $d\vec{v}_{x}$ in the lab frame.
-We should be able to understand the operational interpretation of the succession $L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})$.
-We begin by making an important point: the laws of physics in any one inertial reference frame should be able to account for all physical phenomena, including the observations made by moving observers.
-The lab observer sees the time dilation in the Lorentz frame which moves with respect to the lab frame with velocity $\vec{v}_{z}$: $dt/\gamma=d\tau$.
-What velocity increment $d\vec{v}_{R}$ is measured by moving observer?
-As viewed from the lab frame the moving observer measures the increment $d\vec{v}_{R}=\gamma d\vec{v}_{x}$.
-Because of time dilation in the moving frame, the velocity increment in the lab frame $dv_{x}$ corresponds to a velocity $\gamma dv_{x}$ and $-\gamma dv_{x}$ in the frames $R(\tau)$ and $R(\tau+d\tau)$ respectively.
-The resulting boost compositions can be represented as $X_{R}=L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})X=L(\vec{v}_{z})L(d\vec{v}_{x})X$.
-In other words, Lorentz boosts in different direction do not commute: $L(\vec{v}_{z})L(d\vec{v}_{x})\neq L(d\vec{v}_{x})L(\vec{v}_{z})$.
-17.4 Wigner Rotation
-17.4.1 Expression for the Wigner Rotation in the Case of an Arbitrary Velocity
-A large number of incorrect expressions for the Wigner rotation can be found in the literature.
-Therefore, there is a need to consider the Wigner rotation in detail.
-Rather than working out all the transformation matrices for four-vector components, we would like to show a geometrical approach that is very useful for this problem.
-[…]
-We derived the exact relation $d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$ using only rudimentary knowledge of special relativity.
-In textbooks on the theory of relativity, the spatial rotation associated with the composition of two Lorentz boosts in non-parallel directions is often introduced using the algebraic approach.
-This is one of the reason why authors of textbooks obtained an incorrect expression for the Wigner rotation.
-They describe the rotation of a moving object without geometrical meaning of such rotation and encounter serious difficulties in the interpretation of the applied calculations and of the results.
-Figure 74: The interpretation of the Wigner rotation about the proper frame axes.
-Lab frame view of the observation of the proper observer.
-The lab observer is able to account for the observation the velocity rotation angle in the proper frame ($\gamma dv_{x}/v_{z}$) and the observation of the rotation angle made in the lab frame ($dv_{x}/v_{z}$).
-In 3D space the proper frame moves with respect to the lab frame along the line motion.
-Using the line motion as reference line, the lab observer sees that the observer in the proper frame measures the rotating angle of the lab frame axes with respect to proper frame axes $\delta\theta_{w}=\delta\theta(\gamma-1)$.
-17.5 How to Measure a Wigner rotation?
-The Wigner rotation is a relativistic kinematic effect, which consists in that a coordinate axes of a reference frame, moves along a curvalinear trajectory rotating about the axes of a Lorentz lab frame.
-In the Lorentz lab frame, for infinitely small transformations (due to acceleration) we obtained the formula
-$d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}=\left(1-\frac{1}{\gamma}\right)\vec{\delta\theta}~{},$ (120)
-where $d\vec{v}$ is the vector of small velocity change due to acceleration, $\Phi$ is the Wigner rotation angle of the spatial coordinate axes of the proper system relative to Lorentz lab frame, and $\theta$ is the orbital angle.
-But how to measure this orientation?
-A moving coordinate system changes its position in time.
-The question arises whether it is possible to give an experimental interpretation of the rotation of a moving coordinate system.
-We illustrate the problem of how to represent orientation of the moving coordinate system with a simple example.
-The execution of successive transformations from $K(\tau+d\tau)$ to $K(\tau)$ at velocity $d\vec{v_{x}}$ and from from $K(\tau)$ to $R$ at velocity $\vec{v_{z}}$ equivalent to the composition of boost $L(\vec{v}+d\vec{v})$ and rotation.
-The Wigner rotation Eq.(120) is performed additionally to the Lorentz boost at velocity $\vec{v_{z}}+d\vec{v_{x}}$.
-The interpretation of this rotation about the lab frame of reference is closely associated with the length contraction.
-Figure 75: The interpretation of the Wigner rotation about the lab frame.
-A rod directed along $x$-axis in the comoving frame.
-After the first boost the motion takes place along the $z$-axis.
-The lab frame $(x_{r},z_{r})$ coordinate system rotated with respect to the initial lab frame $(x,z)$ coordinate system on the angle $\delta\theta$.
-The projection of the moving rod on the $z_{r}$-axis is simply $L\delta\theta$.
-After the second boost at velocity $dv_{x}$ along the $x$-axis of the lab frame this projection will be contracted down to $L\delta\theta/\gamma$.
-It is assumed that infinitesimal angle $\delta\theta=dv_{x}/v_{z}$.
-According to the contracted projection the comoving frame is rotated with respect to the initial lab frame axes by the Wigner angle equal to $\delta\theta(1-1/\gamma)$.
-Suppose that the lab observer after the Lorentz boost at velocity $v_{z}$ rotates the coordinate system on the angle $\vec{\delta\theta}=\vec{v_{z}}\times d\vec{v_{x}}/v_{z}^{2}$ and now $x_{r}$ locates orthogonally to the vector $\vec{v_{z}}+d\vec{v_{x}}$.
-Similarly, the axis $z_{r}$ is parallel to the vector $\vec{v_{z}}+d\vec{v_{x}}$.
-Consider a rod directed along $x$-axis in the comoving frame.
-The motion takes place in the plane $(x,z)$ and the rod located perpendicularly to the velocity $\vec{v_{z}}$.
-After the rotation of the lab frame axes, the projection of the rod on the $z_{r}$ axes will be simply $l\delta\theta$, where $l$ is the rod length in the $R$ frame and also in the lab frame after the first boost along the $z$-axis.
-After the second Lorentz boost at velocity $d\vec{v_{x}}$ this projection will be contracted down to $L\delta\theta/\gamma$ (Fig.
-75).
-Let the observer in the lab frame fix the position of the axes of the comoving frame.
-In ultrarelativistic limit $\gamma\to\infty$ these axes will be parallel to the rotated lab frame axes $(x_{r},z_{r})$.
-In fact, projection of the rod on the $z_{r}$ axis will be zero.
-In the case of an arbitrary velocity, axes of the comoving frame are not parallel to the rotated lab frame axes $(x_{r},z_{r})$.
-According to contracted projection, the angle will be $-\vec{\delta\theta}/\gamma$.
-And one can verify directly that the axes of the comoving frame are actually rotated with respect to the initial lab frame axes $(x,z)$ by the angle equal to $\delta\theta-\delta\theta/\gamma$, which is just the Wigner rotation angle in accordance with equation Eq.(120).
-Here we only wished to show how naturally Lorentz transformations lead to the Wigner rotation phenomenon.
-We have come to the conclusion that what are usually considered advanced parts of the theory of relativity are, in fact, quite simple.
-Indeed, we demonstrated that the Wigner rotation results directly from the length contraction.
-17.6 Wigner Rotation and the Reciprocity
-There are several points to be made about the Wigner rotation theory.
-We wish to remark that the expression for the Wigner rotation angle in the proper frame (as viewed from the Lorentz lab frame), $d\vec{\Phi}_{R}=(\gamma-1)\vec{v}\times d\vec{v}/v^{2}$, can be presented in the form $d\vec{\Phi}_{R}=(1-1/\gamma)\vec{v}_{R}\times d\vec{v}_{R}/v_{R}^{2}$.
-Thus, the Wigner rotation angle in the proper frame is expressed in terms of the lab frame velocity $\vec{v}_{R}=\vec{v}$ and its increment $d\vec{v}_{R}=\gamma d\vec{v}$ in the proper frame.
-We have just demonstrated that the reciprocity is the correct concept in the case of infinitesimal Wigner rotation, as it must be.
-17.7 Bibliography and Notes
-1.
-As known, a composition of noncollinear Lorentz boosts does not results in a different boost but in a Lorentz transformation involving a boost and a spatial rotation, the latter being known as Wigner rotation [67, 68].
-Wigner rotation is sometimes called Thomas rotation (see e.g.
-[13, 69]).
-2.
-The correct expression for the Thomas (Wigner) rotation was first obtained by V.
-Ritus [70].
-In deriving expressions for the Thomas (Wigner) rotation, the majority of authors (see e.g.
-[69]) were supposedly guided by the incorrect expression for Thomas (Wigner) rotation from Moeller’s monograph [13].
-The expression obtained by Moeller is given by $\vec{\delta\Phi}=(1-\gamma)\vec{v}\times d\vec{v}/v^{2}=(1-\gamma)\vec{\delta\theta}$ (and subsequently expression for Thomas precession $\Omega_{\mathrm{T}}=(1-\gamma)\omega_{0}$).
-It should be note that, in his monograph, Moeller stated several times that this expression valid in the lab Lorentz frame.
-Clearly, this expression and correct result Eq.
-(120) differ both in sign and in magnitude.
-An analysis of the reason why Moeller obtained an incorrect expression for the Thomas (Wigner) rotation in the lab frame is the focus of Ritus paper [71].
-As shown in [71], the Moeller’s mistake is not computational, but conceptual in nature.
-In review [72] it is shown that the correct result was obtained in the works of several authors, which were published more than half century ago but remained unnoticed against the background of numerous incorrect works.
-The authors of some papers believe that the incorrect result for Wigner rotation in the lab frame presented in textbooks $d\vec{\Phi}=-(\gamma-1)\vec{v}\times d\vec{v}/v^{2}$ is only incorrectly interpreted with the understanding that it should be reinterpreted as a Wigner rotation of the lab frame in the proper frame.
-We note that such reinterpreted expression for Wigner rotation in the proper frame $d\vec{\Phi}\to d\vec{\Phi}_{R}$ is also incorrect in sign.
-3.
-We note that in 1986, M.
-Stranberg obtained an expression for the Thomas (Wigner) rotation correct both in the lab inertial frame and the comoving reference frame [73].
-It is noteworthy that [73] is one of the few papers that explicitly states that the angle of the Thomas (Wigner) rotation in the comoving reference frame is $\gamma$ times higher than in the lab frame.
-18 Relativistic Spin Dynamics
-18.1 Introductory Remarks
-In 1959, a paper by Bargmann, Michel, and Telegdi was published, which dealt with the motion of elementary charged spinning particles with an anomalous magnetic moment in electromagnetic field (1).
-The extremely precise measurements of the magnetic-moment anomaly of the electron made on highly relativistic electrons are based on the BMT equation.
-The anomalous magnetic moment can be calculated by use of quantum electrodynamics.
-The theoretical result agrees with experiments to within a very high accuracy.
-This can be regarded as a direct test of BMT equation.
-The existing textbooks suggest that the experimental test of the BMT equation is a direct test of what we consider the incorrect expression for Wigner rotation in the Lorentz lab frame.
-We claim that the inclusion of this incorrect expression as an integral part of the BMT equation in most texts is based on an incorrect physical argument.
-In this chapter we will investigate in detail the reason why this is the case.
-18.2 Magnetic Dipole at Rest in an Electromagnetic Field
-Let us consider at first the spin precession for a non relativistic charge particle.
-The proportionality of magnetic moment $\vec{\mu}$ and angular momentum $\vec{s}$ has been confirmed in many ”gyromagnetic” experiments on many different systems.
-The constant of proportionality is one of the parameters charactering a particular system.
-It is normally specified by giving the gyromagnetic ratio or $g$ factor, defined by $\vec{\mu}=ge\vec{s}/(2mc)$.
-This formula says that the magnetic moment is parallel to the angular momentum and can have any magnitude.
-For an electron $g$ is very nearly 2.
-Suppose that a particle is at rest in an external magnetic field $\vec{B}_{R}$.
-The equation of motion for the angular momentum in its rest frame is $d\vec{s}/d\tau=\vec{\mu}\times\vec{B}_{R}=eg\vec{s}\times\vec{B}_{R}/(2mc)=\vec{\omega}_{s}\times\vec{s}$.
-In other words, the spin precesses around the direction of magnetic field with the frequency $\omega_{s}=-eg\vec{B}_{R}/(2mc)$.
-In the same non relativistic limit the velocity processes around the direction of $\vec{B}_{R}$ with the frequency $\omega_{p}=-(e/mc)\vec{B}_{R}$: $d\vec{v}/d\tau=(e/mc)\vec{v}\times\vec{B}_{R}$.
-Thus, for $g=2$ spin and velocity precess with the same frequency, so that the angle between them is conserved.
-18.3 Derivation of the Covariant (BMT) Equation of Motion of Spin
-Spin dynamics equations can be expressed as tensor equations in Minkowski space-time.
-We shall limit ourselves to the case of a particle with a magnetic moment $\vec{\mu}$ in a microscopically homogeneous electromagnetic field.
-Evidently the torque affects only the spin and the force affects only the momentum.
-It follows that the motion of the system as a whole in any frame is determined entirely by its charge, independent of magnetic dipole moment.
-This part of the motion has been treated in the Chapter 12.
-We need now only consider the spin motion.
-In seeking the equation for the spin motion, we shall be guided by the known dynamics in the rest frame and the known relativistic transformation laws.
-We emphasize that spin is defined in a particular frame (the rest frame).
-Therefore, to form expressions with known transformation behavior, we need to introduce a four-quantity related to the spin.
-A convenient choice is a four- (pseudo)-vector $S$ defined by the requirement that in the rest frame its space-like components are the spin components, while the time-like component is zero.
-We shall call $S$ four-spin; when normalized by dividing by its invariant length, it will be called polarization four-vector.
-It is space-like, and therefore in no frame does it space-like part vanish.
-Let the spin of the particle be represented in the rest frame by $\vec{s}$.
-The four-vector $S^{\alpha}$ is by definition required to be purely spatial at time $\tau$ in an instantaneous Lorentz rest frame $R(\tau)$ of the particle and to coincide at this time with the spin $\vec{s}(\tau)$ of the particle; that is $S_{R}^{\alpha}(\tau)=(0,\vec{S}_{R}(\tau))=(0,\vec{s}(\tau))$.
-At a later instant $\tau+\Delta\tau$ in an instantaneous inertial rest frame $R(\tau+\Delta\tau)$, we have similarly $S_{R}^{\alpha}(\tau+\Delta\tau)=(0,\vec{S}_{R}(\tau+\Delta\tau))=(0,\vec{s}(\tau+\Delta\tau))$.
-The BMT equation is manifestly covariant equation of motion for a four-vector spin $S^{\alpha}$ in an electromagnetic field $F^{\alpha\beta}$:
-$\frac{dS^{\alpha}}{d\tau}=\frac{ge}{2mc}\left[F^{\alpha\beta}S_{\beta}+\frac{1}{c^{2}}u^{\alpha}\left(S_{\lambda}F^{\lambda\mu}u_{\mu}\right)\right]-\frac{1}{c^{2}}u^{\alpha}\left(S_{\lambda}\frac{du^{\lambda}}{d\tau}\right)~{},$ (121)
-where $u_{\mu}=dx_{\mu}/d\tau$ is the four-dimensional particle velocity vector. With Eq.(20), one has (1)
-$\frac{dS^{\alpha}}{d\tau}=\frac{e}{mc}\left[\frac{g}{2}F^{\alpha\beta}S_{\beta}+\frac{g-2}{2c^{2}}u^{\alpha}\left(S_{\lambda}F^{\lambda\mu}u_{\mu}\right)\right]~{},$ (122)
-The BMT equation is valid for any given inertial frame, and consistently describes, together with the covariant-force law, the motion of a charged particle with spin and magnetic moment.
-If $F^{\mu\nu}\neq 0$, even with $g=0$, we see that $dS^{\mu}/d\tau\neq 0$.
-Thus, a spinning charged particle will precess in an electromagnetic field even if it has no magnetic moment.
-This precession is pure relativistic effect.
-The covariant equation of spin motion for a relativistic particle under the action of the four-force $Q^{\mu}=eF^{\mu\nu}u_{\nu}$ in the Lorentz lab frame, Eq.(121), is a relativistic ”generalization” of the equation of motion for a particle angular momentum in its rest frame.
-Relativistic ”generalization” means that the three independent equations expressing the Larmor spin precession are be embedded into the four-dimensional Minkowski space-time.
-The idea of embedding is based on the principle of relativity i.e.
-on the fact that the classical equatuion of motion for particle angular momentum $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$ can always be used in any Lorentz frame where the particle, whose motion we want to describe, is at rest.
-In other words, if an instantaneously comoving Lorentz frame is given at some instant, one can precisely predict the evolution of the particle spin in this frame during an infinitesimal time interval.
-In Lorentz coordinates there is a kinematics constraint $S^{\mu}u_{\mu}=0$, which is orthogonality condition of four-spin and four-velocity.
-Because of this constraint, the four-dimensional dynamics law, Eq.(121), actually includes only three independent equations of motion.
-Using the explicit expression for Lorentz force we find that the four equations Eq.(121) automatically imply the constraint $S^{\mu}u_{\mu}=0$ as it must be.
-To prove this we may point out that one has in every Lorentz frame $S^{0}=\vec{S}\cdot\vec{v}$.
-While $S^{0}$ vanishes in the rest frame, $dS^{0}/d\tau$ need not.
-In fact $d(S^{\mu}u_{\mu})/d\tau=0$ implies $dS^{0}/d\tau=\vec{S}\cdot d\vec{v}/d\tau$.
-The immediate generalization of $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$ and $dS^{0}/d\tau=\vec{S}\cdot d\vec{v}/d\tau$ to arbitrary Lorentz frames is Eq.(121) as can be checked by reducing to the rest frame.
-A methodological analogy with the relativistic generalization of the Newton’s second law emerges.
-In order to fully understand the meaning of embedding of the spin dynamics law in the Minkowski space-time, one must keep in mind that, above, we characterized the spin dynamics equation in the Lorentz comoving frame as a phenomenological law.
-The microscopic interpretation of the magnetic moment of a particle is not given.
-In other words, it is generally accepted that the spin dynamics law is a phenomenological law and the magnetic moment is introduced in an ad hoc manner.
-The system of coordinates in which the classical equations of motion for particle angular momentum are valid can be defined as Lorentz rest frame.
-The relativistic generalization of the three-dimensional equation $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$ to any Lorentz frame permits us to make correct predictions.
-18.4 Change Spin Variables
-When Bargman, Michel and Telegdi first discovered the correct laws of spin dynamics, they wrote a manifestly covariant equation in Minkowski space-time, Eq.(122), which describes the motion of the four spin $S^{\mu}$.
-The derivation of this equation was very similar to the four-tensor equations that were already known to relativistic particle dynamics.
-The equation Eq.(122) is more complex than one might think.
-In fact, it is composed by a set of coupled differential equations.
-To find solution directly from the system seems quite difficult, even for a very symmetric, uniform magnetic field setup.
-How to solve this four-tensor equation is an interesting question.
-In relativistic spin dynamics it is done in one particular way, which is very convenient.
-In order to apply Eq.(122) to specific problems it is convenient to introduce a three vector $\vec{s}~{}$ by the equation $\vec{s}=\vec{S}+S^{0}\vec{p}c/(\mathcal{E}+mc^{2})$.
-With the help of this relation one can work out the equation of motion for $\vec{s}~{}$.
-In the important case of a uniform magnetic field with no electric field in the lab frame one has, after a somewhat lengthly calculations:
-$\frac{d\vec{s}}{d\tau}=-\frac{e}{2m}\left[\left(g-2+\frac{2}{\gamma}\right)\gamma\vec{B}-\frac{(g-2)\gamma}{\gamma+1}\frac{\vec{v}}{c}\left(\frac{\vec{v}}{c}\cdot\gamma\vec{B}\right)\right]\times\vec{s}~{},$ (123)
-What must be recognized is that in the accepted covariant approach (indeed, Eq.(122) is obviously manifestly covariant), the solution of the dynamics problem for the spin in the lab frame makes no reference to the three-dimensional velocity.
-In fact, the Eq.(123) includes relativistic factor $\gamma$ and vector $\vec{v}/c$, which are actually notations: $\gamma=\mathcal{E}/(mc^{2})$, $\vec{v}/c=\vec{p}c/\mathcal{E}$.
-All quantities $\mathcal{E},\vec{p},\vec{B}$ are measured in the lab frame and have exact objective meaning i.e.
-they are convention-independent.
-The evolution parameter $\tau$ is also measured in the lab frame and has exact objective meaning .
-For instance, it is not hard to demonstrate that $d\tau=mdl/|\vec{p}|$, where $dl$ is the differential of the path length.
-Spin vector $\vec{s}~{}$ is not part of a four-vector, and depends on both $\vec{S}$ and $S^{0}$.
-While not being a four-vector, it is effectively a three-dimensional object (having zero time component in the inertial frame in question) and the spatial part of this object undergoes pure rotation with constant rate for the example of motion along a circle in special relativity.
-If we perform an arbitrary velocity mapping, $\vec{s}$ will have to be recomputed from the transformed values $S^{\mu}$ and $p^{\mu}$.
-However, this new $\vec{s}$ will satisfy an equation of the form Eq.(123), with $\vec{B}$ computed from the transformed $F^{\mu\nu}$.
-Let us restrict our treatment of spinning particle dynamics to purely transverse magnetic fields.
-This means that the magnetic field vector $\vec{B}$ is oriented normal to the particle line motion.
-If the field is transverse, then equation Eq.(123) is reduced to
-$\frac{d\vec{s}}{d\tau}=\vec{\Omega}\times\vec{s}=-\frac{e}{2mc}\left[\left(g-2+\frac{2}{\gamma}\right)\gamma\vec{B}\right]\times\vec{s}~{},$ (124)
-Now we have an equation in the most convenient form to be solved.
-Suppose we let the charged spinning particle in the lab frame through a bending magnet with the length $dl$.
-We know that $d\theta=-eBdl/(|\vec{p}|c)$ is the orbital angle of the particle in the lab frame.
-Note that $d\tau=mdl/|\vec{p}|$.
-Then, Eq.(124) tells us that we may write the spin rotation angle with respect to the lab frame axes $\Omega d\tau$ as $\Omega d\tau=[(g/2-1)\gamma d\theta+d\theta]$.
-This tell us that in the lab frame the spin of a particle $\vec{s}$ changes the angle $\phi$ with its line motion.
-The rate of change of the angle $\phi$ with the orbital angle is $(g/2-1)\gamma$, so we can write $d\phi=(g/2-1)\gamma d\theta$.
-We would like to discuss the following question: Is the vector $\vec{s}$ merely a device which is useful in making calculations - or is it a real quantity ( i.e.
-a quantity which has direct physical meaning)?
-The starting point of Bargman, Michel and Telegdi was the particle rest frame and the equation of motion for particle angular momentum, which they generalized to the Lorentz lab frame and then transformed back to the rest frame, $\vec{s}=\vec{S}+S^{0}\vec{p}c/(\mathcal{E}+mc^{2})$.
-The spin vector $\vec{s}$ directly gives the spin as perceived in a comoving system.
-If we say that in the lab frame the spin of a particle makes the angle $\phi$ with its velocity, we mean that in the particle’s rest frame the spin makes this angle with the line motion of the lab frame.
-This brings up an interesting question: Why it is convenient to transform equation Eq.(122) to the rest frame as of that instant?
-This is understandable.
-The approach in which we deal with the proper spin is much preferred in the experimental practice due to clear physical meaning of the spin vector $\vec{s}$.
-Unlike (3D) momentum vector $\vec{p}$, which has definite components in each reference frame, (3D) angular momentum vector $\vec{s}$ is defined only in one particular reference frame.
-It does not transform.
-Any statement about it refers to the rest frame as of that instant.
-We must conclude that the conventional method used to explain the spin dynamics in the lab frame is very unusual.
-Indeed, the spin rotation measurement in the lab frame is interpreted with viewpoint of the proper observer as viewing this of the lab observer.
-18.5 An Alternative Approach to the BMT Theory
-18.5.1 BMT Equation Transformed to the Rest Frame
-We want to emphasize that the equation Eq.(124) for the proper spin $\vec{s}$ and the BMT equation Eq.(121) for the four spin $S^{\mu}$ are completely equivalent, they both determine the behaviour of the spin from the point of view of the lab frame.
-With Eq.(124) have what we need to know - the evolution of the proper spin vector $\vec{s}$ with respect to the lab frame axes.
-Starting from the classical equation $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$, which describes the Larmor precession with respect to the proper frame axes, we have derived the equation Eq.(123), which describes the spin motion with respect to the lab frame axes in the proper frame and reduced to Eq.(124) in the case of purely transverse magnetic fields.
-That means that we know the orientation of the proper spin with respect to the lab coordinate system which is moving with velocity $-\vec{v}$ and acceleration $-\gamma d\vec{v}/d\tau$ in the proper frame.
-Above we described the BMT equation, Eq.(124), in the standard manner.
-It uses a spin quantity defined in the proper frame but observed with respect to the lab frame axes.
-Let’s look at what the equation Eq.(124) says in a little more detail.
-It will be more convenient if we rewrite this equation as
-$d\vec{s}=\vec{\Omega}d\tau\times\vec{s}=-eg\gamma\vec{B}d\tau/(2mc)\times\vec{s}+e(\gamma-1)\vec{B}d\tau/(mc)\times\vec{s}~{}.$ (125)
-18.5.2 Relativistic Kinematic Addition to the Larmor Rotation
-Now let’s see how we can write the right-hand side of Eq.(125) .
-The first term is that we would expect for the spin rotation due to a torque with respect to the proper frame axes $d\vec{\phi}_{L}=-eg\gamma\vec{B}d\tau/(2mc)=(g/2)\gamma d\vec{\theta}$.
-Here $d\vec{\theta}=-eBdl/(|\vec{p}|c)$ is the angle of the velocity rotation in the lab frame.
-It has also been made evident by our analysis in the previous Chapter 16 that angle of rotation $d\vec{\phi}_{W}=-e(\gamma-1)\vec{B}d\tau/(mc)=(\gamma-1)d\vec{\theta}$ corresponds to the Wigner rotation of the lab frame axes with respect to the proper frame axes.
-With this definitions, we have
-$d\vec{s}=\vec{\Omega}d\tau\times\vec{s}=d\vec{\phi}_{L}\times\vec{s}-d\vec{\phi}_{W}\times\vec{s}~{},$ (126)
-which begins to look interesting.
-18.5.3 Wigner Rotation in the Proper Frame. First Practical Application
-Now we introduce our new approach to the BMT theory, finding another way in which our complicated problem can be solved.
-We know that $d\vec{\phi}_{L}$ and $d\vec{\phi}_{W}$ are the rotations with respect to the proper frame axes.
-Actually we only need to find the spin motion with respect to the lab frame axes.
-Now we must be careful about signs of rotations.
-There is a good mnemonic rule to learn the signs of different rotations.
-The rule says, first, that the direction of the velocity rotation in the proper frame is the same as the direction of the velocity rotation in the lab frame.
-Second, the direction of the lab frame rotation in the proper frame is the same as the direction of the velocity rotation in the proper frame.
-Third, the sign of the spin rotation due to a torque at $g>0$ (it is handy to remember that for an electron $g$ is positive and very nearly 2) means that the direction of the rotation in the proper frame is the same as the direction of the velocity rotation in the proper frame.
-We now ask about the proper spin rotation with respect to the lab frame axes.
-This is easy to find.
-The relative rotation angle is $d\vec{\phi}_{L}-d\vec{\phi}_{W}$.
-So we begin to understand the basic machinery behind spin dynamics.
-We see why the Wigner rotation of the lab frame axes in the proper frame must be taken into account if we need to know the proper spin dynamics with respect to the lab frame axes.
-Why the new derivation of the BMT equation is so simple?
-The reason is that the splitting of the particle spin motion with respect to the lab frame axes into the dynamic and kinematic parts can only be realized in the proper frame.
-In the proper frame, we do not need to know any more about a relativistic ”generalization” of the (phenomenological) classical equation of motion for the particle angular momentum.
-In this case, it is possible to separate the spin dynamics problem into the trivial (Larmor) dynamic problem and into the kinematic problem of Wigner rotation of the lab frame axes in the proper frame.
-18.6 Spin Tracking
-Having written down the spin motion equation in a 4-vector form, Eq.(122), and determined the components of the 4-force, we satisfied the principle of relativity for one thing, and, for another, we obtained the four components of the equation for the spin motion.
-This is a covariant relativistic generalization of the usual three dimensional equation of magnetic moment motion, which is based on the particle proper time as the evolution parameter.
-We next wish to describe the spin motion with respect to the Lorentz lab frame using the lab time $t$ as the evolution parameter.
-18.6.1 Conventional spin tracking. Hidden absolute time coordinatization
-When going from the proper time $\tau$ to the lab time $t$, the frequency of spin precession with respect to the lab frame can be obtained using the well-known formula $d\tau=dt/\gamma$. We then find
-$\frac{d\vec{s}}{dt}=\vec{\varpi}\times\vec{s}=-\frac{e}{2mc}\left[\left(g-2+\frac{2}{\gamma}\right)\vec{B}\right]\times\vec{s}~{}.$ (127)
-The frequency of spin precession can be written in the form
-$\varpi=\omega_{0}[1+\gamma(g/2-1)]~{},$ (128)
-where $\omega_{0}$ is the particle revolution frequency.
-Now the time-like part of the four-velocity is decomposed to $c\gamma=c/\sqrt{1-v^{2}/c^{2}}$ and the trajectory does not include relativistic kinematics effects.
-In particular, the Galilean vectorial law of addition of velocities is actually used.
-So we must have made a jump to the absolute time coordinatization.
-The previous commonly accepted derivation of the equations for the spin precession in the lab frame from the covariant equation Eq.(121) has the same delicate point as the derivation of the equation of particle motion from the covariant equation Eq.(20).
-The four-velocity cannot be decomposed into $u=(c\gamma,\vec{v}\gamma)$ when we deal with a particle accelerating along a curved trajectory in the Lorentz lab frame.
-One of the consequences of non-commutativity of non-collinear Lorentz boosts is the unusual momentum-velocity relation.
-In this case there is a difference between covariant and non-covariant particle trajectories.
-The old kinematics comes from the relation $d\tau=dt/\gamma$.
-The presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization that provides the time coordinate $t$ in the lab frame is based on the use of the absolute time convention.
-18.6.2 Convention-Invariant Spin Tracking
-In the Chapter 12 we saw that the particle path $\vec{x}(l)$ has an exact objective meaning i.e.
-it is convention-invariant.
-The spin orientation $\vec{s}$ at each point of the particle path $\vec{x}(l)$ has also exact objective meaning.
-In contrast to this, and consistently with the conventionality of the three-velocity, the function $\vec{s}(t)$ describing the spinning particle in the lab frame has no exact objective meaning.
-We now want to describe how to determine the spin orientation along the path $\vec{s}(l)$ in covariant spin tracking.
-Using the covariant equation Eq.(121) we obtain Eq.(124).
-If we use the relation $d\tau=mdl/|\vec{p}|$ our convention-invariant equation of spin motion reads
-$\frac{d\vec{s}}{dl}=-\frac{e\mathcal{E}}{m|\vec{p}|c^{3}}\left[\left(\frac{g}{2}-1+\frac{mc^{2}}{\mathcal{E}}\right)\vec{B}\right]\times\vec{s}=\left[\left(\frac{g}{2}-1\right)\frac{\mathcal{E}}{mc^{2}}+1\right]\frac{d\vec{\theta}}{dl}\times\vec{s}~{},$ (129)
-which is based on the path length $l$ as the evolution parameter.
-These three equations corresponds exactly to the equations for components of the proper spin vector that can be found from the non-covariant spin tracking equation Eq.(127).
-We want to emphasize that there are two different (covariant and non covariant) approaches that produce the same spin orientation $\vec{s}(l)$ along the path.
-The point is that both approaches describe correctly the same physical reality and the orientation of the proper spin $\vec{s}$ at any point of particle path in the magnetic field has obviously an objective meaning, i.e.
-is convention-invariant.
-Now we take an example, so it can be seen that we do not need to ask questions about the function $\vec{s}(t)$ of a spinning particle experimentally.
-Just think of experiments related with accelerator physics.
-Suppose we want to calibrate the beam energy in a storage ring based on measurement of spin precession frequency of polarized electrons.
-To measure the precession frequency $\varpi$, a method of beam resonance depolarization by an oscillating electromagnetic field can be used (2).
-There are many forms of depolarizers, but we will mention just one, which especially simple.
-It is a depolarizer whose operation depends on the radio-frequency longitudinal magnetic field which is produced by a current-curring loop around a ceramic section of the vacuum chamber.
-Suppose the observer in the lab frame performs the beam energy measurement.
-We should examine what parts of the measured data depends on the choice of synchronization convention and what parts do not.
-Clearly, physically meaningful results must be convention-invariant.
-One might think that this is a typical time-depending measurement of function $\vec{s}(t)$.
-However, we state that the precession frequency $\varpi$ has no intrinsic meaning - its meaning is only being assigned by a convention.
-It is not possible to determine the precession frequency $\varpi$ uniquely, because there is always some arbitrariness in the $\vec{s}(t)$.
-For instance, it is always possible to make an arbitrary change in the rhythm of the clocks (i.e.
-scale of the time).
-But our problem is to determine the energy for an electron beam.
-So one needs to measure also the revolution frequency $\omega_{0}$ by using the same space-time grid.
-What this all means physically is very interesting.
-The ratio $\varpi/\omega_{0}$ is convention independent i.e it does not depend on the distant clocks synchronization or on the rhythm of the clocks.
-It means, for example, that if we observe the dimensionless frequency $\varpi/\omega_{0}$, we can find out the value of the convention-invariant beam energy $\mathcal{E}$.
-The $(g/2-1)$ factor can be calculated by use of quantum electrodynamics.
-[…]
-18.7 Spin Rotation in the Limit $g\to 0$
-It is generally accepted that spin dynamics law is a phenomenological law and that the magnetic moment is introduced in an ad hoc manner.
-Let us consider the special case with $g\to 0$.
-The BMT equation for a particle with small $g$ factor is
-$\frac{dS^{\alpha}}{d\tau}=-\frac{1}{c^{2}}u^{\alpha}\left(S_{\lambda}\frac{du^{\lambda}}{d\tau}\right)=-\frac{e}{mc^{3}}u^{\alpha}\left(S_{\lambda}F^{\lambda\mu}u_{\mu}\right)~{}.$ (130)
-It is often more convenient to write this equation as the equation of motion for $\vec{s}$. If the field is transverse, then the equation Eq.(130) is reduced to
-$\frac{d\vec{s}}{d\tau}=\left[\left(\frac{\mathcal{E}}{mc^{2}}-1\right)\frac{e}{mc}\vec{B}\right]\times\vec{s}~{},$ (131)
-Note that the equation Eq.(131) for the proper spin $\vec{s}$ and the BMT equation Eq.(130) for four spin $S^{\mu}$ are completely equivalent.
-Eq.(131) is the result of transformation to new spin variables.
-It will be more convenient if we rewrite this equation as
-$d\vec{s}=-(\gamma-1)d\vec{\theta}\times\vec{s}~{}.$ (132)
-Here $d\vec{\theta}$ is the angle of the velocity rotation in the lab frame, $(\gamma-1)d\vec{\theta}$ is the Wigner rotation of the lab frame axes with respect to the proper frame axes.
-BMT equation uses a spin quantity defined in the proper frame but observed with respect to the lab frame axes.
-Conventional spin tracking treats the space-time continuum in a non relativistic format, as a (3+1) manifold.
-In the conventional spin tracking, we assign absolute time coordinate and we have no mixture of positions and time.
-This approach to relativistic spin dynamics relies on the use of three equations for the spin motion
-$\frac{d\vec{s}}{dt}=\left[\left(1-\frac{1}{\gamma}\right)\frac{e}{mc}\vec{B}\right]\times\vec{s}=-\left[\left(\gamma-1\right)\vec{\omega_{0}}\right]\times\vec{s}~{},$ (133)
-which are based on the use of the absolute time $t$ as the evolution parameter.
-Here, $\vec{\omega_{0}}=-e\vec{B}/(mc\gamma)$ is the particle angular frequency in the lab frame.
-Now the time-like part of the four-velocity is decomposed to $c\gamma=c/\sqrt{1-v^{2}/c^{2}}$.
-This decomposition is a manifestation of the absolute time convention.
-It should be note that trick with $dt$ in Eq.
-(133) is only technical; this equation includes the differential $dt$ in the left and right parts.
-We can eliminate $dt$ from Eq.(133).
-If we make this elimination we find Eq.(132).
-There are two different (covariant and non covariant) approaches that produce the same spin orientation $\vec{s}(l)$ along the path. Using the Eq.(131) or Eq.(133) we obtain
-$\frac{d\vec{s}}{dl}=-\left[\frac{\mathcal{E}}{mc^{2}}-1\right]\frac{d\vec{\theta}}{dl}\times\vec{s}~{},$ (134)
-Both approaches describe correctly the same physical reality, and the rotation of the proper spin $\vec{s}$ with respect to the lab frame axes at $g\to 0$ is convention-invariant.
-The relativistic kinematic effects such as Wigner rotation, Lorentz-Fitzgerald contraction, time dilation and relativistic corrections to the law of composition of velocities are coordinate (i.e.
-convention-dependent) effects and have no exact objective meaning.
-In the case of the Lorentz coordinatization, one will experience e.g.
-the Wigner rotation phenomenon.
-In contrast to this, in the case of absolute time coordinatization there are no relativistic kinematics effects, and no Wigner rotation will be found.
-However, the spin orientation at each point of the particle path has exact objective meaning.
-In fact, Eq.(134) is convention-invariant i.e includes only quantities which have exact objective meaning.
-Understanding this result of the theory of relativity is similar to understanding the previously discussed result for relativistic mass correction.
-We find that the evolution of a particle along its path is still given by the corrected Newton’s second law even though the relativistic mass correction has no kinematical origin.
-A methodological analogy with the spin dynamics equation Eq.(134) emerge by itself.
-The spin rotation in the lab frame at $g\to 0$ is relativistic effect (as the relativistic mass correction) but it has no kinematical origin.
-18.7.1 Proper frame view of observations of the lab observer
-Now we wish to continue in our analysis a little further.
-We will look for a different way of calculating the spin rotation.
-We found earlier that the easiest way to derive the BMT equation is to use the Lorentz proper frame.
-The Wigner rotation of the lab frame axes in the proper frame must be taken into account if we need to know the proper spin dynamics with respect to the lab frame axes.
-In contrast, in the case of the absolute time coordinatization in the proper frame, there is no kinematic Wigner rotation of the lab frame axes with respect to the proper frame axes.
-The two approaches give, in fact, a different result for spin rotation with respect to the lab frame axes, which must be, however, convention-invariant.
-This glaring conflict between results of covariant and non covariant approaches in the proper frame can actually explained.
-We will see that it is a dynamical line of arguments that explains this paradoxical situation with the relativistic spin rotation.
-Suppose that an observer in the lab frame performs a spin rotation measurement.
-To measure the spin direction, a polarimeter at rest in the lab frame can be used.
-Suppose we put a charged spinning particle with small $g$ factor through a bending magnet.
-The lab observer can directly measure the angle of spin rotation at the magnet exit using the polarimeter.
-The result he observes is consistent with the spin dynamics equation Eq.(134).
-How does it happen that the construction of the polarimeter never came into discussion before?
-In the lab frame the polarimeter is at rest and the field theory involved in the polarimeter operation description is isotropic.
-We do not need to know any more about the polarimeter operation.
-In this sense we can discuss in the lab frame about the spin orientation with respect to the lab frame axes and any detail about the polarimeter is not needed.
-The proper observer sees that polarimeter is moving with a given acceleration and the lab observer, moving with the polarimeter, performs the spin direction measurement.
-Then, when the polarimeter measurement is analyzed, the proper observer finds that the measured spin rotation angle is nonzero and consistent with the spin dynamics equation Eq.(134), as must be.
-How shell we describe the polarimeter operation after Galilean transformations?
-Suppose that the operation of the polarimeter depends on the electromagnetic field.
-After the Galilean transformations of the field equations we obtain the complicated anisotropic electrodynamics equations.
-The new terms that have to be put into the field equations due to the use of Galilean transformations lead to the same prediction as concerns experimental results: the spin of the particle is rotated with respect to the lab frame axes according to Eq.(134).
-In order to predict the result of the moving polarimeter measurement one does not need to have access to the detailed dynamics of the particle into the polarimeter.
-It is enough to assume Lorentz covariance of the field theory involved in the description of the polarimeter operation.
-As before, we use a mathematical trick for solving the electromagnetic field equations with anisotropic terms: in order to eliminate these terms in the transformed field equations, we make a change of the variables.
-Using new variables we obtain the phenomenon of spin rotation with respect to the lab frame axes.
-18.8 Incorrect Expression for Wigner Rotation. Myth About Experimental Test
-18.8.1 Terminology. Thomas Precession: Correct and Incorrect Solutions
-The expression for the Wigner rotation in the lab frame obtained by authors of textbooks is given by $\vec{\delta\Phi}=(1-\gamma)\vec{v}\times d\vec{v}/v^{2}=(1-\gamma)\vec{\delta\theta}$, which often presented as $\vec{\omega}_{Th}=d\vec{\Phi}/dt=(1-\gamma)[\vec{v}\times d\vec{v}/dt]/v^{2}$.
-In other words, the proper frame coordinate performs a precession relative to the lab frame with the velocity of precession $\vec{\omega}_{Th}$, where $d\vec{v}/dt$ is the acceleration of the spinning particle in the lab frame.
-This precession phenomenon is called Thomas precession.
-From the viewpoint of the generally accepted terminology, Thomas precession is actually a manifestation of the Wigner (Thomas) rotation.
-According to expression for Thomas precession in the lab frame presented in textbooks, the comoving frame precesses in the opposite direction with respect to the direction of the angular velocity of the precession $\vec{\omega}_{0}=\vec{v}\times d\vec{v}/dt/v^{2}$ and $\omega_{Th}\to-\infty$ in the limit $\gamma\longrightarrow\infty$.
-The theory of relativity shows us that the textbook expression for the Thomas precession in the lab frame and correct result $\vec{\omega}_{Th}=(1-1/\gamma)[\vec{v}\times d\vec{v}/dt]/v^{2}$ actually differ both in sign and magnitude.
-18.8.2 Incorrect Interpretation of the Correct BMT Result
-The existence of the usual incorrect expression for the Thomas precession in the lab frame has led to incorrect interpretation of the BMT result and, in particular, of the spin dynamics equation Eq.(133).
-Using the incorrect result for the Thomas precession, the BMT result for a small $g$ factor, Eq.(133), is usually presented as
-$\frac{d\vec{s}}{dt}=-\left[\left(\gamma-1\right)\vec{\omega_{0}}\right]\times\vec{s}=\vec{\omega}_{Th}\times\vec{s}~{},$ (135)
-Frequently, the first stumbling blocks in the process of understanding and accepting the correct Wigner (Thomas) rotation theory is a widespread belief that the experimental test of the BMT equation is a direct test of the incorrect expression for Thomas precession(3).
-The interpretation of Eq.(133) as the Thomas precession Eq.(135) is presented in textbooks as alternative approach to the already developed BMT theory.
-Authors of textbooks got the correct BMT result by using the incorrect expression for the Thomas precession and an incorrect physical argument.
-According to textbooks, equation Eq.
-(135) gives the rotation of an elementary particle spin in the lab frame $d\vec{s}=-(\gamma-1)d\vec{\theta}\times\vec{s}$.
-In other words, the spin rotation is recorded in the lab frame.
-In contrast, equation Eq.
-(132) $d\vec{s}=-(\gamma-1)d\vec{\theta}\times\vec{s}$ gives the rotation of the spin vector $\vec{s}$ with respect to the lab frame axes in the comoving frame.
-We emphasize once again: according to the BMT theory, the spin rotation is recorded in the comoving frame (as viewing this of the lab observer).
-The spin rotation measurement in the lab frame is interpreted with viewpoint of the proper observer due to clear physical meaning of the proper spin vector.
-Angular momentum vector $\vec{s}$ is defined only in one particular ( proper) reference frame, and any statement about it refers to the rest frame as of that instant.
-18.9 Bibliography and Notes
-1.
-The motion of the classical spin in an external electromagnetic field is presented by the Bargmann-Michel-Telegdi (BMT) equation [74].
-The BMT equation is manifestly covariant and can be used in any inertial frame.
-It is the law of motion of the four-spin for a particle in a uniform electric and magnetic fields.
-2. A method for measuring the particle energy in an electron-positron storage ring by means of resonance depolarization by a radio-frequency longitudinal magnetic field is described in [75].
-3.
-The results in the Bargmann-Michel-Telegdi paper [74] were obtained by the method of semi-classical approximation of the Dirac equation.
-The Wigner rotation was not considered in [74] at all, because the Dirac equation allow obtaining the solution for the total particle’s spin motion without an explicit splitting it into the Larmor and Wigner parts.
-Appendix I. Radiation by Moving Charges
+21 Appendix
+21.1 A1. Radiation by Moving Charges
 We start with the solution of Maxwell’s equation in the space-time domain, the well-known Lienard-Wiechert expression, and we subsequently apply a Fourier transformation.
@@
 (133) is that edge terms in the integration by parts can be dropped.
-This assumption can be justified by means of physical arguments in the most general situation accounting for the fact that the integral in $dt^{\prime}$ has to be performed over the entire history of the particle and that at $t^{\prime}=-\infty$ and $t^{\prime}=+\infty$ the electron does not contribute to the field anymore.
+This assumption can be justified by means of physical arguments in the most general situation accounting for the fact that the integral in $dt^{\prime}$ has to be performed over the entire history of the particle and that at $t^{\prime}=-\infty$ and $t^{\prime}=+\infty$, the electron does not contribute to the field anymore.
 Let us give a concrete example for an ultra-relativistic electron.
@@
 $L_{\mathrm{fb}}\sim(c\rho^{2}/\omega)^{1/3}$, where $\rho$ is the bending radius.
-In this way, intuitively, the magnets act like switches: the first magnet switches the radiation on, the second switches it off.
+In this way, intuitively, the magnets act like switches: the first magnet switches the radiation on, and the second switches it off.
 Then, what precedes the upstream bend and what follows the downstream bend does not contribute to the field detected at the screen position.
-With these caveat Eq.
+With this caveat Eq.
 (133) is completely equivalent to Eq.
@@
 (133) is more interesting with respect to that in Eq.
-(138) (although equivalent to it) because the magnitude of the $1/R^{2}$-term in Eq.
+(127) (although equivalent to it) because of the magnitude of the $1/R^{2}$-term in Eq.
 (133) can directly be compared with the magnitude of the $1/R$-term inside the integral sign.
@@
 This is no mistake, as different integrands can lead to the same integral.
-If the position of the observer is far away enough from the charge, one can make the expansion Eq. (30). Using Eq. (148), we obtain Eq. (LABEL:revwied).
-Appendix II. Undulator Radiation in Resonance Approximation. Far Zone
+If the position of the observer is far away enough from the charge, one can make the expansion $\left|\vec{r}_{0}-\vec{r}(t)\right|=r_{0}-\vec{n}\cdot\vec{r}(t)$.
+Using Eq.
+(137), we obtain Eq.
+(LABEL:revwied).
+21.2 A2. Undulator Radiation in Resonance Approximation. Far Zone
 Calculations pertaining undulator radiation are well established see e.g. [80]. In all generality, the field in Eq. (51) can be written as
@@
 […]
-Within the resonance approximation we further select frequencies such that
-$\frac{|\Delta\omega|}{\omega_{r}}\ll 1~{},~{}~{}~{}~{}\mathrm{i.e.}~{}~{}|C|\ll k_{w}~{}.$ (161)
+Within the resonance approximation we further select frequencies such that $|\Delta\omega|/\omega_{r}\ll 1$ (i.e.
+$|C|\ll k_{w}$).
 Note that this condition on frequencies automatically selects observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$.
@@
 As a result, one obtains zero transverse field, $\vec{\widetilde{E}}=0$, with accuracy $1/N_{w}$.
-Under the constraint imposed by (161), independently of the value of $K$ and for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, we have
+Under the constraint imposed by $|\Delta\omega|/\omega_{r}\ll 1$ , independently of the value of $K$ and for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, we have
 $|v|={K|\theta_{x}|\over{\gamma}}{\omega\over{k_{w}c}}=\left(1+\frac{\Delta\omega}{\omega_{r}}\right)\frac{2\sqrt{2}K}{\sqrt{2+K^{2}}}\bar{\gamma}_{z}|\theta_{x}|\lesssim\bar{\gamma}_{z}|\theta_{x}|\ll 1~{}.$ (150)
-This means that, independently of $K$, $|v|\ll 1$ and we may expand $J_{n}(v)$ in Eq. (158) according to $J_{n}(v)\simeq[2^{-n}/\Gamma(1+n)]~{}v^{n}$, $\Gamma(\cdot)$ being the Euler gamma function
 […]
@@
 However, it scales as $\vec{\theta}v$.
-Now, using condition (161) we see that, for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, $|\vec{\theta}|~{}|v|\sim(\sqrt{2}~{}K~{}/\sqrt{2+K^{2}}~{})~{}\bar{\gamma}_{z}\theta^{2}\ll K/\gamma$.
+Now, using condition $|\Delta\omega|/\omega_{r}\ll 1$ we see that, for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, $|\vec{\theta}|~{}|v|\sim(\sqrt{2}~{}K~{}/\sqrt{2+K^{2}}~{})~{}\bar{\gamma}_{z}\theta^{2}\ll K/\gamma$.
 Therefore, the term in $\vec{\theta}$ is negligible with respect to the term in $\vec{e}_{x}$ for $n=0$, that scales as $K/\gamma$.
@@
 […]
-Appendix III. Approximation for the Electron Path
+21.3 A3. Approximation for the Electron Path
 Let us now discuss the case of the radiation from a single electron with an arbitrary angular deflection $\vec{\eta}$ and an arbitrary offset $\vec{l}$ with respect to a reference orbit defined as the path through the origin of the coordinate system, that is $x(0)=y(0)=0$.
@@
 $\left.-2(\theta_{y}-\eta_{y})y_{r}(z^{\prime})+z^{\prime}\left((\theta_{x}-\eta_{x})^{2}+(\theta_{y}-\eta_{y})^{2}\right)\right]~{}.$ (167)
-Appendix IV. Self-Electromagnetic Fields of the Modulated Electron Beam
+21.4 A4. Self-Electromagnetic Fields of the Modulated Electron Beam
 In general, the electrodynamical theory is based on the exploitation, for the ultra-relativistic particles, of the small parameter $1/\gamma^{2}$.